A Complete Bibliography of IEEE Transactions on Computers (2010–2019)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 December 2023 Version 1.94

Title word cross-reference

(2^{n} − 1, 2^{n+p}, 2^{n} + 1) [Hia17]. (G(2^{m}, 4)) [KP13]. (t, k) [CH13, Cha10a, CH11]. + [BKV12]. 0 [XHX^{+}17]. 0 < l < r [KCK16]. 1 [HWG^{+}14, JSC10, MSK15, XHX^{+}17]. 10 [VAB14]. 2 [CMB13, CND^{+}18, EJ15, LJY^{+}15, RKZ16, SG12, TLP18, WHL17]. 2^{n} [Sou15]. 2^{n} + 1 [HMC11, VD12]. 3 [AD10, ASS^{+}18, AVS^{+}14, CTH14, CCK^{+}16a, CCL^{+}18, CWTT13, DYW15, EDL^{+}14, EYBK15, HCY18, JKD^{+}19, KAH^{+}15, KKC15b, LJY^{+}15, LLW^{+}17, MWWT13, RVL^{+}14, SKEB18, SPC^{+}18, TMS^{+}14, TCHL18, WJY^{+}17, WZL^{+}17, XCF16, YEY^{+}16, YMG16, ZDYZ14]. 4 [PP16, Sou15, TAM^{+}16]. 5 [FGS^{+}13]. 8 × 8 [CND^{+}18]. ^{4} [ZLN11]. B^{+} [FYSK14]. d [PRM16]. D^{3} [YTD^{+}18]. η_{T} [BDE^{+}11]. F_{3}^{m} [AH10]. GF (2) [Ose11]. GF(2^{m}) [CLL^{+}14, Cil13, DJA14, WF12]. K [FG10, AD10, AD12, Amm14, Fen14, FEM^{+}18, HK13b, NTR14, YUGD14, YLA^{+}15, ZCL^{+}16]. L [CMB13]. LF SR [Pom16a]. F_{q} [KCK16]. N [ADJG19, AMVOS^{+}15, AVS^{+}16, FG10, HK13b, Ose11, YM11, YUGD14, Zot10]. n × k(k ≥ n/2) [MC11]. P [BCTV15, GT19]. q ≡ lr^{s}(modr^{s+1}) [KCK16]. r [KCK16]. s [KCK16]. T [KMM16, MCT19]. t/k [LXZH16, ZLXW15]. τ [ADJ12]. Z_{N} [LCwW10]. Z_{q} [EBE13].

-Approximation [SG12]. -Ary [HK13b, FG10]. -Atomicity [WHL17]. -Automated [SBMP18]. -Based [Pom16a]. -Bit [YUGD14]. -Circuits [BCTV15].

-Connected [SG12]. -count [MCT19]. -Coverage [AD10, AD12]. -Covered [Amm14]. -Cubes [HK13b, FG10]. -D [MSK15]. -Diagnosability [Cha10a, CH13, LXZH16, ZLXW15]. -Diagnosis [CH11]. -Dimensional [Zot10, AMVOS^{+}15, AVS^{+}16]. -Encoding [XHX^{+}17]. -Extra [CTH14]. -Moduli [Sou15]. -Networks [CMB13]. -NN [FEM^{+}18, ZCL^{+}16]. -overlap [ADJG19]. -Partition [NTR14]. -Pipeline [PRM16]. -Term [Ose11]. -th [KCK16]. -Times [YLA^{+}15]. -Tree [FYSK14].

/Many [SNM16].

10Gbps [PRGBSAC19]. 128-Bit [GV14]. 16 [ZAG19]. 1687 [ZNL18].

2 [XYF^{+}15]. 2.0 [PC16]. 2.5D [DYHX16]. 256 [IDG^{+}17].

3D [ALW11, DSPB13, SKEB16, SVAB14, ZDYZ13]. 3D-ICE [SVAB14]. 3D-NoCs [DSPB13, SKEB16].

4-Bit [GM12].

5 [ZZS10, ZZL14]. 512 [GV14].

6 [FSL^{+}17, ROGHNB^{+}18, ZLWZ15].

802.11n [CKH15, GY16, GY14]. 802.15.4 [HXVQ15, NBZP17]. 802.16 [CTS13]. 802.16-Based [CTS13].

Abort [EFGT18, IGLM15]. Abort-on-Fail [IGLM15]. Absorbing [MTFK19]. Abstract [KN11a]. Abstraction [BFP11, HSH^{+}10, LVF19, YCK16, ZYY10]. Accelerate [RS10, SJVR19, ZLWZ15]. Accelerated [BEHL^{+}19, MTFK19, SCSL12]. Accelerating [CMO^{+}16, CYHC14, DOS15, LLCC13, RWZZ14, WZW^{+}19, XZL^{+}19, YEG^{+}15]. Acceleration [AAR19, GIKR19, KCRG15, KN13, LK18, TLL^{+}13, XYF^{+}15, ZWH^{+}15]. Accelerator [AKTB18, BQP^{+}16, BCMJ10, DW10, LJG^{+}19, LMB17, LLD19, MSPK12, MRL^{+}18, ÖDSS17, PC16, PGvdG14, SDP^{+}15, YMG15, YAG19]. Accelerator-Based [AKTB18]. Accelerators [BZ15, DMK^{+}15, MGW14, SKPK10, VMB19]. Access [Ano13e, CFL^{+}18, CS15, HK16, HHW^{+}18, KP15, LSG^{+}18, LHYZ13, NH10, RC14, SPTC15, SCJ^{+}16b, SCJ^{+}16a, TLH^{+}16, WHZ^{+}15, XKT^{+}15, YLA^{+}15, YAGB17, ZICL12, ZMY11]. Access-Time [HK16]. Accesses [CVPS19, CLW^{+}19, CLM^{+}19, LK15a, LXW^{+}19]. Accounting [LMC^{+}12]. Accumulate [CCRL19]. Accumulation [KGD16]. Accumulators [BGHL19]. Accuracy [Pom16b]. Accurate [AFC10, BBK10, CTL^{+}17, HCL^{+}14, ISC15, Iko15, JCY^{+}13, KGD16, LJ13, MNFA14, MWWT13, MD16, RMB^{+}13, SQJ^{+}15, TKT16, ZMR^{+}13]. Achieving [HHW^{+}18, RRS^{+}16, TLH^{+}16, ZC13]. ACO [CHC^{+}15]. ACO-Based [CHC^{+}15]. Acoustic [UVG16]. Across [LQD^{+}16, CLS14, JSE14]. Activate [LYT^{+}16]. Activation [RSN^{+}18, SCJ^{+}16a]. Active [HV14a, JRP^{+}14, LYT^{+}16, ST18a, WL13, YYC12, DHC^{+}16]. Activities [SJD^{+}18]. Activity [Pom12d, Pom13a]. Actor [AEKT15, ASTU10, MB12b, ZM17]. Actor-Critic [ZM17]. Actors [DSB13]. Actuator [SPC^{+}16]. Ad [CWZ11, CS15, CWTT13, CWY13, DLL^{+}12, FS10, GDY15, LYCT10, RDEN10, TH11, XWY10]. Adaptable [KA19]. Adaptation [Ano13f, GSH^{+}14, RDEN10, YyHL11]. Adapting [WGLL13]. Adaptive [AO12a, AKJ^{+}13, ACW^{+}11, AVS^{+}16, BKPMC13, CHC^{+}15, CCE^{+}18, CKD^{+}17, CGL^{+}13, CYL^{+}14, DYHX16, DY12, EDL^{+}14, FFCB14, FYSK14, GLXY13, Ged14, GV15, HCL15, HCSW15, HLWV17, IBH^{+}13, JT15, JLLH19, KSEG15, LOH17, LSL15, LFH^{+}16, LHH17, LWH^{+}16, MTBB10, NY15, NY19, OKY^{+}19, OGH^{+}14, PPKW12, PM14, RBG14, RVL^{+}14, RS10, RXC^{+}15, SKEB18, SOM^{+}13, SRK^{+}17, SXCL14, SDL^{+}19, TLP17, TRA18, TLGM17, WWM16, WW16, WYL^{+}15, WAK^{+}17, XHLW19, YyHL11, YWW^{+}16, YZ15, YHV13, ZWH^{+}15, ZL16]. Adaptive-Acceleration [ZWH^{+}15]. Adapts [WTBT13]. Adas [HHC^{+}18]. adBoost [khR^{+}18]. Add [WF17, ZCK19]. Adder [CL10, KT19, Nan19, PSL17, SJS^{+}14, VD12]. Adders [DDNT19, HVZ13, Kor15, LHL13b, LHL15b, LCW^{+}16, MHH^{+}17, WLG^{+}19]. Adding [LLLP14]. Addition [GVGNCVM16, JPG10, JDA15, KBP13, MLH12, VMHGN18]. Addition/Subtraction [KBP13]. Additions [LJ13]. Additive [TM18]. Address [AJH15, CKKS14, CQW^{+}15, CC11, KCKL19, LYS10, LLHC15, ML16, SCZ^{+}16, SRHC12, YCKH16]. Addressable [ALBP14, PO13, SMRM17, CCC^{+}18]. Addressing [LXW^{+}19]. ADDSEN [WAK^{+}17]. Adjacent [NL18, Pom13a]. Adjustment [Yam10]. Admission [NZLK14]. Admissions [XWL10]. Advanced [KK18, SBP^{+}14, XL16, MKRM10]. Advances [LLK18]. Advancing [ZCW18]. Advocacy [ST11b]. Advocacy-Free [ST11b]. AES [JL11, LB13, MMP13, MKRM12]. AES-GCM [JL11, MKRM12]. AES/PCLMULQDQ [JL11]. Affine [BLB^{+}19, RKT19, ZYHZ16]. Against [FRB^{+}18, GDLL18, MSS^{+}18, SBMP18, YLY^{+}15a, ASBdS16, GSF^{+}10, KASZ13, LFJ^{+}13, PLZW14, SKZS13, SBM15, SEY14, SWWC11]. Age [GEN^{+}17]. Agent [LLCH13, SD13, TKL^{+}14, ZCYX15].

Agent-Based [ZCYX15]. Aggregate [PSM17, CLW16b]. Aggregation [Ano13g, HXVQ15, KLT16, LWW11, QQW^{+}17, RWZZ14, VBR^{+}13, WJL^{+}14, ZHW^{+}16]. Aggressive [AS12, AS14, ARS16]. Aggressiveness [Pan16]. Aging [AEGH19, KAH18b, LSA18, SKH16, VCSG^{+}19, ZBK^{+}17]. Aging-Aware [KAH18b, LSA18, VCSG^{+}19]. Agnostic [JKD^{+}19]. AH [GYC^{+}16]. AH-Tree [GYC^{+}16]. Aid [NL14]. Algebra [DSB13, FGS^{+}13, FGS^{+}15, PvdGG12]. Algebraic [CND^{+}18, HIJ^{+}19, JLMP11, LYK19]. Algorithm [AT16, AHNT16, AK14, Bai17, CS11b, CZ16, CPL16, DZD^{+}16, DSPB13, FG10, Fuj11, GWZ^{+}10, GNSR14, HC13b, HWL^{+}14, HGEG11, JL11, JGHD11, KS14, KO14, KCK16, LBSK17, LPL^{+}13, LYOB15, LT14, LMC^{+}15, LY11, LBN14, LPW10, LYCT10, LLL11, LLCC13, LZ14, LSZ^{+}15, LCW^{+}15, MJW^{+}14, MRL^{+}18, OGH^{+}14, PGvdG14, PB11, QML^{+}15, Red14, SKEB16, SKEB18, SRR^{+}16, SG13, SL10, SR14, SKA10, SJS10, SCNS10, TLZV11, TLP18, TC16, Tsa13, VVMAZ12, VB13, WBZ^{+}15, WW16, XCW^{+}10, Yan14, YLL16, YLA10, ZWLS15, ZFJ^{+}17, dAJM14]. Algorithmic [CAGM14, DDNP11, GLP^{+}12, HWG^{+}14, JSC10, LKT13, NL14]. Algorithms [ADOKM10, AD13, BJ10, CCK10, CCH^{+}15a, CB15, CCR^{+}17, DALD18, EJ15, FMP19, GY14, GGL^{+}14, HT16, HMA^{+}10, HWG^{+}14, JWH^{+}15, JSC10, JMMP16, KLT16, Kůr12, LHC^{+}14, LB15a, LSX13, LCwW10, LLLP14, LMT13, MMH14, ML16, NZC11, PN16, PP16, Pip11, RHC^{+}14, RT14, SG12, ST11b, TKL^{+}14, WTY^{+}14, WWM16, XLL15, ZHM14, ZMRQ11]. Alignment [BEHL^{+}19, SKPK10]. Alignments [BCMJ10]. All-Optical [KH14]. All-to-All [ZGY13]. Allocating [MFG14]. Allocation [AF14, AQPMS15, BSM^{+}14, CLS14, CPL17, DKG13, GO10, GDY15, HCCG10, HK17, IHR^{+}16, LZA^{+}16, LGF^{+}15, MNGV16, PLP^{+}13, PCLN15, PAP13, PAC^{+}12, PCZB11, PBE17, RCN11, UMN18, VTW16, WLT^{+}16, XLL^{+}18]. Allocation/Deallocation [PCLN15]. Allocations [XLTZ11]. Allocator [PKY19]. Almost [WHL17]. Alternating [HFZ13]. Alternatives [YLGE14]. ALU [AC11, HK15a, RTRM19]. ALV [ZZS10]. Always [TBC^{+}17]. Always-On [TBC^{+}17]. Ambient [JCY^{+}13]. Amdahl [CA12a, YMG16]. Amplification [SQJ^{+}15]. Analysable [KAQC14]. Analyses [LSSE15]. Analysis [ADJG19, AXS^{+}10, ABEP16, AS14, BM13a, BBK10, BRN^{+}15, BLB^{+}19, BS14, BBB16, BDB18, BMZ17, BTW13, CS11b, CLW^{+}15, CSW^{+}15, CTL^{+}17, CJ12, DMXY14, DDNT19, FEM^{+}18, FHL^{+}18, GW16, GGL^{+}14, HB11, HTA10, HMA^{+}10, HGW^{+}17, HL10b, Iko15, JRW^{+}14, JKY10, JJC14, KE19, KGP15, KAH^{+}15, KKY^{+}16, LHC^{+}14, LYK19, Lee17, LLL16, LDB^{+}17, LHL13a, LCW^{+}16, LGS^{+}18, LCY^{+}19, MKT^{+}11, MTGM12, MMTM15, MHHS17, MBD^{+}17, MBB^{+}17, MMAC19, MCXZ18, MHML15, NDC^{+}13, NRG15, OP15, PC10, PFGB14, RZZ^{+}15, RHC^{+}14, RM15a, RBG^{+}19, SXLC15, SKEB16, SMB^{+}15, SX12, SZDL14, SCNS10, TS11, UMN18, VSC^{+}19, VALK19, WCM^{+}16, WLZ10, XWLX17, YYW^{+}16, YLH10, YTM16, ZICL12, ZJH^{+}14, ZT15, ZWD^{+}16, ZAG19]. Analysis-Based [RZZ^{+}15]. Analytic [BD15, Fin10, JAD^{+}18]. Analytical [EYBK15, GPN11, KM11, LLZ^{+}17, LHL15b, MYHL16, MKM14, SV18, SKEB18, SMA19, VED^{+}16]. Analytics [HZW^{+}12, KRP18, XYHD17, ZQZ^{+}19]. Analyze [DS14]. Analyzing [WF14]. Anchored [BGHL19]. Android [CXLX15]. ANGEL [ZCYX15]. Angle [RS10]. Angular [YASS14]. Annual [Ano11a, Ano11b, Ano12a, Ano13a, Ano14a].

Anomalies [KKT15]. Anomaly [VSF^{+}17]. Anomaly-Based [VSF^{+}17]. Anonymization [ZDP^{+}15]. Anonymized [PLZW14]. Anonymizer [LFJ^{+}13]. Anonymous [FHH10, HLT^{+}15, YLA^{+}15]. Answers [SLLG15]. Ant [HCSW15]. Antecedence [SD13]. ANTELOPE [HLJ14]. Antenna [CKH15]. Antennas [GY16, GGL^{+}14]. Anti [YGS15]. Anti-Attacks [YGS15]. Any [Wal19]. AP [GY14]. APC [WS14]. APIs [SWWC11]. AppATP [LSL15]. Application [AKL14, AK14, BRN^{+}15, CCW^{+}10, CH13, CNJ14, cCWS14, DAS14, Fin10, GKB^{+}10, GCF^{+}16, JCY^{+}13, JKD^{+}19, JRC14, KKP^{+}16, KCS^{+}13, KL13, LKYC12, LSA^{+}17, LCHC14, LGMP10, MMCS18, MGW14, MD16, MAG^{+}17, MRW^{+}15, MY10, RBG14, SP16, SIB13, SRK^{+}17, WMW12, XJW^{+}16, YLGD19, ZCZ^{+}19, ZYY18]. Application-Adaptive [RBG14]. Application-Agnostic [JKD^{+}19]. Application-Aware [KCS^{+}13]. Application-Dependent [AKL14, KL13]. Application-Driven [LGMP10]. Application-Guided [SRK^{+}17]. Application-Level [CCW^{+}10]. Application-Specific [JRC14, LSA^{+}17, SP16]. Application-Support [LKYC12]. Application/System [JCY^{+}13]. Application/System-Dependent [JCY^{+}13]. Applications [ABB17, ALW11, AF14, ABEP16, AEP18, BQP^{+}16, BMP^{+}10, BMM11, BDB18, CLX14, CHLL16, CG18, DA12, FBE^{+}18, GJ14, GSX^{+}13, GPR^{+}19, HV12, HV13, KTAvdS16, KKC17, KN13, LKYC12, LGH15, LHH14a, MVB10, ML18, ÖDSS17, PWTS16, PAC^{+}12, QJM^{+}10, RKR15, RQ14, RNS13, SAR^{+}11, SIVH16, TLGM17, VTA16, WZZ10, WHL^{+}12, WLQS13, YCCJ15, YG10, YRG13, YHV13, YAGB17, ZCZL16, ZCS16, ZCW18, ZT15, ZYL15, ZCY^{+}16].

Applying [YY14]. Approach [AD14, ABSK15, BR13, BC16, CWX^{+}14, Cha10b, CLL^{+}14, CFW14, CRK10, CJ12, CH14, DDNP11, DMA^{+}15, DYCG16, DSY^{+}15, DRS^{+}16, DJO11, Fan16, GWMB13, GLXY13, GC16, GBF^{+}19, HRM^{+}16, HCL^{+}14, HF15, HMNN12, LP13a, LBWH11, LKT13, LWL^{+}16, MOMT12, MMB14, MKRM11, ML16, NL14, PCHS17, PR14, RCM^{+}16, RBIQ15, RM15b, SKC^{+}14, STR15, SD13, STK16, SQJ^{+}15, UHSA17, VECD13, VYEB18, VBR^{+}13, WEH^{+}19, WF14, WLS18, YMT13, ZZS10, ZWW^{+}16, ZCR16]. Approaches [DLL^{+}12, NR15, ORBM13, OPV^{+}17, YEY^{+}16]. Appropriate [ZRS^{+}16]. Approximate [AKL18, CHLL16, CHL17, DDNT19, GVPS19, HXVF12, JHQL16, KBO^{+}19, LHL13b, LHL15b, LQW^{+}17, LCY^{+}19, LJVJ18, MHHS17, MHH^{+}17, MHML15, NL16a, RSJR17, VALK19, WLG^{+}19]. Approximated [BM11]. Approximating [BKV12, dRV12]. Approximation [JLLH19, KMM16, KLT16, LHPH15, LSX13, LJ15, MMH14, SG12, SG13, WEX14, XTF^{+}12, XLL15, ZL11]. Approximations [SDP11]. ApproxSSD [ZWW19]. Apps [ZLW^{+}17]. Arb [Joh17, RVL^{+}14]. Arbiter [GSX^{+}13, SMCN18]. Arbitrary [Bin15, BGPV10, FHR14, Joh17, Lef17]. Arbitrary-Precision [Joh17, Lef17]. Arbitrary-State [FHR14]. Arbitrated [GGA^{+}17]. Arbitration [FJA^{+}17]. Architecting [CLM^{+}19, KASZ13]. Architectural [GLP^{+}12, HNB^{+}12, LR16, LGMP10, OKD^{+}16, RM15a]. Architectural-Level [LR16]. Architecture [ADJ12, AYC16, AFC10, Ano10c, BSS14, BSS15, BBB^{+}17, CWZC13, CYC^{+}16, CLOL18, DCCK17, DEE17, DCY^{+}13, DMK^{+}15, DJO11, EKA17, EE10, FMTK19, FZL^{+}14, FM16, FLP^{+}13, GRM16, GM11, GDJZ18, GIKR19, HK16, HLY14, IRMM^{+}16, JC12, JJZ^{+}16, KAH18a, KC14, KK10, KBH^{+}10, KKC15b, KH14, KAK18, KT12, LSC11, LCC10, LK15b, LR16, LS10a, LXW^{+}19, LT15, LHCL13, LYS14, LYL^{+}19, MWZ^{+}17, MSK15, MKAY11, MS12, MKLW14, MC11, MH19, NKEM11, PGvdG14, PMH^{+}14, QLR^{+}11, RVL^{+}14, ROGHNB^{+}18, SSGB19, SNY^{+}10, SWM^{+}10, ST18b, SRHC12, Tho12, TWTT11, THGT13, VED^{+}16, VB13, WTBT13, WLV^{+}14, WLZ^{+}15, WJY^{+}17, WhCCC12, WZLS16, YyHL11, YP12, YMG15, YYC12, YLH13, YEG^{+}15, ZL18, ZZJ^{+}19, ZOK^{+}19, ZM10]. Architecture-Aware [LSC11]. Architecture-Based [HLY14]. Architecture-Centric [DJO11]. Architectures [ARM16, AT16, ARM13, BLN^{+}15, BBI^{+}13, BMM11, BDE^{+}11, BS10, CMO^{+}16, CWF14, CHTD19, CCR^{+}17, DPO17, DKLB15, ERRM16, GD17, GCD^{+}11, GBO^{+}16, HRM^{+}16, HGEG11, JP13, LK10, LP17, LP12, LBN14, LR18, LW15, MGdC^{+}18, MKRM12, NWA12, OWP16, OOD^{+}17, PvdGG12, RMC^{+}15, RMERM19, RKN^{+}18, SLPB18, SLZX15, SL10, Yan14, YHT^{+}16, ZYW^{+}16, ZBK^{+}17, ZBW17]. Area [ABH^{+}13, BWV15, DCCK17, DJA11, FAK16, GKB^{+}10, HC17, LSC11, LYOB15, LFH^{+}16, SH12, VALK19, WF17, YXWL16]. Area-Based [LYOB15]. Area-Efficient [DJA11, FAK16, LFH^{+}16]. Area-Time [ABH^{+}13, DCCK17, GKB^{+}10, HC17, LSC11]. Areas [YKK^{+}15]. Arithmetic [AO12b, Ano11e, AHI12, BLB^{+}19, BF19, BCS11, BdD19, CATB19, FMP19, FVV12, FM19, FML10, GKS14, HSA14, HMO^{+}17, IDG^{+}17, JLMP11, JCK15, Joh17, JMMP16, KMP11, LOC^{+}16, LMB17, NST14, RMC^{+}15, UHSA17, XMH13, ZCW18, ZAG19]. Arithmetic-Based [BLB^{+}19]. Arithmetical [Kůr12, MEBS17]. ARM [BBA19]. Armed [KTAvdS16]. ARMOR [GDJZ18]. ARMv8 [SD18]. ARMv8-A [SD18]. Array [Gor14, LLOS13, NZC11, SBW^{+}16, ZWW19]. Arrays [ASS^{+}18, CVGZ15, DSW^{+}14, DRC14, JWH^{+}15, LB13, SP12, VSC^{+}19, WNCH17]. Art [LLK18]. Ary [FG10, HK13b]. ASAP [BEHL^{+}19]. ASIC [MMP13]. Aspects [HWG^{+}14, JSC10]. ASSER [YTD^{+}17]. Assertions [VA11]. Assessment [CCD12, EYBK15, XWL^{+}16b]. Assigning [MW13]. Assignment [CWZ13, HCH15, LXL^{+}13, LDL^{+}17, RCM^{+}16]. Assisted [ADC11, ACGP13, CJSM17, JAS^{+}15, LHL^{+}15a, LYT^{+}16, LLX^{+}17, LLLJ13, LLKA19, RTL^{+}18]. Associated [Ibr16, RGK15]. Association [GY14, SBH11]. Associative [DPS11, YMG15]. Assume [LH11]. Assume-Guarantee [LH11]. Assumptions [HMZ^{+}14, SBM15]. Asymmetric [GGSPM18, HCY18, KH18, Li12a, LCLC19, RBK^{+}12, YSZ^{+}14, ZFJ^{+}17]. Asymmetry [WBGI19]. Asynchronous [AF14, CCL^{+}13, HKWC14, KC13, LK18, OMFH14, RZPX19, YHML16, ZLJ^{+}17, ZYY10, ZCR16]. Atmosphere [JSE14]. Atmospheric [XYF^{+}15]. Atomic [WHL17]. Atomicity [LJL13, WHL17]. ATP [VECD13]. Attack [DP13, GDLL18, GV14, MLW12, MSS17, MSS^{+}18, MMP13, PLZW14, SBM15, XJWW13, YWM19]. Attacks [AS14, BPBBL13, BK12, CKM15, CSS13, CTL^{+}17, FRB^{+}18, HIJ^{+}19, KSN^{+}15, KASZ13, LFJ^{+}13, OGPK14, PDJ^{+}19, RM15a, SKZS13, SEY14, SBMP18, TJH^{+}15, TIHM18, YLY^{+}15a, YZF^{+}10, YGS15]. Attestation [MGdC^{+}18, MBS^{+}12]. Attribute [FHR14, RVH^{+}16, SX12, WHZ^{+}15, XHX^{+}17, YLA^{+}15, ZPM^{+}15, ZHW15]. Attribute-Based [FHR14, RVH^{+}16, WHZ^{+}15, XHX^{+}17, YLA^{+}15, ZPM^{+}15]. Auctions [ZJL^{+}16, ZGWC15]. Audio [CJG16]. Audio-on-Demand [CJG16]. Auditable [HRK17]. Auditing [JCM16, LLXC16, LRY^{+}15, WCW^{+}13, WCH^{+}15]. Augmentation [NRG15]. Authentic [HLT^{+}15]. Authenticated [BDMLN16, HL10a, LCCJ13, YRT^{+}16]. Authentication [CCW^{+}10, CJ13, Har13, HBCC13, Kim15, KH10, LCLL15, MSKRJ17, SZDL14, ZHW^{+}16]. Auto [ABSK15, YXWL16]. Auto-Correction [ABSK15]. Auto-Focus [YXWL16]. Automata [GO10, KKS14, LW13, MSK15, RSG^{+}19, XXBL17]. Automated [CCD12, FM19, MSK15, SBMP18, XKT^{+}15]. Automatic [BRN^{+}15, BFP11, CLX14, CYA13, GJ14, GAFN15, Sri10]. Automating [MRW^{+}15]. Automaton [LLL11]. Automotive [BCD^{+}16]. Autonomic [VECD13]. Autonomous [JGHD11]. Autotuning [GVPS19]. Availability [CS15, WJF^{+}11, WJM15, ZHM^{+}19]. Available [Ano11g, Ano11h]. Avoidance [CCH^{+}15a, RVH^{+}16, WL13]. Avoiding [CRG^{+}13]. Aware [AQPMS15, ARS16, BLN^{+}15, BMS11, CFL^{+}18, CSPC12, CYCC11, CNJ14, CZP^{+}16, CTD^{+}16, CPL16, CPL17, CKD^{+}17, CG18, CWTT13, CRJZ16, CGJ^{+}10, DSKH15, FYSK14, FSPD16, FSPD17, FBWMM13, GVPS19, GKD^{+}17, GGSPM18, GBA18a, GHL17, GBD^{+}15, GCF^{+}16, HRM^{+}16, HMR^{+}17, HWZ^{+}12, HV12, HWS^{+}17, HWSX17, HS18, HV13, HWZ^{+}17, HLJ14, HLF14, IPS17, IS11, JSA17, JJK^{+}11, JSC^{+}17, JZLD10, JRJ^{+}18, JWC12, KSS12, KLK18, KIJ14, KS14, KPS^{+}17, KAH18b, KAH19, KSJ^{+}12, KkC15a, KLC18, KCS^{+}13, LKJ15, LSC11, LSC10, LY11, LMNP11, LSK13, LBN14, LKS^{+}14, LK16a, LK16b, LSA18, LYJ^{+}18, LWH^{+}16, LCLC19, Man16, MMCS18, MOYB12, MWY^{+}16, MKM14, OOD^{+}17, PVKA14, PAC^{+}12, PBT13, PBE17, QJM^{+}10, QLH^{+}16, RKZ16, SKPC15, SBP16, SJVR19, SRR^{+}16, SKH16, SLC^{+}15b, SZZ^{+}19, SYK14, TFCY16, VCSG^{+}19, WLK15, WJY^{+}17, WZL^{+}17, WWT^{+}18, WBGI19, WSXZ13, WFY^{+}17, XJFT16, XCF16, XWH14, XWL^{+}16a, XLL^{+}14]. Aware [XLJ16, YCLH16, YCW^{+}19, YTM16, ZLBB19, ZJS14, ZDP^{+}15, ZDYZ13, ZDYZ14, ZV14, ZYL15, ZWW19, ZHQX19, ZQQ11, ZMRQ11, khR^{+}18, JYL^{+}17]. Awareness [YHML16]. Axiomatization [AGCD16].

B [WLC^{+}15]. B&B [BMT14]. B-Tree [WLC^{+}15]. Back [XSR15]. Backbone [ZWX12]. Backlight [LHH14a]. Backoff [KCRG14]. Backup [BR13, LXJD15, MYW11, ZFJ^{+}17]. Backward [JRW^{+}14, LLL11, LLW^{+}18, SKZS13]. Bag [BMP^{+}10, HV13]. Bag-of-Tasks [BMP^{+}10, HV13]. BAGC [LSK13]. Balance [HHW^{+}18]. Balanced [LBS15, ZWYY15, ZCY^{+}16]. Balancing [AO12a, ADOKM10, BR13, CHC^{+}15, HC13b, JR17, KRP18, PBL16, QJM^{+}10, RKZ16, SMTK12, SLS^{+}12, Tse12, XAYL15, ZV14]. Bandits [KTAvdS16]. Bandwidth [CLM^{+}19, FSPD16, JPLP13, LKH16, LYS14, NH10, VC10, YYW^{+}16, YCCWC15, YYP^{+}16, ZRS^{+}16, ZGWC15]. Bandwidth-Aware [FSPD16]. Bank [SJC^{+}17a]. Bank-Group [SJC^{+}17a]. Banked [vdBGLGL^{+}16]. Bare [QPG10, YXZZ14]. Bare-Metal [YXZZ14]. Barrett [KVV10]. Barrier [WCLY16, Zot10]. Base [Wal19]. Based [AMR18, AF14, AK15, ABSK15, AKL14, AT16, AHNT16, ADC11, AKTB18, AS16, BLB^{+}19, BBA19, BDM^{+}19, BWV15, BGMR13, BMS11, BF19, BBB^{+}17, BBH12, BDL^{+}13, CCM14, CCK10, CHN14, CMLS15, CJSM17, Cha10b, CK11, CHH^{+}13, CKKS14, CHC^{+}15, CCE^{+}18, CKM15, CMRH17, CHCK12, CHLT14, CCC15, CSW^{+}15, CYC^{+}16, CLOL18, CGL^{+}18, CLCS19, CC11, CYHL14, CYA13, CAGM14, CTS13, CYL^{+}14, CND^{+}18, CH14, CWCS15, DYW15, DYHX16, DP13, DCCK17, DCCK18, DYJ19, DCM16, DA12, DPS11, DJA11, DZ10, DHC^{+}16, DRS^{+}16, DW10, DKK16, DDNT19, EDL^{+}14, EKA17, EYBK15, EGVFC^{+}12, EFGT18, EFPC16, FMTK19, FHH10, FHR14, FAK16, FEM^{+}18, FBR^{+}12, FGS^{+}13, FHW18, GT19, GRM16, GDLL18, GEN^{+}17, GBA18a, GV14, GP14, GNTS13, GTRMG18, GBF^{+}19, GCR^{+}19, GWZ^{+}10, GLP15, GBGI18, GABK11, HTH15, HWZ^{+}12, HSM14, HF15, HMZ^{+}14]. Based [HIJ^{+}19, HL10a, HMNN12, HSA14, HKR^{+}18, HLY14, HCSW15, HQLX15, Ima18, JPG10, JDA^{+}16, JC12, JGG^{+}14, JSH^{+}17, JPLP13, JKD^{+}19, KAH18a, KTAvdS16, KM11, KSN^{+}15, KP15, KLK17, KN11a, KN13, KVV10, KASZ13, KAK18, KL13, LBSK17, LRC10, LYOB15, LP12, LMC^{+}15, LSK13, LLCH13, LK16a, LRP16, LW17, LRP^{+}18, LTL14, LOX^{+}13, LS13, LLL15, LLC^{+}15, LLZ^{+}17, LSG^{+}18, LCwW10, LBWH11, LP13b, LPW10, LKLT12, LCH13, LHH14a, LFJ^{+}13, LTW^{+}12, LHYZ13, LCCJ13, LRY^{+}15, LLW^{+}17, LZZ^{+}17b, LMB17, LYJ^{+}18, LLD19, LJ15, LZS^{+}13, LLHC15, LSG^{+}15, LSGZ16, LB15b, LSXP14, LGF^{+}15, MLW12, MGdC^{+}18, MSPK12, MM16, MMCS18, MFT^{+}17, MTFK19, MFG16, MMAC19, MLE14, MB16, MSKRJ17, ML16, MH19, NZLK14, OCK17, OGH^{+}14, OWP16, OKD^{+}16, PN16, PCHS17, PCHS18, PAP13, PTD^{+}12, PB11, Pom12d, Pom15a, Pom16a, PROM15, PMH^{+}14, PdG13, RVL^{+}14, RVH^{+}16, RZZ^{+}15, RKR15, RTRM19, RZPX19]. Based [RBMO11, RC14, ROGHNB^{+}18, RM15c, RJV^{+}18, SMP16, SMCN18, SKEB18, SBB18, SC18, SBP16, SDP^{+}12, SCZ^{+}16, SX12, SLC15a, SZG^{+}18, SL10, SKM14, SNM16, SKM^{+}13, STK16, SBMP18, TJH^{+}15, TB15, TWTT11, TIHM18, THM^{+}14, TGNSC11, TAM^{+}16, TKT16, USP^{+}13, VCSG^{+}19, VSF^{+}17, WF12, WL13, WXW^{+}14, WLC^{+}15, WHZ^{+}15, WWY^{+}16, WW16, WTZ^{+}19, WCL^{+}18, WLM15, WW14, WZCG16, WLS18, WLG^{+}19, WZW^{+}19, XAYL15, XJFH15, XCF16, XMH13, XLX^{+}14, XKT^{+}15, XJW^{+}16, XHX^{+}17, YLGE14, YW12, Yan14, YLGD19, YSZ^{+}14, YTD^{+}17, YYC12, YLH13, YLA^{+}15, Yun12, ZSP^{+}19, ZZ17, ZPM^{+}15, ZLY15, ZOK^{+}19, ZC13, ZM10, ZHW15, ZYY18, ZXX^{+}14, ZCYX15, ZLYS15, JJZ^{+}16]. Based-Encryption [ZHW15]. BasedX [KCY18]. Bases [ABH^{+}13, ARM13, ERRM16, MH15]. Basic [APP12]. Basis [AH10, CLL^{+}14, CCR^{+}17, DJA14, ERRMG15, Fan16, Gio12, HF15, KEK16, LLCH13, NWA11, NWA12, RM15b, RMERM19]. Bayesian [WP16]. BC [YL14, ZWC13]. BCD [CDL^{+}17, VAB14]. BDDs [FLS16]. Be [HLC^{+}19, LOC^{+}16]. Beaconless [RS13]. Beginning [Ano10f]. Behavior [DEE17, QQW^{+}17, SB10, SIB13]. Behavioral [SBH11]. Belief [ACGP13]. Belief-Propagation-Assisted [ACGP13]. Belt [WXLL13, WXLY15, WCLY16, XLW14]. Benchmark [DSKH15]. Benchmarking [YLY15b]. Berlekamp [Red14]. Bernstein [MRD19]. Best [RMB^{+}13]. Better [FP19, ZZL14]. Between [ADJG19, LK18, ADOKM10, BLMM16, CT13, GHK15, HXL11, LYL^{+}19, PR10, RWZZ14, SCJ^{+}16b]. Beyond [BK12, CA12a, IPS17]. Bi [ML18]. Bi-Objective [ML18]. Bias [LVMS18]. Bias-Variance [LVMS18]. Biased [SMK^{+}16, USH19]. Bidirectional [ZHM14]. BiFI [SBMP18]. Big [CLOL18, GZG^{+}16, JYL^{+}17, KTAvdS16, KLC18, KRP18, KYZC19, LRY^{+}15, MSG14, UMN18, WNCH17, ZDP^{+}15, ZYC16]. Big-Data [KTAvdS16]. Bin [SXCL14]. Binary [ADI11, APP12, AK14, BNP10, BT16, BZ14, BLMM16, CNH13, CHN14, CJSM17, CLC^{+}16, FP19, FNS16, GNTS13, GBF^{+}19, HRM11, HN11, HMNN12, Ima18, JDA15, LCY^{+}19, MBF18, NL15a, TGNSC11, WF17]. Binary-Ternary [ADI11, BT16]. Binary-to-Decimal [BZ14]. Binary/Decimal [WF17]. Binary128 [HMS^{+}12]. Binary64 [Nan16]. Binning [GKC19, OPV^{+}17]. Bio [PBL16]. Bio-Inspired [PBL16]. BioAura [MSKRJ17]. Bioinformatics [WZZ10]. Biological [BCMJ10, SKPK10]. Biology [LSZ^{+}15]. Biology-Inspired [LSZ^{+}15]. Biomass [CKN14]. Biomedical [BBB^{+}17]. Biophysically [ZMR^{+}13]. BIRDS [YXZZ14]. Bisection [LZ14]. Bistatic [WCLY16]. Bit [ARM16, CCH^{+}15a, CK15, Fan16, GRM16, GV14, GM12, Ima18, Jes15, LAAM11, LLQ^{+}14, PCZB11, RLRL19, SKZS13, SMK^{+}16, WNKL16, YFJ^{+}14, YUGD14, Zot10]. Bit-Level [ARM16]. Bit-Parallel [Fan16, Ima18]. Bit-Slice [Zot10]. Bit-Stuffing [CCH^{+}15a]. Bit-Width [CK15, LAAM11]. Bits [FKMK16, NL18, NM10, RXC^{+}15]. Bitstream [SBMP18]. Bivariate [JKMR11]. Blackbox [TIHM18]. BLAKE [GV14]. BLAKE-512-Based [GV14]. Block [BFMT16, CHN14, CQW^{+}15, CCY^{+}16, GWM^{+}17, HMNN12, JC12, JW16, KKJH19, LK15a, LYK19, PCHS17, PDJ^{+}19, SNY^{+}10, SZG^{+}18, TMS^{+}14, TCYH15, WLC^{+}15, WZK^{+}19, WLG^{+}19, YSZ^{+}14]. Block-Based [WLC^{+}15, WLG^{+}19, YSZ^{+}14]. Block-Level [CQW^{+}15, GWM^{+}17]. Block-Mapped [SNY^{+}10]. Block-Precise [LK15a]. Blocking [DVUS14, HWE^{+}16, SRR^{+}16]. Blocks [CRG^{+}13, MCM16, RPM16, SKM^{+}13]. Bloom [ADC11, HXVF12, LLLP14, ML16, PO13, QZC15, SMRML17, YM11]. Blue [WGW^{+}15]. Bluetooth [HWK15]. Board [Zom12a]. Boards [BC16]. Boolean [BCTV15, BCTV19, FD16, GBA18a, TSA^{+}19, ZYHZ16]. Boost [CVPS19, LKK^{+}19, SSW12]. Boosted [LDP10]. Boosting [CCL^{+}18, KAH18b, ZLLX15, khR^{+}18]. Booth [JHQL16, LQW^{+}17, VALK19]. Booting [Cha14]. Borrowing [CCAM14]. Both [LXW^{+}19, WHC^{+}15b, ZHM^{+}19]. Bound [BMT14, Bin15, Fuj11, WLZ10]. Boundary [LKLT12]. Boundary-Recognition-Free [LKLT12]. Bounded [CCK10, HRK17, SC18, XXBL17, XLW14, ZY12]. Bounding [Fuj11]. Bounds [FJA^{+}17, FS10, HCCG10, ISC15, MRD19, RMB^{+}13, ST18a, Tse12, TKL^{+}14, YLH10]. Box [BTBB14, MKRM11, SWF^{+}19]. Boxes [MM17, ST18a]. BPR [YKK^{+}15]. Branch [BMT14, DEE17, Fuj11]. Branch-and-Bound [Fuj11]. Branching [STR15]. Breakpoint [MSPK12]. Breakthrough [CVGZ15]. Bridging [FD16, TLL12]. Brief [Liu11]. Broadcast [BS14, GYC^{+}16, KDEV19, KH10, LTP^{+}14, LWF13, PSM17, QSYS16, WXS12, WQZ^{+}16, XJW^{+}16, ZGY13, ZHW15]. Broadcasting [KS10a, YMTV14]. Broadcasts [LK15a]. Broadside [Pom12c, Pom12b, Pom13a, Pom13b, Pom14, Pom15a, Pom16a]. Brokerage [CWTT13, LKLT12]. Bruijn [MG16, YMAG17]. BRW [CMLRHS13]. BSM [YMTV14]. BTI [GBA18a]. BTI-Induced [GBA18a]. Bubble [MWLJ15, WSL^{+}18]. Buddy [PCLN15]. Budget [AF14]. Budgeting [PKC^{+}17, WZM^{+}16, WSL^{+}18, WTZ^{+}19, ZR15a]. Buffer [Ano13f, CLZ19, LSK13, LKLM15, LKBS16, MBD11, MD13, MBGS10, OGH^{+}14, PPND17, WW16, OGH^{+}14]. Buffer-Aware [LSK13]. Bufferless [KC14, ZGY14]. Buffers [CVMA10, DKLB15, HGCT13]. Bugs [FM19, ZYW^{+}16]. Build [PDXZ13, WSXZ13]. Building [AHNT16, AVS^{+}14, LFH^{+}16, LZZ^{+}17b, SWZG15, TMS^{+}14]. Building-Block [TMS^{+}14]. Buildings [WZY16]. Built [AK16, LLW^{+}17, Pom14]. Built-In [AK16, Pom14]. Built-Off [LLW^{+}17]. Bulk [SBI12, YZH^{+}15, ZWD^{+}16]. Bullet [Ano10g]. Burst [HXVQ15, NL15b, NL16a, RV13, ASE17]. Burstiness [CMS10]. Bus [CYA13, EE17, HHLK12, RVL^{+}14]. Bus-Based [CYA13]. Buses [BPC12, ST12]. Buy [SBI12]. Buy-at-Bulk [SBI12]. BWLOCK [YAGB17]. By-Passing [YKK^{+}15]. Bypass [KRP18, MMB14]. Bypassing [PSND16, SZL^{+}16]. Byte [CCC^{+}18, RV13, SBB18]. Byte-addressable [CCC^{+}18]. Byzantine [DCK16, KDEV19, LLKA19, VCB^{+}13]. Byzantine-Resilient [KDEV19]. ByzCast [KDEV19].

C [CZS^{+}19, KLJ^{+}14, MF14, RWC18]. C-Lock [KLJ^{+}14]. C1G2 [SDZ15]. CA [SBB18]. CA-Based [SBB18]. CABA [MSKRJ17]. CABE [XHX^{+}17]. CACC [CWX^{+}14]. Cache [AK19, AYC16, ACW^{+}11, AGFM11, ADC11, BLN^{+}15, CVPS19, CWX^{+}14, CDQB15, CYCC11, CA12b, CKD^{+}17, CLCS19, DPS11, DW10, EF12, FSGAB^{+}16, GKD^{+}17, GGFPG15, HHM11, HK16, HK17, HK15b, HZX^{+}14, IPS17, JAKD18, JSC^{+}19, JZLD10, JKY10, KIJ14, KSEG15, KS14, KLK17, KCL19, KASZ13, KKC15b, KAQC14, KA19, LYH11, LKJ15, LK15b, LWKA15, LKLM15, LKBS16, LHL^{+}15a, LZZ16, LCLC19, MWLJ15, MHK15, MAD14, MMAC19, OOD^{+}17, PBV11, PCZB11, RM15a, RCFP^{+}12, RXC^{+}15, SV18, SZZ^{+}19, SYK14, SZL^{+}16, TAH^{+}16, VKS^{+}16, VYEB17, WMW12, WLT^{+}16, YMG15, ZGR13, ZDYZ13, ZDYZ14, KMC17]. Cache-Aware [IPS17]. Cache-Based [DW10, KASZ13, MMAC19]. Cache-Coherent [MWLJ15]. Cache-Enabled [LZZ16]. Caches [AVG^{+}15, CXLL16, CFMA19, CRG^{+}13, FKMK16, HWZ^{+}17, KIJ14, LLX^{+}17, MD16, MUMB11, VPS^{+}12, VSLD15, ZJS14]. Caching [CDQB15, HK15b, KJL11, KRP18, LOH17, MCC12]. CaCo [ZWW^{+}16]. CAIF [SDP^{+}15]. Calculating [WLG^{+}19]. Calculation [GPN11, SV18]. Calibration [DAPS14]. Call [Ano10c, Ano11c, Ano11d, Ano11e, Ano11f, Ano13d, CH14]. CAM [FAK16]. Camera [YASS14]. Can [HLC^{+}19, YGS15]. Candidate [Pom16b]. Capabilities [WJX^{+}19]. Capability [ABSK15]. Capable [LYH11]. Capacitance [FM16]. Capacitated [LSX13]. Capacitive [YEY^{+}16]. Capacitor [CLZ19]. Capacitor-Powered [CLZ19]. Capacity [CMS10, LK14, NZLK14, OP15, PDXZ13, SRCbL^{+}15, WJL^{+}12, WJL^{+}14, WLYY16, XHLW19, YCKH16, ZJS14]. Capacity-Based [NZLK14]. Capacity-Independent [YCKH16]. Capital [ZMW15]. Cards [AVS^{+}14]. Carlo [KN13, ZOD13]. Carry [DHM16, KT19, Lee17]. Carry-In [Lee17]. Carry-Save [DHM16]. Cascaded [CHC^{+}15, JWC12]. Cascading [HWSN15]. Case [AR17, BBK10, CCLH10, DZ10, FS10, HHC^{+}18, RCRK13, SD14, UVG16, VIDH19, WZLX12, WJL^{+}14, WRW16]. Categorical [PZZQ19]. Cauchy [CJK15, ZWW^{+}16]. Caused [HWK15]. Cayley [LGF^{+}15]. CC [RBG^{+}19]. CCA2 [GV14]. CCLS [LCW^{+}15]. CD [ZMW^{+}19]. CD-Xbar [ZMW^{+}19]. CDP [YCL^{+}12]. CDT [AMR18]. CDT-Based [AMR18]. CEDA [VA11]. Cell [ALBP14, ASS^{+}18, KO14, NL19, ST16, WZL^{+}17, ZZX^{+}15]. Cells [ACM^{+}16, CJK19, Li12a, PRM19]. Cellular [KKS14, MSK15, YZH^{+}15]. Center [GWMB13, GY15a, GY15b, HLJ14, JRS^{+}15, LXL^{+}14, LW15, LSHC15, SLZX15, ZWH^{+}15, ZMW15]. Centers [AQPMS15, CLS14, CPL17, GZB^{+}15, GZG^{+}16, GCL^{+}13, JSE14, LGF^{+}15, Man16, PAP13, SMTK12, SHH^{+}16, WJM15, XLF15, YHT^{+}16, ZR15a]. Centralized [AD12]. Centralizing [HPR16]. Centric [BQP^{+}16, DJO11, HWX15, LOH17, SLC^{+}15b, WZZ10, WLT^{+}16, YPB^{+}16]. Certification [Ano10d, Ano10e]. Certified [LB15b, MRD19]. Certifying [dDLM11]. CGRAs [JDA^{+}16]. Chain [KCY18, MTFK19, QZL^{+}16, TLL12, ZOD13]. Chained [IRMM^{+}16]. Chains [JDA15]. Challenges [Jun16]. Change [DY14, JSC^{+}19, LYB15, LKK^{+}19, NL19, PLM16, PP11, QML^{+}15, WSXZ13, ZZYZ14]. Channel [Bar16, BMZ17, BK12, CWZ13, CSS13, CYC11, CATB19, DP13, DKLB15, GYC^{+}16, HMZ^{+}14, HLY14, KASZ13, LGH15, LWY15, LFJ^{+}13, LLW^{+}17, LJL13, NDC^{+}13, RBG^{+}19, SKZS13, SMB^{+}15, STE17, SPTC15, WJL^{+}12, XWL^{+}16a, YMTV14, ZMB18]. Channel-Based [LLW^{+}17]. Channel-Diverse [LWY15]. Channel-Hopping [LWY15]. Channels [AP14, Cao12, KW14, LMB^{+}16]. Character [LPCW14]. Character-Set [LPCW14]. Characterising [HT16]. Characteristic [AK16, BDE^{+}11, LSG^{+}18, NR15]. Characteristics [HMZ^{+}14, LHL15b, VED^{+}16]. Characterization [DGC^{+}15, HWSN15, SKC^{+}14, TAH^{+}16, ZLH^{+}15]. Characterizing [DEE17, IS14, LLLJ13]. Charge [CCL^{+}18, NL15a]. Charging [ZWL15]. Chasing [LKLM15]. Chebyshev [PCHS14, ACO12, Gio12, LCwW10, LPW10]. Checking [CYHC14, FLS16, GCLC11, HSH^{+}10, HMC11, LYY16, LH12b, NS13, SP10, Sri10, WNCH17, XXBL17, ZYY10, ZZM^{+}15, CTS13]. Checkpoint [BTW13, KwPK^{+}15, LL11, LSA^{+}17]. Checkpoint/Restart [LL11]. Checkpointing [BCL^{+}17, ECJ^{+}16, HC13a, HCC^{+}18, JT15, LXDV17, SD13, ZYL15]. Checkpointing/Restart [ZYL15]. Checks [GBGI18]. Checksums [NC11, Red18].

CHERI [WJX^{+}19]. Chief [Lou19, Mon15a, Mon19]. Chinese [Fan16]. Chip [ABB17, Ano11d, Ano11f, ARS16, BKH^{+}13, BPT10, CKKS14, CHC^{+}15, CCE^{+}18, CNJ14, CDK^{+}18, CRK10, DMXY14, DAS14, DKLB15, DKG13, EDL^{+}14, EYBK15, FBWMM13, FTP13, GCD^{+}11, GC14, HMD^{+}17, HJBM14, HCSW15, HWE^{+}16, JKY10, JJZ^{+}16, JWC12, JRC14, KC14, KSEG15, KK10, KLK^{+}14, KGGJ14, KLC^{+}16, KKY^{+}16, KCL^{+}16, KH14, LKS^{+}14, LMJ14, LLX^{+}17, LYJ^{+}18, LKMSA16, MWW14, MNFA14, MMCS18, MKAY11, MD13, MKLW14, NVB16, OHCK17, PVKA14, PSND16, RMB^{+}13, RVC^{+}15, RRS^{+}16, RKZ16, SKPK10, SDE^{+}17, SRR^{+}16, SIVH16, SMN^{+}17, SC11, ST17, ST18b, STK16, VYEB18, VCG^{+}12, WMW12, WXW^{+}14, WWM16, WZM^{+}16, WWT^{+}18, WEH^{+}19, WZCG16, WLS18, XCF16, YMK^{+}17, YYC12, ZNL18, ZGY13, ZCY^{+}16, ZMS13]. Chip-Level [KLC^{+}16]. Chip-Multiprocessor [KGGJ14]. Chip-Multiprocessors [FBWMM13, LMJ14]. Chips [BMM11, TW10, YCK16, YLGD19]. Chunking [WLM15, ZFJ^{+}17]. CICQ [JPLP13]. Cipher [BFMT16, CMLS15, GCS^{+}13, HZ11, LYK19, ZAG19]. Ciphers [KE19, PDJ^{+}19]. Ciphertext [ZHW15]. Ciphertext-Policy [ZHW15]. Ciphertexts [WQZ^{+}16]. Circuit [CSS13, CCAM14, FNS16, JLLH19, LJ15, MCT19, NCD^{+}17, NZ15, PSL17, RMKR12, YUGD14]. Circuits [AO12b, BCTV15, CL10, CPRH16, DSR15, FM19, FP19, FFL18, GM15, GM12, ISC15, KMM16, LB15a, LHL13a, MVB10, MBF18, NLRB17, NI11, SRCK10, UHSA17, WMG18]. Circulant [IRMM^{+}16]. Circulants [KP13]. Circular [HK15b]. City [DHC^{+}16, LDMQ16, WHBR16]. Clairvoyant [ZQZ^{+}19]. ClamAV [OWP16].

Class [RLRL19]. Classes [ZYHZ16].

Classification

[BBH12, CXZ13, CW15, CW16, LMJ14, OPZ15, SDP^{+}15, ST11b, WP16, YFJ^{+}14]. Classified [HY12]. Classifier [YM11]. Classifiers [BMS12, BSS14, BSS15, KGV16]. Cleancache [VTW16]. Client [BSM^{+}14, MAG^{+}17]. Client-Side [MAG^{+}17]. Client/Server [BSM^{+}14]. Clients [GY14]. Clifford [FGS^{+}13, KMM16]. CLIM [JRJ^{+}18]. Clock [KAH18a, MOMT12, NL14, PPND17, Yam10, CXLL16, LBN14]. CLOCK-DWF [LBN14]. Clocking [NLRB17]. Clockless [BDDL18]. Clos [KC14, HMD^{+}17]. Closed [KGC14]. Closed-Loop [KGC14]. Closer [Jun16]. Cloud [BDL^{+}13, CLX14, CHLT14, CXYC16, CLW16b, DKW15, EFPC16, FLL14, GSG^{+}15, HLJ14, HLF14, JSE14, JCM16, KKJH19, LLC^{+}15, LYY16, LLXC16, LHH14a, LSL15, LRY^{+}15, LLS^{+}16, MSG14, Man16, MJWT16, MLOL15, ML13, PLP^{+}13, PSM17, PR14, QML^{+}15, RSNK17, SXCL14, SZW^{+}16, SHH^{+}16, TLH^{+}16, VP14, WCW^{+}13, WCH^{+}15, WLK15, WRW16, XL16, XLL^{+}14, XJW^{+}16, XLJ16, YCCJ15, YCLH16, YTD^{+}17, YLY15b, YLA^{+}15, ZDP^{+}15, ZLY15, ZWW^{+}16, ZJL^{+}16, ZYC16, ZLX^{+}16, ZL16, ZL18, ZWC^{+}18, ZGWC15, ZLYS15, Avr13, CSPC12]. Cloud-Based [LHH14a]. Cloud-of-Clouds [CHLT14]. CloudGenius [MRW^{+}15]. Cloudlets [RSN^{+}18]. CloudMon [WLLZ16]. Clouds [ALZ16, CLS14, CHLT14, GHL17, JT15, MNGV16, MRW^{+}15, VP14, WBZ^{+}15, WZL15, WLLZ16, ZCYX15]. CLU [ZJS14]. Cluster [LTVL15, QWB^{+}13, WZK^{+}19, YZHX12]. Clustered [AD12, BPG16, GSL10, USP^{+}13, Yan14]. ClusterFetch [RLSK18]. Clustering [LCL15, PZZQ19, RD18, SH12]. Clusterized [GBO^{+}16]. Clusters [ALZ16, CLOL18, DMK^{+}15, HV14b, HQLX15, LZ15, ML18, MRW^{+}15, QJM^{+}10, ZQQ11, ZMRQ11]. CMOL [ALW11]. CMOS [DCY^{+}13, PN16, SB16, SRCK10, WWY^{+}16, ZCW18]. CMOS-Compatible [DCY^{+}13]. CMOS-Type [ZCW18]. CMP [IB10, WSZ^{+}16]. CMPs [BGM^{+}13, FAA10, GFAM11, JAKD18]. CNN [LJG^{+}19]. cNV [WWY^{+}16]. Co [CI16, LH16, LDL^{+}17, MBD^{+}17, MRL^{+}18, MAHD18, SPTC15, ZJS14]. Co-Channel [SPTC15]. Co-Design [LH16, MRL^{+}18]. Co-Optimizing [ZJS14]. Co-Scheduling [LDL^{+}17, MBD^{+}17, MAHD18]. Co-Transformation [CI16]. Coalescing [WZK^{+}19]. Coarse [JLKR19]. Coarse-Grained [JLKR19]. Code [ADJG19, AFH^{+}10, BKH^{+}13, CJ12, DLC^{+}13, EKJ^{+}10, FKMK16, FSL^{+}17, HT12, KSN^{+}15, NL18, OGPK14, PLM16, SWWC11, VGF16, XLX^{+}14, YLP15, ZXX^{+}14]. Codec [SBB18]. Coded [HQLX15, LS10a, TYY^{+}16, ZLLX15]. Codes [ABA07, CDL^{+}17, DRM16, EBE13, HHKW12, HC17, HBAD14, Jha13, KW14, KLLK11, LSXP14, MNK11, NL15a, NL15b, NL16a, NL16b, PROM15, RV13, Red14, Red19, RLRL19, SEY14, TW10, VAB14, YCW11, YW12]. Codesign [PvdGG12, PGvdG14]. Coding [BBH12, CHLT14, CXYC16, CJ13, LCLL15, LLL15, LLZ^{+}17, NL19, SSKL16, SRK^{+}17, SZG^{+}18, TYY^{+}16, YY14, YCK10, ZWW^{+}16, Kim15]. Coding-Based [LLZ^{+}17]. Coefficient [ADJG19]. Coexistence [AVG^{+}15, HWK15]. Coflow [ZWL^{+}19]. Cognitive [BBVL14, YCCWC15]. Cognizant [KMJ^{+}11]. Coherence [AVG^{+}15, ADC11, FBWMM13, GGFPG15, KSEG15, LHH17, RCFP^{+}12, ST17, VYEB17, YRG13]. Coherent [MWLJ15]. Cold [LXJD15]. Collaborative [ZWL15, ZLYS15]. Collapsing [PR10]. Collection [CW10, DSW^{+}14, LSK13, LWW11, LDMQ16, LTW^{+}12, RLX15, UMN18]. Collections [CYJ^{+}10]. Collective [HWS^{+}17]. Collectors [ZMY11]. Collision [BK12, CWZ13, MMP13]. Collocated [HWK15]. Collusion [LSS13, RVH^{+}16]. Colony [HCSW15]. Column [LW17, LXW^{+}19]. Column-Based [LW17]. Combat [LSS13]. Combined [BK12, Jes15, PDJ^{+}19, SMRM17, WF17, WE12]. Combining [SLL15]. COMeT [RCC14]. Command [GCAG16]. Comment [GJ15, LCLL15, SCNS10, Tho12]. Comments [HWG^{+}14, Jha13, Kim15, Lee12, PCHS14, PCHS16, RM15b, ZM10]. Commercial [NY15, NYHB16]. Commit [HPR16, JLKR19]. Commits [HPR16]. Commodity [CJA^{+}16, RTL^{+}18]. Common [Pom12c]. Communication [AKTB18, BR13, BS16, CCM14, CWF14, CCW^{+}10, CDK^{+}18, DYHX16, GZG^{+}16, HXL11, KW14, KGP15, LHH14b, MTFK19, MAHD18, QJM^{+}10, RS13, SXLC15, SMN^{+}17, VECD13, VYEB17, WLQS13, XWL^{+}16a, YMT13, ZHW^{+}16]. Communication-Aware [QJM^{+}10]. Communications [CWY13, LGH15, LT15, SSKL16]. Community [FLJ14, LDB^{+}17, XAYL15, XWH14]. Community-Aware [XWH14]. Community-Based [XAYL15]. Compact [CCC15, CJK15, OMFH14, SBB18, SVAB14, YP12]. Compaction [Pom12d, Pom15c, RPM16]. Comparable [WHZ^{+}15, XHX^{+}17]. Comparative [AE11, HMA^{+}10]. Comparing [Hie13]. Comparison [BLMM16, CCH15b, CCLH10, HK13b, Li12a, PTD^{+}12, YLL16]. Comparisons [CGT^{+}15]. Compatibility [LTP^{+}14]. Compatible [DCY^{+}13, WWY^{+}16]. Competes [CI16]. Competition [FLL14]. Competitive [MFG16, MFG14]. Compilation [LK10].

Compiler [LHL^{+}15a, MB12a, OKC13]. Compiler-Assisted [LHL^{+}15a]. Compiler-Directed [OKC13]. Complement [LAAM11]. Complementary [Bai17]. Complete [AGCD16, BS14, BJ12]. Completely [LLM^{+}15]. Completeness [KLT16, SP10]. Complex [AC11, BC16, PRBM13, SZW^{+}16, WE12, WEX14, XLS^{+}12]. Complexity [AH10, AO12b, ADOKM10, ARM13, BNP10, BTW13, CHN14, GNSR14, HN11, HMNN12, JCK15, KCRG14, LWW11, OPAGS14, PCHS17, PCHS18, PRBM13, PSL17, WTY^{+}14, WS10, ZM10]. Complexity-Effective [KCRG14]. Component [CH11, CH13, FEM^{+}18]. Component-Composition [CH11, CH13]. Components [FS10, MF14]. Composability [GRL^{+}14]. Composing [LTLC12]. Composite [DKW15, LPD^{+}16, MKRM11, ZL19, ZM10]. Composited [MG16, YMAG17]. Composition [AK15, CT13, CH11, CH13, LH12a, LSSE15, SKPC15, SMCN18]. Compositional [VYEB18, ZZM^{+}15]. Compound [SX12]. Comprehensive [BGPV10, DZLP14, TAH^{+}16]. Compressed [BLN^{+}15, LKK^{+}17, LVJ16, ST17, WJX^{+}19]. Compressing [PBV11]. Compression [AKL18, DN11, DY12, Ged14, HL10b, JSA17, KO14, KWC^{+}16, KN11b, LYH11, LKK^{+}17, XDZ11, ZSP^{+}19, ZLW^{+}17, ZWL^{+}19, dRV12]. Compression-Induced [KWC^{+}16]. Compressor [KK18]. Compressors [MHML15]. Computable [LGH^{+}17]. Computation [ARM15, AK16, ARH14, AAR19, AH10, BG12, BBVL14, CLW^{+}16a, CCR^{+}17, CND^{+}18, CATB19, DNSS11, GRM16, HSH^{+}10, HXL11, HC17, HLF14, KW14, KLLM12, LAAM11, LLQ^{+}14, LQD^{+}16, LJ13, LJVJ18, MTFK19, MAHD18, Pom12c, QLR^{+}11, RS17, Rus13, TX16, TM18, THGT13, WRW16, XP10, XYHD17, YZHX12, YCCJ15, YLGD19, ZYC16]. Computational [BR13, CC16, CFW14, DFP^{+}13, FTP13, HCL^{+}14, HMZ^{+}14, KN11a, LLW^{+}18, MOS14, RMC^{+}15, SH12]. Computations [HTA17, KHPP16, RT14, WL13]. Compute [DS14, WGR^{+}14]. Compute-Intensive [WGR^{+}14]. Computer [Ano10c, Ano10d, Ano10e, Ano10b, Ano11e, AHI12, BCS11, BdD19, GM11, HMO^{+}17, LNCX18, MFT^{+}17, MOS14, NST14, NVB16, SLPB18, TJH^{+}15, TSK16, Yam10, YMG15]. Computers [Ano11g, Ano11h, BD15, CGT^{+}15, CG18, Li12b, Liu11, Ano15a, Ano16a, Ano17a, BKP16, Ano18a, Ano19a, CHTD19, DPO17]. Computing [AKL18, AXS^{+}10, Ano11c, AISA16, BDDL18, CLX14, CHL17, CLOL18, CAGM14, DPO17, EM12, FMTK19, FJA^{+}17, GVPS19, HV16, HLA^{+}17, IBH^{+}13, JAJK15, JKMR11, KFB^{+}15, KYZC19, LT14, LXJD15, LLC^{+}15, LZZ16, LQW^{+}17, LMT13, MSG14, MGdC^{+}18, MFT^{+}17, MSC12, MLOL15, MCXZ18, PDXZ13, PLP^{+}13, PR14, QWB^{+}13, RMB^{+}13, RSNK17, RSN^{+}18, RM15c, RD18, SMTK12, SDMM12, SYD18, SG12, SG13, SCSL12, TLH^{+}16, VIDH19, WCH^{+}15, WHBR16, WRW16, XTF^{+}12, YLA^{+}15, ZGG^{+}16, ZYHZ16, ZJL^{+}16, ZLX^{+}16, ZOK^{+}19, ZV14, ZLYS15, Ano13d, Ano13c, DPO17]. Concatenation [Pom12a]. Concentrate [WJX^{+}19]. Concentrators [RO11]. Concept [ZZ19]. Concertina [FSGAB^{+}16]. Concise [LJG^{+}19]. Concrete [BS14]. Concurrency [RWC18, ZYW^{+}16]. Concurrent [BMP^{+}10, BPC12, CT13, GRM16, GSX^{+}13, GTRMG18, HRM11, KMJ^{+}11, KT19, LLL15, MLW12, MKFM13, MG11a, MKRM10, MKRM11, PWTS16, PRBM13, XYHD17, ZYW^{+}16, ZCL^{+}16, ZZM^{+}15]. Condition [YSLL16]. Conditional [CCH15b, DCCK18, HFZ13, HBCC13, HK13b, HTC13, LH12a, LKT13, MBB^{+}17, MY10, XJW^{+}16, ZGW14]. Conditional-Diagnosis [CCH15b]. Conditional-Fault [LKT13]. Conditions [JGG^{+}14, KN12, RDEN10, SMB^{+}15]. Confidence [NL16c]. Configurable [CLM^{+}19, RSJR17, SKH16, VMB19, WHYS16, vdBGLGL^{+}16]. Configurable-ECC [CLM^{+}19]. Configurations [SB16]. Confirmation [MTFK19]. Conflict [CFL^{+}18, RXC^{+}15]. Conformal [FGS^{+}15]. ConformalALU [FGS^{+}15]. Congestion [BKV12, FBR^{+}12, JRW^{+}14, JRS^{+}15]. Connected [Amm14, Ano11i, DSPB13, Gor14, SKEB16, SG12, SG13]. Connecting [LY11]. Connection [JWC12, JRC14, SMG14]. Connection-Aware [JWC12]. Connectivity [AD10, CTH14, RRS^{+}16, YL14, YLA10]. Connectivity-Guaranteed [RRS^{+}16]. Conquer [CK15]. Conscious [PB16]. Consensus [LLKA19]. CONSER [MBM11]. Conservation [LBWH11, YZHX12]. Conserving [LSL15, SC18]. Considerations [BF19]. Considered [SL13]. Considering [GSK12, HL10b, LKLM15]. Consistency [AD14, CWX^{+}14, HCC^{+}12, LXL^{+}13, LCX^{+}16, LSGZ16, SLC15a, WHL17, WFT^{+}19]. Consistent [BMS12, CLCS19, RWC18, RBIQ15, SJD^{+}18, YWW^{+}16]. Consolidated [CGJ^{+}10, JJK^{+}11]. Consolidation [MJW^{+}14]. Constant [GSF^{+}10, KRR^{+}18, KHPP16, KHZ17]. Constant-Time [KRR^{+}18]. Constants [UdDG^{+}17]. Constrained [CZP^{+}16, FK15, GDY15, KM11, Li12b, LTVL15, MMH14, SN16, SDP11, TLZV11, TCHL18, TLGM17, WMW12, WJL^{+}12, XTF^{+}12, YZH^{+}15, ZHM14]. Constraints [GHK15, GZB^{+}15, WLYY16, ZL15]. Constructing [ALZ16, GFAM11].

Construction [KLT16, MM17, NZC11, SJS10, WCL^{+}18, XHX^{+}17, ZWX12, ZL19]. Constructions [AP14]. Consumer [KSEG15]. Consumption [AO12a, CGJ^{+}10, Dar15, HT12, KLK17, VPS^{+}12, Yan14]. Contact [WW14]. Contagion [KKH^{+}14]. Containers [CT13]. Containing [FFL18]. Contemporary [ZZ10]. Content [ALBP14, CWZC13, FSGAB^{+}16, KL16, MCC12, MBS^{+}12, PO13, SMRM17, SKH16, THM^{+}14, WLT^{+}16, WBGI19, WLM15, ZFJ^{+}17]. Content-Addressable [ALBP14, SMRM17]. Content-Aware [SKH16]. Content-Based [WLM15]. Content-Centric [WLT^{+}16]. Contention [BD15, BPT10, CA12b, CG18, FJA^{+}17, GGSPM18, KCRG14]. Contention-Aware [GGSPM18]. Context [FFCB14, SRK^{+}17, YHML16]. Context-Awareness [YHML16]. Contiguous [CH14, PKY19]. Continued [Bra10]. Continuous [CCV^{+}11, MSKRJ17, RCC14, WXS12, YCL^{+}12]. Contributory [WQZ^{+}16]. Control [ABEP16, BIP^{+}17, BGMR13, BDB18, CYCC11, CBTU14, CP10, DSR15, DZD^{+}16, DRS^{+}16, HCZW13, HDYS16, JRS^{+}15, KMLH11, KKY^{+}16, LZ15, LZZV16, MWW14, MWLJ15, MBM11, MBD11, MD13, NZLK14, NC11, RSNK17, RSN^{+}18, SCK10, STR15, TLH^{+}16, TSK16, VA11, WMW12, WHZ^{+}15, XKT^{+}15, YTND12, YLA^{+}15, YAGB17]. Control-Driven [DZD^{+}16]. Control-Flow [STR15, VA11]. Controllable [Hie11]. Controlled [ASTU10, PdG13]. Controller [JSC^{+}17, MKT^{+}11, NKEM11, Pan16]. Controllers [EE17, LMNP11, MKFM13, ZJH^{+}14]. Converge [ZMW^{+}19]. Converge-Diverge [ZMW^{+}19]. Conversion [ADJ12, BZ14, BJ10, LJ15]. Converter [CHCK12]. Convolution [DWZ18, LJG^{+}19, RBMO11, ZYY18]. Convolutional [CDK^{+}18, HHKW12, KBO^{+}19, LYL^{+}19]. Convolutions [LPL12]. Cood [SMTK12]. Cool [CZ14]. Cooled [SVAB14]. Cooling [ASS^{+}18, HV14a]. Cooperation [CWZ11, OKY^{+}19, YZ15]. Cooperative [CWX^{+}14, CWY13, JZLD10, KJL11, LGH15, QSYS16, SL13, SKM14, XWL^{+}16a]. Coordinated [LZ15, XLL^{+}18, ZYL15]. Coordinating [DSW^{+}14]. Coordination [CB15, LBS15]. Coprime [GL19]. Coprocessor [CWZC13, FGS^{+}13, FGS^{+}15, KCS14, NVB16]. Copula [SD14]. Copy [DW10]. CORDIC [PP16, RS10, SR14, VVMAZ12]. Core [BLKM^{+}18, BD15, BCD^{+}16, BBB^{+}17, CZ14, CvdBC18, DYW15, DYHX16, DMK^{+}15, EF12, GP14, HRM^{+}16, IHR^{+}16, JAD^{+}18, KIJ14, KAH19, KKC17, LKH16, LRP^{+}18, LR18, LB13, MB16, Pan16, PCLN15, PBE17, PM14, RVC^{+}15, RTL^{+}18, SJVR19, SNM16, TFCY16, VTA16, WSL^{+}18, WTZ^{+}19, WhCCC12, YSLL16, YYP^{+}16, ZLBB19, ZZ19, ZCY^{+}16]. Core-Level [YSLL16]. CoreRank [YSLL16]. Cores [CCK^{+}16b, HMR^{+}17, IPS17, LKS^{+}14, LPD^{+}16, MMP13, OCK17, RRK11, WSL^{+}18]. coreSNP [GAC14]. Corner [PMH^{+}14]. Correctable [MAD14]. Correcting [FKMK16, NL15b, NL16a, RV13, Red14, SBB18]. Correction [ABSK15, CJA^{+}16, DHM16, DRM16, NL16b, NL18, PO13, PROM15, RBMO11, RLRL19, TC16, WZL^{+}17]. Correctly [BHR17, KLLM12, Lef17]. Correctly-Rounded [BHR17]. Correctness [CL10]. Correlation [SKZS13, SD14]. Correlations [LR10]. Corruptions [LS10b]. Cosine [RMC^{+}15]. Cost [AH13, BR13, BCK^{+}16, CMLS15, CCE^{+}18, CJA^{+}16, CPL17, COLK18, DVUS14, GZG^{+}16, GCL^{+}13, HS18, HK15a, HMS^{+}12, HLT^{+}15, JT15, JK15, KS14, KO14, KLT16, LYH11, LK15b, LXJD15, LOC^{+}16, MKFM13, MAD14, MUMB11, ORBM13, OGH^{+}14, SC11, SP12, TKT16, WCLY16, WFT^{+}19, YCLH16, YTD^{+}17, YLY15b, ZC13]. Cost-Aware [CPL17]. Cost-Based [OGH^{+}14]. Cost-Effective [BCK^{+}16, CCE^{+}18, GCL^{+}13, HS18, HLT^{+}15, MUMB11, YTD^{+}17]. Cost-Efficient [JK15, LYH11, LOC^{+}16]. Cost-Sensitive [KS14]. Costs [BTW13, CYC11, KHPP16, WLYY16]. COTS [BBP^{+}13, HHC^{+}18]. Count [RC14, MCT19]. Counter [EE10]. Counterexample [LH11]. Counterexample-Guided [LH11]. Countering [AS14]. Countermeasure [MLW12]. Countermeasures [BRN^{+}15, BMZ17, GSF^{+}10, YZF^{+}10, ZMB18]. Counters [DJN17]. Coupled [DMK^{+}15, PBL16]. Coupling [TMS^{+}14]. Cours [RCFP^{+}12]. Cover [XLW14, XLL15]. Cover1 [Ano12c]. Cover2 [Ano12d]. Cover3 [Ano12e]. Cover4 [Ano12f]. Coverage [AD10, AD12, BKH^{+}13, CYHC14, DLL^{+}12, GLTC16, Pom12a, Pom15b, SBH11, WXLL13, WXLY15, WCLY16, XCW^{+}10, YASS14, ZLH^{+}15]. Coverage-Preserving [GLTC16]. Covered [Amm14, Yun12]. Covering [YHH^{+}12]. Covers [KP13]. Covert [LMB^{+}16, LFJ^{+}13]. CPS [ZGB^{+}15]. CPU [AF14, CLOL18, GD17, Jun16, KkC15a, LLD19, LMC^{+}12, WGLL13, XYF^{+}15, ZYW^{+}16]. CPU-Budget [AF14]. CPU-MIC [XYF^{+}15]. CPU/GPU [GD17]. CPUs [MHRARG^{+}14, MB16, YLML15]. CRAM [ZCZ^{+}19]. Crash [WFT^{+}19]. CRAT [XLL^{+}18]. CRC [GRM16]. Creation [DRC14]. Credit [KP15]. Credit-Based [KP15]. Criteria [KKT15, Tse12]. Critic [ZM17]. Critical [ARGT14, BM13b, Ibr16, NL18, ST11a]. Criticality [BBD^{+}12, BDBB18, CAbZM18, CGL^{+}18, GGA^{+}17, LRP^{+}18, LLX^{+}17, LGS^{+}18].

Cross [CZL^{+}17, JRJ^{+}18, KCW^{+}17, RCK^{+}16, SVD18, VSC^{+}19, ZLH^{+}15]. Cross-Layer [KCW^{+}17, RCK^{+}16, SVD18, VSC^{+}19]. Cross-Level [JRJ^{+}18]. Cross-Platform [CZL^{+}17]. Crossbar [BGMR13, JWC12, PVKA14, RO11, ZMW^{+}19]. Crossbar-Based [BGMR13]. Crosstalk [CCH^{+}15a]. Crowdsensing [HZL^{+}16]. Cryptanalysis [Bar16]. Cryptographic [ARH^{+}18, BKL^{+}13, CATB19, HSA14, LLK18, SEY14, XZL^{+}19]. Cryptographically [MC11, NDG^{+}17]. Cryptography [BF19, BJ10, Cil11, GT19, HKR^{+}18, KAK18, LGH^{+}17]. Cryptography-Related [Cil11]. Cryptoprocessor [GV14, SWM^{+}10]. Cryptosystem [SWM^{+}10]. Cryptosystems [ADI11, MEBS17, PSM17, ZSP^{+}19]. CSDA [Ano10d]. CSDP [Ano10e]. CSMT [GSL10]. CTDaaS [DHC^{+}16]. Cube [AH10, HIJ^{+}19, HLJ14, SJS10, TIHM18, YWM19]. Cubes [FG10, HK13b]. Cubing [AH10]. Cuckoo [PRM16]. Current [PdG13, SRCK10]. Currents [GSK12]. Curve [ARM15, ADI11, BJ10, GKB^{+}10, GT19, LGH^{+}17, LJL13, NR15, ZWC^{+}18]. Curves [ADJ12, AK14, BDE^{+}11, CMRH17, DJJ^{+}08, FVV12, LT14, Lee12, TX16]. Custom [LSC11, LMB17, ÖDSS17]. Customized [SDMM12]. Customizing [HMD^{+}17]. Cut [LXK12]. Cyber [HWSN15, LcZLG19, SLC^{+}15b, YLY^{+}15a]. Cyber-Physical [HWSN15, LcZLG19, SLC^{+}15b, YLY^{+}15a]. Cyberspace [YGS15]. Cycle [GHG^{+}14, Iko15, LCH13, XLL15]. Cycle-Accurate [Iko15]. Cycle-Efficient [LCH13]. Cycled [WCM^{+}16]. Cycles [AB16, HBAD14]. Cyclic [SN16]. Cyclic-Random [SN16].

D [HWG^{+}14, KAH^{+}15, LJY^{+}15, TMS^{+}14, WJY^{+}17, WZL^{+}17, ZDYZ14, AD10, ASS^{+}18, AVS^{+}14, CMB13, CCK^{+}16a, CCL^{+}18, CWTT13, CND^{+}18, DYW15, EDL^{+}14, EYBK15, FGS^{+}13, HCY18, JSC10, JKD^{+}19, KKC15b, LJY^{+}15, LLW^{+}17, MWWT13, MSK15, PP16, RVL^{+}14, RKZ16, SKEB18, SPC^{+}18, TLP18, TCHL18, XCF16, YEY^{+}16, YMG16, ZLN11]. D-MAPS [KAH^{+}15]. D-Memory [HCY18]. D-Mesh [RKZ16]. D-NoC [SKEB18]. D-Stacked [SPC^{+}18]. D/ [LJY^{+}15]. DACO [Tho12, LS10a]. DaDianNao [LLL^{+}17]. dAElite [SMG14]. DaemonGuard [SNM16]. DAGs [SF17]. Damage [SPC^{+}16]. Dapper [XHLW19]. DARE [XJFT16]. Dark [EKA17, HMR^{+}17, PKC^{+}17, WSL^{+}18, WTZ^{+}19, khR^{+}18]. DART [WLV^{+}14]. Data [AD14, AQPMS15, AMG17, Ano13f, Ano13g, BLKM^{+}18, CLS14, CCV^{+}11, CT13, CDQB15, CMLS15, CHK10, CCW^{+}10, CLW^{+}15, CPL16, CPL17, CLOL18, CLW16b, DHC^{+}16, FFCB14, FLS16, GWMB13, GYC^{+}16, Ged14, GAFN15, GZB^{+}15, GZG^{+}16, GCL^{+}13, GY15a, GDY15, GY15b, GAC14, HXVQ15, HSM14, HNV19, HWSX17, HS18, HK17, HHY11, HLJ14, HLF14, HLT^{+}15, HHW^{+}18, JGG^{+}14, JSE14, JP13, JRS^{+}15, JCM16, JC11, KTAvdS16, KGV16, KLKL13, KP15, KLK17, KLC18, KRP18, KN11a, KN11b, KYZC19, KKT15, KLT16, LBSK17, LHC^{+}14, LK15b, LKLM15, LKK^{+}17, LS10b, LWW11, LMJ14, LXL^{+}14, LW15, LSHC15, LYY16, LDMQ16, LLXC16, LWF^{+}17, LcZLG19, LTP^{+}14, LZYL13, LRY^{+}15, LLS^{+}16, LWF13, LLHC15, LGF^{+}15, Man16, ML18, MJWT16, MBGS10, NCD^{+}17, OYP^{+}18, PZZQ19, PWTS16, PP10, PSM17, Pom12c, QQW^{+}17, RHC^{+}14, Red11, RWZZ14, RLX15, ROGHNB^{+}18, SMRM17, SST12, SMTK12, SHGW15, SKC^{+}14]. Data [SMK^{+}16, SLZX15, ST12, SZW^{+}16, SBW^{+}16, SHH^{+}16, UMN18, UVL^{+}13, VPS^{+}12, VCSG^{+}19, VCG^{+}12, WZZ10, WXS12, WHZ^{+}15, WLYY16, WSZ^{+}16, WBGI19, WSXZ13, WJM15, WAK^{+}17, XJFH15, XJFT16, XLS^{+}12, XHZC16, XDZ11, XLF15, YY10, YCL^{+}12, YLY^{+}15a, YCK16, YXZZ14, YWQX15, YHT^{+}16, ZR15a, ZZS10, ZZL14, ZWH^{+}15, ZDP^{+}15, ZYC16, ZLX^{+}16, ZFJ^{+}17, ZQZ^{+}19, ZMY11, ZY12, ZMW15, ZWD^{+}16, ZRL15, ZWW19, dRV12]. Data-Allocation [UMN18]. Data-Classifiers [KGV16]. Data-Dependence [KLKL13]. Data-Driven [PP10]. Data-Flow [CCV^{+}11, GAFN15, MBGS10]. Data-Intensive [WSZ^{+}16]. Data-Mining [SKC^{+}14]. Data-Parallel [ML18]. Database [CLW^{+}16a, DYCG16, KSS12, WLC^{+}15, WCH^{+}15, WCL^{+}18, XTW15]. Datacenter [ZWL^{+}19]. Datacenters [CCC^{+}17, CRJZ16, TJX^{+}17]. Datapath [ABSK15]. Datapaths [GAFN15, RD18]. Datasets [SSGB19, YLY15b]. DC [CJK19]. DC-PCM [CJK19]. DCCS [HK15b]. DCT [AKL18, CND^{+}18, XMH13]. DDoS [BPBBL13]. DDRx [FZL^{+}14]. Dead [CSJ^{+}11]. Dead-End [CSJ^{+}11]. Deadline [Bin15, BGRH15, HCH15, LXL^{+}13, LDL^{+}17, YZH^{+}15, ZWLS15]. Deadline-Constrained [YZH^{+}15]. Deadline-Floor [BGRH15]. Deadlock [DSPB13, FG10, RRS^{+}16, VYEB17, WL13]. Deadlock-Free [DSPB13, FG10, RRS^{+}16]. Deallocation [PCLN15]. Deanonymization [PLZW14]. Debiasing [USH19]. Debug [COLK18, DN11, DRS^{+}16, MVB10, OHCK17, OCK17, PBV11]. Debugging [ABSK15, CCL^{+}13, FM19, KN12, NZ14, WhCCC12]. Decentralized [DNSS11, HGML11, RVH^{+}16, SPC^{+}16, YMK^{+}17]. Decimal [APP12, BZ14, BLMM16, CHCK12, CDL^{+}17, GVGNCVM16, GNTS13, GJ15, HK13a, TGNSC11, VAM10, VVMAZ12, WF17]. Decimal-Based [GNTS13, TGNSC11]. Decision [AXS^{+}10, CJSM17, CCO^{+}14, CW15, MRW^{+}15, SDP^{+}15, SJS10]. Decodable [NL15b, NL16a, RLRL19]. Decoders [CMM15]. Decoding [LLC^{+}16, NL18, YCW11]. Decomposable [BGM^{+}13]. Decomposer [WDSP12]. Decomposition [GAFN15, JC12, KEK16, KYZC19, LVMS18, LZZ17a, XHZC16]. Decoupled [PVKA14, SCJ^{+}16a, XYHD17]. Decoupling [STR15]. DECS [CTS13]. Dedup [WWY^{+}18]. Deduplicatable [HCD^{+}16]. Deduplicating [LLXC16]. Deduplication [FHL^{+}18, LCH^{+}15, MJWT16, MYW11, WWY^{+}18, XJFH15, XJFT16, YTD^{+}18, ZFJ^{+}17]. Deduplication-Aware [XJFT16]. Deep [CDK^{+}18, FMTK19, LW13, LPCW14, MFT^{+}17, SSGB19, SJVR19, SSJ^{+}18, ZYC16]. Defect [BKP16, CSW^{+}15, Pom16b]. Defects [BGPV10, HTH15]. Deferrable [LHC^{+}14, RHC^{+}14]. Defined [HGL^{+}15, KKP^{+}16, LSHC15, WLJ^{+}16, ZLG^{+}15, ZGG^{+}16, ZFJ^{+}17]. Degradation [HGW^{+}17, ORBM13]. Degraded [FSL^{+}17, ZLLX15]. Degree [CMRH17, Ste14]. Degrees [Cil13, UHSA17]. Delay [CA12a, CFMS14, DY12, GDY15, HCZW13, JRW^{+}14, LYT^{+}16, MOYB12, NI11, NL14, OMFH14, PROM15, PSL17, RSNK17, SXLC15, WGZ^{+}15, XTF^{+}12, ZRS^{+}16]. Delay-Constrained [GDY15]. Delay-Insensitive [OMFH14]. Delay-Tolerant [CFMS14, HCZW13]. Delays [GSK12]. Delivery [BLKM^{+}18, FS10, KL16, RS13]. Demand [CAbZM18, CQW^{+}15, CJG16, DYW15, LLLJ13, WZY16, WZ14, XLF15]. Demand-Supply [DYW15]. Demands [LLLJ13, ZJL^{+}16]. Demarcated [YMK^{+}17]. Demotion [LHTG15]. Demotions [LWH^{+}16]. Denial [TJH^{+}15]. Denial-of-Service [TJH^{+}15]. Denoising [LHCL13, dRV12]. Dense [JAKD18].

Density [KPS^{+}17, KCE^{+}18]. Density-Aware [KPS^{+}17]. DEP [ASE17]. Dependability [CCD12, RCK^{+}16]. Dependable [Ano10c, GM11, GFAM11, IS11, IBH^{+}13]. Dependence [KLKL13]. Dependency [JLMH10]. Dependent [AKL14, JCY^{+}13, KL13, LR13, ZM17]. Deployability [TC14]. Deployed [WLJ^{+}16]. Deploying [BWCW15]. Deployment [SZS14, XLW14]. Derivation [YLH13]. Derived [DRM16]. Deriving [CCK10, XXBL17]. Descriptions [BFP11]. DeSIgn [GEN^{+}17, ABB17, ARH^{+}18, ACW^{+}11, AD16, AS10, ABEP16, Ano11f, BKL^{+}13, BS10, BMS11, CCO^{+}14, CHH^{+}13, CKKS14, CCC15, CHLL16, CYA13, DCY^{+}13, DZLP14, DJO11, EKA17, FFISC13, FAK16, FGS^{+}13, GH11, GJ14, GZC^{+}17, GEvS10, GSX^{+}13, GTRMG18, GSF^{+}10, HFG^{+}17, HHM11, HSA14, HMC11, HHCH11, HKWC14, HWK15, HMS^{+}12, JAKD18, JKY10, JKD^{+}19, JAD^{+}18, Jun16, KSS12, KC14, KW14, KAH^{+}15, KKY^{+}16, KCL^{+}16, KH18, KKS14, LYB15, LH16, LJ18, LCL17, LTLC12, LCW^{+}16, LOC^{+}16, LQW^{+}17, LCY^{+}19, LVF19, LLOS13, MJW^{+}14, MMCS18, MSK15, MOMT12, MRL^{+}18, MLE14, MHML15, MF14, MCT19, NBZP17, PC16, PBT13, PR14, RQ14, SBP^{+}14, SVD18, SCZ^{+}16, SJD^{+}18, SRR^{+}16, SJS^{+}14, ST11a, SZZ^{+}19, SVAB14, SZDL14, TAH^{+}16, TS11, VPS^{+}12, VSLD15, VCSG^{+}19, VAM10, VALK19, VKS^{+}16, VD12, WZBB15, WLT^{+}16, WKB16, WEH^{+}19, WSXZ13, YSZ^{+}14, YCCWC15]. Design [ZCZ^{+}19, ZD13, ZL18, ZV14, ZMS13]. Design-Exploration [SVD18]. Design-for-Test [WEH^{+}19]. Design-Space [JAD^{+}18]. Design-Stage [TS11]. Designed [LS10a]. Designing [AO12b, CWZ13, FBWMM13, HK16, HHY11, LMB^{+}16, LCHX11, Red18].

Designs [ABSK15, AS12, AS14, BKH^{+}13, CFR^{+}14, CCAM14, FML10, JLLH19, KAQC14, LLC^{+}16, LKS^{+}14, NS13, PSL17, TGNSC11, VTA16, WZCG16, ZYY10, ZZ10]. Destination [TC14]. Destination-Oriented [TC14]. Detailed [Fin10]. Details [Bai17]. Detect [LXK12, OWP16]. Detectable [KT19]. Detecting [EBE13, GDJZ18, KW14, Red18]. Detection [AHNT16, CVMA10, CJK19, CJ12, CH14, GRM16, GTRMG18, GBF^{+}19, HRM11, HHC^{+}18, HBR11, KMJ^{+}11, KC13, KT12, MLW12, MKFM13, MKRM10, MKRM11, NDC^{+}13, OHCK17, OCK17, ORBM13, OKD^{+}16, PNKI13, PO13, PRBM13, PBT13, PMH^{+}14, RBK^{+}12, RSU17, RBMO11, SPC^{+}16, SRR^{+}16, ST12, TJH^{+}15, TC16, TM18, VA11, VSF^{+}17, WF14, WhCCC12, XJFT16, XCW^{+}10, YHML16, ZZ17, ZYW^{+}16, ZLN11, ZCR16]. Detector [Hia16]. Detectors [NY15, UVG16]. Determination [BBK10, KN11a]. Determining [ZRS^{+}16]. Deterministic [AK16, CB15, KN11b, RTL^{+}18, YZGG16]. Determinizing [CCL^{+}13]. Development [BCD^{+}16, MOS14, SAR^{+}11]. Device [DA12, JKJ^{+}10, JW16, TLB^{+}17]. Devices [CHTD19, CXLX15, CKH15, CYL^{+}14, DPO17, JRP^{+}14, KCRG15, KkC15a, LK14, LZZ^{+}17b, LLD19, OGH^{+}14, SWF^{+}19, SYH17, SHH^{+}16, TCYH15, WKB16, WW16, YCKH16, YCK16, ZLW^{+}17]. DeyPoS [HCD^{+}16]. DFA [LPCW14, PDJ^{+}19]. DFT [CCR^{+}17, Red14, Red19]. DHT [SX12]. Diagnosability [Cha10a, CL12, CH13, HFZ13, HK13b, HTC13, LXZH16, ZLXW15, ZGW14]. Diagnosing [Li12a]. Diagnosis [AKL14, AD16, BC16, BGPV10, CH11, CCH15b, HK13b, HWL^{+}14, LVMS18, LKT13, PB16, PR10, Pom16b, SDE^{+}17, TW10, TLL12, Tsa13, YLL16]. Diagram [CJSM17]. Diagrams [AXS^{+}10, SJS10]. DIALIGN [BCMJ10]. Dickson [HN11]. Die-Stacked [ZDYZ13]. Difference [BS14, BS16]. Different [CLM^{+}19]. Differential [Bar16, CTL^{+}17, KCW^{+}17, LYK19, LSGZ16, MSS17, SBM15]. Differentially [ST18a]. Differentiated [CCM14, ZLN11]. Differentiation [WMW12]. Diffie [FHLOJRH18, LNL^{+}19]. Digit [CLL^{+}14, DALD18, EJ15, ERRM16, FBE^{+}18, JPG10, Kor15, RM15b, Rus13, SJS^{+}14, TC16, TAM^{+}16]. Digit-Level [ERRM16]. Digit-Serial [CLL^{+}14, RM15b]. Digital [BIP^{+}17, CK15, KBP13, MG11a, NC11]. Digraph [ZLBB19]. Dilation [BKV12, KTA^{+}14]. Dimension [SBI12]. Dimensional [AMVOS^{+}15, AVS^{+}16, CCE^{+}18, JWH^{+}15, KLC18, MEBS17, MKLW14, Pom15c, Ste14, TWTT11, WEX14, Zot10]. Direct [IRMM^{+}16, VIDH19]. Directed [CM11, CVH^{+}13, LTLC12, OKC13, THGT13]. Directional [SZS14]. Directories [LFH^{+}16]. Directory [CRG^{+}13, LCC10, ST17, ST18b]. Dirty [LKLM15, ZZJ^{+}19]. Disaggregated [KKJH19]. Disaggregation [SHH^{+}16, TJX^{+}17]. Disaster [GSG^{+}15, MTFK19, WYL^{+}15]. Disciplined [LWKA15]. Disclosure [FRB^{+}18]. Discontiguous [CH14]. Discovery [MCC12, SH12, WYL^{+}15]. Discrete [HKR^{+}18, HHCH11, KRR^{+}18, LZZ17a, TWTT11, YFJ^{+}14, CTS13]. Discrete-Error-Checking [CTS13]. Discriminative [Ged14]. Disjoint [AB16, HBAD14, KP13, LY11, SKA10]. Disjunction [AD16]. Disk [KBH^{+}10, LS10a, LBWH11, RLSK18, SNY^{+}10, Tho12, VC10, WZW^{+}19, XLX^{+}14, ZLWZ15, ZXX^{+}14]. Disk-Based [WZW^{+}19]. Diskless [HC13a]. Disks [Cha10b, JR17, KSJ^{+}12]. Disparity [THGT13, TKT16]. Dispatch [JR17, LBS15]. Dispatching [YCLH16]. Dispersal [CWL^{+}17]. Dissemination [DKH^{+}13, WAK^{+}17, XAYL15, ZWD^{+}16]. Distortion [dRV12]. Distributed [ASTU10, BBPQ15, CWZC13, CPL17, DZD^{+}16, DLL^{+}12, DSPB13, DKK16, FEP^{+}12, FSR^{+}18, GBO^{+}16, GZB^{+}15, GZG^{+}16, GW16, HRK17, HSM14, Hie13, HCH15, HC13b, HXL11, HHW^{+}18, LLCH13, LCL15, LS13, LCH^{+}15, LcZLG19, LXK12, MM16, MB12a, SSKL16, SH12, SLC^{+}15b, SWF^{+}19, TLZV11, TPR16, WZK^{+}19, WHL17, XMH13, YY10, YZ15, YTD^{+}18, ZLN11, ZT15, ZCR16, Zot10]. Distributed-Block [WZK^{+}19]. Distributed-Memory [GBO^{+}16]. Distribution [CLS14, HCY18, HS18, LHYZ13, RNS13, THM^{+}14, VC10]. Distribution-Aware [HS18]. Distributor [Zhe10]. Disturbance [CCK^{+}16a, CJK19, JSC^{+}19]. Diverge [ZMW^{+}19]. Diverse [LWY15]. Diversity [JRC14, WGLL13]. Divide [CK15, XLW14]. Divide-and-Conquer [CK15]. Divide-and-Cover [XLW14]. Divider [JLLH19]. Dividers [AS10, CHLL16, KKS14]. Dividing [Yan14]. Divisible [CC16, ZR15b]. Division [BZ14, CTS13, EJ15, FBE^{+}18, GABK11, HIJ^{+}19, KS10b, LN12, MPZ15, MG11b, Nan16, ST18a, TIHM18, UdDG^{+}17, WE12, YWM19, ZMY11]. Division-Free [BZ14]. Division/Square [WE12]. DMA [VCG^{+}12]. DMR [RBMO11]. Domain [HXVQ15, VKS^{+}16, ZCW18]. Domains [CWZ13, LQD^{+}16]. Dominance [PR10]. Don’t [DJN17]. Dot [KKS14]. Double [ARM16, AMR18, ARM13, AK14, BNP10, CS11a, CLL^{+}14, DRM16, DJA14, DS14, ERRM16, HK15b, LKLT12, MH15, RLRL19, RM15b, ZGWC15]. Double-Loop [CS11a]. Double-Ruling-Based [LKLT12]. Double-Scalar [DS14]. Doubling [SBI12].

Doubling-Dimension [SBI12]. Down [LRY^{+}15]. Downloading [LHH14b]. Downtime [DSY^{+}15]. DPA [BK12, GSF^{+}10, LRY^{+}15, MM17]. DPPC [MWW14]. DR [CLW^{+}19]. DRAM [ACM^{+}16, BCC^{+}16, CLW^{+}19, CJA^{+}16, FZL^{+}14, GC16, HK15b, Iko15, JYL^{+}17, KCKL19, LBN14, LK16b, LZZZ13, LHTG15, LWH^{+}16, OCK17, RSJR17, SMA19, SCJ^{+}16b, SCJ^{+}16a, SJC^{+}17b, SPC^{+}18, SD14, WZL^{+}17, WBGI19, ZZ10]. DRAM-Based [OCK17]. DRAM-Latency [SCJ^{+}16b]. DRAM/PRAM [HK15b]. DRAMs [SSJ^{+}18, ST16]. DRINA [VBR^{+}13]. Drive [KLK18]. Driven [AD13, BR13, BM13a, BM13b, CMS10, DZD^{+}16, HWX15, LK15b, LGMP10, PP10, PCZB11, RM15a, SAR^{+}11, TLZV11, TS11, YTND12, YHV13]. Driver [JW16]. Drivers [JKJ^{+}10]. Drives [CDQB15, CLZ19, DSW^{+}14, Jun16, LRP16, MLE14, PDXZ13, TAH^{+}16, WFY^{+}17, XWL^{+}16b, YCW^{+}19]. Drone [WAK^{+}17]. Drought [JGG^{+}14]. DSRC [YMT13, YMTV14]. DT [SDP^{+}15]. DT-CAIF [SDP^{+}15]. DTNs [LS13, YZH^{+}15]. Dual [GSF^{+}10, GCL^{+}13, HGW^{+}17, KAH18a, KwPK^{+}15, LPL^{+}13, LW15, LXW^{+}19, LW13, PPND17, YTD^{+}18]. Dual-Addressing [LXW^{+}19]. Dual-Clock [KAH18a, PPND17]. Dual-Level [HGW^{+}17]. Dual-Modular [KwPK^{+}15]. Dual-Phase [YTD^{+}18]. Dual-Port [GCL^{+}13, LW15]. Dual-Rail [GSF^{+}10]. DuCNoC [KAH18a]. Due [MD16]. DUOS [BSS14]. Duplication [WMG18, ZMRQ11]. Durability [CDQB15]. During [LS10b, UMN18, DN11, KN12, XXBL17]. Duty [GHG^{+}14, WCM^{+}16]. Duty-Cycled [WCM^{+}16]. DVFS [ASE17, EE10, GHK15, GZB^{+}15, HV12, HV14a, KkC15a, LSC10, LY17, LY18]. DVM [MSG14]. DwarfCode [ZCS16]. DWF [LBN14]. Dynamic [ABSK15, CBB19, CLS10, CKH15, CFW14, CCP^{+}13, DCV^{+}12, DKK16, FHR14, FKMK16, FCB^{+}19, GVPS19, HCCG10, HCD^{+}16, HHY11, HH17, HV13, HV14b, HCG^{+}16, HLWV17, IHR^{+}16, JSH^{+}17, JCM16, JR17, KAH19, KLKL13, KCRG14, KCL19, KKT15, LKYC12, LK16a, LCL15, LXDV17, LHH14a, LHYZ13, LRY^{+}15, LZA^{+}16, LZS^{+}13, LHTG15, LWH^{+}16, LPL10, MWW14, MSC12, NM10, NH10, OKC13, RBG14, RF14, RDEN10, SKZS13, SJSLD11, WSL^{+}18, WTZ^{+}19, XLF15, YZHX12, YFJ^{+}14, YTD^{+}18, YHV13, YZGG16, YLY15b, YAGB17, ZWC^{+}18, ZLN11]. Dynamically [CW15, GLXY13, KGC14, LYJ^{+}18, PPP13, RSU17, TLGM17, YSLL16]. Dynamics [JWWZ16, LLL16].

E-MACs [AP14]. E-Shadow [TZL^{+}14]. EAD [ZMRQ11]. Early [SVAB14, VSC^{+}19]. Early-Stage [SVAB14]. Easy [KT19]. ECC [CSCW13, CLM^{+}19, FKMK16, GBGI18, HK17, HCL15, PN16, PCZB11]. ECC-Based [PN16]. ECM [IDG^{+}17]. Ecosystem [Cro14]. eCryptfs [XZL^{+}19]. EDF [BGRH15, CBB19, CQ14, LXL^{+}13, SL14b]. Edge [AB16, CTH14, DHW^{+}19, HBAD14, PMH^{+}14, RSNK17, RSN^{+}18, THGT13, YL14]. Edge-Connectivity [YL14]. Edge-Directed [THGT13]. Edge-Disjoint [HBAD14]. Edges [WHC^{+}15b]. Editor [Lou19, Mon15a, Mon19]. Editor-in-Chief [Lou19, Mon19]. Editorial [BKP16, CHTD19, DPO17, Lom11, Lou19, Mon15a, Mon19, WHBR16, XL16, Zom15a, Zom12a]. Editors [AHI12, AISA16, Avr13, BKPMC13, BMM11, BS10, BCS11, BdD19, EM12, GC14, GM11, LLK18, MG11a, MOS14, NST14, ST11a, VP14, ZMS13]. eDRAM [FHW18, JJZ^{+}16, VPS^{+}12]. eDRAM-Based [JJZ^{+}16]. eDRAM/SRAM [VPS^{+}12]. Edwards [LT14]. Effect [BD15, GC16, YMG16]. Effective [BCK^{+}16, BTBB14, CXZ13, CCE^{+}18, DCV^{+}12, GCL^{+}13, HS18, HLT^{+}15, JT15, KCRG14, KCKL19, LYOB15, LY11, LK16a, LVF19, ML13, MUMB11, SST12, SL14a, SP12, WLK15, YTD^{+}17, ZLY15]. Effectiveness [CRG^{+}13, SLZX15]. Effects [AEGH19, SRCK10]. Efficiency [AEGH19, CCL^{+}18, CKH15, Fen14, IPS17, JDA^{+}16, KKC17, LKK^{+}17, LR18, LYCT10, MYHL16, SLL15, ST17]. Efficient [ALBP14, AO11, AYC16, ASE17, AP14, AMG17, ASBdS16, Ano11c, BBPQ15, BS15, BSS15, BPG16, BBB^{+}17, BBH12, CVPS19, CFR^{+}14, CHN14, CXZ13, CDQB15, CMLRHS13, CHH^{+}13, CYJ^{+}10, CM11, CS15, CZP^{+}16, CXLL16, CLCS19, CCRL19, CJ13, CDK^{+}18, CWCS15, DCCK17, DHW^{+}19, DCY^{+}13, DZD^{+}16, DKLB15, DJA11, DCK16, DCL^{+}11, DCV^{+}12, DSY^{+}15, DZLP14, DNSS11, EKA17, ECJ^{+}16, EM12, FVV12, FZL^{+}14, FAK16, FAA10, FSL^{+}17, GKB^{+}10, GH11, GBO^{+}16, GKS14, HB11, HCL^{+}14, HV14a, HNV19, HBCC13, Hia17, HMC11, HC17, HNB^{+}12, HQLX15, HDYS16, HLA^{+}17, ISC15, IDG^{+}17, IBH^{+}13, JAKD18, JK15, JP13, JC11, JJZ^{+}16, Joh17, KMC17, KJL11, KLKL13, KLJ^{+}14, KO14, Kim15, KHPP16, KBO^{+}19, KKC15b, KH14, KAQC14, KA19, KCS14, KH10, LYH11, LPL^{+}13, LSC11, LP13a, LK15b, LKLM15, LDP10, LXL^{+}13, LCLL15, LGH15, LWF^{+}17, LHCL13, LCH13, LZ14, LCT11, LN12]. Efficient [LHYZ13, LLM^{+}15, LSW15, LXZ^{+}15, LFH^{+}16, LOC^{+}16, LLD19, LKMSA16, LJ13, LJ15, LJVJ18, LSXP14, LCW^{+}15, MWZ^{+}17, MAG^{+}17, MB12a, MH15, MYW11, MBSSA19, MS12, MKRM12, MC11, ML16, NZC11, OPZ15, OPAGS14, PKC^{+}17, PP14, PCHS18, PAC^{+}12, PP10, RMKR12, RURM18, RBK^{+}12, RS17, SRCK10, SVD18, SDMM12, SJD^{+}18, SRK^{+}17, SG12, SZG^{+}18, SWZG15, TLH^{+}16, TH11, TWTT11, TLB^{+}17, TCYH15, TM18, UMN18, USH19, VALK19, VCB^{+}13, VSF^{+}17, WF17, WF12, WHZ^{+}15, WCM^{+}16, WW16, WEH^{+}19, WDSP12, WQZ^{+}16, WLG^{+}19, XL16, XMH13, XHZ14, XLTZ11, XLF15, YY10, YCW11, YMAG17, YTD^{+}17, YM11, YWQX15, YUGD14, Yun12, YYP^{+}16, ZD13, ZWX12, ZGY13, ZWW^{+}16, ZCL^{+}16, ZLJ^{+}17, ZCK19, ZFHC19, ZMY11, ZYY10, ZHW15]. Efficiently [GJ14, LGH^{+}17, OGPK14]. Effort [RMB^{+}13]. Eight [SG13]. Eight-Approximation [SG13]. Eisenstein [CCR^{+}17]. EJ [FB13]. Elaborate [SSJ^{+}18]. Elastic [CRJZ16, JT15, KKJH19, MBD11, MD13, WBZ^{+}15, YMK^{+}17]. Electrical [FM16]. Electronic [BC16]. Element [MTGM12]. Elementary [AFC10, FS10, LJ13, SDP11, dDLM11]. Elevation [CZS^{+}19]. Elevator [DSPB13]. Elevator-First [DSPB13]. Eliminating [LKBS16, UMN18]. Elimination [RKN^{+}18, XJFT16]. Elliptic [ARM15, ADI11, AK14, BDE^{+}11, CMRH17, CZ16, GKB^{+}10, LGH^{+}17, LJL13, NR15]. ELmD [BDMLN16]. Email [XJW^{+}16]. Embedded [ABB17, AAR19, ACM^{+}16, AEP18, ARGT14, BQP^{+}16, BCSR14, BM13b, BGRH15, Cha14, CSS13, CPRH16, DA12, DSB13, DLC^{+}13, EKJ^{+}10, FSR^{+}18, FGS^{+}13, FRB^{+}18, GBO^{+}16, HHLK12, HC17, HT12, HLA^{+}17, JLC10, KSS12, KMLH11, KLJ^{+}14, LSA18, LJVJ18, MW10, MS15, MUMB11, MNK11, OKC13, PAC^{+}12, PC10, PBE17, QLH^{+}16, SWF^{+}19, TB15, TKT16, VSF^{+}17, WLC^{+}15, ZGG^{+}16]. Embedding [CMRH17, CS11a]. EMC [SZZ^{+}19]. Emergence [HJBM14]. Emerging [Ano13d, Ano13c, AISA16, BMM11, CHTD19, DPO17, OOD^{+}17]. Empirical [DJO11]. Employing [MPZ15]. Empty [MWLJ15]. Emulation [AHK10, EGVFC^{+}12]. Emulation-Based [EGVFC^{+}12]. En-Route [YLY^{+}15a]. Enable [ACW^{+}11]. Enabled [KCL^{+}16, LP17, LZZ16, QZL^{+}16].

Enabling [CLW^{+}19, DYCG16, GLTC16, LLL15, SNM16, WHL17, XLL^{+}18, YHML16]. Enciphering [CMLRHS13, MLCH10]. Encoded [TAM^{+}16]. Encoder [HHCH11]. Encoders [HHKW12]. Encoding [CND^{+}18, LCA10, LSXP14, SKZS13, TAM^{+}16, XHX^{+}17, YCW11, Yun12]. Encoding/Decoding [YCW11]. Encodings [GJ15, HK13a, MVB10]. Encrypted [KGV16, LQD^{+}16, NBZP17, TM18]. Encryption [AEP18, BS14, BDMLN16, CMO^{+}16, CLW16b, DOS15, FHH10, FHR14, HZ11, HC17, JSA17, KHPP16, LLC^{+}15, LB13, MRL^{+}18, MBF18, MKRM10, ÖDSS17, PRGBSAC19, RVH^{+}16, RZZ^{+}15, SWF^{+}19, WHC^{+}15a, WQZ^{+}16, XJWW13, XJW^{+}16, XHX^{+}17, ZPM^{+}15, ZHW15]. End [CCV^{+}11, CSJ^{+}11, NLP^{+}14, SXLC15, SRCbL^{+}15, YLH10]. End-Link [SRCbL^{+}15]. End-to-End [CCV^{+}11, NLP^{+}14, SXLC15, YLH10]. Endomorphisms [AK14, LGH^{+}17]. Ends [PPB^{+}14]. Endurance [DY14, FYSK14, HJF^{+}13, JSA17, PLM16, SYK14]. Endurance-Aware [FYSK14, JSA17, SYK14]. Energy [AO12a, AYC16, ASE17, AE11, Amm14, Ano11c, AS16, BZ15, BPG16, BG12, BBB^{+}17, CFL^{+}18, CSPC12, CA12a, CTD^{+}16, CKD^{+}17, CKH15, CQ14, CKN14, CWCS15, DHW^{+}19, DCY^{+}13, EM12, FAA10, GH11, GKS14, GBD^{+}15, HWZ^{+}12, HV14a, HNV19, HT12, HDYS16, HLA^{+}17, IPS17, IDG^{+}17, JDA^{+}16, JJK^{+}11, JAJK15, JJZ^{+}16, KMC17, KLJ^{+}14, KGGJ14, KKC17, KLK17, KCL19, KSC^{+}14, KKC15b, KA19, KLT16, LK15b, LK16b, LKK^{+}17, LR18, LTL14, Li12b, LGH15, LBWH11, LHH14a, LSL15, LLD19, LZW^{+}15, LJVJ18, LBS15, LCLC19, MHK15, ML18, MWZ^{+}17, MBSSA19, MMH14, NYHB16, NY19, OPZ15, OPAGS14, OKC13, PAC^{+}12, PP10, RMKR12, SC18, SDMM12, SZZ^{+}19, TH11, TFCY16, VPS^{+}12, VSF^{+}17, WWY^{+}16, WLJ^{+}16, WCM^{+}16, XTW15, YZHX12, Yan14, ZLBB19, ZLG^{+}15, ZMW15, ZMRQ11]. Energy-Aware [CSPC12, CKD^{+}17, SZZ^{+}19, TFCY16, ZLBB19, ZMRQ11]. Energy-Balanced [LBS15]. Energy-Efficiency [IPS17]. Energy-Efficient [AYC16, ASE17, BPG16, BBB^{+}17, DHW^{+}19, DCY^{+}13, FAA10, GH11, GKS14, HV14a, HNV19, HDYS16, HLA^{+}17, IDG^{+}17, JJZ^{+}16, KKC15b, KA19, LK15b, LGH15, LLD19, LJVJ18, MWZ^{+}17, MBSSA19, OPZ15, OPAGS14, PAC^{+}12, PP10, RMKR12, TH11, VSF^{+}17]. Energy-Harvesting [AS16]. Energy-Saving [LHH14a]. Energy-Scalable [LJVJ18]. Energy-Time [BG12]. Enforcing [WZL15]. Engine [BDB18, CLS10, PWW^{+}11, SLC^{+}15b, SWZG15, VSF^{+}17]. Engineering [HGL^{+}15, LLK18]. Engines [LCHX11, LB13]. Enhance [CCY^{+}16, FSPD17]. Enhanced [CSCW13, JRW^{+}14, KLC^{+}16, ML13, SMCN18, SC11, WLS18]. Enhancement [CHH^{+}13, HCY18, JSA17, LBWH11, LLHC15, SMP16, ZZYZ14]. Enhancing [JC11, KKC17, PCLN15, SMAR^{+}19, ZOD13]. Enough [JAJK15]. Ensuring [LYY16, MHK15]. Enterprise [XZL^{+}19]. Enterprise-Level [XZL^{+}19]. Entropy [DEE17, GLH^{+}19, KKT15, LB15b]. Environment [KFB^{+}15, LDL^{+}17, WSZ^{+}16]. Environmentally [MTBB10]. Environments [BMT14, CGJ^{+}10, GW16, HCD^{+}16, HGCT13, KC13, LP13b, LWF13, MFG14, WXS12, YHML16]. EPC [SDZ15]. EQAR [LY11]. Equal [DDNT19]. Equalization [TLGM17]. Equalized [WMG18]. Equilibrium [BBVL14, Cro14]. Equivalence [PR10, ZYHZ16]. Equivalent [BFP11]. Era [EKA17, HMR^{+}17, YMG16]. Erase [CCL^{+}18, JSH^{+}17]. Erasure [CJA^{+}16, HQLX15, LS10a, LSXP14, SZG^{+}18, ZLLX15]. Erasure-Coded [HQLX15, LS10a, ZLLX15]. Errata [ZDYZ14]. Error [AS12, Bai17, BM11, CHLL16, CJA^{+}16, CZS^{+}19, CFMA19, CCAM14, CTS13, CMM15, DRM16, DZLP14, EBE13, EGVFC^{+}12, FKMK16, GPN11, GTRMG18, HRM11, HHCH11, HBR11, JRJ^{+}18, KMJ^{+}11, KW14, KTA^{+}14, KT19, LR16, LHL13a, LHL15b, LCY^{+}16, LOC^{+}16, LQW^{+}17, MLW12, MRD19, MKFM13, MAD14, MHHS17, MHH^{+}17, MG11a, NEE18, NL15b, NL16a, NL16b, NL18, NC11, OHCK17, OCK17, PNKI13, PCZB11, PO13, PRBM13, PROM15, RSU17, RBMO11, RLRL19, SB16, SBB18, SMRM17, SVD18, TC16, VTA16, VA11, WZL^{+}17, WLG^{+}19, XH16, YW12]. Error-Detection [RSU17]. Error-Tolerance [HHCH11]. Error-Tolerant [CHLL16, LQW^{+}17]. Error/Fault [MG11a]. Errors [DRS^{+}16, HCC^{+}18, JSC^{+}19, KK10, KCKL19, RV13, Red18, SMAR^{+}19, TM18, UVG16, VSC^{+}19, WBGI19, dOPSR16]. Essential [Ano11g, Ano11h]. Esterel [LvH12]. Estimate [PB11]. Estimating [Dar15, MYHL16]. Estimation [BJ12, CYHC14, DAS14, DAPS14, EE10, ISC15, LPL10, RM15c, SRCbL^{+}15, SMA19, SRCK10, SDP^{+}12, TKT16, WWT^{+}18, XTW15, YLH13]. Ethernet [CWF14, HGML11, JRW^{+}14, JRS^{+}15, SME^{+}17]. Evaluating [CPS^{+}10, LHL15b, LOC^{+}16, MBD11, RSG^{+}19]. Evaluation [BLB^{+}19, CWF14, CCO^{+}14, EGVFC^{+}12, FTP13, GEvS10, GSF^{+}10, HCL^{+}14, HHM11, HWCH17, JKMR11, JWL^{+}16, JRP^{+}14, KSS12, KCL^{+}16, KKT15, LJ18, LCHX11, MNK11, ROH17, RQ14, RJV^{+}18, dLSGDR17, ST11b, TSK16, WGW^{+}15, WLT^{+}16, WFY^{+}17, YL14, YMT13, YMTV14, ZCL^{+}16, ZWC13, dOPSR16]. Evaluations [GLH^{+}19]. Even [ARH14, PCHS18, WF12]. Even-Type [WF12]. Event [CVH^{+}13, HWX15, SMRML17, WNKL16, XAYL15, XLS^{+}12]. Event-Driven [HWX15]. Event/Multiple [WNKL16]. Events [BJ12, LBS15, MB12b]. Evidence [EFPC16]. Evolution [YZ15]. Evolutionary [AD13, HSH^{+}10, RM15c]. Evolvable [SOM^{+}13]. Evolving [HCZW13, JWWZ16]. EvolvingSpace [WZZ10]. Exact [BM11, JLMP11, MBD^{+}17, RCRK13, dLSGDR17, XP10]. Exascale [YWXZ12]. Exchange [EFGT18, FHLOJRH18, GDLL18, SD18, YRT^{+}16]. Exclusive [LSHC15]. Executing [WLY^{+}14]. Execution [ASE17, BBK10, DZ10, GLXY13, GPRS17, KLLK11, KCRG15, LK10, LKK^{+}17, LMB13, Sri10, WLZ^{+}15, WA10, XLC14, ZLSI17]. Executions [LKLK13]. Existing [FNS16, YTND12]. Expandable [GCL^{+}13]. Expansion [AVS^{+}14, RCRK13]. Expansions [JMMP16, MRD19, RMB^{+}12]. Expenses [ZMW15]. Expensive [GBGI18]. Experience [MBM11]. Experiences [LHH14b]. Experiments [BM13a, DN11, dRV12]. Experts [RF14]. Exploiting [AKKH12, CSPC12, CZ14, CCC^{+}17, CLZ19, CWY13, CYL^{+}14, CLMM11, DSR15, DSG^{+}19, EF12, FMTK19, GC16, HCY18, HIJ^{+}19, HJBM14, HK15a, HJF^{+}13, IS14, JCY^{+}13, JRC14, KWC^{+}16, LK14, LWKA15, LR16, LR10, LS13, LWF^{+}17, LJG^{+}19, LCY^{+}16, NLRB17, SSW12, SPC^{+}18, WSL^{+}18, WBGI19, WGZ^{+}15]. Exploits [GDJZ18]. Exploration [DJO11, JLMH10, JAD^{+}18, KBH^{+}10, Nan16, SMP16, SVD18, SBW^{+}16]. Explorer [SQJ^{+}15]. Exploring [Cil11, GY15a, HXL11, HJF^{+}13, Jun16, KH18, LPD^{+}16, VMB19, WHC^{+}15a, YXWL16]. Explosively [YCKH16]. Exponential [BHR17, LP17, VB13]. Exponentiation [DCCK18, ERRM16, GLP^{+}12, HMA^{+}10]. Expression [OWP16, YP12]. Expressive [LFH^{+}16]. EXR [LSHC15]. Extended [BFMT16, Hia17, JMMP16, LPCW14, LWK11, Sou15, WJL^{+}14]. Extending [FKMK16, JSH^{+}17, PPB^{+}14, RCFP^{+}12, SF17, XWLX17]. Extension [ARH14, HRM11, RCC14, Red14]. Extensions [RS17, SJVR19]. External [LBSK17, LRP16]. Extra [CTH14, YL14]. Extraction [KCK16, LJY^{+}15, VB13]. Extractor [USH19]. EXTREME [SPC^{+}18]. Extremely [MAD14]. Extremum [ZFJ^{+}17]. EZ [PDXZ13].

Fabric [GFAM11]. Fabrics [AD16, DPO17, VYEB17]. Face [FS10]. Facilitating [cCWS14]. Factor [HL10b, LLS^{+}16]. Factoring [GKB^{+}10]. Factorization [CND^{+}18, PGvdG14]. Factors [MPZ15]. Fading [QSYS16]. Fail [HCC^{+}18, IGLM15]. Fail-Stop [HCC^{+}18]. Failure [CVMA10, CSW^{+}15, HL10b, LLL15, LWL^{+}16, MD16, ST12, TS11, WCL^{+}18, XLX^{+}14, XWL^{+}16b, ZXX^{+}14]. Failure-to-Fault [HL10b]. Failures [FEP^{+}12, HK16, HWSN15]. Fair [GGSPM18, KH18, TSK16, VC10, FSPD17]. Fairness [FSPD17, LMC^{+}15, TSK16, WMW12]. Fairness-Based [LMC^{+}15]. Faithfully [DRC14]. False [YLY^{+}15a]. Family [ARH^{+}18, GPN11, SBM15]. Farewell [Zom15a]. Fast [ADJ12, AJH15, ASM^{+}16, AD16, BLKM^{+}18, BDE^{+}11, BCMJ10, CLS10, CLW^{+}15, Cil13, DJA14, GNSR14, HNV19, Ima18, IDG^{+}17, JDA15, JAD^{+}18, Kůr12, LNCX18, LL11, LCHX11, LW13, LPCW14, MNFA14, NL18, Pom12b, Red18, SYH17, SMG14, VAB14, WF17, WWT^{+}18, WZ14, YFCV14, YLL16, YUGD14, ZHW^{+}16, ZFJ^{+}17, ZWL^{+}19, GTRMG18]. Fast-Write-and-Rewrite [WZ14]. Faster [FHLOJRH18, KVV10, ZSP^{+}19]. Fat [GY15b, SJSLD11, WXW^{+}14]. Fat-Tree [GY15b]. Fat-Tree-Based [WXW^{+}14].

Fault [AE11, BMT14, BWCW15, BKP16, CKM15, CL12, DCK16, EYBK15, EGVFC^{+}12, GRM16, GV15, HL10b, JK15, JWH^{+}15, JKJ^{+}10, KCRG15, LYK19, LCC10, LH12a, LW17, LKT13, LCY^{+}13, LZS^{+}13, MLW12, MSS17, MG11a, MOMT12, MBSSA19, MKRM10, MKRM11, NI11, PNKI13, PPP13, PR10, Pom12a, Pom15b, QLR^{+}11, RVL^{+}14, RZZ^{+}15, RRS^{+}16, RKZ16, RZPX19, SBM15, SEY14, SDE^{+}17, SJSLD11, SRK^{+}17, SD13, SPH13, SBMP18, TLB^{+}17, TLL12, VCB^{+}13, WBZ^{+}15, WZL^{+}17, ZNL18, ZBK^{+}17, ZBW17, ZL15, ZJXL11, ZQQ11]. Fault-Aware [RKZ16]. Fault-Tolerance [BWCW15, JWH^{+}15, PPP13, RRS^{+}16, VCB^{+}13]. Fault-Tolerant [AE11, BMT14, JK15, LCC10, LW17, LZS^{+}13, QLR^{+}11, RVL^{+}14, SDE^{+}17, WBZ^{+}15, ZBW17, ZQQ11]. Faults [Ano13g, DKK16, EFGT18, FD16, HK15a, Li12a, MMC15, MCM16, MKT^{+}11, MD16, PB16, Pom12b, Pom16b, SBM15, ST16, ZHM^{+}19]. Faulty [AGFM11]. FD [OGH^{+}14]. FD-Buffer [OGH^{+}14]. Feasibility [ACM^{+}16, WHC^{+}15a, ZD13]. Feasible [YZGG16]. Feature [AHNT16, LJVJ18, WW14, ZYC16]. Features [OWP16, OKD^{+}16, PTD^{+}12, ZMB18]. Featuring [RRK11]. Feedback [CVH^{+}13, FD16, HZ11, MG16]. Feistel [BFMT16]. Femtocell [SPTC15]. FESTAL [WBZ^{+}15]. Few [CLZ19, KKH^{+}14, SBM15]. FFT [CYC^{+}16, CCR^{+}17, DCCK17, DCCK18, DALD18, SS12]. FFT-Based [CYC^{+}16, DCCK17, DCCK18]. Fi [HWK15]. Fidelity [SLL15]. Field [ABH^{+}13, BNP10, ERRM16, GKB^{+}10, HMNN12, HSA14, LCwW10, MLS19, MKRM11, NWA11, UHSA17, ZM10, ZAG19]. Fields [ARH14, HRM11, HN11, JDA15, NR15, ZL19]. FIFO [FJA^{+}17]. File [CS15, CCY^{+}16, CCC^{+}18, CLCS19, HWSX17, HZW^{+}12, HHW^{+}18, JZLD10, KL16, LYB15, LKBS16, LSW15, MWZ^{+}17, MLE14, PP11, RURM18, SCZ^{+}16, SL13, SLC15a, SYH17, VCSG^{+}19, WLK15, WLM15, XZL^{+}19, YY14, YSZ^{+}14, ZFHC19]. Files [RRK11]. Filter [AHNT16, EF12, HXVF12, LKYC12, LK15b, ML16, QZC15, SMRML17, ZL11]. Filter-Based [AHNT16]. Filtering [CWZC13, FEM^{+}18, RM15c, SL14a, TKT16, YLY^{+}15a]. Filters [ADC11, KBP13, LLLP14, NC11, PO13, VIDH19, YM11]. Financial [APP12]. Finding [Fen14, FK15, YUGD14]. Fine [Ged14, LSA^{+}17, LPD^{+}16, PSND16, SSJ^{+}18, SNM16, WZM^{+}16]. Fine-Grained [Ged14, LSA^{+}17, LPD^{+}16, PSND16, SNM16, WZM^{+}16]. Finely [GT19]. Finely-Pipelined [GT19]. FinFET [ACM^{+}16]. Finite [AWFV13, CWZ11, Hie11, Hie13, LLQ^{+}14, LCwW10, LW13, NWA11, SP10, ZL19, ZM10]. Finite-State [LLQ^{+}14]. Finite-Time [CWZ11]. FIPS [RBG^{+}19]. Firewall [YCZ10]. Firewalls [YCZ10]. First [CCE^{+}18, DSPB13, LPL12, PC16, RBG^{+}19]. First-Last [CCE^{+}18]. First-Order [LPL12]. FITS [CWZ11]. Fixed [BDB18, CK15, JCK15, Lee17, MBD^{+}17, NRG15, ZZ19]. Fixed-Point [CK15, JCK15]. Fixed-Priority [BDB18, Lee17, ZZ19]. Flash [AKJ^{+}13, Cha10b, CHK10, CK11, CHH^{+}13, CCK^{+}16a, CQW^{+}15, CWL^{+}17, CCL^{+}18, CLZ19, CC11, CYL^{+}14, DKH^{+}13, DSG^{+}19, FYSK14, FAK16, GKD^{+}17, GWM^{+}17, GCF^{+}16, IS11, JSH^{+}17, KLLK11, LSK13, LKLK13, LK14, LKLM15, LK16a, LRP16, LOX^{+}13, LSG^{+}18, LCY^{+}16, LSG^{+}15, LSGZ16, MLE14, NKEM11, OGH^{+}14, PDXZ13, PP11, PPKW12, PRM19, ROGHNB^{+}18, SNY^{+}10, SKM14, SYK14, TCYH15, UMN18, WLC^{+}15, WW16, WLY^{+}14, YCKH16, YCK16]. Flash-Aware [GCF^{+}16, IS11]. Flash-Based [CYL^{+}14, FAK16, JSH^{+}17, LSK13, LK16a, LRP16, LSG^{+}15, LSGZ16, MLE14, SKM14, WLC^{+}15, WW16]. Flash-Cache [SYK14]. Flash-Dissemination [DKH^{+}13]. Flash-Memory [CK11, CHH^{+}13]. FlexCL [LWZ18]. Flexibility [BCTV15, SLL15]. Flexible [CW10, CGL^{+}18, CLM^{+}19, DCL^{+}11, GT19, IB10, KLC^{+}16, NS13]. FlinkCL [CLOL18]. Flip [SMK^{+}16, YTND12]. Flip-Flops [YTND12]. Flit [MWLJ15]. Floating [AMG17, BLMM16, BGHL19, CHCK12, CI16, FKMK16, GH11, GNTS13, GABK11, HMS^{+}12, JPG10, JKMR11, JLMP11, JMMP16, KGD16, LP17, Lef17, LCW^{+}16, MKFM13, Nan19, PGvdG14, SS12, VVMAZ12, VMHGN18, WF17, ZMR^{+}13, ZCK19, dDLM11]. Floating-ECC [FKMK16]. Floating-Point [BLMM16, BGHL19, CHCK12, CI16, GH11, GNTS13, GABK11, HMS^{+}12, JPG10, JKMR11, JLMP11, JMMP16, KGD16, LP17, Lef17, LCW^{+}16, MKFM13, Nan19, PGvdG14, SS12, VVMAZ12, VMHGN18, ZCK19, dDLM11]. Flooding [GHG^{+}14]. Floor [BGRH15]. Floorplan [WXW^{+}14]. Flops [YTND12]. FLOTT [GNSR14]. Flow [AT16, ASS^{+}18, BCD^{+}16, CCV^{+}11, CBZ14, DDN14, DRS^{+}16, GAFN15, JWL^{+}16, JPLP13, KKY^{+}16, MWLJ15, MBD11, MD13, MBGS10, OPAGS14, PAP13, PTD^{+}12, QSYS16, STR15, SLS^{+}12, TSK16, VA11]. Flow-Based [PAP13]. Flow-Level [JPLP13]. Fluid [JWWZ16, KCE^{+}18, LRP^{+}18, MBM11, LRP^{+}18]. Fluid-Based [LRP^{+}18]. Flush [LKBS16]. Fly [DHM16, Pip11, VCG^{+}12, YLY15b]. FMA [BM11, HMS^{+}12]. Focus [YXWL16]. Fog [ZGG^{+}16]. Folded [CTH14, HTC13]. Fool [YGS15]. Forest [Cro14, LSX13]. ForestDB [ASM^{+}16]. Form [ACO12, LMZQ17, LPW10, PCHS14, VIDH19]. Formal [ABSK15, BIP^{+}17, HTA10, HSA14, LJ18, Rus13, SKEB16, UHSA17]. Formalizing [LY17]. Formation [LTP^{+}14]. Formats [AMG17, HV16]. Forming [JWWZ16]. Formulas [CNH13, Ose11]. Formulation [GGL^{+}14, ZD13]. Formulations [PSL17]. Forward [HLT^{+}15]. Forwarding [CSJ^{+}11, FS10, MS12, WGZ^{+}15]. Fourier [GTRMG18]. FP [FVV12]. FP-Arithmetic [FVV12]. FPGA [ALW11, AD13, BBA19, CJSM17, FML10, GP14, JC12, KAH18a, KN13, KL13, LW17, LMB17, LLD19, MLS19, PTD^{+}12, PBT13, PMH^{+}14, ROGHNB^{+}18, RJV^{+}18, SDP^{+}15, SME^{+}17, TB15, VMB19, YP12]. FPGA-Based [BBA19, CJSM17, GP14, JC12, KAH18a, LMB17, PMH^{+}14, ROGHNB^{+}18, RJV^{+}18, TB15]. FPGAs [ABB17, AKL14, EKA17, GT19, GEN^{+}17, GBA18a, GTRMG18, HVZ13, IBH^{+}13, KK18, LSC11, LP17, LWZ18, MB16, NZ15, NY15, NYHB16, RURM18, RTRM19, SBMP18, WLV^{+}14, WMG18, WDSP12, BMS11]. FPTAS [NRG15]. FPU [MKFM13]. Fragile [BIP^{+}17]. Fragmentation [KWC^{+}16]. Frame [DA12]. Frame-Based [DA12]. Frames [GM15]. Framework [ARGT14, BGPV10, CXLL16, DHC^{+}16, Fin10, GVPS19, GZG^{+}16, HCSW15, HDYS16, KPBC17, KCW^{+}17, LKLK13, LR16, LMJ14, LHL15b, LJY^{+}15, LQD^{+}16, LZZ16, LJL13, MYHL16, PBL16, PNKI13, RVC^{+}15, SBP16, SVD18, SCK10, SCZ^{+}16, SMN^{+}17, SPTC15, WZZ10, WLYY16, WLZ10, YTD^{+}18, YZGG16, YAGB17, ZJL^{+}16]. Free [BZ14, CT13, CJG16, CSJ^{+}11, DSPB13, FG10, GDJZ18, HTA17, LKLT12, PCHS17, RRS^{+}16, ST11b, WP16, WMG18]. FREM [LL11]. Frequency [DA12, FCB^{+}19, KkC15a, LHTG15, NYHB16, NY19, WBGI19]. Freshness [LHC^{+}14, LcZLG19, RHC^{+}14]. Friendliness [KIJ14]. Friendliness-Aware [KIJ14]. Friendly [FVV12, SZDL14, ZM17].

Friends [JWWZ16]. FSA [LZS^{+}13]. FTCAM [FAK16]. FTH [BMT14]. FTH-B&B [BMT14]. FTL [HLY14, YLH13]. Full [WWT^{+}18]. Full-Chip [WWT^{+}18]. Fully [CMO^{+}16, DOS15, HFG^{+}17, LXK12, WHC^{+}15a]. Fully-Pipelined [HFG^{+}17]. Function [Bai17, BLB^{+}19, CA12a, CJSM17, DKLB15, HWCH17, RJV^{+}18, Tho15, WEX14, dDLM11]. Functional [BCSR14, BC16, GPRS17, GAFN15, JLMH10, JRJ^{+}18, LQD^{+}16, Pom12a, Pom12b, Pom13a, Pom13b, Pom14, Pom15a, Pom15c, Pom15b, Pom16a, YTND12, ZRL15]. Functionalities [SBP^{+}14]. Functions [AFC10, AR17, BHR17, JLMP11, JRP^{+}14, KM11, KFB^{+}15, LJ13, dLSGDR17, SKM^{+}13, SDP11, TSA^{+}19, ZYHZ16, vdBGLGL^{+}16]. Fused [SS12, WF17, ZYW^{+}16, ZCK19]. Fuzzy [EFPC16, GSH^{+}14, LZ15, PdG13, USH19, XJWW13]. Fuzzy-Controlled [PdG13]. FV [MRL^{+}18, RJV^{+}18].

Gains [CA12a]. Galois [HSA14, UHSA17]. Game [CB15, EFPC16, SCK10]. Games [BSM^{+}14, BKV12, CFW14]. Gaming [DGC^{+}15]. Gappa [dDLM11]. Garbage [CW10, DSW^{+}14, LSK13, LTW^{+}12, UMN18]. Gate [GSK12]. Gates [Ibr16]. Gateway [SME^{+}17]. Gathering [WLYY16, ZMY11, ZY12]. Gating [LW17, RRK11]. Gauges [ROGHNB^{+}18]. Gauss [CCR^{+}17]. Gaussian [AMR18, AB16, ARM13, ERRM16, FB13, GPR^{+}19, HFG^{+}17, HKR^{+}18, KRR^{+}18, RMERM19, WJL^{+}12, WZCG16, ZGY13, ZGY14, ZR15b]. GCM [JL11, MKRM12]. GCMA [PKY19]. GDP [WTZ^{+}19]. Gene [WGW^{+}15]. Gene/Q [WGW^{+}15]. General [GZG^{+}16, HTA17, LKLK13, LL11, LJL13, SHGW15, SJVR19, WLZ10, ZBW17]. General-Purpose [SJVR19]. Generalization [CSCW13, JDA15]. Generalized [BFMT16, GL19, HBAD14, JWL^{+}16, PAP13]. Generalizing [LP13b]. Generate [BGM^{+}13]. Generated [CW15, YLY15b]. Generating [AFH^{+}10, HT16, LB15a, SN16]. Generation [AK16, CM11, CYA13, CCD12, FM19, FD16, GT19, GJ14, GSK12, NZ14, NM10, Pom14, Pom16a, TXL11, USP^{+}13, VK15, ZLY15]. Generator [CLC^{+}16, Jes15]. Generators [MG16, YMAG17]. Generic [WCL^{+}18]. Genetic [CJSM17, LJVJ18, QML^{+}15]. Genomic [KPBC17]. Genomics [GKC19]. Genuine [WJY^{+}17]. Geo [BBPQ15, CPL17, GZB^{+}15, GZG^{+}16]. Geo-Distributed [BBPQ15, CPL17, GZB^{+}15, GZG^{+}16]. Geographic [CSJ^{+}11, CBVL16]. Geometric [EG11, FGS^{+}15, SZS14, WF14]. Geometry [EG11]. Georouting [RS13]. Getting [Jun16]. GF [GT19]. GliFreD [WMG18]. Glitch [FNS16, WMG18]. Glitch-Free [WMG18]. Global [GYC^{+}16, GHK15, MTGM12, YPB^{+}16, ZL15]. Globally [GGA^{+}17]. GNB [PCHS18, WF12]. Goldschmidt [KS10b, KKS14, PB11]. GPGPU [ADP^{+}15, CVPS19, LLC^{+}16, MWZ^{+}17, YEG^{+}15]. GPGPU-MiniBench [YEG^{+}15]. GPGPUs [JJZ^{+}16, LKJ15]. GPS [LKLT12]. GPS-Free [LKLT12]. GPSR [LYCT10]. GPSR-Like [LYCT10]. GPU [CLOL18, DALD18, GD17, GBF^{+}19, KLC18, LR16, LSA18, LMT13, MTFK19, PTD^{+}12, SCSL12, SKYK16, VCSG^{+}19, VAN^{+}18, WZK^{+}19, XHZC16, ZS13, ZYW^{+}16, ZRL15]. GPU-Accelerated [MTFK19, SCSL12]. GPU-Aware [KLC18]. GPU-to-GPU [ZS13]. GPUs [LWKA15, LKK^{+}17, LLCC13, MB16, OYP^{+}18, OKY^{+}19, XLL^{+}18, YLML15, ZMW^{+}19, vdBGLGL^{+}16]. GPUvm [SKYK16]. Gradients [Cro14]. Grain [SBM15]. Grained [CCY^{+}16, Ged14, JLKR19, LSA^{+}17, LPD^{+}16, PSND16, SNM16, WZM^{+}16]. Granular [KKT15, LFH^{+}16]. Granularity [LHH17, QZC15, SSJ^{+}18]. Graph [DHW^{+}19, DLL^{+}12, GZC^{+}17, GWZ^{+}10, PPB^{+}14, SX12, SD13, WZW^{+}19, XYHD17, ZLXW15]. Graph-Based [SX12]. Graphics [CCLH10, GBA^{+}18b, HTA17, LR10, dOPSR16]. Graphs [CH11, CL12, CH13, CCH15b, HFZ13, HLWV17, HNB^{+}12, LMB13, MBB^{+}17, MY10, SBI12, UMN18, ZLJ^{+}17, dRV12]. Gray [ABA07, BTBB14, BBH12, Jha13, HBAD14]. Gray-Box [BTBB14]. Greedy [EG11, GLXY13, WTZ^{+}19]. Green [QML^{+}15, LWKA15]. Greening [LSHC15]. Grid [AD14, WZY16, ZV14]. Grids [BR13, HV13, SH12]. Group [CLW16b, HFZ13, HL10a, Har13, Hia16, JCM16, LHYZ13, LCCJ13, SJC^{+}17a, TKL^{+}14, ZYHZ16]. Grouping [GWZ^{+}10, LBWH11, SDZ15]. Grow [DYJ19]. Grow-and-Prune [DYJ19]. Growing [YCKH16]. GSV [MTGM12]. Guarantee [AD14, LZ15, LH11, WZLX12, ZWC^{+}18]. Guaranteed [CWTT13, GY13, KS10a, MLOL15, PKY19, RRS^{+}16, RS13]. Guaranteeing [NLP^{+}14, ZRS^{+}16]. Guarantees [FS10, HC13b, ZWLS15]. Guardband [KAH19]. Guardbanding [RBG14]. Guessing [XJWW13]. Guest [BKP16, CHTD19, DPO17, WHBR16, XL16, AHI12, AISA16, Avr13, BKPMC13, BMM11, BS10, BCS11, BdD19, EM12, GC14, GM11, LLK18, MG11a, MOS14, NST14, ST11a, VP14, ZMS13]. Guided [LSG^{+}18, LH11, NC11, SRK^{+}17, TKT16].

H [JAS^{+}15]. H-SVM [JAS^{+}15]. H.264 [SRK^{+}17]. HALLS [KA19]. Hamiltonian [AB16, HBAD14]. Handauth [HBCC13]. Handling [CBB19]. Handover [HBCC13, JLKR19]. Hard [AE11, CW10, CYCC11, EKA17, HV12, HK15a, KMC17, KK10, LP17, MBSSA19, RCN11, WLY^{+}14, WA10, XWL^{+}16b]. Hardened [CPRH16]. Hardening [KwPK^{+}15, MTGM12]. Hardness [JWH^{+}15, XTF^{+}12]. Hardware [ADJ12, AVG^{+}15, AS10, ADC11, BQP^{+}16, BEHL^{+}19, BKPMC13, BCSR14, BDMLN16, BCMJ10, CMLRHS13, CVH^{+}13, CCAM14, DJJ^{+}08, DOS15, DW10, DKK16, ERRMG15, FVV12, GKB^{+}10, GBD^{+}15, GLP15, GCS^{+}13, HFG^{+}17, HSA14, HWG^{+}14, HCG^{+}16, HC17, HGEG11, HNB^{+}12, JSC10, JAS^{+}15, KAK18, KT12, LK10, Lee12, LH16, LKH16, LCL17, LLW^{+}11, LTLC12, LGH^{+}17, LLKA19, MGdC^{+}18, MSPK12, MLCH10, MGW14, MW10, MRL^{+}18, MKRM12, MF14, NDC^{+}13, OWP16, OKD^{+}16, PC16, ROH17, RWC18, RCK^{+}16, RTL^{+}18, RM15c, SOM^{+}13, SBP^{+}14, SKPK10, SDP11, TW10, TB15, TCK^{+}18, TGNSC11, THGT13, TS11, USP^{+}13, UdDG^{+}17, VAN^{+}18, VIDH19, WGW^{+}15, WHYS16, XMH13, YCK16, ZCZ^{+}19, ZZ17, ZYW^{+}16, ZL19, ZM17, ZAG19, Zot10]. Hardware-Assisted [ADC11, JAS^{+}15, LLKA19, RTL^{+}18]. Hardware-Based [MGdC^{+}18, OKD^{+}16, RM15c, USP^{+}13]. Hardware-Efficient [XMH13]. Hardware-Friendly [ZM17]. Hardware/Software [HWG^{+}14, JSC10, LK10, LH16, MRL^{+}18, WHYS16]. HARL [HWSX17]. Harmonised [HT16]. Harmony [LYL^{+}19]. Harvesting [AS16, CQ14, MMH14]. Hash [AR17, AS16, HC13b, LRY^{+}15, vdBGLGL^{+}16]. Hash-Based [AS16]. Hashing [ASBdS16, BKL^{+}13, LLL11, PRM16]. HDD [HWS^{+}17]. HDD/SSD [HWS^{+}17]. Healing [CWL^{+}17]. Health [BWCW15, GLTC16, XWL^{+}16b, ZL18]. Healthy [YSLL16]. Heating [CWL^{+}17]. Hellman [FHLOJRH18, LNL^{+}19]. HEPCloud [RJV^{+}18]. Heterogeneity [CCC^{+}17, CNJ14, CCK^{+}16b, CLMM11, HWS^{+}17, HWSX17, LK14, LPD^{+}16, XLJ16]. Heterogeneity-Aware [CNJ14, HWS^{+}17, HWSX17]. Heterogeneous [AQPMS15, AT16, Ano11f, BMP^{+}10, BPT10, CFR^{+}14, CLS14, CLOL18, CRJZ16, CYC11, CDK^{+}18, DSB13, DFP^{+}13, FMTK19, FAA10, GY14, GW16, HLA^{+}17, JKD^{+}19, KPS^{+}17, KFB^{+}15, KKC17, LRC10, LR18, LXDV17, LTVL15, LZZV16, MYHL16, NEE18, PKC^{+}17, PBL16, QWB^{+}13, TLZV11, TCHL18, ZLBB19, ZJL^{+}16, ZQQ11, ZLLX15, ZMS13]. Heterogeneous-ISA [TCHL18]. Heuristic [JWH^{+}15, KCS^{+}13, KL13]. Heuristics [IB10]. Hierarchical [BMT14, BBB16, HWCH17, KKY^{+}16, NH10, PZZQ19, TLH^{+}16, TLB^{+}17, ZHW^{+}16]. Hierarchy [DHW^{+}19, GO10, LKJ15, OOD^{+}17]. Hierarchy-Aware [OOD^{+}17]. High [ARM16, AFC10, AS10, ASBdS16, Ano13f, AS12, BGHL19, CCH^{+}15a, CLM^{+}19, CWY13, CDL^{+}17, FFISC13, FEM^{+}18, FG10, GZC^{+}17, GAFN15, GJ15, GY13, GCS^{+}13, HK13a, HZ11, HLY14, IS11, JDA^{+}16, JC12, JHQL16, KAH18a, KIJ14, KBH^{+}10, KLC18, Kor15, KAK18, KHZ17, LLC^{+}16, LS10a, LMJ14, LCL17, LZZ^{+}17b, LSG^{+}15, MEBS17, MM17, MIS^{+}14, MKRM11, MKRM12, NWA12, ORM10, OPV^{+}17, OMFH14, PLM16, PP11, PvdGG12, PBT13, QWB^{+}13, QLH^{+}16, RVL^{+}14, RSG^{+}19, SMTK12, SCSL12, SME^{+}17, Tho12, TCK^{+}18, VC10, VAM10, WWY^{+}18, WDSP12, XJFH15, YP12, YAG19, ZZ17, ZL16, ZLSI17, ZZJ^{+}19, ZMW^{+}19, ZCC^{+}14]. High-Dimensional [KLC18, MEBS17]. High-Level [GAFN15, MIS^{+}14]. High-Performance [AS12, FG10, GCS^{+}13, HLY14, IS11, JC12, JHQL16, KBH^{+}10, KAK18, LS10a, LZZ^{+}17b, LSG^{+}15, MKRM11, MKRM12, ORM10, PP11, PvdGG12, QWB^{+}13, RVL^{+}14, SCSL12, VAM10, WWY^{+}18, YP12, YAG19, ZZJ^{+}19, ZMW^{+}19]. High-Precision [BGHL19]. High-Radix [AS10, Kor15]. High-Speed [ARM16, Ano13f, CCH^{+}15a, GJ15, GY13, HK13a, HZ11, NWA12, PBT13]. High-Throughput [AFC10, FFISC13, KAH18a, LCL17, OMFH14]. Higher [BMZ17, UHSA17, ZMB18]. Higher-Order [BMZ17, ZMB18]. Highly [CCP^{+}13, GEN^{+}17, IBH^{+}13, KA19, SZDL14]. History [LBN14, Liu11]. Hoc [CWZ11, CS15, CWTT13, CWY13, DLL^{+}12, FS10, GDY15, LYCT10, RDEN10, TH11, XWY10]. Hole [Amm14, PC16]. Holes [MMB14, WS15]. Holistic [MJW^{+}14, STK16]. Homogeneous [ML18, TFCY16, ZMRQ11]. Homomorphic [CMO^{+}16, CJ13, DOS15, KGV16, Kim15, LCLL15, MLW12, MRL^{+}18, MBF18, ÖDSS17, RJV^{+}18, WHC^{+}15a, YCK10]. Hook [MAG^{+}17]. Hop [LWY15, MWY^{+}16, ZY12]. Hope [SD18]. Hopping [LWY15]. Horizontal [LLD^{+}16]. Host [CH14, SRCbL^{+}15, WFY^{+}17, YCW^{+}19, ZS13]. Host-Aware [YCW^{+}19]. Host-Based [CH14]. Host-to-Host [ZS13]. Hosting [PAP13]. HotGraph [ZLJ^{+}17]. Hotspot [STK16]. Hotspots [GBF^{+}19]. HOWP [LHYZ13]. HOWP-Based [LHYZ13]. HPC [HLC^{+}19, SC18]. HRC [WJY^{+}17]. HRT [WLY^{+}14]. HRT-PLRU [WLY^{+}14]. HUB [VMHGN18]. Human [OPZ15, SLC^{+}15b, WW14]. Human-Centric [SLC^{+}15b]. Hundred [DJN17]. HW [DMK^{+}15, MH19]. HW/SW [MH19]. Hybrid [ARM16, ADI11, AYC16, AFH^{+}10, Ano13g, ARM13, BBB^{+}17, Cha10b, CC11, CDL^{+}17, DHW^{+}19, DJA14, ERRM16, FRB^{+}18, GD17, GCD^{+}11, GY13, HNV19, HWS^{+}17, HCL15, HK15b, HV14b, JYL^{+}17, KKY^{+}16, LBN14, LCL15, LLL11, LLW^{+}11, LFH^{+}16, LLD19, LKMSA16, LCLC19, MRW^{+}15, RVL^{+}14, RSNK17, RBG^{+}19, SMA19, SPTC15, SME^{+}17, SR14, TAH^{+}16, VSLD15, WWY^{+}16, WGLL13, WS15, WLS18, XXBL17, YY10]. Hybrid-Double [ARM16, ARM13, DJA14, ERRM16]. Hybrid-Memory [BBB^{+}17]. Hybrid-Switched [LKMSA16]. Hybrid-Systems [AFH^{+}10]. Hybrid-Triple [ERRM16]. Hydra [SNY^{+}10]. Hyper [ADP^{+}15]. Hyper-Real-Time [ADP^{+}15]. Hyperbolic [EG11, dLSGDR17]. Hypercube [SKA10, Tsa13, WW14, YLL16, ZWYY15]. Hypercube-Based [WW14]. Hypercube-Like [Tsa13, YLL16]. Hypercubes [CTH14, HTC13, HBAD14]. Hypermesh [HWL^{+}14]. Hyperperiod [RBR13]. Hypervisor [JAS^{+}15, SKYK16]. HyVE [DHW^{+}19]. HyWin [GD17].

I/O

[BBP^{+}13, DYHX16, GKD^{+}17, HWS^{+}17, HS18, HQLX15, HLC^{+}19, KSJ^{+}12, KRP18, LKBS16, SYH17, SHH^{+}16, TAH^{+}16, ZL16]. I/O-Redirection-Based [HQLX15]. I/Os [WZW^{+}19, ZLWZ15]. IaaS [FLL14, GHL17]. iAware [XLL^{+}14]. IBM [MBC^{+}13]. iBuddy [PCLN15]. ICCI [GGFPG15]. Ice [ADP^{+}15, SVAB14]. ICs [SVAB14, XCF16, YEY^{+}16]. ID2S [YRT^{+}16]. Ideal [RM15a, SMB^{+}15, SWZG11, WCL^{+}18]. Identical [OCK17]. Identifiability [ZC13]. Identification [BTBB14, LWL^{+}16, Pom12b, YXWL16, ZCC^{+}14]. Identifiers [HT16]. Identifying [Ibr16]. Identity [FHH10, HSM14, LLC^{+}15, XJW^{+}16]. Identity-Based [FHH10, HSM14, LLC^{+}15, XJW^{+}16]. Idle [CYL^{+}14]. Idle-Time [CYL^{+}14]. IEEE [Ano10d, Ano10e, Ano10b, Ano13e, Ano15a, Ano16a, Ano17a, BKP16, CTS13, HXVQ15, Liu11, NBZP17, SMB^{+}15, ZICL12, ZNL18, Ano13c, Ano18a, Ano19a, CHTD19, DPO17]. If [YGS15]. IIPS [WZBB15]. IIR [VIDH19]. IMA [NLP^{+}14]. Image [AKL18, FGS^{+}15, PTD^{+}12, SOM^{+}13, TKT16, ZGG^{+}16]. Images [LHCL13]. Imaging [KN13]. Impact [CBTU14, GBGI18, KBP13, MKT^{+}11, PC10]. Implement [CXLL16]. Implementation [AFC10, Bai17, BM13a, BDMLN16, CFR^{+}14, CS11b, DJJ^{+}08, FVV12, FHLOJRH18, FGS^{+}13, GKB^{+}10, GLP15, GCS^{+}13, IDG^{+}17, JRC14, KSS12, KMLH11, KGV16, Lee12, LYB15, LCH13, MLE14, MNK11, PRM16, QWB^{+}13, RM15a, SDP^{+}15, SJS^{+}14, SDP11, SS12, YSZ^{+}14, Yun12, ZPM^{+}15, ZL18, dDLM11]. Implementations [BJ10, CMLRHS13, CCRL19, ERRMG15, LGH^{+}17, MLCH10, MG16, SMRML17, STE17]. Implemented [GTRMG18, RURM18]. Implementing [BMS11, CCV^{+}11]. Implication [Tho15]. Implications [SLLG15, WCM^{+}16, ZWD^{+}16]. Importance [YRG13]. Imprecise [LGS^{+}18]. Improve [CJA^{+}16, LYCT10, LHH14b, MJWT16, YCK16]. Improved [ABH^{+}13, BLB^{+}19, CNH13, CHCK12, DS14, Fuj11, GPR^{+}19, HIJ^{+}19, HJF^{+}13, LT14, LHPH15, Lee17, LCH^{+}15, LCCJ13, MWW14, MLS19, MM17, MG11b, Ose11, Pom12a, RCN11, SJS10, VAM10, WJM15, ZGB^{+}15]. Improvement [CK11, CZ16, MEBS17, Pom15b, WCM^{+}16, ZWD^{+}16]. Improving [AK19, CDQB15, CHK10, Fen14, HGCT13, JZLD10, KLK18, LK14, LKK^{+}17, LR18, LLW^{+}11, LCY^{+}16, Pom16b, RKR15, SMK^{+}16, WZL^{+}17, WJF^{+}11, XLC14, YyHL11, YCW^{+}19, ZHM^{+}19]. Impulse [LHCL13]. In-Cache [GGFPG15]. In-Line [ROGHNB^{+}18]. In-Memory [CLOL18, GIKR19, SSGB19, SCZ^{+}16, ZCZ^{+}19, ZHQX19]. In-Memory-Computation [AAR19]. In-Network [VBR^{+}13]. In-Order [RRK11]. In-Place [LNCX18]. In-Situ [NY15].

In-System [SN16]. In-the-Field [MLS19]. Incentive [FLJ14]. Incoming [Lou19]. Incomplete [NS13]. Incorporating [SRCK10]. Increased [PRM16]. Increasing [CRG^{+}13, DY14, MMAC19, NLP^{+}14]. Incremental [BC16, CLW^{+}16a, DLC^{+}13, TC14]. Independent [DEE17, MPZ15, MKRM10, TYWC10, Tse12, USP^{+}13, VED^{+}16, YCKH16]. Index [AKJ^{+}13, Ano11a, Ano12a, Ano13a, Ano14a, Ano15a, Ano16a, Ano17a, Ano18a, Ano19a, DALD18, KLK17, KLC18, SJD^{+}18]. Index-Based [KLK17]. Index-Digit [DALD18]. Indexing [GYC^{+}16, RXC^{+}15, WYL^{+}15, XJFH15]. Individual [dRV12]. Induced [DSG^{+}19, GBA18a, HK16, KWC^{+}16, ST16, dOPSR16]. Inductions [LDP10]. Inductive [TMS^{+}14, YEY^{+}16]. Inductive-Coupling [TMS^{+}14]. Industrial [SXLC15]. Inefficiency [LZW^{+}15]. Inexact [AKL18, LCW^{+}16]. Infected [YKK^{+}15]. Inference [LJVJ18, NY19, XP10]. Information [CWTT13, GGFPG15, HCSW15, LOH17, LKLT12, OPAGS14, RKN^{+}18, SLL15, WHC^{+}15b]. Information-Centric [LOH17]. Infrastructure [HLF14, MCC12, WZBB15, XH16]. Inherently [SKA10]. Inheritance [BGRH15]. Inherited [HH17]. Injection [EGVFC^{+}12, PNKI13, YLY^{+}15a]. Injections [SBMP18]. Inline [WWY^{+}18]. Innovation [DPO17]. Input [ACGP13, BGMR13, Ibr16, NCD^{+}17, SJS10, ZWLS15]. Input-Queued [ACGP13, BGMR13]. Input/Multi [TWTT11]. Inputs [BCK^{+}16]. Insensitive [OMFH14, XSR15]. Insertion [JSC^{+}19, YTND12]. Inspection [LW13, LPCW14, ROGHNB^{+}18]. Inspired [LSZ^{+}15, PBL16, SCJ^{+}16b]. Instance [JT15]. Instant [YXZZ14]. Instantiating [CMRH17]. Instruction [DZ10, GBF^{+}19, JLKR19, LYH11, LSA^{+}17, MKT^{+}11, MIS^{+}14, RS17, SMAR^{+}19, TCK^{+}18, ZYY18]. Instruction-Based [ZYY18]. Instruction-Level [MKT^{+}11]. Instruction-Set [LSA^{+}17]. Instructions [IS14, JL11, LSC11, USP^{+}13]. Instrumentation [GBF^{+}19, GDJZ18]. Instrumentation-Free [GDJZ18]. Integer [ADJ12, CL10, CZS^{+}19, CND^{+}18, GNTS13, MCT19, RV13, ROH17, TGNSC11, UdDG^{+}17, WHL^{+}12]. Integers [MG11b]. Integral [ZOK^{+}19]. Integrated [ASS^{+}18, CWZC13, CKN14, DYHX16, DAPS14, GWMB13, LSW15, TS11, ZLYS15]. Integrating [HSH^{+}10, WZZ10]. Integration [ALW11, DFP^{+}13, LKK^{+}19, LVF19, VGF16]. Integrity [JCM16]. Intelligence [JRP^{+}14, SLC15a]. Intelligent [FMTK19, MFT^{+}17, STK16]. Intensive [RLSK18, WSZ^{+}16, WGR^{+}14, YZHX12]. Inter [cCWS14, SMN^{+}17]. Inter-Application [cCWS14]. Interacting [YMT13]. Interaction [TZL^{+}14]. Interactions [cCWS14]. Interactive [ZT15]. Interconnect [KL13, ZGY14]. Interconnected [LKT13]. Interconnecting [LW15]. Interconnection [CMB13, CTD^{+}16, FB13, SRR^{+}16, SMN^{+}17, Ste14]. Interconnects [AKL14, DCY^{+}13, FAA10, HJBM14, PVKA14, SC18]. Intercore [WLQS13]. Interdependent [HWSN15]. Interface [DDN14, DRS^{+}16, SBP^{+}14]. Interface-Based [DRS^{+}16]. Interfaces [Hie13]. Interference [HWK15, LGF^{+}15, PC10, XWL^{+}16a, XLL^{+}14, XLJ16]. Interference-Aware [XWL^{+}16a, XLL^{+}14, XLJ16]. Interlaced [FF16]. Interleaved [KVV10]. Interleaving [CVGZ15, KS10a]. Internal [GJ15, HK13a, KWC^{+}16]. Internet [CLX14, CAGM14, LHH14b, LGH^{+}17, PPB^{+}14, XLF15, YZF^{+}10].

Internet-Based [CAGM14]. Internode [YLA10]. Interplay [DA12, GHK15, HXL11, ZZX^{+}15]. Interpolation [WEX14]. Interrupt [LLW^{+}11]. Interrupts [LLW^{+}11]. Interval [Joh17, RT14]. Intra [SRR^{+}16, SMN^{+}17, WZK^{+}19]. Intra-Chip [SRR^{+}16, SMN^{+}17]. Intra-Cluster [WZK^{+}19]. Introducing [SAR^{+}11]. Introduction [AHI12, AISA16, BKPMC13, BMM11, BS10, BCS11, BdD19, EM12, GC14, GM11, HMO^{+}17, LLK18, MG11a, MOS14, NST14, ST11a, VP14, ZMS13, Avr13]. Intrusion [AHNT16, CH14, PBT13, VSF^{+}17]. Intrusive [TJX^{+}17]. Invalidations [ADC11]. Invariant [MG11b]. Invasive [ADJG19]. Inventory [CKN14]. Inverse [MKRM11, PCLN15, RMERM19]. Inverses [Dum14, Wal19]. Inversion [BT16, DJA14, LLHC15]. Inversions [JDA15]. Investigating [Amm14]. Investigation [CJ12]. Invocation [RQ14]. IO [SKC^{+}14]. IoD [DHC^{+}16]. IOMMU [VMB19]. IP [CLS10, CKKS14, HY12, HH17, JP13, LYS10, LS10c, ML16, WZBB15, YZGG16]. IPs [NDG^{+}17, BFP11]. IPv4 [KCS14, LP12]. IPv4/IPv6 [KCS14]. IPv4/v6 [LP12]. IPv6 [KCS14]. Iran [JGG^{+}14]. Irreducible [Fan16, HF15]. ISA [GGSPM18, MIS^{+}14, TCHL18]. Ising [ZCW18]. Islands [OKC13]. Iso [SRHC12]. Iso-Power [SRHC12]. Isogeny [BF19, FHLOJRH18, KAK18, LNL^{+}19, ZSP^{+}19]. Isogeny-Based [BF19, KAK18, ZSP^{+}19]. Isolation [HGCT13, MGdC^{+}18, WZL15, YYP^{+}16]. Isomorphism [ZYHZ16]. ISRA [LBWH11]. ISRA-Based [LBWH11]. Issue [GC14, HMO^{+}17, LLK18, VP14]. Issues [RT14, TKL^{+}14]. Iteration [Dum14, FMTK19, PP16, Wal19]. Iterations [BBK10]. Iterative [CG18, FTP13, TC16, VB13]. Itinerary [LLCH13]. Itinerary-Based [LLCH13]. Ivan [Zom15b].

Java [CCH11]. JEM [JAJK15]. Jitter [MAHD18]. Job [SL14b, YLH10]. Jobs [BBD^{+}12, NL16c]. Join [LOX^{+}13]. Joining [Ano10b]. Joint [BKP16, DPO17, GDY15, HKWC14, HGL^{+}15, ZGG^{+}16]. Journal [Lom10, Mon15b, Mon16, MMC^{+}16, Mon17, Mon18, Mon19, Zom11b, Zom11a, Zom12b, Zom12c, Zom13, Zom15a]. Journaling [CCC^{+}18, LYB15, ZFHC19]. JPEG [HHCH11]. Just [HZW^{+}12, JAJK15, VIDH19]. Just-in-Time [HZW^{+}12].

Kalman [Red14]. Kalray [IDG^{+}17]. Karatsuba [LMZQ17, LPW10, MRL^{+}18, Ose11]. Karatsuba-Based [LPW10]. Karatsuba-Like [Ose11]. KASE [CLW16b]. Keccak [RS17]. Kernel [BDM^{+}19, JCK15, SWWC11]. Kernels [XP10]. Key [ASM^{+}16, AFC10, CJ13, EFGT18, FHLOJRH18, GDLL18, HL10a, HWZ^{+}17, Kim15, LCLL15, LCwW10, LHYZ13, LCCJ13, PSM17, RVH^{+}16, RNS13, SMRM17, SWM^{+}10, SD18, WCL^{+}18, XJWW13, YRT^{+}16, ZSP^{+}19, CLW16b]. Key-Aggregate [PSM17, CLW16b]. Key-Policy [RVH^{+}16]. Key-Value [ASM^{+}16]. Keyed [KE19]. Keys [ASM^{+}16, PSM17]. Keyword [XJWW13, ZLX^{+}16]. Kiel [LvH12]. Kleene [DSB13]. Knowledge [SLLG15, SDZ15]. Koblitz [Lee12, ADJ12, BJ10, DJJ^{+}08, TX16]. Krivine [MRD19].

L0 [LK15b]. L1 [EF12, HK16, SV18, VPS^{+}12]. L2 [SV18]. Label [LCL15]. LACS [KS14]. Lags [CFMS14]. Lanczos [JCK15]. Land [DAPS14, LMB17]. Land-Use [DAPS14]. Lane [HHC^{+}18]. Language [ASE17]. LANs [GY16, XHZ14]. Large [AISA16, BMT14, CYJ^{+}10, CL12, CQW^{+}15, CXLL16, CLW^{+}16a, DALD18, FFCB14, Fin10, GDC^{+}16, GV15, GY16, HZW^{+}12, JKY10, LBSK17, LP13a, LS10a, LDB^{+}17, LXK12, LXZ^{+}15, LQD^{+}16, MCXZ18, MC11, NM10, PDXZ13, ROH17, SSGB19, WS15, WJM15, XHLW19, ZCZL16, ZWC^{+}18, ZYY10]. Large-Capacity [PDXZ13, XHLW19]. Large-Scale [CQW^{+}15, FFCB14, GY16, JKY10, LP13a, LS10a, LXK12, LQD^{+}16, MCXZ18, WJM15, ZCZL16, ZWC^{+}18]. Last [CCE^{+}18, KIJ14, KKC15b, KA19, LCLC19, YMG15, ZJS14]. Last-Level [KIJ14, KKC15b, LCLC19, YMG15]. Latch [SB16, ZZ17]. Latch-Based [ZZ17]. Latches [ORM10]. Late [KKH^{+}14]. Latencies [LLW^{+}11]. Latency [ADOKM10, CLL^{+}14, CYL^{+}14, FFISC13, GY13, KGP15, LR13, LYS14, MKAY11, NLP^{+}14, NL19, PLM16, PB11, QSYS16, RZPX19, RM15b, SL10, SCJ^{+}16b, SR14, WMW12, XSR15, YCCJ15, ZLW^{+}17]. Latent [CJ12]. Latin [NL15a, DRM16]. Lattice [EFGT18, GLP15, HKR^{+}18]. Lattice-Based [EFGT18, GLP15, HKR^{+}18]. Lattices [AR12, GPR^{+}19, MEBS17]. Law [CA12a, YMG16]. LAWC [GKD^{+}17]. Laws [WJL^{+}12]. Layer [CC11, GY16, KCW^{+}17, LLW^{+}18, PRGBSAC19, RCK^{+}16, SVD18, SDE^{+}17, VSC^{+}19, YCK16]. Layered [BS14]. LayeredTrees [CKKS14]. Layers [RWZZ14]. Layout [AKJ^{+}13, GKD^{+}17, HWSX17, HT12, LZZ17a, PVKA14, SWZG11, ZZL14, ZWW19]. Layout-Aware [GKD^{+}17, PVKA14]. LCP [WNCH17]. LDet [CCL^{+}13]. LDPC [CMM15, LLC^{+}16]. LEAD [SKEB18]. Leakage [Bar16, CYCC11, GDLL18, LVMS18, LGMP10, MKM14, SRCK10, ST16, WWY^{+}16, WWT^{+}18]. Leakage-Aware [MKM14, WWT^{+}18]. Leaks [CATB19].

Learning [Bar16, CM11, DYHX16, FMTK19, FCB^{+}19, GBA^{+}18b, JKD^{+}19, LPL^{+}13, LK18, LH11, MFT^{+}17, WP16, ZYC16, ZM17]. Learning-Based [JKD^{+}19]. Learning-to-Rank [LPL^{+}13]. Least [LR13]. Least-Latency [LR13]. Leaving [MWLJ15]. Lee [Jha13, ABA07]. Legacy [TJX^{+}17]. Length [ASM^{+}16, Fen14, Pom12a, RS13, SRK^{+}17, WTY^{+}14, XMH13, YLP15]. LEO [LZS^{+}13]. Less [HLC^{+}19, WBGI19]. Level [ARM16, AJH15, AE11, Ano11f, BCL^{+}17, BPT10, BS10, BM13b, CCC^{+}17, CCW^{+}10, CQW^{+}15, cCWS14, ERRM16, GAFN15, GWM^{+}17, HWSX17, HGW^{+}17, JRJ^{+}18, JPLP13, KIJ14, KGP15, KO14, KLC^{+}16, KKC15b, KA19, LK10, LR16, LLW^{+}11, LCLC19, MKT^{+}11, MIS^{+}14, NL15b, NL16a, NWA11, NL16c, OKD^{+}16, PNKI13, SJC^{+}17a, SJC^{+}17b, VSLD15, WP16, XP10, XZL^{+}19, XLL^{+}18, YSLL16, YMG15, ZGR13, ZJS14, ZMS13, NL19]. Leveling [CHK10, DY14]. Levelled [LRY^{+}15]. Leveraging [KSC^{+}14, LSS13, MGW14, MJWT16, QPG10, RTL^{+}18, SX12, SL14a]. LFSR [AK16, LCH13]. Lifetime [AK19, CBTU14, FKMK16, GGL^{+}14, JSH^{+}17, KLK18, LK16a, SMA19, WWM16, YyHL11]. Lifetime-Aware [LK16a]. Lifting [TWTT11]. Lifting-Based [TWTT11]. Light [SWF^{+}19]. Light-Weight [SWF^{+}19]. Lightweight [BFMT16, BDM^{+}19, BKL^{+}13, CXLX15, KAH18a, KE19, LSG^{+}15, RLSK18, STE17, SL13, VBR^{+}13, VAN^{+}18]. Like [DJN17, LYCT10, LJY^{+}15, Ose11, Tsa13, YLL16]. Likelihood [DAPS14, LCT11]. Lilliput [BFMT16, ST18a]. Limitation [Lee17]. Limited [EBE13, RF14, TCK^{+}18]. Limits [BK12]. Line [BCSR14, BCD^{+}16, FSPD16, GY14, LJY^{+}15, MG11a, NZLK14, ROGHNB^{+}18]. Line-Like [LJY^{+}15]. Linear [BCMJ10, CC16, DP13, DEE17, HWL^{+}14, HCC^{+}12, KO14, LLQ^{+}14, NZC11, PvdGG12, WHL^{+}12, WRW16, XXBL17]. Linearly [ST18a]. Lines [AGFM11, DSR15]. Link [GY16, GLH^{+}19, LGF^{+}15, SRCbL^{+}15, WTY^{+}14, YCCWC15, ZC13]. Link-Layer [GY16]. Link-Length [WTY^{+}14]. Links [CA12a, GHG^{+}14, PRGBSAC19, TMS^{+}14, YMK^{+}17]. Liquid [SVAB14]. Liquid-Cooled [SVAB14]. List [Ano10a, Ano11b, Ano12b, Ano13b, Ano14b, Ano15b, Ano17b, Ano18b, Ano19b, Ano16b]. Lithography [LZZ17a]. Little [JYL^{+}17]. Live [ECJ^{+}16, XLL^{+}14, ZRS^{+}16]. Lizard [MSS^{+}18]. LNS [CI16]. Load [ADOKM10, BR13, CLS14, HC13b, HHW^{+}18, JLC10, JC11, JR17, KRP18, PBL16, Pom12c, QJM^{+}10, RKZ16, SKPC15, SMTK12, SLS^{+}12, XAYL15, XLF15, YCLH16, ZV14]. Load-Balancing [PBL16, RKZ16, SLS^{+}12]. Load-Demand [XLF15]. Loaded [JC11]. Loading [SRCK10]. Loads [CC16, ZR15b]. Local [AVG^{+}15, BWV15, CFL^{+}18, HCH15, LKT13, LWL^{+}16, LCW^{+}15, LMT13, PTD^{+}12, ZDP^{+}15, ZL15]. Local-Deadline [HCH15]. Local-Recoding [ZDP^{+}15]. Locality [CG18, FBWMM13, GZC^{+}17, HWZ^{+}17, HXVF12, JZLD10, KS14, KGGJ14, LCY^{+}16, QZC15, QGPZ13, SH12, XJFH15, ZJS14]. Locality-Aware [CG18, HWZ^{+}17, JZLD10, KS14]. Locality-Preserving [SH12]. Locality-Sensitive [HXVF12, QZC15, QGPZ13]. Localization [AEKT15, JGHD11, LYOB15, MWWT13, XLW14, YZF^{+}10]. Localized [MMB14, YyHL11, YLA10]. Location [CWTT13, KLK^{+}14, LCT11, LHY13, ZZX^{+}15, ZLYS15]. Location-Aware [CWTT13]. Location-Based [ZLYS15]. Location-Oblivious [KLK^{+}14]. Locations [CTS13]. Lock [CT13, CWCS15, KLJ^{+}14]. Lock-Free [CT13]. Locking [CWCS15]. Loeffler [CND^{+}18]. Log [GWM^{+}17, KBO^{+}19, MLE14]. Log-Block [GWM^{+}17]. Log-Structured [MLE14]. Logarithm [Bra10, LP17, VB13]. Logarithmic [AC11, CHCK12, FML10, KBP13, LOC^{+}16, LJ15]. Logarithms [LOC^{+}16]. Logging [YHT^{+}16]. Logic [AR12, AD16, AGCD16, BDDL18, BGPV10, Cil11, CCLH10, EKA17, EFPC16, ISC15, Ibr16, LWK11, NI11, NYHB16, Pom14, PSL17, QLR^{+}11, RMKR12, RZZ^{+}15, RLRL19, ST12, SKM^{+}13, Tho15, TLL12, ZMR^{+}13, ZJH^{+}14]. Logic-Chain [TLL12]. Logical [LLQ^{+}14]. Logics [FLS16, GSF^{+}10]. Long [WXLL13, WXLY15, XLW14]. Long-Bounded [XLW14]. Longest [CWZC13, LLLP14]. Look [DJN17]. Lookup [HY12, JP13, LP12, LYS10, ML16, dLSGDR17]. Lookups [CLS10, CKKS14]. Loop [BBK10, CS11a, DZ10, EFGT18, GLXY13, KGC14, QLH^{+}16]. Loop-Abort [EFGT18]. Loop-Based [DZ10]. Loosely [PBL16]. Loss [KCY18, SRR^{+}16]. Loss-Aware [SRR^{+}16]. Lossless [XDZ11]. Lossy [DN11, GDY15, LLZ^{+}17, dRV12]. Low [AH10, ACW^{+}11, AVS^{+}14, AS12, ARM13, BDDL18, BR13, CMLS15, CSCW13, CJA^{+}16, CLL^{+}14, CJK19, CYL^{+}14, FFISC13, FHW18, GC16, GNSR14, GHG^{+}14, HN11, HK15a, HMS^{+}12, JCK15, JHQL16, JLLH19, KLK^{+}14, KBP13, KHZ17, LK15a, LYS14, LCL17, LOC^{+}16, MKT^{+}11, MKFM13, MAD14, MKAY11, MKRM11, NCD^{+}17, NC11, ORBM13, OPAGS14, OKD^{+}16, PLM16, PvdGG12, PPP13, PRBM13, PROM15, QSYS16, QLH^{+}16, RZPX19, RBMO11, RM15b, SP16, SKH16, SL10, SR14, SBI12, TW10, TKT16, WWY^{+}16, XJFT16, YAG19, ZOK^{+}19, ZM10]. Low-Complexity [ARM13, OPAGS14]. Low-Cost [HK15a, HMS^{+}12, MKFM13, TKT16]. Low-Duty-Cycle [GHG^{+}14]. Low-Error [LOC^{+}16]. Low-Latency [CLL^{+}14, FFISC13, LYS14, MKAY11, QSYS16, RZPX19, RM15b, SL10]. Low-Level [MKT^{+}11, OKD^{+}16]. Low-Memory [LYS14]. Low-Overhead [KLK^{+}14, PPP13]. Low-Power [GC16, JHQL16, JLLH19, KBP13, LK15a, LCL17, MKRM11, NC11, PvdGG12, SP16, SKH16, YAG19]. Low-Profile [AVS^{+}14]. Low-Voltage [ACW^{+}11]. Lower [Fuj11]. LRW [CFMA19]. LS [QGPZ13]. LS-Sig [QGPZ13]. LU [JC12, WDSP12]. Lubricating [TZL^{+}14]. Lyra2 [ASBdS16].

m* [AKJ^{+}13]. m*-Tree [AKJ^{+}13]. MAC [Kim15, LCLL15, CCRL19, CJ13, DMA^{+}15, HWK15, HWX15, LMC^{+}15, SMB^{+}15]. Machine [JJK^{+}11, LT15, MSG14, Man16, RSN^{+}18, VTW16, XLJ16, ZRS^{+}16, ZL18, ZJXL11]. Machines [APP12, DKW15, ECJ^{+}16, Hie11, Hie13, JAS^{+}15, JKJ^{+}10, KJL11, KP15, KT12, LLQ^{+}14, Pip11, SCSL12, SP10, WGLL13, WZL15, WLLZ16, XLL^{+}14]. Macho [MHK15]. Macro [JC11]. Macrocell [SBP^{+}14]. Macrocells [VPS^{+}12]. Macrochip [ZGY14]. Macroprogramming [PP10]. MACs [AP14]. Made [SD18]. Magnetic [WFY^{+}17]. Magnitude [EBE13, KN11a]. Magny [RCFP^{+}12]. Magny-Cours [RCFP^{+}12]. Main [GBGI18, HNV19, HZX^{+}14, JYL^{+}17, SMA19]. Maintaining [LHC^{+}14, LXL^{+}13, LcZLG19, RHC^{+}14]. Maintenance [CSJ^{+}11, FEP^{+}12, LCX^{+}16, SLC15a]. Majority [AGCD16, PSL17, RLRL19, TSA^{+}19]. Majority-Logic [AGCD16]. Making [XLL^{+}14]. Malicious [SWWC11, TM18]. Malleability [MBC^{+}13]. Malleable [SIVH16]. Malware [CXZ13, OKD^{+}16]. Malwise [CXZ13]. Managed [ASE17]. Management [AO11, AE11, ARS16, BLN^{+}15, BBP^{+}13, CVPS19, CK11, CCY^{+}16, CKH15, CGL^{+}13, CCP^{+}13, Cro14, DYW15, DA12, DGC^{+}15, EKJ^{+}10, FFCB14, FBR^{+}12, FAA10, GWMB13, GBD^{+}15, GWM^{+}17, HMR^{+}17, HWZ^{+}12, HNV19, HWZ^{+}17, JAJK15, KIJ14, KPS^{+}17, KFB^{+}15, KSJ^{+}12, KCRG14, KKJH19, KL16, LKLM15, LSA18, LZZZ13, LLD^{+}16, LHTG15, LWH^{+}16, LPL10, MM16, MOYB12, ML13, MTBB10, RDEN10, SSJ^{+}18, SIVH16, SYK14, SBW^{+}16, WP16, WSZ^{+}16, WW16, WJY^{+}17, WGLL13, XLS^{+}12, YYP^{+}16, ZJS14, ZL16, ZDYZ13, ZDYZ14]. Management/Monitoring [CCP^{+}13]. Manager [XHLW19]. Maneuvering [WF14]. Manipulation [VGF16]. Many [BLKM^{+}18, DYW15, DYHX16, GP14, HRM^{+}16, HMR^{+}17, KP13, LYT^{+}16, LB13, MMC15, RVC^{+}15, WSL^{+}18, WhCCC12, ZCY^{+}16]. Many-Core [BLKM^{+}18, DYW15, DYHX16, GP14, HRM^{+}16, LB13, RVC^{+}15, SNM16, WTZ^{+}19, WhCCC12, ZCY^{+}16]. Many-Cores [HMR^{+}17]. Many-to-Many [KP13]. Manycore [CA12b, CDK^{+}18, JKD^{+}19, LY18, PKC^{+}17]. Manycores [MYHL16]. Map [THGT13]. Mapped [SNY^{+}10]. Mapping [CAbZM18, CQW^{+}15, CCK^{+}16b, GSG^{+}15, HNB^{+}12, JK15, KN12, LKJ15, LSC11, LYJ^{+}18, MMCS18, OOD^{+}17, PP10, TSA^{+}19, YCKH16, YLGD19, ZCZ^{+}19, ZCY^{+}16]. Mappings [MC11]. MapReduce [CSPC12, CZL^{+}17, JSE14, LZA^{+}16, XLC14, YWQX15, ZDP^{+}15]. MAPS [KAH^{+}15]. MAR [WSZ^{+}16]. Marathon [DJN17]. March [CSW^{+}15]. Marginal [LY17]. Margins [CTL^{+}17]. mARGOt [GVPS19]. Market [FLL14]. Markets [BBVL14]. Marking [FBR^{+}12, YZGG16]. Markov [MTFK19, ZOD13]. Mashup [ZCL^{+}16]. Masking [KCY18, KN11b, LHL13a, ORBM13, USH19]. Massey [Red14]. Massive [BSM^{+}14, CLW^{+}15, WGR^{+}14]. Massively [KAH^{+}15]. Master [CAGM14]. Master-Worker [CAGM14]. Mastrovito [ARM16, LMZQ17]. Match [CW16]. Matching [CWZC13, DYW15, GBA18a, LP13a, LH12a, LLLP14, LLCC13, LBS15, MGW14, RSG^{+}19, YP12, Yun12, ZS13, ZLN11, ZYY18]. Matrices [CJK15, IRMM^{+}16]. Matrix [BFMT16, CNH13, CLL^{+}14, HF15, HMNN12, HN13, IRMM^{+}16, KEK16, KHZ17, NZ15, PCHS16, PCHS18, PGvdG14, RM15b]. Matrix-Vector [HF15, PCHS18, RM15b]. Max [LZ14, XLL15]. Max-Bisection [LZ14]. Maximization [LMC^{+}15, MLOL15, RLX15]. Maximizing [AGFM11, CS15, GSK12, WL13]. Maximum [AT16, DAPS14, GPN11, LCT11, LCW^{+}15, YCZ10, YUGD14]. Maximum-Likelihood [DAPS14]. Maxterm [YHH^{+}12]. MBs [CLZ19]. MBU [WNKL16]. MC [LRP^{+}18]. MC-Fluid [LRP^{+}18]. McEliece [GV14, SWM^{+}10]. McLaughlin [DCCK18, GL19]. MCMC [LMB17, MB16]. MDPC [HC17]. MDRUs [MTFK19]. MDS [FSL^{+}17]. Me [YGS15, CXLL16]. Me-CLOCK [CXLL16]. Mean [GPN11]. Meandering [AEKT15]. Measure [SRCbL^{+}15]. Measurement [NL14, SQJ^{+}15, WS14]. Measurements [KGC14]. Measures [AD10]. Measuring [CFMS14, GNSR14]. Mechanism [DKK16, HK15a, JC11, KLKL13, LL11, LLS^{+}16, MNGV16, NZ14, PR14, SWWC11, WZLS16, Zot10]. Mechanisms [BPBBL13, BCC^{+}16, CAGM14, FLJ14, HZL^{+}16, KSEG15, MFG16, MFG14]. Media [KLLK11, YW12]. Medical [FGS^{+}15]. Meets [CXYC16, RBG^{+}19, MOS14]. Membership [FHR14, HXVF12]. Memetic [LZ14]. MemFlex [ZLSI17]. Memoriam [Zom15b]. Memories [AVG^{+}15, CMM15, DPS11, GBGI18, HTH15, HZX^{+}14, JSA17, LLW^{+}17, LGMP10, NL16a, PO13, SMA19, SMRM17, SKH16, VCG^{+}12, WNKL16, ZZYZ14, vdBGLGL^{+}16]. Memory [ALBP14, AAR19, AKJ^{+}13, ASBdS16, AISA16, AKTB18, AH13, BQP^{+}16, BD15, BPG16, BBB^{+}17, CFL^{+}18, CVPS19, CVMA10, CK11, CHH^{+}13, Cha14, CCK^{+}16a, CHTD19, CXLL16, CWL^{+}17, CLOL18, CLM^{+}19, CLZ19, CC11, CRG^{+}13, DYHX16, DHW^{+}19, DMK^{+}15, DSG^{+}19, DCV^{+}12, DY14, DW10, EKJ^{+}10, FYSK14, FZL^{+}14, Fin10, GBO^{+}16, GGA^{+}17, GBD^{+}15, GDJZ18, GWM^{+}17, GNSR14, GIKR19, HCY18, HCCG10, HPR16, HNV19, HGCT13, HHKW12, HCC^{+}12, HWZ^{+}17, IS11, IS14, JJK^{+}11, JSC^{+}17, JSC^{+}19, JYL^{+}17, KLLK11, KCRG14, KO14, KAH^{+}15, KKJH19, KCS14, LP13a, LCC10, LMNP11, LKLK13, LBN14, LK14, LYB15, LWKA15, LH16, LK16b, LKH16, LKK^{+}19, LR10, LXW^{+}19, LYS14, LWL^{+}16, LLD^{+}16, LZZ^{+}17b, LJ13, MMC15, MWZ^{+}17, MB12a, MBD^{+}17, NKEM11, NL15a, NL19, OGH^{+}14, PLM16, PDXZ13, PN16, PPKW12, PCLN15, PKY19, PRM19, QML^{+}15, QGPZ13, RCC14, ROGHNB^{+}18, SSGB19, SNY^{+}10, SCZ^{+}16, SJD^{+}18]. Memory [SPC^{+}18, SP12, SRHC12, SBW^{+}16, SZL^{+}16, TPR16, UMN18, VSC^{+}19, VKS^{+}16, VTW16, VAN^{+}18, VMB19, WZLX12, WS14, WGW^{+}15, WWY^{+}16, WWY^{+}18, WLY^{+}14, WFT^{+}19, WDSP12, WSXZ13, XZL^{+}19, XHLW19, YWW^{+}16, YPB^{+}16, YYW^{+}16, YYC12, YYP^{+}16, YAGB17, ZCZ^{+}19, ZL16, ZLSI17, ZFHC19, ZOK^{+}19, ZHQX19, OWP16]. Memory-Aware [JJK^{+}11]. MEMORY-Based [OWP16]. Memory-Centric [BQP^{+}16, YPB^{+}16]. Memory-Efficient [CXLL16, KCS14, LP13a, LJ13]. Memory-Processor [MBD^{+}17]. Memristive [RMKR12]. Memristor [HTH15, RKR15, SKM^{+}13]. Memristor-Based [HTH15, RKR15, SKM^{+}13]. Memristors [RKR15]. Merge [LBSK17]. Merged [DKG13]. Mergesort [LRP16]. Merkle [LRY^{+}15]. Mesh [CHC^{+}15, GSH^{+}14, GP14, RKZ16, SZS14, TLP18, WLS18, ZR15b, ZBW17]. Mesh-Based [CHC^{+}15, WLS18]. Meshes [CS11a, Zot10]. Mesoscale [XYF^{+}15]. Message [AEP18, FAA10, LCT11, WGZ^{+}15, Zom12a]. Message-Efficient [LCT11]. Metadata [GWZ^{+}10, PP11]. Metaheuristic [LMT13]. Metal [YXZZ14]. Metastability [FFL18]. Metastability-Containing [FFL18]. Meter [DJN17]. Method [BCK^{+}16, CYL^{+}14, DAS14, GKB^{+}10, HN13, JCK15, KRP18, KEK16, KS10b, KL13, LZZ17a, LXK12, LJ13, MRW^{+}15, MH19, NTR14, OHCK17, OCK17, PP14, PCHS16, PB11, RSU17, RSNK17, SB10, SL13, ST12, SZDL14, USH19, WWT^{+}18, WTZ^{+}19, WYL^{+}15, WLG^{+}19]. Methodology [BGM^{+}13, BMS11, CSCW13, GBA^{+}18b, Iko15, JAJK15, JJC14, LKJ15, MNFA14, PWTS16, RTRM19, RRS^{+}16, RBG^{+}19, WS14, ZGR13]. Methods [AE11, AS16, DS14, EDL^{+}14, FBE^{+}18, KVV10, LCY^{+}13, ROH17, WNCH17, WLS18, ZOD13]. Metric [ABA07, Jha13, OP15, Pom13a, SIB13, WS14, ZWX12]. Metrics [EYBK15, GSF^{+}10, LHL13b]. MIC [XYF^{+}15]. Micro [DEE17, KAH18a, RM15a, VED^{+}16, YEG^{+}15]. Micro-Architectural [RM15a]. Micro-Architecture [DEE17, KAH18a, VED^{+}16, YEG^{+}15]. Microarchitectural [CVMA10, DJO11, JCY^{+}13, LDP10]. Microarchitecture [CPRH16, LCH13, PVKA14]. Microarchitectures [BPG16]. Microarray [GAC14]. Microbiome [LDB^{+}17]. Microcontroller [PRM19]. Microcontrollers [BCD^{+}16]. Microeconomic [NH10]. Microprocessor [BCSR14, CPRH16, DYW15, DYHX16, KMJ^{+}11, MKT^{+}11, SDP^{+}12, VSC^{+}19]. Microprocessor-Based [SDP^{+}12]. Microprocessors [EGVFC^{+}12, FRB^{+}18, KKC15b, MTGM12, MMTM15, OPV^{+}17, SNM16]. Microprotocol [VECD13]. Midpoint [Joh17]. Midpoints [JLMP11]. Migrating [YLH10]. Migration [GBO^{+}16, HV14a, LR18, LLX^{+}17, LCLC19, MRW^{+}15, RSNK17, RSN^{+}18, TKL^{+}14, WJM15, XLL^{+}14, ZRS^{+}16]. Migrations [LWH^{+}16]. MIHST [BCSR14]. Miller [LT14]. Millimeter [DCY^{+}13]. Millimeter-Wave [DCY^{+}13]. MIMD [NVB16]. Mimicking [YGS15]. MIMO [CWY13, LZ15]. Min [GLH^{+}19, XLL15]. Min-Entropy [GLH^{+}19]. Min-Max [XLL15]. Mini [FZL^{+}14]. Mini-Rank [FZL^{+}14]. Miniaturized [GJ14]. MiniBench [YEG^{+}15]. Minimal [ARH^{+}18, EDL^{+}14, HHKW12, MKLW14, RBR13, SBM15, SJD^{+}18, SG12, ZBW17]. Minimal-Memory [HHKW12]. MINIME [DSKH15]. Minimization [BCTV19, Cha14, LHL^{+}15a, LGMP10, XSR15, ZLG^{+}15]. Minimizing [DSG^{+}19, HT12, RSNK17, YCZ10, ZC13]. Minimum [DVUS14, FEP^{+}12, GPN11, KHPP16, KLT16, LYY16, LSX13, QPG10, SG13, WTY^{+}14, WCLY16, YLY15b]. Mining [HBR11, RGK15, SKC^{+}14]. Minus [BT16]. Miss [LK15b, SV18]. Misses [RXC^{+}15]. Missing [LLM^{+}15]. Mission [LXJD15]. Mitchell [KBO^{+}19, LJ15]. Mitchell-Based [LJ15]. Mitigate [GBA18a, JSC^{+}19, RBG14]. Mitigating [CJK19, GBGI18, HWK15, KCY18, KCKL19]. Mitigation [AS12, MLS19, SKH16, dOPSR16]. Mixed [ABA07, BBD^{+}12, BM13b, BDBB18, CAbZM18, CGL^{+}18, GGA^{+}17, GIW18, LRP^{+}18, LGS^{+}18, RMB^{+}12, ZCK19]. Mixed-Critical [BM13b].

Mixed-Criticality [BBD^{+}12, BDBB18, CAbZM18, CGL^{+}18, LRP^{+}18, LGS^{+}18]. Mixed-Precision [ZCK19]. Mixed-Radix [ABA07, GIW18]. Mixed-Time-Criticality [GGA^{+}17]. Mixing [Ged14]. Mixture [HFG^{+}17, RF14]. Mixture-of-Experts [RF14]. MLC [NL19, CK11, CCC^{+}17, CKD^{+}17, GWM^{+}17, LK14, ZZJ^{+}19]. MLC-Based [CK11]. MLC/SLC [CCC^{+}17]. MM* [CH11]. MMU [EKJ^{+}10]. Mobile [BWV15, Cao12, CS15, CKH15, FFCB14, GDY15, GCF^{+}16, HZL^{+}16, KCRG15, KkC15a, KKC17, LLCH13, LH16, LSX13, LHH14a, LCT11, LSL15, LZZ^{+}17b, MCC12, RSN^{+}18, SMP16, SKM14, SD13, TH11, TZL^{+}14, WLYY16, XWY10, XWH14, YCCJ15, YCLH16, ZWL15, ZLW^{+}17, ZMY11, ZY12, ZCC^{+}14, ZLYS15]. Mobile-Cloud [LSL15]. Mobility [AEKT15, ASTU10, OPZ15, SKPC15]. Möbius [Kůr12]. Modal [LT15]. Mode [AHK10, HZ11, JRS^{+}15, QLH^{+}16, SWWC11, ZLBB19, PVKA14, DFP^{+}13]. MoDe-X [PVKA14]. Model [Ano10f, BFR^{+}15, CMB13, CMS10, CH11, CYHC14, CCH15b, CC16, Dar15, DAPS14, FLS16, GLH^{+}19, HXVQ15, HSH^{+}10, HK13b, HWL^{+}14, JRJ^{+}18, JAD^{+}18, KGP15, LH12a, LWKA15, LK18, LLZ^{+}17, LWZ18, LH12b, LKT13, MD16, MTFK19, MY10, NH10, NS13, SMA19, SAR^{+}11, SIB13, SVAB14, TLP17, Tho15, Tsa13, WJL^{+}12, WP16, WHYS16, XKT^{+}15, XYHD17, YLL16, ZLBB19, ZYC16, ZYY10, ZZM^{+}15]. Model-Based [XKT^{+}15]. Model-Driven [CMS10, SAR^{+}11]. Model-Free [WP16]. Modeling [ADP^{+}15, BTBB14, BG12, CJSM17, CA12b, CSW^{+}15, DMXY14, DDNT19, GBA^{+}18b, HL10b, IPS17, KM11, KMLH11, LZZZ13, LMB^{+}16, MBM11, MHH^{+}17, MKM14, NS13, RKT19, SC11, VED^{+}16, WZCG16, XYF^{+}15, YZF^{+}10, ZGB^{+}15]. Models [AFH^{+}10, BD15, BGM^{+}13, CYA13, CCD12, FNS16, HFG^{+}17, LK10, LCY^{+}13, LHTG15, SD14, YMT13, YHT^{+}16, ZYL15, ZRL15]. Modern [KMJ^{+}11, LLD^{+}16, MKT^{+}11, MTGM12, MMTM15, MYW11, WS14]. Modes [AR17, PC16]. Modified [AO12b, CLC^{+}16, Red14, SJS^{+}14]. Modify [LKK^{+}19]. Modular [BT16, CYC^{+}16, FFISC13, GL19, HMA^{+}10, HGEG11, IGLM15, IDG^{+}17, KwPK^{+}15, KS12, KVV10, LP13a, LYS14, LNL^{+}19, PVKA14, SMRM17, SL10, YFCV14, ZYY10]. Modulation [ZM17]. Module [LKK^{+}19]. Modules [CZP^{+}18]. Moduli [GL19, Hia16, Hia17, Sou15]. Modulo [Dum14, HMC11, VD12, Wal19]. Modulus [SEY14]. Moments [LPL12]. Monitoring [BWCW15, CCP^{+}13, GEN^{+}17, GDJZ18, GLTC16, KKP^{+}16, ML13, OYP^{+}18, RVL^{+}14, RVC^{+}15, WLLZ16, ZNL18, ZL18, ZC13]. Monitoring-as-a-Service [ML13]. Monitors [HCG^{+}16, YZF^{+}10]. Monotone [TSA^{+}19]. Monotonic [PP14]. Monte [ZOD13, KN13]. Montgomery [CS11b, CYC^{+}16, DCCK17, DCCK18, GLP^{+}12, HRM11, HGEG11, KVV10, LYS14, NTR14, SL10, WF12]. Montgomery-Based [WF12]. MONTRES [LBSK17]. Morphable [CCC^{+}17, SZZ^{+}19]. MORUS [YWM19]. Most [CKKS14, XTF^{+}12]. Motion [RM15c]. Movable [ASTU10]. Moveable [YLA10]. Movement [WLYY16, YWQX15]. Moving [CT13, CHK10, LTL14]. MPPA [IDG^{+}17]. MPPA-256 [IDG^{+}17]. MPSoC [CYA13]. MPSoCs [ASS^{+}18, CCM14, MB12a, WLQS13, WLZ^{+}15]. MRAM [CFMA19, GBGI18, KCW^{+}17, ZOK^{+}19]. mRT [LKLK13]. mRT-PLRU [LKLK13]. MUCH [WLM15]. Muller [RMB^{+}12]. Multi [AMR18, ARS16, BD15, BCL^{+}17, BBB^{+}17, BSM^{+}14, CZ14, CLW^{+}15, CCY^{+}16, CCK^{+}16b, CG18, CvdBC18, CW16, DDN14, DMK^{+}15, DKLB15, DY14, DVUS14, FK15, GYC^{+}16, HCD^{+}16, HLY14, HH17, IPS17, JAD^{+}18, KTAvdS16, KIJ14, KAH19, KO14, KKC17, LKS^{+}14, LKH16, LRP^{+}18, LGH15, LT15, LWY15, LRY^{+}15, LFH^{+}16, LHH17, LLW^{+}18, LZW^{+}15, MS15, MWY^{+}16, MB16, MCXZ18, NL16a, Pan16, PCLN15, PMH^{+}14, PBE17, PM14, QZC15, QSYS16, RCC14, RTL^{+}18, SVD18, SDE^{+}17, SX12, SZG^{+}18, SNM16, SZW^{+}16, TWTT11, TFCY16, TLGM17, WLC^{+}15, WTZ^{+}19, XWL^{+}16a, YCCJ15, YMTV14, YYP^{+}16, ZLBB19, ZLG^{+}15, ZHM14, ZLX^{+}16, ZGWC15, ZRL15, NL19]. Multi- [SNM16]. Multi-/Many-Core [SNM16]. Multi-Armed [KTAvdS16]. Multi-Attribute [SX12]. Multi-Block [SZG^{+}18]. Multi-Channel [GYC^{+}16, HLY14, LGH15, XWL^{+}16a, YMTV14]. Multi-Cloud [SZW^{+}16, ZGWC15]. Multi-Constrained [FK15, ZHM14]. Multi-Core [BD15, BBB^{+}17, CZ14, CvdBC18, DMK^{+}15, JAD^{+}18, KIJ14, KAH19, KKC17, LKH16, LRP^{+}18, MB16, Pan16, PCLN15, PBE17, PM14, RTL^{+}18, TFCY16, YYP^{+}16, ZLBB19]. Multi-Cores [CCK^{+}16b, IPS17, LKS^{+}14]. Multi-Flow [QSYS16]. Multi-Function [DKLB15]. Multi-GPU [ZRL15]. Multi-Grained [CCY^{+}16]. Multi-Granular [LFH^{+}16]. Multi-Granularity [QZC15]. Multi-Hop [LWY15, MWY^{+}16]. Multi-Inherited [HH17]. Multi-Input [TWTT11]. Multi-Input/Multi-Output [TWTT11]. Multi-Interface [DDN14]. Multi-Keyword [ZLX^{+}16]. Multi-Layer [LLW^{+}18, SDE^{+}17]. Multi-Level [BCL^{+}17, NL16a, NL19]. Multi-Match [CW16]. Multi-Modal [LT15]. Multi-Mode [ZLBB19]. Multi-Objective [SVD18]. Multi-Output [TWTT11]. Multi-Path [DVUS14]. Multi-Player [BSM^{+}14]. Multi-Radio [LWY15, XWL^{+}16a]. Multi-Replica [LRY^{+}15]. Multi-Resolution [PMH^{+}14]. Multi-Socket [CG18, LHH17].

Multi-State [MCXZ18]. Multi-Task [ZLG^{+}15]. Multi-Tasking [CvdBC18]. Multi-Tenant [ZGWC15]. Multi-Threaded [TLGM17]. Multi-Threading [CvdBC18, MS15, RCC14]. Multi-Tier [LZW^{+}15]. Multi-User [HCD^{+}16, YCCJ15]. Multi-Version [WLC^{+}15]. Multi-Way [CLW^{+}15, DY14]. Multi/Many [WTZ^{+}19]. Multi/Many-Core [WTZ^{+}19]. Multiagent [Ano13g, KMLH11]. Multibeam [GGL^{+}14]. Multibyte [AMG17]. Multicast [ADOKM10, FG10, GY16, GY13, GGL^{+}14, GY15a, GY15b, LXL^{+}14, LHYZ13, SMG14, TH11, TC14, WJL^{+}12, WS15, WLS18, XCF16]. Multicast-Based [XCF16]. Multicasting [SO10, XWY10]. Multichannel [LWF13, XLTZ11]. Multichip [SMN^{+}17]. Multicore [BTBB14, CLS14, CCH11, CS11b, CWCS15, DSKH15, DW10, FJA^{+}17, GJ14, GCD^{+}11, GGSPM18, HWZ^{+}12, HV12, HV14a, HBR11, IHR^{+}16, JLC10, JJC14, KLJ^{+}14, KCL^{+}16, KH18, KCE^{+}18, KA19, LNCX18, LMC^{+}12, Man16, ML18, MHRARG^{+}14, MBB^{+}17, NEE18, OOD^{+}17, RCM^{+}16, RC14, RRK11, RJV^{+}18, SC11, TCHL18, YLML15, YPB^{+}16, YYW^{+}16, YMG16, YRG13, YAGB17, ZDYZ13, ZDYZ14, ZZ10, ZLN11, ZRL15, ZYY18, ZHM^{+}19]. Multicore-Aware [Man16]. Multicore/Multithreaded [RCM^{+}16]. Multicores [BZ15, FSPD16, FSPD17, KPS^{+}17, STK16]. Multicycle [Pom12d]. Multidimensional [TYWC10]. Multidomain [BPC12]. Multifactor [SL13]. Multiformat [GVGNCVM16, MLH12]. Multihop [CTS13]. Multilayer [HHLK12]. Multilevel [EGVFC^{+}12, HC13a, HJF^{+}13, NL15a, WSXZ13, WYL^{+}15]. Multimedia [KKC17, LKYC12, LGH15, MM16, MSC12, PAC^{+}12]. Multioperand [HVZ13, MLH12]. Multipartite [HWCH17]. Multipath [SLS^{+}12, WW14]. Multipattern [Yun12, ZS13]. Multiple [ALBP14, CLS14, CWZ13, CP10, DJA11, FM19, FLL14, FK15, GRM16, Hie11, Jha13, LQD^{+}16, LWK11, MKAY11, NDC^{+}13, NL16c, OCK17, PCZB11, Pom16b, PPND17, RWZZ14, TC16, WNKL16, YCZ10, ZYW^{+}16, ZLX^{+}16, ZCK19, ZMY11, ZCY^{+}16]. Multiple-Bit [GRM16]. Multiple-Parameter [NDC^{+}13]. Multiple-Precision [ZCK19]. Multiple-Queue [PPND17]. Multiple-Radix [DJA11, Jha13]. Multiple-Valued [LWK11]. Multiplexed [GCLC11]. Multiplexer [SMCN18]. Multiplexer-Based [SMCN18]. Multiplexing [DYW15]. Multiplication [ARM15, ARM16, ABH^{+}13, ACO12, AK14, CMO^{+}16, CS11b, CYC^{+}16, Cil13, DCCK17, DJJ^{+}08, FP19, Gio12, GNTS13, GJ15, GL19, GIW18, HK13a, HRM11, HN11, HMNN12, HGEG11, IRMM^{+}16, KEK16, KVV10, KHZ17, Lee12, LPW10, LYS14, LNL^{+}19, LJL13, MH15, MHML15, NWA12, NR15, NTR14, PCHS14, PCHS17, PRBM13, ROH17, SL10, VAB14, YFCV14]. Multiplications [ARM13, DS14, ERRM16]. Multiplicative [Dum14, MPZ15, RMERM19, Wal19]. Multiplier [AS10, ARM13, BNP10, CLL^{+}14, DHM16, FF16, HMC11, KS10b, LMZQ17, MCT19, NWA11, PCHS18, RM15b, WF12, WS10, ZM10]. Multiplier-Dividers [AS10]. Multipliers [ARM16, CHN14, CLC^{+}16, CDL^{+}17, DJA11, DJA14, DRC14, Fan16, GT19, GPN11, HF15, Ima18, JHQL16, KBO^{+}19, LAAM11, LQW^{+}17, LCY^{+}19, MHHS17, SP16, TAM^{+}16, VAM10, VALK19]. Multiply [CCRL19, WF17, ZCK19]. Multiply-Add [ZCK19]. Multiplying [PP16]. Multiprefix [HHY11].

Multiprocessor [BKH^{+}13, CA12a, Fuj11, HTC13, HWL^{+}14, KGGJ14, LSSE15, Lee17, Li12b, LCX^{+}16, LLX^{+}17, LKT13, LYJ^{+}18, MWW14, MW13, OP15, Tsa13, YHC19, YHV13]. Multiprocessors [Ano11f, BPT10, CNJ14, DAS14, FBWMM13, KSEG15, KK10, LMJ14, LKMSA16, ST17, ST18b, WMW12, WXW^{+}14, ZCY^{+}16, ZMS13]. Multiprogrammed [CPS^{+}10, CA12b, HGCT13]. Multiradio [CWZ13]. Multireceiver [FHH10]. Multiresource [GSX^{+}13]. Multiscale [NL16b]. Multiscale-Symbol [NL16b]. Multisize [LPL10]. Multispeculative [DHM16]. Multistate [AXS^{+}10]. Multisuffix [HY12]. Multitask [LKLK13]. Multitasking [CGL^{+}13]. Multithreaded [AT16, JJC14, LvH12, RCM^{+}16, RRK11, WLM15]. Multithreading [CCH11, CCK^{+}16b, GSL10, YG10]. Multithreshold [ST12]. Multituple [YFJ^{+}14]. Multivalued [AXS^{+}10, LB15a]. Multivariate [dAJM14]. Multiway [CLS10, HN13, PCHS16, SWZG11]. MuR [LRY^{+}15]. MuR-DPA [LRY^{+}15]. MuSA [dAJM14]. Mutation [PFGB14]. Mutually [LKS^{+}14, ZV14].

N [KS12]. N-Modular [KS12]. NAF [ADJ12, TX16]. NAND [AKJ^{+}13, Cha10b, CQW^{+}15, CWL^{+}17, CCL^{+}18, CC11, CYL^{+}14, GWM^{+}17, JSH^{+}17, KLLK11, LKLK13, LK16a, LCY^{+}16, PDXZ13, PPKW12, ROGHNB^{+}18, SKM14, UMN18, WW16, WLY^{+}14]. NAND-Flash-Based [Cha10b]. Nano [LT15]. Nano-Machine [LT15]. Nano-Scale [LT15]. Nanophotonic [MKLW14]. Nanoscaled [SRCK10]. Nanotechnology [BKP16, BKP16]. Narrow [HK15a]. Narrow-Width [HK15a]. Nationwide [JGG^{+}14]. Native [FGS^{+}13, SWWC11]. Navigation [CYHL14]. NBTI [ORBM13]. NCCloud [CHLT14]. Near [FSGAB^{+}16, MHK15, RVC^{+}15, SSGB19, TYY^{+}16]. Near-Memory [SSGB19]. Near-Optimal [RVC^{+}15, TYY^{+}16]. Near-Threshold [FSGAB^{+}16, MHK15]. Nearest [HV16, KMP11]. Nearly [ZWLS15]. Necklace [HHKW12]. Need [LYT^{+}16]. Negative [DSR15, WHC^{+}15b]. Neighbouring [ST16]. NEON [SD18]. NeST [DYJ19]. Nesting [TPR16]. Net [CZ16, YLH13, RVL^{+}14]. Net-Based [YLH13, RVL^{+}14]. Netlist [AD13]. Netlist-Driven [AD13]. Nets [CCK10, HB11, XKT^{+}15]. Network [AAR19, ASTU10, Ano13g, BPBBL13, CHC^{+}15, CHLT14, CC16, CXYC16, CPL16, CJ13, CDK^{+}18, DYJ19, DKLB15, DKG13, EYBK15, FSR^{+}18, GCD^{+}11, GEvS10, GC14, GHL17, GSG^{+}15, GW16, GCL^{+}13, GIKR19, HMD^{+}17, HCSW15, HCG^{+}16, HGL^{+}15, HWE^{+}16, IHR^{+}16, JWL^{+}16, JWC12, JRC14, KC14, KGC14, KLK^{+}14, Kim15, KCRG15, KKY^{+}16, KH14, LKS^{+}14, LCLL15, LW15, LSHC15, LSW15, LYJ^{+}18, LKMSA16, LLL^{+}17, MMCS18, MFT^{+}17, MKLW14, NY19, PAP13, PP10, PBT13, PSND16, RDEN10, RKZ16, SKPK10, SLZX15, SLLG15, SPTC15, STK16, TYY^{+}16, TLP18, UVL^{+}13, VBR^{+}13, WZLX12, WJL^{+}14, WWM16, WZCG16, XCF16, YY14, YLGD19, YG10, YRG13, YYC12, YCZ10, YCCWC15, YCK10, ZWX12, ZR15b, ZL16, ZMW^{+}19, ZC13]. Network-Based [WZCG16]. Network-Coding-Based [CHLT14]. Network-on-Chip [CHC^{+}15, DKLB15, DKG13, EYBK15, GCD^{+}11, GC14, HMD^{+}17, HCSW15, HWE^{+}16, JRC14, KC14, KLK^{+}14, KKY^{+}16, LYJ^{+}18, LKMSA16, MMCS18, MKLW14, PSND16, SKPK10, WWM16, WZCG16, XCF16]. Network-on-Chip-Based [STK16, YYC12]. Network-on-Chips [YLGD19]. Networked [DLC^{+}13, JZLD10, Yam10, YLY^{+}15a]. Networking [SBP^{+}14, WLT^{+}16]. Networks [AO11, AO12a, ABB17, AEKT15, ASTU10, AK15, AB16, AD10, AD12, Amm14, Ano11d, BFMT16, BDP15, BWCW15, BWV15, BKV12, BBVL14, CMB13, CZP^{+}18, CFMS14, Cha10a, CCE^{+}18, CS11a, CWZ11, CBZ14, CS15, CTD^{+}16, CJG16, CWTT13, CLR13, CYC11, CDK^{+}18, CSJ^{+}11, CWY13, CGL^{+}13, CTS13, CCD12, CCP^{+}13, CRK10, CBTU14, CBVL16, CJK15, DDN14, DMXY14, DY12, DCL^{+}11, DLL^{+}12, EDL^{+}14, FB13, FTP13, FS10, GD17, GSH^{+}14, GDC^{+}16, GHG^{+}14, GY15a, GDY15, GY15b, GLTC16, HXVQ15, HBCC13, HMD^{+}17, HKWC14, HLJ14, HCZW13, HWX15, JK15, JGHD11, JRW^{+}14, JRS^{+}15, JWL^{+}16, JKY10, KGGJ14, KBO^{+}19, KKT15, KLT16, KL16, KH10, LRC10, LR13, LYOB15, LMC^{+}15, LH12a, LCL15, LWW11, LSS13, LS13, LXL^{+}14, LGH15, LYT^{+}16, LLZ^{+}17, LOH17, LSX13, LCHC14, LYCT10, LKLT12, LWY15, LXZH16, LCT11, LXK12, LZYL13, LSZ^{+}15, LJY^{+}15, LZS^{+}13, LYL^{+}19, MNFA14, MCC12]. Networks [MM16, MB12b, MHRARG^{+}14, MWY^{+}16, MD13, MMH14, MMB14, NH10, PLP^{+}13, PLZW14, QSYS16, RMB^{+}13, RL13, RGK15, RDEN10, RLX15, RRS^{+}16, RS13, RNS13, SKPC15, SMP16, SXLC15, SBH11, SPC^{+}16, SDE^{+}17, SSGB19, SJVR19, SCK10, SRR^{+}16, SG12, SG13, SL14a, SPTC15, SO10, SZS14, SKA10, Ste14, TYWC10, TH11, TSK16, THM^{+}14, VYEB18, VBR^{+}13, WJL^{+}12, WXLL13, WJL^{+}14, WXW^{+}14, WXLY15, WZM^{+}16, WLYY16, WLJ^{+}16, WCM^{+}16, WEH^{+}19, WGZ^{+}15, WHC^{+}15b, WS15, WW14, WLS18, XWY10, XCW^{+}10, XWH14, XWL^{+}16a, XWL10, XLTZ11, XWL^{+}16b, YKK^{+}15, YL14, YMK^{+}17, YZ15, YLL16, YASS14, YLA10, ZNL18, ZLG^{+}15, ZGY13, ZMY11, ZY12, ZWD^{+}16, ZLXW15, ZWL^{+}19, ZWC13, ZGW14, ZLYS15, dAJM14]. Networks-on-Chip [ABB17, Ano11d, CCE^{+}18, CRK10, DMXY14, EDL^{+}14, FTP13, KGGJ14, RMB^{+}13, SDE^{+}17, WXW^{+}14, WZM^{+}16, WEH^{+}19, WLS18, YMK^{+}17]. Neural [AAR19, CLW^{+}15, DYJ19, GIKR19, HMD^{+}17, KBO^{+}19, LYL^{+}19, LLL^{+}17, NY19, SSGB19, SJVR19, SKM^{+}13, XWL^{+}16b]. Neuromorphic [QWB^{+}13]. Neuroprocessors [ZMR^{+}13]. Neurosynaptic [TBC^{+}17]. Newton [Dum14, Wal19]. NextCell [ZZX^{+}15]. NFRA [PAP13]. NICO [WFT^{+}19]. Niederreiter [HC17]. NIPD [TJX^{+}17]. NN [FEM^{+}18, ZCL^{+}16]. NNPIM [GIKR19]. NO2 [WGR^{+}14]. NoC [CCM14, DCY^{+}13, GD17, JKD^{+}19, KAH18a, KCL^{+}16, KN13, LDB^{+}17, LY17, MSPK12, OMFH14, PB16, RVL^{+}14, RC14, SKEB18, SBP^{+}14, SMG14, TMS^{+}14, WLV^{+}14, WJY^{+}17, WZK^{+}19]. NoC-Based [CCM14, KN13, MSPK12, RC14]. NoC-Bus [RVL^{+}14]. NoCs [DSPB13, FCB^{+}19, LY18, MJW^{+}14, MWLJ15, SKEB16, XWLX17]. Node [FEP^{+}12, LY11, MMH14, SKA10, WXLL13, WXLY15, XP10, XLW14, YTM16]. Node-Disjoint [SKA10]. Node-Level [XP10]. Node-To-Node [SKA10]. Nodes [CCP^{+}13]. Noise [LHCL13, ZLY15]. Noisy [Cao12]. Non [AK19, BCK^{+}16, BIP^{+}17, Cha14, CHTD19, CC16, GL19, HZX^{+}14, HWE^{+}16, JAKD18, JSA17, JW16, KO14, LKBS16, Lee17, LTL14, LMZQ17, LXW^{+}19, LZZ^{+}17b, NL15a, RD18, SMB^{+}15, SJD^{+}18, SRR^{+}16, STR15, ST18b, SZZ^{+}19, SZL^{+}16, TJX^{+}17, TAH^{+}16, TC16, TIHM18, TAM^{+}16, WWY^{+}16, WWY^{+}18, WNKL16, WhCCC12, XZL^{+}19, YWW^{+}16, ZFHC19]. Non-Binary [NL15a]. Non-Blackbox [TIHM18]. Non-Blocking [HWE^{+}16, SRR^{+}16]. Non-Coprime [GL19]. Non-Fragile [BIP^{+}17]. Non-Ideal [SMB^{+}15]. Non-Intrusive [TJX^{+}17]. Non-Iterative [TC16]. Non-Linear [CC16, KO14]. Non-Numeric [BCK^{+}16]. Non-Parametric [LTL14]. Non-Preemptive [Lee17]. Non-Recursive [LMZQ17]. Non-Redundant [TAM^{+}16]. Non-Speculative [STR15]. Non-Uniform [ST18b, WhCCC12]. Non-Volatile [AK19, Cha14, CHTD19, HZX^{+}14, JAKD18, JSA17, JW16, LKBS16, LXW^{+}19, LZZ^{+}17b, RD18, SJD^{+}18, SZZ^{+}19, SZL^{+}16, TAH^{+}16, WWY^{+}16, WWY^{+}18, WNKL16, XZL^{+}19, YWW^{+}16, ZFHC19]. Nonblocking [GY15a, GY15b, Zhe10]. Noncoherent [CRG^{+}13]. Noncooperative [CWZ13]. Nondeterministic [Hie11]. Nonindexed [LOX^{+}13]. Nonintrusive [WhCCC12]. Nonlinear [GO10, KW14]. Nonlinearity [MM17]. Nonlinearly [SKM^{+}13]. Nonuniform [ZDYZ13, ZDYZ14]. NOR [PRM19]. Normal [ABH^{+}13, ARM13, DJA14, ERRM16, KEK16, NWA11, NWA12, RMERM19]. Normalization [ADJG19]. Normally [CFL^{+}18, RD18]. Normally-Off [CFL^{+}18, RD18]. Note [CL12, CQ14, MW13]. Notification [BBPQ15, JRW^{+}14]. Novel [BMS11, BC16, CSW^{+}15, CC16, CC11, GWZ^{+}10, IBH^{+}13, KCL19, KL13, LLCC13, LFJ^{+}13, LPL12, NL14, SSKL16, SWM^{+}10, SPH13, TLZV11, WXLL13, WS14, WSZ^{+}16, WNKL16, WZLS16, YLY^{+}15a, YFCV14]. NP [KLT16, XTF^{+}12]. NP-Completeness [KLT16]. NP-Hardness [XTF^{+}12]. NPAM [PBE17]. NROM [LLHC15]. NROM-Based [LLHC15]. NTRUEncrypt [DWZ18]. NUDA [ST18b, WhCCC12]. Null [YCW11, YLP15]. NUMA [KP15]. Number [ADI11, AC11, GKB^{+}10, KBP13, Kůr12, MG16, NM10]. Numbers [BLMM16, HV16, YUGD14]. Numeric [BCK^{+}16]. Numerical [RT14]. NV [RD18, WWY^{+}18, XZL^{+}19, YWW^{+}16].

NV-Clustering [RD18]. NV-Dedup [WWY^{+}18]. NV-eCryptfs [XZL^{+}19]. NV-Tree [YWW^{+}16]. NVM [LXW^{+}19, PBE17, SMA19]. NVM-Aware [PBE17]. NVMM [CLCS19]. NVRAM [BDM^{+}19, CCC^{+}18]. NVRAM-Based [BDM^{+}19]. nvramdisk [JW16]. Nyquist [LVJ16]. Nyquist-Sampled [LVJ16].

O [BBP^{+}13, DYHX16, GKD^{+}17, HWS^{+}17, HS18, HLC^{+}19, KSJ^{+}12, KRP18, LKBS16, SNM16, SYH17, SHH^{+}16, TAH^{+}16, ZL16]. O-Redirection-Based [HQLX15]. O/S [SNM16]. O/S-Orchestrated [SNM16]. O3 [NKEM11]. Obfuscation [FRB^{+}18, ZL19]. Object [KT12, YTD^{+}17]. Object-Based [YTD^{+}17]. Objective [CA12a, ML18, SVD18]. Objectives [CCC^{+}17]. Objects [CT13, LCT11, RKN^{+}18]. Oblivious [HMZ^{+}14, KLK^{+}14, KCS^{+}13, RL13, SBI12, WCL^{+}18]. Observability [KCY18]. Obtained [Ste14]. Off [CFL^{+}18, KWC^{+}16, LLW^{+}17, MAG^{+}17, MEBS17, RD18]. Off-Based [SC18]. Off-the-Hook [MAG^{+}17]. Offloading [XLF15, YZH^{+}15]. Offs [ADOKM10, ASBdS16, BCC^{+}16, BS14, BG12, GCAG16, KCL^{+}16, KN11b, ZZ10]. Offset [RBMO11]. OFWAR [WZ14]. OLAP [HNV19]. Oligopoly [FLL14]. Omission [Pom15b]. On-Chip [CKKS14, CDK^{+}18, HJBM14, JKY10, JJZ^{+}16, JWC12, LKS^{+}14, MNFA14, MKAY11, MD13, OHCK17, PVKA14, RRS^{+}16, RKZ16, SIVH16, VYEB18, VCG^{+}12, ZNL18, ZGY13]. On-Demand [CAbZM18, CQW^{+}15, LLLJ13, WZ14]. On-Line [BCSR14, BCD^{+}16, FSPD16, GY14, MG11a, NZLK14]. On-the-Fly [DHM16, Pip11, YLY15b]. ON-the-Run [LBSK17]. On/Off [SC18]. On/Off-Based [SC18]. One [CMRH17, MWLJ15, MMP13, PC16, RLRL19, TYY^{+}16]. One-Sided [TYY^{+}16]. Ones [LYT^{+}16]. Online [ASE17, AD16, BSM^{+}14, CVMA10, DY12, DRS^{+}16, EE10, GCLC11, GVGNCVM16, GBA^{+}18b, HRM^{+}16, HMC11, JWL^{+}16, KL16, MNGV16, MLH12, PSM17, QQW^{+}17, RCC14, SLLG15, ST12, SXCL14, Tse12, WHC^{+}15b, XWL10, XTW15, ZJL^{+}16]. OnlinePlus [Ano10f]. onto [HNB^{+}12, KN12, LSC11]. Ontology [MH19]. Ontology-Based [MH19]. Open [Ano13e, HTH15, TPR16]. OpenCL [CLOL18, HHC^{+}18, LWKA15, LWZ18]. OpenCL-Based [CLOL18]. OpenFlow [JPLP13]. OpenFlow-Based [JPLP13]. OpenMP [MB12a]. Operand [LK15a, WE12]. Operate [FSGAB^{+}16]. Operating [DCL^{+}11, IBH^{+}13, LMB^{+}16, LLW^{+}11, LLD^{+}16, ZMW15]. Operation [ACW^{+}11, HV14a, JHQL16, MYW11, YLH13]. Operational [RBIQ15]. Operations [APP12, Cil11, GC16, IRMM^{+}16, KPBC17, KMP11, SEY14, SYH17, SS12, XTW15, ZLWZ15]. Opinions [JWWZ16]. Opportunistic [CBVL16, GHG^{+}14, HKWC14, LCHC14, SKPC15, SMP16, XWH14, YZH^{+}15, ZGB^{+}15]. Optical [GY13, KH14, PTD^{+}12, PRGBSAC19, ZGY14]. Optimal [ABH^{+}13, AT16, BCL^{+}17, CFL^{+}18, CLS14, CCC15, CPL16, CYC11, CTS13, GM12, GZB^{+}15, GKC19, GDY15, HT12, KwPK^{+}15, KCS^{+}13, KEK16, LCA10, LP13b, LZZV16, LWF13, MBGS10, PLP^{+}13, RL13, RVC^{+}15, TYY^{+}16, TX16, VK15, WXLY15, WLT^{+}16, XAYL15, YLGD19, YASS14, ZWLS15, Zhe10, ZV14, ZL15]. Optimality [TKL^{+}14]. Optimally [WLQS13]. Optimised [CMO^{+}16]. Optimization [AEP18, AR17, AWFV13, CPL16, CvdBC18, CK15, DSR15, DZD^{+}16, DKW15, DDNT19, GYC^{+}16, GAFN15, GZG^{+}16, GW16, GLP15, HCSW15, HHLK12, HGL^{+}15, JP13, KM11, KSEG15, KAH19, KHZ17, LYT^{+}16, LHH14a, LSZ^{+}15, ML18, NYHB16, PWTS16, QML^{+}15, RC14, SKM14, SCJ^{+}16b, SZW^{+}16, TLGM17, WXW^{+}14, WWM16, WRW16, WSL^{+}18, XLL^{+}18, XSR15, YLML15, ZD13, ZGG^{+}16]. Optimization-Based [HCSW15]. Optimizations [BZ15, CVH^{+}13, GHL17, LHTG15, YHT^{+}16, ZAG19]. Optimize [HZX^{+}14]. Optimized [FML10, IRMM^{+}16, LNL^{+}19, MCT19, Yam10]. Optimizing [DWZ18, DEE17, FSL^{+}17, GKD^{+}17, HWSX17, HWZ^{+}17, LOX^{+}13, LTLC12, LLLJ13, NEE18, NL16c, RS13, SYH17, ZJS14, ZFHC19, ZZ19, ZJXL11, Avr13]. Optional [PC16]. Orchestrated [SNM16]. Order [BMZ17, CVMA10, DP13, LPL12, MY10, NKEM11, RRK11, RBG^{+}19, Sri10, WLZ^{+}15, WA10, YCW11, YLP15, ZMB18]. Ordered [AKJ^{+}13]. Ordering [BBPQ15]. Organization [KCL19, LR10, SBW^{+}16]. Organized [DKK16]. Organizing [GEvS10]. Oriented [GZC^{+}17, TC14, YHC19, ZL15]. Orthogonal [DRM16, NL15a]. OS-Level [cCWS14]. OS/Apps [ZLW^{+}17]. Oscillator [LB15b]. OSGi [LLCH13]. Other [FLS16]. Out-of-Order [NKEM11, Sri10, WLZ^{+}15, WA10]. Outgoing [Mon19]. Output [AHK10, TWTT11]. Output-Queued [AHK10]. Outsourced [LLC^{+}15, LQD^{+}16, WCH^{+}15]. Outsourcing [WRW16, WJF^{+}11]. Over-Collection [LDMQ16]. Over-Redundant [EJ15]. Over/Under [LY17]. Overclocking [KSC^{+}14]. OverCome [JLKR19]. Overhead [AS12, BS16, CFR^{+}14, CCW^{+}10, CJK19, EE17, KLK^{+}14, LKBS16, PPP13, RBMO11, RS13, TW10, WLQS13, ZLW^{+}17]. Overheads [LKH16, SC18, XJFT16]. Overlap [PCHS17, RLRL19, ADJG19]. Overlap-Free [PCHS17]. Overlay [AK15, LXK12]. Overscaling [AEGH19]. Oversubscribed [KFB^{+}15]. Overview [FLP^{+}13]. Owners [ZLX^{+}16]. Ozone [NKEM11].

P [GT19]. P1687 [ZICL12]. P2P [CS15, CLMM11, LSW15, PPB^{+}14, SL13, SLC15a]. Packed [CXZ13, JL11]. Packer [LTW^{+}12]. Packet [AHK10, BMS12, BSS14, BSS15, CW15, CW16, FBR^{+}12, GY13, IHR^{+}16, LFJ^{+}13, LW13, LPCW14, MFT^{+}17, MOYB12, RO11, SRCbL^{+}15, ST11b, XHZ14, YFJ^{+}14, YM11, YZGG16, ZWLS15]. Packet-Mode [AHK10]. Packets [KKH^{+}14]. Packing [CGJ^{+}10, SXCL14]. Page [AKJ^{+}13, GCF^{+}16, LBN14, LH16, LZZ^{+}17b, PBE17, SPC^{+}18]. Paging [KLLK11, WLY^{+}14]. Pair [PR10]. Paired [KP13]. Pairing [BDE^{+}11, CKM15, CMRH17, FVV12, KHPP16]. Pairing-Based [CKM15, CMRH17]. Pairing-Friendly [FVV12]. Pairings [LT14]. Pairs [PR10, VK15]. Pairwise [RNS13]. Pancyclicity [LH12a]. PAnDA [TLB^{+}17, WTBT13]. Papers [Ano10c, Ano11c, Ano11d, Ano11e, Ano11f, Ano13d]. Paradigm [BDDL18, DYJ19, KYZC19, MFT^{+}17]. Parallel [AR17, CS11b, CZP^{+}16, Cil13, CDL^{+}17, DN15, DYCG16, Fan16, Fin10, FM16, GRM16, GJ15, GAC14, HK13a, HWS^{+}17, HWSX17, HS18, HT16, HHW^{+}18, Ima18, KGD16, KAH^{+}15, KLC18, KT12, LR16, LR10, LLCC13, LWL^{+}16, LTW^{+}12, LB13, LMT13, ML18, MHRARG^{+}14, MBB^{+}17, MKRM12, MC11, NL15b, NL16a, NZ15, NR15, NTR14, PZZQ19, PRM16, PSL17, QJM^{+}10, QWB^{+}13, RKR15, RT14, RS10, RQ14, SNY^{+}10, TRA18, VAM10, WGR^{+}14, WZLS16, XP10, XLX^{+}14, YSZ^{+}14, ZLBB19, ZCZL16, ZCS16, ZYL15, ZM10, ZYY18, ZXX^{+}14, ZMRQ11]. Parallelism [HJF^{+}13, JDA^{+}16, LR16, LWF^{+}17, SHGW15, SJC^{+}17a, SJC^{+}17b, XLL^{+}18]. Parallelization [YLGE14]. Parameter [CYC^{+}16, DZLP14, KSJ^{+}12, NDC^{+}13, YFCV14]. Parameter-Aware [KSJ^{+}12]. Parameterized [WLZ10]. Parameters [MBF18]. Parametric [LTL14, YLH13]. Parity [GRM16, KLC^{+}16, MAD14]. Parity-Based [GRM16]. PaRS [ZHQX19]. Parsing [TLL^{+}13]. Partial [CYHC14, CLC^{+}16, DHM16, FSL^{+}17, JWC12, JRC14, LHPH15, MY10, XHZ14]. Partial-Terminal [LHPH15]. Partially [DSPB13, DKK16, Pom16a, RSU17, SKEB16, WLJ^{+}16]. Partially-Functional [Pom16a]. Participation [WZY16]. Participatory [LZZV16]. Particle [LKYC12, RM15c]. Partition [FF16, NTR14, SDL^{+}19]. Partitioned [LCX^{+}16, LS10c, YLML15, KMC17]. Partitioning [ASTU10, CTD^{+}16, EDL^{+}14, HWG^{+}14, JSC10, KCY18, KKT15, LP12, LZA^{+}16, MWW14, SJS10, WHYS16, YCCJ15, YHC19, ZRL15]. Partitions [LC16a]. Party [KKH^{+}14, QZL^{+}16]. Pass [ZGR13]. Passing [YKK^{+}15]. Passive [JGHD11]. Password [ASBdS16, LLD19, YRT^{+}16]. Password-Authenticated [YRT^{+}16]. Path [BDP15, DSY^{+}15, DVUS14, EDL^{+}14, Fen14, GEN^{+}17, KP13, NI11, RS13, SBP16, WTY^{+}14, XTF^{+}12, Yan14]. Path-Based [EDL^{+}14, GEN^{+}17]. Path-Dividing [Yan14]. Paths [Fen14, FK15, GDC^{+}16, SKA10]. Patient [LDP10]. Pattern [AK16, DSKH15, GSK12, LP13a, LLCC13, LWL^{+}16, MGW14, PPKW12, RSG^{+}19, ZLY15, ZLN11]. Pattern-Aware [DSKH15]. Patterning [LZZ17a, khR^{+}18]. Patterns [CCC15, CH14, LWF^{+}17, Pom13b, RGK15, SBH11]. Payload [SPH13]. PBC [RPM16]. PC [BSS14, BSS15]. PC-DUOS [BSS14]. PC-TRIO [BSS15]. PCLMULQDQ [JL11]. PCM [NL15a, NL19, CCY^{+}16, CJK19, LBN14, LWF^{+}17, MCM16, QLH^{+}16, ZZJ^{+}19]. PCs [KBH^{+}10]. PDG GEN [MNFA14]. PEAF [EKA17]. Pearl [HHKW12]. Pearl-Necklace [HHKW12]. PEBD [ZMRQ11]. Peer [CYJ^{+}10, HHM11, LSS13, LLLJ13, YY10, YY14]. Peer-Assisted [LLLJ13]. Peer-to-Peer [CYJ^{+}10, HHM11, LSS13, YY10, YY14]. Penalties [BCC^{+}16]. Pentanomials [HF15, Ima18]. Per-Job [YLH10]. Per-Resource [YLH10]. Perceptron [LLW^{+}18]. Perf [FSPD17]. PerfBound [SC18]. PERFECTORY [LCC10]. Performability [AXS^{+}10, MCXZ18, SCNS10]. Performance [ALZ16, ASE17, AS12, BR13, BD15, BZ15, BJ12, CVPS19, CA12a, CPS^{+}10, CK11, CCH11, CGT^{+}15, CCK^{+}16b, CJK19, CWY13, CTS13, CCD12, CCLH10, CDL^{+}17, FSPD17, FTP13, FG10, GZC^{+}17, GHL17, GCAG16, GRL^{+}14, GW16, GCS^{+}13, GBA^{+}18b, HMR^{+}17, HGCT13, HC13b, HLY14, HJF^{+}13, HGW^{+}17, Iko15, IS11, JAKD18, JC12, JZLD10, JHQL16, JJC14, KLK18, KM11, KIJ14, KBH^{+}10, KP15, KCRG15, KCL^{+}16, KSC^{+}14, KAK18, LLC^{+}16, LZ15, LK14, LS10a, LMJ14, LWZ18, LBWH11, LCY^{+}16, LZZ^{+}17b, LSG^{+}15, LY17, MWW14, ML18, MJWT16, MLS19, MKRM11, MKRM12, Nan16, ORM10, OPV^{+}17, PP11, PCLN15, PvdGG12, QWB^{+}13, QLH^{+}16, RVL^{+}14, RMB^{+}13, RSG^{+}19, SMP16, SMA19, SMB^{+}15, SMTK12, SSW12, SX12, SCSL12, SME^{+}17, Tho12, TCK^{+}18, TLGM17, VPS^{+}12, VED^{+}16, VAM10, WZLX12, WWY^{+}18, WDSP12, WFY^{+}17, XJFH15, XLC14, XLW14, XLJ16, YP12, YLML15]. Performance [YAG19, YMT13, YMTV14, YYP^{+}16, ZCZL16, ZWX12, ZCS16, ZL16, ZLSI17, ZZJ^{+}19, ZMW^{+}19, ZZ10, ZRL15, ZOD13, ZJXL11, khR^{+}18]. Performance-Driven [BR13]. Performance-Energy [BZ15]. Performance-Predictable [ALZ16]. Performance-Varying [ALZ16].

Performance/Power [Nan16].

Performance/Reliability [HMR^{+}17]. Performance/Reliability-Aware [HMR^{+}17]. Performing [KMP11]. Period [DZD^{+}16, LXL^{+}13, LDL^{+}17, RBR13]. Periodic [HGW^{+}17, LKBS16, LXJD15, LP13b, RBR13, WXS12]. Peripheral [PC10]. Peripheral-Processor [PC10]. Peripherals [BBP^{+}13]. Permanent [DKK16, MBSSA19, PB16, ZHM^{+}19]. Permutation [CJK15]. Permutations [ARH^{+}18]. Persistent [WFT^{+}19, XHLW19]. Perspective [KCW^{+}17, KSC^{+}14, WHBR16, WSXZ13]. Perturbed [PRM19, SST12]. Pervasive [KC13]. Pessimistic [HWL^{+}14, Tsa13, YLL16]. PETCAM [MS12]. Petri [HB11, CCK10, YLH13]. PHAETON [SBP16]. Phantasy [RZPX19]. Phase [DY14, JSC^{+}19, LYB15, LKK^{+}19, NBZP17, NL15a, PLM16, PP11, PTD^{+}12, QML^{+}15, WSXZ13, YTD^{+}18, ZZYZ14, NL19]. Phase-Based [PTD^{+}12]. Phase-Change [DY14, LYB15, PP11, QML^{+}15]. Phase-Encrypted [NBZP17]. Phases [CZ14]. Phishing [MAG^{+}17]. Phone [ZZX^{+}15]. Phones [TZL^{+}14]. Photonic [KC14, SRR^{+}16]. Phylogeny [MSPK12]. Physarum [LSZ^{+}15]. Physical [FBWMM13, HWSN15, LcZLG19, PRGBSAC19, QPG10, SLC^{+}15b, WTBT13, YLY^{+}15a]. Physically [CJSM17]. Piecewise [Pom15a, SDP11]. Piecewise-Functional [Pom15a]. Piecewise-Polynomial [SDP11]. PIM [LK18]. Pin [RC14]. Pin-Count [RC14]. Pinpointing [LLM^{+}15]. Pipeline [PRM16, PSND16, ROGHNB^{+}18]. Pipelineable [BDMLN16]. Pipelined [CLS10, CKKS14, GT19, HFG^{+}17, HZ11, KGP15, SDP^{+}15, WZLX12]. Pipelines [CP10, SF17, Sri10]. Pipelining [ZWL^{+}19]. Pixel [GKS14]. Pixel-Arithmetic [GKS14].

Place [LNCX18]. Placement [BDP15, CPL16, CYC11, FB13, GZB^{+}15, HGL^{+}15, KLK^{+}14, KP15, KLC^{+}16, LY11, LCLC19, Man16, MMH14, SWZG11, SZW^{+}16, WXLL13, WXLY15, WCLY16, YCLH16, YASS14, ZGG^{+}16]. Placement-Ideal [SWZG11]. Plagiarism [CJ12]. Planes [UVL^{+}13]. Planning [CMS10, CKN14, Cro14]. Plantlet [MSS17]. Plasticity [ZM17]. Platform [CZL^{+}17, RVL^{+}14, SPH13, VGF16, ZLN11]. Platforms [AKTB18, AS16, BMP^{+}10, DMA^{+}15, FMTK19, LCX^{+}16, YYP^{+}16, YAGB17, ZRL15]. Player [BSM^{+}14]. Players [KLLK11]. PLRU [KLLK11, LKLK13, WLY^{+}14]. Plus [BT16]. Plus-Minus [BT16]. PMC [HWL^{+}14, LKT13, Tsa13]. Podcast [Ano10g]. Point [AMG17, AK14, BLMM16, BGHL19, CHCK12, CK15, CI16, DJJ^{+}08, GH11, GNTS13, GABK11, HMS^{+}12, JPG10, JKMR11, JLMP11, JCK15, JMMP16, KGD16, LP17, Lee12, Lef17, LCW^{+}16, MKFM13, Nan19, PGvdG14, SBH11, SS12, VVMAZ12, VMHGN18, WF17, YTND12, ZMR^{+}13, ZCK19, dDLM11]. Pointers [RKN^{+}18]. Points [JLMP11, YTND12]. Policies [CXLL16, XLF15]. Policy [CWZC13, CFMA19, Cro14, LXDV17, LCHX11, RVH^{+}16, ZHW15]. Polymorphic [CXZ13, TXL11]. Polynomial [AH10, ACO12, BNP10, BLB^{+}19, CNH13, CHN14, Cil13, DWZ18, ERRMG15, Fan16, FP19, GAFN15, Gio12, HF15, HN11, JKMR11, LPW10, PCHS14, PCHS17, SDP11]. Polynomials [AK16, CMLRHS13, LCwW10, TIHM18]. Polysynchronous [NLRB17]. Popularity [ZHQX19]. Popularity-Aware [ZHQX19]. Population [MB16]. Population-Based [MB16]. Pore [MHRARG^{+}14]. Port [GCL^{+}13, LW15]. Portable [KLLK11]. Ports [Hie11]. Positive [WHC^{+}15b]. Post [BKH^{+}13, COLK18, DSR15, KN12, LLK18, NZ14, NL14, OHCK17, OCK17, PN16, SD18]. Post-CMOS [PN16]. Post-Quantum [LLK18, SD18]. Post-Silicon [BKH^{+}13, COLK18, NZ14, NL14, OHCK17, OCK17, KN12]. Post-Synthesis [DSR15]. Postsilicon [CCL^{+}13, PBV11]. Potential [CLW^{+}19, Cil11]. Power [AQPMS15, AS14, ARS16, BDDL18, BSS15, BRN^{+}15, BGM^{+}13, BGMR13, BJ12, CLS14, CSCW13, CRJZ16, CGJ^{+}10, CBTU14, DYW15, DMXY14, Dar15, DA12, DKLB15, DGC^{+}15, EKA17, FZL^{+}14, FHW18, GWMB13, GC16, GCAG16, GBD^{+}15, GRL^{+}14, GSF^{+}10, GDY15, HRM^{+}16, HMA^{+}10, IPS17, JAKD18, JP13, JHQL16, JLLH19, KPS^{+}17, KWC^{+}16, KH14, KBP13, KHZ17, LK15a, LZ15, LW17, LWZ18, LCL17, LN12, LGMP10, LHTG15, LWH^{+}16, LY17, LPL10, MWW14, MYHL16, MOYB12, MS12, MTBB10, MKRM11, Nan16, NC11, OPV^{+}17, PKC^{+}17, PvdGG12, QLH^{+}16, RSNK17, RSN^{+}18, RRK11, SP16, SCK10, SYD18, SKH16, SIVH16, SRK^{+}17, SLZX15, SPC^{+}18, SRHC12, TJX^{+}17, TLGM17, VED^{+}16, VALK19, WMW12, WJL^{+}12, WZM^{+}16, WP16, WSZ^{+}16, WTZ^{+}19, WMG18, XLS^{+}12, YAG19, YM11, YTM16, YKDEV19, ZR15a, ZOK^{+}19, ZZ10]. Power-Aware [CRJZ16, CGJ^{+}10, HRM^{+}16, YTM16]. Power-Constrained [TLGM17, WMW12, WJL^{+}12]. Power-Efficient [DKLB15, EKA17, FZL^{+}14, JAKD18, JP13, KH14, SRK^{+}17]. Power-Equalized [WMG18]. Power-Off [KWC^{+}16]. Power/Performance [GCAG16]. POWER5 [MBC^{+}13]. POWER6 [MBC^{+}13]. PowerCool [ASS^{+}18]. Powered [BDM^{+}19, CLZ19, MTBB10]. Powering [ASS^{+}18, HZL^{+}16, VB13]. Powers [Dum14, Wal19]. PowerTracer [LZW^{+}15]. Practical [DCM16, GDLL18, GC16, GPR^{+}19, HSA14, HKR^{+}18, KMM16, LMJ14, MBF18, PDJ^{+}19, SQJ^{+}15, WJX^{+}19, YFJ^{+}14]. Practice [LLL15]. PRAM [HK15b]. Pre [ARM15, TX16, TAM^{+}16]. Pre-Computation [ARM15, TX16]. Pre-Encoded [TAM^{+}16]. Precedence [GHK15, Li12b, LTVL15, TLZV11]. Precedence-Constrained [TLZV11]. Precise [CVH^{+}13, LK15a]. Precision [AMR18, BLB^{+}19, BGHL19, CCRL19, CZS^{+}19, FEM^{+}18, Joh17, JMMP16, KS10b, LP17, Lef17, LMB17, ZCK19, NVB16]. Precomputation [AS16]. Predicate [KHPP16, KC13, XKT^{+}15, YHML16, ZCR16]. Predicate/Transition [XKT^{+}15]. Predictable [ALZ16, ARGT14, DCL^{+}11, HCZW13, WA10, XLJ16]. Predicting [BD15, DSY^{+}15, SB10, ZZX^{+}15]. Prediction [AF14, AYC16, ASE17, BWV15, CGJ^{+}10, Fen14, JGG^{+}14, JWWZ16, JRJ^{+}18, LR16, LTL14, LDP10, LCT11, MKAY11, SB10, SIB13, XWL^{+}16b, ZCZL16, ZCS16]. Prediction-Based [AF14, BWV15]. Predictive [CNJ14, LSC10]. Predictors [MKAY11, MUMB11]. Preempt [SL14b]. Preemption [IGLM15]. Preemptive [Lee17]. Prefetch [LMNP11]. Prefetch-Aware [LMNP11]. Prefetched [RPM16]. Prefetcher [LKS^{+}14, Pan16, RLSK18]. Prefetching [DZ10, GWZ^{+}10, OYP^{+}18, OKY^{+}19, RZPX19, TLL^{+}13]. Prefix [CWZC13, CKKS14, LP12, LLLP14]. Prefix-Based [CKKS14]. Prescaling [WE12]. Preservation [ZDP^{+}15]. Preserving [CYHL14, GLTC16, LQD^{+}16, MB12b, QQW^{+}17, RVH^{+}16, SH12, SZDL14, WCW^{+}13, ZYC16, ZLX^{+}16, ZHW^{+}16, ZHW15]. Pressure [WZK^{+}19, XSR15]. Preventing [KKH^{+}14, LS10b, LDMQ16]. Prevention [MAG^{+}17, STK16]. Price [FLL14]. Pricing [JT15, MNGV16, ZV14]. Primal [LPL^{+}13]. Primal-Dual [LPL^{+}13].

Primary

[BR13, FHL^{+}18, MJWT16, YTD^{+}18]. Primary-Backup [BR13]. Prime [Dum14, XMH13]. Prime-Length [XMH13]. Primitives [AAR19]. Principal [FEM^{+}18]. Prioritized [LH12b]. Priority [BM13a, BDB18, Lee17, LYS10, MBD^{+}17, NRG15, ZZ19]. Privacy [CYHL14, HBCC13, LDMQ16, LQD^{+}16, QQW^{+}17, RVH^{+}16, SKZS13, SZDL14, WCW^{+}13, ZDP^{+}15, ZLY15, ZYC16, ZLX^{+}16, ZHW^{+}16, ZHW15]. Privacy-Preserving [CYHL14, LQD^{+}16, QQW^{+}17, RVH^{+}16, SZDL14, WCW^{+}13, ZHW^{+}16, ZHW15]. Private [QZL^{+}16, WCL^{+}18]. Privatization [AH13]. Proactive [LYY16, LZYL13, WZY16, WJM15]. Proactively [CHK10]. Probabilistic [LJ18, LHL13b, MHHS17, MHH^{+}17, NC11, WNCH17, ZCR16]. Probabilistically [KAQC14, WHL17]. Probabilistically-Atomic [WHL17]. Probabilities [LB15a]. Probability [NCD^{+}17, WF14]. Probable [XTF^{+}12]. Probe [ZC13]. Probe-Based [ZC13]. Probing [ZC13]. Problem [Amm14, CPS^{+}10, DALD18, DVUS14, Fuj11, LHPH15, LZ14, LSZ^{+}15, WTY^{+}14]. Problems [DDN14, GO10, LYT^{+}16, LSX13, WHL^{+}12, XLL15]. Procedure [BTBB14, LOC^{+}16, XXBL17]. Process [DSG^{+}19, GV15, HK16, KSC^{+}14, KKC15b, LSA18, MLS19, PR10, ZDYZ13, ZDYZ14]. Processes [Cao12, NZLK14]. Processing [BBB^{+}17, CK15, DHW^{+}19, DKW15, FGS^{+}15, GZC^{+}17, GZG^{+}16, GIKR19, GAC14, KLC18, LOX^{+}13, LVJ16, MW10, SRCbL^{+}15, SOM^{+}13, SSW12, SPH13, WZLX12, WGR^{+}14, WAK^{+}17, WZW^{+}19, XLS^{+}12, XYHD17, ZCZ^{+}19, ZLJ^{+}17, dOPSR16]. Processor [BJ12, Dar15, Gor14, GRL^{+}14, HCG^{+}16, IDG^{+}17, JWH^{+}15, JAD^{+}18, KAH^{+}15, LvH12, Li12b, LB13, LOC^{+}16, MBD^{+}17, MFG14, MIS^{+}14, PBV11, PCZB11, PC10, PPP13, RJV^{+}18, RSG^{+}19, SDMM12, TBC^{+}17, VED^{+}16, YMG15, YYC12, ZDYZ13, ZDYZ14]. Processors [ARS16, BLKM^{+}18, CLS14, CA12b, CCLH10, DZLP14, FHW18, GJ14, GSL10, GBA^{+}18b, HTA17, HV12, HV14a, HBR11, IHR^{+}16, JLC10, JJC14, KKC17, LK15a, LKH16, LSA^{+}17, LMC^{+}12, MMAC19, MBC^{+}13, NEE18, OPAGS14, RCM^{+}16, RURM18, RTRM19, RCN11, RTL^{+}18, RRK11, SZZ^{+}19, TCHL18, USP^{+}13, YG10, YRG13, ZR15b, ZZ10]. Procurement [PR14]. Produced [Jes15]. Producer [KSEG15]. Producer-Consumer [KSEG15]. Product [CLL^{+}14, CLC^{+}16, HF15, HMNN12, HN13, PCHS16, PCHS18, Red19, RM15b]. Production [CKN14]. Products [BCTV19, CNH13]. Profile [AVS^{+}14, SV18]. Profiled [Bar16]. Profiling [AKTB18, Fin10, KLKL13, LVMS18]. Profit [MLOL15]. Profitability [EE10]. Program [CZ14, KLLK11, LYH11, MUMB11, PNKI13, RKT19, SB10]. Program-Level [PNKI13]. Programmable [BEHL^{+}19, WLV^{+}14, ZJH^{+}14]. Programming [BFR^{+}15, CJSM17, HTA17, LvH12, LJVJ18, WHL^{+}12, WRW16]. Programs [BCK^{+}16, Fin10, GPRS17, GCR^{+}19, WLY^{+}14]. Progress [FSPD17, OYP^{+}18, TLGM17]. Progress-Aware [FSPD17]. ProgressFace [LYCT10]. Progressive [FBR^{+}12]. Projected [BCTV19]. Projections [SST12]. Promise [LT15]. Promoting [LYL^{+}19]. Proof [HCD^{+}16, SDZ15, XXBL17, ZGWC15]. Proofs [CL10]. Propagating [LS10b]. Propagation [ACGP13, FNS16, LLW^{+}18, WHC^{+}15b]. Properties [CL10, HIJ^{+}19, LHL13a, WGZ^{+}15]. Property [CM11, HIJ^{+}19, MLW12, ST18a, TIHM18, YWM19, ZWYY15]. Property-Based [MLW12]. Proportional [NY19]. Propose [BFMT16]. Protect [CSS13, FRB^{+}18]. Protected [MAD14]. Protection [CMM15, DRM16, DCV^{+}12, HCG^{+}16, KSN^{+}15, LDMQ16, LLS^{+}16, MAD14, NDG^{+}17, RURM18, RTRM19, Red11, SKZS13, SMRM17, SEY14, SP12, SWWC11, YCL^{+}12, YW12, ZLY15]. Protein [MGW14]. Protocol [AVG^{+}15, BGRH15, CSJ^{+}11, FHLOJRH18, FBWMM13, HL10a, HWX15, JZLD10, LCCJ13, LZYL13, LSL15, LXZ^{+}15, RQ14, SMB^{+}15, SDZ15, TC14]. Protocols [AD12, CKM15, CMRH17, CWZ13, DVUS14, EFGT18, KKP^{+}16, LLZ^{+}17, LYCT10, PFGB14, VYEB17, VYEB18, YRG13, YRT^{+}16]. Prototype [Bar16, CS11b]. Prototyping [CCAM14, JRP^{+}14]. Provably [DJJ^{+}08, Lee12, PSM17, XJWW13]. Providers [FLL14]. Providing [GPN11, YASS14, ZCY^{+}16]. Proving [HTA10]. Provisioning [CRJZ16, HGL^{+}15, JPLP13, KL16, LY17, PAC^{+}12, SXCL14, XLF15, XLJ16, ZRS^{+}16, ZT15]. Proxies [GJ14]. Proximity [ZDP^{+}15]. Proximity-Aware [ZDP^{+}15]. Proxy [HHM11, XJW^{+}16]. Prune [DYJ19]. Pruning [CZP^{+}16, Ste14]. PSBS [DCM16]. Pseudo [FD16, KCW^{+}17]. Pseudo-Boolean [FD16]. Pseudo-Differential [KCW^{+}17]. Pseudorandom [MG16]. PSRAM [CSCW13]. Pub [BBPQ15]. Pub/Sub [BBPQ15]. Public [JCM16, LCwW10, LRY^{+}15, MRW^{+}15, SWM^{+}10, WCW^{+}13, XJWW13]. Public-Key [LCwW10, SWM^{+}10, XJWW13]. Publication [Ano10f]. Publish [BS15]. Publish/Subscribe [BS15]. Publishing [Ano13e]. Pubsub [BS15]. PUF [GLH^{+}19, SMCN18]. PUFs [USH19].

Purpose [HTA17, SJVR19]. Push [LP13b]. Push-Based [LP13b].

Q [DYHX16, SCJ^{+}16a, WGW^{+}15]. Q-DRAM [SCJ^{+}16a]. Q-Learning [DYHX16]. QBF [MVB10]. QC [HC17]. QC-MDPC [HC17]. QCA [LLOS13]. QoE [THM^{+}14]. QoE-Based [THM^{+}14]. QoS [AK15, CP10, GZB^{+}15, HGL^{+}15, HLC^{+}19, KS10a, KSS12, KP15, LY11, SMG14, ZWC^{+}18, ZQQ11]. QoS-Aware [KSS12, LY11, ZQQ11]. QoS-Based [AK15]. QoS-Less [HLC^{+}19]. Quadratic [Bin15, SEY14]. Quadruple [FGS^{+}13]. Quadruple-Based [FGS^{+}13]. Qualification [PFGB14]. Quality [BKV12, Jes15, MLOL15, PAC^{+}12, RSJR17, SC11, YHV13]. Quality-Driven [YHV13]. Quantifying [LY17, YSLL16]. Quantitative [ZLH^{+}15]. Quantum [GM15, HHKW12, KMM16, KKS14, LLK18, LWK11, MSK15, MCT19, RO11, SO10, SD18, WCL^{+}18]. Quantum-Dot [KKS14]. Quantum-Oblivious-Key-Transfer-Based [WCL^{+}18]. Quasi [LC16b, PDXZ13]. Quasi-EZ-NAND [PDXZ13]. Quasi-Random [LC16b]. Quaternions [PP16]. Quatrain [RWZZ14]. Qubit [KMM16]. Queries [SX12, WXS12, ZCL^{+}16]. Query [CLR13, HXVF12, OOD^{+}17, WCL^{+}18]. Quest [RM15a]. Queue [PPND17]. Queued [ACGP13, AHK10, BGMR13, ZWLS15]. Queues [TB15]. Queuing [FHW18, SKEB18]. Queuing-Based [FHW18]. Queuing-Theory [SKEB18]. Quick [SCJ^{+}16a, Tsa13, ZZJ^{+}19]. Quick-Access [SCJ^{+}16a]. Quick-and-Dirty [ZZJ^{+}19]. Quorum [LRC10]. Quorum-Based [LRC10]. Quotient [Rus13].

R2 [DLC^{+}13]. R3TOS [IBH^{+}13]. Race [DJN17, WhCCC12]. Races [RBK^{+}12]. Racetrack [MWZ^{+}17, SBW^{+}16]. Radar [WCLY16]. Radiation [CPRH16, dOPSR16]. Radiation-Induced [dOPSR16]. Radio [BBVL14, LWY15, XWL^{+}16a, YCCWC15]. Radius [Joh17]. Radix [ARM15, AS10, ABA07, DJA11, EJ15, GIW18, Jha13, JHQL16, Kor15, LQW^{+}17, MLH12, RMB^{+}12, TAM^{+}16, VAB14, WE12]. Radix- [EJ15, TAM^{+}16, VAB14]. Radix-16 [WE12]. Radix-2 [MLH12]. Radix-4 [LQW^{+}17]. Radix-8 [ARM15, JHQL16]. RAHMM [JYL^{+}17]. RAID [FSL^{+}17, IS11, KLC^{+}16, LLL16, ROGHNB^{+}18, WJF^{+}11, ZZS10, ZZL14, ZLWZ15]. RAID-5 [ZZS10, ZZL14]. RAID-6 [FSL^{+}17, ZLWZ15]. RAID-Structured [WJF^{+}11]. Rail [GSF^{+}10, NI11]. Railway [HDYS16]. RAM [AYC16, CKD^{+}17, FKMK16, JW16, KA19, LHL^{+}15a, PP11, WW16, YLGD19]. RAMS [LK16b, AAR19]. Random [BCK^{+}16, CB15, EF12, Jes15, LLD^{+}16, LC16b, SST12, SMK^{+}16, SN16, TM18, YCW11, ZOD13]. Randomization [DSY^{+}15]. Randomized [CATB19, GDLL18, GIW18, LC16b, RL13, XCW^{+}10]. Range [KN11a, SX12]. Rank [LPL^{+}13, LK16b, LWH^{+}16, SJC^{+}17b, FZL^{+}14]. Rank-Aware [LK16b, LWH^{+}16]. Rank-Level [SJC^{+}17b]. Ranked [ZLX^{+}16]. Ranking [ABSK15]. Raphson [Dum14, Wal19]. Rapid [HGML11, JRP^{+}14]. Raptor [MNK11]. Rate [CFMA19, GDY15, LOH17, LB15b, MBGS10, PP14, SV18, ZCC^{+}14, dRV12]. Rate-Distortion [dRV12]. Rate-Optimal [MBGS10]. Rate-Selective [LOH17]. Rateless [YW12]. Rating [JWWZ16]. Ratio [LHPH15]. Ray [XHZC16]. RB [SQJ^{+}15]. RB-Explorer [SQJ^{+}15]. RC [HWX15, LXW^{+}19]. RC-MAC [HWX15]. RC-NVM [LXW^{+}19]. RC4 [GCS^{+}13]. RD [SV18]. RDIS [MMC15]. RDP [ZLWZ15].

Re [XJW^{+}16]. Re-Encryption [XJW^{+}16]. Reachability [HB11, XXBL17]. Reachable [Pom15a]. Reactive [LvH12, RBRL15]. Read [BEHL^{+}19, CLW^{+}19, GC16, KCL19, LKK^{+}19, LSG^{+}18, LCY^{+}16, HCY18]. Read-Modify-Write [LKK^{+}19]. Reads [FSL^{+}17, RLSK18, ZLLX15]. Real [AF14, ADP^{+}15, ABEP16, AEP18, AE11, AC11, BBD^{+}12, BBP^{+}13, Bin15, BBB16, BPC12, BGRH15, CWF14, CBB19, CW10, Cha14, CYCC11, CGL^{+}18, CQ14, CLR13, DZD^{+}16, DA12, EE17, FM16, GCAG16, HWZ^{+}12, HV12, HCH15, HV16, HHLK12, HXL11, HV14b, HGW^{+}17, IBH^{+}13, KS10a, KSS12, KM11, KGP15, KMC17, KLLK11, KCE^{+}18, KAQC14, KDEV19, KT12, LYH11, LHC^{+}14, LKLK13, LSSE15, LXDV17, LC16a, LLW^{+}11, LGS^{+}18, MUMB11, MBSSA19, MFG14, MKM14, MW13, NRG15, NZLK14, OPZ15, PTD^{+}12, PC10, PMH^{+}14, RHC^{+}14, RF14, SYD18, TB15, TH11, TFCY16, TCHL18, THGT13, TKT16, WXS12, WBZ^{+}15, WLY^{+}14, WA10, WLZ10, XWLX17, YPB^{+}16, YYW^{+}16, YHV13, YTM16, YAGB17, ZLBB19, ZD13, ZLJ^{+}17, ZCW18, ZZ19, ZYY18, ZHM^{+}19, ZQQ11, ZCYX15]. Real-Numbers [HV16]. Real-Power [SYD18]. Real-Time [ABEP16, AEP18, AE11, BBD^{+}12, BBP^{+}13, Bin15, BBB16, BPC12, BGRH15, CWF14, CBB19, CW10, Cha14, CYCC11, CGL^{+}18, CQ14, CLR13, DZD^{+}16, DA12, EE17, FM16, GCAG16, HWZ^{+}12, HV12, HCH15, HHLK12, HXL11, HV14b, HGW^{+}17, IBH^{+}13, KS10a, KSS12, KM11, KGP15, KMC17, KLLK11, KCE^{+}18, KAQC14, KDEV19, KT12, LYH11, LHC^{+}14, LKLK13, LSSE15, LXDV17, LC16a, LLW^{+}11, LGS^{+}18, MUMB11, MBSSA19, MFG14, MKM14, MW13, NRG15, NZLK14, OPZ15, PTD^{+}12, PC10, PMH^{+}14, RHC^{+}14, RF14, TB15, TH11, TFCY16, TCHL18, THGT13, TKT16, WXS12, WBZ^{+}15, WLY^{+}14, WA10, WLZ10, XWLX17, YPB^{+}16, YYW^{+}16, YHV13, YTM16, YAGB17, ZLBB19, ZD13, ZZ19, ZYY18, ZHM^{+}19, ZQQ11, ZCYX15]. Real-World [ZLJ^{+}17, ZCW18]. Real/Complex [AC11]. Realignment [VCG^{+}12]. Realizable [LT15]. Realization [LK10, XMH13]. Realizing [WKB16, ZLW^{+}17]. Reallocation [Tse12]. Rearranging [SJS10]. Reasoning [DSB13]. Reasons [Ano10b]. Receiver [HWX15]. Receiver-Centric [HWX15]. Receiving [Ano13f, ST12]. Rechargeable [WLYY16]. Recoding [RS10, ZDP^{+}15]. Recognition [AWFV13, LKLT12, QWB^{+}13, TBC^{+}17, YAG19]. Recombination [CHN14, HMNN12, PCHS17]. Recompilation [GDJZ18]. Reconfigurable [AD16, BQP^{+}16, BBI^{+}13, CZP^{+}18, CCLH10, DPO17, DKK16, EKA17, GKB^{+}10, GP14, GLP15, GFAM11, HNB^{+}12, IBH^{+}13, KMLH11, KGC14, LDB^{+}17, LYJ^{+}18, MLCH10, NVB16, PPP13, RWC18, RSU17, SOM^{+}13, SPH13, TLB^{+}17, TBC^{+}17, UVL^{+}13, WTBT13, WZCG16, ZMR^{+}13, ZBK^{+}17, ZYY18]. Reconfiguration [MKLW14, PB16, YZHX12]. Reconfiguring [JWH^{+}15, ZNL18]. Reconstructing [GDC^{+}16]. Reconstruction [HQLX15, LS10b, MG16, MH19, SST12]. Recording [WFY^{+}17]. Recoverability [YCK16]. Recoverable [DDNP11]. Recovery [ASTU10, AD16, DKK16, GSG^{+}15, HGML11, JSC^{+}17, KWC^{+}16, LLL15, LSA^{+}17, LLD19, TLB^{+}17, XHZ14, XLX^{+}14, YXZZ14, ZXX^{+}14]. ReCREW [DKH^{+}13]. Rectilinear [WTY^{+}14]. Recurrent [XWL^{+}16b]. Recursive [KP13, LMZQ17, MHHS17, Red14]. Recursively [LS10c]. Redeeming [YSLL16]. Redio [WZW^{+}19]. Redirection [HQLX15]. Redistribution [ZZS10]. ReDO [SVD18]. Reduce [COLK18, DZ10, GPRS17, KGGJ14, KS10b, LXK12, LZW^{+}15, OHCK17, OCK17, QPG10, WBGI19, WZK^{+}19]. Reduced [CCRL19, HLC^{+}19, LK15a, Li12b, Pom12a, PSL17, WS10]. Reducing [BS16, CTD^{+}16, CLZ19, KLK17, KCL19, LAAM11, LKH16, LCLC19, RS13, SPC^{+}18, WFT^{+}19, WZ14, WZW^{+}19, YCW11, Yan14, YLP15, ZLWZ15, ZMW15]. Reduction [AKKH12, CCC15, CCC^{+}18, CFMA19, GBA18a, JJK^{+}11, JT15, KVV10, LK15b, NL19, OKC13, SMK^{+}16, TLL12, WS10, XJFT16]. Redundancy [CAbZM18, CCRL19, FHL^{+}18, GY15a, GY15b, HBR11, KwPK^{+}15, KS12, LW17, SSW12, VCSG^{+}19, ZHQX19]. Redundant [AO12b, Bra10, CCK^{+}16b, CvdBC18, CLC^{+}16, EJ15, GJ15, HK13a, HVZ13, JPG10, LCY^{+}19, TAM^{+}16, VVMAZ12, VAB14]. Redundant-Digit [JPG10]. Reed [PROM15, RMB^{+}12, TW10]. Reference [ZZ17]. Refine [LVMS18]. Refinement [HSH^{+}10, Sri10, ZYY10]. Refresh [BCM14, BCC^{+}16, CLW^{+}19, DSG^{+}19, GC16, JSC^{+}17, LHL^{+}15a, SSJ^{+}18, SCJ^{+}16b, SPC^{+}18, WBGI19]. Refresh-Aware [JSC^{+}17]. Refreshing [FHW18]. Refurbishment [AD13]. Regenerating [LLL15]. Regenerating-Coding-Based [LLL15]. Regime [SMB^{+}15]. Region [CCK10, HWSX17, HCSW15, MSC12, ZD13]. Region-Based [CCK10]. Region-Level [HWSX17]. Regional [CHC^{+}15]. Regions [DNSS11]. Register [CCV^{+}11, JLKR19, MWZ^{+}17, NCD^{+}17, QPG10, RURM18, RRK11, VCSG^{+}19, XLL^{+}18]. Registers [QPG10]. Regression [DP13, LDP10]. Regular [ARM15, CMB13, Cha10a, LKT13, LXZH16, OWP16, YP12]. Regularity [LYL^{+}19]. Regulated [YYW^{+}16]. Regulation [LSG^{+}18]. Reinforcement [FCB^{+}19, WP16, ZM17]. Reinforcing [YCK16]. Rejuvenation [BDL^{+}13]. Rejuvenations [HGW^{+}17]. Related [Cil11]. Relating [ZGW14]. Relational [XTW15]. Relations [BCTV15, BCTV19, PR10]. Relationship [HL10b, SCJ^{+}16b]. Relationships [MB12b]. Relaunching [GCF^{+}16]. Relaxation [CCK^{+}16a, LZZ17a]. Relaxed [LYL^{+}19]. Relay [CWY13, LY11, LB15a, MMH14, ZY12]. Releasing [CLW^{+}19]. Reliability [BMS11, BM13b, Cao12, CK11, CHH^{+}13, CCK^{+}16b, CJA^{+}16, CvdBC18, FHL^{+}18, HCL^{+}14, HCY18, HTA10, HL10b, ISC15, Ibr16, KLC^{+}16, LXJD15, LCH^{+}15, LLL16, LYY16, LHL13b, MHK15, MMAC19, NEE18, PCZB11, SMCN18, SKEB16, SDP^{+}12, VSC^{+}19, WWM16, XH16, YyHL11, YWXZ12, YL14, YMTV14, ZYL15, ZWYY15, ZWC13]. Reliability-Aware [BMS11, HMR^{+}17, ZYL15]. Reliability-Driven [BM13b, PCZB11]. Reliable [ACW^{+}11, AFH^{+}10, AS14, ARS16, BBI^{+}13, BS10, CHL17, CAGM14, DKH^{+}13, HLY14, IBH^{+}13, KDEV19, LXL^{+}14, LKLT12, VBR^{+}13, WKB16, XL16, YTD^{+}17, ZL15]. Relinquishment [ST17]. Relocatable [DLC^{+}13]. Relocation [KK10]. Remainder [Fan16, PB11]. Remapping [JDA^{+}16, KCKL19]. Remote [HPR16, JGG^{+}14, KCRG15, SRCbL^{+}15]. Removal [LHCL13]. Removing [RXC^{+}15, WLQS13]. Renaming [JLKR19]. Renewable [CFW14, RLX15]. Reordered [NWA12]. Reordering [KKH^{+}14]. Reorganization [LBWH11]. Reorganizing [AGFM11]. Repair [LLW^{+}17, PN16]. Repairing [RSU17, TW10]. Repeatable [DN11]. Repetitive [TLP17]. Replacement [CXLL16, CFMA19, KS14, LBN14, OGH^{+}14]. Replacing [YMG15]. Replay [RTL^{+}18, ZCZL16]. Replica [AT16, LYY16, LRY^{+}15]. Replicas [LK18]. Replication [AD14, BR13, CS15, GV15, HS18, LYY16, LLX^{+}17, SWZG11, SL13, SLC15a, Tse12].

Reporting [LZYL13]. Repositioning [FKMK16]. Represent [LCA10]. Representation [AO12b, BNP10, BFMT16, Bra10, HN11, LFH^{+}16, LWK11]. Representations [DJA11, KMP11, MRD19]. Representative [ZCZL16]. Reproducibility [RT14]. Reproducible [BGHL19, DN15]. Reprogramming [DLC^{+}13, LLZ^{+}17]. RepuCoin [YKDEV19]. Reputation [CWZ11, LSS13, RDEN10, YKDEV19]. Request [LR10]. Requester [CWCS15]. Requester-Based [CWCS15]. Requests [LZZV16, LZW^{+}15]. Required [KS10b]. Requirement [HV13, LYS14]. Requirement-Aware [HV13]. Requirements [HHKW12, MAHD18, SAR^{+}11, SZW^{+}16]. Resemblance [XJFT16]. Reservation [ZGWC15]. Residue [TC16]. Resilience [CCAM14, CATB19, HV13, HLWV17, MG11a, PCZB11, SMAR^{+}19, VSC^{+}19, ZBK^{+}17]. Resilient [DKLB15, DZLP14, HK16, KDEV19, SKEB16, SVD18, YEY^{+}16]. Resistance [STE17]. Resistant [GV14]. Resistive [MMC15]. Resistivity [MM17]. Resolution [PMH^{+}14, ZZ17]. Resource [BSM^{+}14, BGRH15, CZP^{+}16, CRJZ16, CZL^{+}17, COLK18, DCK16, Fin10, FB13, GO10, HMR^{+}17, HGW^{+}17, HLA^{+}17, KGP15, KPS^{+}17, KFB^{+}15, KL16, LC16a, LZA^{+}16, LGF^{+}15, MNGV16, NRG15, PB16, PAP13, PR14, SLL15, SPTC15, SXCL14, WHYS16, YHC19, YLH10]. Resource-Efficient [DCK16]. Resource-Oriented [YHC19]. Resource-Saving [KL16]. Resources [ALZ16, BBB16, CFW14, CP10, FJA^{+}17, PLP^{+}13, RF14, SH12, WL13, YHC19]. Respect [CATB19]. Response [BDB18, MBD^{+}17, NRG15, WZ14, YLH10, ZHW^{+}16]. Response-Time [BDB18]. Responses [KN11b]. Responsive [BGM^{+}13]. Responsiveness [YCW^{+}19].

Restart [LL11, ZYL15]. Restoration [HRK17, YXZZ14]. Restore [CKD^{+}17]. Restoring [CHLL16, LHL13a, SCJ^{+}16a, YLA10]. Restricted [OWP16]. Restriction [AK19]. Result [LK15a]. Retention [DSG^{+}19, SSJ^{+}18, WBGI19, JYL^{+}17]. Retention-Aware [JYL^{+}17]. Retentive [RRK11]. Rethinking [LLD^{+}16, ZZL14]. Retrieval [BCMJ10, CWTT13, LWF13]. Retrieval-Guaranteed [CWTT13]. Reusability [MWY^{+}16]. Reusability-Aware [MWY^{+}16]. Reuse [JC11, KSN^{+}15, OGPK14]. Reused [COLK18]. Reversible [DSR15, GM12]. Reviewers [Ano10a, Ano11b, Ano12b, Ano13b, Ano14b, Ano15b, Ano17b, Ano18b, Ano19b, Ano16b]. Reviewing [Kor15]. Revisited [CKM15, CB15, MBGS10, SKC^{+}14]. Revisiting [MMTM15]. ReviveNet [YyHL11]. Revocable [SZDL14]. Revocation [JCM16, LLC^{+}15]. Reward [MFG16]. Reward-Based [MFG16]. Rewarding [SPTC15]. Rewrite [WZ14]. RFID [LLM^{+}15, LXZ^{+}15, MWWT13, QZL^{+}16, SKZS13, SSKL16, SDZ15, YXWL16, ZCC^{+}14]. RFID-Enabled [QZL^{+}16]. Right [VIDH19]. Ring [BPC12, HGML11, KKY^{+}16]. RLWE [GDLL18]. RLWE-Based [GDLL18]. RNS [BT16, CATB19, GLP^{+}12, GL19, Hia17, Sou15, YFCV14]. Road [JGHD11]. ROBDD [LCA10]. Robin [FJA^{+}17]. Robot [LBS15]. Robotic [SAR^{+}11]. Robots [LBS15]. Robust [BDDL18, BDBB18, DDNP11, LMB13, MVB10, MC11, NI11, ORM10, SB16, WZLX12, ZWX12]. Role [XKT^{+}15]. Role-Based [XKT^{+}15]. ROM [AAR19]. ROM-Embedded [AAR19]. ROMs [LLHC15]. Roofline [IPS17]. Root [ARH14, AH10, FBE^{+}18, JLLH19, KCK16, LP13b, LN12, Rus13, VB13, WE12].

Rooted [SG13]. Roots [JKMR11]. ROSE [CC11]. Rotated [YHT^{+}16]. Rotation [MM17]. Rotations [RS10]. Round [FJA^{+}17, HV16, KMP11, MEBS17]. Round-Off [MEBS17]. Round-Robin [FJA^{+}17]. Round-to-Nearest [HV16, KMP11]. Rounded [BHR17, DRC14, KLLM12, Lef17]. Rounding [CHCK12, JPG10, KS10b, PB11, TGNSC11, VMHGN18]. Roundoff [MRD19]. Route [YLY^{+}15a]. Routed [KH14]. Router [BM13a, HHY11, HH17, KAH18a, LS10c, MKAY11, OMFH14, PSND16, YLGD19]. Router-Tables [HHY11, HH17]. Routers [JRC14, YMK^{+}17]. Routing [AVS^{+}16, BKV12, CFR^{+}14, CHC^{+}15, CCE^{+}18, CBVL16, DVUS14, DSPB13, EDL^{+}14, EG11, FMTK19, FG10, FS10, GDC^{+}16, HCSW15, HKWC14, KCS^{+}13, KCS14, LR13, LS13, LSHC15, LYCT10, LWY15, LZS^{+}13, MJW^{+}14, MFT^{+}17, MWY^{+}16, MMB14, RL13, RO11, RKZ16, SKEB16, SKEB18, SZS14, SKA10, TLP17, TLP18, TH11, VBR^{+}13, WWM16, WS15, WW14, WLS18, XWH14, ZGB^{+}15, ZHM14, ZGY14, ZBW17]. Row [HGCT13, LXW^{+}19, SCJ^{+}16b, SCJ^{+}16a]. Row-Access [SCJ^{+}16b]. Row-Activation [SCJ^{+}16a]. Rowhammer [KCKL19]. RPR [CCRL19]. RRAM [CSW^{+}15, LW17]. RRNS [TC16]. RRR [HGML11]. RSA [MLW12]. RSS [LTL14]. RT [KLLK11, KDEV19]. RT-ByzCast [KDEV19]. RT-PLRU [KLLK11]. RTL [ABSK15, BFP11]. Ruining [KKH^{+}14]. Rule [HGL^{+}15, LDP10, LP13b, MMP13, RCRK13, SWZG15, YCZ10]. Rules [DZLP14, LJG^{+}19, SBH11]. Ruling [LKLT12]. Run [DKK16, DJN17, LBSK17]. Run-Time [DKK16]. Running [JSE14, VYEB18]. Runtime [BBI^{+}13, DGC^{+}15, KK10, KP15, OYP^{+}18, RBK^{+}12, SIB13, SPH13, WZM^{+}16, WJY^{+}17, ZBK^{+}17].

S [MM17, MKRM11, ST18a]. S-Box [MKRM11]. S-Boxes [MM17, ST18a]. S-Orchestrated [SNM16]. Safe [FJA^{+}17, PKC^{+}17, SHGW15]. Safety [ARGT14, MTFK19, SAR^{+}11, ST11a, YMT13]. Safety-Critical [ARGT14]. SALSA [XWL10]. Same [CS11a]. Sample [DSY^{+}15]. Sampled [GCR^{+}19, LVJ16]. Samplers [HKR^{+}18]. Samples [CVH^{+}13]. Sampling [AMR18, EF12, GPR^{+}19, KRR^{+}18, ZWW19, dAJM14]. Sandboxed [GD17]. Sanitizer [GBGI18]. SAT [JLMH10, SBP16]. SAT-Based [SBP16]. Satellite [LZS^{+}13, SPH13]. Satisfiability [LCW^{+}15, YHH^{+}12]. Saturated [SMB^{+}15]. Save [DHM16]. Saving [KL16, LK16b, LTL14, LHH14a, WLJ^{+}16, XDZ11]. Savings [GBD^{+}15]. SCA [PDJ^{+}19]. SCADFA [PDJ^{+}19]. Scalability [CCH11, ZZL14]. Scalable [ABSK15, AKL14, AD13, CLW^{+}15, CW16, CWCS15, DAS14, FEM^{+}18, GCD^{+}11, GEN^{+}17, JC12, JLMH10, KGV16, KAK18, LP12, LMJ14, LT15, LYS14, LCHX11, LLKA19, LJVJ18, MBS^{+}12, PP11, RBRL15, RSN^{+}18, SSGB19, SIVH16, SL10, ST18b, TC14, UVL^{+}13, XP10, XYF^{+}15, YMK^{+}17, ZDP^{+}15, ZLN11, Zot10]. Scalar [ARM15, DS14, GIW18, LJL13, MH15, NR15]. Scale [AISA16, BMT14, CQW^{+}15, FFCB14, GDC^{+}16, GV15, GY16, JKY10, LP13a, LS10a, LDB^{+}17, LT15, LXK12, LQD^{+}16, MYHL16, MCXZ18, SRR^{+}16, WS15, WJM15, ZCZL16, ZWC^{+}18]. Scale-Out [MYHL16]. Scale-Up [MYHL16]. Scalers [Hia17, Sou15]. Scaling [CLX14, DA12, FCB^{+}19, GRL^{+}14, JSH^{+}17, JCK15, LKYC12, LHH14a, LHTG15, MHK15, NY15, NYHB16, NY19, OKC13, PdG13, WJL^{+}12, YTM16, ZZS10, ZLWZ15]. Scan [CZP^{+}18, CCC15, KCY18, Pom12d]. Scan-Based [Pom12d]. Scanning [LLL11, PWW^{+}11]. Scenario [XLW14]. Scenarios [MTFK19]. Schedulability [BBB16, HGW^{+}17, LHC^{+}14, LSSE15, Lee17, MBB^{+}17, PP14, RHC^{+}14, WLZ10, YYW^{+}16]. Scheduled [BGRH15]. Scheduler [CZ14, FSPD17, KMJ^{+}11, TB15]. Schedules [AO12a, LSXP14]. Scheduling [ACGP13, BBD^{+}12, BMP^{+}10, BDB18, BTW13, BPC12, CFL^{+}18, CAbZM18, CBB19, CNJ14, CZP^{+}16, CGL^{+}18, CZL^{+}17, CQ14, CLR13, DCM16, FSPD16, Fuj11, GKD^{+}17, GGSPM18, GHK15, GCAG16, GY13, GLTC16, HRM^{+}16, HZL^{+}16, HKWC14, HXL11, HV13, HV14b, HZX^{+}14, HLWV17, IGLM15, IHR^{+}16, JJK^{+}11, JR17, KTAvdS16, KH18, KCE^{+}18, LRC10, LHC^{+}14, LK16b, Lee17, LRP^{+}18, Li12b, LTVL15, LC16a, LDL^{+}17, LP13b, LGS^{+}18, LMB13, LWF13, MFG16, MBD^{+}17, MAHD18, MBGS10, NEE18, OKY^{+}19, PM14, RHC^{+}14, RWC18, RF14, RLX15, RC14, SZG^{+}18, SL14b, TLZV11, TYY^{+}16, TCK^{+}18, TCHL18, VC10, WXS12, WBZ^{+}15, WZK^{+}19, XCF16, XCW^{+}10, Yan14, YLGD19, YPB^{+}16, YHV13, YTM16, ZGG^{+}16, ZWLS15, ZR15b, ZWC^{+}18, ZQZ^{+}19, ZZ19, ZWL^{+}19, ZQQ11, ZCYX15, ZLYS15, ZMRQ11]. Scheme [ARM15, AKJ^{+}13, BS14, BS16, CMLS15, CCW^{+}10, CSW^{+}15, CLM^{+}19, CWTT13, COLK18, CWCS15, GYC^{+}16, GTRMG18, GWM^{+}17, HS18, HCL15, HK15b, HHCH11, HLJ14, HQLX15, JPG10, KLLK11, KL16, LTL14, LCL15, LKLT12, LWY15, LLW^{+}17, MLOL15, MRL^{+}18, RVH^{+}16, RSN^{+}18, SSKL16, SRCbL^{+}15, SWF^{+}19, SZS14, UMN18, WLY^{+}14, WNKL16, XJFT16, XJWW13, YLY^{+}15a, YTD^{+}17, ZPM^{+}15, ZFHC19, ZHQX19, CTS13]. Schemes [CMLRHS13, CJA^{+}16, CKD^{+}17, CHL17, EFGT18, HSM14, MLCH10, MBF18, MKRM10, XTF^{+}12, YMTV14]. Science [MOS14, ST11a]. Scientific [KN11a, SDMM12, YLY15b]. SCM [JYL^{+}17]. SCPS [SLC^{+}15b]. Scrambling [LLHC15]. Scratchpad [EKJ^{+}10, LGMP10, MB12a, VAN^{+}18, vdBGLGL^{+}16]. SD3 [KLKL13]. SDR [DMA^{+}15]. SDRAM [EE17, GCAG16]. Seamless [SMN^{+}17]. Search [CLS10, CYJ^{+}10, CSW^{+}15, FEM^{+}18, HH17, HWG^{+}14, JSC10, LCL15, LWL^{+}16, LCW^{+}15, LMT13, SLC^{+}15b, XJWW13, ZLX^{+}16]. Searchable [CLW16b]. Searching [CTL^{+}17, LXZ^{+}15, PWW^{+}11]. Second [DP13, VSLD15, YCW11, YLP15]. Second-Level [VSLD15]. Second-Order [DP13, YCW11, YLP15]. Secret [HL10a, LCCJ13, LLKA19, WKB16]. Section [Ano10c, Ano11e, AHI12, AISA16, BKPMC13, BMM11, BS10, BKP16, BCS11, BdD19, CHTD19, DPO17, EM12, GM11, MG11a, MOS14, NST14, SLPB18, ST11a, XL16, ZMS13, Avr13]. Secure [AP14, CSS13, CXYC16, CYHL14, GV14, HSM14, JAS^{+}15, KW14, KYZC19, KH10, LCH^{+}15, LLXC16, LRY^{+}15, MW10, NBZP17, NDG^{+}17, PSM17, QZL^{+}16, SSKL16, SLPB18, SZDL14, TLH^{+}16, WCW^{+}13, WZBB15, WKB16, WRW16, XJWW13]. Securing [CMLS15, OGPK14, SWF^{+}19]. Security [ASBdS16, Ano10g, BQP^{+}16, BBA19, CTL^{+}17, DY14, GSF^{+}10, HMZ^{+}14, HLT^{+}15, JSA17, LKH16, LCwW10, LLS^{+}16, NDG^{+}17, SMCN18, TLZV11, WCL^{+}18, ZL18]. Security-Driven [TLZV11]. SEDUM [LS13]. SEED [GLXY13]. Segment [CLS10, DDNT19]. Segmented [KCL19, TYY^{+}16]. Select [KT19]. Selecting [ZL15]. Selection [AT16, AHNT16, CHCK12, CCP^{+}13, DZLP14, EFPC16, KwPK^{+}15, KAH19, LWW11, RBR13, RXC^{+}15, Rus13, SEY14, THM^{+}14, TCYH15, YFJ^{+}14, YFCV14]. Selections [DCCK18]. Selective [ADC11, JSA17, KRP18, LOH17, MTGM12, OYP^{+}18, QQW^{+}17, SNM16].

Selective-Testing [SNM16]. Self [ADC11, BCSR14, BCD^{+}16, CCV^{+}11, CWL^{+}17, CRJZ16, DYHX16, DPS11, DKK16, FFCB14, GVPS19, GEvS10, HZ11, HMC11, LLW^{+}17, PRGBSAC19, RO11, RSU17, SOM^{+}13, SRCK10, TW10, YyHL11, YZ15, ZNL18, ZZ17]. Self-Adaptive [DYHX16, FFCB14, YyHL11, YZ15]. Self-Aware [GVPS19]. Self-Checking [HMC11]. Self-Diagnosis [TW10]. Self-Healing [CWL^{+}17]. Self-Invalidations [ADC11]. Self-Loading [SRCK10]. Self-Organized [DKK16]. Self-Organizing [GEvS10]. Self-Reconfigurable [SOM^{+}13]. Self-Reconfiguring [ZNL18]. Self-Reference [ZZ17]. Self-Repair [LLW^{+}17]. Self-Repairing [RSU17, TW10]. Self-Routing [RO11]. Self-Sustainable [CRJZ16]. Self-Synchronized [PRGBSAC19]. Self-Synchronizing [HZ11]. Self-Test [BCSR14, BCD^{+}16, CCV^{+}11, DPS11]. Semantic [CJ12, CH14, HLJ14]. Semantic-Aware [HLJ14]. Semi [WF12, ZQZ^{+}19, KMC17]. Semi-Clairvoyant [ZQZ^{+}19]. Semi-PArtitioned [KMC17]. Semi-Systolic [WF12]. Sense [Zhe10]. SENSIBle [GEN^{+}17]. Sensing [FLJ14, JGG^{+}14, KCW^{+}17, LCHC14, LZZV16, WAK^{+}17]. Sensitive [DY12, HXVF12, KS14, QZC15, QGPZ13, YCCJ15]. Sensitivity [EGVFC^{+}12]. Sensitization [SBP16]. SenSmart [CGL^{+}13]. Sensor [AO11, AO12a, AEKT15, ASTU10, AD10, AD12, Amm14, BWCW15, BWV15, CJG16, CLR13, CSJ^{+}11, CGL^{+}13, CCD12, CBTU14, CBVL16, DY12, DCL^{+}11, DLL^{+}12, FS10, GDC^{+}16, GHG^{+}14, GLTC16, HXVQ15, HKWC14, HWX15, JGHD11, KKT15, KLT16, KH10, LRC10, LR13, LY11, LWW11, LLZ^{+}17, LSX13, LCHC14, LKLT12, LCT11, LZYL13, LJY^{+}15, MM16, MB12b, MMH14, MMB14, PP10, RGK15, RCN11, RLX15, RS13, RNS13, SBH11, SPC^{+}16, SCK10, VBR^{+}13, WJL^{+}14, WLYY16, WS15, XCW^{+}10, YKK^{+}15, YASS14, ZLG^{+}15, ZMY11, ZY12, ZWD^{+}16, ZLYS15, dAJM14, GEN^{+}17]. Sensor/Actor [ASTU10]. Sensors [WCLY16, YLA10]. SenSpire [DCL^{+}11]. Separable [SKM^{+}13]. Sequence [BCMJ10, LBS15, MGW14, SKPK10, YMAG17, YLH13]. Sequences [Jes15, MG16, Pom15c, Pom15b, SN16]. Sequential [LHL13a, MVB10, Pip11]. SER [HCY18]. Serial [ARM16, CLL^{+}14, FBE^{+}18, RM15b]. Serial-Out [ARM16]. Series [DGC^{+}15, ZLY15, Ano10g]. Serpent [PC16]. Server [BSM^{+}14, CLS14, DSY^{+}15, GY15a, GY15b, LZ15, MBM11, PBL16, THM^{+}14, XLF15, ZT15]. Servers [ABEP16, GCL^{+}13, HWS^{+}17, JJK^{+}11, LW15, SYK14]. Service [AK15, CCC^{+}17, DKW15, DHC^{+}16, EFPC16, LHH14a, LHY13, MLOL15, ML13, NZLK14, PAC^{+}12, RSNK17, SKPC15, TJH^{+}15, WYL^{+}15, YZHX12, YCLH16, ZCL^{+}16, ZWC^{+}18, ZL15]. Service-Level [CCC^{+}17]. Service-Oriented [ZL15]. Services [CCM14, CLX14, LLCH13, LZW^{+}15, SLL15, YXZZ14]. Set [CYJ^{+}10, DAS14, DPS11, EJ15, Hia17, HGW^{+}17, LCA10, LSA^{+}17, LPCW14, NM10, RS17, SMG14, SJS10, YCZ10]. Set-Associative [DPS11]. Set-Up [SMG14]. Sets [CL12, EF12, GL19, Hia16, HT16, MIS^{+}14, NI11, Pom16b, Sou15, Ano11g, Ano11h]. SEU [WNKL16]. SEU/MBU [WNKL16]. Shadow [TZL^{+}14]. Shamir [WKB16]. Shared [BZ15, BBB16, CFL^{+}18, DMK^{+}15, JCM16, KIJ14, PPND17, VMB19, YHC19, ZJS14, ZL16, ZLSI17]. Shared-Memory [DMK^{+}15]. Sharer [ST17]. Sharing [AKKH12, BGRH15, CS15, CLW16b, HL10a, HLT^{+}15, KSEG15, LCCJ13, LSW15, LLKA19, NH10, PSM17, Pom14, SLLG15, SLC15a, SPTC15, SHH^{+}16, WKB16, YY10, YY14]. Shave [ZMW15]. Shell [XHZC16]. Sherlock [ADJG19]. Shield [NDG^{+}17, KGV16]. Shift [MFT^{+}17]. Shifted [HF15]. Shifting [MCM16]. Shingled [WFY^{+}17]. Shoot [YXWL16]. Short [BEHL^{+}19, FSL^{+}17, LAAM11, WQZ^{+}16, ZZJ^{+}19]. Short-Read [BEHL^{+}19]. Shortest [Fen14, FK15, SKA10]. ShortPath [PSND16]. Shuffling [YWQX15]. Shuttle [cCWS14]. SIC [VK15]. Sick [YSLL16]. Side [Bar16, BMZ17, BK12, CSS13, CATB19, DP13, KASZ13, LJL13, MAG^{+}17, NDC^{+}13, RBG^{+}19, STE17, ZMB18]. Side-Channel [Bar16, BK12, CSS13, CATB19, KASZ13, LJL13, NDC^{+}13, ZMB18]. Sided [TYY^{+}16]. Sidewalk [PPB^{+}14]. Sieve [GKB^{+}10]. SIFT [OPAGS14]. Sig [QGPZ13]. Sign [Hia16]. Signal [BBB^{+}17, CK15, GDLL18, LVJ16, MTGM12, Pom13b]. Signal-Transition [Pom13b]. Signals [LVJ16]. Signature [EFGT18, KSN^{+}15, TXL11]. Signature-Based [KSN^{+}15]. Signatures [AS16, GLP15, OWP16, QGPZ13, YCK10]. Signed [Kor15, SJS^{+}14, TAM^{+}16]. Signed-Digit [Kor15, SJS^{+}14, TAM^{+}16]. Silent [LS10b]. Silicon [BKH^{+}13, COLK18, DN11, EKA17, HMR^{+}17, NZ14, NL14, OHCK17, OCK17, PKC^{+}17, WTZ^{+}19, YSLL16, khR^{+}18, KN12]. Silver [Ano10g]. SIMD [HMS^{+}12, NVB16, YMG15]. SIMD/MIMD [NVB16]. Similarity [ADJG19, HLF14, PR10, XJFH15]. Similarity-Aware [HLF14]. SIMON [LYK19, STE17]. Simple [Fen14, LVMS18, SD18, TLH^{+}16]. Simulation [ADP^{+}15, ASS^{+}18, BPT10, FHL^{+}18, GM15, GCR^{+}19, GABK11, JCY^{+}13, LR16, LZZ16, MNFA14, MHRARG^{+}14, NZ15, Tho15, WLV^{+}14, WZLS16, YEG^{+}15, ZGR13]. Simulation-Based [GABK11]. Simulations [KN13]. Simulator [KAH18a]. Simultaneous [GSL10, KCRG15, LR18, YG10]. Single [ARM13, BPT10, ERRM16, GGSPM18, KP15, KMM16, LP17, LCA10, MFG14, MBGS10, NL16b, NL18, PROM15, SBB18, SMRML17, SBI12, TKL^{+}14, WNKL16, XLX^{+}14, ZGR13, ZXX^{+}14]. Single-Chip [BPT10]. Single-ISA [GGSPM18]. Single-Pass [ZGR13]. Single-Qubit [KMM16]. Single-Rate [MBGS10]. Single-Sink [SBI12]. Sink [Amm14, CJG16, LSX13, SBI12]. Sink-Free [CJG16]. Sink-Hole [Amm14]. SinkTrail [LZYL13]. SiPs [TMS^{+}14]. SISO [NCD^{+}17]. Situ [NY15]. Six [AFC10]. Size [BLN^{+}15, CS11a, CJ13, DAS14, DCM16, FLS16, Kim15, KLC^{+}16, LCLL15, LFJ^{+}13]. Size-Aware [BLN^{+}15]. Size-Based [DCM16, LFJ^{+}13]. Sized [CLM^{+}19]. Sizes [DALD18]. Sizing [MBGS10, VTW16]. Skeleton [LJY^{+}15]. Skew [LZA^{+}16, NLRB17]. Skewed [Pom12c]. Skewed-Load [Pom12c]. SLA [KGP15]. SLC [CCC^{+}17]. Sleep [AO12a, CFL^{+}18, HKWC14, ZLYS15]. Sleep-Aware [CFL^{+}18]. Sleep-Wake [HKWC14]. Slice [SLS^{+}12, Zot10]. Slicing [ABSK15]. Sliding [JRS^{+}15, LPL10]. sLiSCP [ARH^{+}18]. Slot [MWLJ15, RLX15]. Small [AO12b, BDE^{+}11, CFR^{+}14, CJ13, IS14, Kim15, KCK16, LK15b, LCLL15, NR15, UdDG^{+}17, VTA16, WGZ^{+}15]. Small-Characteristic [BDE^{+}11]. Small-Value [IS14]. Smart [DYCG16, EFPC16, GY16, HK17, HDYS16, LVF19, SWZG15, WZY16, XLC14, LDMQ16, WHBR16]. Smartphones [LTL14, OPZ15]. SMR [WFY^{+}17, YCW^{+}19]. SMT [FSPD16, FSPD17, MBC^{+}13, OPAGS14]. Snakes [PC16]. Snapshots [YCL^{+}12]. SNM [GBA18a]. Snoop [AKKH12]. Snooping [AKKH12]. SNOW [PC16]. SoC [BBA19, CFR^{+}14, LYH11, VTA16, WZBB15]. Social [JWL^{+}16, LCL15, LSS13, LS13, LSW15, PLZW14, SL14a, SLLG15, SLC^{+}15b, TZL^{+}14, WHC^{+}15b, WW14, XWH14, ZZX^{+}15]. Social-Aware [SLC^{+}15b]. Society [Ano10d, Ano10e, Ano10b]. Socket [CG18, LHH17]. SOCs [IGLM15, RC14]. Soft [AF14, AS12, DZLP14, EGVFC^{+}12, HBR11, KS12, NEE18, RURM18, RTRM19, RCN11, RBMO11, SB16, SMAR^{+}19, SVD18, UVG16, VSC^{+}19, XH16, YAGB17, dOPSR16]. Soft-Errors [UVG16]. Soft-Processors [RURM18]. Soft-Real-Time [AF14]. Softcore [PPP13]. Software [AVG^{+}15, ADC11, AH13, BDL^{+}13, CSS13, DPS11, FHLOJRH18, GBO^{+}16, HPR16, HWG^{+}14, HL10b, HGL^{+}15, JSC10, KKP^{+}16, KASZ13, LK10, LH16, LSHC15, MRL^{+}18, MMAC19, MF14, RCK^{+}16, SNM16, TS11, WGW^{+}15, WLJ^{+}16, WHYS16, WFT^{+}19, XHZ14, YLGE14, ZPM^{+}15, ZLG^{+}15, ZGG^{+}16]. Software-Based [ADC11, DPS11, SNM16, YLGE14]. Software-Defined [ZLG^{+}15, ZGG^{+}16]. Software-Driven [TS11]. Software-Transparent [WFT^{+}19]. Solid [Cha10b, CLZ19, DSW^{+}14, JR17, Jun16, KLK18, KBH^{+}10, KSJ^{+}12, LRP16, MLE14, PDXZ13, SNY^{+}10, TAH^{+}16]. Solid-State [Cha10b, CLZ19, DSW^{+}14, KBH^{+}10, PDXZ13, SNY^{+}10, TAH^{+}16]. Solomon [PROM15, TW10]. Solution [CCE^{+}18, CXLX15, GSH^{+}14, NZ15, RBRL15, SKPC15]. Solutions [CCO^{+}14, PLM16]. Solve [MTFK19]. Solvers [DALD18]. Solving [DALD18, Fuj11, GO10]. Some [JLMP11]. Somewhat [MBF18, RJV^{+}18].

Sophisticated [BPBBL13]. Sorting [FFISC13, LNCX18, LCL17, LBSK17]. SOSEMANUK [PC16]. Sound [AGCD16]. Source [ADJG19, CJ12]. Source-Code [CJ12]. SPAC [Pan16]. Space [BNP10, BKL^{+}13, BCMJ10, BBH12, CHN14, CCY^{+}16, CYC^{+}16, DYW15, DJO11, Fen14, HN11, HMNN12, HK17, JAD^{+}18, KH18, LWF^{+}17, Nan16, PCHS17, PCHS18, SN16, XDZ11, ZMY11, KMC17]. Space-Division [ZMY11]. Space-Efficient [BBH12, LWF^{+}17]. Space-Time [DYW15]. Spaces [LTW^{+}12]. SPaCS [ZGR13]. Spam [SL14a, ZL11]. Spanning [FEP^{+}12, SBI12, TYWC10, WTY^{+}14]. SparCE [SJVR19]. Spare [AGFM11, SC11, AAR19]. Spare-Enhanced [SC11]. Spark [CZL^{+}17]. Sparse [JSC^{+}19, LPL^{+}13, LJG^{+}19, MRD19, NZ15, PP16, RO11]. Sparse-Insertion [JSC^{+}19]. Sparse-Iteration [PP16]. Sparsity [LYL^{+}19, SJVR19]. Spatial [DNSS11, KGGJ14, MWY^{+}16, SSW12, XHZC16, ZCL^{+}16]. Spatio [CSPC12, DYCG16]. Spatio-Temporal [CSPC12, DYCG16]. Spatiotemporal [Cro14]. Special [Ano10c, Ano11e, AHI12, AISA16, Avr13, BKPMC13, BMM11, BS10, BKP16, BCS11, BdD19, CHTD19, DPO17, EM12, GC14, GM11, HMO^{+}17, LLK18, MG11a, MOS14, NST14, SLPB18, ST11a, VP14, XL16, ZMS13]. Specially [LS10a]. Specific [AK15, CKKS14, JRC14, LSA^{+}17, MGW14, SP16]. Specifications [MIS^{+}14]. Specified [YXWL16]. SPECT [KN13]. Spectral [YCW11, YLP15, ZMB18]. Spectral-Null [YCW11, YLP15]. Spectrum [BBVL14, NH10, XWL10, XLTZ11]. Speculation [KGGJ14]. Speculative [GLXY13, LR18, STR15, XLC14, YLGE14]. Speculatively [Bra10]. Speech [AWFV13, TBC^{+}17, YAG19]. Speed [ARM16, Ano13f, CCH^{+}15a, GJ15, GY13, GCF^{+}16, HK13a, HZ11, NWA12, PBT13, TFCY16, ZL11]. Speed-Up [GCF^{+}16]. Speeding [WGR^{+}14]. Speedup [ZYL15]. Spike [ZM17]. Spike-Timing-Dependent [ZM17]. Spiking [AAR19]. Spin [CWCS15]. SpiNNaker [BFR^{+}15, FLP^{+}13]. SpiNNaker-Programming [BFR^{+}15]. Spintronic [BDDL18, ZCZ^{+}19]. Split [CNH13, DMA^{+}15]. Splitting [HN13, PCHS16]. SpMV [YLML15]. SPONGENT [BKL^{+}13]. Sporadic [MFG14]. Spot [JT15]. Square [ARH14, DRM16, FBE^{+}18, GPN11, JKMR11, JLLH19, LP13b, LN12, NL15a, Rus13, WE12]. Squashing [YLGE14]. Squeeze [CSW^{+}15]. Squeeze-Search [CSW^{+}15]. SqueezeFlow [LJG^{+}19]. Squeezing [FSGAB^{+}16]. SRAM [AKL14, BMS11, CW16, EKA17, GBA18a, GTRMG18, KL13, RTRM19, SKH16, SBMP18, VPS^{+}12, WWY^{+}16]. SRAM-Based [AKL14, BMS11, EKA17, GBA18a, GTRMG18, KL13, RTRM19, SBMP18]. SRAMs [LCY^{+}13]. SRT [Rus13]. SSD [CDQB15, CLCS19, HWS^{+}17, HLY14, HCL15, IS11, KLC^{+}16, KWC^{+}16, KRP18, LBSK17, LLL16, WZ14]. SSD-Based [CLCS19]. SSDs [CLZ19, CCC^{+}17, HJF^{+}13, JSH^{+}17, KSJ^{+}12, LK16a, LCY^{+}16, LSG^{+}15, LSGZ16, SQJ^{+}15, ZWW19]. ST [YCL^{+}12]. ST-CDP [YCL^{+}12]. Stability [AKKH12, BIP^{+}17, Yam10]. Stability-Optimized [Yam10]. Stabilizer [GM15]. Stabilizing [KkC15a, SKA10]. STABLE [GBA18a]. Stack [CGL^{+}13]. Stacked [KAH^{+}15, MKLW14, SPC^{+}18, XCF16, ZDYZ13, ZDYZ14]. Stackless [MS15]. Stage [KGP15, PLZW14, SVAB14, TS11, ZWL^{+}19]. Stage-Level [KGP15]. Staged [KTAvdS16]. StageNet [GFAM11]. StaleLearn [LK18]. Standard [Red19, MKRM10]. Standby [LXJD15, LXDV17]. Star [CL12, ZLXW15, ZGWC15]. Start [LVMS18]. State [AWFV13, Cha10b, CLZ19, DSW^{+}14, FHR14, Hie11, Hie13, HT16, JR17, Jun16, KLK18, KBH^{+}10, KSJ^{+}12, LRP16, LLQ^{+}14, LLK18, Lom10, MTGM12, MLE14, MCXZ18, Mon15b, Mon16, MMC^{+}16, Mon17, Mon18, Mon19, OPZ15, PDXZ13, PBV11, RRK11, SNY^{+}10, SP10, TAH^{+}16, Yun12, Zom11b, Zom11a, Zom12b, Zom12c, Zom13, Zom15a]. State-Retentive [RRK11]. Stateless [XWY10]. States [Pom15a, PRM19]. Static [ABSK15, CHK10, CTD^{+}16, LWY15, LS10c, NM10, Pom12d, Pom15c, RBG14, TCK^{+}18]. Statically [GLXY13]. Statistical [CGT^{+}15, CTL^{+}17, CRK10, HZ11, RCM^{+}16, SIB13, SDP^{+}12, SZL^{+}16, ZL11]. Statistics [WLG^{+}19]. Status [XWL^{+}16b]. Stay [Ano11i]. Stealing [CG18]. Steiner [LHPH15, LSZ^{+}15, SG12, SG13]. STEM [TCHL18]. Stengle [MRD19]. Step [RLRL19, LXZ^{+}15]. Stereo [PTD^{+}12, TKT16]. STES [CMLS15]. Stochastic [BPT10, CBZ14, Dar15, GO10, HXVQ15, HCL^{+}14, LB15a, LLQ^{+}14, LTVL15, LZZV16, LLW^{+}18, NLRB17, QLR^{+}11, YMT13, ZGB^{+}15]. Stojmenovic [Zom15b]. Stop [HCC^{+}18]. Storage [ASM^{+}16, AT16, AISA16, BDP15, CK11, CHH^{+}13, CHLT14, CQW^{+}15, CXYC16, CYC11, CYL^{+}14, CLW16b, EFPC16, FYSK14, FHL^{+}18, GKD^{+}17, GWM^{+}17, HSM14, HCD^{+}16, HQLX15, HLC^{+}19, JJZ^{+}16, JRP^{+}14, LBSK17, LSK13, LK14, LS10a, LOX^{+}13, LLL15, LSG^{+}18, LRY^{+}15, LLS^{+}16, MJWT16, NL16b, PP11, PPKW12, ROGHNB^{+}18, SWZG11, SZG^{+}18, SYH17, SZW^{+}16, SYK14, TRA18, TCYH15, WL13, WCW^{+}13, WLK15, WHL17, WJF^{+}11, WSXZ13, WJM15, XL16, XDZ11, XLX^{+}14, YCKH16, YTD^{+}17, YTD^{+}18, YHT^{+}16, ZWW^{+}16, ZFJ^{+}17, ZWC^{+}18, ZMW15, ZXX^{+}14, ZLLX15]. Storage-Based [LOX^{+}13]. Storages [CCY^{+}16]. Store [LK15a]. Stored [CMLS15, JPG10, UMN18]. Stores [ZHQX19]. Storing [YLY15b]. Strategies [BBI^{+}13, CTL^{+}17, CBTU14, MW13, TLB^{+}17, ZV14]. Strategy [CK11, CCL^{+}18, JLC10, MMCS18, XLC14, YZ15, ZL15, ZGWC15]. Strategy-Proof [ZGWC15]. Strategyproof [XWL10, XLTZ11]. Stream [CMLS15, GZG^{+}16, GCS^{+}13, HZ11, KE19, LTP^{+}14]. StreamCiphers [ERRMG15]. Streaming [CPL17, CLMM11, KTAvdS16, LOH17, LHH14a, LLLJ13, MLH12, PAC^{+}12, RBRL15, SHGW15, WLQS13, YW12]. Streams [Ged14, LLQ^{+}14]. Stress [GBA18a]. Stress-Aware [GBA18a]. String [ASM^{+}16, GNSR14, LP13a, PWW^{+}11, ZS13]. Stripe [FSL^{+}17, KLC^{+}16]. Strong [HTC13, WHL17, ZGW14]. Strongly [KW14]. Structural [BWCW15, GLTC16, TLL12]. Structure [AKJ^{+}13, CZP^{+}16, HHY11, JP13, KS10a, MKRM10, PPKW12, SJD^{+}18, XHZC16, YWW^{+}16, ZZ17]. Structure-Aware [CZP^{+}16]. Structure-Independent [MKRM10]. Structured [CYC11, MLE14, SL13, SLC15a, WJF^{+}11]. Structures [ALW11, FLS16, GCL^{+}13, LHYZ13, PWTS16]. STT [AYC16, CKD^{+}17, CFMA19, FKMK16, GBGI18, KCW^{+}17, KA19, LHL^{+}15a, YLGD19, ZOK^{+}19]. STT-MRAM [CFMA19, GBGI18, KCW^{+}17, ZOK^{+}19]. STT-RAM [AYC16, CKD^{+}17, FKMK16, KA19, LHL^{+}15a]. Stuck [MMC15]. Stuck-At [MMC15]. Studies [WHC^{+}15b]. Study [AE11, AD10, AR17, CTS13, CCLH10, GLP^{+}12, GM12, HHC^{+}18, SD14, VIDH19, WRW16]. Stuffing [CCH^{+}15a]. Sub [BBPQ15, GD17]. Sub-Networks [GD17]. Subarea [XLW14]. Subbanking [LGMP10]. Subclass [WHL^{+}12]. Subcodes [Red11]. Sublinear [DJJ^{+}08, Lee12]. Submicron [SSJ^{+}18]. Submillisecond [HGML11]. Subnetworks [LY11]. Suboptimal [TLP18]. Subpage [KLK18]. Subpage-Aware [KLK18]. Subquadratic [BNP10, CHN14, CLL^{+}14, HF15, HMNN12, PCHS17, PCHS18, RM15b]. Subscribe [BS15]. Subsequences [Pom12a]. Subset [BS14, BS16]. Subspace [AKKH12, PZZQ19]. Substitution [CJK15]. Substitution-Permutation [CJK15]. Substrate [WTBT13]. Subsystem [CVPS19]. Subthreshold [LCY^{+}13]. Subtraces [USP^{+}13]. Subtraction [KBP13]. Succinct [EG11]. Suffering [ZHM^{+}19]. Suffix [LNCX18, NZC11, WNCH17]. Suitable [JL11]. Summation [BGHL19, DN15, Lef17]. Sums [BCTV19, BHR17, KLLM12]. Supercomputer [CTD^{+}16, LLL^{+}17]. Supercomputing [YWXZ12]. Superior [ST17]. Superpeer [GEvS10]. Superpoly [HIJ^{+}19]. Supersingular [BDE^{+}11, FHLOJRH18, LNL^{+}19]. Supervised [FCB^{+}19]. Supplies [LLLJ13]. Supply [DYW15, DMXY14, GSK12, PdG13, QZL^{+}16]. Support [CCO^{+}14, CLM^{+}19, DMK^{+}15, FGS^{+}13, GSG^{+}15, HMS^{+}12, KLK^{+}14, KT12, LKYC12, LKH16, LSG^{+}15, MJW^{+}14, MWWT13, MW10, MRW^{+}15, RKN^{+}18, SAR^{+}11, SBP^{+}14, WLZ^{+}15, WGW^{+}15, WJY^{+}17, ZYW^{+}16, ZCK19]. Supported [LCL15, ZGG^{+}16]. Supporting [CT13, KPBC17, LXW^{+}19, LSGZ16, SMG14]. Supports [KH18]. Survivable [Ano13g, GSG^{+}15]. Sustainability [CFW14, MOS14]. Sustainable [CCO^{+}14, CRJZ16, MOS14]. SVM [JAS^{+}15]. SW [MH19]. Swapper [ZLSI17]. Swapping [GCF^{+}16, LZZ^{+}17b]. Swarm [SLC15a]. Swarms [WAK^{+}17]. Switch [DKG13, KKH^{+}14, SRR^{+}16]. Switched [CWF14, LKMSA16]. Switches [ACGP13, AHK10, BGMR13, DKG13, GY13, JPLP13, MOYB12, ZWLS15].

Switching [AR12, CCH^{+}15a, GSK12, Pom12d, Pom13a, SLS^{+}12, SO10, Tho15]. Sword [WHC^{+}15b]. Symbol [MCM16, NL16b, PROM15]. Symbolic [BMZ17, NS13, PNKI13, ZJH^{+}14]. Symmetric [HC13b, MM17, ZWYY15]. SymPLFIED [PNKI13]. Synchronization [BBB^{+}17, HWZ^{+}12, KLJ^{+}14, LK18, TFCY16, WZLS16, ZJXL11]. Synchronization-Aware [HWZ^{+}12]. Synchronization-Based [BBB^{+}17]. Synchronized [PRGBSAC19]. Synchronizer [Zot10]. Synchronizing [HZ11]. Synchronous [LYL^{+}19]. Syndrome [Red14]. Synergistic [Pan16]. Synergizing [LHH14b]. Synergy [LHTG15]. Synthesis [AR12, BM13b, CBZ14, DYJ19, DSR15, FSR^{+}18, GAFN15, GM12, IB10, JWC12, JRC14, KK18, LH11, MY10, MIS^{+}14]. Synthesizer [DSKH15]. Synthesizing [RBIQ15]. Synthetic [GJ14]. SyRA [VSC^{+}19]. System [ADI11, ASM^{+}16, AHNT16, AE11, Ano11f, AC11, BKH^{+}13, BS15, BBP^{+}13, BJ12, BS10, BM13b, BSM^{+}14, CMS10, CXZ13, CCO^{+}14, CWZ11, CHLT14, CCC^{+}18, CLCS19, CH14, DKH^{+}13, DSB13, DFP^{+}13, DALD18, DCL^{+}11, FLP^{+}13, IBH^{+}13, JZLD10, KM11, KGC14, KP15, KCE^{+}18, KBP13, LK10, LYB15, LKK^{+}19, LSW15, LLD^{+}16, LLS^{+}16, LYJ^{+}18, LVF19, LSGZ16, MBM11, MMCS18, MLE14, MKM14, NLP^{+}14, PP11, PBT13, QWB^{+}13, ROGHNB^{+}18, RM15c, SOM^{+}13, SPC^{+}16, SCZ^{+}16, SWZG11, SLS^{+}12, SN16, SKM14, TKT16, VSC^{+}19, WZLX12, WLK15, WWY^{+}16, WP16, WSXZ13, XZL^{+}19, YY10, YSZ^{+}14, YXZZ14, ZGG^{+}16, ZZ10, ZMS13]. System-Dependent [JCY^{+}13]. System-Level [AE11, Ano11f, BS10, BM13b, LK10, WP16, ZMS13]. System-on-Chip [BKH^{+}13, LYJ^{+}18]. System-Wide [KCE^{+}18, SKM14]. Systematic [BGM^{+}13, DRC14, PWTS16, PC16, Red11, SZW^{+}16]. SystemC [PFGB14]. Systems [AJH15, AXS^{+}10, AEP18, AE11, AFH^{+}10, ARGT14, Ano13g, AISA16, BBPQ15, BKPMC13, BDM^{+}19, BIP^{+}17, BBB16, BMS11, BM13b, BKP16, BDL^{+}13, BGRH15, BDBB18, CAbZM18, CW10, CK11, CHH^{+}13, Cha14, CZ14, CHC^{+}15, CHTD19, CCH11, CS11b, CYCC11, CQW^{+}15, CCY^{+}16, CJA^{+}16, CvdBC18, CLM^{+}19, CQ14, CDK^{+}18, CK15, CH14, CWCS15, DZD^{+}16, DSB13, DYCG16, DLC^{+}13, DSY^{+}15, DW10, DKK16, EKJ^{+}10, FFCB14, FYSK14, FSR^{+}18, GCD^{+}11, GGSPM18, GV15, GGA^{+}17, GWM^{+}17, HRK17, HWZ^{+}12, HNV19, HWS^{+}17, HWSX17, HS18, Hie13, HCH15, HTC13, HHLK12, HWL^{+}14, HZW^{+}12, HT12, HWSN15, HLA^{+}17, HHW^{+}18, HLC^{+}19, JAJK15, JKY10, JKD^{+}19, KS10a, KIJ14, KMC17, KMLH11, KAH19, KCRG14, KLJ^{+}14, KH18, KAQC14, KA19, Kůr12, LBSK17, LSK13, LSSE15, LXJD15, LXDV17, LS10a, LSS13, LOX^{+}13, LLL15, LCH^{+}15, LTVL15, LSG^{+}18, LcZLG19, LH12b, LKT13, LZZZ13, LMB^{+}16, LLW^{+}11, LXZ^{+}15]. Systems [LZZV16, LZZ16, LHH17, LJVJ18, MW10, MJWT16, MS15, MBB^{+}17, MG11a, MUMB11, MBSSA19, MNK11, MCXZ18, MTBB10, MW13, NL16a, NL16b, NVB16, OGPK14, OKC13, PKC^{+}17, PDXZ13, Pan16, PCLN15, PBL16, PC10, PBE17, QZL^{+}16, QLH^{+}16, RVC^{+}15, RF14, RDEN10, RSU17, RBR13, SSKL16, SDP^{+}12, SVD18, SIVH16, SMN^{+}17, SSW12, SL13, SLC15a, SZG^{+}18, SKM14, SL14b, ST11a, SD13, SYH17, SWZG15, TLZV11, TB15, TSK16, TRA18, Tsa13, TFCY16, TS11, VSC^{+}19, VSF^{+}17, WS14, WLC^{+}15, WZM^{+}16, WSZ^{+}16, WSL^{+}18, WTZ^{+}19, WHL17, WhCCC12, WA10, WLZ10, WJF^{+}11, XDZ11, XLX^{+}14, XTW15, XSR15, YLY^{+}15a, YCLH16, YHC19, YPB^{+}16, YYW^{+}16, YTD^{+}17, YHV13, ZLBB19, ZWW^{+}16, ZFJ^{+}17, ZFHC19, ZQZ^{+}19, ZZ19, ZV14, ZZM^{+}15, ZL15, ZHM^{+}19, ZXX^{+}14, ZLLX15, ZCR16]. Systems-on-Chip [NVB16, RVC^{+}15]. Systolic [CVGZ15, CLL^{+}14, LLOS13, RM15b, WF12]. Sytare [BDM^{+}19].

T [GNSR14, ZGR13]. T-SPaCS [ZGR13]. T-Transform [GNSR14]. TA-LRW [CFMA19]. Table [KCS14, SPC^{+}18]. Tables [HC13b, HHY11, HH17, LJ13, LS10c, MS12, Rus13, dLSGDR17]. Tables-and-Additions [LJ13]. Tackling [EE17, USH19]. TACLC [CYCC11]. Tag [KCL19, LXZ^{+}15, YXWL16]. Tagged [DSB13, RKN^{+}18]. Tagging [KLK17]. Tags [LLM^{+}15, SDZ15]. Tailored [NCD^{+}17]. Tainting [DCV^{+}12]. Taming [CVH^{+}13]. TAP [LCLC19]. Target [BWV15, GJ14, SBH11]. Targets [WF14]. Task [AO11, CFL^{+}18, CCK^{+}16b, GCR^{+}19, GZB^{+}15, HV14a, HLWV17, HGW^{+}17, HNB^{+}12, KLK^{+}14, LC16a, LYJ^{+}18, LMB13, MBB^{+}17, PAC^{+}12, PP10, PM14, RBR13, TB15, TFCY16, WLZ^{+}15, WJY^{+}17, ZLBB19, ZLG^{+}15, ZGG^{+}16]. Task-Based [GCR^{+}19]. Tasking [CvdBC18]. Tasks [BMP^{+}10, Bin15, CGL^{+}18, CP10, GHK15, HV13, HV14b, Li12b, LTVL15, LGS^{+}18, MFG14, MW13, NRG15, TLZV11, WBZ^{+}15, WGR^{+}14, YTM16, ZQQ11, ZCYX15, ZMRQ11]. TC [Ano13a, Ano13b]. TCAM [BSS14, BSS15, BBH12, CWZC13, CW16, KCS14, MS12, RCRK13, SP12, Yun12]. TCAM-Based [BBH12, Yun12]. TCAMs [Li12a]. TCP [YCCWC15, ZWH^{+}15]. TDM [SMG14]. TE [ZMW15]. TE-Shave [ZMW15]. Technique [BCSR14, DSR15, FTP13, GSK12, KTA^{+}14, LSC11, LCT11, PPP13, RBMO11, SRCK10, TW10, ZZ17, ZMY11]. Techniques [AS12, CM11, CZP^{+}16, EKJ^{+}10, FBR^{+}12, FCB^{+}19, Fuj11, GPRS17, GTRMG18, IS11, LLHC15, MYW11, NM10, RCN11, SP12, TJH^{+}15, TLL12, XL16, ZR15a]. Technologies [AISA16, BMM11, CHTD19, LHH14b]. Technology [ACM^{+}16, WWY^{+}16]. Telephone [CB15]. Temperature [AEGH19, HV12, JCY^{+}13, KkC15a, LSC10, YMG16]. Temperature-Aware [HV12, KkC15a, LSC10]. Template [ZYY18]. Temporal [AKKH12, CSPC12, CFMS14, DYCG16, FLS16, LXL^{+}13, LCX^{+}16, LC16a, MB12b, SSW12]. Temporary [ZZJ^{+}19]. Tenant [ZGWC15]. Tensor [KYZC19, Red19]. Term [CYJ^{+}10, Ose11]. Terminal [LHPH15]. Terms [YL14]. Ternary [ADI11, BT16, FAK16]. Test [AK16, BBI^{+}13, BCSR14, BCD^{+}16, CZP^{+}18, CCV^{+}11, CCC15, CSW^{+}15, COLK18, DPS11, DRS^{+}16, FM19, FD16, GPRS17, HRM^{+}16, IGLM15, KN11b, LCY^{+}13, MVB10, NI11, NM10, PN16, PP14, Pom12c, Pom12a, Pom12d, Pom15c, Pom15b, RC14, RBG^{+}19, WEH^{+}19, XCF16, YTND12]. Testability [KT19, MOMT12]. Testbench [PFGB14]. Testing [BCK^{+}16, CVMA10, HTH15, Hie11, HL10b, HWE^{+}16, KL13, Li12a, LC16b, MG11a, MOMT12, NI11, RCC14, ST16, SNM16, SDL^{+}19, XCF16, XKT^{+}15, ZOD13]. Tests [CM11, Pom12c, Pom12b, Pom12d, Pom13a, Pom13b, Pom14, Pom15a, Pom16a, SP10]. Text [QWB^{+}13]. Textual [CYJ^{+}10]. th [KCK16]. Thank [Mon19]. Thank-You [Mon19]. Their [GM12, KBP13, LK10]. Them [MMP13]. Theorem [Fan16, HTA10]. Theoretic [SCK10]. Theoretical [GLH^{+}19]. Theory [AS10, EFPC16, LLL15, LCHC14, SKEB18]. Thermal [AQPMS15, BTBB14, KAH19, LZZZ13, LYJ^{+}18, MMCS18, MKM14, PKC^{+}17, RVC^{+}15, RCN11, SVAB14, TCHL18, WJY^{+}17, WWT^{+}18, XCF16, ZMW15, khR^{+}18]. Thermal-Aware [AQPMS15, KAH19, LYJ^{+}18, MMCS18, WJY^{+}17, XCF16]. Thermal-Constrained [TCHL18]. Things [LGH^{+}17]. Third [QZL^{+}16]. Third-Party [QZL^{+}16]. though [CS15]. Thrashing [LCLC19]. Thread [HGCT13, KK10, LKJ15, LR18, LLX^{+}17, LY18, RCM^{+}16, TLGM17, XLL^{+}18, ZJS14]. Thread-Aware [ZJS14]. Thread-Level [XLL^{+}18]. Threaded [TLGM17]. Threading [CvdBC18, MS15, RCC14]. Threat [YZF^{+}10]. Three [CNH13, CCE^{+}18, Hia16, Hia17, JWH^{+}15, MKLW14, PC16, SG13, Ste14]. Three-Dimensional [CCE^{+}18, JWH^{+}15, MKLW14]. Three-Moduli [Hia16, Hia17]. Three-Vertex [SG13]. Three-Way [CNH13]. Threshold [Cil11, FSGAB^{+}16, MHK15, RMKR12, STE17]. Throttling [LK16a]. Throughput [AFC10, CVGZ15, CA12b, FFISC13, HGCT13, KAH18a, KHZ17, LMC^{+}15, LCL17, OMFH14, PRM16, RL13, VC10, WSL^{+}18, YM11, ZZYZ14]. Throughput-Efficient [YM11]. Throughput-Optimal [RL13]. Throwbox [LYT^{+}16]. Throwbox-Assisted [LYT^{+}16]. Throwboxes [LYT^{+}16]. Tianhe [XYF^{+}15]. Tianhe-2 [XYF^{+}15]. Tier [LZW^{+}15]. Tiered [SRHC12]. Tight [ST18a]. Tighter [YLH10]. Tightly [DMK^{+}15]. Tightly-Coupled [DMK^{+}15]. Tiled [KPS^{+}17]. Tiling [QLH^{+}16]. Time [ABH^{+}13, AF14, ADP^{+}15, ABEP16, AEP18, AE11, ASBdS16, BM13a, BBK10, BBD^{+}12, BBP^{+}13, BG12, Bin15, BBB16, BDB18, BPC12, BGRH15, CWF14, CBB19, CW10, Cha14, CWZ11, CYCC11, CCC15, CGL^{+}18, CQ14, CLR13, CCAM14, CYL^{+}14, DYW15, DCCK17, DZD^{+}16, DA12, DGC^{+}15, DZ10, DKK16, EE17, FM16, GKB^{+}10, GCLC11, GPRS17, GGA^{+}17, GCAG16, GSF^{+}10, HB11, HXVQ15, HWZ^{+}12, HV12, HCH15, HK16, HHLK12, HWL^{+}14, HXL11, HCC^{+}12, HV14b, HC17, HGW^{+}17, HZW^{+}12, HCZW13, IBH^{+}13, JAJK15, JSH^{+}17, JWWZ16, KS10a, KSS12, KM11, KRR^{+}18, KGP15, KTA^{+}14, KMC17, KLLK11, KCE^{+}18, KAQC14, KDEV19, KT12, LYH11, LR13, LSC11, LHC^{+}14, LAAM11, LMC^{+}15, LKLK13, LSSE15, LXJD15, LXDV17, LC16a, LWF^{+}17, LLW^{+}11, LLM^{+}15, LXZ^{+}15, LGS^{+}18, LMB13, MBD^{+}17, MUMB11, MAHD18, MBSSA19, MFG14, MKM14, MW13, NRG15, NZLK14, NZC11, OHCK17, OCK17]. Time [OPZ15, PTD^{+}12, PC10, PMH^{+}14, RHC^{+}14, RF14, RM15a, RLX15, SRCbL^{+}15, SCJ^{+}16b, TB15, TH11, TFCY16, TCHL18, THGT13, TKT16, VK15, WXS12, WBZ^{+}15, WLY^{+}14, WA10, WLZ10, WSXZ13, WZ14, XWLX17, Yam10, YPB^{+}16, YYW^{+}16, YHV13, YTM16, YAGB17, ZICL12, ZLBB19, ZD13, ZLY15, ZZ19, ZYY18, ZHM^{+}19, ZQQ11, ZCYX15]. Time-Aware [WSXZ13]. Time-Borrowing [CCAM14]. Time-Dependent [LR13]. Time-Domain [HXVQ15]. Time-Driven [BM13a, RM15a]. Time-Efficient [LLM^{+}15, LXZ^{+}15]. Time-Evolving [HCZW13, JWWZ16]. Time-Memory [ASBdS16]. Time-Multiplexed [GCLC11]. Time-Predictable [WA10]. Time-Series [ZLY15]. Time-Slot [RLX15]. Time-Triggered [MAHD18]. Time/Just [JAJK15]. Time/Utility [KM11]. Timed [LH12b]. Timely [STR15]. Times [YLH10, YLA^{+}15]. Timetables [DDNP11]. Timing [CYCC11, CCAM14, GV14, JRJ^{+}18, KTA^{+}14, LR16, MMAC19, MMP13, SBP16, SCJ^{+}16b, VTA16, YyHL11, ZM17]. Timing-Aware [CYCC11, SBP16]. TLB [KLK17, SMAR^{+}19]. TLM [BFP11, PFGB14]. TMDTO [MSS^{+}18]. Toeplitz [PCHS16, RM15b, CNH13, CLL^{+}14, HF15, HMNN12, HN13, PCHS18]. Toffoli [GM12]. Toggle [KCY18]. Toggle-BasedX-Masking [KCY18]. Tolerance [BWCW15, BKP16, DCK16, EYBK15, GV15, HHCH11, JWH^{+}15, JKJ^{+}10, KTA^{+}14, MSC12, MBSSA19, NLRB17, PPP13, RKR15, RRS^{+}16, RZPX19, SJSLD11, SRK^{+}17, SD13, VTA16, VCB^{+}13, WZL^{+}17, ZBK^{+}17, ZL15, ZJXL11]. Tolerant [ALBP14, AE11, BMT14, CFMS14, CHLL16, HCZW13, JK15, KCRG15, KKC15b, LCC10, LW17, LYT^{+}16, LQW^{+}17, LZS^{+}13, QLR^{+}11, RVL^{+}14, SB16, SMRML17, SDE^{+}17, SPH13, WBZ^{+}15, WGZ^{+}15, ZBW17, ZQQ11]. Tolerating [CZS^{+}19, HK15a, KK10, MMC15, MCM16, WNKL16]. Toleration [RBK^{+}12]. Tology [CPL16]. Tology-Aware [CPL16]. Tomography [FM16]. Tool [DYJ19, LVMS18, ZCS16]. Tools [DMXY14]. Top [LRY^{+}15]. Top-Down [LRY^{+}15]. Topics [Ano13d, Ano13c, DPO17]. Topological [CMB13, DLL^{+}12]. Topologies [RKZ16, YUGD14]. Topology [AMVOS^{+}15, CTD^{+}16, CSJ^{+}11, DNSS11, HCZW13, JWC12, JRC14, MM16]. Topology-Aware [CTD^{+}16]. Tori [CS11a, Ste14]. Torus [AMVOS^{+}15, AVS^{+}14, AVS^{+}16, MWLJ15, RL13, RKZ16, TYWC10, ZR15b]. Tower [ZAG19]. Trace [LYH11]. Trace-Capable [LYH11]. Traceability [SAR^{+}11]. Traceback [YZGG16]. Traceroutes [PPB^{+}14]. Traces [RKT19, ZZX^{+}15]. Tracing [LZW^{+}15, MUMB11]. Track [WF14]. Tracking [BWV15, CRG^{+}13, LLL11, LFH^{+}16, OPAGS14, PdG13, RCN11, ST17]. Trade [ADOKM10, ASBdS16, BCC^{+}16, BS14, BG12, GCAG16, KCL^{+}16, KN11b, ZZ10]. Trade-Offs [ADOKM10, ASBdS16, BCC^{+}16, BS14, BG12, GCAG16, KCL^{+}16, KN11b, ZZ10]. Tradeoff [KCE^{+}18, MTFK19]. Tradeoffs [ABH^{+}13, CSPC12, PvdGG12]. Traffic [CHC^{+}15, CPL16, CLZ19, DHC^{+}16, GSH^{+}14, HHM11, HGL^{+}15, KGC14, PBT13, SMB^{+}15, YZH^{+}15, ZWL^{+}19]. Traffic-Aware [PBT13]. Traffics [GDY15]. TRAID [SSW12]. Train [HDYS16]. Training [CDK^{+}18, SSGB19]. Transaction [HPR16, LSG^{+}15, SSW12]. Transactional [AH13, HPR16, JW16, KCRG14, QGPZ13, RQ14, TPR16, VAN^{+}18, WGW^{+}15]. Transactions [Ano13c, Ano15a, Ano16a, Ano17a, Ano18a, Ano19a, BKP16, BPC12, CHTD19, DPO17, LDL^{+}17, Liu11, LSGZ16, Ano11g, Ano11h, Ano13d]. Transceiver [NBZP17]. Transcendent [VTW16]. Transducer [AWFV13, CYA13]. Transfer [Ano13f, CCL^{+}13, HMZ^{+}14, HL10a, LCCJ13, Tho15, WCL^{+}18]. Transform [GNSR14, HHCH11, RMC^{+}15, RBMO11, TWTT11, GTRMG18]. Transform-Based [RBMO11]. Transformation [CI16, JDA^{+}16, RBIQ15]. Transformations [LTLC12]. Transforms [Red18]. Transient [DSY^{+}15, DKK16, FEP^{+}12, MD16, SMRML17, WWT^{+}18, ZHM^{+}19]. Transiently [BDM^{+}19]. Transiently-Powered [BDM^{+}19]. Transients [CBB19]. TranSim [LZZ16]. Transition [NCD^{+}17, Pom12b, Pom13b, ST16, XKT^{+}15]. Translation [AJH15, CC11, LH16, LVF19]. TransMap [JDA^{+}16]. Transmission [CBTU14, LSL15, MFT^{+}17, RSNK17, RSN^{+}18, XLS^{+}12]. Transparent [CCW^{+}10, DHC^{+}16, JKJ^{+}10, KYZC19, LC16a, LZZ16, WFT^{+}19, ZLW^{+}17]. Transportation [DYCG16, DAPS14, HDYS16]. Transpose [IRMM^{+}16]. Transversals [WF14]. Trap [CCL^{+}18, YCL^{+}12]. Trapdoor [GPR^{+}19]. TRAST [MM16]. Traversal [DKG13, XHZC16]. Tree [AKJ^{+}13, AR17, BS14, CLS10, CCC15, CYC11, FYSK14, FEP^{+}12, GYC^{+}16, GGA^{+}17, GY15b, HH17, KK18, KLT16, LP12, LHPH15, LRY^{+}15, LSZ^{+}15, SDP^{+}15, SBI12, TXL11, WXW^{+}14, WLC^{+}15, YWW^{+}16]. Tree-Based [LP12].

Tree-Structured [CYC11]. TreeFTL [WW16]. Trees [CW15, DDNP11, LDP10, SJSLD11, TYWC10]. Tridiagonal [DALD18]. Trie [HHY11, HY12, ML16]. Trie-Based [ML16]. Tries [LYS10]. Trigger [KN12, NZ14]. Triggered [MAHD18]. Trigonometric [dLSGDR17]. Trinomials [Fan16, LMZQ17]. TRIO [BSS15]. Triple [ERRM16, FMP19, LZZ17a, RNS13, TFCY16]. Triple-Word [FMP19]. TRNG [LB15b]. TRNGs [PRM19]. Trojan [NDC^{+}13, ZZ17]. TrueNorth [TBC^{+}17]. Truncated [DRC14, GPN11]. Trust [JWL^{+}16, MM16]. Trust-Based [MM16]. Trusted [JWWZ16, MGdC^{+}18]. Trustworthy [LSW15]. TrustZone [BBA19]. Truthful [HZL^{+}16, MFG16, MFG14, ZJL^{+}16]. TSAC [WZL15]. TSP [PKC^{+}17]. TSS [CYJ^{+}10]. TSV [CCE^{+}18, EYBK15, YEY^{+}16]. TSV-Based [CCE^{+}18, EYBK15]. TSV-to-TSV [YEY^{+}16]. Tunable [Nan19]. Tuning [NL14]. Turn [TLP17]. Turnaround [EE17]. Tweakable [CMLRHS13, MLCH10]. Twin [AMVOS^{+}15, AVS^{+}16]. Two [AJH15, CHL17, LAAM11, LLW^{+}11, LLS^{+}16, NI11, NL15b, NZC11, PLZW14, Pom15c, RLRL19, Ste14, TWTT11, Tse12, WEX14, WHC^{+}15b, ZGR13, ZMRQ11]. Two-Dimensional [Pom15c, Ste14, TWTT11, WEX14]. Two-Factor [LLS^{+}16]. Two-Level [AJH15, LLW^{+}11, NL15b, ZGR13]. Two-Rail [NI11]. Two-Stage [PLZW14]. Twofold [GO10]. TXOP [LHY13]. Type [PCHS18, RKN^{+}18, WF12, ZCW18, Ima18]. Type-I [Ima18]. Types [KN11a].

Ultimate [PN16]. Ultra [BDDL18, FHW18, LT15, WWY^{+}16, XYF^{+}15]. Ultra-Low [BDDL18, FHW18, WWY^{+}16]. Ultra-Scalable [LT15, XYF^{+}15]. Unbalancing [JLC10]. Unbiased [LMB17, VMHGN18]. Unbounded [XXBL17]. Uncertainty [BIP^{+}17, LMB13, WHYS16]. Unclonable [CJSM17]. Underwater [CBVL16]. Undetectable [Pom12b]. Unequal [CMM15, DRM16, NL15b, NL16a, YW12]. Unfaithful [FNS16]. Unicast [RBRL15]. Unified [AFC10, LLC^{+}16, LJY^{+}15, LJ15, MF14, SB10]. Uniform [NTR14, ST18b, WhCCC12]. Uniprocessor [SL14b]. Uniqueness [GLH^{+}19]. Unit [BPG16, GH11, HMS^{+}12, LCL17, LN12, MKFM13, PGvdG14, WF17, WE12, ZCW18]. Unitaries [KMM16]. Unitary [Red18]. Units [FFISC13, JRJ^{+}18, KN12, Nan16, dOPSR16]. UNIVERCM [DFP^{+}13]. Universal [DZLP14, DFP^{+}13]. Unix [SF17]. Unknowns [NS13]. Unobtrusive [MUMB11]. Unreliable [BMT14, GHG^{+}14, KAH19, RCK^{+}16]. Unschedulability [ZZ19]. Unsigned [JLLH19]. Unspecified [NM10]. Unstable [YCCWC15]. Unstable-Bandwidth [YCCWC15]. Update [HY12, KE19, KCS14, LDL^{+}17]. Updates [BMS12, CLW^{+}16a, SZG^{+}18]. Uplink [CKH15]. Upon [LCX^{+}16]. Upper [Bin15, HCCG10]. Upsets [ALBP14, WNKL16]. Urban [WAK^{+}17]. Usable [MAG^{+}17]. Use [DAPS14, Fin10, MB12a]. Used [LOC^{+}16]. User [HCD^{+}16, JCM16, QQW^{+}17, RVH^{+}16, SZDL14, YCCJ15, ZJL^{+}16]. User-Friendly [SZDL14]. Users [PPB^{+}14]. Using [ARM16, ABH^{+}13, AAR19, ADP^{+}15, AKL18, AD16, AXS^{+}10, AHNT16, AMG17, AVS^{+}14, ADC11, AD13, AS14, BNP10, BD15, BFMT16, BCTV15, BJ12, BBH12, CSS13, CCH15b, CZP^{+}16, CTD^{+}16, CJA^{+}16, CW15, CW16, CLL^{+}14, CJK19, CCR^{+}17, CND^{+}18, CJ12, CH14, Cro14, CP10, CDL^{+}17, DN11, DJA14, DLC^{+}13, DCV^{+}12, DAPS14, EKA17, ERRMG15, ERRM16, EG11, GKD^{+}17, GAFN15, GO10, GZB^{+}15, GCL^{+}13, HTA10, HCSW15, HNB^{+}12, IRMM^{+}16, JSA17, JZLD10, JWWZ16, JWL^{+}16, JLLH19, JMMP16, KAH18a, KYZC19, KKT15, LP12, LW17, Lee17, LDP10, LLCC13, LCT11, LPL12, LHYZ13, MRD19, MWZ^{+}17, MSK15, MHRARG^{+}14, MRL^{+}18, MF14, ML16, NWA11, NWA12, NM10, OPZ15, OKD^{+}16, OKC13, PDXZ13, PP11, PCHS18, PO13, PPB^{+}14, PRM19, RVL^{+}14, RKR15, RCN11, RM15b, RMERM19, RS10, RD18, SIB13, ST12, SXCL14, SDP11]. Using [TW10, TAH^{+}16, TZL^{+}14, Tho15, TBC^{+}17, VED^{+}16, VAB14, VA11, WEX14, WSXZ13, WZ14, XLS^{+}12, XLC14, XDZ11, XKT^{+}15, YKK^{+}15, YFJ^{+}14, YLML15, YWM19, Yun12, ZNL18, ZCZL16, ZZX^{+}15, ZRL15, dDLM11, GV15]. Utility [KM11, KFB^{+}15, LS13, LTLC12, SMP16, ZJS14]. Utility-Based [LS13, SMP16]. Utility-Directed [LTLC12]. Utilization [AGFM11, ARS16, Bin15, CCY^{+}16, CGL^{+}18, HZX^{+}14, HLA^{+}17, JDA^{+}16, WLK15]. Utilization-Based [CGL^{+}18]. Utilizing [HK17, PWTS16].

v6 [LP12]. Validation [Ano11f, FBR^{+}12, KN12, SN16, ZMS13]. Validations [PDJ^{+}19]. Value [ASM^{+}16, FMTK19, HWZ^{+}17, IS14, JPG10]. Valued [LWK11]. Values [HK15a]. VANET [CYHL14]. VANET-Based [CYHL14]. VANETs [RBRL15]. Variability [GBD^{+}15, HLC^{+}19, MLS19, PAC^{+}12, RBG14]. Variability-Aware [GBD^{+}15, PAC^{+}12]. Variable [ASM^{+}16, BTW13, PB11, SRK^{+}17]. Variable-Length [ASM^{+}16]. Variance [LVMS18]. Variants [HNB^{+}12]. Variation [DSG^{+}19, HK16, KCRG15, KSC^{+}14, KKC15b, LSA18, ZDYZ13, ZDYZ14]. Variation-Aware [ZDYZ13, ZDYZ14]. Variation-Induced [HK16].

Variation-Tolerant [KKC15b]. Variations [DMXY14, RKR15, WTBT13]. Varying [ALZ16]. vCUDA [SCSL12]. Vector [BZ15, CNH13, CLL^{+}14, HF15, HMNN12, HN13, KN11a, KT12, PCHS16, PCHS18, Pom15b, RS17, RM15b]. Vector-Magnitude-Based [KN11a]. Vectorization [AMG17]. Vectors [Ibr16]. VEGa [SME^{+}17]. Vehicle [WLYY16]. Vehicular [SME^{+}17, YZH^{+}15, YMT13, ZGB^{+}15, ZHW^{+}16]. Verifiable [CLW^{+}16a, WCH^{+}15]. Verification [BIP^{+}17, GABK11, HCC^{+}12, LJ18, MVB10, Rus13, SKEB18, VYEB17, ZL18]. Verified [Wal19]. Verifying [Hie13, UHSA17, VYEB18]. VERsatile [DFP^{+}13, CVH^{+}13, IRMM^{+}16]. Version [CHH^{+}13, WLC^{+}15]. Version-Based [CHH^{+}13]. versus [KCE^{+}18, MMP13]. Vertex [DHW^{+}19, SG13, YL14]. Vertex- [YL14]. Vertex-Edge [DHW^{+}19]. Vertical [LLD^{+}16]. Vertically [DSPB13]. Vertices [LXK12]. Very [AISA16, SBM15]. VFI [HWZ^{+}12, KCL^{+}16]. VFI-Based [HWZ^{+}12]. VFI-Enabled [KCL^{+}16]. via [ASTU10, BCTV19, BKV12, CZS^{+}19, CLW16b, DZLP14, GM15, GRL^{+}14, HT12, JKMR11, JWWZ16, KCY18, LPL^{+}13, LLKA19, LCLC19, RLX15, RZPX19, YyHL11, YLH13, ZLN11, ZWL^{+}19]. Victim [TCYH15]. Video [CLMM11, KL16, RBRL15, YASS14]. ViPZonE [GBD^{+}15]. Virtual [CSS13, DKW15, DSY^{+}15, ECJ^{+}16, GBD^{+}15, JJK^{+}11, JK15, JAS^{+}15, JKJ^{+}10, KJL11, KP15, LTW^{+}12, LGF^{+}15, MSG14, Man16, PLP^{+}13, PP11, RSN^{+}18, SCZ^{+}16, SCSL12, VTW16, VGF16, VMB19, WGLL13, WZL15, WLLZ16, XLL^{+}14, XLJ16, YCK16, YWQX15, ZRS^{+}16, ZL18, ZJXL11, Zot10]. Virtualization [CXLX15, cCWS14, LH16, RTL^{+}18, RZPX19, SKYK16, UVL^{+}13, XH16]. Virtualization-Based [RZPX19].

Virtualized [AJH15, LZ15, WBZ^{+}15, ZL16, ZCYX15]. Virtually [AGFM11]. Virus [LLL11, OWP16, PWW^{+}11]. Vision [NVB16, TJH^{+}15, TKT16]. VLIW [GSL10, Yan14]. VLSI [BG12, BKP16, CCR^{+}17, FM16, RMC^{+}15, TWTT11]. VM [RSNK17, ZLSI17]. Vol [Ano15a, Ano16a, Ano17a, Ano18a, Ano19a]. Volatile [AK19, Cha14, CHTD19, HZX^{+}14, JAKD18, JSA17, JW16, LKBS16, LHL^{+}15a, LXW^{+}19, LZZ^{+}17b, QLH^{+}16, RD18, SJD^{+}18, SZZ^{+}19, SZL^{+}16, TAH^{+}16, WWY^{+}16, WWY^{+}18, WNKL16, XZL^{+}19, YWW^{+}16, ZFHC19]. Voltage [ACW^{+}11, AEGH19, DA12, FSGAB^{+}16, FCB^{+}19, JSH^{+}17, KkC15a, LKYC12, MHK15, NY15, NYHB16, NY19, OKC13, PdG13]. Voltage/Frequency [DA12]. Volumes [LBSK17]. Voting [CHL17, LY18]. vs [Red19]. VSPN [CYHL14]. Vulnerabilities [LDP10]. Vulnerability [BPBBL13, MTGM12, MMTM15]. Vulnerable [JAS^{+}15, WBGI19].

WaFS [WLK15]. Wait [HTA17]. Wait-Free [HTA17]. Wake [HKWC14, LRC10]. Wake-Up [LRC10]. Walks [YCW11]. Wall [SRHC12, VKS^{+}16, YWXZ12]. Wallace [WS10]. WANETs [GGL^{+}14]. Warp [OYP^{+}18, OKY^{+}19]. WASP [OYP^{+}18]. Wastewater [CCO^{+}14]. Wave [DCY^{+}13, UVG16]. Wavelength [KH14]. Wavelength-Routed [KH14]. Wavelet [HHCH11, Red11, TWTT11]. Way [CNH13, CLW^{+}15, DEE17, DY14, LLM^{+}15]. WBSP [WZLS16]. WCET [PC10, YLH13]. Weak [WZL^{+}17]. Weakly [RWC18]. Wear [CHK10, DY14]. Wear-Leveling [CHK10, DY14]. Web [MRW^{+}15, MBS^{+}12, RQ14]. WEC [CDQB15]. Weight [SWF^{+}19]. Weighted [AWFV13, GWZ^{+}10, LCL15, LCW^{+}15]. Weighted-Graph-Based [GWZ^{+}10]. Well [Gor14]. Well-Connected [Gor14]. WG [ERRMG15, ZAG19]. WG- [ERRMG15]. WG-16 [ZAG19]. Where [MOS14, PPB^{+}14]. Which [LYT^{+}16]. while [RS13]. White [SWF^{+}19]. White-Box [SWF^{+}19]. Whole [KKH^{+}14]. Wi [HWK15]. Wi-Fi [HWK15]. Wide [KCE^{+}18, SH12, SKM14, Zhe10]. Wide-Area [SH12]. Wide-Sense [Zhe10]. Width [CK15, HK15a, LAAM11]. Window [LPL10, TX16]. Windows [QPG10]. Wireless [AO11, AO12a, AEKT15, AD10, AD12, Amm14, BWCW15, BWV15, CWZ11, CJG16, CWTT13, CLR13, CSJ^{+}11, CCD12, CBTU14, DCY^{+}13, DY12, DCL^{+}11, DLL^{+}12, DVUS14, GD17, GSH^{+}14, GCD^{+}11, GYC^{+}16, GY16, GHG^{+}14, GLTC16, HXVQ15, HMZ^{+}14, HBCC13, HKWC14, HWX15, KCL^{+}16, KKT15, KLT16, KH10, LRC10, LR13, LYOB15, LY11, LWW11, LGH15, LLZ^{+}17, LDB^{+}17, LSX13, LCHC14, LP13b, LYCT10, LKLT12, LHH14b, LWY15, LCT11, LZYL13, LWF13, LGF^{+}15, MM16, MB12b, MWY^{+}16, MMH14, MMB14, QSYS16, RGK15, RS13, RNS13, SXLC15, SBH11, SPC^{+}16, SCK10, SMN^{+}17, SZS14, VBR^{+}13, WJL^{+}12, WXLL13, WJL^{+}14, WXLY15, WLYY16, XCW^{+}10, XLS^{+}12, XHZ14, XWL^{+}16a, XWL10, XLTZ11, YKK^{+}15, YASS14, ZWX12, ZMY11, ZY12, ZWD^{+}16, ZLYS15, dAJM14]. within [RV13]. without [ARM15, CCW^{+}10, DCCK18, EKJ^{+}10, HMZ^{+}14, IGLM15]. WLAN [LHY13]. WLANs [GY14]. WMNs [CWX^{+}14]. WOM [PLM16]. WOM-Code [PLM16]. Word [FMP19, LCH13, NWA11, SL10]. Word-Based [LCH13, SL10]. Word-Level [NWA11]. Work [CG18]. Work-Stealing [CG18]. Worker [CAGM14]. Workflow [CPL17, WL13, WLK15]. Workflow-Aware [WLK15]. Workflow-Based [WL13].

Workflows [GW16, HCC^{+}18]. Working [DAS14, EF12]. Workload [AQPMS15, BDL^{+}13, CMS10, CBB19, DGC^{+}15, JRJ^{+}18, KMJ^{+}11, LSA18, LXL^{+}13, LPL10, MJW^{+}14, NLP^{+}14, SKC^{+}14, WW16, WJF^{+}11, YWW^{+}16, ZL16]. Workload-Adaptive [WW16, YWW^{+}16]. Workload-Aware [JRJ^{+}18]. Workload-Based [BDL^{+}13]. Workload-Cognizant [KMJ^{+}11]. Workload-Efficient [LXL^{+}13]. Workloads [CPS^{+}10, CRJZ16, LWZ18, WSL^{+}18]. World [LLK18, WGZ^{+}15, ZLJ^{+}17, ZCW18]. Wormhole [MBD11, XWLX17]. Worms [TXL11]. Worst [BBK10, DZ10, FS10, Ibr16, RCRK13, WZLX12]. Worst-Case [BBK10, DZ10, FS10]. Write [AK19, CDQB15, CCC^{+}17, CCC^{+}18, CLZ19, CJK19, GKD^{+}17, JSC^{+}17, JSC^{+}19, KRP18, LBN14, LKK^{+}19, LWF^{+}17, LSG^{+}18, NL19, QLH^{+}16, SJD^{+}18, SQJ^{+}15, WZ14, YCW^{+}19, ZFHC19]. Write-Efficient [CDQB15, ZFHC19]. Write-History-Aware [LBN14]. Write-Reduction [CCC^{+}18]. Write-Through [KRP18]. Writes [FSL^{+}17, ZZJ^{+}19]. wrJFS [CCC^{+}18].

X [KN11b, PVKA14, XLX^{+}14, ZXX^{+}14]. X-Code-Based [XLX^{+}14, ZXX^{+}14]. X-Masking [KN11b]. XACML [LCHX11]. Xbar [ZMW^{+}19]. XHive [KJL11]. Xilinx [RURM18]. XML [TLL^{+}13]. XOR [LSXP14, ZLWZ15, vdBGLGL^{+}16]. XOR-Based [LSXP14]. XQDD [LWK11].

Yahoo [SLLG15]. YARN [CZL^{+}17]. Yield [LLHC15, SC11].

Zero [SDZ15, WCL^{+}18, ZLW^{+}17]. Zombie [LKLM15].

References

Agrawal:2019:SSN

[AAR19] A. Agrawal, A. Ankit, and K. Roy. SPARE: Spiking neural network acceleration using ROMembedded RAMs as inmemory-computation primitives. IEEE Transactions on Computers, 68(8):1190– 1200, August 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Albader:2016:EDH

[AB16] B. Albader and B. Bose. Edge disjoint Hamiltonian cycles in Gaussian networks. IEEE Transactions on Computers, 65(1):315–321, January 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anantha:2007:MRG

[ABA07] M. Anantha, B. Bose, and B. F. AlBdaiwi. Mixedradix gray codes in Lee metric. IEEE Transactions on Computers, 56(10): 1297–1307, October 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                                      http://ieeexplore.ieee.
                                                      org/stamp/stamp.jsp?tp=
                                                      &arnumber=4302703. See

comments [Jha13].

Abdelfattah:2017:DAE

[ABB17] Mohamed S. Abdelfattah, Andrew Bitar, and Vaughn Betz. Design and applications for embedded networks-on-chip on FP-GAs. IEEE Transactions on Computers, 66(6):1008– 1021, June 2017. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7676325/.

Aminifar:2016:ADR

[ABEP16] A. Aminifar, E. Bini, P. Eles, and Z. Peng. Analysis and design of real-time servers for control applications. IEEE Transactions on Computers, 65(3):834–846, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Adikari:2013:IAT

[ABH^{+}13] J. Adikari, A. Barsoum, M. A. Hasan, A. H. Namin, and C. Negre. Improved area-time tradeoffs for field multiplication using optimal normal bases. IEEE Transactions on Computers, 62 (1):193–199, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Alizadeh:2015:SFD

[ABSK15] B. Alizadeh, P. Behnam, and S. Sadeghi-Kohan. A scalable formal debugging approach with auto-correction capability based on static slicing and dynamic ranking for RTL datapath designs. IEEE Transactions on Computers, 64(6):1564– 1578, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Arnold:2011:RCL

[AC11] Mark G. Arnold and Sylvain Collange. A real/ complex logarithmic number system ALU. IEEE Transactions on Computers, 60 (2):202–213, February 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Atalla:2013:BPA

[ACGP13] Shadi Atalla, Davide Cuda, Paolo Giaccone, and Marco Pretti. Belief-propagation-assisted scheduling in inputqueued switches. IEEE Transactions on Computers, 62(10):2101–2107, October 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Amat:2016:FED

[ACM^{+}16] E. Amat, A. Calomarde, F. Moll, R. Canal, and A. Rubio. Feasibility of embedded DRAM cells on FinFET technology. IEEE Transactions on Computers, 65(4):1068–1074, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Akleylek:2012:PMC

[ACO12] Sedat Akleylek, Murat Cenk, and Ferruh Ozbudak. On the polynomial multiplication in Chebyshev form. IEEE Transactions on Computers, 61(4):584–587, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See comments [PCHS14].

Alameldeen:2011:ACD

[ACW^{+}11] Alaa R. Alameldeen, Zeshan Chishti, Chris Wilkerson, Wei Wu, and Shih-Lien Lu. Adaptive cache design to enable reliable lowvoltage operation. IEEE Transactions on Computers, 60(1):50–63, January 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ammari:2010:SCM

[AD10] H. M. Ammari and S. K.

Das. A study of k-coverage and measures of connectivity in 3D wireless sensor networks. IEEE Transactions on Computers, 59 (2):243–257, February 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.

                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5313797.

Ammari:2012:CCC

[AD12] Habib M. Ammari and Sajal K. Das. Centralized and clustered k-coverage protocols for wireless sensor networks. IEEE Transactions on Computers, 61 (1):118–133, January 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ashraf:2013:SFR

[AD13] Rizwan A. Ashraf and Ronald F. DeMara. Scalable FPGA refurbishment using netlist-driven evolutionary algorithms. IEEE Transactions on Computers, 62 (8):1526–1541, August 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Abawajy:2014:DRA

[AD14] J. H. Abawajy and M. M. Deris. Data replication approach with consistency guarantee for data grid. IEEE Transactions on Computers, 63(12):2975–2987, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Alzahrani:2016:FOD

[AD16] Ahmad Alzahrani and Ronald F. DeMara. Fast online diagnosis and recovery of reconfigurable logic fabrics using design disjunction. IEEE Transactions on Computers, 65(10):3055–3069, October 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ashby:2011:SBC

[ADC11] Thomas J. Ashby, Pedro Diaz, and Marcelo Cintra. Software-based cache coherence with hardware-assisted selective self-invalidations using bloom filters. IEEE Transactions on Computers, 60(4):472–483, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Adikari:2011:HBT

[ADI11] Jithra Adikari, Vassil S.

Dimitrov, and Laurent Imbert. Hybrid binary-ternary number system for elliptic curve cryptosystems. IEEE Transactions on Computers, 60(2):254–265, February 2011. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Adikari:2012:FHA

[ADJ12] Jithra Adikari, Vassil S.

Dimitrov, and Kimmo U. Jarvinen. A fast hardware architecture for integer to τNAF conversion for Koblitz curves. IEEE Transactions on Computers, 61(5):732–737, May 2012.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Allyson:2019:SOI

[ADJG19] F. B. Allyson, M. L. Danilo, S. M. José, and B. C. Giovanni. Sherlock N-overlap: Invasive normalization and overlap coefficient for the similarity analysis between source code. IEEE Transactions on Computers, 68(5): 740–751, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Al-Dubai:2010:TOB

[ADOKM10] A. Y. Al-Dubai, M. Ould-Khaoua, and L. M. Mackenzie. Trade-offs between latency, complexity, and load balancing with multicast algorithms. IEEE Transactions on Computers, 59 (2):159–173, February 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5184806.

Alawneh:2015:HRT

[ADP^{+}15] S. Alawneh, R. Dragt, D. Peters, C. Daley, and S. Bruneau. Hyper-real-time ice simulation and modeling using GPGPU. IEEE Transactions on Computers, 64(12):3475–3487, December 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Aminzadeh:2011:CSS

[AE11] Soheil Aminzadeh and Alireza Ejlali. A comparative study of system-level energy management methods for fault-tolerant hard realtime systems. IEEE Transactions on Computers, 60(9): 1288–1299, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5710892.

Amrouch:2019:EVO

[AEGH19] H. Amrouch, S. B. Ehsani, A. Gerstlauer, and J. Henkel. On the efficiency of voltage overscaling under temperature and aging effects. IEEE Transactions on Computers, 68(11):1647–1662, November 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Akbas:2015:LWS

[AEKT15] M. I. Akbas, M. Erol-Kantarci, and D. Turgut. Localization for wireless sensor and actor networks with meandering mobility. IEEE Transactions on Computers, 64(4):1015–1028, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Aminifar:2018:OME

[AEP18] Amir Aminifar, Petru Eles, and Zebo Peng. Optimization of message encryption for real-time applications in embedded systems. IEEE Transactions on Computers, 67(5):748–754, May 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8125122/.

Ahmed:2014:PBA

[AF14] Safayet Ahmed and Bonnie H. Ferri. Predictionbased asynchronous CPUbudget allocation for softreal-time applications. IEEE Transactions on Computers, 63(9):2343–2355, September 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Alimohammad:2010:UAA

[AFC10] A. Alimohammad, S. F. Fard, and B. F. Cockburn. A unified architecture for the accurate and high-throughput implementation of six key elementary functions. IEEE Transactions on Computers, 59 (4):449–456, April 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5313801.

Anand:2010:GRC

[AFH^{+}10] M. Anand, S. Fischmeister, Y. Hur, Jesung Kim, and Insup Lee. Generating reliable code from hybrid-systems models. IEEE Transactions on Computers, 59(9): 1281–1294, September 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5453343.

Amaru:2016:SCA

[AGCD16] Luca Amarú, Pierre-Emmanuel Gaillardon, Anupam Chattopadhyay, and Giovanni De Micheli. A sound and complete axiomatization of majority-logic. IEEE Transactions on Computers, 65(9): 2889–2895, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ansari:2011:MSU

[AGFM11] Amin Ansari, Shantanu Gupta, Shuguang Feng, and Scott Mahlke. Maximizing spare utilization by virtually reorganizing faulty cache lines. IEEE Transactions on Computers, 60 (1):35–49, January 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ahmadi:2010:LCC

[AH10] O. Ahmadi and F. R.

Henríquez. Low complexity cubing and cube root computation over F_{3}^{m} in polynomial basis. IEEE Transactions on Computers, 59(10): 1297–1308, October 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5374372.

Attiya:2013:CPS

[AH13] Hagit Attiya and Eshcar Hillel. The cost of privatization in software transactional memory. IEEE Transactions on Computers, 62(12):2531–2543, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Antelo:2012:GEI

[AHI12] Elisardo Antelo, David Hough, and Paolo Ienne. Guest Editors’ introduction: Special section on computer arithmetic. IEEE Transactions on Computers, 61 (8):1057–1058, August 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Attiya:2010:PME

[AHK10] H. Attiya, D. Hay, and I. Keslassy. Packet-mode emulation of output-queued switches. IEEE Transactions on Computers, 59(10): 1378–1391, October 2010.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5374375.

Ambusaidi:2016:BID

[AHNT16] Mohammed A. Ambusaidi, Xiangjian He, Priyadarsi Nanda, and Zhiyuan Tan. Building an intrusion detection system using a filterbased feature selection algorithm. IEEE Transactions on Computers, 65(10): 2986–2998, October 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Asadi:2016:GEI

[AISA16] H. Asadi, P. Ienne, and H. Sarbazi-Azad. Guest Editors’ introduction: Special section on emerging memory technologies in very large scale computing and storage systems. IEEE Transactions on Computers, 65 (4):1006–1009, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ahn:2015:FTL

[AJH15] J. Ahn, S. Jin, and J. Huh.

Fast two-level address translation for virtualized systems. IEEE Transactions on Computers, 64(12): 3461–3474, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Azarderakhsh:2014:NDP

[AK14] R. Azarderakhsh and K. Karabina. A new double point multiplication algorithm and its application to binary elliptic curves with endomorphisms. IEEE Transactions on Computers, 63(10): 2614–2619, October 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

AlRidhawi:2015:QBC

[AK15] Y. Al Ridhawi and A. Karmouch. QoS-based composition of service specific overlay networks. IEEE Transactions on Computers, 64(3):832–846, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Acevedo:2016:CLC

[AK16] O. Acevedo and D. Kagaris. On the computation of LFSR characteristic polynomials for builtin deterministic test pattern generation. IEEE Transactions on Computers, 65 (2):664–669, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Agarwal:2019:ILN

[AK19] S. Agarwal and H. K. Kapoor. Improving the lifetime of non-volatile cache by write restriction. IEEE Transactions on Computers, 68(9):1297–1312, September 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ahn:2013:MTO

[AKJ^{+}13] Jung-Sang Ahn, Dongwon Kang, Dawoon Jung, Jin-Soo Kim, and Seungryoul Maeng. m*-tree: An ordered index structure for NAND flash memory with adaptive page layout scheme. IEEE Transactions on Computers, 62(4):784–797, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ahn:2012:SSE

[AKKH12] Jeongseob Ahn, Daehoon Kim, Jaehong Kim, and Jaehyuk Huh. Subspace snooping: Exploiting temporal sharing stability for snoop reduction. IEEE Transactions on Computers, 61(11): 1624–1637, November 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Almurib:2014:SAD

[AKL14] Haider A. F. Almurib, T. Nandha Kumar, and Fabrizio Lombardi. Scalable application-dependent diagnosis of interconnects of SRAM-based FPGAs. IEEE Transactions on Computers, 63(6):1540–1550, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Almurib:2018:ADI

[AKL18] Haider A. F. Almurib, Thulasiraman Nandha Kumar, and Fabrizio Lombardi. Approximate DCT image compression using inexact computing. IEEE Transactions on Computers, 67(2): 149–159, February 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7990539/.

Ashraf:2018:MCP

[AKTB18] Imran Ashraf, Nader Khammassi, Mottaqiallah Taouil, and Koen Bertels. Memory and communication profiling for accelerator-based platforms. IEEE Transactions on Computers, 67(7): 934–948, July 2018. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8234629/.

Abbas:2014:EMC

[ALBP14] Syed Mohsin Abbas, Soonyoung Lee, Sanghyeon Baeg, and Sungju Park. An efficient multiple cell upsets tolerant content-addressable memory. IEEE Transactions on Computers, 63(8): 2094–2098, August 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Abid:2011:ICS

[ALW11] Z. Abid, Ming Liu, and Wei Wang. 3D integration of CMOL structures for FPGA applications. IEEE Transactions on Computers, 60(4):463–471, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Adam:2016:CPP

[ALZ16] Omer Y. Adam, Young Choon Lee, and Albert Y. Zomaya. Constructing performancepredictable clusters with performance-varying resources of clouds. IEEE Transactions on Computers, 65(9): 2709–2724, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anderson:2017:EMF

[AMG17] Andrew Anderson, Servesh Muralidharan, and David Gregg. Efficient multibyte floating point data formats using vectorization. IEEE Transactions on Computers, 66(12):2081–2096, December 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7950938/.

Ammari:2014:IES

[Amm14]  H. M. Ammari.  Inves-

tigating the energy sinkhole problem in connected k-covered wireless sensor networks. IEEE Transactions on Computers, 63(11): 2729–2742, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Aguilar-Melchor:2018:CBG

[AMR18] Carlos Aguilar-Melchor and Thomas Ricosset. CDTbased Gaussian sampling: From multi to double precision. IEEE Transactions on Computers, 67(11): 1610–1621, November 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8295226/.

Andujar-Munoz:2015:DTT

[AMVOS^{+}15] F. J. Andujar-Munoz, J. A. Villar-Ortiz, J. L. Sanchez, F. J. Alfaro, and J. Duato. N-dimensional twin torus topology. IEEE Transactions on Computers, 64(10): 2847–2861, October 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2010:RL

[Ano10a] Anonymous. 2009 reviewers list. IEEE Transactions on Computers, 59 (1):139–143, January 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5341480.

Anonymous:2010:RJI

[Ano10b] Anonymous. 7 reasons for joining the IEEE Computer Society. IEEE Transactions on Computers, 59(3): 431, March 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5398798.

Anonymous:2010:CPS

[Ano10c] Anonymous. Call for papers: Special section on dependable computer architecture. IEEE Transactions on Computers, 59(1):144, January 2010. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2010:ICSa

[Ano10d] Anonymous. IEEE Computer Society CSDA certification. IEEE Transactions on Computers, 59 (4):576, April 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5420343.

Anonymous:2010:ICSb

[Ano10e] Anonymous. IEEE Computer Society CSDP certification. IEEE Transactions on Computers, 59 (11):1584, November 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5586717.

Anonymous:2010:NOP

[Ano10f] Anonymous. New Online-Plus publication model beginning in 2010. IEEE Transactions on Computers, 59(8):1152, August 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5494720.

Anonymous:2010:SBS

[Ano10g] Anonymous. Silver Bullet Security Podcast series. IEEE Transactions on Computers, 59(3):432, March 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5398799.

Anonymous:2011:AI

[Ano11a] Anonymous. 2010 annual index. IEEE Transactions on Computers, 60(1):INDEX:1– INDEX:18, January 2011.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2011:ARL

[Ano11b] Anonymous. 2010 annual reviewers list. IEEE Transactions on Computers, 60 (1):139–143, January 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2011:CPE

[Ano11c] Anonymous. Call for papers: Energy efficient computing. IEEE Transactions on Computers, 60(10):1518, October 2011. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2011:CPN

[Ano11d] Anonymous. Call for papers:

Networks-on-chip. IEEE Transactions on Computers, 60(10):1517, October 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2011:CPSa

[Ano11e] Anonymous. Call for papers: Special section on computer arithmetic. IEEE Transactions on Computers, 60(6):910, June 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2011:CPSb

[Ano11f] Anonymous. Call for papers: System-level design and validation of heterogeneous chip multiprocessors. IEEE Transactions on Computers, 60(10):1519, October 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2011:NTCa

[Ano11g] Anonymous. New Transactions on Computers Essential Sets available. IEEE Transactions on Computers, 60(10):1520, October 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=6008542.

Anonymous:2011:NTCb

[Ano11h] Anonymous. New Transactions on Computers Essential Sets available. IEEE Transactions on Computers, 60(11):1664, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=6029372.

Anonymous:2011:SC

[Ano11i] Anonymous. Stay connected. IEEE Transactions on Computers, 60(9):1376, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5962405.

Anonymous:2012:AI

[Ano12a] Anonymous. 2011 annual index. IEEE Transactions on Computers, 61(1):WebOnly, January 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2012:RL

[Ano12b] Anonymous. 2011 reviewers list. IEEE Transactions on Computers, 61 (1):139–144, January 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2012:Ca

[Ano12c] Anonymous. Cover1. IEEE Transactions on Computers, 61(9):c1, September 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2012:Cb

[Ano12d] Anonymous. Cover2. IEEE Transactions on Computers, 61(9):c2, September 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2012:Cc

[Ano12e] Anonymous. Cover3. IEEE Transactions on Computers, 61(9):c3, September 2012.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2012:Cd

[Ano12f] Anonymous. Cover4. IEEE Transactions on Computers, 61(9):c4, September 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2013:TAI

[Ano13a] Anonymous. 2012 TC annual index. IEEE Transactions on Computers, 62(1): not in print, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2013:TRL

[Ano13b] Anonymous. 2012 TC reviewers list. IEEE Transactions on Computers, 62 (1):201–208, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2013:ITE

[Ano13c] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Computers, 62(3):628, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2013:CPN

[Ano13d] Anonymous. Call for papers for new Transactions on Emerging Topics in Computing. IEEE Transactions on Computers, 62(2):416, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2013:IOA

[Ano13e] Anonymous. IEEE open access publishing. IEEE Transactions on Computers, 62(3):627, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2013:RBA

[Ano13f]  Anonymous.  Receiving

buffer adaptation for highspeed data transfer. IEEE Transactions on Computers, 62(11):2278–2291, November 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2013:SDA

[Ano13g] Anonymous. Survivable data aggregation in multiagent network systems with hybrid faults. IEEE Transactions on Computers, 62(10): 2054–2068, October 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2014:AI

[Ano14a] Anonymous. 2013 annual index. IEEE Transactions on Computers, 63(1): not in print, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2014:RL

[Ano14b] Anonymous. 2013 reviewers list. IEEE Transactions on Computers, 63 (1):256–262, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2015:IIT

[Ano15a] Anonymous. 2014 index IEEE Transactions on Computers vol. 63. IEEE Transactions on Computers, 64(1): 1–30, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2015:RL

[Ano15b] Anonymous. 2014 reviewers list. IEEE Transactions on Computers, 64 (1):295–302, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2016:IIT

[Ano16a] Anonymous. 2015 index IEEE Transactions on Computers vol. 64. IEEE Transactions on Computers, 65(1): 1–34, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2016:RL

[Ano16b] Anonymous. 2015 reviewers list. IEEE Transactions on Computers, 65 (1):334–342, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2017:IIT

[Ano17a] Anonymous. 2016 index IEEE Transactions on Computers vol. 65. IEEE Transactions on Computers, 66(1): 1–37, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2017:RLa

[Ano17b] Anonymous. 2016 reviewers list. IEEE Transactions on Computers, 66 (1):178–182, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2018:IIT

[Ano18a] Anonymous. 2017 index IEEE Transactions on Computers vol. 66. IEEE Transactions on Computers, 67(1): 1–23, January 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2018:RL

[Ano18b] Anonymous. 2017 reviewers list. IEEE Transactions on Computers, 67 (1):144–148, January 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2019:IIT

[Ano19a] Anonymous. 2018 index IEEE Transactions on Computers vol. 67. IEEE Transactions on Computers, 68 (1):157–176, January 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2019:RL

[Ano19b] Anonymous. 2018 reviewers list. IEEE Transactions on Computers, 68 (1):152–156, January 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

AbdelSalam:2011:TET

[AO11] Hady S. AbdelSalam and Stephan Olariu. Toward efficient task management in wireless sensor networks. IEEE Transactions on Computers, 60(11): 1638–1651, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5669286.

AbdelSalam:2012:TAS

[AO12a] Hady S. AbdelSalam and Stephan Olariu. Toward adaptive sleep schedules for balancing energy consumption in wireless sensor networks. IEEE Transactions on Computers, 61(10): 1443–1458, October 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Akleylek:2012:MRR

[AO12b] Sedat Akleylek and Ferruh Ozbudak. Modified redundant representation for designing arithmetic circuits with small complexity. IEEE Transactions on Computers, 61(3):427–432, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Alomair:2014:MTM

[AP14] Basel Alomair and Radha Poovendran. E-MACs: Toward more secure and more efficient constructions of secure channels. IEEE Transactions on Computers, 63 (1):204–217, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Aswal:2012:BFD

[APP12] Abhilasha Aswal, M. Ganesh Perumal, and G. N. Srinivasa Prasanna. On basic financial decimal operations on binary machines. IEEE Transactions on Computers, 61 (8):1084–1096, August 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Al-Qawasmeh:2015:PTA

[AQPMS15] A. M. Al-Qawasmeh, S. Pasricha, A. A. Maciejewski, and H. J. Siegel. Power and thermal-aware workload allocation in heterogeneous data centers. IEEE Transactions on Computers, 64 (2):477–491, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Altun:2012:LSS

[AR12] Mustafa Altun and Marc D. Riedel. Logic synthesis for switching lattices. IEEE Transactions on Computers, 61(11):1588–1600, November 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Atighehchi:2017:OTM

[AR17] Kevin Atighehchi and Robert Rolland. Optimization of tree modes for parallel hash functions: A case study. IEEE Transactions on Computers, 66(9): 1585–1598, September 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7896557/.

Andalam:2014:PFS

[ARGT14] Sidharta Andalam, Partha S.

Roop, Alain Girault, and Claus Traulsen. A predictable framework for safetycritical embedded systems. IEEE Transactions on Computers, 63(7):1600–1612, July 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Adj:2014:SRC

[ARH14] G. Adj and F. Rodriguez-Henriquez. Square root computation over even extension fields. IEEE Transactions on Computers, 63(11): 2829–2841, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

AlTawy:2018:TCM

[ARH^{+}18] Riham AlTawy, Raghvendra Rohit, Morgan He, Kalikinkar Mandal, Gangqiang Yang, and Guang Gong. Towards a cryptographic minimal design: The sLiSCP family of permutations. IEEE Transactions on Computers, 67(9):1341–1358, September 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8305605/.

Azarderakhsh:2013:LCM

[ARM13] R. Azarderakhsh and A. Reyhani-Masoleh. Low-complexity multiplier architectures for single and hybrid-double multiplications in Gaussian normal bases. IEEE Transactions on Computers, 62 (4):744–757, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Abdulrahman:2015:NRR

[ARM15] E. A. H. Abdulrahman and A. Reyhani-Masoleh. New regular radix-8 scheme for elliptic curve scalar multiplication without precomputation. IEEE Transactions on Computers, 64 (2):438–451, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Abdulrahman:2016:HSH

[ARM16] Ebrahim A. Hasan Abdulrahman and Arash Reyhani-Masoleh. High-speed hybriddouble multiplication architectures using new serial-out bit-level Mastrovito multipliers. IEEE Transactions on Computers, 65(6):1734– 1747, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Avirneni:2016:UAP

[ARS16] N. D. P. Avirneni, P. K. Ramesh, and A. K. Somani. Utilization aware power management in reliable and aggressive chip multi processors. IEEE Transactions on Computers, 65(3):979–991, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Amin:2010:HRM

[AS10] Alaaeldin Amin and Waleed Shinwari. High-radix multiplierdividers: Theory, design, and hardware. IEEE Transactions on Computers, 59 (8):1009–1022, August 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5453337.

Avirneni:2012:LOS

[AS12] Naga Durga Prasad Avirneni and Arun K. Somani. Low overhead soft error mitigation techniques for highperformance and aggressive designs. IEEE Transactions on Computers, 61(4):488– 501, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Avirneni:2014:CPA

[AS14] Naga Durga Prasad Avirneni and Arun K. Somani. Countering power analysis attacks using reliable and aggressive designs. IEEE Transactions on Computers, 63(6):1408–1420, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Aysu:2016:PMH

[AS16] Aydin Aysu and Patrick Schaumont. Precomputation methods for hash-based signatures on energy-harvesting platforms. IEEE Transactions on Computers, 65(9): 2925–2931, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Andrade:2016:LEP

[ASBdS16] Ewerton R. Andrade, Marcos A. Simplicio, Paulo S. L. M. Barreto, and Paulo C. F. dos Santos. Lyra2: Efficient password hashing with high security against timememory trade-offs. IEEE Transactions on Computers, 65(10):3096–3108, October 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Akram:2017:DOD

[ASE17] Shoaib Akram, Jennifer B. Sartor, and Lieven Eeckhout. DEP+BURST: Online DVFS performance prediction for energy-efficient managed language execution. IEEE Transactions on Computers, 66(4):601– 615, April 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ahn:2016:FFK

[ASM^{+}16] J. Ahn, C. Seo, R. Mayuram, R. Yaseen, J. Kim, and S. Maeng. ForestDB: A fast key-value storage system for variable-length string keys. IEEE Transactions on Computers, 65 (3):902–915, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Andreev:2018:PSC

[ASS^{+}18] Artem Aleksandrovich Andreev, Arvind Sridhar, Mohamed M. Sabry, Marina Zapater, Patrick Ruch, Bruno Michel, and David Atienza. PowerCool: Simulation of cooling and powering of 3D MPSoCs with integrated flow cell arrays. IEEE Transactions on Computers, 67(1):73–85, January 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7967719/.

Akkaya:2010:DRN

[ASTU10] K. Akkaya, F. Senel, A. Thimmapuram, and S. Uludag. Distributed recovery from network partitioning in movable sensor/actor networks via controlled mobility. IEEE Transactions on Computers, 59(2):258–271, February 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5184822.

Altiparmak:2016:MMF

[AT16] N. Altiparmak and A. S. Tosun. Multithreaded maximum flow based optimal replica selection algorithm for heterogeneous storage architectures. IEEE Transactions on Computers, 65 (5):1543–1557, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Alvarez:2015:HSC

[AVG^{+}15] L. Alvarez, L. Vilanova, M. Gonzalez, X. Martorell, N. Navarro, and E. Ayguade. Hardware–software coherence protocol for the coexistence of caches and local memories. IEEE Transactions on Computers, 64 (1):152–165, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Avresky:2013:GEI

[Avr13]  D. R. Avresky.  Guest

Editors’ introduction: Special section on optimizing the cloud. IEEE Transactions on Computers, 62 (6):1058–1059, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Andujar:2014:BTU

[AVS^{+}14] F. J. Andujar, J. A. Villar, J. L. Sanchez, F. J. Alfaro, and J. Duato. Building 3D torus using low-profile expansion cards. IEEE Transactions on Computers, 63(11):2701–2715, November 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Andujar:2016:ARD

[AVS^{+}16] Francisco J. Andújar, Juan A.

Villar, José L. Sánchez, Francisco J. Alfaro, and José Duato. Adaptive routing for N-dimensional twin torus. IEEE Transactions on Computers, 65(12): 3780–3786, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Aubert:2013:OWF

[AWFV13] Louis-Marie Aubert, Roger Woods, Scott Fischaber, and Richard Veitch. Optimization of weighted finite state transducer for speech recognition. IEEE Transactions on Computers, 62 (8):1607–1615, August 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Amari:2010:PAM

[AXS^{+}10] S. V. Amari, Liudong Xing, A. Shrestha, J. Akers, and K. S. Trivedi. Performability analysis of multistate computing systems using multivalued decision diagrams. IEEE Transactions on Computers, 59(10): 1419–1433, October 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5374373.

Ahn:2016:PHC

[AYC16] J. Ahn, S. Yoo, and K. Choi. Prediction hybrid cache: An energy-efficient STT-RAM cache architecture. IEEE Transactions on Computers, 65(3):940–951, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Baikov:2017:AID

[Bai17] Nikita Baikov. Algorithm and implementation details for complementary error function. IEEE Transactions on Computers, 66 (7):1106–1118, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07792222-
abs.html.

Bartkewitz:2016:LPL

[Bar16] Timo Bartkewitz. Leakage prototype learning for profiled differential side-channel cryptanalysis. IEEE Transactions on Computers, 65 (6):1761–1774, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Benhani:2019:SAT

[BBA19] E. M. Benhani, L. Bossuet, and A. Aubert. The security of ARM trustzone in a FPGA-based SoC. IEEE Transactions on Computers, 68(8):1238–1248, August 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Biondi:2016:SAH

[BBB16] A. Biondi, G. C. Buttazzo, and M. Bertogna. Schedulability analysis of hierarchical real-time systems under shared resources. IEEE Transactions on Computers, 65(5):1593–1605, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Braojos:2017:SBH

[BBB^{+}17] Rubén Braojos, Daniele Bortolotti, Andrea Bartolini, Giovanni Ansaloni, Luca Benini, and David Atienza. A synchronizationbased hybrid-memory multicore architecture for energyefficient biomedical signal processing. IEEE Transactions on Computers, 66 (4):575–585, April 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Baruah:2012:SRT

[BBD^{+}12] Sanjoy Baruah, Vincenzo Bonifaci, Gianlorenzo D’Angelo, Haohan Li, Alberto Marchetti-Spaccamela, Nicole Megow, and Leen Stougie. Scheduling real-time mixed-criticality jobs. IEEE Transactions on Computers, 61(8):1140– 1152, August 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bremlerr-Barr:2012:SET

[BBH12] Anat Bremlerr-Barr and Danny Hendler. Spaceefficient TCAM-based classification using gray coding. IEEE Transactions on Computers, 61(1):18–30, January 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bauer:2013:TSR

[BBI^{+}13] Lars Bauer, Claus Braun, Michael E. Imhof, Michael A. Kochte, Eric Schneider, Hongyan Zhang, Jorg Henkel, and Hans-Joachim Wunderlich. Test strategies for reliable runtime reconfigurable architectures. IEEE Transactions on Computers, 62 (8):1494–1507, August 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bartlett:2010:ADL

[BBK10] M. Bartlett, I. Bate, and D. Kazakov. Accurate determination of loop iterations for worst-case execution time analysis. IEEE Transactions on Computers, 59(11): 1520–1532, November 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5432156.

Betti:2013:RTM

[BBP^{+}13] Emiliano Betti, Stanley Bak, Rodolfo Pellizzoni, Marco Caccamo, and Lui Sha. Realtime I/O management system with COTS peripherals. IEEE Transactions on Computers, 62(1):45–58, January 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Baldoni:2015:ENO

[BBPQ15] R. Baldoni, S. Bonomi, M. Platania, and L. Querzoni. Efficient notification ordering for geo-distributed pub/sub systems. IEEE Transactions on Computers, 64(10):2796–2808, October 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Byun:2014:CES

[BBVL14] Sang-Seon Byun, Ilangko Balashingham, Athanasios V. Vasilakos, and Heung-No Lee. Computation of an equilibrium in spectrum markets for cognitive radio networks. IEEE Transactions on Computers, 63(2):304– 316, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bolchini:2016:NAI

[BC16] C. Bolchini and L. Cassano. A novel approach to incremental functional diagnosis for complex electronic boards. IEEE Transactions on Computers, 65 (1):42–52, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bhati:2016:DRM

[BCC^{+}16] I. Bhati, M. Chang, Z. Chishti, S. Lu, and B. Jacob. DRAM refresh mechanisms, penalties, and trade-offs. IEEE Transactions on Computers, 65(1):108–121, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bernardi:2016:DFL

[BCD^{+}16] P. Bernardi, R. Cantoro, S. De Luca, E. Sanchez, and A. Sansonetti. Development flow for on-line core self-test of automotive microcontrollers. IEEE Transactions on Computers, 65 (3):744–754, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Barus:2016:CER

[BCK^{+}16] Arlinta C. Barus, Tsong Yueh Chen, Fei-Ching Kuo, Huai Liu, Robert Merkel, and Gregg Rothermel. A costeffective random testing method for programs with non-numeric inputs. IEEE Transactions on Computers, 65(12):3509–3523, December 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Benoit:2017:TOM

[BCL^{+}17] Anne Benoit, Aurelien Cavelan, Valentin Le Fevre, Yves Robert, and Hongyang Sun. Towards optimal multi-level checkpointing. IEEE Transactions on Computers, 66 (7):1212–1226, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07795220-
abs.html.

Baek:2014:RNT

[BCM14] S. Baek, Sangyeun Cho, and R. Melhem. Refresh now and then. IEEE Transactions on Computers, 63(12): 3114–3126, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Boukerche:2010:HAF

[BCMJ10] A. Boukerche, J. M. Correa, A. Melo, and R. P. Jacobi. A hardware accelerator for the fast retrieval of DIALIGN biological sequence alignments in linear space. IEEE Transactions on Computers, 59(6):808–821, June 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5416685.

Bruguera:2011:GEI

[BCS11] Javier Bruguera, Marius Cornea, and Debjit Das Sarma. Guest Editors’ introduction: Special section on computer arithmetic. IEEE Transactions on Computers, 60(2):145–147, February 2011. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bernardi:2014:MHT

[BCSR14] P. Bernardi, L. M. Ciganda, E. Sanchez, and M. S. Reorda. MIHST: A hardware technique for embedded microprocessor functional online self-test. IEEE Transactions on Computers, 63(11): 2760–2771, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bernasconi:2015:UFC

[BCTV15] A. Bernasconi, V. Ciriani, G. Trucco, and T. Villa. Using flexibility in P -circuits by Boolean relations. IEEE Transactions on Computers, 64(12):3605–3618, December 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bernasconi:2019:BMP

[BCTV19] A. Bernasconi, V. Ciriani, G. Trucco, and T. Villa. Boolean minimization of projected sums of products via Boolean relations. IEEE Transactions on Computers, 68(9):1269–1282, September 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bardhan:2015:PEM

[BD15] S. Bardhan and A. Daniel.

Predicting the effect of memory contention in multicore computers using analytic performance models. IEEE Transactions on Computers, 64(8):2279–2292, August 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Biondi:2018:RTA

[BDB18] Alessandro Biondi, Marco Di Natale, and Giorgio Buttazzo. Response-time analysis of engine control applications under fixedpriority scheduling. IEEE Transactions on Computers, 67(5):687–703, May 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8119938/.

Burns:2018:RMC

[BDBB18] Alan Burns, Robert I. Davis, Sanjoy Baruah, and Iain Bate. Robust mixedcriticality systems. IEEE Transactions on Computers, 67(10):1478–1491, October 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8352522/.

Bruguera:2019:GEI

[BdD19] J. D. Bruguera and F. de Dinechin. Guest Editors introduction: Special section on computer arithmetic. IEEE Transactions on Computers, 68(7): 951–952, July 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bai:2018:CSL

[BDDL18] Yu Bai, Ronald F. DeMara, Jia Di, and Mingjie Lin. Clockless spintronic logic: A robust and ultra-low power computing paradigm. IEEE Transactions on Computers, 67(5):631–645, May 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8118192/.

Beuchat:2011:FAP

[BDE^{+}11] Jean-Luc Beuchat, Jeremie Detrey, Nicolas Estibals, Eiji Okamoto, and Francisco Rodriguez-Henriquez. Fast architectures for the η_{T} pairing over smallcharacteristic supersingular elliptic curves. IEEE Transactions on Computers, 60 (2):266–281, February 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bruneo:2013:WBS

[BDL^{+}13] Dario Bruneo, Salvatore Distefano, Francesco Longo, Antonio Puliafito, and Marco Scarpa. Workload-based software rejuvenation in cloud systems. IEEE Transactions on Computers, 62 (6):1072–1085, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Berthou:2019:SLK

[BDM^{+}19] G. Berthou, T. Delizy, K. Marquet, T. Risset, and G. Salagnac. Sytare: A lightweight kernel for NVRAM-based transientlypowered systems. IEEE Transactions on Computers, 68(9):1390–1403, September 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bossuet:2016:EPA

[BDMLN16] Lilian Bossuet, Nilanjan Datta, Cuauhtemoc Mancillas-López, and Mridul Nandi. ELmD: A pipelineable authenticated encryption and its hardware implementation. IEEE Transactions on Computers, 65(11):3318– 3331, November 2016. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bertossi:2015:SPP

[BDP15] A. A. Bertossi, D. Diodati, and C. M. Pinotti. Storage placement in path networks. IEEE Transactions on Computers, 64 (4):1201–1207, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Banerjee:2019:AAS

[BEHL^{+}19] Subho Sankar Banerjee, Mohamed El-Hadedy, Jong Bin Lim, Zbigniew T. Kalbarczyk, Deming Chen, Steven S. Lumetta, and Ravishankar K. Iyer. ASAP: Accelerated short-read alignment on programmable hardware. IEEE Transactions on Computers, 68(3):331–346, March 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8490591/.

Bos:2019:ACI

[BF19] J. W. Bos and S. J. Friedberger. Arithmetic considerations for isogeny-based cryptography. IEEE Transactions on Computers, 68(7): 979–990, July 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Berger:2016:EGF

[BFMT16] Thierry P. Berger, Julien Francq, Marine Minier, and Gaël Thomas. Extended generalized Feistel networks using matrix representation to propose a new lightweight block cipher: Lilliput. IEEE Transactions on Computers, 65(7):2074–2089, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bombieri:2011:AAR

[BFP11] Nicola Bombieri, Franco Fummi, and Graziano Pravadelli. Automatic abstraction of RTL IPs into equivalent TLM descriptions. IEEE Transactions on Computers, 60(12):1730–1743, December 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5560642.

Brown:2015:SPM

[BFR^{+}15] A. D. Brown, S. B. Furber, J. S. Reeve, J. D. Garside, K. J. Dugan, L. A. Plana, and S. Temple. SpiNNakerprogramming model. IEEE Transactions on Computers, 64(6):1769–1782, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bingham:2012:MET

[BG12] Brad D. Bingham and Mark R. Greenstreet. Modeling energy-time trade-offs in VLSI computation. IEEE Transactions on Computers, 61(4):530–547, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Burgess:2019:HPA

[BGHL19] N. Burgess, C. Goodyer, C. N. Hinds, and D. R. Lutz. High-precision anchored accumulators for reproducible floating-point summation. IEEE Transactions on Computers, 68(7):967–978, July 2019. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic).

Bertran:2013:SMG

[BGM^{+}13] Ramon Bertran, Marc Gonzelez, Xavier Martorell, Nacho Navarro, and Eduard Ayguade. A systematic methodology to generate decomposable and responsive power models for CMPs. IEEE Transactions on Computers, 62 (7):1289–1302, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bianco:2013:PCC

[BGMR13] Andrea Bianco, Paolo Giaccone, Guido Masera, and Marco Ricca. Power control for crossbar-based inputqueued switches. IEEE Transactions on Computers, 62(1):74–82, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bosio:2010:CFL

[BGPV10] B. Bosio, P. Girard, S. Pravossoudovitch, and A. Virazel. A comprehensive framework for logic diagnosis of arbitrary defects. IEEE Transactions on Computers, 59 (3):289–300, March 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5342412.

Burns:2015:DFI

[BGRH15] A. Burns, M. Gutierrez, M. Aldea Rivas, and M. Gonzalez Harbour. A deadlinefloor inheritance protocol for EDF scheduled embedded real-time systems with resource sharing. IEEE Transactions on Computers, 64 (5):1241–1253, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Brisebarre:2017:ESC

[BHR17] Nicolas Brisebarre, Guillaume Hanrot, and Olivier Robert. Exponential sums and correctly-rounded functions. IEEE Transactions on Computers, 66(12): 2044–2057, December 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7891945/.

Bini:2015:QUU

[Bin15] E. Bini. The quadratic utilization upper bound for arbitrary deadline real-time tasks. IEEE Transactions on Computers, 64(2):593– 599, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bessa:2017:FNF

[BIP^{+}17] Iury Bessa, Hussama Ismail, Reinaldo Palhares, Lucas Cordeiro, and João Edgar Chaves Filho. Formal non-fragile stability verification of digital control systems with uncertainty. IEEE Transactions on Computers, 66(3):545–552, March 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Brumley:2010:CAI

[BJ10] B. B. Brumley and K. U.

Jarvinen. Conversion algorithms and implementations for Koblitz curve cryptography. IEEE Transactions on Computers, 59 (1):81–92, January 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5255226.

Bircher:2012:CSP

[BJ12] W. Lloyd Bircher and Lizy K. John. Complete system power estimation using processor performance events. IEEE Transactions on Computers, 61(4):563– 577, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bogdanov:2012:BLD

[BK12] Andrey Bogdanov and Ilya Kizhvatov. Beyond the limits of DPA: Combined sidechannel collision attacks.

IEEE Transactions on Computers, 61(8):1153–1164, August 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Balston:2013:PSC

[BKH^{+}13] Kyle Balston, Mehdi Karimibiuki, Alan J. Hu, Andre Ivanov, and Steven J. E. Wilton. Post-silicon code coverage for multiprocessor system-on-chip designs. IEEE Transactions on Computers, 62(2):242– 246, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bogdanov:2013:SDS

[BKL^{+}13] Andrey Bogdanov, Miroslav Knezevic, Gregor Leander, Deniz Toz, Kerem Varici, and Ingrid Verbauwhede. SPONGENT: The design space of lightweight cryptographic hashing. IEEE Transactions on Computers, 62(10):2041–2053, October 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bolchini:2016:GEI

[BKP16] C. Bolchini, S. Kundu, and S. Pontarelli. Guest editorial: IEEE Transactions on Computers and IEEE Transactions on Nanotechnology joint special section on defect and fault tolerance in VLSI and nanotechnology systems. IEEE Transactions on Computers, 65(3):677–678, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Benkrid:2013:GEI

[BKPMC13] Khaled Benkrid, Didier Keymeulen, Umeshkumar D. Patel, and David Merodio-Codinachs. Guest Editors’ introduction: Special section on adaptive hardware and systems. IEEE Transactions on Computers, 62 (8):1478–1480, August 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Busch:2012:ACD

[BKV12] Costas Busch, Rajgopal Kannan, and Athanasios V. Vasilakos. Approximating congestion + dilation in networks via “quality of routing” games. IEEE Transactions on Computers, 61(9): 1270–1283, September 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bellal:2019:IAA

[BLB^{+}19] R. Bellal, E. Lamini, H. Belbachir, S. Tagzout, and A. Belouchrani. Improved affine arithmetic-based precision analysis for polynomial function evaluation. IEEE Transactions on Computers, 68(5):702–712, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bakhshalipour:2018:FDD

[BLKM^{+}18] Mohammad Bakhshalipour, Pejman Lotfi-Kamran, Abbas Mazloumi, Farid Samandi, Mahmood Naderan-Tahan, Mehdi Modarressi, and Hamid Sarbazi-Azad. Fast data delivery for many-core processors. IEEE Transactions on Computers, 67(10): 1416–1429, October 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8331120/.

Brisebarre:2016:CBB

[BLMM16] Nicolas Brisebarre, Christoph Lauter, Marc Mezzarobba, and Jean-Michel Muller. Comparison between binary and decimal floating-point numbers. IEEE Transactions on Computers, 65 (7):2032–2044, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Baek:2015:SAC

[BLN^{+}15] Seungcheol Baek, Hyung Gyu Lee, C. Nicopoulos, Junghee Lee, and Jongman Kim. Size-aware cache management for compressed cache architectures. IEEE Transactions on Computers, 64 (8):2337–2352, August 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Boldo:2011:EAE

[BM11] Sylvie Boldo and Jean-Michel Muller. Exact and approximated error of the FMA. IEEE Transactions on Computers, 60(2):157– 164, February 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Baldi:2013:TDP

[BM13a] Mario Baldi and Guido Marchetto. Time-driven priority router implementation: Analysis and experiments. IEEE Transactions on Computers, 62 (5):1017–1030, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bolchini:2013:RDS

[BM13b] Cristiana Bolchini and Antonio Miele. Reliabilitydriven system-level synthesis for mixed-critical embedded systems. IEEE Transactions on Computers, 62(12):2489–2502, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Benso:2011:GEI

[BMM11] Alfredo Benso, Yiorgos Makris, and Pinaki Mazumder. Guest Editors’ introduction: Special section on chips and architectures for emerging technologies and applications. IEEE Transactions on Computers, 60(4):450– 451, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Benoit:2010:SCB

[BMP^{+}10] A. Benoit, L. Marchal, J.- F. Pineau, Y. Robert, and F. Vivien. Scheduling concurrent bag-of-tasks applications on heterogeneous platforms. IEEE Transactions on Computers, 59 (2):202–217, February 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5184819.

Bolchini:2011:NDM

[BMS11] Cristiana Bolchini, Antonio Miele, and Chiara Sandionigi. A novel design methodology for implementing reliability-aware systems on SRAM-based FPGAs. IEEE Transactions on Computers, 60(12): 1744–1758, December 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5674027.

Banerjee-Mishra:2012:CUP

[BMS12] Tania Banerjee-Mishra and Sartaj Sahni. Consistent updates for packet classifiers. IEEE Transactions on Computers, 61(9): 1284–1295, September 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bendjoudi:2014:FBF

[BMT14] Ahcene Bendjoudi, Nouredine Melab, and El-Ghazali Talbi. FTH-B&B: A faulttolerant hierarchical branch and bound for large scale unreliable environments. IEEE Transactions on Computers, 63(9):2302–2315, September 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bisi:2017:SAH

[BMZ17] Elia Bisi, Filippo Melzani, and Vittorio Zaccaria. Symbolic analysis of higherorder side channel countermeasures. IEEE Transactions on Computers, 66 (6):1099–1105, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7769194/.

Bajard:2010:SSC

[BNP10] J.-C. Bajard, C. Negre, and T. Plantard. Subquadratic space complexity binary field multiplier using double polynomial representation. IEEE Transactions on Computers, 59(12):1585–1597, December 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5487504.

Ben-Porat:2013:VNM

[BPBBL13] Udi Ben-Porat, Anat Bremler-Barr, and Hanoch Levy. Vulnerability of network mechanisms to sophisticated DDoS attacks. IEEE Transactions on Computers, 62(5):1031– 1043, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bui:2012:RTS

[BPC12] Bach Duy Bui, Rodolfo Pellizzoni, and Marco Caccamo. Real-time scheduling of concurrent transactions in multidomain ring buses. IEEE Transactions on Computers, 61(9): 1311–1324, September 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bieschewski:2016:EEM

[BPG16] Stefan Bieschewski, Joan-Manuel Parcerisa, and Antonio González. An energyefficient memory unit for clustered microarchitectures. IEEE Transactions on Computers, 65(8):2631–2637, August 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bobrek:2010:SCL

[BPT10] A. Bobrek, J. M. Paul, and D. E. Thomas. Stochastic contention level simulation for single-chip heterogeneous multiprocessors. IEEE Transactions on Computers, 59(10):1402–1418, October 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5396328.

Babecki:2016:EMC

[BQP^{+}16] Christopher Babecki, Wenchao Qian, Somnath Paul, Robert Karam, and Swarup Bhunia. An embedded memory-centric reconfigurable hardware accelerator for security applications. IEEE Transactions on Computers, 65(10):3196– 3202, October 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Balasangameshwara:2013:PDL

[BR13] Jasma Balasangameshwara and Nedunchezhian Raju.

Performance-driven load balancing with a primarybackup approach for computational grids with low communication cost and replication cost. IEEE Transactions on Computers, 62 (5):990–1003, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Brabec:2010:SRC

[Bra10] T. Brabec. Speculatively redundant continued logarithm representation. IEEE Transactions on Computers, 59(11):1441–1454, November 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5467052.

Bayrak:2015:AAP

[BRN^{+}15] A. G. Bayrak, F. Regazzoni, D. Novo, P. Brisk, F.- X. Standaert, and P. Ienne. Automatic application of power analysis countermeasures. IEEE Transactions on Computers, 64(2):329– 341, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bolchini:2010:GEI

[BS10] Cristiana Bolchini and Donatella Sciuto. Guest Editors’ introduction: Special section on system-level design of reliable architectures.

IEEE Transactions on Computers, 59(5):577–578, May 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5437559.

Bhattacherjee:2014:CAT

[BS14] Sanjay Bhattacherjee and Palash Sarkar. Concrete analysis and trade-offs for the (complete tree) layered subset difference broadcast encryption scheme. IEEE Transactions on Computers, 63(7):1709–1722, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Banerjee:2015:PEP

[BS15] T. Banerjee and S. Sahni.

Pubsub: An efficient publish/subscribe system. IEEE Transactions on Computers, 64(4):1119–1132, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bhattacherjee:2016:RCO

[BS16] Sanjay Bhattacherjee and Palash Sarkar. Reducing communication overhead of the subset difference scheme. IEEE Transactions on Computers, 65(8):2575–2587, August 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Briceno:2014:RAC

[BSM^{+}14] L. D. Briceno, H. J. Siegel, A. A. Maciejewski, Ye Hong, B. Lock, C. Panaccione, F. Wedyan, M. N. Teli, and Chen Zhang. Resource allocation in a client/server system for massive multiplayer online games. IEEE Transactions on Computers, 63(12):3127–3142, December 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Banerjee:2014:PDT

[BSS14] Tania Banerjee, Sartaj Sahni, and Gunasekaran Seetharaman. PC-DUOS+: A TCAM architecture for packet classifiers. IEEE Transactions on Computers, 63(6):1527–1540, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Banerjee:2015:PTP

[BSS15] T. Banerjee, S. Sahni, and G. Seetharaman. PC-TRIO: A power efficient TCAM architecture for packet classifiers. IEEE Transactions on Computers, 64(4):1104– 1118, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bigou:2016:BTP

[BT16] Karim Bigou and Arnaud Tisserand. Binary-ternary plus-minus modular inversion in RNS. IEEE Transactions on Computers, 65(11): 3495–3501, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Benini:2014:EGB

[BTBB14] Luca Benini, Andrea Tilli, Andrea Bartolini, and Francesco Beneventi. An effective gray-box identification procedure for multicore thermal modeling. IEEE Transactions on Computers, 63 (5):1097–1110, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bouguerra:2013:CAC

[BTW13] Mohamed-Slim Bouguerra, Denis Trystram, and Frederic Wagner. Complexity analysis of checkpoint scheduling with variable costs. IEEE Transactions on Computers, 62(6):1269– 1275, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bhuiyan:2015:DWS

[BWCW15] M. Z. A. Bhuiyan, Guojun Wang, Jiannong Cao, and Jie Wu. Deploying wireless sensor networks with fault-tolerance for structural health monitoring. IEEE Transactions on Computers, 64(2):382–395, February 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bhuiyan:2015:LAP

[BWV15] M. Z. A. Bhuiyan, Guojun Wang, and A. V. Vasilakos. Local area predictionbased mobile target tracking in wireless sensor networks. IEEE Transactions on Computers, 64(7):1968– 1982, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bouvier:2014:DFB

[BZ14] Cyril Bouvier and Paul Zimmermann. Divisionfree binary-to-decimal conversion. IEEE Transactions on Computers, 63(8): 1895–1901, August 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Beldianu:2015:PEO

[BZ15] S. F. Beldianu and S. G. Ziavras. Performance-energy optimizations for shared vector accelerators in multicores. IEEE Transactions on Computers, 64(3):805– 817, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cassidy:2012:BAL

[CA12a] Andrew S. Cassidy and Andreas G. Andreou. Beyond Amdahl’s Law: An objective function that links multiprocessor performance gains to delay and energy. IEEE Transactions on Computers, 61(8):1110–1126, August 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2012:MCC

[CA12b] Xi E. Chen and Tor M.

Aamodt. Modeling cache contention and throughput of multiprogrammed manycore processors. IEEE Transactions on Computers, 61(7): 913–927, July 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Caplan:2018:MSM

[CAbZM18] Jonah Caplan, Zaid Albayati, Haibo Zeng, and Brett H. Meyer. Mapping and scheduling mixedcriticality systems with ondemand redundancy. IEEE Transactions on Computers, 67(4):582–588, April 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8066347/.

Christoforou:2014:AMR

[CAGM14] Evgenia Christoforou, Antonio Fernandez Anta, Chryssis Georgiou, and Miguel A. Mosteiro. Algorithmic mechanisms for reliable masterworker Internet-based computing. IEEE Transactions on Computers, 63(1):179– 195, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cao:2012:RMP

[Cao12] Yongzhi Cao. Reliability of mobile processes with noisy channels. IEEE Transactions on Computers, 61(9): 1217–1230, September 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Courtois:2019:RRR

[CATB19] J. Courtois, L. Abbas-Turki, and J. Bajard. Resilience of randomized RNS arithmetic with respect to side-channel leaks of cryptographic computation. IEEE Transactions on Computers, 68(12): 1720–1730, December 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2015:TCG

[CB15] Lin Chen and Kaigui Bian. The telephone coordination game revisited: From random to deterministic algorithms. IEEE Transactions on Computers, 64(10): 2968–2980, October 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Casini:2019:HTD

[CBB19] D. Casini, A. Biondi, and G. Buttazzo. Handling transients of dynamic realtime workload under EDF scheduling. IEEE Transactions on Computers, 68 (6):820–835, June 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cotuk:2014:ITP

[CBTU14] H. Cotuk, K. Bicakci, B. Tavli, and E. Uzun. The impact of transmission power control strategies on lifetime of wireless sensor networks. IEEE Transactions on Computers, 63(11): 2866–2879, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Coutinho:2016:GOR

[CBVL16] R. W. L. Coutinho, A. Boukerche, L. F. M. Vieira, and A. A. F. Loureiro. Geographic and opportunistic routing for underwater sensor networks. IEEE Transactions on Computers, 65 (2):548–561, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2014:SSF

[CBZ14] Ho-Lin Chen, Jehoshua Bruck, and Hongchao Zhou. Synthesis of stochastic flow networks. IEEE Transactions on Computers, 63 (5):1234–1247, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chiao:2011:RNF

[CC11] Mong-Ling Chiao and Da-Wei Chang. ROSE: a novel flash translation layer for NAND flash memory based on hybrid address translation. IEEE Transactions on Computers, 60(6):753– 766, June 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2016:NCM

[CC16] C. Chen and C. Chu. A novel computational model for non-linear divisible loads on a linear network. IEEE Transactions on Computers, 65(1):53–65, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Choudhury:2014:TBC

[CCAM14] Mihir R. Choudhury, Vikas Chandra, Robert C. Aitken, and Kartik Mohanram.

Time-borrowing circuit designs and hardware prototyping for timing error resilience. IEEE Transactions on Computers, 63 (2):497–509, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2015:DOS

[CCC15] L. Chen, A. Cui, and C. Chang. Design of optimal scan tree based on compact test patterns for test time reduction. IEEE Transactions on Computers, 64(12): 3417–3429, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chang:2017:EWH

[CCC^{+}17] Che-Wei Chang, Geng-You Chen, Yi-Jung Chen, Chia-Wei Yeh, Pei Yin Eng, Ana Cheung, and Chia-Lin Yang. Exploiting write heterogeneity of morphable MLC/SLC SSDs in datacenters with service-level objectives. IEEE Transactions on Computers, 66(8): 1457–1463, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7869253/.

Chen:2018:WWR

[CCC^{+}18] Tseng-Yi Chen, Yuan-Hao Chang, Shuo-Han Chen, Chih-Ching Kuo, Ming-Chang Yang, Hsin-Wen Wei, and Wei-Kuan Shih. wr-JFS: A write-reduction journaling file system for byteaddressable NVRAM. IEEE Transactions on Computers, 67(7):1023–1038, July 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8260880/.

Cinque:2012:AGP

[CCD12] Marcello Cinque, Domenico Cotroneo, and Catello Di Martino. Automated generation of performance and dependability models for the assessment of wireless sensor networks. IEEE Transactions on Computers, 61(6):870–884, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Charif:2018:FLC

[CCE^{+}18] Amir Charif, Alexandre Coelho, Masoumeh Ebrahimi, Nader Bagherzadeh, and Nacer-Eddine Zergainoh. First-last: A cost-effective adaptive routing solution for TSV-based three-dimensional networks-on-chip. IEEE Transactions on Computers, 67(10):1430–1444, October 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8333738/.

Chen:2011:MJP

[CCH11] Kuo-Yi Chen, J. Morris Chang, and Ting-Wei Hou. Multithreading in Java: Performance and scalability on multicore systems. IEEE Transactions on Computers, 60(11):1521–1534, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5661769.

Chang:2015:BSA

[CCH^{+}15a] C. Chang, J. Cheng, T. Huang, X. Huang, D. Lee, and C. Chen. Bit-stuffing algorithms for crosstalk avoidance in high-speed switching. IEEE Transactions on Computers, 64(12):3404–3416, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2015:CDG

[CCH15b] Chun-An Chen, Guey-Yun Chang, and Sun-Yuan Hsieh. Conditional-diagnosis in graphs by using the comparison diagnosis model. IEEE Transactions on Computers, 64 (6):1622–1632, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Carmona:2010:NRB

[CCK10] J. Carmona, J. Cortadella, and M. Kishinevsky. New region-based algorithms for deriving bounded Petri nets. IEEE Transactions on Computers, 59(3):371– 384, March 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5235141.

Chang:2016:DRF

[CCK^{+}16a] Y. M. Chang, Y. H. Chang, T. W. Kuo, Y. C. Li, and H. P. Li. Disturbance relaxation for 3D flash memory. IEEE Transactions on Computers, 65(5):1467– 1483, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2016:TMR

[CCK^{+}16b] Kuan-Hsun Chen, Jian-Jia Chen, Florian Kriebel, Semeen Rehman, Muhammad Shafique, and Jörg Henkel. Task mapping for redundant multithreading in multi-cores with reliability and performance heterogeneity. IEEE Transactions on Computers, 65(11): 3441–3455, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2013:LDA

[CCL^{+}13] Yunji Chen, Tianshi Chen, Ling Li, Lei Li, Liang Yang, Menghao Su, and Weiwu Hu. LDet: Determinizing asynchronous transfer for postsilicon debugging. IEEE Transactions on Computers, 62(9): 1732–1744, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2018:EEB

[CCL^{+}18] Shuo-Han Chen, Yuan-Hao Chang, Yu-Pei Liang, Hsin-Wen Wei, and Wei-Kuan Shih. An erase efficiency boosting strategy for 3D charge trap NAND flash. IEEE Transactions on Computers, 67(9): 1246–1258, September 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8322264/.

Cope:2010:PCG

[CCLH10] B. Cope, P. Y. K. Cheung, W. Luk, and L. Howes. Performance comparison of graphics processors to reconfigurable logic: a case study. IEEE Transactions on Computers, 59(4):433–448, April 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5374368.

Carara:2014:DCS

[CCM14] Everton Alceu Carara, Ney Laert Vilar Calazans, and Fernando Gehm Moraes. Differentiated communication services for NoC-based MP-SoCs. IEEE Transactions on Computers, 63(3):595– 608, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chamberlain:2014:DSS

[CCO^{+}14] Brent C. Chamberlain, Giuseppe Carenini, Gunilla Oberg, David Poole, and Hamed Taheri. A decision support system for the design and evaluation of sustainable wastewater solutions. IEEE Transactions on Computers, 63(1):129–141, January 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Clegg:2013:SMM

[CCP^{+}13] Richard G. Clegg, Stuart Clayman, George Pavlou, Lefteris Mamatas, and Alex Galis. On the selection of management/monitoring nodes in highly dynamic networks. IEEE Transactions on Computers, 62 (6):1207–1220, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Coelho:2017:DCU

[CCR^{+}17] Diego F. G. Coelho, Renato J. Cintra, Nilanka Rajapaksha, Gihan J. Mendis, Arjuna Madanayake, and Vassil S. Dimitrov. DFT computation using Gauss– Eisenstein basis: FFT algorithms and VLSI architectures. IEEE Transactions on Computers, 66(8): 1442–1448, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7869325/.

Chen:2019:EIR

[CCRL19] K. Chen, L. Chen, P. Reviriego, and F. Lombardi. Efficient implementations of reduced precision redundancy (RPR) multiply and accumulate (MAC). IEEE Transactions on Computers, 68(5):784–790, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Carretero:2011:IEE

[CCV^{+}11] Javier Carretero, Pedro Chaparro, Xavier Vera, Jaume Abella, and Antonio Gonzalez. Implementing end-to-end register data-flow continuous self-test. IEEE Transactions on Computers, 60 (8):1194–1206, August 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2010:ALD

[CCW^{+}10] Songqing Chen, Shiping Chen, Xinyuan Wang, Zhao Zhang, and Sushil Jajodia. An application-level data transparent authentication scheme without communication overhead. IEEE Transactions on Computers, 59(7):943–954, July 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5453339.

Chiueh:2014:SFI

[cCWS14] Tzi cker Chiueh, Xin Wang, and Zhiyong Shan. Shuttle: Facilitating inter-application interactions for OS-level virtualization. IEEE Transactions on Computers, 63 (5):1220–1233, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2016:MGB

[CCY^{+}16] Tseng-Yi Chen, Yuan-Hao Chang, Ming-Chang Yang, Yun-Jhu Chen, Hsin-Wen Wei, and Wei-Kuan Shih. Multi-grained block management to enhance the space utilization of file systems on PCM storages. IEEE Transactions on Computers, 65 (6):1831–1845, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Choi:2018:CCN

[CDK^{+}18] Wonje Choi, Karthi Duraisamy, Ryan Gary Kim, Janardhan Rao Doppa, Partha Pratim Pande, Diana Marculescu, and Radu Marculescu. On-chip communication network for efficient training of deep convolutional networks on heterogeneous manycore systems. IEEE Transactions on Computers, 67(5):672– 686, May 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8119941/.

Cui:2017:HPP

[CDL^{+}17] Xiaoping Cui, Wenwen Dong, Weiqiang Liu, Earl E. Swartzlander, and Fabrizio Lombardi. High performance parallel decimal multipliers using hybrid BCD codes. IEEE Transactions on Computers, 66(12): 1994–2004, December 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7931610/.

Chai:2015:WID

[CDQB15] Y. Chai, Z. Du, X. Qin, and D. A. Bader. WEC: Improving durability of SSD cache drives by caching write-efficient data. IEEE Transactions on Computers, 64(11):3304–3316, November 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Calinescu:2018:EOT

[CFL^{+}18] Gruia Calinescu, Chenchen Fu, Minming Li, Kai Wang, and Chun Jason Xue. Energy optimal task scheduling with normally-off local memory and sleep-aware shared memory with access conflict. IEEE Transactions on Computers, 67 (8):1121–1135, August 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8290729/.

Cheshmikhani:2019:LRP

[CFMA19] Elham Cheshmikhani, Hamed Farbeh, Seyed Ghassem Miremadi, and Hossein Asadi. TA-LRW: A replacement policy for error rate reduction in STT-MRAM caches. IEEE Transactions on Computers, 68 (3):455–470, March 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8489928/.

Casteigts:2014:MTL

[CFMS14] Arnaud Casteigts, Paola Flocchini, Bernard Mans, and Nicola Santoro. Measuring temporal lags in delaytolerant networks. IEEE Transactions on Computers, 63(2):397–410, February 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cano:2014:ERH

[CFR^{+}14] Jose Cano, Jose Flich, Antoni Roca, Jose Duato, Marcello Coppola, and Riccardo Locatelli. Efficient routing in heterogeneous SoC designs with small implementation overhead. IEEE Transactions on Computers, 63(3):557–569, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chung:2014:DSG

[CFW14] Sung Hoon Chung, Terry L.

Friesz, and Robert D. Weaver. Dynamic sustainability games for renewable resources — a computational approach. IEEE Transactions on Computers, 63 (1):155–166, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2018:CLA

[CG18] Quan Chen and Minyi Guo.

Contention and localityaware work-stealing for iterative applications in multisocket computers. IEEE Transactions on Computers, 67(6):784–798, June 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8214252/.

Choi:2010:PCP

[CGJ^{+}10] Jeonghwan Choi, S. Govindan, Jinkyu Jeong, B. Urgaonkar, and A. Sivasubramaniam. Power consumption prediction and power-aware packing in consolidated environments. IEEE Transactions on Computers, 59(12): 1640–1654, December 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5453350.

Chu:2013:SAS

[CGL^{+}13] Rui Chu, Lin Gu, Yunhao Liu, Mo Li, and Xicheng Lu. SenSmart: Adaptive stack management for multitasking sensor networks. IEEE Transactions on Computers, 62(1):137–150, January 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2018:UBS

[CGL^{+}18] Gang Chen, Nan Guan, Di Liu, Qingqiang He, Kai Huang, Todor Stefanov, and Wang Yi. Utilization-based scheduling of flexible mixedcriticality real-time tasks.

IEEE Transactions on Computers, 67(4):543–558, April 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8068215/.

Chen:2015:SPC

[CGT^{+}15] Tianshi Chen, Qi Guo, O. Temam, Yue Wu, Yungang Bao, Zhiwei Xu, and Yunji Chen. Statistical performance comparisons of computers. IEEE Transactions on Computers, 64 (5):1442–1455, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2011:DCC

[CH11] Chun-An Chen and Sun-Yuan Hsieh. (t, k)-diagnosis for component-composition graphs under the MM* model. IEEE Transactions on Computers, 60(12): 1704–1717, December 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5601693.

Chen:2013:CCG

[CH13] Chun-An Chen and Sun-Yuan Hsieh. Componentcomposition graphs: (t, k)- diagnosability and its application. IEEE Transactions on Computers, 62 (6):1097–1110, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Creech:2014:SAH

[CH14] Gideon Creech and Jiankun Hu. A semantic approach to host-based intrusion detection systems using contiguous and discontiguous system call patterns. IEEE Transactions on Computers, 63(4):807–819, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chang:2010:DRN

[Cha10a] Guey-Yun Chang. (t, k)- diagnosability for regular networks. IEEE Transactions on Computers, 59(9): 1153–1157, September 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5396325.

Chang:2010:HAN

[Cha10b] Li-Pin Chang. A hybrid approach to NAND-flash-based solid-state disks. IEEE Transactions on Computers, 59(10):1337–1349, October 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5383350.

Chang:2014:BTM

[Cha14] Yuan-Hao Chang. Booting time minimization for realtime embedded systems with non-volatile memory. IEEE Transactions on Computers, 63(4):847–859, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chang:2015:RAB

[CHC^{+}15] En-Jui Chang, Hsien-Kai Hsin, Chih-Hao Chao, Shu-Yen Lin, and An-Yeu Wu. Regional ACO-based cascaded adaptive routing for traffic balancing in meshbased network-on-chip systems. IEEE Transactions on Computers, 64(3):868– 875, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2012:IDF

[CHCK12] Dongdong Chen, Liu Han, Younhee Choi, and Seok-Bum Ko. Improved decimal floating-point logarithmic converter based on selection by rounding. IEEE Transactions on Computers, 61(5):607–621, May 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chang:2013:REF

[CHH^{+}13] Yuan-Hao Chang, Po-Chun Huang, Pei-Han Hsu, Lue-Jane Lee, Tei-Wei Kuo, and David Hung-Chang Du. Reliability enhancement of flashmemory storage systems: An efficient version-based design. IEEE Transactions on Computers, 62(12):2503– 2515, 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chang:2010:IFW

[CHK10] Yuan-Hao Chang, Jen-Wei Hsieh, and Tei-Wei Kuo. Improving flash wear-leveling by proactively moving static data. IEEE Transactions on Computers, 59(1):53– 65, January 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5255228.

Chen:2017:TAV

[CHL17] Ke Chen, Jie Han, and Fabrizio Lombardi. Two approximate voting schemes for reliable computing. IEEE Transactions on Computers, 66(7):1227–1239, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07819499-
abs.html.

Chen:2016:DAR

[CHLL16] Linbin Chen, Jie Han, Weiqiang Liu, and Fabrizio Lombardi. On the design of approximate restoring dividers for error-tolerant applications. IEEE Transactions on Computers, 65 (8):2522–2533, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2014:NNC

[CHLT14] Henry C. H. Chen, Yuchong Hu, Patrick P. C. Lee, and Yang Tang. NC-Cloud: A network-coding-based storage system in a cloud-of-clouds. IEEE Transactions on Computers, 63(1):31–44, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cenk:2014:ESS

[CHN14] Murat Cenk, M. Anwar Hasan, and Christophe Negre. Efficient subquadratic space complexity binary polynomial multipliers based on block recombination. IEEE Transactions on Computers, 63(9): 2273–2287, September 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chang:2019:GEI

[CHTD19] Y. Chang, J. Hu, M. B.

Tahoori, and R. F. De-Mara. Guest editorial: IEEE Transactions on Computers special section on emerging non-volatile memory technologies: From devices to architectures and systems. IEEE Transactions on Computers, 68(8):1111–1113, August 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Coleman:2016:LCT

[CI16] J. N. Coleman and R. Che Ismail. LNS with cotransformation competes with floating-point. IEEE Transactions on Computers, 65(1):136–146, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cilardo:2011:EPT

[Cil11] Alessandro Cilardo. Exploring the potential of threshold logic for cryptographyrelated operations. IEEE Transactions on Computers, 60(4):452–462, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cilardo:2013:FPP

[Cil13] Alessandro Cilardo. Fast parallel GF(2^{m}) polynomial multiplication for all degrees. IEEE Transactions on Computers, 62(5):929– 943, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cosma:2012:ASC

[CJ12] Georgina Cosma and Mike Joy. An approach to sourcecode plagiarism detection and investigation using latent semantic analysis. IEEE Transactions on Computers, 61(3):379–394, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheng:2013:EHM

[CJ13] Chi Cheng and Tao Jiang.

An efficient homomorphic MAC with small key size for authentication in network coding. IEEE Transactions on Computers, 62(10): 2096–2100, October 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See comments [LCLL15, Kim15].

Chen:2016:ULC

[CJA^{+}16] Hsing-Min Chen, Supreet Jeloka, Akhil Arunkumar, David Blaauw, Carole-Jean Wu, Trevor Mudge, and Chaitali Chakrabarti. Using low cost erasure and error correction schemes to improve reliability of commodity DRAM systems. IEEE Transactions on Computers, 65(12):3766–3779, December 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2016:SFA

[CJG16] H. Chen, H. Jin, and L. Guo. Sink-free audio-on-demand over wireless sensor networks. IEEE Transactions on Computers, 65(5):1606– 1618, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cui:2015:CCM

[CJK15] Ting Cui, Chenhui Jin, and Zhiyin Kong. On compact Cauchy matrices for substitution-permutation networks. IEEE Transactions on Computers, 64 (7):2098–2102, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Choi:2019:DPM

[CJK19] J. Choi, J. Jang, and L. Kim. DC-PCM: Mitigating PCM write disturbance with low performance overhead by using detection cells. IEEE Transactions on Computers, 68(12):1741–1754, December 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chakraborty:2017:BDD

[CJSM17] Rajat Subhra Chakraborty, Ratan Rahul Jeldi, Indrasish Saha, and Jimson Mathew. Binary decision diagram assisted modeling of FPGA-based physically unclonable function by genetic programming. IEEE Transactions on Computers, 66(6):971–981, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7553573/.

Chang:2011:MSR

[CK11] Yuan-Hao Chang and Tei-Wei Kuo. A management strategy for the reliability and performance improvement of MLC-based flash-memory storage systems. IEEE Transactions on Computers, 60(3):305– 320, March 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chung:2015:BWO

[CK15] J. Chung and L. Kim.

Bit-width optimization by divide-and-conquer for fixedpoint digital signal processing systems. IEEE Transactions on Computers, 64(11): 3091–3101, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2017:EAA

[CKD^{+}17] X. Chen, N. Khoshavi, R. F.

DeMara, J. Wang, D. Huang, W. Wen, and Y. Chen. Energy-aware adaptive restore schemes for MLC STT-RAM cache. IEEE Transactions on Computers, 66(5): 786–798, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheng:2015:DAM

[CKH15] Sheng-Wei Cheng, Ling-Chia Ku, and Pi-Cheng Hsiu. Dynamic antenna management for uplink energy efficiency on 802.11n mobile devices. IEEE Transactions on Computers, 64(10): 2767–2780, October 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chang:2014:LMS

[CKKS14] Yeim-Kuan Chang, Fang-Chen Kuo, Han-Jhen Kuo, and Cheng-Chien Su. LayeredTrees: Most specific prefix-based pipelined design for on-chip IP address lookups. IEEE Transactions on Computers, 63(12): 3039–3052, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chatterjee:2015:FAP

[CKM15] S. Chatterjee, K. Karabina, and A. Menezes. Fault attacks on pairing-based protocols revisited. IEEE Transactions on Computers, 64 (6):1707–1714, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chiarandini:2014:IPB

[CKN14] Marco Chiarandini, Niels H.

Kjeldsen, and Napoleao Nepomuceno. Integrated planning of biomass inventory and energy production. IEEE Transactions on Computers, 63(1):102–114, January 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2010:PCP

[CL10] Gang Chen and Feng Liu.

Proofs of correctness and properties of integer adder circuits. IEEE Transactions on Computers, 59 (1):134–136, January 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5255231.

Chen:2012:NDL

[CL12] Y-Chuang Chen and Shun-Fu Liu. A note on diagnosability of large fault sets on star graphs. IEEE Transactions on Computers, 61(6):911–912, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cui:2016:MPP

[CLC^{+}16] X. Cui, W. Liu, X. Chen, E. E. Swartzlander, and F. Lombardi. A modified partial product generator for redundant binary multipliers. IEEE Transactions on Computers, 65 (4):1165–1171, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2019:ECN

[CLCS19] Y. Chen, Y. Lu, P. Chen, and J. Shu. Efficient and consistent NVMM cache for SSD-based file system. IEEE Transactions on Computers, 68(8):1147–1158, August 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chiou:2014:LLD

[CLL^{+}14] Che Wun Chiou, Wen-Yo Lee, Chiou-Yng Lee, Jim-Min Lin, Mehran Mozaffari Kermani, Reza Azarderakhsh, and Jeng-Shyang Pan. Low-latency digitserial systolic double basis multiplier over GF(2^{m}) using subquadratic Toeplitz matrix–vector product approach. IEEE Transactions on Computers, 63 (5):1169–1181, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See comments [RM15b].

Chen:2019:CEA

[CLM^{+}19] H. Chen, S. Lee, T. Mudge, C. Wu, and C. Chakrabarti. Configurable-ECC: Architecting a flexible ECC scheme to support different sized accesses in high bandwidth memory systems. IEEE Transactions on Computers, 68(5):646–659, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

CoutodaSilva:2011:EHP

[CLMM11] Ana Paula Couto da Silva, Emilio Leonardi, Marco Mellia, and Michela Meo. Exploiting heterogeneity in P2P video streaming. IEEE Transactions on Computers, 60(5):667–679, May 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2018:FOB

[CLOL18] Cen Chen, Kenli Li, Aijia Ouyang, and Keqin Li. FlinkCL: An OpenCLbased in-memory computing architecture on heterogeneous CPU–GPU clusters for big data. IEEE Transactions on Computers, 67(12): 1765–1779, December 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8362980/.

Chipara:2013:RTQ

[CLR13] Octav Chipara, Chenyang Lu, and Gruia-Catalin Roman. Real-time query scheduling for wireless sensor networks. IEEE Transactions on Computers, 62(9): 1850–1865, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chang:2010:DMS

[CLS10] Yeim-Kuan Chang, Yung-Chieh Lin, and Cheng-Chien Su. Dynamic multiway segment tree for IP lookups and the fast pipelined search engine. IEEE Transactions on Computers, 59 (4):492–506, April 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5278660.

Cao:2014:OPA

[CLS14] Junwei Cao, Keqin Li, and Ivan Stojmenovic. Optimal power allocation and load distribution for multiple heterogeneous multicore server processors across clouds and data centers. IEEE Transactions on Computers, 63 (1):45–58, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2015:FSM

[CLW^{+}15] Dan Chen, Xiaoli Li, Lizhe Wang, S. U. Khan, Juan Wang, Ke Zeng, and Chang Cai. Fast and scalable multi-way analysis of massive neural data. IEEE Transactions on Computers, 64(3):707–719, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2016:VCL

[CLW^{+}16a] Xiaofeng Chen, Jin Li, Jian Weng, Jianfeng Ma, and Wenjing Lou. Verifiable computation over large database with incremental updates. IEEE Transactions on Computers, 65(10): 3184–3195, October 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cui:2016:KAS

[CLW16b] Baojiang Cui, Zheli Liu, and Lingyu Wang. Key-Aggregate Searchable Encryption (KASE) for group data sharing via cloud storage. IEEE Transactions on Computers, 65(8):2374– 2385, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cao:2019:DRR

[CLW^{+}19] Y. Cao, C. Li, J. Wang, W. Zhang, Q. Chen, J. Leng, B. Yao, Y. Shen, and M. Guo. DR refresh: Releasing DRAM potential by enabling read accesses under refresh. IEEE Transactions on Computers, 68(11): 1584–1596, November 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2014:ASI

[CLX14] Qi Chen, Haipeng Luo, and Zhen Xiao. Automatic scaling of Internet applications for cloud computing services. IEEE Transactions on Computers, 63 (5):1111–1123, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2019:RFM

[CLZ19] Xubin Chen, Yin Li, and Tong Zhang. Reducing flash memory write traffic by exploiting a few MBs of capacitor-powered write buffer inside solid-state drives (SSDs). IEEE Transactions on Computers, 68 (3):426–439, March 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8470129/.

Chen:2011:PLT

[CM11] Mingsong Chen and Prabhat Mishra. Property learning techniques for efficient generation of directed tests. IEEE Transactions on Computers, 60(6):852–864, June 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Camarero:2013:NTM

[CMB13] Cristobal Camarero, Carmen Martinez, and Ramon Beivide. L-networks: A topological model for regular 2D interconnection networks. IEEE Transactions on Computers, 62(7):1362– 1375, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chakraborty:2013:EHI

[CMLRHS13] Debrup Chakraborty, Cuauhtemoc Mancillas-Lopez, Francisco Rodriguez-Henriquez, and Palash Sarkar. Efficient hardware implementations of BRW polynomials and tweakable enciphering schemes. IEEE Transactions on Computers, 62 (2):279–294, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chakraborty:2015:SSC

[CMLS15] D. Chakraborty, C. Mancillas-Lopez, and P. Sarkar. STES: A stream cipher based low cost scheme for securing stored data. IEEE Transactions on Computers, 64(9): 2691–2707, September 2015.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Condo:2015:UEP

[CMM15] C. Condo, G. Masera, and P. Montuschi. Unequal error protection of memories in LDPC decoders. IEEE Transactions on Computers, 64(10):2981–2993, October 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cao:2016:OMA

[CMO^{+}16] Xiaolin Cao, Ciara Moore, Máire O’Neill, Elizabeth O’Sullivan, and Neil Hanley. Optimised multiplication architectures for accelerating fully homomorphic encryption. IEEE Transactions on Computers, 65(9): 2794–2806, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chatterjee:2017:IPB

[CMRH17] Sanjit Chatterjee, Alfred Menezes, and Francisco Rodríguez-Henríquez. On instantiating pairing-based protocols with elliptic curves of embedding degree one. IEEE Transactions on Computers, 66(6):1061–1070, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/
document/7762073/.

Casale:2010:MDS

[CMS10] G. Casale, Ningfang Mi, and E. Smirni. Model-driven system capacity planning under workload burstiness. IEEE Transactions on Computers, 59(1):66–80, January 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5255229.

Coelho:2018:CDB

[CND^{+}18] Diego F. G. Coelho, Sushmabhargavi Nimmalapalli, Vassil S. Dimitrov, Arjuna Madanayake, Renato J. Cintra, and Arnaud Tisserand. Computation of 2D 8 × 8 DCT based on the Loeffler factorization using algebraic integer encoding. IEEE Transactions on Computers, 67(12): 1692–1702, December 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8360454/.

Cenk:2013:ITW

[CNH13] Murat Cenk, Christophe Negre, and M. Anwar Hasan. Improved three-way split formulas for binary polynomial and Toeplitz matrix vector products. IEEE Transactions on Computers, 62 (7):1345–1361, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2014:PHA

[CNJ14] Jian Chen, Arun Arvind Nair, and Lizy K. John. Predictive heterogeneity-aware application scheduling for chip multiprocessors. IEEE Transactions on Computers, 63(2):435–447, February 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Choi:2018:TRR

[COLK18] Inhyuk Choi, Hyunggoy Oh, Young-Woo Lee, and Sungho Kang. Test resource reused debug scheme to reduce the post-silicon debug cost. IEEE Transactions on Computers, 67(12): 1835–1839, December 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8359333/.

Cucinotta:2010:QCP

[CP10] T. Cucinotta and L. Palopoli. QoS control for pipelines of tasks using multiple resources. IEEE Transactions on Computers, 59 (3):416–430, March 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5184818.

Chen:2016:TAO

[CPL16] Wuhui Chen, Incheon Paik, and Zhenni Li. Tology-aware optimal data placement algorithm for network traffic optimization. IEEE Transactions on Computers, 65 (8):2603–2617, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2017:CAS

[CPL17] Wuhui Chen, Incheon Paik, and Zhenni Li. Cost-aware streaming workflow allocation on geo-distributed data centers. IEEE Transactions on Computers, 66(2): 256–271, February 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Clark:2016:EMR

[CPRH16] L. T. Clark, D. W. Patterson, C. Ramamurthy, and K. E. Holbert. An embedded microprocessor radiation hardened by microarchitecture and circuits. IEEE Transactions on Computers, 65(2):382–395, February 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cazorla:2010:PEP

[CPS^{+}10] F. J. Cazorla, A. Pajuelo, O. J. Santana, E. Fernandez, and M. Valero. On the problem of evaluating the performance of multiprogrammed workloads. IEEE Transactions on Computers, 59(12): 1722–1728, December 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5432159.

Chetto:2014:NES

[CQ14] Maryline Chetto and Audrey Queudet. A note on EDF scheduling for realtime energy harvesting systems. IEEE Transactions on Computers, 63(4):1037– 1040, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2015:DBL

[CQW^{+}15] Renhai Chen, Zhiwei Qin, Yi Wang, Duo Liu, Zili Shao, and Yong Guan. On-demand block-level address mapping in large-scale NAND flash storage systems. IEEE Transactions on Computers, 64(6):1729–1741, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cuesta:2013:IED

[CRG^{+}13] Blas Cuesta, Alberto Ros, Maria E. Gomez, Antonio Robles, and Jose Duato. Increasing the effectiveness of directory caches by avoiding the tracking of noncoherent memory blocks. IEEE Transactions on Computers, 62(3):482–495, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheng:2016:EPA

[CRJZ16] D. Cheng, J. Rao, C. Jiang, and X. Zhou. Elastic poweraware resource provisioning of heterogeneous workloads in self-sustainable datacenters. IEEE Transactions on Computers, 65(2): 508–521, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cohen:2010:SAN

[CRK10] I. Cohen, O. Rottenstreich, and I. Keslassy. Statistical approach to networkson-chip. IEEE Transactions on Computers, 59(6): 748–761, June 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5416678.

Crowley:2014:UEP

[Cro14] Mark Crowley. Using equilibrium policy gradients for spatiotemporal planning in forest ecosystem management. IEEE Transactions on Computers, 63(1):142– 154, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2011:EMT

[CS11a] Yawen Chen and Hong Shen. Embedding meshes and tori on double-loop networks of the same size. IEEE Transactions on Computers, 60 (8):1157–1168, August 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2011:PIM

[CS11b] Zhimin Chen and Patrick Schaumont. A parallel implementation of Montgomery multiplication on multicore systems: Algorithm, analysis, and prototype. IEEE Transactions on Computers, 60(12):1692–1703, December 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5669278.

Chen:2015:MPF

[CS15] Kang Chen and Haiying Shen. Maximizing P2P file access availability in mobile ad hoc networks though replication for efficient file sharing. IEEE Transactions on Computers, 64 (4):1029–1042, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2013:GEE

[CSCW13] Po-Yuan Chen, Chin-Lung Su, Chao-Hsun Chen, and Cheng-Wen Wu. Generalization of an enhanced ECC methodology for low power PSRAM. IEEE Transactions on Computers, 62 (7):1318–1331, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chou:2011:DEF

[CSJ^{+}11] Chih-Hsun Chou, Kuo-Feng Ssu, Hewijin Christine Jiau, Wei-Tong Wang, and Chao Wang. A dead-end free topology maintenance protocol for geographic forwarding in wireless sensor networks. IEEE Transactions on Computers, 60(11): 1610–1621, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5601700.

Cardosa:2012:EST

[CSPC12] Michael Cardosa, Aameek Singh, Himabindu Pucha, and Abhishek Chandra. Exploiting spatio-temporal tradeoffs for energy-aware MapReduce in the Cloud.

IEEE Transactions on Computers, 61(12):1737–1751, December 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2013:UVS

[CSS13] Zhimin Chen, Ambuj Sinha, and Patrick Schaumont. Using virtual secure circuit to protect embedded software from side-channel attacks. IEEE Transactions on Computers, 62(1):124– 136, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2015:RDM

[CSW^{+}15] Ching-Yi Chen, Hsiu-Chuan Shih, Cheng-Wen Wu, Chih-He Lin, Pi-Feng Chiu, Shyh-Shyuan Sheu, and F. T. Chen. RRAM defect modeling and failure analysis based on march test and a novel squeeze-search scheme. IEEE Transactions on Computers, 64(1):1–11, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cederman:2013:SLF

[CT13] Daniel Cederman and Philippas Tsigas. Supporting lockfree composition of concurrent data objects: Moving data between containers. IEEE Transactions on Computers, 62(9):1866– 1878, September 2013. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2016:RSE

[CTD^{+}16] Juan Chen, Yuhua Tang, Yong Dong, Jingling Xue, Zhiyuan Wang, and Wenhao Zhou. Reducing static energy in supercomputer interconnection networks using topology-aware partitioning. IEEE Transactions on Computers, 65(8): 2588–2602, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chang:2014:ECE

[CTH14] Nai-Wen Chang, Cheng-Yen Tsai, and Sun-Yuan Hsieh. On 3-extra connectivity and 3-extra edge connectivity of folded hypercubes. IEEE Transactions on Computers, 63(6):1594–1600, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2017:TAS

[CTL^{+}17] Jiageng Chen, Jesen Teh, Zhe Liu, Chunhua Su, Azman Samsudin, and Yang Xiang. Towards accurate statistical analysis of security margins: New searching strategies for differential attacks. IEEE Transactions on Computers, 66(10): 1763–1777, October 2017.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7914659/.

Chuang:2013:PSD

[CTS13] Yue-Ru Chuang, Hsueh-Wen Tseng, and Shiann-Tsong Sheu. A performance study of Discrete-Error-Checking Scheme (DECS) with the optimal division locations for IEEE 802.16-based multihop networks. IEEE Transactions on Computers, 62(12): 2354–2365, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2018:ROM

[CvdBC18] Kuan-Hsun Chen, Georg von der Brüggen, and Jian-Jia Chen. Reliability optimization on multi-core systems with multi-tasking and redundant multi-threading. IEEE Transactions on Computers, 67(4):484–497, April 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8094023/.

Causapruno:2015:ISA

[CVGZ15] G. Causapruno, M. Vacca, M. Graziano, and M. Zamboni. Interleaving in systolic arrays: A throughput breakthrough. IEEE Transactions on Computers, 64(7):1940–1953, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2013:THE

[CVH^{+}13] Dehao Chen, Neil Vachharajani, Robert Hundt, Xinliang Li, Stephane Eranian, Wenguang Chen, and Weimin Zheng. Taming hardware event samples for precise and versatile feedback directed optimizations. IEEE Transactions on Computers, 62 (2):376–389, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Carretero:2010:MOT

[CVMA10] J. Carretero, X. Vera, P. C.

Monferrer, and J. Abella. Microarchitectural online testing for failure detection in memory order buffers. IEEE Transactions on Computers, 59(5):623–637, May 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5255233.

Candel:2019:EMC

[CVPS19] F. Candel, A. Valero, S. Petit, and J. Sahuquillo. Efficient management of cache accesses to boost GPGPU memory subsystem performance. IEEE Transactions on Computers, 68(10):

1442–1454, October 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chang:2010:GCF

[CW10] Yang Chang and Andy Wellings. Garbage collection for flexible hard realtime systems. IEEE Transactions on Computers, 59 (8):1063–1075, August 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5383349.

Cheng:2015:PCU

[CW15] Yu-Chieh Cheng and Pi-Chung Wang. Packet classification using dynamically generated decision trees. IEEE Transactions on Computers, 64(2):582– 586, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheng:2016:SMM

[CW16] Yu-Chieh Cheng and Pi-Chung Wang. Scalable multi-match packet classification using TCAM and SRAM. IEEE Transactions on Computers, 65(7):2257– 2269, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cui:2015:RBS

[CWCS15] Yan Cui, Yingxin Wang, Yu Chen, and Yuanchun Shi. Requester-based spin lock: A scalable and energy efficient locking scheme on multicore systems. IEEE Transactions on Computers, 64 (1):166–179, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Carvajal:2014:ECA

[CWF14] Gonzalo Carvajal, Chun Wah Wu, and Sebastian Fischmeister. Evaluation of communication architectures for switched real-time Ethernet. IEEE Transactions on Computers, 63(1):218– 229, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2017:HDS

[CWL^{+}17] Renhai Chen, Yi Wang, Duo Liu, Zili Shao, and Song Jiang. Heating dispersal for self-healing NAND flash memory. IEEE Transactions on Computers, 66 (2):361–367, February 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheng:2013:RGL

[CWTT13] Yuan-Po Cheng, Chia-Yi Wu, Yao-Jen Tang, and Ming-Jer Tsai. Retrievalguaranteed location-aware information brokerage scheme in 3D wireless ad hoc networks. IEEE Transactions on Computers, 62(4):798– 812, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cao:2014:CCA

[CWX^{+}14] Jiannong Cao, Hejun Wu, Wenzheng Xu, Weigang Wu, and Xiaola Lin. CACC: A cooperative approach to cache consistency in WMNs. IEEE Transactions on Computers, 63(4):860–873, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chu:2013:ECR

[CWY13] Shan Chu, Xin Wang, and Yuanyuan Yang. Exploiting cooperative relay for high performance communications in MIMO ad hoc networks. IEEE Transactions on Computers, 62(4):716– 729, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2011:FFT

[CWZ11] Tingting Chen, Fan Wu, and Sheng Zhong. FITS: a finite-time reputation system for cooperation in wireless ad hoc networks. IEEE Transactions on Computers, 60(7):1045–1056, July 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2013:DPN

[CWZ13] Tingting Chen, Fan Wu, and Sheng Zhong. On designing protocols for noncooperative, multiradio channel assignment in multiple collision domains. IEEE Transactions on Computers, 62(7):1403–1416, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cai:2013:DTC

[CWZC13] Zhiping Cai, Zhijun Wang, Kai Zheng, and Jiannong Cao. A distributed TCAM coprocessor architecture for integrated longest prefix matching, policy filtering, and content filtering. IEEE Transactions on Computers, 62(3):417–427, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2016:MCM

[CXLL16] Zhiguang Chen, Nong Xiao, Yutong Lu, and Fang Liu. Me-CLOCK: A memoryefficient framework to implement replacement policies for large caches. IEEE Transactions on Computers, 65 (8):2665–2671, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2015:LVS

[CXLX15] Wenzhi Chen, Lei Xu, Guoxi Li, and Yang Xiang. A lightweight virtualization solution for Android devices. IEEE Transactions on Computers, 64(10):2741–2751, October 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2016:SCS

[CXYC16] Fei Chen, Tao Xiang, Yuanyuan Yang, and Sherman S. M. Chow. Secure cloud storage meets with secure network coding. IEEE Transactions on Computers, 65(6):1936–1948, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cesare:2013:MEE

[CXZ13] Silvio Cesare, Yang Xiang, and Wanlei Zhou. Malwise — an effective and efficient classification system for packed and polymorphic malware. IEEE Transactions on Computers, 62 (6):1193–1206, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cho:2013:AGT

[CYA13] Hansu Cho, Lochi Yu, and Samar Abdi. Automatic generation of transducer models for bus-based MPSoC design. IEEE Transactions on Computers, 62(2):211–224, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chiu:2011:OSP

[CYC11] Ge-Ming Chiu, Li-Hsing Yen, and Tai-Lin Chin. Optimal storage placement for tree-structured networks with heterogeneous channel costs. IEEE Transactions on Computers, 60(10): 1431–1444, October 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5661768.

Chen:2016:PSA

[CYC^{+}16] D. D. Chen, G. X. Yao, R. C. Cheung, D. Pao, and C. K. Koc. Parameter space for the architecture of FFTbased Montgomery modular multiplication. IEEE Transactions on Computers, 65 (1):147–160, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2011:TTA

[CYCC11] Yi-Jung Chen, Chia-Lin Yang, Jaw-Wei Chi, and Jian-Jia Chen. TACLC: Timing-aware cache leakage control for hard real-time systems. IEEE Transactions on Computers, 60(6):767– 782, June 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2014:ACE

[CYHC14] Yean-Ru Chen, Jia-Jen Yeh, Pao-Ann Hsiung, and Sao-Jie Chen. Accelerating coverage estimation through partial model checking. IEEE Transactions on Computers, 63(7):1613–1625, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chim:2014:VVB

[CYHL14] T. W. Chim, S. M. Yiu, Lucas C. K. Hui, and Victor O. K. Li. VSPN: VANETbased secure and privacypreserving navigation. IEEE Transactions on Computers, 63(2):510–524, February 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2010:TET

[CYJ^{+}10] Hanhua Chen, Jun Yan, Hai Jin, Yunhao Liu, and L. M. Ni. TSS: Efficient term set search in large peer-to-peer textual collections. IEEE Transactions on Computers, 59(7):969–980, July 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5453340.

Chung:2014:AIT

[CYL^{+}14] Eui-Young Chung, Sungjoo Yoo, Hyuk-Jun Lee, Kwanhu Bang, Dong gun Kim, and Sang-Hoon Park. An adaptive idle-time exploiting method for low latency NAND flash-based storage devices. IEEE Transactions on Computers, 63(5):1085– 1096, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chang:2014:CSM

[CZ14] J. Morris Chang and Zhiming Zhang. A cool scheduler for multi-core systems exploiting program phases. IEEE Transactions on Computers, 63(5):1061– 1073, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2016:IEN

[CZ16] Binglong Chen and Chang-An Zhao. An improvement of the elliptic net algorithm. IEEE Transactions on Computers, 65(9): 2903–2909, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheng:2017:CPR

[CZL^{+}17] Dazhao Cheng, Xiaobo Zhou, Palden Lama, Jun Wu, and Changjun Jiang.

Cross-platform resource scheduling for Spark and MapReduce on YARN. IEEE Transactions on Computers, 66 (8):1341–1353, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7857034/.

Chen:2016:ERC

[CZP^{+}16] Mingsong Chen, Xinqian Zhang, Geguang Pu, Xin Fu, and Prabhat Mishra. Efficient resource constrained scheduling using parallel structure-aware pruning techniques. IEEE Transactions on Computers, 65 (7):2059–2073, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cantoro:2018:TRM

[CZP^{+}18] Riccardo Cantoro, Farrokh Ghani Zadegan, Marco Palena, Paolo Pasini, Erik Larsson, and Matteo Sonza Reorda. Test of reconfigurable modules in scan networks. IEEE Transactions on Computers, 67(12): 1806–1817, December 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8357895/.

Cheng:2019:TCI

[CZS^{+}19] Xi Cheng, Min Zhou, Xiaoyu Song, Ming Gu, and Jiaguang Sun. Tolerating C integer error via precision elevation. IEEE Transactions on Computers, 68 (2):270–286, February 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8443077/.

Devadas:2012:IVF

[DA12] Vinay Devadas and Hakan Aydin. On the interplay of voltage/frequency scaling and device power management for frame-based realtime embedded applications. IEEE Transactions on Computers, 61(1):31–44, January 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

deAlbuquerque:2014:MMS

[dAJM14] Edler Lins de Albuquerque, Orlando S. Junior, and Raquel A. F. Mini. MuSA: Multivariate sampling algorithm for wireless sensor networks. IEEE Transactions on Computers, 63(4):968– 978, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dieguez:2018:SLP

[DALD18] Adrián Pérez Diéguez, Margarita Amor, Jacobo Lobeiras, and Ramón Doallo. Solving large problem sizes of indexdigit algorithms on GPU: FFT and tridiagonal system solvers. IEEE Transactions on Computers, 67 (1):86–101, January 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7970194/.

Dutta:2014:CIL

[DAPS14] Parikshit Dutta, Elise Arnaud, Emmanuel Prados, and Mathieu Saujot. Calibration of an integrated land-use and transportation model using maximumlikelihood estimation. IEEE Transactions on Computers, 63(1):167–178, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dargie:2015:SME

[Dar15]  W. Dargie.  A stochas-

tic model for estimating the power consumption of a processor. IEEE Transactions on Computers, 64 (5):1311–1322, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dani:2014:TSW

[DAS14] Aparna Mandke Dani, Bharadwaj Amrutur, and Y. N. Srikant. Toward a scalable working set size estimation method and its application for chip multiprocessors. IEEE Transactions on Computers, 63 (6):1567–1579, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dai:2017:ATE

[DCCK17] Wangchen Dai, Donald Donglong Chen, Ray C. C. Cheung, and Çetin Kaya Koç. Area-time efficient architecture of FFT-based Montgomery multiplication. IEEE Transactions on Computers, 66(3):375–388, March 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dai:2018:FBM

[DCCK18] Wangchen Dai, Donglong Chen, Ray C. C. Cheung, and Çetin Kaya Koç. FFT-based McLaughlin’s Montgomery exponentiation without conditional selections. IEEE Transactions on Computers, 67(9): 1301–1314, September 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8307235/.

Distler:2016:REB

[DCK16] Tobias Distler, Christian Cachin, and Rüdiger Kapitza. Resource-efficient Byzantine fault tolerance. IEEE Transactions on Computers, 65(9):

2807–2819, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dong:2011:SPF

[DCL^{+}11] Wei Dong, Chun Chen, Xue Liu, Yunhao Liu, Jiajun Bu, and Kougen Zheng. Sen-Spire OS: a predictable, flexible, and efficient operating system for wireless sensor networks. IEEE Transactions on Computers, 60(12): 1788–1801, December 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

DellAmico:2016:PPS

[DCM16] Matteo Dell’Amico, Damiano Carra, and Pietro Michiardi. PSBS: Practical size-based scheduling. IEEE Transactions on Computers, 65(7):2199–2212, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Doudalis:2012:EEM

[DCV^{+}12] Ioannis Doudalis, James Clause, Guru Venkataramani, Milos Prvulovic, and Alessandro Orso. Effective and efficient memory protection using dynamic tainting. IEEE Transactions on Computers, 61(1):87–100, January 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Deb:2013:DEE

[DCY^{+}13] Sujay Deb, Kevin Chang, Xinmin Yu, Suman Prasad Sah, Miralem Cosic, Amlan Ganguly, Partha Pratim Pande, Benjamin Belzer, and Deukhyoun Heo. Design of an energy-efficient CMOScompatible NoC architecture with millimeter-wave wireless interconnects. IEEE Transactions on Computers, 62(12):2382–2396, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

deDinechin:2011:CFP

[dDLM11] Florent de Dinechin, Christoph Lauter, and Guillaume Melquiond. Certifying the floating-point implementation of an elementary function using Gappa. IEEE Transactions on Computers, 60(2):242–253, February 2011. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

DAngelo:2014:FPM

[DDN14] Gianlorenzo D’Angelo, Gabriele Di Stefano, and Alfredo Navarra. Flow problems in multi-interface networks. IEEE Transactions on Computers, 63(2):361– 374, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

DAngelo:2011:RRT

[DDNP11] Gianlorenzo D’Angelo, Gabriele Di Stefano, Alfredo Navarra, and Cristina M. Pinotti. Recoverable robust timetables: An algorithmic approach on trees. IEEE Transactions on Computers, 60(3):433–446, March 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dutt:2019:AMO

[DDNT19] Sunil Dutt, Satyabrata Dash, Sukumar Nandi, and Gaurav Trivedi. Analysis, modeling and optimization of equal segment based approximate adders. IEEE Transactions on Computers, 68(3):314–330, March 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8468108/.

DePestel:2017:LBE

[DEE17] Sander De Pestel, Stijn Eyerman, and Lieven Eeckhout. Linear branch entropy: Characterizing and optimizing branch behavior in a micro-architecture independent way. IEEE Transactions on Computers, 66 (3):458–472, March 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

DiGuglielmo:2013:UUV

[DFP^{+}13] Luigi Di Guglielmo, Franco Fummi, Graziano Pravadelli, Francesco Stefanni, and Sara Vinco. UNIVERCM: The UNIversal VERsatile Computational Mode for heterogeneous system integration. IEEE Transactions on Computers, 62(2):225–241, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dietrich:2015:TSC

[DGC^{+}15] B. Dietrich, D. Goswami, S. Chakraborty, A. Guha, and M. Gries. Time series characterization of gaming workload for runtime power management. IEEE Transactions on Computers, 64 (1):260–273, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Du:2016:ACD

[DHC^{+}16] Bowen Du, Runhe Huang, Xi Chen, Zhipu Xie, Ye Liang, Weifeng Lv, and Jianhua Ma. Active CTDaaS: A data service framework based on transparent IoD in city traffic. IEEE Transactions on Computers, 65(12): 3524–3536, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

DelBarrio:2016:PCS

[DHM16] Alberto A. Del Barrio, Román Hermida, and Seda Ogrenci Memik. A partial carry-save on-the-fly correction multispeculative multiplier. IEEE Transactions on Computers, 65(11):3251–3264, November 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dai:2019:HHV

[DHW^{+}19] G. Dai, T. Huang, Y. Wang, H. Yang, and J. Wawrzynek. HyVE: Hybrid vertex-edge memory hierarchy for energyefficient graph processing. IEEE Transactions on Computers, 68(8):1131–1146, August 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dimitrov:2011:AEM

[DJA11] Vassil S. Dimitrov, Kimmo U. Jarvinen, and Jithra Adikari. Area-efficient multipliers based on multiple-radix representations. IEEE Transactions on Computers, 60 (2):189–201, February 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dimitrov:2014:FIF

[DJA14] Vassil Dimitrov, Kimmo Jarvinen, and Reza Azarderakhsh. Fast inversion in GF(2^{m}) with normal basis using hybrid-double multipliers. IEEE Transactions on Computers, 63(4):1041– 1047, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dimitrov:2008:PSP

[DJJ^{+}08] V. S. Dimitrov, K. U.

Jarvinen, M. J. Jacobson, W. Chan, and Zhun Huang. Provably sublinear point multiplication on Koblitz curves and its hardware implementation. IEEE Transactions on Computers, 57 (11):1469–1481, November 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=4487060. See

comments [Lee12].

Dutta:2017:NLC

[DJN17] Avijit Dutta, Ashwin Jha, and Mridul Nandi. A new look at counters: Don’t run like Marathon in a hundred meter race. IEEE Transactions on Computers, 66(11): 1851–1864, November 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7936672/.

Dubach:2011:EAC

[DJO11] Christophe Dubach, Timothy M. Jones, and Michael F. P. O’Boyle. An empirical architecture-centric approach to microarchitectural design space exploration. IEEE Transactions on Computers, 60(10): 1445–1458, October 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5674026.

Dimitrakopoulos:2013:MSA

[DKG13] Giorgos Dimitrakopoulos, Emmanouil Kalligeros, and Kostas Galanopoulos. Merged switch allocation and traversal in network-on-chip switches. IEEE Transactions on Computers, 62(10):2001–2012, October 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Deshpande:2013:RRF

[DKH^{+}13] Mayur Deshpande, Kyungbaek Kim, Bijit Hore, Sharad Mehrotra, and Nalini Venkatasubramanian. ReCREW: A reliable flash-dissemination system. IEEE Transactions on Computers, 62(7):1432– 1446, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dumitriu:2016:RTR

[DKK16] Victor Dumitriu, Lev Kirischian, and Valeri Kirischian. Runtime recovery mechanism for transient and permanent hardware faults based on distributed, self-organized dynamic partially reconfigurable systems. IEEE Transactions on Computers, 65(9): 2835–2847, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

DiTomaso:2015:RPE

[DKLB15] D. DiTomaso, A. K. Kodi, A. Louri, and R. Bunescu. Resilient and power-efficient multi-function channel buffers in network-on-chip architectures. IEEE Transactions on Computers, 64(12): 3555–3568, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Di:2015:OCC

[DKW15] Sheng Di, D. Kondo, and Cho-Li Wang. Optimization of composite cloud service processing with virtual machines. IEEE Transactions on Computers, 64 (6):1755–1768, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dong:2013:RIR

[DLC^{+}13] Wei Dong, Yunhao Liu, Chun Chen, Jiajun Bu, Chao Huang, and Zhiwei Zhao. R2: Incremental reprogramming using relocatable code in networked embedded systems. IEEE Transactions on Computers, 62(9): 1837–1849, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dong:2012:DCW

[DLL^{+}12] Dezun Dong, Xiangke Liao, Kebin Liu, Yunhao Liu, and Weixia Xu. Distributed coverage in wireless ad hoc and sensor networks by topological graph approaches. IEEE Transactions on Computers, 61(10):1417–1428, October 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Saint-Genies:2017:ELT

[dLSGDR17] Hugues de Lassus Saint-Geniès, David Defour, and Guillaume Revy. Exact lookup tables for the evaluation of trigonometric and hyperbolic functions. IEEE Transactions on Computers, 66(12):2058–2071, December 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7927421/.

DiFrancesco:2015:SMA

[DMA^{+}15] P. Di Francesco, S. McGettrick, U. K. Anyanwu, J. C. O’Sullivan, A. B. MacKenzie, and L. A. DaSilva. A split MAC approach for SDR platforms. IEEE Transactions on Computers, 64 (4):912–924, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dehyadegari:2015:AST

[DMK^{+}15] M. Dehyadegari, A. Marongiu, M. R. Kakoee, S. Mohammadi, N. Yazdani, and L. Benini. Architecture support for tightly-coupled multi-core clusters with shared-memory HW accelerators. IEEE Transactions on Computers, 64(8): 2132–2144, August 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dahir:2014:MTP

[DMXY14] Nizar S. Dahir, Terrence Mak, Fei Xia, and Alexandre Yakovlev. Modeling and tools for power supply variations analysis in networks-on-chip. IEEE Transactions on Computers, 63(3):679–690, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Daoud:2011:ULC

[DN11] Ehab Anis Daoud and Nicola Nicolici. On using lossy compression for repeatable experiments during silicon debug. IEEE Transactions on Computers, 60(7):937– 950, July 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Demmel:2015:PRS

[DN15] J. Demmel and Hong Diep Nguyen. Parallel reproducible summation. IEEE Transactions on Computers, 64(7):2060–2070, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Duckham:2011:EDC

[DNSS11] Matt Duckham, Doron Nussbaum, Jorg-Rudiger Sack, and Nicola Santoro. Efficient, decentralized computation of the topology of spatial regions. IEEE Transactions on Computers, 60 (8):1100–1113, August 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

deOliveira:2016:EMR

[dOPSR16] D. A. G. de Oliveira, L. L. Pilla, T. Santini, and P. Rech. Evaluation and mitigation of radiation-induced soft errors in graphics processing units. IEEE Transactions on Computers, 65 (3):791–804, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Doroz:2015:AFH

[DOS15] Y. Doroz, E. Ozturk, and B. Sunar. Accelerating fully homomorphic encryption in hardware. IEEE Transactions on Computers, 64 (6):1509–1521, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dabosville:2013:NSO

[DP13] Guillaume Dabosville and Emmanuel Prouff. A new second-order side channel attack based on linear regression. IEEE Transactions on Computers, 62(8): 1629–1640, August 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

DeMara:2017:GEI

[DPO17] Ronald F. DeMara, Marco Platzner, and Marco Ottavi. Guest editorial: IEEE Transactions on Computers and IEEE Transactions on Emerging Topics in Computing joint special section on innovation in reconfigurable computing fabrics from devices to architectures. IEEE Transactions on Computers, 66(6):927–929, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

DiCarlo:2011:SBS

[DPS11] Stefano Di Carlo, Paolo Prinetto, and Alessandro Savino. Software-based selftest of set-associative cache memories. IEEE Transactions on Computers, 60 (7):1030–1044, July 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Drane:2014:SCF

[DRC14] T. A. Drane, T. M. Rose, and G. A. Constantinides. On the systematic creation of faithfully rounded truncated multipliers and arrays. IEEE Transactions on Computers, 63(10):2513– 2525, October 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Demirci:2016:UEP

[DRM16] Mustafa Demirci, Pedro Reviriego, and Juan Antonio Maestro. Unequal error protection codes derived from double error correction orthogonal Latin square codes. IEEE Transactions on Computers, 65(9): 2932–2938, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Du:2016:OTC

[DRS^{+}16] Boyang Du, Matteo Sonza Reorda, Luca Sterpone, Luis Parra, Marta Portela-García, Almudena Lindoso, and Luis Entrena. Online test of control flow errors: A new debug interface-based approach. IEEE Transactions on Computers, 65 (6):1846–1855, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

deRooij:2012:ARD

[dRV12] Steven de Rooij and Paul M. B. Vitanyi. Approximating rate-distortion graphs of individual data: Experiments in lossy compression and denoising. IEEE Transactions on Computers, 61 (3):395–407, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Doche:2014:NIM

[DS14] Christophe Doche and Daniel Sutantyo. New and improved methods to analyze and compute double-scalar multiplications. IEEE Transactions on Computers, 63 (1):230–242, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dey:2013:KAT

[DSB13] Soumyajit Dey, Dipankar Sarkar, and Anupam Basu. A Kleene algebra of tagged system actors for reasoning about heterogeneous embedded systems. IEEE Transactions on Computers, 62(10): 1917–1931, October 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Di:2019:MRI

[DSG^{+}19] Yejia Di, Liang Shi, Congming Gao, Qiao Li, Chun Jason Xue, and Kaijie Wu. Minimizing retention induced refresh through exploiting process variation of flash memory. IEEE Transactions on Computers, 68(1):83–98, January 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8423091/.

Deniz:2015:MPA

[DSKH15] E. Deniz, A. Sen, B. Kahne, and J. Holt. MINIME: Pattern-aware multicore benchmark synthesizer. IEEE Transactions on Computers, 64(8):2239–2252, August 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dubois:2013:EFD

[DSPB13] Florentine Dubois, Abbas Sheibanyrad, Frederic Petrot, and Maryam Bahmani. Elevator-first: A deadlock-free distributed routing algorithm for vertically partially connected 3D-NoCs. IEEE Transactions on Computers, 62 (3):609–615, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Datta:2015:PSO

[DSR15] K. Datta, I. Sengupta, and H. Rahaman. A postsynthesis optimization technique for reversible circuits exploiting negative control lines. IEEE Transactions on Computers, 64(4):1208– 1214, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dillow:2014:CGC

[DSW^{+}14] David A. Dillow, Galen M.

Shipman, Feiyi Wang, Sarp Oral, Junghee Lee, and Youngjae Kim. Coordinating garbage collection for arrays of solid-state drives. IEEE Transactions on Computers, 63(4):888–901, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Du:2015:PTD

[DSY^{+}15] A. Y. Du, S. D. Smith, Z. Yang, C. Qiao, and R. Ramesh. Predicting transient downtime in virtual server systems: An efficient sample path randomization approach. IEEE Transactions on Computers, 64(12): 3541–3554, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dumas:2014:NRI

[Dum14] Jean-Guillaume Dumas. On Newton–Raphson iteration for multiplicative inverses modulo prime powers. IEEE Transactions on Computers, 63(8):2106–2109, August 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See corrections [Wal19].

Duan:2014:MCB

[DVUS14] Qi Duan, Mohit Virendra, Shambhu Upadhyaya, and Ameya Sanzgiri. Minimum cost blocking problem in multi-path wireless routing protocols. IEEE Transactions on Computers, 63(7):1765–1777, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Duarte:2010:CBM

[DW10] F. Duarte and S. Wong. Cache-based memory copy hardware accelerator for multicore systems. IEEE Transactions on Computers, 59(11):1494–1507, November 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5416684.

Dai:2018:OPC

[DWZ18] Wei Dai, William Whyte, and Zhenfei Zhang. Optimizing polynomial convolution for NTRUEncrypt. IEEE Transactions on Computers, 67(11):1572–1583, November 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8303667/.

Deng:2012:OAC

[DY12] Xi Deng and Yuanyuan Yang. Online adaptive compression in delay sensitive wireless sensor networks. IEEE Transactions on Computers, 61(10): 1429–1442, October 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Du:2014:IES

[DY14] Yuyang Du and Hongliang Yu. Increasing endurance and security of phase-change memory with multi-way wear-leveling. IEEE Transactions on Computers, 63 (5):1157–1168, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ding:2016:EST

[DYCG16] Z. Ding, B. Yang, Y. Chi, and L. Guo. Enabling smart transportation systems: A parallel spatiotemporal database approach. IEEE Transactions on Computers, 65(5):1377–1391, May 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

D:2016:QLB

[DYHX16] S. M. P. D., H. Yu, H. Huang, and D. Xu. A Qlearning based self-adaptive I/O communication for 2.5D integrated many-core microprocessor and memory. IEEE Transactions on Computers, 65(4):1185–1196, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dai:2019:NNN

[DYJ19] X. Dai, H. Yin, and N. K. Jha. NeST: A neural network synthesis tool based on a grow-and-prune paradigm. IEEE Transactions on Computers, 68(10):1487–1497, October 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

D:2015:MCM

[DYW15] S. Manoj P. D., H. Yu, and K. Wang. 3D manycore microprocessor power management by space-time multiplexing based demandsupply matching. IEEE Transactions on Computers, 64(11):3022–3036, November 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ding:2010:LBI

[DZ10] Yiqiang Ding and Wei Zhang. Loop-based instruction prefetching to reduce the worst-case execution time. IEEE Transactions on Computers, 59 (6):855–864, June 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5416687.

Deng:2016:ECD

[DZD^{+}16] Peng Deng, Qi Zhu, Abhijit Davare, Anastasios Mourikis, Xue Liu, and Marco Di Natale. An efficient control-driven period optimization algorithm for distributed real-time systems. IEEE Transactions on Computers, 65(12): 3552–3566, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Duan:2014:CED

[DZLP14] Lide Duan, Ying Zhang, Bin Li, and Lu Peng. Comprehensive and efficient design parameter selection for soft error resilient processors via universal rules. IEEE Transactions on Computers, 63(9): 2201–2214, September 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Elarief:2013:LME

[EBE13] Noha Elarief, Bella Bose, and Samir Elmougy. Limited magnitude error detecting codes over Z_{q}. IEEE Transactions on Computers, 62(5):984–989, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Egger:2016:ECL

[ECJ^{+}16] Bernhard Egger, Younghyun Cho, Changyeon Jo, Eunbyun Park, and Jaejin Lee. Efficient checkpointing of live virtual machines. IEEE Transactions on Computers, 65(10):3041–3054, October 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ebrahimi:2014:PBP

[EDL^{+}14] Masoumeh Ebrahimi, Masoud Daneshtalab, Pasi Liljeberg, Juha Plosila, Jose Flich, and Hannu Tenhunen. Path-based partitioning methods for 3D networks-on-chip with minimal adaptive routing. IEEE Transactions on Computers, 63(3):718–733, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Eyerman:2010:CAO

[EE10] S. Eyerman and L. Eeckhout. A counter architecture for online DVFS profitability estimation. IEEE Transactions on Computers, 59(11): 1576–1583, November 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5432162.

Ecco:2017:TBT

[EE17] Leonardo Ecco and Rolf Ernst. Tackling the bus turnaround overhead in realtime SDRAM controllers. IEEE Transactions on Computers, 66(11):1961–1974, November 2017. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7946168/.

Etsion:2012:ECW

[EF12] Yoav Etsion and Dror G.

Feitelson. Exploiting core working sets to filter the L1 cache with random sampling. IEEE Transactions on Computers, 61(11):1535–1550, November 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Espitau:2018:LAF

[EFGT18] Thomas Espitau, Pierre-Alain Fouque, Benoît Gérard, and Mehdi Tibouchi. Loopabort faults on lattice-based signature schemes and key exchange protocols. IEEE Transactions on Computers, 67(11):1535–1549, November 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8354897/.

Esposito:2016:SCS

[EFPC16] Christian Esposito, Massimo Ficco, Francesco Palmieri, and Aniello Castiglione. Smart cloud storage service selection based on fuzzy logic, theory of evidence and game theory. IEEE Transactions on Computers, 65 (8):2348–2362, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Eppstein:2011:SGG

[EG11] David Eppstein and Michael T. Goodrich. Succinct greedy geometric routing using hyperbolic geometry. IEEE Transactions on Computers, 60(11):1571–1580, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5669279.

Entrena:2012:SES

[EGVFC^{+}12] L. Entrena, M. Garcia-Valderas, R. Fernandez-Cardenal, A. Lindoso, M. Portela Garcia, and C. Lopez-Ongil.

Soft error sensitivity evaluation of microprocessors by multilevel emulation-based fault injection. IEEE Transactions on Computers, 61 (3):313–322, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ebergen:2015:RDA

[EJ15] J. Ebergen and N. Jamadagni. Radix-2 division algorithms with an overredundant digit set. IEEE Transactions on Computers, 64(9):2652–2663, September 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ebrahimi:2017:PPE

[EKA17] Zahra Ebrahimi, Behnam Khaleghi, and Hossein Asadi. PEAF: A power-efficient architecture for SRAM-based FPGAs using reconfigurable hard logic design in dark silicon era. IEEE Transactions on Computers, 66 (6):982–995, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7775010/.

Egger:2010:SMM

[EKJ^{+}10] Bernhard Egger, Seungkyun Kim, Choonki Jang, Jaejin Lee, Sang Lyul Min, and Heonshik Shin. Scratchpad memory management techniques for code in embedded systems without an MMU. IEEE Transactions on Computers, 59(8): 1047–1062, August 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5374377.

Elnozahy:2012:GEI

[EM12] Mootaz Elnozahy and Rami Melhem. Guest Editors’ introduction: Special section on energy efficient computing. IEEE Transactions on Computers, 61(12): 1666–1667, December 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

El-Razouk:2016:NAD

[ERRM16] Hayssam El-Razouk and Arash Reyhani-Masoleh. New architectures for digitlevel single, hybrid-double, hybrid-triple field multiplications and exponentiation using Gaussian normal bases. IEEE Transactions on Computers, 65(8):2495–2509, August 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

El-Razouk:2015:NHI

[ERRMG15] H. El-Razouk, A. Reyhani-Masoleh, and Guang Gong.

New hardware implementations of WG and WG-StreamCiphers using polynomial basis. IEEE Transactions on Computers, 64 (7):2020–2035, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Eghbal:2015:AFT

[EYBK15] A. Eghbal, P. M. Yaghini, N. Bagherzadeh, and M. Khayambashi. Analytical fault tolerance assessment and metrics for TSV-based 3D network-on-chip. IEEE Transactions on Computers, 64(12):3591–3604, December 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Flores:2010:HIE

[FAA10] A. Flores, J. L. Aragon, and M. E. Acacio. Heterogeneous interconnects for energyefficient message management in CMPs. IEEE Transactions on Computers, 59 (1):16–28, January 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5235139.

Fedorov:2016:FAE

[FAK16] Viacheslav V. Fedorov, Monther Abusultan, and Sunil P. Khatri. FTCAM: An areaefficient flash-based ternary CAM design. IEEE Transactions on Computers, 65 (8):2652–2658, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fan:2016:CRT

[Fan16] H. Fan. A Chinese remainder theorem approach to bit-parallel polynomial basis multipliers for irreducible trinomials. IEEE Transactions on Computers, 65 (2):343–352, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Flahive:2013:RPG

[FB13] Mary Flahive and Bella Bose. On resource placement in Gaussian and EJ interconnection networks. IEEE Transactions on Computers, 62(3):623–626, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ferguson:2018:DSM

[FBE^{+}18] Warren E. Ferguson, Jesse Bingham, Levent Erkök, John R. Harrison, and Joe Leslie-Hurd. Digit serial methods with applications to division and square root. IEEE Transactions on Computers, 67 (3):449–456, March 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8060979/.

Ferrer:2012:PCM

[FBR^{+}12] Joan-Lluis Ferrer, Elvira Baydal, Antonio Robles, Pedro Lopez, and Jose Duato. Progressive congestion management based on packet marking and validation techniques. IEEE Transactions on Computers, 61(9): 1296–1310, September 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fensch:2013:DPL

[FBWMM13] Christian Fensch, Nick Barrow-Williams, Robert D. Mullins, and Simon Moore. Designing a physical locality aware coherence protocol for chip-multiprocessors. IEEE Transactions on Computers, 62(5):914–928, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fettes:2019:DVF

[FCB^{+}19] Quintin Fettes, Mark Clark, Razvan Bunescu, Avinash Karanth, and Ahmed Louri. Dynamic voltage and frequency scaling in NoCs with supervised and reinforcement learning techniques. IEEE Transactions on Computers, 68 (3):375–389, March 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8489913/.

Favalli:2016:BPB

[FD16] M. Favalli and M. Dalpasso. Boolean and pseudo-Boolean test generation for feedback bridging faults. IEEE Transactions on Computers, 65 (3):706–715, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feng:2018:PCA

[FEM^{+}18] Huan Feng, David Eyers, Steven Mills, Yongwei Wu, and Zhiyi Huang. Principal component analysis based filtering for scalable, high precision k-NN search. IEEE Transactions on Computers, 67(2):252–267, February 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8024082/.

Feng:2014:ISE

[Fen14] Gang Feng. Improving space efficiency with path length prediction for finding k shortest simple paths. IEEE Transactions on Computers, 63(10):2459–2472, October 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Flocchini:2012:DMS

[FEP^{+}12] Paola Flocchini, T. Mesa Enriquez, Linda Pagli, Giuseppe Prencipe, and Nicola Santoro. Distributed minimum spanning tree maintenance for transient node failures. IEEE Transactions on Computers, 61(3):408– 414, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fritz:2016:IPM

[FF16] Christopher Fritz and Adly T.

Fam. Interlaced partition multiplier. IEEE Transactions on Computers, 65 (8):2672–2677, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fanelli:2014:SAC

[FFCB14] M. Fanelli, L. Foschini, A. Corradi, and A. Boukerche. Self-adaptive context data management in largescale mobile systems. IEEE Transactions on Computers, 63(10):2549–2562, October 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Farmahini-Farahani:2013:MDH

[FFISC13] Amin Farmahini-Farahani, Henry J. Duwe III, Michael J. Schulte, and Katherine Compton. Modular design of high-throughput, lowlatency sorting units. IEEE Transactions on Computers, 62(7):1389–1402, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Friedrichs:2018:MCC

[FFL18] Stephan Friedrichs, Matthias Függer, and Christoph Lenzen. Metastabilitycontaining circuits. IEEE Transactions on Computers, 67(8):1167–1183, August 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8314764/.

Francalanci:2010:HPD

[FG10] C. Francalanci and P. Giacomazzi. A high-performance deadlock-free multicast routing algorithm for K-ary N-cubes. IEEE Transactions on Computers, 59 (2):174–187, February 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5156496.

Franchini:2013:DIE

[FGS^{+}13] Silvia Franchini, Antonio Gentile, Filippo Sorbello, Giorgio Vassallo, and Salvatore Vitabile. Design and implementation of an embedded coprocessor with native support for 5D, quadruplebased Clifford algebra. IEEE Transactions on Computers, 62(12):2366–2381, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Franchini:2015:CCG

[FGS^{+}15] S. Franchini, A. Gentile, F. Sorbello, G. Vassallo, and S. Vitabile. ConformalALU: A conformal geometric algebra coprocessor for medical image processing. IEEE Transactions on Computers, 64(4):955–970, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fan:2010:AMI

[FHH10] Chun-I Fan, Ling-Ying Huang, and Pei-Hsiu Ho. Anonymous multireceiver identity-based encryption. IEEE Transactions on Computers, 59(9):1239–1249, September 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5396332.

Fu:2018:SAR

[FHL^{+}18] Min Fu, Shujie Han, Patrick P. C. Lee, Dan Feng, Zuoning Chen, and Yu Xiao. A simulation analysis of redundancy and reliability in primary storage deduplication. IEEE Transactions on Computers, 67(9): 1259–1272, September 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8300656/.

Faz-Hernandez:2018:FSI

[FHLOJRH18] Armando Faz-Hernández, Julio López, Eduardo Ochoa-Jiménez, and Francisco Rodríguez-Henríquez. A faster software implementation of the supersingular isogeny Diffie–Hellman key exchange protocol. IEEE Transactions on Computers, 67(11):1622–1636, November 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8100879/.

Fan:2014:ASA

[FHR14] Chun-I Fan, Vincent Shi-Ming Huang, and He-Ming Ruan. Arbitrarystate attribute-based encryption with dynamic membership. IEEE Transactions on Computers, 63(8): 1951–1961, August 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Frankel:2018:QBE

[FHW18] Binyamin Frankel, Roi Herman, and Shmuel Wimer.

Queuing-based eDRAM refreshing for ultra-low power processors. IEEE Transactions on Computers, 67(9): 1331–1340, September 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8310027/.

Finkler:2010:AFD

[Fin10] U. Finkler. An analytic framework for detailed resource profiling in large and parallel programs and its application for memory use. IEEE Transactions on Computers, 59 (3):358–370, March 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5276794.

Fernandez:2017:CSC

[FJA^{+}17] Gabriel Fernandez, Javier Jalle, Jaume Abella, Eduardo Quiñones, Tullio Vardanega, and Francisco J. Cazorla. Computing safe contention bounds for multicore resources with round-robin and FIFO arbitration. IEEE Transactions on Computers, 66(4):586–600, April 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feng:2015:FMC

[FK15] Gang Feng and T. Korkmaz. Finding multiconstrained multiple shortest paths. IEEE Transactions on Computers, 64(9): 2559–2572, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Farbeh:2016:FED

[FKMK16] Hamed Farbeh, Hyeonggyu Kim, Seyed Ghassem Miremadi, and Soontae Kim. Floating-ECC: Dynamic repositioning of error correcting code bits for extending the lifetime of STT-RAM caches. IEEE Transactions on Computers, 65(12): 3661–3675, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Faltings:2014:IMC

[FLJ14] Boi Faltings, Jason Jingshi Li, and Radu Jurca. Incentive mechanisms for community sensing. IEEE Transactions on Computers, 63 (1):115–128, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feng:2014:PCO

[FLL14] Yuan Feng, Baochun Li, and Bo Li. Price competition in an oligopoly market with multiple IaaS cloud providers. IEEE Transactions on Computers, 63 (1):59–73, January 2014.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Furber:2013:OSS

[FLP^{+}13] Steve B. Furber, David R.

Lester, Luis A. Plana, Jim D. Garside, Eustace Painkras, Steve Temple, and Andrew D. Brown. Overview of the SpiNNaker system architecture. IEEE Transactions on Computers, 62(12): 2454–2467, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ferrara:2016:SBO

[FLS16] Andrea Ferrara, Paolo Liberatore, and Marco Schaerf. The size of BDDs and other data structures in temporal logics model checking. IEEE Transactions on Computers, 65(10):3148–3156, October 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Firdaus:2016:NPV

[FM16] A. Fajar Firdaus and M. Meribout. A new parallel VLSI architecture for real-time electrical capacitance tomography. IEEE Transactions on Computers, 65 (1):30–41, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Farahmandi:2019:ATG

[FM19] Farimah Farahmandi and Prabhat Mishra. Automated test generation for debugging multiple bugs in arithmetic circuits. IEEE Transactions on Computers, 68 (2):182–197, February 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8453904/.

Fu:2010:FDO

[FML10] Haohuan Fu, O. Mencer, and W. Luk. FPGA designs with optimized logarithmic arithmetic. IEEE Transactions on Computers, 59(7):1000–1006, July 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5416693.

Fabiano:2019:ATW

[FMP19] Nicolas Fabiano, Jean-Michel Muller, and Joris Picot. Algorithms for tripleword arithmetic. IEEE Transactions on Computers, 68(11):1573–1583, November 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /hal.archives-ouvertes.
                                        fr/hal-01869009v2.

Fadlullah:2019:VIA

[FMTK19] Z. M. Fadlullah, B. Mao, F. Tang, and N. Kato. Value iteration architecture based deep learning for intelligent routing exploiting heterogeneous computing platforms. IEEE Transactions on Computers, 68(6):939– 950, June 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fugger:2016:UGP

[FNS16] M. Fugger, T. Nowak, and U. Schmid. Unfaithful glitch propagation in existing binary circuit models. IEEE Transactions on Computers, 65(3):964–978, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Find:2019:BCB

[FP19] Magnus Gaudal Find and René Peralta. Better circuits for binary polynomial multiplication. IEEE Transactions on Computers, 68(4):624–630, April 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8494787/.

Fyrbiak:2018:HOP

[FRB^{+}18] Marc Fyrbiak, Simon Rokicki, Nicolai Bissantz, Russell Tessier, and Christof Paar. Hybrid obfuscation to protect against disclosure attacks on embedded microprocessors. IEEE Transactions on Computers, 67 (3):307–321, March 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7809080/.

Frey:2010:DGW

[FS10] H. Frey and I. Stojmenovic. On delivery guarantees and worst-case forwarding bounds of elementary face routing components in ad hoc and sensor networks. IEEE Transactions on Computers, 59(9): 1224–1238, September 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5467049.

Ferreron:2016:CSC

[FSGAB^{+}16] A. Ferreron, D. Suarez-Gracia, J. Alastruey-Benede, T. Monreal-Arnal, and P. Ibanez. Concertina: Squeezing in cache content to operate at nearthreshold voltage. IEEE Transactions on Computers, 65(3):755–769, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fu:2017:SCE

[FSL^{+}17] Yingxun Fu, Jiwu Shu, Xianghong Luo, Zhirong Shen, and Qingda Hu. Short code: An efficient RAID-6 MDS code for optimizing degraded reads and partial stripe writes. IEEE Transactions on Computers, 66 (1):127–137, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feliu:2016:BAL

[FSPD16] J. Feliu, J. Sahuquillo, S. Petit, and J. Duato. Bandwidth-aware on-line scheduling in SMT multicores. IEEE Transactions on Computers, 65(2):422– 434, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feliu:2017:PFP

[FSPD17] J. Feliu, J. Sahuquillo, S. Petit, and J. Duato. Perf Fair: A progress-aware scheduler to enhance performance and fairness in SMT multicores. IEEE Transactions on Computers, 66(5):905–911, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fraccaroli:2018:NSD

[FSR^{+}18] Enrico Fraccaroli, Francesco Stefanni, Romeo Rizzi, Davide Quaglia, and Franco Fummi. Network synthesis for distributed embedded systems. IEEE Transactions on Computers, 67(9): 1315–1330, September 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8307094/.

Foroutan:2013:ICT

[FTP13] Sahar Foroutan, Yvain Thonnart, and Frederic Petrot. An iterative computational technique for performance evaluation of networks-on-chip. IEEE Transactions on Computers, 62(8):1641–1655, August 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fujita:2011:BBA

[Fuj11] Satoshi Fujita. A branchand-bound algorithm for solving the multiprocessor scheduling problem with improved lower bounding techniques. IEEE Transactions on Computers, 60(7):1006– 1016, July 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fan:2012:EHI

[FVV12] Junfeng Fan, Frederik Vercauteren, and Ingrid Verbauwhede. Efficient hardware implementation of FP-arithmetic for pairingfriendly curves. IEEE Transactions on Computers, 61(5): 676–685, May 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fang:2014:AEA

[FYSK14] Hua-Wei Fang, Mi-Yen Yeh, Pei-Lun Suei, and Tei-Wei Kuo. An adaptive endurance-aware B^{+}-tree for flash memory storage systems. IEEE Transactions on Computers, 63(11): 2661–2673, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fang:2014:MRP

[FZL^{+}14] Kun Fang, Hongzhong Zheng, Jiang Lin, Zhao Zhang, and Zhichun Zhu. Mini-Rank: A power-efficient DDRx DRAM memory architecture. IEEE Transactions on Computers, 63 (6):1500–1512, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guralnik:2011:SBV

[GABK11] Elena Guralnik, Merav Aharoni, Ariel J. Birnbaum, and Anatoly Koyfman. Simulation-based verification of floating-point division. IEEE Transactions on Computers, 60(2):176– 188, February 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guzzi:2014:CPP

[GAC14] P. H. Guzzi, G. Agapito, and M. Cannataro. coreSNP: Parallel processing of microarray data. IEEE Transactions on Computers, 63 (12):2961–2974, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ghandali:2015:AHL

[GAFN15] S. Ghandali, B. Alizadeh, M. Fujita, and Z. Navabi. Automatic high-level dataflow synthesis and optimization of polynomial datapaths using functional decomposition. IEEE Transactions on Computers, 64 (6):1579–1593, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ghaderi:2018:SSA

[GBA18a] Zana Ghaderi, Nader Bagherzadeh, and Ahmad Albaqsami. STABLE: Stress-aware Boolean matching to mitigate BTIinduced SNM reduction in SRAM-based FPGAs. IEEE Transactions on Computers, 67(1):102–114, January 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7974802/.

Gupta:2018:OLM

[GBA^{+}18b] Ujjwal Gupta, Manoj Babu, Raid Ayoub, Michael Kishinevsky, Francesco Paterna, Suat Gumussoy, and Umit Y. Ogras. An online learning methodology for performance modeling of graphics processors. IEEE Transactions on Computers, 67(12): 1677–1691, December 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8365819/.

Gottscho:2015:VHP

[GBD^{+}15] M. Gottscho, L. A. D. Bathen, N. Dutt, A. Nicolau, and P. Gupta. ViPZonE: Hardware power variabilityaware virtual memory management for energy savings. IEEE Transactions on Computers, 64(5):1483– 1496, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gorshkov:2019:GIH

[GBF^{+}19] A. V. Gorshkov, M. Berezalsky, J. Fedorova, K. Levit-Gurevich, and N. Itzhaki. GPU instruction hotspots detection based on binary instrumentation approach. IEEE Transactions on Computers, 68(8):1213–1224, August 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2018:SMI

[GBGI18] Xiaochen Guo, Mahdi Nazm Bojnordi, Qing Guo, and Engin Ipek. Sanitizer: Mitigating the impact of expensive ECC checks on STT-MRAM based main memories. IEEE Transactions on Computers, 67(6):847–860, June 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8126226/.

Garibotti:2016:EES

[GBO^{+}16] Rafael Garibotti, Anastasiia Butko, Luciano Ost, Abdoulaye Gamatié, Gilles Sassatelli, and Chris Adeniyi-Jones. Efficient embedded software migration towards clusterized distributed-memory architectures. IEEE Transactions on Computers, 65 (8):2645–2651, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ginosar:2014:GEI

[GC14] Ran Ginosar and Karam S. Chatha. Guest Editors’ introduction — special issue on network-on-chip. IEEE Transactions on Computers, 63(3):527–528, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gong:2016:ERE

[GC16] Y. H. Gong and S. W.

Chung. Exploiting refresh effect of DRAM read operations: A practical approach to low-power refresh. IEEE Transactions on Computers, 65(5):1507–1517, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Goossens:2016:PPT

[GCAG16] Sven Goossens, Karthik Chandrasekar, Benny Akesson, and Kees Goossens. Power/ performance trade-offs in real-time SDRAM command scheduling. IEEE Transactions on Computers, 65 (6):1882–1895, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ganguly:2011:SHW

[GCD^{+}11] Amlan Ganguly, Kevin Chang, Sujay Deb, Partha Pratim Pande, Benjamin Belzer, and Christof Teuscher. Scalable hybrid wireless networkon-chip architectures for multicore systems. IEEE Transactions on Computers, 60(10):1485–1502, October 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5551123.

Guo:2016:MAR

[GCF^{+}16] W. Guo, K. Chen, H. Feng, Y. Wu, R. Zhang, and W. Zheng. Mobile application relaunching speedup through flash-aware page swapping. IEEE Transactions on Computers, 65 (3):916–928, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2013:ECE

[GCL^{+}13] Deke Guo, Tao Chen, Dan Li, Mo Li, Yunhao Liu, and Guihai Chen. Expandable and cost-effective network structures for data centers using dual-port servers. IEEE Transactions on Computers, 62(7):1303–1317, July 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gao:2011:TMO

[GCLC11] Ming Gao, Hsiu-Ming (Sherman) Chang, Peter Lisherness, and Kwang-Ting (Tim) Cheng. Time-multiplexed online checking. IEEE Transactions on Computers, 60(9):1300–1312, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5710883.

Grass:2019:SST

[GCR^{+}19] Thomas Grass, Trevor E.

Carlson, Alejandro Rico, Germán Ceballos, Eduard Ayguadé, Marc Casas, and Miquel Moreto. Sampled simulation of task-based programs. IEEE Transactions on Computers, 68(2): 255–269, February 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8424416/.

Gupta:2013:HPH

[GCS^{+}13] Sourav Sen Gupta, A. Chattopadhyay, K. Sinha, S. Maitra, and B. P. Sinha. Highperformance hardware implementation for RC4 stream cipher. IEEE Transactions on Computers, 62(4):730– 743, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gade:2017:HHW

[GD17] Sri Harsha Gade and Sujay Deb. HyWin: Hybrid wireless NoC with sandboxed sub-networks for CPU/GPU architectures. IEEE Transactions on Computers, 66 (7):1145–1158, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07795230-
abs.html.

Gao:2016:TRR

[GDC^{+}16] Y. Gao, W. Dong, C. Chen, J. Bu, and X. Liu. Towards reconstructing routing paths in large scale sensor networks. IEEE Transactions on Computers, 65 (1):281–293, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Grieve:2018:ARI

[GDJZ18] Alex Grieve, Michael Davies, Phillip H. Jones, and Joseph Zambreno. AR-MOR: A recompilation and instrumentation-free monitoring architecture for detecting memory exploits. IEEE Transactions on Computers, 67(8):1092–1104, August 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8295231/.

Gao:2018:PRR

[GDLL18] Xinwei Gao, Jintai Ding, Lin Li, and Jiqiang Liu. Practical randomized RLWE-based key exchange against signal leakage attack. IEEE Transactions on Computers, 67(11):1584–1593, November 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8300634/.

Guo:2015:JOD

[GDY15] Songtao Guo, Changyin Dang, and Yuanyuan Yang. Joint optimal data rate and power allocation in lossy mobile ad hoc networks with delay-constrained traffics. IEEE Transactions on Computers, 64(3):747– 762, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gedik:2014:DFG

[Ged14] Bugra Gedik. Discriminative fine-grained mixing for adaptive compression of data streams. IEEE Transactions on Computers, 63(9): 2228–2244, September 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ghaderi:2017:SHS

[GEN^{+}17] Z. Ghaderi, M. Ebrahimi, Z. Navabi, E. Bozorgzadeh, and N. Bagherzadeh. SEN-SIBle: A highly scalable SENsor DeSIgn for Path-Based Age Monitoring in FPGAs. IEEE Transactions on Computers, 66(5):919– 926, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Garbacki:2010:DES

[GEvS10] P. Garbacki, D. H. J. Epema, and M. van Steen. The design and evaluation of a self-organizing superpeer network. IEEE Transactions on Computers, 59 (3):317–331, March 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5291690.

Gupta:2011:SRF

[GFAM11] Shantanu Gupta, Shuguang Feng, Amin Ansari, and Scott Mahlke. StageNet: a reconfigurable fabric for constructing dependable CMPs. IEEE Transactions on Computers, 60(1):5–19, January 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gomony:2017:GAM

[GGA^{+}17] Manil Dev Gomony, Jamie Garside, Benny Akesson, Neil Audsley, and Kees Goossens. A globally arbitrated memory tree for mixed-time-criticality systems. IEEE Transactions on Computers, 66(2):212– 225, February 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Garcia-Guirado:2015:ICC

[GGFPG15] A. Garcia-Guirado, R. Fernandez-Pascual, and J. M. Garcia. ICCI: In-cache coherence information. IEEE Transactions on Computers, 64(4):995–1014, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2014:MLW

[GGL^{+}14] Song Guo, Minyi Guo, Victor C. M. Leung, Shui Yu, and Yong Xiang. On the multicast lifetime of WANETs with multibeam antennas: Formulation, algorithms, and analysis. IEEE Transactions on Computers, 63(8):1988–2001, August 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Garcia-Garcia:2018:CAF

[GGSPM18] Adrian Garcia-Garcia, Juan Carlos Saez, and Manuel Prieto-Matias. Contention-aware fair scheduling for asymmetric single-ISA multicore systems. IEEE Transactions on Computers, 67(12): 1703–1719, December 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8360522/.

Galal:2011:EEF

[GH11] Sameh Galal and Mark Horowitz. Energy-efficient floating-point unit design. IEEE Transactions on Computers, 60(7):913–922, July 2011. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2014:OFL

[GHG^{+}14] Shuo Guo, Liang He, Yu Gu, Bo Jiang, and Tian He. Opportunistic flooding in low-duty-cycle wireless sensor networks with unreliable links. IEEE Transactions on Computers, 63(11): 2787–2802, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gerards:2015:IBG

[GHK15] M. E. T. Gerards, J. L. Hurink, and J. Kuper. On the interplay between global DVFS and scheduling tasks with precedence constraints. IEEE Transactions on Computers, 64 (6):1742–1754, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gong:2017:NPA

[GHL17] Yifan Gong, Bingsheng He, and Dan Li. Network performance aware optimizations on IaaS clouds. IEEE Transactions on Computers, 66(4):672–687, April 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gupta:2019:NPM

[GIKR19] S. Gupta, M. Imani, H. Kaur, and T. S. Rosing. NNPIM:

A processing in-memory architecture for neural network acceleration. IEEE Transactions on Computers, 68(9): 1325–1337, September 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Giorgi:2012:PMC

[Gio12] Pascal Giorgi. On polynomial multiplication in Chebyshev basis. IEEE Transactions on Computers, 61(6):780–789, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guerrini:2018:RMR

[GIW18] Eleonora Guerrini, Laurent Imbert, and Théo Winterhalter. Randomized mixed-radix scalar multiplication. IEEE Transactions on Computers, 67 (3):418–431, March 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8031048/.

Ganesan:2014:AGM

[GJ14] Karthik Ganesan and Lizy Kurian John. Automatic generation of miniaturized synthetic proxies for target applications to efficiently design multicore processors. IEEE Transactions on Computers, 63(4):833–846, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gorgin:2015:CXH

[GJ15] S. Gorgin and G. Jaberipur. Comment on “High-Speed Parallel Decimal Multiplication With Redundant Internal Encodings”. IEEE Transactions on Computers, 64(1):293–294, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [HK13a].

Gaj:2010:ATE

[GKB^{+}10] K. Gaj, Soonhak Kwon, P. Baier, P. Kohlbrenner, Hoang Le, M. Khaleeluddin, R. Bachimanchi, and M. Rogawski. Area-time efficient implementation of the elliptic curve method of factoring in reconfigurable hardware for application in the number field sieve. IEEE Transactions on Computers, 59(9): 1264–1280, September 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5374380.

Gulino:2019:OBG

[GKC19] Andrea Gulino, Abdulrahman Kaitoua, and Stefano Ceri. Optimal binning for genomics. IEEE Transactions on Computers, 68 (1):125–138, January 2019.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8410020/.

Ganesh:2017:LOW

[GKD^{+}17] Kalidas Ganesh, Youngjae Kim, Monobrata Debnath, Sungyong Park, and Junghee Lee. LAWC: Optimizing write cache using layoutaware I/O scheduling for all flash storage. IEEE Transactions on Computers, 66(11): 1890–1902, November 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7932926/.

Gilani:2014:EEP

[GKS14] Syed Zohaib Gilani, Nam Sung Kim, and Michael Schulte. Energy-efficient pixel-arithmetic. IEEE Transactions on Computers, 63(8):1882–1894, August 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gu:2019:GRM

[GL19] Z. Gu and S. Li. A generalized RNS Mclaughlin modular multiplication with noncoprime moduli sets. IEEE Transactions on Computers, 68(11):1689–1696, November 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gu:2019:TML

[GLH^{+}19] Chongyan Gu, Weiqiang Liu, Neil Hanley, Robert Hesselbarth, and Máire O’Neill. A theoretical model to link uniqueness and min-entropy for PUF evaluations. IEEE Transactions on Computers, 68(2):287–293, February 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8444682/.

Gandino:2012:AAS

[GLP^{+}12] Filippo Gandino, Fabrizio Lamberti, Gianluca Paravati, Jean-Claude Bajard, and Paolo Montuschi. An algorithmic and architectural study on Montgomery exponentiation in RNS. IEEE Transactions on Computers, 61(8):1071–1083, August 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guneysu:2015:LBS

[GLP15] T. Guneysu, V. Lyubashevsky, and T. Poppelmann. Lattice-based signatures: Optimization and implementation on reconfigurable hardware. IEEE Transactions on Computers, 64(7):1954–1967, July 2015.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2016:ECP

[GLTC16] Peng Guo, Xuefeng Liu, Shaojie Tang, and Jiannong Cao. Enabling coveragepreserving scheduling in wireless sensor networks for structural health monitoring. IEEE Transactions on Computers, 65(8):2456– 2469, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gao:2013:SSG

[GLXY13] Lin Gao, Lian Li, Jingling Xue, and Pen-Chung Yew. SEED: A statically greedy and dynamically adaptive approach for speculative loop execution. IEEE Transactions on Computers, 62 (5):1004–1016, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gizopoulos:2011:GEI

[GM11] Dimitris Gizopoulos and Shubhendu Mukherjee. Guest Editors’ introduction: Special section on dependable computer architecture. IEEE Transactions on Computers, 60(1):3–4, January 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Golubitsky:2012:SOB

[GM12] Oleg Golubitsky and Dmitri Maslov. A study of optimal 4-bit reversible Toffoli circuits and their synthesis. IEEE Transactions on Computers, 61(9): 1341–1353, September 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Garcia:2015:SQC

[GM15] H. J. Garcia and I. L. Markov. Simulation of quantum circuits via stabilizer frames. IEEE Transactions on Computers, 64(8): 2323–2336, August 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gulliver:2014:FFL

[GNSR14] T. Aaron Gulliver, Stephen W. Neville, Ulrich Speidel, and Niko Rebenich. FLOTT — a fast, low memory T-transform algorithm for measuring string complexity. IEEE Transactions on Computers, 63(4):917–926, April 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gonzalez-Navarro:2013:BID

[GNTS13]  Sonia  Gonzalez-Navarro,

Charles Tsen, and Michael J. Schulte. Binary integer decimal-based floating-point multiplication. IEEE Transactions on Computers, 62 (7):1460–1466, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Granmo:2010:SSN

[GO10] O.-C. Granmo and B. J.

Oommen. Solving stochastic nonlinear resource allocation problems using a hierarchy of twofold resource allocation automata. IEEE Transactions on Computers, 59(4):545–560, April 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5374378.

Gordon:2014:WCP

[Gor14] Dan Gordon. The wellconnected processor array. IEEE Transactions on Computers, 63(5):1287– 1295, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Giefers:2014:FBR

[GP14] H. Giefers and M. Platzner.

An FPGA-based reconfigurable mesh many-core. IEEE Transactions on Computers, 63(12):2919–2932, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Garofalo:2011:ACM

[GPN11] Valeria Garofalo, Nicola Petra, and Ettore Napoli. Analytical calculation of the maximum error for a family of truncated multipliers providing minimum mean square error. IEEE Transactions on Computers, 60(9): 1366–1371, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5669274.

Gur:2019:PAI

[GPR^{+}19] Kamil D. Gür, Yuriy Polyakov, Kurt Rohloff, Gerard W. Ryan, Hadi Sajjadpour, and Erkay Savaş. Practical applications of improved Gaussian sampling for trapdoor lattices. IEEE Transactions on Computers, 68 (4):570–584, April 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8493319/.

Gaudesi:2017:NTR

[GPRS17] Marco Gaudesi, Irith Pomeranz, Matteo Sonza Reorda, and Giovanni Squillero. New techniques to reduce the execution time of functional test programs. IEEE Transactions on Computers, 66 (7):1268–1273, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07795264-
abs.html.

Govindan:2014:SPP

[GRL^{+}14] Madhu Saravana Sibi Govindan, Behnam Robatmili, Dong Li, Bertrand A. Maher, Aaron Smith, Stephen W. Keckler, and Doug Burger. Scaling power and performance via processor composability. IEEE Transactions on Computers, 63 (8):2025–2038, August 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gangopadhyay:2016:MBP

[GRM16] Dipanwita Gangopadhyay and Arash Reyhani-Masoleh. Multiple-bit parity-based concurrent fault detection architecture for parallel CRC computation. IEEE Transactions on Computers, 65 (7):2143–2157, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guilley:2010:EPC

[GSF^{+}10] S. Guilley, L. Sauvage, F. Flament, Vinh-Nga Vong, P. Hoogvorst, and R. Pacalet. Evaluation of power constant dual-rail logics countermeasures against DPA with design time security metrics. IEEE Transactions on Computers, 59(9): 1250–1263, September 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5467046.

Gu:2015:SCN

[GSG^{+}15] Feng Gu, K. Shaban, N. Ghani, S. Khan, M. R. Naeini, M. M. Hayat, and C. Assi. Survivable cloud network mapping for disaster recovery support. IEEE Transactions on Computers, 64(8):2353–2366, August 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gaiti:2014:TFT

[GSH^{+}14] Dominique Gaiti, Ahmad Sardouk, Abdelhakim Hafid, Ali El Masri, and Lyes Khoukhi. Toward fuzzy traffic adaptation solution in wireless mesh networks. IEEE Transactions on Computers, 63(5):1296–1308, May 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ganeshpure:2012:PGT

[GSK12] Kunal P. Ganeshpure, Alodeep Sanyal, and Sandip Kundu. A pattern generation technique for maximizing switching supply currents considering gate delays. IEEE Transactions on Computers, 61(7):986–998, July 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gupta:2010:CSM

[GSL10] M. Gupta, F. Sanchez, and J. Llosa. CSMT: Simultaneous multithreading for clustered VLIW processors. IEEE Transactions on Computers, 59 (3):385–399, March 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5161255.

Golubcovs:2013:CMA

[GSX^{+}13] Stanislavs Golubcovs, Delong Shang, Fei Xia, Andrey Mokhov, and Alex Yakovlev. Concurrent multiresource arbiter: Design and applications. IEEE Transactions on Computers, 62(1):31–44, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gallin:2019:GFP

[GT19] G. Gallin and A. Tisserand. Generation of finelypipelined GF(P P) multipliers for flexible curve based cryptography on FPGAs. IEEE Transactions on Computers, 68(11):1612–1622, November 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gonzalez-Toral:2018:SDC

[GTRMG18] Ricardo González-Toral, Pedro Reviriego, Juan Antonio Maestro, and Zhen Gao. A scheme to design concurrent error detection techniques for the Fast Fourier Transform implemented in SRAM-based FPGAs. IEEE Transactions on Computers, 67(7):1039–1045, July 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8258993/.

Ghosh:2014:BBB

[GV14] Santosh Ghosh and Ingrid Verbauwhede. BLAKE-512-based 128-bit CCA2 secure timing attack resistant McEliece cryptoprocessor. IEEE Transactions on Computers, 63(5):1124– 1133, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

George:2015:FTL

[GV15] C. George and S. Vadhiyar. Fault tolerance on large scale systems using adaptive process replication. IEEE Transactions on Computers, 64(8):2213–2225, August 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Garcia-Vega:2016:DMO

[GVGNCVM16] Carlos Garcia-Vega, Sonia Gonzalez-Navarro, Pedro Balboa-La Chica, and Julio Villalba-Moreno. Decimal multiformat online addition. IEEE Transactions on Computers, 65(10): 3203–3209, October 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gadioli:2019:MDA

[GVPS19] D. Gadioli, E. Vitali, G. Palermo, and C. Silvano. mARGOt: A dynamic autotuning framework for self-aware approximate computing. IEEE Transactions on Computers, 68(5):713–728, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gu:2016:PAO

[GW16] Y. Gu and C. Q. Wu. Performance analysis and optimization of distributed workflows in heterogeneous network environments. IEEE Transactions on Computers, 65(4):1266–1282, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guan:2017:BLL

[GWM^{+}17] Yong Guan, Guohui Wang, Chenlin Ma, Renhai Chen, Yi Wang, and Zili Shao. A block-level log-block management scheme for MLC NAND flash memory storage systems. IEEE Transactions on Computers, 66(9): 1464–1477, September 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7874142/.

Ganesh:2013:IAD

[GWMB13] Lakshmi Ganesh, Hakim Weatherspoon, Tudor Marian, and Ken Birman. Integrated approach to data center power management. IEEE Transactions on Computers, 62(6):1086–1096, June 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gu:2010:NWG

[GWZ^{+}10] Peng Gu, Jun Wang, Yifeng Zhu, Hong Jiang, and Pengju Shang. A novel weighted-graph-based grouping algorithm for metadata prefetching. IEEE Transactions on Computers, 59(1):1–15, January 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=
&arnumber=5184817.

Guo:2013:HSM

[GY13] Zhiyang Guo and Yuanyuan Yang. High-speed multicast scheduling in hybrid optical packet switches with guaranteed latency. IEEE Transactions on Computers, 62(10): 1972–1987, October 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gong:2014:LAA

[GY14] Dawei Gong and Yuanyuan Yang. On-line AP association algorithms for 802.11n WLANs with heterogeneous clients. IEEE Transactions on Computers, 63(11): 2772–2786, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2015:ESR

[GY15a] Zhiyang Guo and Yuanyuan Yang. Exploring server redundancy in nonblocking multicast data center networks. IEEE Transactions on Computers, 64(7):1912– 1926, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2015:NMF

[GY15b] Zhiyang Guo and Yuanyuan Yang. On nonblocking multicast fat-tree data center networks with server redundancy. IEEE Transactions on Computers, 64 (4):1058–1073, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gong:2016:LLM

[GY16] Dawei Gong and Yuanyuan Yang. Link-layer multicast in large-scale 802.11n wireless LANs with smart antennas. IEEE Transactions on Computers, 65 (7):2118–2133, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gao:2016:GOM

[GYC^{+}16] Xiaofeng Gao, Yongtian Yang, Guihai Chen, Xin Lu, and Jiaofei Zhong. Global optimization for multi-channel wireless data broadcast with AH-tree indexing scheme. IEEE Transactions on Computers, 65 (7):2104–2117, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gu:2015:OTP

[GZB^{+}15] Lin Gu, Deze Zeng, A. Barnawi, Song Guo, and I. Stojmenovic. Optimal task placement with QoS constraints in geo-distributed data centers using DVFS. IEEE Transactions on Computers, 64(7):2049–2059, July 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gao:2017:HPG

[GZC^{+}17] Pin Gao, Mingxing Zhang, Kang Chen, Yongwei Wu, and Weimin Zheng. High performance graph processing with locality oriented design. IEEE Transactions on Computers, 66 (7):1261–1267, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07815299-
abs.html.

Gu:2016:GCC

[GZG^{+}16] L. Gu, D. Zeng, S. Guo, Y. Xiang, and J. Hu. A general communication cost optimization framework for big data stream processing in geo-distributed data centers. IEEE Transactions on Computers, 65(1):19–29, January 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Harn:2013:GA

[Har13] Lein Harn. Group authentication. IEEE Transactions on Computers, 62(9): 1893–1898, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hadjidj:2011:ERA

[HB11] Rachid Hadjidj and Hanifa Boucheneb. Efficient reachability analysis for time petri nets. IEEE Transactions on Computers, 60 (8):1085–1099, August 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hussain:2014:GHE

[HBAD14] Zaid A. Hussain, Bella Bose, and Abdullah Al-Dhelaan. Generalized hypercubes: Edge-disjoint Hamiltonian cycles and Gray codes. IEEE Transactions on Computers, 63(2):375– 382, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

He:2013:HEH

[HBCC13] Daojing He, Jiajun Bu, Sammy Chan, and Chun Chen. Handauth: Efficient handover authentication with conditional privacy for wireless networks. IEEE Transactions on Computers, 62(3):616–622, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hyman:2011:RMS

[HBR11] Ransford Hyman, Koustav Bhattacharya, and Nagarajan Ranganathan. Redundancy mining for soft error detection in multicore processors. IEEE Transactions on Computers, 60 (8):1114–1125, August 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hakkarinen:2013:MDC

[HC13a] D. Hakkarinen and Zizhong Chen. Multilevel diskless checkpointing. IEEE Transactions on Computers, 62(4):772–783, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hsiao:2013:SLB

[HC13b] Hung-Chang Hsiao and Che-Wei Chang. A symmetric load balancing algorithm with performance guarantees for distributed hash tables. IEEE Transactions on Computers, 62(4):662– 675, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2017:ATE

[HC17] Jingwei Hu and Ray C. C.

Cheung. Area-time efficient computation of Niederreiter encryption on QC-MDPC codes for embedded hardware. IEEE Transactions on Computers, 66 (8):1313–1325, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7862221/.

Hu:2012:LTM

[HCC^{+}12] Weiwu Hu, Yunji Chen, Tianshi Chen, Cheng Qian, and Lei Li. Linear time memory consistency verification. IEEE Transactions on Computers, 61(4):502– 516, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2018:CWF

[HCC^{+}18] Li Han, Louis-Claude Canon, Henri Casanova, Yves Robert, and Frédéric Vivien. Checkpointing workflows for failstop errors. IEEE Transactions on Computers, 67 (8):1105–1120, August 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8279499/.

Hasan:2010:UBD

[HCCG10] Y. Hasan, Wei-Mei Chen, J. M. Chang, and B. M. Gharaibeh. Upper bounds for dynamic memory allocation. IEEE Transactions on Computers, 59 (4):468–477, April 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5282494.

He:2016:DDD

[HCD^{+}16] Kun He, Jing Chen, Ruiying Du, Qianhong Wu, Guoliang Xue, and Xiang Zhang. DeyPoS: Deduplicatable dynamic proof of storage for multi-user environments. IEEE Transactions on Computers, 65(12): 3631–3645, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2016:DHM

[HCG^{+}16] K. Hu, H. K. Chandrikakutty, Z. Goodman, R. Tessier, and T. Wolf. Dynamic hardware monitors for network processor protection. IEEE Transactions on Computers, 65 (3):860–872, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hong:2015:LDA

[HCH15] Shengyan Hong, T. Chantem, and Xiaobo Sharon Hu. Local-deadline assignment for distributed real-time systems. IEEE Transactions on Computers, 64(7):1983– 1997, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2014:SCA

[HCL^{+}14] Jie Han, Hao Chen, Jinghang Liang, Peican Zhu, Zhixi Yang, and Fabrizio Lombardi. A stochastic computational approach for accurate and efficient reliability evaluation. IEEE Transactions on Computers, 63 (6):1336–1350, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hsieh:2015:AES

[HCL15] J. Hsieh, C. Chen, and H. Lin. Adaptive ECC scheme for hybrid SSD’s. IEEE Transactions on Computers, 64(12):3348–3361, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hsin:2015:ACO

[HCSW15] Hsien-Kai Hsin, En-Jui Chang, Kuan-Yu Su, and An-Yeu Wu. Ant colony optimization-based adaptive network-on-chip routing framework using network information region. IEEE Transactions on Computers, 64(8):2119–2131, August 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2018:RRE

[HCY18] Hyunseung Han, Jaeyong Chung, and Joon-Sung Yang. READ: Reliability enhancement in 3D-memory exploiting asymmetric SER distribution. IEEE Transactions on Computers, 67 (8):1193–1201, August 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8283794/.

Huang:2013:TCT

[HCZW13] Minsu Huang, Siyuan Chen, Ying Zhu, and Yu Wang. Topology control for timeevolving and predictable delay-tolerant networks. IEEE Transactions on Computers, 62(11):2308–2321, November 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2016:EET

[HDYS16] J. Huang, Y. Deng, Q. Yang, and J. Sun. An energyefficient train control framework for smart railway transportation. IEEE Transactions on Computers, 65 (5):1407–1417, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2015:SPB

[HF15] Jiangtao Han and Haining Fan. Shifted polynomial basis multipliers based on subquadratic Toeplitz matrix-vector product approach for all irreducible pentanomials. IEEE Transactions on Computers, 64 (3):862–867, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

He:2017:FPH

[HFG^{+}17] Conghui He, Haohuan Fu, Ce Guo, Wayne Luk, and Guangwen Yang. A fullypipelined hardware design for Gaussian mixture models. IEEE Transactions on Computers, 66(11):1837– 1850, November 2017. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7938761/.

Hao:2013:CDA

[HFZ13] Rong-Xia Hao, Yan-Quan Feng, and Jin-Xin Zhou. Conditional diagnosability of alternating group graphs. IEEE Transactions on Computers, 62(4):827–831, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Herrero:2013:TRB

[HGCT13] Enric Herrero, Jose Gonzalez, Ramon Canal, and Dean Tullsen. Thread row buffers: Improving memory performance isolation and throughput in multiprogrammed environments. IEEE Transactions on Computers, 62(9): 1879–1892, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2011:NHA

[HGEG11] Miaoqing Huang, Kris Gaj, and Tarek El-Ghazawi. New hardware architectures for Montgomery modular multiplication algorithm. IEEE Transactions on Computers, 60(7):923–936, July 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2015:JOR

[HGL^{+}15] H. Huang, S. Guo, P. Li, B. Ye, and I. Stojmenovic. Joint optimization of rule placement and traffic engineering for QoS provisioning in software defined network. IEEE Transactions on Computers, 64(12): 3488–3499, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huynh:2011:RRR

[HGML11] Minh Huynh, Stuart Goose, Prasant Mohapatra, and Raymond Liao. RRR: Rapid ring recovery submillisecond decentralized recovery for Ethernet ring. IEEE Transactions on Computers, 60 (11):1561–1570, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.

                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5560640.

Hua:2017:SAR

[HGW^{+}17] Xiayu Hua, Chunhui Guo, Hao Wu, Douglas Lautner, and Shangping Ren. Schedulability analysis for realtime task set on resource with performance degradation and dual-level periodic rejuvenations. IEEE Transactions on Computers, 66 (3):553–559, March 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hsu:2017:MIS

[HH17] Po-Cheng Hsu and Sun-Yuan Hsieh. Multi-inherited search tree for dynamic IP router-tables. IEEE Transactions on Computers, 66 (1):114–126, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2018:ACO

[HHC^{+}18] Kai Huang, Biao Hu, Long Chen, Alois Knoll, and Zhihua Wang. Adas on Cots with OpenCL: A case study with lane detection. IEEE Transactions on Computers, 67(4):559–565, April 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8057795/.

Hsu:2011:DET

[HHCH11] Chun-Lung Hsu, Yu-Sheng Huang, Ming-Da Chang, and Hung-Yen Huang. Design of an error-tolerance scheme for discrete wavelet transform in JPEG 2000 encoder. IEEE Transactions on Computers, 60(5):628–638, May 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Houshmand:2012:MMR

[HHKW12] Monireh Houshmand, Saied Hosseini-Khayat, and Mark M. Wilde. Minimal-memory requirements for pearl-necklace encoders of quantum convolutional codes. IEEE Transactions on Computers, 61 (3):299–312, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hsiu:2012:MBO

[HHLK12] Pi-Cheng Hsiu, Cheng-Kang Hsieh, Der-Nien Lee, and Tei-Wei Kuo. Multilayer bus optimization for real-time embedded systems. IEEE Transactions on Computers, 61(11):1638–1650, November 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hefeeda:2011:DEP

[HHM11] Mohamed Hefeeda, Cheng-Hsin Hsu, and Kianoosh Mokhtarian. Design and evaluation of a proxy cache for peer-to-peer traffic. IEEE Transactions on Computers, 60(7):964–977, July 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2018:ALB

[HHW^{+}18] Dan Huang, Dezhi Han, Jun Wang, Jiangling Yin, Xunchao Chen, Xuhong Zhang, Jian Zhou, and Mao Ye. Achieving load balance for parallel data access on distributed file systems. IEEE Transactions on Computers, 67(3):388–402, March 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8027054/.

Hsieh:2011:MTN

[HHY11] Sun-Yuan Hsieh, Yi-Ling Huang, and Ying-Chi Yang. Multiprefix trie: a new data structure for designing dynamic router-tables. IEEE Transactions on Computers, 60(5):693–706, May 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hiasat:2016:SDG

[Hia16] Ahmad Hiasat. A sign detector for a group of threemoduli sets. IEEE Transactions on Computers, 65(12): 3580–3590, December 2016.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hiasat:2017:ERS

[Hia17] Ahmad Hiasat. Efficient RNS scalers for the extended three-moduli set (2^{n} − 1, 2^{n+p}, 2^{n} + 1). IEEE Transactions on Computers, 66(7):1253–1260, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07815319-
abs.html.

Hierons:2011:CTN

[Hie11] Robert M. Hierons. Controllable testing from nondeterministic finite state machines with multiple ports. IEEE Transactions on Computers, 60(12):1818–1822, December 2011. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hierons:2013:VCF

[Hie13] Robert M. Hierons. Verifying and comparing finite state machines for systems that have distributed interfaces. IEEE Transactions on Computers, 62 (8):1673–1683, August 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hao:2019:IDP

[HIJ^{+}19] Y. Hao, T. Isobe, L. Jiao, C. Li, W. Meier, Y. Todo, and Q. Wang. Improved division property based cube attacks exploiting algebraic properties of Superpoly. IEEE Transactions on Computers, 68(10):1470–1486, October 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hollis:2014:EEC

[HJBM14] Simon J. Hollis, Chris Jackson, Paul Bogdan, and Radu Marculescu. Exploiting emergence in on-chip interconnects. IEEE Transactions on Computers, 63 (3):570–582, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2013:EEM

[HJF^{+}13] Yang Hu, Hong Jiang, Dan Feng, Lei Tian, Hao Luo, and Chao Ren. Exploring and exploiting the multilevel parallelism inside SSDs for improved performance and endurance. IEEE Transactions on Computers, 62 (6):1141–1155, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2013:HSP

[HK13a] Liu Han and Seok-Bum Ko. High-speed parallel decimal multiplication with redundant internal encodings. IEEE Transactions on Computers, 62(5):956– 968, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See comment [GJ15].

Hsieh:2013:CDK

[HK13b] Sun-Yuan Hsieh and Chi-Ya Kao. The conditional diagnosability of k-ary ncubes under the comparison diagnosis model. IEEE Transactions on Computers, 62(4):839–843, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hong:2015:LCM

[HK15a] Seokin Hong and Soontae Kim. A low-cost mechanism exploiting narrow-width values for tolerating hard faults in ALU. IEEE Transactions on Computers, 64(9): 2433–2446, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hsieh:2015:DDC

[HK15b] J. Hsieh and Y. Kuan.

DCCS: Double circular caching scheme for DRAM/ PRAM hybrid cache. IEEE Transactions on Computers, 64(11):3115–3127, November 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hong:2016:DRL

[HK16] Seokin Hong and Soontae Kim. Designing a resilient L1 cache architecture to process variation-induced access-time failures. IEEE Transactions on Computers, 65(10):2999–3012, October 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hong:2017:SEA

[HK17] Jeongkyu Hong and Soontae Kim. Smart ECC allocation cache utilizing cache data space. IEEE Transactions on Computers, 66 (2):368–374, February 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Howe:2018:PDG

[HKR^{+}18] James Howe, Ayesha Khalid, Ciara Rafferty, Francesco Regazzoni, and Máire O’Neill. On practical discrete Gaussian samplers for latticebased cryptography. IEEE Transactions on Computers, 67(3):322–334, March 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7792671/.

Hsu:2014:JDA

[HKWC14] Chih-Cheng Hsu, Ming-Shing Kuo, Shi-Chen Wang, and Cheng-Fu Chou. Joint design of asynchronous sleepwake scheduling and opportunistic routing in wireless sensor networks. IEEE Transactions on Computers, 63(7):1840–1846, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Harn:2010:AGK

[HL10a] L. Harn and Changlu Lin.

Authenticated group key transfer protocol based on secret sharing. IEEE Transactions on Computers, 59 (6):842–846, June 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5416683.

Huang:2010:ASR

[HL10b] Chin-Yu Huang and Chu-Ti Lin. Analysis of software reliability modeling considering testing compression factor and failure-to-fault relationship. IEEE Transactions on Computers, 59 (2):283–288, February 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5372166.

Huang:2017:EER

[HLA^{+}17] Jing Huang, Renfa Li, Jiyao An, Derrick Ntalasha, Fan Yang, and Keqin Li. Energyefficient resource utilization for heterogeneous embedded computing systems. IEEE Transactions on Computers, 66(9):1518–1531, September 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7896542/.

Huang:2019:CVR

[HLC^{+}19] D. Huang, Q. Liu, J. Choi, N. Podhorszki, S. Klasky, J. Logan, G. Ostrouchov, X. He, and M. Wolf. Can I/O variability be reduced on QoS-less HPC storage systems? IEEE Transactions on Computers, 68(5):631– 645, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hua:2014:DSA

[HLF14] Yu Hua, Xue Liu, and Dan Feng. Data similarityaware computation infrastructure for the cloud. IEEE Transactions on Computers, 63(1):3–16, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hua:2014:ASA

[HLJ14] Yu Hua, Xue Liu, and Hong Jiang. ANTELOPE:

A semantic-aware data cube scheme for cloud data center networks. IEEE Transactions on Computers, 63(9): 2146–2159, September 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2015:CEA

[HLT^{+}15] Xinyi Huang, J. K. Liu, Shaohua Tang, Yang Xiang, Kaitai Liang, Li Xu, and Jianying Zhou. Costeffective authentic and anonymous data sharing with forward security. IEEE Transactions on Computers, 64(4):971–983, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2017:AST

[HLWV17] Menglan Hu, Jun Luo, Yang Wang, and Bharadwaj Veeravalli. Adaptive scheduling of task graphs with dynamic resilience. IEEE Transactions on Computers, 66(1):17–23, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hsieh:2014:MCA

[HLY14] Jen-Wei Hsieh, Han-Yi Lin, and Dong-Lin Yang. Multichannel architecture-based FTL for reliable and highperformance SSD. IEEE Transactions on Computers, 63(12):3079–3091, December 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Homma:2010:CPA

[HMA^{+}10] N. Homma, A. Miyamoto, T. Aoki, A. Satoh, and A. Samir. Comparative power analysis of modular exponentiation algorithms. IEEE Transactions on Computers, 59(6):795–807, June 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5342411.

Hong:2011:EOS

[HMC11] Wonhak Hong, Rajashekhar Modugu, and Minsu Choi. Efficient online self-checking modulo 2^{n} + 1 multiplier design. IEEE Transactions on Computers, 60(9): 1354–1365, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5962404.

Hojabr:2017:CCN

[HMD^{+}17] Reza Hojabr, Mehdi Modarressi, Masoud Daneshtalab, Ali Yasoubi, and Ahmad Khonsari. Customizing Clos network-on-chip for neural networks. IEEE Transactions on Computers, 66(11): 1865–1877, November 2017.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7948744/.

Hasan:2012:BRA

[HMNN12] M. Anwar Hasan, Nicolas Meloni, Ashkan Hosseinzadeh Namin, and Christophe Negre. Block recombination approach for subquadratic space complexity binary field multiplication based on Toeplitz matrix– vector product. IEEE Transactions on Computers, 61 (2):151–163, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hormigo:2017:ISI

[HMO^{+}17] Javier Hormigo, Jean-Michel Muller, Stuart Oberman, Nathalie Revol, Arnaud Tisserand, and Julio Villalba-Moreno. Introduction to the special issue on computer arithmetic. IEEE Transactions on Computers, 66(12): 1991–1993, December 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Haghbayan:2017:PRA

[HMR^{+}17] Mohammad-Hashem Haghbayan, Antonio Miele, Amir M. Rahmani, Pasi Liljeberg, and Hannu Tenhunen. Performance/reliability-aware resource management for many-cores in dark silicon era. IEEE Transactions on Computers, 66(9): 1599–1612, September 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7892847/.

Huang:2012:LCB

[HMS^{+}12] Libo Huang, Sheng Ma, Li Shen, Zhiying Wang, and Nong Xiao. Low-cost Binary128 floating-point FMA unit design with SIMD support. IEEE Transactions on Computers, 61(5):745– 751, May 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hao:2014:TWS

[HMZ^{+}14] Zhuo Hao, Yunlong Mao, Sheng Zhong, Li Erran Li, Haifan Yao, and Nenghai Yu. Toward wireless security without computational assumptions — oblivious transfer based on wireless channel characteristics. IEEE Transactions on Computers, 63(6):1580– 1593, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hasan:2011:LSC

[HN11] M. Anwar Hasan and Christophe Negre. Low space complexity multiplication over binary fields with Dickson polynomial representation. IEEE Transactions on Computers, 60(4):602–607, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hasan:2013:MSM

[HN13] M. Anwar Hasan and Christophe Negre. Multiway splitting method for Toeplitz matrix vector product. IEEE Transactions on Computers, 62 (7):1467–1471, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See comment [PCHS16].

Huang:2012:EMT

[HNB^{+}12] Miaoqing Huang, Vikram K.

Narayana, Mohamed Bakhouya, Jaafar Gaber, and Tarek El-Ghazawi. Efficient mapping of task graphs onto reconfigurable hardware using architectural variants. IEEE Transactions on Computers, 61(9):1354–1360, September 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hassan:2019:FEE

[HNV19] A. Hassan, D. Nikolopoulos, and H. Vandierendonck. Fast and energy-efficient OLAP data management on hybrid main memory systems. IEEE Transactions on Computers, 68(11):1597–1611, November 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hassan:2016:RTC

[HPR16] Ahmed Hassan, Roberto Palmieri, and Binoy Ravindran. Remote transaction commit: Centralizing software transactional memory commits. IEEE Transactions on Computers, 65 (7):2228–2240, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2015:ERB

[HQLX15] J. Huang, X. Qin, X. Liang, and C. Xie. An efficient I/O-redirection-based reconstruction scheme for erasurecoded storage clusters. IEEE Transactions on Computers, 64(11):3037–3050, November 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hajisheykhi:2017:BAR

[HRK17] Reza Hajisheykhi, Mohammad Roohitavaf, and Sandeep S. Kulkarni. Bounded auditable restoration of distributed systems. IEEE Transactions on Computers, 66(2):240–255, February 2017. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hariri:2011:CED

[HRM11] Arash Hariri and Arash Reyhani-Masoleh. Concurrent error detection in Montgomery multiplication over binary extension fields. IEEE Transactions on Computers, 60(9):1341–1353, September 2011. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5669280.

Haghbayan:2016:PAA

[HRM^{+}16] M. Haghbayan, A. Rahmani, A. Miele, M. Fattah, J. Plosila, P. Liljeberg, and H. Tenhunen. A poweraware approach for online test scheduling in manycore architectures. IEEE Transactions on Computers, 65(3):730–743, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

He:2018:CED

[HS18] Shuibing He and Xian-He Sun. A cost-effective distribution-aware data replication scheme for parallel I/O systems. IEEE Transactions on Computers, 67(10): 1374–1387, October 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8352821/.

Homma:2014:TFD

[HSA14] N. Homma, K. Saito, and T. Aoki. Toward formal design of practical cryptographic hardware based on Galois field arithmetic. IEEE Transactions on Computers, 63(10):2604–2613, October 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

He:2010:IEC

[HSH^{+}10] Fei He, Xiaoyu Song, W. N. N. Hung, Ming Gu, and Jiaguang Sun. Integrating evolutionary computation with abstraction refinement for model checking. IEEE Transactions on Computers, 59(1):116– 126, January 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5184807.

Han:2014:IBS

[HSM14] Jinguang Han, Willy Susilo, and Yu Mu. Identity-based secure distributed data storage schemes. IEEE Transactions on Computers, 63 (4):941–953, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2012:MEC

[HT12] Chen-Wei Huang and Shiao-Li Tsao. Minimizing energy consumption of embedded systems via optimal code layout. IEEE Transactions on Computers, 61 (8):1127–1139, August 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hierons:2016:PAG

[HT16] Robert M. Hierons and Uraz Cengiz Türker. Parallel algorithms for generating harmonised state identifiers and characterising sets. IEEE Transactions on Computers, 65(11): 3370–3383, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hasan:2010:FRA

[HTA10] O. Hasan, S. Tahar, and N. Abbasi. Formal reliability analysis using theorem proving. IEEE Transactions on Computers, 59(5): 579–592, May 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5306067.

Ha:2017:WFP

[HTA17] Phuong Hoai Ha, Philippas Tsigas, and Otto J. Anshus. Wait-free programming for general purpose computations on graphics processors. IEEE Transactions on Computers, 66 (8):1407–1420, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/6357187/.

Hsieh:2013:SDC

[HTC13] Sun-Yuan Hsieh, Cheng-Yen Tsai, and Chun-An Chen. Strong diagnosability and conditional diagnosability of multiprocessor systems and folded hypercubes. IEEE Transactions on Computers, 62(7):1472–1477, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hamdioui:2015:TOD

[HTH15] S. Hamdioui, M. Taouil, and N. Z. Haron. Testing open defects in memristor-based memories. IEEE Transactions on Computers, 64 (1):247–259, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hanumaiah:2012:TAD

[HV12] Vinay Hanumaiah and Sarma Vrudhula. Temperatureaware DVFS for hard realtime applications on multicore processors. IEEE Transactions on Computers, 61(10):1484–1494, October 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2013:RAS

[HV13] Menglan Hu and Bharadwaj Veeravalli. Requirementaware scheduling of bag-of-tasks applications on grids with dynamic resilience. IEEE Transactions on Computers, 62(10):2108–2114, October 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hanumaiah:2014:EEO

[HV14a] Vinay Hanumaiah and Sarma Vrudhula. Energy-efficient operation of multicore processors by DVFS, task migration, and active cooling. IEEE Transactions on Computers, 63(2):349–360, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2014:DSH

[HV14b] Menglan Hu and B. Veeravalli. Dynamic scheduling of hybrid real-time tasks on clusters. IEEE Transactions on Computers, 63(12): 2988–2997, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hormigo:2016:NFC

[HV16] Javier Hormigo and Julio Villalba. New formats for computing with realnumbers under round-to-nearest. IEEE Transactions on Computers, 65(7):2158– 2168, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hormigo:2013:MRA

[HVZ13] Javier Hormigo, Julio Villalba, and Emilio L. Zapata. Multioperand redundant adders on FPGAs. IEEE Transactions on Computers, 62(10):2013–2025, October 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hsiao:2017:HMF

[HWCH17] Shen-Fu Hsiao, Chia-Sheng Wen, Yi-Hau Chen, and Kuei-Chun Huang. Hierarchical multipartite function evaluation. IEEE Transactions on Computers, 66 (1):89–99, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2016:NBT

[HWE^{+}16] L. Huang, J. Wang, M. Ebrahimi, M. Daneshtalab, X. Zhang, G. Li, and A. Jantsch. Nonblocking testing for networkon-chip. IEEE Transactions on Computers, 65(3):679– 692, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2014:CAA

[HWG^{+}14] Ruimin Hu, Xiaochen Wang, Jichang Guo, Qiang Liu, Tao Zhang, and Haojun Quan. Comments on “Algorithmic Aspects of Hardware/ Software Partitioning: 1D Search Algorithms”. IEEE Transactions on Computers, 63(4):1055–1056, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [JSC10].

Hsu:2015:CWF

[HWK15] A. C.-C. Hsu, D. S. L. Wei, and C.-C. J. Kuo. Coexistence Wi-Fi MAC design for mitigating interference caused by collocated Bluetooth. IEEE Transactions on Computers, 64 (2):342–352, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hsu:2014:LTP

[HWL^{+}14] Hong-Chun Hsu, Kuang-Shyr Wu, Cheng-Kuan Lin, Chiou-Yng Lee, and Chien-Ping Chang. A linear time pessimistic diagnosis algorithm for hypermesh multiprocessor systems under the PMC model. IEEE Transactions on Computers, 63(12): 2894–2904, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

He:2017:HAC

[HWS^{+}17] Shuibing He, Yang Wang, Xian-He Sun, Chuanhe Huang, and Chenzhong Xu. Heterogeneity-aware collective I/O for parallel I/O systems with hybrid HDD/ SSD servers. IEEE Transactions on Computers, 66 (6):1091–1098, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7778155/.

Huang:2015:CCF

[HWSN15] Zhen Huang, Cheng Wang, M. Stojmenovi, and A. Nayak. Characterization of cascading failures in interdependent cyber-physical systems. IEEE Transactions on Computers, 64(8):2158–2168, August 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

He:2017:HOP

[HWSX17] Shuibing He, Yang Wang, Xian-He Sun, and Chengzhong Xu. HARL: Optimizing parallel file systems with heterogeneity-aware regionlevel data layout. IEEE Transactions on Computers, 66(6):1048–1060, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7779011/.

Huang:2015:RMR

[HWX15] Pei Huang, Chen Wang, and Li Xiao. RC-MAC: A receiver-centric MAC protocol for event-driven wireless sensor networks. IEEE Transactions on Computers, 64(4):1149–1161, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2012:SAE

[HWZ^{+}12] Jian-Jun Han, Xiaodong Wu, Dakai Zhu, Hai Jin, Laurence T. Yang, and Jean-Luc Gaudiot. Synchronizationaware energy management for VFI-based multicore realtime systems. IEEE Transactions on Computers, 61 (12):1682–1696, December 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2017:OLA

[HWZ^{+}17] X. Hu, X. Wang, L. Zhou, Y. Luo, C. Ding, S. Jiang, and Z. Wang. Optimizing locality-aware memory management of key–value caches. IEEE Transactions on Computers, 66(5):862– 875, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2011:EIB

[HXL11] Xinfa Hu, Guoliang Xing, and Joseph Y.-T. Leung. Exploring the interplay between computation and communication in distributed real-time scheduling. IEEE Transactions on Computers, 60(12):1759–1771, December 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5560641.

Hua:2012:LSB

[HXVF12] Yu Hua, Bin Xiao, Bharadwaj Veeravalli, and Dan Feng. Locality-sensitive Bloom filter for approximate membership query. IEEE Transactions on Computers, 61(6):817–830, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Haghighi:2015:STD

[HXVQ15] M. S. Haghighi, Yang Xiang, V. Varadharajan, and B. Quinn. A stochastic time-domain model for burst data aggregation in IEEE 802.15.4 wireless sensor networks. IEEE Transactions on Computers, 64 (3):627–639, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hsieh:2012:CMT

[HY12] Sun-Yuan Hsieh and Ying-Chi Yang. A classified multisuffix trie for IP lookup and update. IEEE Transactions on Computers, 61(5): 726–731, May 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Heys:2011:PSC

[HZ11] Howard M. Heys and Liang Zhang. Pipelined statistical cipher feedback: a new mode for high-speed self-synchronizing stream encryption. IEEE Transactions on Computers, 60(11): 1581–1595, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5499465.

Han:2016:TSM

[HZL^{+}16] K. Han, C. Zhang, J. Luo, M. Hu, and B. Veeravalli. Truthful scheduling mechanisms for powering mobile crowdsensing. IEEE Transactions on Computers, 65 (1):294–307, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2012:JTA

[HZW^{+}12] H. Howie Huang, Nan Zhang, Wei Wang, Gautam Das, and Alexander S. Szalay. Just-in-time analytics on large file systems. IEEE Transactions on Computers, 61(11):1651–1664, November 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2014:SOC

[HZX^{+}14] Jingtong Hu, Qingfeng Zhuge, Chun Jason Xue, Wei-Che Tseng, Shouzhen Gu, and Edwin H.-M. Sha. Scheduling to optimize cache utilization for non-volatile main memories. IEEE Transactions on Computers, 63(8):2039–2051, August 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ishebabi:2010:HFC

[IB10] Harold Ishebabi and Christophe Bobda. Heuristics for flexible CMP synthesis. IEEE Transactions on Computers, 59 (8):1091–1104, August 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5453336.

Iturbe:2013:RNR

[IBH^{+}13] Xabier Iturbe, Khaled Benkrid, Chuan Hong, Ali Ebrahim, Raul Torrego, Imanol Martinez, Tughrul Arslan, and Jon Perez. R3TOS: A novel reliable reconfigurable realtime operating system for highly adaptive, efficient, and dependable computing on FPGAs. IEEE Transactions on Computers, 62 (8):1542–1556, August 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ibrahim:2016:IWR

[Ibr16] Walid Ibrahim. Identifying the worst reliability input vectors and the associated critical logic gates. IEEE Transactions on Computers, 65(6):1748–1760, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ishii:2017:FMA

[IDG^{+}17] Masahiro Ishii, Jérémie Detrey, Pierrick Gaudry, Atsuo Inomata, and Kazutoshi Fujikawa. Fast modular arithmetic on the Kalray MPPA-256 processor for an energy-efficient implementation of ECM. IEEE Transactions on Computers, 66(12): 2019–2030, December 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7927487/.

Ingelsson:2015:AFT

[IGLM15] U. Ingelsson, S. K. Goel, E. Larsson, and E. J.

Marinissen. Abort-on-fail test scheduling for modular SOCs without and with preemption. IEEE Transactions on Computers, 64(12): 3335–3347, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Iqbal:2016:DCA

[IHR^{+}16] Muhammad Faisal Iqbal, Jim Holt, Jee Ho Ryoo, Gustavo de Veciana, and Lizy K. John. Dynamic core allocation and packet scheduling in multicore network processors. IEEE Transactions on Computers, 65(12): 3646–3660, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ikodinovic:2015:MCA

[Iko15] I. Ikodinovic. Methodology for cycle-accurate DRAM performance analysis. IEEE Transactions on Computers, 64(7):2084–2091, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Imana:2018:FBP

[Ima18] José L. Imaña. Fast bitparallel binary multipliers based on Type-I pentanomials. IEEE Transactions on Computers, 67(6):898–904, June 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8125152/.

Ilic:2017:BRC

[IPS17] Aleksandar Ilic, Frederico Pratas, and Leonel Sousa. Beyond the roofline: Cacheaware power and energyefficiency modeling for multicores. IEEE Transactions on Computers, 66(1):52–58, January 2017. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Iakymchuk:2016:VDT

[IRMM^{+}16] Taras Iakymchuk, Alfredo Rosado-Muñoz, Manuel Bataller Mompéan, José Vicente Francés Víllora, and Emmanuel Ovie Osimiry. Versatile direct and transpose matrix multiplication with chained operations: An optimized architecture using circulant matrices. IEEE Transactions on Computers, 65(11): 3470–3479, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Im:2011:FAR

[IS11] Soojun Im and Dongkun Shin. Flash-aware RAID techniques for dependable and high-performance flash memory SSD. IEEE Transactions on Computers, 60 (1):80–92, January 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Islam:2014:CES

[IS14] Mafijul Md. Islam and Per Stenstrom. Characterizing and exploiting small-value memory instructions. IEEE Transactions on Computers, 63(7):1640–1655, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ibrahim:2015:AEE

[ISC15] W. Ibrahim, M. Shousha, and J. W. Chinneck. Accurate and efficient estimation of logic circuits reliability bounds. IEEE Transactions on Computers, 64 (5):1217–1229, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jongerius:2018:AMC

[JAD^{+}18] Rik Jongerius, Andreea Anghel, Gero Dittmann, Giovanni Mariani, Erik Vermij, and Henk Corporaal. Analytic multi-core processor model for fast designspace exploration. IEEE Transactions on Computers, 67(6):755–770, June 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8168422/.

Jawad:2015:JJT

[JAJK15] M. Jawad, S. M. Ali, J. A.

Jorgenson, and S. U. Khan. JEM: Just in time/just enough energy management methodology for computing systems. IEEE Transactions on Computers, 64 (6):1798–1804, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jadidi:2018:PPE

[JAKD18] Amin Jadidi, Mohammad Arjomand, Mahmut T. Kandemir, and Chita R. Das. Performance and power-efficient design of dense non-volatile cache in CMPs. IEEE Transactions on Computers, 67 (7):1054–1061, July 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8265206/.

Jin:2015:HSH

[JAS^{+}15] Seongwook Jin, Jeongseob Ahn, Jinho Seol, Sanghoon Cha, Jaehyuk Huh, and Seungryoul Maeng. H-SVM: Hardware-assisted secure virtual machines under a vulnerable hypervisor. IEEE Transactions on Computers, 64(10):2833– 2846, October 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jin:2011:MDL

[JC11] Lei Jin and Sangyeun Cho. Macro data load: An efficient mechanism for enhancing loaded data reuse. IEEE Transactions on Computers, 60(4):526–537, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jaiswal:2012:FBH

[JC12] Manish Kumar Jaiswal and Nitin Chandrachoodan. FPGA-based high-performance and scalable block LU decomposition architecture. IEEE Transactions on Computers, 61(1):60–72, January 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jerez:2015:LCS

[JCK15] J. L. Jerez, G. A. Constantinides, and E. C. Kerrigan. A low complexity scaling method for the Lanczos kernel in fixed-point arithmetic. IEEE Transactions on Computers, 64 (2):303–315, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2016:PIA

[JCM16] Tao Jiang, Xiaofeng Chen, and Jianfeng Ma. Public integrity auditing for shared dynamic cloud data with group user revocation. IEEE Transactions on Computers, 65(8):2363–2373, August 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jang:2013:EAS

[JCY^{+}13] Hyung Beom Jang, Jinhang Choi, Ikroh Yoon, Sung-Soo Lim, Seungwon Shin, Naehyuck Chang, and Sung Woo Chung. Exploiting application/system-dependent ambient temperature for accurate microarchitectural simulation. IEEE Transactions on Computers, 62(4):705– 715, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jarvinen:2015:GAC

[JDA15] K. Jarvinen, V. Dimitrov, and R. Azarderakhsh. A generalization of addition chains and fast inversions in binary fields. IEEE Transactions on Computers, 64(9): 2421–2432, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jafri:2016:TTB

[JDA^{+}16] Syed M. A. H. Jafri, Masoud Daneshtalab, Naeem Abbas, Guillermo Serrano Leon, and Ahmed Hemani. TransMap: Transformation based remapping and parallelism for high utilization and energy efficiency in CGRAs. IEEE Transactions on Computers, 65(11): 3456–3469, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jessa:2015:QRS

[Jes15] M. Jessa. On the quality of random sequences produced with a combined random bit generator. IEEE Transactions on Computers, 64(3):791–804, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jalili:2014:NPD

[JGG^{+}14] Mahdi Jalili, Joobin Gharibshah, Seyed Morsal Ghavami, Mohammadreza Beheshtifar, and Reza Farshi. Nationwide prediction of drought conditions in Iran based on remote sensing data. IEEE Transactions on Computers, 63(1):90–101, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jeong:2011:APL

[JGHD11] Jaehoon Jeong, Shuo Guo, Tian He, and David H. C. Du. Autonomous passive localization algorithm for road sensor networks. IEEE Transactions on Computers, 60(11):1622–1637, November 2011. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5669282.

Jha:2013:CXM

[Jha13] Pranava K. Jha. Comments on “Multiple-Radix Gray Codes in Lee Metric”. IEEE Transactions on Computers, 62(1):200, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [ABA07].

Jiang:2016:ARB

[JHQL16] Honglan Jiang, Jie Han, Fei Qiao, and Fabrizio Lombardi. Approximate radix-8 Booth multipliers for lowpower and high-performance operation. IEEE Transactions on Computers, 65 (8):2638–2644, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ju:2014:PAM

[JJC14] Miao Ju, Hun Jung, and Hao Che. A performance analysis methodology for multicore, multithreaded processors. IEEE Transactions on Computers, 63(2): 276–289, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jang:2011:ERC

[JJK^{+}11] Jae-Wan Jang, Myeongjae Jeon, Hyo-Sil Kim, Heeseung Jo, Jin-Soo Kim, and Seungryoul Maeng. Energy reduction in consolidated servers through memory-aware virtual machine scheduling. IEEE Transactions on Computers, 60(4):552–564, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jing:2016:EEE

[JJZ^{+}16] N. Jing, L. Jiang, T. Zhang, C. Li, F. Fan, and X. Liang. Energy-efficient eDRAM-Based on-chip storage architecture for GPGPUs. IEEE Transactions on Computers, 65(1):122–135, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jarray:2015:CEM

[JK15] A. Jarray and A. Karmouch. Cost-efficient mapping for fault-tolerant virtual networks. IEEE Transactions on Computers, 64 (3):668–681, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Joardar:2019:LBA

[JKD^{+}19] B. K. Joardar, R. G. Kim, J. R. Doppa, P. P. Pande, D. Marculescu, and R. Marculescu. Learning-based application-agnostic 3D NoC design for heterogeneous manycore systems. IEEE Transactions on Computers, 68(6):852–866, June 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jo:2010:TFT

[JKJ^{+}10] Heeseung Jo, Hwanju Kim, Jae-Wan Jang, Joonwon Lee, and Seungryoul Maeng. Transparent fault tolerance of device drivers for virtual machines. IEEE Transactions on Computers, 59(11): 1466–1479, November 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5432158.

Jeannerod:2011:CFP

[JKMR11] Claude-Pierr Jeannerod, Hervé Knochel, Christophe Monat, and Guillaume Revy. Computing floating-point square roots via bivariate polynomial evaluation. IEEE Transactions on Computers, 60(2):214–227, February 2011. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jin:2010:DAC

[JKY10] Yuho Jin, Eun Jung Kim, and Ki Hwan Yum. Design and analysis of onchip networks for large-scale cache systems. IEEE Transactions on Computers, 59 (3):332–344, March 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5235140.

Jankowski:2011:PAG

[JL11] Krzysztof Jankowski and Pierre Laurent. Packed AES-GCM algorithm suitable for AES/PCLMULQDQ instructions. IEEE Transactions on Computers, 60 (1):135–138, January 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jeon:2010:LUS

[JLC10] Hyeran Jeon, Woo Hyong Lee, and Sung Woo Chung. Load unbalancing strategy for multicore embedded processors. IEEE Transactions on Computers, 59(10): 1434–1440, October 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5374370.

Jeong:2019:OCG

[JLKR19] I. Jeong, C. Lee, K. Kim, and W. W. Ro. OverCome: Coarse-grained instruction commit with handover register renaming. IEEE Transactions on Computers, 68 (12):1802–1816, December 2019. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2019:LPU

[JLLH19] H. Jiang, L. Liu, F. Lombardi, and J. Han. Lowpower unsigned divider and square root circuit designs using adaptive approximation. IEEE Transactions on Computers, 68(11):1635– 1646, November 2019. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2010:SSS

[JLMH10] J.-H. R. Jiang, Chih-Chun Lee, A. Mishchenko, and Chung-Yang Huang. To SAT or not to SAT: Scalable exploration of functional dependency. IEEE Transactions on Computers, 59(4):457–467, April 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5374381.

Jeannerod:2011:MEP

[JLMP11] Claude-Pierre Jeannerod, Nicolas Louvet, Jean-Michel Muller, and Adrien Panhaleux. Midpoints and exact points of some algebraic functions in floating-point arithmetic. IEEE Transactions on Computers, 60 (2):228–241, February 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Joldes:2016:AAE

[JMMP16] M. Joldes, O. Marty, J. M. Muller, and V. Popescu. Arithmetic algorithms for extended precision using floating-point expansions. IEEE Transactions on Computers, 65(4):1197–1210, April 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Johansson:2017:AEA

[Joh17] Fredrik Johansson. Arb: Efficient arbitrary-precision midpoint–radius interval arithmetic. IEEE Transactions on Computers, 66 (8):1281–1292, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7891956/.

Jiang:2013:DSO

[JP13] Weirong Jiang and Viktor K. Prasanna. Data structure optimization for powerefficient IP lookup architectures. IEEE Transactions on Computers, 62(11): 2169–2182, November 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jaberipur:2010:RDF

[JPG10] G. Jaberipur, B. Parhami, and S. Gorgin. Redundantdigit floating-point addition scheme based on a stored rounding value. IEEE Transactions on Computers, 59(5):694–706, May 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5278659.

Jin:2013:OBF

[JPLP13] Hao Jin, Deng Pan, Jason Liu, and Niki Pissinou. OpenFlow-based flow-level bandwidth provisioning for CICQ switches. IEEE Transactions on Computers, 62(9): 1799–1812, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jo:2017:DLB

[JR17] Myung Hyun Jo and Won Woo Ro. Dynamic load balancing of dispatch scheduling for solid state disks. IEEE Transactions on Computers, 66(6):1034–1047, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7776864/.

Jun:2014:EID

[JRC14] Minje Jun, Won W. Ro, and Eui-Young Chung. Exploiting implementation diversity and partial connection of routers in applicationspecific network-on-chip topology synthesis. IEEE Transactions on Computers, 63 (6):1434–1445, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiao:2018:CCL

[JRJ^{+}18] Xun Jiao, Abbas Rahimi, Yu Jiang, Jianguo Wang, Hamed Fatemi, Jose Pineda de Gyvez, and Rajesh K. Gupta. CLIM: A crosslevel workload-aware timing error prediction model for functional units. IEEE Transactions on Computers, 67(6):771–783, June 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8207606/.

Joo:2014:RPE

[JRP^{+}14] Yongsoo Joo, Junhee Ryu, Sangsoo Park, Heonshik Shin, and Kang G. Shin. Rapid prototyping and evaluation of intelligence functions of active storage devices. IEEE Transactions on Computers, 63(9): 2356–2368, September 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2015:SMC

[JRS^{+}15] Wanchun Jiang, Fengyuan Ren, Ran Shu, Yongwei Wu, and Chuang Lin. Sliding mode congestion control for data center Ethernet networks. IEEE Transactions on Computers, 64(9): 2675–2690, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2014:ABC

[JRW^{+}14] Wanchun Jiang, Fengyuan Ren, Yongwei Wu, Chuang Lin, and I. Stojmenovic. Analysis of backward congestion notification with delay for enhanced Ethernet networks. IEEE Transactions on Computers, 63(11): 2674–2684, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jalili:2017:EAS

[JSA17] Majid Jalili and Hamid Sarbazi-Azad. Enduranceaware security enhancement in non-volatile memories using compression and selective encryption. IEEE Transactions on Computers, 66(7):1132–1144, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07792116-
abs.html.

Jigang:2010:AAH

[JSC10] Wu Jigang, T. Srikanthan, and Guang Chen. Algorithmic aspects of hardware/ software partitioning: 1D search algorithms. IEEE Transactions on Computers, 59(4):532–544, April 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5332226. See

comments [HWG^{+}14].

Jang:2017:RAW

[JSC^{+}17] Jaemin Jang, Wongyu Shin, Jungwhan Choi, Jinwoong Suh, Yongkee Kwon, Yongju Kim, and Lee-Sup Kim. Refresh-aware write recovery memory controller. IEEE Transactions on Computers, 66(4):688–701, April 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jang:2019:SIW

[JSC^{+}19] J. Jang, W. Shin, J. Choi, Y. Kim, and L. Kim. Sparseinsertion write cache to mitigate write disturbance errors in phase change memory. IEEE Transactions on Computers, 68(5):752– 764, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jayalath:2014:CAR

[JSE14] Chamikara Jayalath, Julian Stephen, and Patrick Eugster. From the cloud to the atmosphere: Running MapReduce across data centers. IEEE Transactions on Computers, 63(1):74–87, January 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jeong:2017:DEV

[JSH^{+}17] Jaeyong Jeong, Youngsun Song, Sangwook Shane Hahn, Sungjin Lee, and Jihong Kim. Dynamic erase voltage and time scaling for extending lifetime of NAND flash-based SSDs. IEEE Transactions on Computers, 66(4):616–630, April 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jangjaimon:2015:ECR

[JT15] I. Jangjaimon and Nian-Feng Tzeng. Effective cost reduction for elastic clouds under spot instance pricing through adaptive checkpointing. IEEE Transactions on Computers, 64 (2):396–409, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jung:2016:EDC

[Jun16] M. Jung. Exploring design challenges in getting solid state drives closer to CPU. IEEE Transactions on Computers, 65(4):1103– 1115, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jung:2016:NTB

[JW16] J. Jung and Y. Won. nvramdisk: A transactional block device driver for nonvolatile RAM. IEEE Transactions on Computers, 65 (2):589–600, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jun:2012:PCA

[JWC12] Minje Jun, Deumji Woo, and Eui-Young Chung. Partial connection-aware topology synthesis for on-chip cascaded crossbar network. IEEE Transactions on Computers, 61(1):73–86, January 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2015:RTD

[JWH^{+}15] Guiyuan Jiang, Jigang Wu, Yajun Ha, Yi Wang, and Jizhou Sun. Reconfiguring three-dimensional processor arrays for fault-tolerance: Hardness and heuristic algorithms. IEEE Transactions on Computers, 64(10): 2926–2939, October 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2016:TEO

[JWL^{+}16] W. Jiang, J. Wu, F. Li, G. Wang, and H. Zheng. Trust evaluation in online social networks using generalized network flow. IEEE Transactions on Computers, 65(3):952–963, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2016:FOT

[JWWZ16] W. Jiang, J. Wu, G. Wang, and H. Zheng. Forming opinions via trusted friends: Time-evolving rating prediction using fluid dynamics. IEEE Transactions on Computers, 65(4):1211– 1224, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jing:2017:RAH

[JYL^{+}17] W. Jing, K. Yang, Y. Lin, B. Lee, S. Yoon, Y. Ye, Y. Du, and B. Chen. Retention-Aware Hybrid Main Memory (RAHMM): Big DRAM and little SCM. IEEE Transactions on Computers, 66(5):912–918, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2010:INF

[JZLD10] Song Jiang, Xuechen Zhang, Shuang Liang, and K. Davis. Improving networked file system performance using a locality-aware cooperative cache protocol. IEEE Transactions on Computers, 59 (11):1508–1519, November 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5416680.

Kuan:2019:HEE

[KA19] K. Kuan and T. Adegbija. HALLS: An energy-efficient highly adaptable last level STT-RAM cache for multicore systems. IEEE Transactions on Computers, 68(11): 1623–1634, November 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2015:DAM

[KAH^{+}15] Dae Hyun Kim, K. Athikulwongse, M. B. Healy, M. M. Hossain, Moongon Jung, I. Khorosh, G. Kumar, Young-Joon Lee, D. L. Lewis, Tzu-Wei Lin, Chang Liu, S. Panth, M. Pathak, Minzhen Ren, Guanhao Shen, Taigon Song, Dong Hyuk Woo, Xin Zhao, Joungho Kim, Ho Choi, G. H. Loh, H.-H. S. Lee, and Sung Kyu Lim. Design and analysis of 3D-MAPS (3D Massively Parallel Processor with Stacked Memory). IEEE Transactions on Computers, 64 (1):112–125, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kamali:2018:DHT

[KAH18a] Hadi Mardani Kamali, Kimia Zamiri Azar, and Shaahin Hessabi. DuCNoC: A highthroughput FPGA-based NoC simulator using dualclock lightweight router micro-architecture. IEEE Transactions on Computers, 67(2):208–221, February 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8000664/.

Khdr:2018:AAB

[KAH18b] Heba Khdr, Hussam Amrouch, and Jörg Henkel. Aging-aware boosting. IEEE Transactions on Computers, 67(9):1217–1230, September 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8319494/.

Khdr:2019:DGS

[KAH19] Heba Khdr, Hussam Amrouch, and Jörg Henkel. Dynamic guardband selection: Thermal-aware optimization for unreliable multi-core systems. IEEE Transactions on Computers, 68(1):53–66, January 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8395346/.

Koziel:2018:HPS

[KAK18] Brian Koziel, Reza Azarderakhsh, and Mehran Mozaffari Kermani. A highperformance and scalable hardware architecture for isogeny-based cryptography. IEEE Transactions on Computers, 67(11):1594–1609, November 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8315051/.

Kosmidis:2014:ECD

[KAQC14] L. Kosmidis, J. Abella, E. Quinones, and F. J. Cazorla. Efficient cache designs for probabilistically analysable real-time systems. IEEE Transactions on Computers, 63(12): 2998–3011, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kong:2013:AAS

[KASZ13] Jingfei Kong, Onur Aciicmez, Jean-Pierre Seifert, and Huiyang Zhou. Architecting against software cache-based side-channel attacks. IEEE Transactions on Computers, 62(7):1276– 1288, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2010:AEH

[KBH^{+}10] Dong Kim, Kwanhu Bang, Seung-Hwan Ha, Sungroh Yoon, and Eui-Young Chung. Architecture exploration of high-performance PCs with a solid-state disk. IEEE Transactions on Computers, 59(7):878–890, July 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5438986.

Kim:2019:EMA

[KBO^{+}19] M. S. Kim, A. A. D. Barrio, L. T. Oliveira, R. Hermida, and N. Bagherzadeh. Efficient Mitchell’s approximate log multipliers for convolutional neural networks. IEEE Transactions on Computers, 68(5):660–675, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kouretas:2013:LPL

[KBP13] Ioannis Kouretas, Charalambos Basetas, and Vassilis Paliouras. Low-power logarithmic number system addition/subtraction and their impact on digital filters. IEEE Transactions on Computers, 62(11):2196–2209, November 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kshemkalyani:2013:PDA

[KC13] Ajay D. Kshemkalyani and Jiannong Cao. Predicate detection in asynchronous pervasive environments. IEEE Transactions on Computers, 62(9):1823–1836, September 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kao:2014:DBP

[KC14] Yu-Hsiang Kao and H. Jonathan Chao. Design of a bufferless photonic clos networkon-chip architecture. IEEE Transactions on Computers, 63(3):764–776, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2018:SWT

[KCE^{+}18] Kang-Wook Kim, Youngeun Cho, Jeongyoon Eo, Chang-Gun Lee, and Junghee Han. System-wide time versus density tradeoff in real-time multicore fluid scheduling. IEEE Transactions on Computers, 67 (7):1007–1022, July 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8259245/.

Koo:2016:TRE

[KCK16] N. Koo, G. H. Cho, and S. Kwon. On r-th root extraction algorithm in F_{q} for q ≡ lr^{s}(modr^{s+1}) with 0 < l < r and small s. IEEE Transactions on Computers, 65(1):322–325, January 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2019:EDA

[KCKL19] M. Kim, J. Choi, H. Kim, and H. Lee. An effective DRAM address remapping for mitigating rowhammer errors. IEEE Transactions on Computers, 68(10): 1428–1441, October 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2016:WNV

[KCL^{+}16] R. G. Kim, W. Choi, G. Liu, E. Mohandesi, P. P. Pande, D. Marculescu, and R. Marculescu. Wireless NoC for VFI-enabled multicore chip design: Performance evaluation and design tradeoffs. IEEE Transactions on Computers, 65(4):1323– 1336, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2019:STC

[KCL19] M. Kim, I. Chang, and H. Lee. Segmented tag cache: A novel cache organization for reducing dynamic read energy. IEEE Transactions on Computers, 68(10): 1546–1552, October 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2014:CEC

[KCRG14] Seung Hun Kim, Dongmin Choi, Won Woo Ro, and Jean-Luc Gaudiot. Complexity-effective contention management with dynamic backoff for transactional memory systems. IEEE Transactions on Computers, 63(7):1696–1708, July 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2015:NVF

[KCRG15] Keunsoo Kim, B. Y. Cho, Won Woo Ro, and J.-L. Gaudiot. Network variation and fault tolerant performance acceleration in mobile devices with simultaneous remote execution. IEEE Transactions on Computers, 64(10):2862–2874, October 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kinsy:2013:OHA

[KCS^{+}13] Michel A. Kinsy, Myong Hyon Cho, Keun Sup Shim, Mieszko Lis, G. Edward Suh, and Srinivas Devadas. Optimal and heuristic application-aware oblivious routing. IEEE Transactions on Computers, 62 (1):59–73, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kuo:2014:MET

[KCS14] Fang-Chen Kuo, Yeim-Kuan Chang, and Cheng-Chien Su. A memory-efficient TCAM coprocessor for IPv4/IPv6 routing table update. IEEE Transactions on Computers, 63(9):2110–2121, September 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kang:2017:PDS

[KCW^{+}17] Wang Kang, Liang Chang, Zhaohao Wang, Weifeng Lv, Guanyu Sun, and Weisheng Zhao. Pseudo-differential sensing framework for STT-MRAM: A cross-layer perspective. IEEE Transactions on Computers, 66 (3):531–544, March 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2018:MOL

[KCY18] Sae-Eun Kim, Jaeyong Chung, and Joon-Sung Yang. Mitigating observability loss of toggle-BasedX-masking via scan chain partitioning. IEEE Transactions on Computers, 67(8): 1184–1192, August 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8280565/.

Kozhaya:2019:RBB

[KDEV19] David Kozhaya, Jérémie Decouchant, and Paulo Esteves-Verissimo. RT-ByzCast: Byzantine-resilient real-time reliable broadcast. IEEE Transactions on Computers, 68(3):440– 454, March 2019. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8470958/.

Kara:2019:ALS

[KE19] Orhun Kara and Muhammed F. Esgin. On analysis of lightweight stream ciphers with keyed update. IEEE Transactions on Computers, 68(1):99–110, January 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8400392/.

Kizilkale:2016:MDM

[KEK16] Can Kızılkale, Ömer Eğecioğlu, and Çetin Kaya Koç. A matrix decomposition method for optimal normal basis multiplication. IEEE Transactions on Computers, 65 (11):3239–3250, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Khemka:2015:UFR

[KFB^{+}15] B. Khemka, R. Friese, L. D. Briceno, A. A. Maciejewski, G. A. Koenig, G. Okonski, M. M. Hilton, R. Rambharos, S. Poole, and C. Groer. Utility functions and resource management in an oversubscribed heterogeneous computing environment. IEEE Transactions on Computers, 64 (8):2394–2407, August 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Khan:2014:DRS

[KGC14] Faisal Khan, Soheil Ghiasi, and Chen-Nee Chuah. A dynamically reconfigurable system for closed-loop measurements of network traffic. IEEE Transactions on Computers, 63(2):263–275, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kadric:2016:APF

[KGD16] Edin Kadric, Paul Gurniak, and André DeHon. Accurate parallel floating-point accumulation. IEEE Transactions on Computers, 65(11): 3224–3238, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2014:SLS

[KGGJ14] Hyungjun Kim, Boris Grot, Paul V. Gratz, and Daniel A. Jimenez. Spatial locality speculation to reduce energy in chip-multiprocessor networks-on-chip. IEEE Transactions on Computers, 63(3):543–556, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kashif:2015:SSL

[KGP15] H. Kashif, S. Gholamian, and H. Patel. SLA: A stage-level latency analysis for real-time communication in a pipelined resource model. IEEE Transactions on Computers, 64 (4):1177–1190, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Khedr:2016:SSH

[KGV16] Alhassan Khedr, Glenn Gulak, and Vinod Vaikuntanathan. SHIELD: Scalable homomorphic implementation of encrypted dataclassifiers. IEEE Transactions on Computers, 65(9): 2848–2858, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kwon:2010:SEB

[KH10] Taekyoung Kwon and Jin Hong. Secure and efficient broadcast authentication in wireless sensor networks. IEEE Transactions on Computers, 59 (8):1120–1133, August 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5313804.

Koohi:2014:AOW

[KH14] Somayyeh Koohi and Shaahin Hessabi. All-optical wavelengthrouted architecture for a power-efficient network on chip. IEEE Transactions on Computers, 63(3):777– 792, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2018:EDS

[KH18] Changdae Kim and Jaehyuk Huh. Exploring the design space of fair scheduling supports for asymmetric multicore systems. IEEE Transactions on Computers, 67 (8):1136–1152, August 2018.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8265024/.

Kim:2016:EPE

[KHPP16] Intae Kim, Seong Oun Hwang, Jong Hwan Park, and Chanil Park. An efficient predicate encryption with constant pairing computations and minimum costs. IEEE Transactions on Computers, 65(10): 2947–2958, October 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

kanduri:2018:ATA

[khR^{+}18] Anil kanduri, Mohammad-Hashem haghbayan, Amir M. Rahmani, Muhammad Shafique, Axel Jantsch, and Pasi Liljeberg. adBoost: Thermal aware performance boosting through dark silicon patterning. IEEE Transactions on Computers, 67 (8):1062–1077, August 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8292829/.

Kumm:2017:OCM

[KHZ17] Martin Kumm, Martin Hardieck, and Peter Zipf. Optimization of constant matrix multiplication with low power and high throughput. IEEE Transactions on Computers, 66(12):2072– 2080, December 2017. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7919250/.

Kaseridis:2014:CFA

[KIJ14] Dimitris Kaseridis, Muhammad Faisal Iqbal, and Lizy Kurian John. Cache friendliness-aware management of shared last-level caches for high performance multi-core systems. IEEE Transactions on Computers, 63(4):874–887, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2015:CEH

[Kim15] Y. Kim. Comments on “An Efficient Homomorphic MAC with Small Key Size for Authentication in Network Coding. IEEE Transactions on Computers, 64(12): 3619–3620, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [CJ13].

Kim:2011:XEC

[KJL11] Hwanju Kim, Heeseung Jo, and Joonwon Lee. XHive: Efficient cooperative caching for virtual machines. IEEE Transactions on Computers, 60(1):106–119, January 2011. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic).

Khan:2010:TRR

[KK10] O. Khan and S. Kundu. Thread relocation: a runtime architecture for tolerating hard errors in chip multiprocessors. IEEE Transactions on Computers, 59(5):651–665, May 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5010431.

Kumm:2018:ACT

[KK18] Martin Kumm and Johannes Kappauf. Advanced compressor tree synthesis for FPGAs. IEEE Transactions on Computers, 67(8): 1078–1091, August 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8263391/.

Kim:2015:SCF

[KkC15a] Jae Min Kim, Young Geun kim, and Sung Woo Chung. Stabilizing CPU frequency and voltage for temperatureaware DVFS in mobile devices. IEEE Transactions on Computers, 64(1):286– 292, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kong:2015:EEL

[KKC15b] Joonho Kong, F. Koushanfar, and Sung Woo Chung. An energy-efficient last-level cache architecture for process variation-tolerant 3D microprocessors. IEEE Transactions on Computers, 64(9):2460–2475, September 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2017:EEE

[KKC17] Young Geun Kim, Minyong Kim, and Sung Woo Chung. Enhancing energy efficiency of multimedia applications in heterogeneous mobile multicore processors. IEEE Transactions on Computers, 66 (11):1878–1889, November 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7937919/.

Kalyanaraman:2014:SRC

[KKH^{+}14] Shivkumar Kalyanaraman, Isaac Keslassy, Inbal Horev, Pu Li, and Ori Rottenstreich. The switch reordering contagion: Preventing a few late packets from ruining the whole party. IEEE Transactions on Computers, 63(5):1262–1276, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Koh:2019:DCM

[KKJH19] Kwangwon Koh, Kangho Kim, Seunghyub Jeon, and Jaehyuk Huh. Disaggregated cloud memory with elastic block management. IEEE Transactions on Computers, 68(1):39–52, January 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8400405/.

Kekely:2016:SDM

[KKP^{+}16] L. Kekely, J. Kucera, V. Pus, J. Korenek, and A. V. Vasilakos. Software defined monitoring of application protocols. IEEE Transactions on Computers, 65 (2):615–626, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kong:2014:DGD

[KKS14] Inwook Kong, Seong-Wan Kim, and E. E. Swartzlander. Design of Goldschmidt dividers with quantum-dot cellular automata. IEEE Transactions on Computers, 63(10):2620–2625, October 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kumarage:2015:GEA

[KKT15] H. Kumarage, I. Khalil, and Z. Tari. Granular evaluation of anomalies in wireless sensor networks using dynamic data partitioning with an entropy criteria. IEEE Transactions on Computers, 64(9): 2573–2585, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2016:DAH

[KKY^{+}16] H. Kim, G. Kim, H. Yeo, J. Kim, and S. Maeng. Design and analysis of hybrid flow control for hierarchical ring network-on-chip. IEEE Transactions on Computers, 65(2):480– 494, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kumar:2013:NHM

[KL13] T. Nandha Kumar and Fabrizio Lombardi. A novel heuristic method for application-dependent testing of a SRAM-based FPGA interconnect. IEEE Transactions on Computers, 62 (1):163–172, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kuo:2016:RSF

[KL16] Wen-Hsing Kuo and Yung-Hsuan Lin. Resource-saving file management scheme for online video provisioning on content delivery networks. IEEE Transactions on Computers, 65(6):1910– 1920, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2016:CLR

[KLC^{+}16] J. Kim, E. Lee, J. Choi, D. Lee, and S. H. Noh. Chip-level RAID with flexible stripe size and parity placement for enhanced SSD reliability. IEEE Transactions on Computers, 65 (4):1116–1130, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2018:GAP

[KLC18] Mincheol Kim, Ling Liu, and Wonik Choi. A GPUaware parallel index for processing high-dimensional big data. IEEE Transactions on Computers, 67(10): 1388–1402, October 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8332539/.

Kim:2014:CLE

[KLJ^{+}14] Seung Hun Kim, Sang Hyong Lee, Minje Jun, Byunghoon Lee, Won Woo Ro, Eui-Young Chung, and Jean-Luc Gaudiot. C-Lock: Energy efficient synchronization for embedded multicore systems. IEEE Transactions on Computers, 63 (8):1962–1974, August 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2014:LON

[KLK^{+}14] Gwangsun Kim, Michael Mihn-Jong Lee, John Kim, Jae W. Lee, Dennis Abts, and Michael Marty. Lowoverhead network-on-chip support for location-oblivious task placement. IEEE Transactions on Computers, 63(6):1487–1500, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2017:TIB

[KLK17] Jesung Kim, Jongmin Lee, and Soontae Kim. TLB index-based tagging for reducing data cache and TLB energy consumption. IEEE Transactions on Computers, 66(7):1200–1211, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07805198-
abs.html.

Kang:2018:SAS

[KLK18] Mincheol Kang, Wonyoung Lee, and Soontae Kim. Subpage-aware solid state drive for improving lifetime and performance. IEEE Transactions on Computers, 67(10):1492–1505, October 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8338090/.

Kim:2013:SED

[KLKL13] Minjang Kim, Nagesh B. Lakshminarayana, Hyesoon Kim, and Chi-Keung Luk. SD3: An efficient dynamic data-dependence profiling mechanism. IEEE Transactions on Computers, 62(12): 2516–2530, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2011:RPN

[KLLK11] Jong-Chan Kim, Duhee Lee, Chang-Gun Lee, and Kanghee Kim. RT-PLRU: a new paging scheme for real-time execution of program codes on NAND flash memory for portable media players. IEEE Transactions on Computers, 60 (8):1126–1141, August 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kornerup:2012:CCR

[KLLM12] Peter Kornerup, Vincent Lefèvre, Nicolas Louvet, and Jean-Michel Muller. On the computation of correctly rounded sums. IEEE Transactions on Computers, 61 (3):289–298, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kuo:2016:CDA

[KLT16] Tung-Wei Kuo, Kate Ching-Ju Lin, and Ming-Jer Tsai. On the construction of data aggregation tree with minimum energy cost in wireless sensor networks: NPcompleteness and approximation algorithms. IEEE Transactions on Computers, 65(10):3109–3121, October 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kargahi:2011:POB

[KM11] Mehdi Kargahi and Ali Movaghar. Performance optimization based on analytical modeling in a real-time system with constrained time/utility functions. IEEE Transactions on Computers, 60(8):1169–1181, August 2011. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kedar:2017:SSP

[KMC17] Gil Kedar, Avi Mendelson, and Israel Cidon. SPACE: Semi-PArtitioned CachE for energy efficient, hard realtime systems. IEEE Transactions on Computers, 66 (4):717–730, April 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Karimi:2011:WCC

[KMJ^{+}11] Naghmeh Karimi, Michail Maniatakos, Abhijit Jas, Chandra Tirumurti, and Yiorgos Makris. Workloadcognizant concurrent error detection in the scheduler of a modern microprocessor. IEEE Transactions on Computers, 60(9): 1274–1287, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5669287.

Khalgui:2011:RME

[KMLH11] Mohamed Khalgui, Olfa Mosbahi, ZhiWu Li, and Hans-Michael Hanisch. Reconfigurable multiagent embedded control systems: From modeling to implementation. IEEE Transactions on Computers, 60(4):538– 551, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kliuchnikov:2016:PAS

[KMM16] V. Kliuchnikov, D. Maslov, and M. Mosca. Practical approximation of singlequbit unitaries by singlequbit quantum Clifford and T circuits. IEEE Transactions on Computers, 65 (1):161–172, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kornerup:2011:PAO

[KMP11] Peter Kornerup, Jean-Michel Muller, and Adrien Panhaleux. Performing arithmetic operations on roundto-nearest representations. IEEE Transactions on Computers, 60(2):282–291, February 2011. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kinsman:2011:CVM

[KN11a] Adam B. Kinsman and Nicola Nicolici. Computational vector-magnitude-based range determination for scientific abstract data types. IEEE Transactions on Computers, 60(11): 1652–1663, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5669288.

Kinsman:2011:TOT

[KN11b] Adam B. Kinsman and Nicola Nicolici. Tradeoffs in test data compression and deterministic Xmasking of responses. IEEE Transactions on Computers, 60(4):498–507, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ko:2012:MTC

[KN12] Ho Fai Ko and Nicola Nicolici. Mapping trigger conditions onto trigger units during post-silicon validation and debugging. IEEE Transactions on Computers, 61(11):1563–1575, November 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kinsman:2013:NBF

[KN13] Phillip J. Kinsman and Nicola Nicolici. NoCbased FPGA acceleration for Monte Carlo simulations with applications to SPECT imaging. IEEE Transactions on Computers, 62(3):524–535, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2014:ENL

[KO14] Jihye Kim and Hyunok Oh. An efficient non-linear cost compression algorithm for multi level cell memory. IEEE Transactions on Computers, 63(4):820– 832, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kornerup:2015:RHR

[Kor15]  P. Kornerup.  Review-

ing high-radix signed-digit adders. IEEE Transactions on Computers, 64(5):1502– 1505, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2013:PMM

[KP13] Sook-Yeon Kim and Jung-Heum Park. Paired manyto-many disjoint path covers in recursive circulants (G(2^{m}, 4)). IEEE Transactions on Computers, 62(12): 2468–2475, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2015:CBR

[KP15] Chulmin Kim and Kyu Ho Park. Credit-based runtime placement of virtual machines on a single NUMA system for QoS of data access performance. IEEE Transactions on Computers, 64(6):1633–1646, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kaitoua:2017:FSG

[KPBC17] Abdulrahman Kaitoua, Pietro Pinoli, Michele Bertoni, and Stefano Ceri. Framework for supporting genomic operations. IEEE Transactions on Computers, 66 (3):443–457, March 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Khdr:2017:PDA

[KPS^{+}17] Heba Khdr, Santiago Pagani, Éricles Sousa, Vahid Lari, Anuj Pathania, Frank Hannig, Muhammad Shafique, Jürgen Teich, and Jörg Henkel. Power densityaware resource management for heterogeneous tiled multicores. IEEE Transactions on Computers, 66(3):488– 501, March 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2018:SBL

[KRP18] Jaehyung Kim, Hongchan Roh, and Sanghyun Park. Selective I/O bypass and load balancing method for write-through SSD caching in big data analytics. IEEE Transactions on Computers, 67(4):589–595, April 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8100898/.

Karmakar:2018:CTD

[KRR^{+}18] Angshuman Karmakar, Sujoy Sinha Roy, Oscar Reparaz, Frederik Vercauteren, and Ingrid Verbauwhede. Constanttime discrete Gaussian sampling. IEEE Transactions on Computers, 67(11): 1561–1571, November 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8314133/.

Kang:2010:ISG

[KS10a] Kyungtae Kang and Lui Sha.

An interleaving structure for guaranteed QoS in real-time broadcasting systems. IEEE Transactions on Computers, 59(5):666–678, May 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5276796.

Kong:2010:RMR

[KS10b] Inwook Kong and E. E.

Swartzlander. A rounding method to reduce the required multiplier precision for Goldschmidt division. IEEE Transactions on Computers, 59(12): 1703–1708, December 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5453345.

Kim:2012:SMR

[KS12] Eric P. Kim and Naresh R.

Shanbhag. Soft N-modular redundancy. IEEE Transactions on Computers, 61 (3):323–336, March 2012.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kharbutli:2014:LLA

[KS14] Mazen Kharbutli and Rami Sheikh. LACS: A localityaware cost-sensitive cache replacement algorithm. IEEE Transactions on Computers, 63(8):1975–1987, August 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kong:2014:LPV

[KSC^{+}14] Joonho Kong, Taeweon Suh, Sung Woo Chung, Junhee Lee, and Hyung Beom Jang. Leveraging process variation for performance and energy: In the perspective of overclocking. IEEE Transactions on Computers, 63 (5):1316–1322, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kayi:2015:ACC

[KSEG15] A. Kayi, O. Serres, and T. El-Ghazawi. Adaptive cache coherence mechanisms with producer-consumer sharing optimization for chip multiprocessors. IEEE Transactions on Computers, 64(2):316–328, February 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2012:PAM

[KSJ^{+}12] Jaehong Kim, Sangwon Seo, Dawoon Jung, Jin-Soo Kim, and Jaehyuk Huh. Parameter-aware I/O management for solid state disks (SSDs). IEEE Transactions on Computers, 61(5):636– 649, May 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kayaalp:2015:SBP

[KSN^{+}15] M. Kayaalp, T. Schmitt, J. Nomani, D. Ponomarev, and N. Abu Ghazaleh. Signature-based protection from code reuse attacks. IEEE Transactions on Computers, 64(2):533–546, February 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kang:2012:DIE

[KSS12] Woochul Kang, Sang Hyuk Son, and John A. Stankovic. Design, implementation, and evaluation of a QoSaware real-time embedded database. IEEE Transactions on Computers, 61 (1):45–59, January 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kyrkou:2012:PHA

[KT12] Christos Kyrkou and Theocharis Theocharides. A parallel hardware architecture for real-time object detection with support vector machines. IEEE Transactions on Computers, 61(6):831– 842, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kito:2019:CED

[KT19] N. Kito and N. Takagi. Concurrent error detectable carry select adder with easy testability. IEEE Transactions on Computers, 68 (7):1105–1110, July 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kavousianos:2014:TDT

[KTA^{+}14] Xrysovalantis Kavousianos, Yiorgos Tsiatouhas, Angela Arapoyanni, Andreas Floros, and Stefanos Valadimas. The time dilation technique for timing error tolerance. IEEE Transactions on Computers, 63(5):1277–1286, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kanoun:2016:BDS

[KTAvdS16] Karim Kanoun, Cem Tekin, David Atienza, and Mihaela van der Schaar. Bigdata streaming applications scheduling based on staged multi-armed bandits. IEEE Transactions on Computers, 65(12):3591–3605, December 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kurka:2012:FAA

[Kůr12] Petr Kůrka. Fast arithmetical algorithms in Möbius number systems. IEEE Transactions on Computers, 61(8):1097–1109, August 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Knezevic:2010:FIM

[KVV10] M. Knezevic, F. Vercauteren, and I. Verbauwhede. Faster interleaved modular multiplication based on Barrett and Montgomery reduction methods. IEEE Transactions on Computers, 59(12): 1715–1721, December 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5453352.

Karpovsky:2014:DSS

[KW14] M. Karpovsky and Zhen Wang. Design of strongly secure communication and computation channels by nonlinear error detecting codes. IEEE Transactions on Computers, 63(11): 2716–2728, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2016:ECI

[KWC^{+}16] Dongwook Kim, Youjip Won, Jaehyuk Cha, Sungroh Yoon, Jongmoo Choi, and Sooyong Kang. Exploiting compression-induced internal fragmentation for poweroff recovery in SSD. IEEE Transactions on Computers, 65(6):1720–1733, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kang:2015:OCS

[KwPK^{+}15] Shin-Haeng Kang, Hae woo Park, Sungchan Kim, Hyunok Oh, and Soonhoi Ha. Optimal checkpoint selection with dualmodular redundancy hardening. IEEE Transactions on Computers, 64(7):2036– 2048, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kuang:2019:STD

[KYZC19] Liwei Kuang, Laurence T. Yang, Qing Zhu, and Jinjun Chen. Secure tensor decomposition for big data using transparent computing paradigm. IEEE Transactions on Computers, 68(4):585–596, April 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8310010/.

Lamberti:2011:RCT

[LAAM11] Fabrizio Lamberti, Nikos Andrikos, Elisardo Antelo, and Paolo Montuschi. Reducing the computation time in (short bit-width) two’s complement multipliers. IEEE Transactions on Computers, 60(2):148–156, February 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2013:PAE

[LB13] Bin Liu and Bevan M. Baas.

Parallel AES encryption engines for many-core processor arrays. IEEE Transactions on Computers, 62 (3):536–547, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2015:AGP

[LB15a] D. T. Lee and J. Bruck.

Algorithms for generating probabilities with multivalued stochastic relay circuits. IEEE Transactions on Computers, 64(12):3376–3388, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lubicz:2015:TOB

[LB15b] D. Lubicz and N. Bochard.

Towards an oscillator based TRNG with a certified entropy rate. IEEE Transactions on Computers, 64 (4):1191–1200, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2014:CDW

[LBN14] Soyoon Lee, Hyokyung Bahn, and Sam H. Noh. CLOCK-DWF: A writehistory-aware page replacement algorithm for hybrid PCM and DRAM memory architectures. IEEE Transactions on Computers, 63(9): 2187–2200, September 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lukic:2015:RCE

[LBS15] M. Lukic, A. Barnawi, and I. Stojmenovic. Robot coordination for energybalanced matching and sequence dispatch of robots to events. IEEE Transactions on Computers, 64 (5):1416–1428, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Laga:2017:MMR

[LBSK17] Arezki Laga, Jalil Boukhobza, Frank Singhoff, and Michel Koskas. MONTRES: Merge ON-the-Run External Sorting algorithm for large data volumes on SSD based storage systems. IEEE Transactions on Computers, 66(10): 1689–1702, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7932112/.

Liao:2011:IBG

[LBWH11] Xue-Liang Liao, Shi Bai, Yu-Ping Wang, and Shi-Min Hu. ISRA-based grouping: a disk reorganization approach for disk energy conservation and disk performance enhancement. IEEE Transactions on Computers, 60 (2):292–304, February 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2016:TRT

[LC16a] Yu Li and Albert Mo Kim Cheng. Transparent realtime task scheduling on temporal resource partitions. IEEE Transactions on Computers, 65(5):1646– 1655, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2016:RQR

[LC16b] Huai Liu and Tsong Yueh Chen. Randomized quasirandom testing. IEEE Transactions on Computers, 65(6):1896–1909, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lhotak:2010:OER

[LCA10] O. Lhotak, S. Curial, and J. N. Amaral. An optimal encoding to represent a single set in an ROBDD. IEEE Transactions on Computers, 59(4):574–575, April 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5184823.

Lee:2010:PFT

[LCC10] Hyunjin Lee, Sangyeun Cho, and B. R. Childers. PER-FECTORY: a fault-tolerant directory memory architecture. IEEE Transactions on Computers, 59(5):638– 650, May 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5255232.

Liu:2013:IAG

[LCCJ13] Yining Liu, Chi Cheng, Jianyu Cao, and Tao Jiang. An improved authenticated group key transfer protocol based on secret sharing. IEEE Transactions on Computers, 62(11):2335–2336, November 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2013:CEL

[LCH13] Jui-Chieh Lin, Sao-Jie Chen, and Yu Hen Hu. Cycleefficient LFSR implementation on word-based microarchitecture. IEEE Transactions on Computers, 62 (4):832–838, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2015:SDD

[LCH^{+}15] J. Li, X. Chen, X. Huang, S. Tang, Y. Xiang, M. M. Hassan, and A. Alelaiwi. Secure distributed deduplication systems with improved reliability. IEEE Transactions on Computers, 64(12): 3569–3579, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liang:2014:OSW

[LCHC14] Qilian Liang, Xiuzhen Cheng, Scott C. H. Huang, and Dechang Chen. Opportunistic sensing in wireless sensor networks: Theory and application. IEEE Transactions on Computers, 63 (8):2002–2010, August 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2011:DFS

[LCHX11] Alex X. Liu, Fei Chen, Jee-Hyun Hwang, and Tao Xie. Designing fast and scalable XACML policy evaluation engines. IEEE Transactions on Computers, 60(12): 1802–1817, December 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Leu:2015:HSS

[LCL15] Jenq-Shiou Leu, Jheng-Huei Chen, and Kuen-Han Li. Hybrid search scheme for social networks supported by dynamic weighted distributed label clustering. IEEE Transactions on Computers, 64(9):2586–2594, September 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2017:HDL

[LCL17] Shih-Hsiang Lin, Pei-Yin Chen, and Yu-Ning Lin. Hardware design of lowpower high-throughput sorting unit. IEEE Transactions on Computers, 66(8): 1383–1395, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7862779/.

Luo:2019:TRE

[LCLC19] J. Luo, H. Cheng, I. Lin, and D. Chang. TAP: Reducing the energy of asymmetric hybrid last-level cache via thrashing aware placement and migration. IEEE Transactions on Computers, 68(12):1704–1719, December 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2015:CEH

[LCLL15] Chen Li, Le Chen, Rongxing Lu, and Hui Li. Comment on “An Efficient Homomorphic MAC with Small Key Size for Authentication in Network Coding”. IEEE Transactions on Computers, 64(3):882–883, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [CJ13].

Liu:2011:MEL

[LCT11] Bing-Hong Liu, Min-Lun Chen, and Ming-Jer Tsai. Message-efficient location prediction for mobile objects in wireless sensor networks using a maximum likelihood technique. IEEE Transactions on Computers, 60(6):865–878, June 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2015:CEL

[LCW^{+}15] Chuan Luo, Shaowei Cai, Wei Wu, Zhong Jie, and Kaile Su. CCLS: An efficient local search algorithm for weighted maximum satisfiability. IEEE Transactions on Computers, 64 (7):1830–1843, July 2015.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2016:DAI

[LCW^{+}16] W. Liu, L. Chen, C. Wang, M. O’Neill, and F. Lombardi. Design and analysis of inexact floating-point adders. IEEE Transactions on Computers, 65(1):308– 314, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liao:2010:SPK

[LCwW10] Xiaofeng Liao, Fei Chen, and Kwok wo Wong. On the security of public-key algorithms based on Chebyshev polynomials over the finite field Z_{N}. IEEE Transactions on Computers, 59(10): 1392–1401, October 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5487511.

Li:2016:TCM

[LCX^{+}16] J. Li, J. J. Chen, M. Xiong, G. Li, and W. Wei. Temporal consistency maintenance upon partitioned multiprocessor platforms. IEEE Transactions on Computers, 65(5):1632–1645, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2013:FMT

[LCY^{+}13] Chen-Wei Lin, Hung-Hsin Chen, Hao-Yu Yang, Chin-Yuan Huang, Mango C.-T. Chao, and Rei-Fu Huang. Fault models and test methods for subthreshold SRAMs. IEEE Transactions on Computers, 62 (3):468–481, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2016:IRP

[LCY^{+}16] R. S. Liu, M. Y. Chuang, C. L. Yang, C. H. Li, K. C. Ho, and H. P. Li. Improving read performance of NAND flash SSDs by exploiting error locality. IEEE Transactions on Computers, 65(4):1090–1102, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2019:DAA

[LCY^{+}19] W. Liu, T. Cao, P. Yin, Y. Zhu, C. Wang, E. E. Swartzlander, and F. Lombardi. Design and analysis of approximate redundant binary multipliers. IEEE Transactions on Computers, 68(6):804–819, June 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2019:MDF

[LcZLG19] G. Li, c. Zhou, J. Li, and B. Guo. Maintaining data freshness in distributed cyber-physical systems. IEEE Transactions on Computers, 68(7):1077– 1090, July 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2017:RWN

[LDB^{+}17] Xian Li, Karthi Duraisamy, Joe Baylon, Turbo Majumder, Guopeng Wei, Paul Bogdan, Deukhyoun Heo, and Partha Pratim Pande. A reconfigurable wireless NoC for large scale microbiome community analysis. IEEE Transactions on Computers, 66(10):1653–1666, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7932130/.

Li:2017:DPA

[LDL^{+}17] Guohui Li, Chenggang Deng, Jianjun Li, Quan Zhou, and Wei Wei. Deadline and period assignment for update transactions in co-scheduling environment. IEEE Transactions on Computers, 66 (7):1119–1131, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07797455-
abs.html.

Li:2016:PPP

[LDMQ16] Y. Li, W. Dai, Z. Ming, and M. Qiu. Privacy protection for preventing data over-collection in Smart City. IEEE Transactions on Computers, 65(5):1339– 1350, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2010:EMV

[LDP10] Bin Li, Lide Duan, and Lu Peng. Efficient microarchitectural vulnerabilities prediction using boosted regression trees and patient rule inductions. IEEE Transactions on Computers, 59(5):593–607, May 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5255234.

Lee:2012:CPS

[Lee12] Mun-Kyu Lee. Comments on “Provably Sublinear Point Multiplication on Koblitz Curves and Its Hardware Implementation”. IEEE Transactions on Computers, 61(4):591–592, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [DJJ^{+}08].

Lee:2017:ISA

[Lee17]  Jinkyu Lee.  Improved

schedulability analysis using carry-in limitation for nonpreemptive fixed-priority multiprocessor scheduling. IEEE Transactions on Computers, 66(10):1816–1823, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/
document/7927414/.

Lefevre:2017:CRA

[Lef17]  Vincent Lefèvre.  Cor-

rectly rounded arbitraryprecision floating-point summation. IEEE Transactions on Computers, 66(12): 2111–2124, December 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7891894/.

Liu:2016:BEA

[LFH^{+}16] P. Liu, L. Fang, M. C. Huang, Q. Hu, and G. Jiang. Building expressive and area-efficient directories with hybrid representation and adaptive multi-granular tracking. IEEE Transactions on Computers, 65(3):847– 859, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ling:2013:NPS

[LFJ^{+}13] Zhen Ling, Xinwen Fu, Weijia Jia, Wei Yu, Dong Xuan, and Junzhou Luo. Novel packet size-based covert channel attacks against anonymizer. IEEE Transactions on Computers, 62(12): 2411–2426, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2015:VRA

[LGF^{+}15] Juan Luo, Yaling Guo, Shan Fu, Keqin Li, and Wenfeng He. Virtual resource allocation based on link interference in Cayley wireless data centers. IEEE Transactions on Computers, 64(10): 3016–3021, October 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2015:EEC

[LGH15] Peng Li, Song Guo, and Jiankun Hu. Energy-efficient cooperative communications for multimedia applications in multi-channel wireless networks. IEEE Transactions on Computers, 64 (6):1670–1679, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2017:ECC

[LGH^{+}17] Z. Liu, J. Großschädl, Z. Hu, K. Järvinen, H. Wang, and I. Verbauwhede. Elliptic curve cryptography with efficiently computable endomorphisms and its hardware implementations for the Internet of Things. IEEE Transactions on Computers, 66(5):773–785, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Loghi:2010:ALP

[LGMP10] M. Loghi, O. Golubeva, E. Macii, and M. Poncino. Architectural leakage power minimization of scratchpad memories by applicationdriven subbanking. IEEE Transactions on Computers, 59(7):891–904, July 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5416686.

Liu:2018:SAI

[LGS^{+}18] Di Liu, Nan Guan, Jelena Spasic, Gang Chen, Songran Liu, Todor Stefanov, and Wang Yi. Scheduling analysis of imprecise mixedcriticality real-time tasks. IEEE Transactions on Computers, 67(7):975–991, July 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8247214/.

Lin:2011:CGA

[LH11] Shang-Wei Lin and Pao-Ann Hsiung. Counterexampleguided assume-guarantee synthesis through learning.

IEEE Transactions on Computers, 60(5):734–750, May 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2012:PMC

[LH12a] Chia-Wei Lee and Sun-Yuan Hsieh. Pancyclicity of matching composition networks under the conditional fault model. IEEE Transactions on Computers, 61 (2):278–183, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2012:MCP

[LH12b] Shang-Wei Lin and Pao-Ann Hsiung. Model checking prioritized timed systems. IEEE Transactions on Computers, 61(6):843– 856, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2016:HSC

[LH16] Yuan-Cheng Lee and Chih-Wen Hsueh. Hardware/ software co-design of memory page translation for mobile virtualization. IEEE Transactions on Computers, 65(10):3070–3082, October 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lam:2014:SAD

[LHC^{+}14] Kam-Yiu Lam, Song Han, Deji Chen, Ming Xiong, Krithi Ramamritham, and Aloysius K. Mok. Schedulability analysis of deferrable scheduling algorithms for maintaining real-time data freshness. IEEE Transactions on Computers, 63 (4):979–994, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lien:2013:EDA

[LHCL13] Chih-Yuan Lien, Chien-Chuan Huang, Pei-Yin Chen, and Yi-Fan Lin. An efficient denoising architecture for removal of impulse noise in images. IEEE Transactions on Computers, 62(4):631–643, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2014:DBS

[LHH14a] Chun-Han Lin, Pi-Cheng Hsiu, and Cheng-Kang Hsieh. Dynamic backlight scaling optimization: A cloud-based energy-saving service for mobile streaming applications. IEEE Transactions on Computers, 63 (2):335–348, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2014:SWC

[LHH14b] Ting-Yu Lin, Tai-Yi Huang, and Chia-Fu Hsu. Synergizing wireless communication technologies to improve Internet downloading experiences. IEEE Transactions on Computers, 63(11): 2851–2865, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2017:ACG

[LHH17] Peng Liu, Qi Hu, and Xingcheng Hua. Adaptive coherence granularity for multi-socket systems. IEEE Transactions on Computers, 66(8):1302–1312, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7867795/.

Liang:2013:AEM

[LHL13a] Jinghang Liang, Jie Han, and Fabrizio Lombardi. Analysis of error masking and restoring properties of sequential circuits. IEEE Transactions on Computers, 62(9):1694–1704, September 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liang:2013:NMR

[LHL13b] Jinghang Liang, Jie Han, and Fabrizio Lombardi. New metrics for the reliability of approximate and probabilistic adders. IEEE Transactions on Computers, 62(9): 1760–1771, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2015:CAR

[LHL^{+}15a] Qingan Li, Yanxiang He, Jianhua Li, Liang Shi, Yiran Chen, and C. J. Xue. Compiler-assisted refresh minimization for volatile STT-RAM cache. IEEE Transactions on Computers, 64(8):2169–2181, August 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2015:AFE

[LHL15b] Cong Liu, Jie Han, and F. Lombardi. An analytical framework for evaluating the error characteristics of approximate adders. IEEE Transactions on Computers, 64(5):1268–1281, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2015:IAR

[LHPH15] Chia-Wei Lee, Chao-Wen Huang, Wen-Hao Pi, and Sun-Yuan Hsieh. An improved approximation ratio to the partial-terminal Steiner tree problem. IEEE Transactions on Computers, 64(1):274–279, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2015:SDF

[LHTG15] Yanchao Lu, Bingsheng He, Xueyan Tang, and Minyi Guo. Synergy of dynamic frequency scaling and demotion on DRAM power management: Models and optimizations. IEEE Transactions on Computers, 64 (8):2367–2381, August 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2013:WLS

[LHY13] Ren Ping Liu, Mark Hedley, and Xun Yang. WLAN location service with TXOP. IEEE Transactions on Computers, 62(3):589–598, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2013:EMK

[LHYZ13] Jing Liu, Qiong Huang, Bo Yang, and Yang Zhang. Efficient multicast key distribution using HOWP-based dynamic group access structures. IEEE Transactions on Computers, 62(8):1656– 1672, August 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2012:TDC

[Li12a] Jin-Fu Li. Testing and diagnosing comparison faults of TCAMs with asymmetric cells. IEEE Transactions on Computers, 61(11): 1576–1587, November 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2012:SPC

[Li12b] Keqin Li. Scheduling precedence constrained tasks with reduced processor energy on multiprocessor computers. IEEE Transactions on Computers, 61(12):1668–1681, December 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2011:BHI

[Liu11] Ming T. Liu. A brief history of the IEEE transactions on computers. IEEE Transactions on Computers, 60(4):449, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Low:2013:MET

[LJ13] Joshua Yung Lih Low and Ching Chuen Jong. A memory-efficient tables-and-additions method for accurate computation of elementary functions. IEEE Transactions on Computers, 62(5):858–872, May 2013.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Low:2015:UMB

[LJ15] J. Y. L. Low and Ching Chuen Jong. Unified Mitchellbased approximation for efficient logarithmic conversion circuit. IEEE Transactions on Computers, 64 (6):1783–1797, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2018:TFE

[LJ18] Nian-Ze Lee and Jie-Hong R.

Jiang. Towards formal evaluation and verification of probabilistic design. IEEE Transactions on Computers, 67(8):1202–1216, August 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8302961/.

Li:2019:SSC

[LJG^{+}19] J. Li, S. Jiang, S. Gong, J. Wu, J. Yan, G. Yan, and X. Li. SqueezeFlow: A sparse CNN accelerator exploiting concise convolution rules. IEEE Transactions on Computers, 68(11): 1663–1677, November 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2013:GFS

[LJL13] Chia-Yu Lu, Shang-Ming Jen, and Chi-Sung Laih. A general framework of sidechannel atomicity for elliptic curve scalar multiplication. IEEE Transactions on Computers, 62 (3):428–438, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2018:GPE

[LJVJ18] Jie Lu, Hongyang Jia, Naveen Verma, and Niraj K. Jha. Genetic programming for energyefficient and energy-scalable approximate feature computation in embedded inference systems. IEEE Transactions on Computers, 67 (2):222–236, February 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8008802/.

Liu:2015:UFL

[LJY^{+}15] Wenping Liu, Hongbo Jiang, Yang Yang, Xiaofei Liao, Hongzhi Lin, and Zemeng Jin. A unified framework for line-like skeleton extraction in 2D/3D sensor networks. IEEE Transactions on Computers, 64(5):1323– 1335, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lange:2010:AEM

[LK10] H. Lange and A. Koch.

Architectures and execution models for hardware/ software compilation and their system-level realization. IEEE Transactions on Computers, 59(10):1363– 1377, October 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5374369.

Lee:2014:IPC

[LK14] Sungjin Lee and Jihong Kim.

Improving performance and capacity of flash storage devices by exploiting heterogeneity of MLC flash memory. IEEE Transactions on Computers, 63(10): 2445–2458, October 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lakshminarayana:2015:BPP

[LK15a] N. B. Lakshminarayana and H. Kim. Block-precise processors: Low-power processors with reduced operand store accesses and result broadcasts. IEEE Transactions on Computers, 64(11): 3102–3114, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2015:FDC

[LK15b] Jongmin Lee and Soontae Kim. Filter data cache: An energy-efficient small L0 data cache architecture driven by miss cost reduction. IEEE Transactions on Computers, 64(7):1927– 1939, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2016:ELA

[LK16a] S. Lee and J. Kim. Effective lifetime-aware dynamic throttling for NAND flashbased SSDs. IEEE Transactions on Computers, 65 (4):1075–1089, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2016:RDR

[LK16b] Yebin Lee and Soontae Kim. RAMS: DRAM rank-aware memory scheduling for energy saving. IEEE Transactions on Computers, 65(10): 3210–3216, October 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2018:SLA

[LK18] Joo Hwan Lee and Hyesoon Kim. StaleLearn: Learning acceleration with asynchronous synchronization between model replicas on PIM. IEEE Transactions on Computers, 67 (6):861–873, June 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8168367/.

Lee:2016:EPF

[LKBS16] E. Lee, H. Kang, H. Bahn, and K. G. Shin. Eliminating periodic flush overhead of file I/O with nonvolatile buffer cache. IEEE Transactions on Computers, 65(4):1145–1157, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2016:RMB

[LKH16] Junghoon Lee, Taehoon Kim, and Jaehyuk Huh. Reducing the memory bandwidth overheads of hardware security support for multi-core processors. IEEE Transactions on Computers, 65(11):3384–3397, November 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lai:2015:CHA

[LKJ15] B.-C. C. Lai, Hsien-Kai Kuo, and Jing-Yang Jou. A cache hierarchy aware thread mapping methodology for GPGPUs. IEEE Transactions on Computers, 64(4):884–898, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2017:IEE

[LKK^{+}17] S. Lee, K. Kim, G. Koo, H. Jeon, M. Annavaram, and W. W. Ro. Improving energy efficiency of GPUs through data compression and compressed execution. IEEE Transactions on Computers, 66(5):834–847, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2019:IBR

[LKK^{+}19] H. Lee, M. Kim, H. Kim, H. Kim, and H. Lee. Integration and boost of a read-modify-write module in phase change memory system. IEEE Transactions on Computers, 68(12): 1772–1784, December 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2013:MPG

[LKLK13] Duhee Lee, Jong-Chan Kim, Chang-Gun Lee, and Kanghee Kim. mRT-PLRU: A general framework for realtime multitask executions on NAND flash memory. IEEE Transactions on Computers, 62(4):758–771, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2015:ZCE

[LKLM15] Youngjae Lee, Jin-Soo Kim, Sang-Won Lee, and Seungryoul Maeng. Zombie chasing: Efficient flash management considering dirty data in the buffer cache. IEEE Transactions on Computers, 64 (2):569–581, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2012:GFB

[LKLT12] Chia-Hung Lin, Jian-Jhih Kuo, Bing-Hong Liu, and Ming-Jer Tsai. GPS-free, boundary-recognition-free, and reliable double-ruling-based information brokerage scheme in wireless sensor networks. IEEE Transactions on Computers, 61 (6):885–898, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lotfi-Kamran:2016:EHS

[LKMSA16] Pejman Lotfi-Kamran, Mehdi Modarressi, and Hamid Sarbazi-Azad. An efficient hybrid-switched network-on-chip for chip multiprocessors. IEEE Transactions on Computers, 65(5):1656– 1662, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2014:MAP

[LKS^{+}14] Junghoon Lee, Hanjoon Kim, Minjeong Shin, John Kim, and Jaehyuk Huh. Mutually aware prefetcher and on-chip network designs for multi-cores. IEEE Transactions on Computers, 63(9): 2316–2329, September 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2013:AAC

[LKT13] Cheng-Kuan Lin, Tzu-Liang Kung, and Jimmy J. M. Tan. An algorithmic approach to conditional-fault local diagnosis of regular multiprocessor interconnected systems under the PMC model. IEEE Transactions on Computers, 62(3):439–451, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2012:ASP

[LKYC12] Jae-Beom Lee, Myoung-Jin Kim, Sungroh Yoon, and Eui-Young Chung. Application-support particle filter for dynamic voltage scaling of multimedia applications. IEEE Transactions on Computers, 61(9): 1256–1269, September 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2011:FFR

[LL11] Yawei Li and Zhiling Lan.

FREM: a fast restart mechanism for general checkpoint/restart. IEEE Transactions on Computers, 60(5): 639–652, May 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2015:IBE

[LLC^{+}15] Jin Li, Jingwei Li, Xiaofeng Chen, Chunfu Jia, and Wenjing Lou. Identitybased encryption with outsourced revocation in cloud computing. IEEE Transactions on Computers, 64 (2):425–437, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lai:2016:UDH

[LLC^{+}16] Bo-Cheng Charles Lai, Chia-Ying Lee, Tsou-Han Chiu, Hsien-Kai Kuo, and Chun-Kai Chang. Unified designs for high performance LDPC decoding on GPGPU. IEEE Transactions on Computers, 65(12):3754–3765, December 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2013:APM

[LLCC13] Cheng-Hung Lin, Chen-Hsiung Liu, Lung-Sheng Chien, and Shih-Chieh Chang. Accelerating pattern matching using a novel parallel algorithm on GPUs. IEEE Transactions on Computers, 62(10):1906–1916, October 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2013:IBM

[LLCH13] Jonathan Lee, Shin-Jie Lee, Hsi-Min Chen, and Kuo-Hsun Hsu. Itinerary-based mobile agent as a basis for distributed OSGi services. IEEE Transactions on Computers, 62(10):1988– 2000, October 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2016:RMM

[LLD^{+}16] Lei Liu, Yong Li, Chen Ding, Hao Yang, and Chengyong Wu. Rethinking memory management in modern operating system: Horizontal, vertical or random? IEEE Transactions on Computers, 65(6):1921–1935, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2019:EEA

[LLD19] Peng Liu, Shunbin Li, and Qingyuan Ding. An energyefficient accelerator based on hybrid CPU–FPGA devices for password recovery. IEEE Transactions on Computers, 68(2):170– 181, February 2019. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8453825/.

Lu:2015:ASD

[LLHC15] Shyue-Kung Lu, Tsu-Lin Li, M. Hashizume, and Jiann-Liang Chen. Address scrambling and data inversion techniques for yield enhancement of NROM-based ROMs. IEEE Transactions on Computers, 64(5):1230– 1240, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2018:GEI

[LLK18] Zhe Liu, Patrick Longa, and Çetin Kaya Koç. Guest Editors’ introduction to the special issue on cryptographic engineering in a postquantum world: State of the art advances. IEEE Transactions on Computers, 67(11): 1532–1534, November 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8485531/.

Liu:2019:SBC

[LLKA19] Jian Liu, Wenting Li, Ghassan O. Karame, and N. Asokan. Scalable Byzantine consensus via hardwareassisted secret sharing. IEEE Transactions on Computers, 68(1):139–151, January 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8419336/.

Lin:2011:HAB

[LLL11] Po-Ching Lin, Ying-Dar Lin, and Yuan-Cheng Lai. A hybrid algorithm of backward hashing and automaton tracking for virus scanning. IEEE Transactions on Computers, 60(4):594– 601, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2015:ECF

[LLL15] Runhui Li, Jian Lin, and P. P. C. Lee. Enabling concurrent failure recovery for regenerating-coding-based storage systems: From theory to practice. IEEE Transactions on Computers, 64(7):1898–1911, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2016:ARD

[LLL16] Y. Li, P. P. C. Lee, and J. C. S. Lui. Analysis of reliability dynamics of SSD RAID. IEEE Transactions on Computers, 65 (4):1131–1144, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2017:DNN

[LLL^{+}17] Tao Luo, Shaoli Liu, Ling Li, Yuqing Wang, Shijin Zhang, Tianshi Chen, Zhiwei Xu, Olivier Temam, and Yunji Chen. DaDian-Nao: A neural network supercomputer. IEEE Transactions on Computers, 66 (1):73–88, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2013:PAD

[LLLJ13] Fangming Liu, Bo Li, Baochun Li, and Hai Jin. Peer-assisted on-demand streaming: Characterizing demands and optimizing supplies. IEEE Transactions on Computers, 62(2): 351–361, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lim:2014:ABF

[LLLP14] Hyesook Lim, Kyuhee Lim, Nara Lee, and Kyong-Hye Park. On adding Bloom filters to longest prefix matching algorithms. IEEE Transactions on Computers, 63 (2):411–423, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2015:CPM

[LLM^{+}15] Xiulong Liu, Keqiu Li, Geyong Min, Yanming Shen, A. X. Liu, and Wenyu Qu. Completely pinpointing the missing RFID tags in a time-efficient way. IEEE Transactions on Computers, 64(1):87–96, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2013:QSA

[LLOS13] Liang Lu, Weiqiang Liu, Maire O’Neill, and Earl E. Swartzlander. QCA systolic array design. IEEE Transactions on Computers, 62(3):548–560, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2014:LCS

[LLQ^{+}14] Peng Li, David J. Lilja, Weikang Qian, Marc D. Riedel, and Kia Bazargan. Logical computation on stochastic bit streams with linear finite-state machines. IEEE Transactions on Computers, 63(6):1474–1486, June 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2016:TFD

[LLS^{+}16] Joseph K. Liu, Kaitai Liang, Willy Susilo, Jianghua Liu, and Yang Xiang. Twofactor data security protection mechanism for cloud storage system. IEEE Transactions on Computers, 65 (6):1992–2004, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2011:IRT

[LLW^{+}11] Miao Liu, Duo Liu, Yi Wang, Meng Wang, and Zili Shao. On improving real-time interrupt latencies of hybrid operating systems with twolevel hardware interrupts. IEEE Transactions on Computers, 60(7):978–991, July 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2017:BSR

[LLW^{+}17] Hsuan-Hung Liu, Bing-Yang Lin, Cheng-Wen Wu, Wan-Ting Chiang, Lee Mincent, Hung-Chih Lin, Ching-Nen Peng, and Min-Jer Wang. A built-off self-repair scheme for channel-based 3D memories. IEEE Transactions on Computers, 66 (8):1293–1301, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7850958/.

Liu:2018:SCM

[LLW^{+}18] Yidong Liu, Siting Liu, Yanzhi Wang, Fabrizio Lombardi, and Jie Han. A stochastic computational multi-layer perceptron with backward propagation. IEEE Transactions on Computers, 67(9):1273–1286, September 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8319953/.

Li:2017:TCA

[LLX^{+}17] Jianhua Li, Minming Li, Chun Jason Xue, Yiming Ouyang, and Fanfan Shen. Thread criticality assisted replication and migration for chip multiprocessor caches. IEEE Transactions on Computers, 66(10): 1747–1762, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7931700/.

Li:2016:SAD

[LLXC16] Jingwei Li, Jin Li, Dongqing Xie, and Zhang Cai. Secure auditing and deduplicating data in cloud. IEEE Transactions on Computers, 65 (8):2386–2396, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2017:AMC

[LLZ^{+}17] Jun-Wei Li, Shi-Ning Li, Yu Zhang, Tao Gu, Yee Wei Law, Zhe Yang, Xingshe Zhou, and Marimuthu Palaniswami. An analytical model for coding-based reprogramming protocols in lossy wireless sensor networks. IEEE Transactions on Computers, 66(1):24–37, January 2017. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lombardi:2013:RST

[LMB13] Michele Lombardi, Michela Milano, and Luca Benini. Robust scheduling of task graphs under execution time uncertainty. IEEE Transactions on Computers, 62 (1):98–111, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2016:DMC

[LMB^{+}16] Yuqi Lin, Saif U. R. Malik, Kashif Bilal, Qiusong Yang, Yongji Wang, and Samee U. Khan. Designing and modeling of covert channels in operating systems. IEEE Transactions on Computers, 65(6):1706–1719, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2017:UMF

[LMB17] S. Liu, G. Mingas, and C. S. Bouganis. An unbiased MCMC FPGA-based accelerator in the land of custom precision arithmetic. IEEE Transactions on Computers, 66(5):745–758, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luque:2012:CAM

[LMC^{+}12] Carlos Luque, Miquel Moreto, Francisco J. Cazorla, Roberto Gioiosa, Alper Buyuktosunoglu, and Mateo Valero. CPU accounting for multicore processors. IEEE Transactions on Computers, 61 (2):251–264, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Le:2015:TFB

[LMC^{+}15] Yuan Le, Liran Ma, Wei Cheng, Xiuzhen Cheng, and Biao Chen. A time fairnessbased MAC algorithm for throughput maximization in 802.11 networks. IEEE Transactions on Computers, 64(1):19–31, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2014:PDC

[LMJ14] Yong Li, R. Melhem, and A. K. Jones. A practical data classification framework for scalable and high performance chip-multiprocessors. IEEE Transactions on Computers, 63(12):2905–2918, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2011:PAM

[LMNP11] Chang Joo Lee, Onur Mutlu, Veynu Narasiman, and Yale N. Patt. Prefetch-aware memory controllers. IEEE Transactions on Computers, 60(10):1406–1430, October 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=6008537.

Luong:2013:GCP

[LMT13] The Van Luong, Nouredine Melab, and El-Ghazali Talbi. GPU computing for parallel local search metaheuristic algorithms. IEEE Transactions on Computers, 62 (1):173–185, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2017:MFN

[LMZQ17] Yin Li, Xingpo Ma, Yu Zhang, and Chuanda Qi. Mastrovito form of non-recursive Karatsuba multiplier for all trinomials. IEEE Transactions on Computers, 66(9): 1573–1584, September 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7870679/.

Liu:2012:PED

[LN12] Wei Liu and Alberto Nannarelli. Power efficient division and square root unit. IEEE Transactions on Computers, 61(8):1059–1070, August 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lao:2018:FPS

[LNCX18] Bin Lao, Ge Nong, Wai Hong Chan, and Jing Yi Xie. Fast in-place suffix sorting on a multicore computer. IEEE Transactions on Computers, 67(12): 1737–1749, December 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8371211/.

Liu:2019:OMM

[LNL^{+}19] W. Liu, J. Ni, Z. Liu, C. Liu, and M. O. Neill. Optimized modular multiplication for supersingular isogeny Diffie– Hellman. IEEE Transactions on Computers, 68 (8):1249–1255, August 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2016:LEC

[LOC^{+}16] C. W. Liu, S. H. Ou, K. C. Chang, T. C. Lin, and S. K. Chen. A low-error, costefficient design procedure for evaluating logarithms to be used in a logarithmic arithmetic processor. IEEE Transactions on Computers, 65(4):1158–1164, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2017:RSC

[LOH17] Wenjie Li, Sharief M. A.

Oteafy, and Hossam S. Hassanein. Rate-selective caching for adaptive streaming over information-centric networks. IEEE Transactions on Computers, 66(9): 1613–1628, September 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7887688/.

Lombardi:2010:SJ

[Lom10] Fabrizio Lombardi. State of the journal. IEEE Transactions on Computers, 59 (2):145–149, February 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5372165.

Lombardi:2011:E

[Lom11] Fabrizio Lombardi. Editorial. IEEE Transactions on Computers, 60(1):1, January 2011. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Louri:2019:EIE

[Lou19]  Ahmed Louri.  Editorial

from the incoming Editorin-Chief. IEEE Transactions on Computers, 68(1): 3, January 2019. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8573058/.

Li:2013:ONJ

[LOX^{+}13] Yu Li, Sai Tung On, Jianliang Xu, Byron Choi, and Haibo Hu. Optimizing nonindexed join processing in flash storage-based systems. IEEE Transactions on Computers, 62(7):1417– 1431, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Le:2012:STB

[LP12] Hoang Le and Viktor K. Prasanna. Scalable treebased architectures for IPv4/ v6 lookup using prefix partitioning. IEEE Transactions on Computers, 61 (7):1026–1039, July 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Le:2013:MEM

[LP13a] Hoang Le and Viktor K. Prasanna. A memoryefficient and modular approach for large-scale string pattern matching. IEEE Transactions on Computers, 62(5):844–857, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liaskos:2013:GSR

[LP13b] Christos K. Liaskos and Georgios I. Papadimitriou. Generalizing the square root rule for optimal periodic scheduling in push-based wireless environments. IEEE Transactions on Computers, 62(5):1044–1050, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Langhammer:2017:SPL

[LP17] Martin Langhammer and Bogdan Pasca. Single precision logarithm and exponential architectures for hard floating-point enabled FPGAs. IEEE Transactions on Computers, 66(12): 2031–2043, December 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7927449/.

Liu:2014:DEC

[LPCW14] Cong Liu, Yan Pan, Ai Chen, and Jie Wu. A DFA with extended character-set for fast deep packet inspection. IEEE Transactions on Computers, 63(8):1925–1937, August 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lukefahr:2016:EFG

[LPD^{+}16] A. Lukefahr, S. Padmanabha, R. Das, F. M. Sleiman, R. G. Dreslinski, T. F. Wenisch, and S. Mahlke. Exploring fine-grained heterogeneity with composite cores. IEEE Transactions on Computers, 65(2):535– 547, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luiz:2010:MSW

[LPL10] S. O. D. Luiz, A. Perkusich, and A. M. N. Lima. Multisize sliding window in workload estimation for dynamic power management. IEEE Transactions on Computers, 59(12):1625–1639, December 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5453349.

Liu:2012:NCU

[LPL12] Jianguo Liu, Chao Pan, and Zhenbing Liu. Novel convolutions using first-order moments. IEEE Transactions on Computers, 61 (7):1050–1056, July 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lai:2013:SLR

[LPL^{+}13] Hanjiang Lai, Yan Pan, Cong Liu, Liang Lin, and Jie Wu. Sparse learning-to-rank via an efficient primaldual algorithm. IEEE Transactions on Computers, 62 (6):1221–1233, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lima:2010:KBA

[LPW10] J. B. Lima, D. Panario, and Qiang Wang. A Karatsubabased algorithm for polynomial multiplication in Chebyshev form. IEEE Transactions on Computers, 59(6):835–841, June 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5416688.

Liu:2016:PPO

[LQD^{+}16] Ximeng Liu, Baodong Qin, Robert H. Deng, Rongxing Lu, and Jianfeng Ma. A privacy-preserving outsourced functional computation framework across largescale multiple encrypted domains. IEEE Transactions on Computers, 65(12): 3567–3579, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2017:DAR

[LQW^{+}17] Weiqiang Liu, Liangyu Qian, Chenghua Wang, Honglan Jiang, Jie Han, and Fabrizio Lombardi. Design of approximate radix-4 Booth multipliers for error-tolerant computing. IEEE Transactions on Computers, 66 (8):1435–1441, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7862783/.

Lentaris:2010:GPM

[LR10] G. Lentaris and D. Reisis. A graphics parallel memory organization exploiting request correlations. IEEE Transactions on Computers, 59(6):762–775, June 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5416691.

Lai:2013:LLR

[LR13] Shouwen Lai and Binoy Ravindran. Least-latency routing over time-dependent wireless sensor networks. IEEE Transactions on Computers, 62(5):969–983, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2016:PGA

[LR16] S. Lee and W. W. Ro. Parallel GPU architecture simulation framework exploiting architectural-level parallelism with timing error prediction. IEEE Transactions on Computers, 65 (4):1253–1265, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2018:SST

[LR18] Changmin Lee and Won Woo Ro. Simultaneous and speculative thread migration for improving energy efficiency of heterogeneous core architectures. IEEE Transactions on Computers, 67 (4):498–512, April 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8097407/.

Lai:2010:HQB

[LRC10] Shouwen Lai, B. Ravindran, and Hyeonjoong Cho. Heterogeneous quorum-based wake-up scheduling in wireless sensor networks. IEEE Transactions on Computers, 59(11):1562–1575, November 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5396329.

Lee:2016:EMF

[LRP16] J. Lee, H. Roh, and S. Park. External Mergesort for flash-based solid state drives. IEEE Transactions on Computers, 65 (5):1518–1527, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2018:MFM

[LRP^{+}18] Jaewoo Lee, Saravanan Ramanathan, Kieu-My Phan, Arvind Easwaran, Insik Shin, and Insup Lee. MC-Fluid: Multi-core fluid-based mixed-criticality scheduling. IEEE Transactions on Computers, 67(4):469–483, April 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8059775/.

Liu:2015:MDT

[LRY^{+}15] Chang Liu, R. Ranjan, Chi Yang, Xuyun Zhang, Lizhe Wang, and Jinjun Chen. MuR-DPA: Top-down levelled multi-replica Merkle hash tree based secure public auditing for dynamic big data storage on cloud. IEEE Transactions on Computers, 64(9):2609–2622, September 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2010:DHP

[LS10a] Mingqiang Li and Jiwu Shu. DACO: a highperformance disk architecture designed specially for large-scale erasure-coded storage systems. IEEE Transactions on Computers, 59(10):1350–1362, October 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5396331. See

comment [Tho12].

Li:2010:PSD

[LS10b] Mingqiang Li and Jiwu Shu. Preventing silent data corruptions from propagating during data reconstruction. IEEE Transactions on Computers, 59(12): 1611–1624, December 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5416679.

Lu:2010:RPS

[LS10c] Wencheng Lu and S. Sahni.

Recursively partitioned static IP router tables. IEEE Transactions on Computers, 59(12):1683–1690, December 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5487500.

Li:2013:SES

[LS13] Ze Li and Haiying Shen. SE-DUM: Exploiting social networks in utility-based distributed routing for DTNs. IEEE Transactions on Computers, 62(1):83–97, January 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2017:FGC

[LSA^{+}17] Tuo Li, Muhammad Shafique, Jude Angelo Ambrose, Jörg Henkel, and Sri Parameswaran. Fine-grained checkpoint recovery for application-specific instruction-set processors. IEEE Transactions on Computers, 66(4):647–660, April 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2018:AAW

[LSA18] Haeseung Lee, Muhammad Shafique, and Mohammad Abdullah Al Faruque. Aging-aware workload management on embedded GPU under process variation. IEEE Transactions on Computers, 67(7):920–933, July 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8247279/.

Lee:2010:PTA

[LSC10] Jong Sung Lee, K. Skadron, and Sung Woo Chung. Predictive temperature-aware DVFS. IEEE Transactions on Computers, 59 (1):127–133, January 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5255230.

Lam:2011:AAT

[LSC11] Siew Kei Lam, Thambipillai Srikanthan, and Christopher T. Clarke. Architectureaware technique for mapping area-time efficient custom instructions onto FP-GAs. IEEE Transactions on Computers, 60(5):680– 692, May 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2015:HPL

[LSG^{+}15] Youyou Lu, Jiwu Shu, Jia Guo, Shuai Li, and O. Mutlu. High-performance and lightweight transaction support in flash-based SSDs. IEEE Transactions on Computers, 64(10):2819–2832, October 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2018:ACG

[LSG^{+}18] Qiao Li, Liang Shi, Congming Gao, Yejia Di, and Chun Jason Xue. Access characteristic guided read and write regulation on flash based storage systems. IEEE Transactions on Computers, 67(12): 1663–1676, December 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8362696/.

Lu:2016:SSC

[LSGZ16] Y. Lu, J. Shu, J. Guo, and P. Zhu. Supporting system consistency with differential transactions in flashbased SSDs. IEEE Transactions on Computers, 65 (2):627–639, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2015:EGD

[LSHC15] Dan Li, Yunfei Shang, Wu He, and Congjie Chen. EXR: Greening data center network with software defined exclusive routing. IEEE Transactions on Computers, 64(9):2534–2544, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2013:BBA

[LSK13] Sungjin Lee, Dongkun Shin, and Jihong Kim. BAGC: Buffer-aware garbage collection for flash-based storage systems. IEEE Transactions on Computers, 62(11): 2141–2154, November 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2015:AEC

[LSL15] F. Liu, P. Shu, and J. C.

Lui. AppATP: An energy conserving adaptive mobile-cloud transmission protocol. IEEE Transactions on Computers, 64(11): 3051–3063, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2013:LSN

[LSS13] Ze Li, Haiying Shen, and Karan Sapra. Leveraging social networks to combat collusion in reputation systems for peer-to-peer networks. IEEE Transactions on Computers, 62(9): 1745–1759, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2015:CSA

[LSSE15] Jinkyu Lee, K. G. Shin, I. Shin, and A. Easwaran. Composition of schedulability analyses for realtime multiprocessor systems. IEEE Transactions on Computers, 64(4):941–954, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2015:ETP

[LSW15] Guoxin Liu, Haiying Shen, and L. Ward. An efficient and trustworthy P2P and social network integrated file sharing system. IEEE Transactions on Computers, 64(1):54–70, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liang:2013:AAC

[LSX13] Weifa Liang, Pascal Schweitzer, and Zichuan Xu. Approximation algorithms for capacitated minimum forest problems in wireless sensor networks with a mobile sink. IEEE Transactions on Computers, 62(10): 1932–1944, October 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2014:EES

[LSXP14] Jianqiang Luo, Mochan Shrestha, Lihao Xu, and James S. Plank. Efficient encoding schedules for XORbased erasure codes. IEEE Transactions on Computers, 63(9):2259–2272, September 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2015:POB

[LSZ^{+}15] Liang Liu, Yuning Song, Haiyang Zhang, Huadong Ma, and A. V. Vasilakos. Physarum optimization: A biology-inspired algorithm for the Steiner tree problem in networks. IEEE Transactions on Computers, 64(3):819–832, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Le:2014:IMX

[LT14] Duc-Phong Le and Chik How Tan. Improved Miller’s algorithm for computing pairings on Edwards curves. IEEE Transactions on Computers, 63(10):2626–2632, October 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liaskos:2015:PRU

[LT15] C. Liaskos and A. Tsioliaridou. A promise of realizable, ultra-scalable communications at nano-scale: A multi-modal nano-machine architecture. IEEE Transactions on Computers, 64 (5):1282–1295, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Leu:2014:NPR

[LTL14] Jenq-Shiou Leu, Nguyen Hai Tung, and Chun-Yao Liu. Non-parametric RSS prediction based energy saving scheme for moving smartphones. IEEE Transactions on Computers, 63(7):1793– 1801, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2012:OHD

[LTLC12] Qiang Liu, Tim Todman, Wayne Luk, and George A. Constantinides. Optimizing hardware design by composing utility-directed transformations. IEEE Transactions on Computers, 61(12): 1800–1812, December 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liaskos:2014:DCB

[LTP^{+}14] Christos K. Liaskos, Angeliki N. Tsioliaridou, Georgios I. Papadimitriou, Petros Nicopolitidis, and Andreas S. Pomportsis. On data compatibility and broadcast stream formation. IEEE Transactions on Computers, 63(9):2369–2375, September 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2015:SPC

[LTVL15] Kenli Li, Xiaoyong Tang, B. Veeravalli, and Keqin Li. Scheduling precedence constrained stochastic tasks on heterogeneous cluster systems. IEEE Transactions on Computers, 64(1):191– 204, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2012:PPG

[LTW^{+}12] Shaoshan Liu, Jie Tang, Ligang Wang, Xiao-Feng Li, and Jean-Luc Gaudiot. Packer: Parallel garbage collection based on virtual spaces. IEEE Transactions on Computers, 61(11): 1611–1623, November 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lora:2019:TAI

[LVF19] M. Lora, S. Vinco, and F. Fummi. Translation, abstraction and integration for effective smart system design. IEEE Transactions on Computers, 68(10): 1525–1538, October 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2012:MRP

[LvH12] Xin Li and Reinhard von Hanxleden. Multithreaded reactive programming — the Kiel Esterel processor. IEEE Transactions on Computers, 61(3):337–349, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2016:CSP

[LVJ16] Jie Lu, Naveen Verma, and Niraj K. Jha. Compressed signal processing on Nyquistsampled signals. IEEE Transactions on Computers, 65(11):3293–3303, November 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lerman:2018:SST

[LVMS18] Liran Lerman, Nikita Veshchikov, Olivier Markowitch, and François-Xavier Standaert. Start simple and then refine: Bias-variance decomposition as a diagnosis tool for leakage profiling. IEEE Transactions on Computers, 67 (2):268–283, February 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7990260/.

Liu:2013:FDP

[LW13] Cong Liu and Jie Wu. Fast deep packet inspection with a dual finite automata. IEEE Transactions on Computers, 62(2):310–321, February 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2015:DCN

[LW15] D. Li and J. Wu. On data center network architectures for interconnecting dual-port servers. IEEE Transactions on Computers, 64(11): 3210–3222, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2017:FTF

[LW17] Kibum Lee and S. Simon Wong. Fault-tolerant FPGA with column-based redundancy and power gating using RRAM. IEEE Transactions on Computers, 66(6):946–956, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7763756/.

Lu:2013:ODR

[LWF13] Zaixin Lu, Weili Wu, and Bin Fu. Optimal data retrieval scheduling in the multichannel wireless broadcast environments. IEEE Transactions on Computers, 62(12):2427–2439, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2017:TSE

[LWF^{+}17] Zheng Li, Fang Wang, Dan Feng, Yu Hua, Jingning Liu, Wei Tong, Yu Chen, and Salah S. Harb. Time and space-efficient write parallelism in PCM by exploiting data patterns. IEEE Transactions on Computers, 66(9): 1629–1644, September 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7870591/.

Lu:2016:RAD

[LWH^{+}16] Y. Lu, D. Wu, B. He, X. Tang, J. Xu, and M. Guo. Rank-aware dynamic migrations and adaptive demotions for DRAM power management. IEEE Transactions on Computers, 65 (1):187–202, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2011:EXR

[LWK11] Chin-Yung Lu, Shiou-An Wang, and Sy-Yen Kuo. An extended XQDD representation for multiplevalued quantum logic. IEEE Transactions on Computers, 60(10):1377–1389, October 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5928328.

Lee:2015:GCE

[LWKA15] J. Lee, D. H. Woo, H. Kim, and M. Azimi. GREEN cache: Exploiting the disciplined memory model of OpenCL on GPUs. IEEE Transactions on Computers, 64(11):3167–3180, November 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2016:LPS

[LWL^{+}16] B. Lin, C. Wu, M. Lee, H. Lin, C. Peng, and M. Wang. A local parallel search approach for memory failure pattern identification. IEEE Transactions on Computers, 65 (3):770–780, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2011:CDC

[LWW11] Xiang-Yang Li, Yajun Wang, and Yu Wang. Complexity of data collection, aggregation, and selection for wireless sensor networks. IEEE Transactions on Computers, 60(3):386–399, March 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2015:CHS

[LWY15] Ting-Yu Lin, Kun-Ru Wu, and Guang-Chuen Yin. Channel-hopping scheme and channel-diverse routing in static multi-radio multihop wireless networks. IEEE Transactions on Computers, 64(1):71–86, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liang:2018:FMP

[LWZ18] Yun Liang, Shuo Wang, and Wei Zhang. FlexCL: A model of performance and power for OpenCL workloads on FPGAs. IEEE Transactions on Computers, 67(12):

1750–1764, December 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8365849/.

Levitin:2017:DCP

[LXDV17] Gregory Levitin, Liudong Xing, Yuanshun Dai, and Vinod M. Vokkarane. Dynamic checkpointing policy in heterogeneous real-time standby systems. IEEE Transactions on Computers, 66(8):1449–1456, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7851034/.

Levitin:2015:MRC

[LXJD15] G. Levitin, Liudong Xing, B. W. Johnson, and Yuanshun Dai. Mission reliability, cost and time for cold standby computing systems with periodic backup. IEEE Transactions on Computers, 64(4):1043–1057, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2012:FDM

[LXK12] Xiaomei Liu, Li Xiao, and Andrew Kreling. A fully distributed method to detect and reduce cut vertices in large-scale overlay networks. IEEE Transactions on Computers, 61(7):969– 985, July 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2013:WED

[LXL^{+}13] Jianjun Li, Ming Xiong, Victor C. S. Lee, LihChyun Shu, and Guohui Li. Workloadefficient deadline and period assignment for maintaining temporal consistency under EDF. IEEE Transactions on Computers, 62 (6):1255–1268, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2014:RMD

[LXL^{+}14] Dan Li, Mingwei Xu, Ying Liu, Xia Xie, Yong Cui, Jingyi Wang, and Guihai Chen. Reliable multicast in data center networks. IEEE Transactions on Computers, 63(8):2011–2024, August 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2019:RND

[LXW^{+}19] Shuo Li, Nong Xiao, Peng Wang, Guangyu Sun, Xiaoyang Wang, Yiran Chen, Hai Helen Li, Jason Cong, and Tao Zhang. RC-NVM: Dual-addressing nonvolatile memory architecture supporting both row and column memory accesses. IEEE Transactions on Computers, 68(2): 239–254, February 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8453833/.

Liu:2015:STE

[LXZ^{+}15] X. Liu, B. Xiao, S. Zhang, K. Bu, and A. Chan. STEP: A time-efficient tag searching protocol in large RFID systems. IEEE Transactions on Computers, 64(11): 3265–3277, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2016:DRN

[LXZH16] Limei Lin, Li Xu, Shuming Zhou, and Sun-Yuan Hsieh. The t/k-diagnosability for regular networks. IEEE Transactions on Computers, 65(10):3157–3170, October 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2011:EEQ

[LY11] Sookyoung Lee and Mohamed F. Younis. EQAR: Effective QoS-aware relay node placement algorithm for connecting disjoint wireless sensor subnetworks. IEEE Transactions on Computers, 60(12):1772–1787, December 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2017:MPF

[LY17] Zhonghai Lu and Yuan Yao.

Marginal performance: Formalizing and quantifying power over/under provisioning in NoC DVFS. IEEE Transactions on Computers, 66(11):1903–1917, November 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7947135/.

Lu:2018:TVD

[LY18] Zhonghai Lu and Yuan Yao.

Thread voting DVFS for manycore NoCs. IEEE Transactions on Computers, 67(10):1506–1524, October 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8338086/.

Lee:2015:DIJ

[LYB15] Eunji Lee, Seung Hoon Yoo, and Hyokyung Bahn. Design and implementation of a journaling file system for phase-change memory. IEEE Transactions on Computers, 64(5):1349–1360, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2010:PAI

[LYCT10] Chia-Hung Lin, Shiao-An Yuan, Shih-Wei Chiu, and Ming-Jer Tsai. Progress-Face: An algorithm to improve routing efficiency of GPSR-like routing protocols in wireless ad hoc networks. IEEE Transactions on Computers, 59(6):822–834, June 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5416690.

Lai:2011:TCI

[LYH11] Chun-Hung Lai, Fu-Ching Yang, and Ing-Jer Huang. A trace-capable instruction cache for cost-efficient realtime program trace compression in SoC. IEEE Transactions on Computers, 60(12): 1665–1677, December 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5601688.

Liu:2018:TAT

[LYJ^{+}18] Weichen Liu, Lei Yang, Weiwen Jiang, Liang Feng, Nan Guan, Wei Zhang, and Nikil Dutt. Thermal-aware task mapping on dynamically reconfigurable networkon-chip based multiprocessor system-on-chip. IEEE Transactions on Computers, 67(12):1818–1834, December 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8373717/.

Le:2019:ADF

[LYK19] D. Le, S. L. Yeo, and K. Khoo. Algebraic differential fault analysis on SIMON block cipher. IEEE Transactions on Computers, 68(11): 1561–1572, November 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2019:PHB

[LYL^{+}19] W. Lu, G. Yan, J. Li, S. Gong, S. Jiang, J. Wu, and X. Li. Promoting the harmony between sparsity and regularity: A relaxed synchronous architecture for convolutional neural networks. IEEE Transactions on Computers, 68(6):867– 881, June 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lasla:2015:EAB

[LYOB15] N. Lasla, M. F. Younis, A. Ouadjaout, and N. Badache. An effective area-based localization algorithm for wireless networks. IEEE Transactions on Computers, 64(8):2103– 2118, August 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lim:2010:PTI

[LYS10] Hyesook Lim, Changhoon Yim, and E. E. Swartzlander. Priority tries for IP address lookup. IEEE Transactions on Computers, 59(6):784–794, June 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5416681.

Lin:2014:SMM

[LYS14] Wen-Ching Lin, Jheng-Hao Ye, and Ming-Der Shieh. Scalable Montgomery modular multiplication architecture with low-latency and low-memory bandwidth requirement. IEEE Transactions on Computers, 63 (2):475–483, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2016:OPT

[LYT^{+}16] Fan Li, Zhiyuan Yin, Shaojie Tang, Yu Cheng, and Yu Wang. Optimization problems in throwboxassisted delay tolerant networks: Which throwboxes to activate? How many active ones I need? IEEE Transactions on Computers, 65(5):1663–1670, May 2016.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2016:ECD

[LYY16] W. Li, Y. Yang, and D. Yuan. Ensuring cloud data reliability with minimum replication by proactive replica checking. IEEE Transactions on Computers, 65(5):1494–1506, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2014:EMA

[LZ14] Geng Lin and Wenxing Zhu.

An efficient memetic algorithm for the max-bisection problem. IEEE Transactions on Computers, 63 (6):1365–1376, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lama:2015:CPP

[LZ15] P. Lama and Xiaobo Zhou.

Coordinated power and performance guarantee with fuzzy MIMO control in virtualized server clusters. IEEE Transactions on Computers, 64(1):97–111, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2016:DRA

[LZA^{+}16] Zhihong Liu, Qi Zhang, Reaz Ahmed, Raouf Boutaba, Yaping Liu, and Zhenghu Gong. Dynamic resource allocation for MapReduce with partitioning skew. IEEE Transactions on Computers, 65(11):3304–3317, November 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2013:DFT

[LZS^{+}13] Yong Lu, Youjian Zhao, Fuchun Sun, Hongbo Li, and Dianjun Wang. Dynamic fault-tolerant routing based on FSA for LEO satellite networks. IEEE Transactions on Computers, 62(10): 1945–1958, October 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2015:PTR

[LZW^{+}15] Gang Lu, Jianfeng Zhan, Haining Wang, Lin Yuan, Yunwei Gao, Chuliang Weng, and Yong Qi. PowerTracer: Tracing requests in multitier services to reduce energy inefficiency. IEEE Transactions on Computers, 64 (5):1389–1401, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2013:SPD

[LZYL13] Xinxin Liu, Han Zhao, Xin Yang, and Xiaolin Li. Sink-Trail: a proactive data reporting protocol for wireless sensor networks. IEEE Transactions on Computers, 62(1):151–162, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2016:TSF

[LZZ16] Jinzhao Liu, Yuezhi Zhou, and Di Zhang. TranSim: A simulation framework for cache-enabled transparent computing systems. IEEE Transactions on Computers, 65(10):3171–3183, October 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2017:DRM

[LZZ17a] Xingquan Li, Ziran Zhu, and Wenxing Zhu. Discrete relaxation method for triple patterning lithography layout decomposition. IEEE Transactions on Computers, 66 (2):285–298, February 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2017:NVM

[LZZ^{+}17b] Duo Liu, Kan Zhong, Xiao Zhu, Yang Li, Lingbo Long, and Zili Shao. Non-volatile memory based page swapping for building highperformance mobile devices. IEEE Transactions on Computers, 66(11): 1918–1931, November 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7938390/.

Liu:2016:SOC

[LZZV16] T. Liu, Y. Zhu, Q. Zhang, and A. Vasilakos. Stochastic optimal control for participatory sensing systems with heterogeneous requests. IEEE Transactions on Computers, 65(5):1619– 1631, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2013:TMM

[LZZZ13] Jiang Lin, Hongzhong Zheng, Zhichun Zhu, and Zhao Zhang. Thermal modeling and management of DRAM systems. IEEE Transactions on Computers, 62(10): 2069–2082, October 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Manoochehri:2014:ELC

[MAD14] M. Manoochehri, M. Annavaram, and M. Dubois. Extremely low cost error protection with correctable parity protected cache. IEEE Transactions on Computers, 63(10):2431–2444, October 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Marchal:2017:HEU

[MAG^{+}17] Samuel Marchal, Giovanni Armano, Tommi Gröndahl, Kalle Saari, Nidhi Singh, and N. Asokan. Off-the-hook: An efficient and usable clientside phishing prevention application. IEEE Transactions on Computers, 66(10): 1717–1733, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7926371/.

Minaeva:2018:TTC

[MAHD18] Anna Minaeva, Benny Akesson, Zdeněk Hanzálek, and Dakshina Dasari. Timetriggered co-scheduling of computation and communication with jitter requirements. IEEE Transactions on Computers, 67 (1):115–129, January 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7967685/.

Mann:2016:MAV

[Man16] Zoltán Ádám Mann. Multicoreaware virtual machine placement in cloud data centers. IEEE Transactions on Computers, 65(11):3357–3369, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Marongiu:2012:OCE

[MB12a] Andrea Marongiu and Luca Benini. An OpenMP compiler for efficient use of distributed scratchpad memory in MPSoCs. IEEE Transactions on Computers, 61 (2):222–236, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Martirosyan:2012:PTR

[MB12b] Anahit Martirosyan and Azzedine Boukerche. Preserving temporal relationships of events for wireless sensor actor networks. IEEE Transactions on Computers, 61(8):1203–1216, August 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mingas:2016:PBM

[MB16] G. Mingas and C. S. Bouganis. Population-based MCMC on multi-core CPUs, GPUs and FPGAs. IEEE Transactions on Computers, 65(4):1283–1296, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Melani:2017:SAC

[MBB^{+}17] Alessandra Melani, Marko Bertogna, Vincenzo Bonifaci, Alberto Marchetti-Spaccamela, and Giorgio Buttazzo. Schedulability analysis of conditional parallel task graphs in multicore systems. IEEE Transactions on Computers, 66 (2):339–353, February 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Morari:2013:SMI

[MBC^{+}13] A. Morari, C. Boneti, F. J.

Cazorla, R. Gioiosa, Chen-Yong Cher, A. Buyuktosunoglu, P. Bose, and M. Valero. SMT malleability in IBM POWER5 and POWER6 processors. IEEE Transactions on Computers, 62(4):813–826, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Michelogiannakis:2011:EEB

[MBD11] George Michelogiannakis, Daniel U. Becker, and William J. Dally. Evaluating elastic buffer and wormhole flow control. IEEE Transactions on Computers, 60(6):896–903, June 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Melani:2017:ERT

[MBD^{+}17] Alessandra Melani, Marko Bertogna, Robert I. Davis, Vincenzo Bonifaci, Alberto Marchetti-Spaccamela, and Giorgio Buttazzo. Exact response time analysis for fixed priority memory-processor co-scheduling. IEEE Transactions on Computers, 66 (4):631–646, April 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Migliore:2018:PPF

[MBF18] Vincent Migliore, Guillaume Bonnoron, and Caroline Fontaine. Practical parameters for somewhat homomorphic encryption schemes on binary circuits. IEEE Transactions on Computers, 67(11):1550–1560, November 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8302942/.

Moreira:2010:BSR

[MBGS10] O. Moreira, T. Basten, M. Geilen, and S. Stuijk. Buffer sizing for rateoptimal single-rate data-flow scheduling revisited. IEEE Transactions on Computers, 59(2):188–201, February 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5282495.

Malrait:2011:ECS

[MBM11] Luc Malrait, Sara Bouchenak, and Nicolas Marchand. Experience with CONSER: a system for server control through fluid modeling. IEEE Transactions on Computers, 60(7):951– 963, July 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Moyer:2012:SWC

[MBS^{+}12] Thomas Moyer, Kevin Butler, Joshua Schiffman, Patrick McDaniel, and Trent Jaeger. Scalable Web content attestation. IEEE Transactions on Computers, 61(5):686– 699, May 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mireshghallah:2019:EEP

[MBSSA19] F. Mireshghallah, M. Bakhshalipour, M. Sadrosadati, and H. Sarbazi-Azad. Energy-efficient permanent fault tolerance in hard real-time systems. IEEE Transactions on Computers, 68(10):1539–1545, October 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mukhopadhyay:2011:PEA

[MC11]  Debdeep  Mukhopadhyay

and Dipanwita Roy Chowdhury. A parallel efficient architecture for large cryptographically robust n × k(k ≥ n/2) mappings. IEEE Transactions on Computers, 60 (3):375–385, March 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Malandrino:2012:CDC

[MCC12] Francesco Malandrino, Claudio Casetti, and Carla-Fabiana Chiasserini. Content discovery and caching in mobile networks with infrastructure. IEEE Transactions on Computers, 61(10): 1507–1520, October 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Maddah:2016:SST

[MCM16] Rakan Maddah, Sangyeun Cho, and Rami Melhem. Symbol shifting: Tolerating more faults in PCM blocks. IEEE Transactions on Computers, 65(7):2270– 2283, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Munoz-Coreas:2019:QCD

[MCT19] E. Muñoz-Coreas and H. Thapliyal. Quantum circuit design of a T -count optimized integer multiplier. IEEE Transactions on Computers, 68(5):729–739, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mo:2018:PAL

[MCXZ18] Yuchang Mo, Lirong Cui, Liudong Xing, and Zhao Zhang. Performability analysis of large-scale multi-state computing systems. IEEE Transactions on Computers, 67(1):59–72, January 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7981361/.

Michelogiannakis:2013:EBF

[MD13] George Michelogiannakis and William J. Dally. Elastic buffer flow control for onchip networks. IEEE Transactions on Computers, 62 (2):295–309, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Manoochehri:2016:AMA

[MD16] Mehrtash Manoochehri and Michel Dubois. Accurate model for application failure due to transient faults in caches. IEEE Transactions on Computers, 65 (8):2397–2410, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Martins:2017:AIR

[MEBS17] Paulo Martins, Julien Eynard, Jean-Claude Bajard, and Leonel Sousa. Arithmetical improvement of the round-off for cryptosystems in high-dimensional lattices. IEEE Transactions on Computers, 66(12):

2005–2018, December 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7891511/.

Muck:2014:TUD

[MF14] T. R. Muck and A. A. Frohlich. Toward unified design of hardware and software components using C++. IEEE Transactions on Computers, 63(11): 2880–2893, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mohammadi:2014:TMA

[MFG14] Anwar Mohammadi, Nathan Fisher, and Daniel Grosu. Truthful mechanisms for allocating a single processor to sporadic tasks in competitive real-time environments. IEEE Transactions on Computers, 63 (8):2066–2079, August 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mashayekhy:2016:TMC

[MFG16] Lena Mashayekhy, Nathan Fisher, and Daniel Grosu. Truthful mechanisms for competitive reward-based scheduling. IEEE Transactions on Computers, 65 (7):2299–2312, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mao:2017:RCP

[MFT^{+}17] Bomin Mao, Zubair Md.

Fadlullah, Fengxiao Tang, Nei Kato, Osamu Akashi, Takeru Inoue, and Kimihiro Mizutani. Routing or computing? The paradigm shift towards intelligent computer network packet transmission based on deep learning. IEEE Transactions on Computers, 66(11):1946– 1960, November 2017. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7935536/.

Metra:2011:GEI

[MG11a] Cecilia Metra and Rajesh Galivanche. Guest Editors’ introduction: Special section on concurrent on-line testing and error/fault resilience of digital systems. IEEE Transactions on Computers, 60(9): 1217–1218, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5962402.

Moller:2011:IDI

[MG11b] Niels Möller and Torbjörn Granlund. Improved division by invariant integers. IEEE Transactions on Computers, 60(2):165–175, February 2011. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mandal:2016:FRI

[MG16] Kalikinkar Mandal and Guang Gong. Feedback reconstruction and implementations of pseudorandom number generators from composited De Bruijn sequences. IEEE Transactions on Computers, 65(9): 2725–2738, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Maene:2018:HBT

[MGdC^{+}18] Pieter Maene, Johannes Götzfried, Ruan de Clercq, Tilo Müller, Felix Freiling, and Ingrid Verbauwhede. Hardware-based trusted computing architectures for isolation and attestation. IEEE Transactions on Computers, 67(3):361–374, March 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7807249/.

Manole:2014:PSP

[MGW14] Sagi Manole, Amit Golander, and Shlomo Weiss. Protein sequence pattern matching: Leveraging application specific hardware accelerators. IEEE Transactions on Computers, 63 (2):448–460, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Meloni:2015:EDB

[MH15] N. Meloni and M. A. Hasan.

Efficient double bases for scalar multiplication. IEEE Transactions on Computers, 64(8):2204–2212, August 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Musavi:2019:OBM

[MH19] S. A. Musavi and M. R.

Hashemi. An ontology-based method for HW/SW architecture reconstruction. IEEE Transactions on Computers, 68(7):1007–1018, July 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mazahir:2017:PEM

[MHH^{+}17] Sana Mazahir, Osman Hasan, Rehan Hafiz, Muhammad Shafique, and Jörg Henkel. Probabilistic error modeling for approximate adders. IEEE Transactions on Computers, 66(3):515–530, March 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mazahir:2017:PEA

[MHHS17] Sana Mazahir, Osman Hasan, Rehan Hafiz, and Muhammad Shafique. Probabilistic error analysis of approximate recursive multipliers. IEEE Transactions on Computers, 66(11):1982–1990, November 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7935435/.

Mahmood:2015:ECR

[MHK15] T. Mahmood, Seokin Hong, and Soontae Kim. Ensuring cache reliability and energy scaling at near-threshold voltage with Macho. IEEE Transactions on Computers, 64(6):1694–1706, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Momeni:2015:DAA

[MHML15] A. Momeni, Jie Han, P. Montuschi, and F. Lombardi. Design and analysis of approximate compressors for multiplication. IEEE Transactions on Computers, 64 (4):984–994, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Matadamas-Hernandez:2014:PSP

[MHRARG^{+}14] J. Matadamas-Hernandez, G. Roman-Alonso, F. Rojas-Gonzalez, M. A. Castro-Garcia, Azzedine Boukerche, M. Aguilar-Cornejo, and S. Cordero-Sanchez. Parallel simulation of pore networks using multicore CPUs. IEEE Transactions on Computers, 63(6):1513–1525, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mokhov:2014:SPI

[MIS^{+}14] Andrey Mokhov, Alexei Iliasov, Danil Sokolov, Maxim Rykunov, Alex Yakovlev, and Alexander Romanovsky. Synthesis of processor instruction sets from high-level ISA specifications. IEEE Transactions on Computers, 63(6):1552–1566, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ma:2014:HRA

[MJW^{+}14] Sheng Ma, Natalie Enright Jerger, Zhiying Wang, Mingche Lai, and Libo Huang. Holistic routing algorithm design to support workload consolidation in NoCs. IEEE Transactions on Computers, 63 (3):529–542, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mao:2016:LDD

[MJWT16] Bo Mao, Hong Jiang, Suzhen Wu, and Lei Tian. Leveraging data deduplication to improve the performance of primary storage systems in the cloud. IEEE Transactions on Computers, 65 (6):1775–1788, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Matsutani:2011:PRL

[MKAY11] Hiroki Matsutani, Michihiro Koibuchi, Hideharu Amano, and Tsutomu Yoshinaga. Prediction router: a lowlatency on-chip router architecture with multiple predictors. IEEE Transactions on Computers, 60(6):783– 799, June 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Maniatakos:2013:LCC

[MKFM13] Michail Maniatakos, Prabhakar Kudva, Bruce M. Fleischer, and Yiorgos Makris. Low-cost concurrent error detection for floatingpoint unit (FPU) controllers. IEEE Transactions on Computers, 62(7):1376– 1388, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Morris:2014:TDS

[MKLW14] Randy W. Morris, Avinash Karanth Kodi, Ahmed Louri, and Ralph D. Whaley. Threedimensional stacked nanophotonic network-on-chip architecture with minimal reconfiguration. IEEE Transactions on Computers, 63 (1):243–255, January 2014.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mohaqeqi:2014:ALA

[MKM14] Morteza Mohaqeqi, Mehdi Kargahi, and Ali Movaghar. Analytical leakage-aware thermal modeling of a realtime system. IEEE Transactions on Computers, 63 (6):1378–1392, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mozaffari-Kermani:2010:CSI

[MKRM10] M. Mozaffari-Kermani and A. Reyhani-Masoleh. Concurrent structure-independent fault detection schemes for the Advanced Encryption Standard. IEEE Transactions on Computers, 59(5): 608–622, May 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5406504.

Mozaffari-Kermani:2011:LPH

[MKRM11] Mehran Mozaffari-Kermani and Arash Reyhani-Masoleh. A low-power high-performance concurrent fault detection approach for the composite field S-box and inverse S-box. IEEE Transactions on Computers, 60(9): 1327–1340, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5962403.

Mozaffari-Kermani:2012:EHP

[MKRM12] Mehran Mozaffari-Kermani and Arash Reyhani-Masoleh. Efficient and high-performance parallel hardware architectures for the AES-GCM. IEEE Transactions on Computers, 61(8):1165–1178, August 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Maniatakos:2011:ILI

[MKT^{+}11] Michail Maniatakos, Naghmeh Karimi, Chandra Tirumurti, Abhijit Jas, and Yiorgos Makris. Instruction-level impact analysis of low-level faults in a modern microprocessor controller. IEEE Transactions on Computers, 60(9):1260–1273, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5432157.

Meng:2013:EMS

[ML13] Shicong Meng and Ling Liu. Enhanced monitoring-as-a-service for effective cloud management. IEEE Transactions on Computers, 62(9): 1705–1720, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mun:2016:NAE

[ML16] J. H. Mun and H. Lim.

New approach for efficient IP address lookup using a Bloom filter in trie-based algorithms. IEEE Transactions on Computers, 65 (5):1558–1565, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Manumachu:2018:BOO

[ML18] Ravindranath Reddy Manumachu and Alexey Lastovetsky. Bi-objective optimization of data-parallel applications on homogeneous multicore clusters for performance and energy. IEEE Transactions on Computers, 67 (2):160–177, February 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8013836/.

Mancillas-Lopez:2010:RHI

[MLCH10] C. Mancillas-Lopez, D. Chakraborty, and F. Rodriguez Henriquez. Reconfigurable hardware implementations of tweakable enciphering schemes. IEEE Transactions on Computers, 59(11):1547–1561, November 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.

                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5432161.

Min:2014:DIL

[MLE14] Changwoo Min, Sang-Won Lee, and Young Ik Eom. Design and implementation of a log-structured file system for flash-based solid state drives. IEEE Transactions on Computers, 63(9): 2215–2227, September 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Moreno:2012:RMM

[MLH12] Julio Villalba Moreno, Tomas Lang, and Javier Hormigo. Radix-2 multioperand and multiformat streaming online addition. IEEE Transactions on Computers, 61 (6):790–803, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mei:2015:PMS

[MLOL15] J. Mei, K. Li, A. Ouyang, and K. Li. A profit maximization scheme with guaranteed quality of service in cloud computing. IEEE Transactions on Computers, 64(11):3064–3078, November 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Maragos:2019:FMP

[MLS19] K. Maragos, G. Lentaris, and D. Soudris. In-the-field mitigation of process variability for improved FPGA performance. IEEE Transactions on Computers, 68 (7):1049–1063, July 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ma:2012:HPB

[MLW12] Kun Ma, Han Liang, and Kaijie Wu. Homomorphic property-based concurrent error detection of RSA: a countermeasure to fault attack. IEEE Transactions on Computers, 61(7):1040– 1049, July 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mali:2016:TTB

[MM16] Goutam Mali and Sudip Misra. TRAST: Trustbased distributed topology management for wireless multimedia sensor networks. IEEE Transactions on Computers, 65(6):1978– 1991, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mazumdar:2017:CRS

[MM17] Bodhisatwa Mazumdar and Debdeep Mukhopadhyay. Construction of rotation symmetric S-boxes with high nonlinearity and improved DPA resistivity. IEEE Transactions on Computers, 66(1):59–72, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Milutinovic:2019:IRS

[MMAC19] S. Milutinovic, E. Mezzetti, J. Abella, and F. J. Cazorla. Increasing the reliability of software timing analysis for cache-based processors. IEEE Transactions on Computers, 68(6):836– 851, June 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mostefaoui:2014:LRA

[MMB14] A. Mostefaoui, M. Melkemi, and A. Boukerche. Localized routing approach to bypass holes in wireless sensor networks. IEEE Transactions on Computers, 63(12): 3053–3065, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Maddah:2015:RTM

[MMC15] R. Maddah, R. Melhem, and Sangyeun Cho. RDIS: Tolerating many stuck-at faults in resistive memory. IEEE Transactions on Computers, 64(3):847–861, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Montuschi:2016:SJb

[MMC^{+}16] Paolo Montuschi, Edward J.

McCluskey, Samarjit Chakraborty, Jason Cong, Ramón M. Rodríguez-Dagnino, Fred Douglis, Lieven Eeckhout, Gernot Heiser, Sushil Jajodia, Ruby B. Lee, Dinesh Manocha, Tomás F. Pena, Isabelle Puaut, Hanan Samet, and Donatella Sciuto. State of the journal. IEEE Transactions on Computers, 65 (7):2014–2018, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Manna:2018:TAA

[MMCS18] Kanchan Manna, Priyajit Mukherjee, Santanu Chattopadhyay, and Indranil Sengupta. Thermal-aware application mapping strategy for network-on-chip based system design. IEEE Transactions on Computers, 67 (4):528–542, April 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8096996/.

Misra:2014:AAC

[MMH14] S. Misra, N. E. Majd, and Hong Huang. Approximation algorithms for constrained relay node placement in energy harvesting wireless sensor networks. IEEE Transactions on Computers, 63(12):2933–2947, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Moradi:2013:OAR

[MMP13] Amir Moradi, Oliver Mischke, and Christof Paar. One attack to rule them all: Collision timing attack versus 42 AES ASIC cores. IEEE Transactions on Computers, 62(9):1786–1798, September 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Maniatakos:2015:RVA

[MMTM15] M. Maniatakos, M. Michael, C. Tirumurti, and Y. Makris. Revisiting vulnerability analysis in modern microprocessors. IEEE Transactions on Computers, 64(9): 2664–2674, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Macdonald:2014:PMF

[MNFA14] Kevin Macdonald, Christopher Nitta, Matthew Farrens, and Venkatesh Akella. PDG GEN: A methodology for fast and accurate simulation of on-chip networks. IEEE Transactions on Computers, 63(3):650– 663, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mashayekhy:2016:OMR

[MNGV16] L. Mashayekhy, M. M. Nejad, D. Grosu, and A. V. Vasilakos. An online mechanism for resource allocation and pricing in clouds. IEEE Transactions on Computers, 65(4):1172–1184, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mladenov:2011:IER

[MNK11] Todor Mladenov, Saeid Nooshabadi, and Keseon Kim. Implementation and evaluation of Raptor codes on embedded systems. IEEE Transactions on Computers, 60(12):1678–1691, December 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5601702.

Metra:2012:NDT

[MOMT12] Cecilia Metra, Martin Omana, T. M. Mak, and Simon Tam. New design for testability approach for clock fault testing. IEEE Transactions on Computers, 61(4):448– 457, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Montuschi:2015:ENE

[Mon15a]  P. Montuschi.  Editorial

from the new Editor in Chief. IEEE Transactions on Computers, 64(1):2, January 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Montuschi:2015:SJ

[Mon15b]  P. Montuschi.  State of

the journal. IEEE Transactions on Computers, 64 (6):1506–1508, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Montuschi:2016:SJa

[Mon16] P. Montuschi. State of the journal. IEEE Transactions on Computers, 65(1): 1–4, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Montuschi:2017:SJ

[Mon17] Paolo Montuschi. State of the journal. IEEE Transactions on Computers, 66(1): 1–2, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Montuschi:2018:SJ

[Mon18] Paolo Montuschi. State of the journal. IEEE Transactions on Computers, 67(1): 1, January 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Montuschi:2019:TYS

[Mon19] Paolo Montuschi. Thank-you state of the journal editorial by the outgoing Editorin-Chief. IEEE Transactions on Computers, 68(1): 1–2, January 2019. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8573061/.

Milano:2014:GEI

[MOS14] Michela Milano, Barry O’Sullivan, and Martin Sachenbacher. Guest Editors’ introduction: Special section on computational sustainability: Where computer science meets sustainable development. IEEE Transactions on Computers, 63(1):88–89, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mastroleon:2012:PDA

[MOYB12] Lykomidis Mastroleon, Daniel C.

O’Neill, Benjamin Yolken, and Nicholas Bambos. Power and delay aware management of packet switches. IEEE Transactions on Computers, 61(12):1789–1799, December 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Matula:2015:MDE

[MPZ15] D. W. Matula, M. T. Panu, and J. Y. Zhang. Multiplicative division employing independent factors. IEEE Transactions on Computers, 64(7):2012–2019, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Magron:2019:CRE

[MRD19] V. Magron, A. Rocca, and T. Dang. Certified roundoff error bounds using Bernstein expansions and sparse Krivine–Stengle representations. IEEE Transactions on Computers, 68(7):953– 966, July 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Migliore:2018:HSC

[MRL^{+}18] Vincent Migliore, Maria Méndez Real, Vianney Lapotre, Arnaud Tisserand, Caroline Fontaine, and Guy Gogniat. Hardware/software codesign of an accelerator for FV homomorphic encryption scheme using Karatsuba algorithm. IEEE Transactions on Computers, 67 (3):335–347, March 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7797469/.

Menzel:2015:CHD

[MRW^{+}15] M. Menzel, R. Ranjan, Lizhe Wang, S. U. Khan, and Jinjun Chen. CloudGenius: A hybrid decision support method for automating the migration of Web application clusters to public clouds. IEEE Transactions on Computers, 64 (5):1336–1348, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mishra:2012:PPE

[MS12] Tania Banerjee Mishra and Sartaj Sahni. PETCAM — a power efficient TCAM architecture for forwarding tables. IEEE Transactions on Computers, 61(1):3–17, January 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

McCartney:2015:SMT

[MS15] W. P. McCartney and N. Sridhar. Stackless multithreading for embedded systems. IEEE Transactions on Computers, 64(10):2940– 2952, October 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Martinez:2012:DTR

[MSC12] Carlos Alvarez Martinez, Jesus Corbal San Adrian, and Mateo Valero Cortes. Dynamic tolerance region computing for multimedia. IEEE Transactions on Computers, 61(5):650–665, May 2012.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ma:2014:DBV

[MSG14] Zhiqiang Ma, Zhonghua Sheng, and Lin Gu. DVM: A big virtual machine for cloud computing. IEEE Transactions on Computers, 63(9): 2245–2258, September 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mardiris:2015:ADA

[MSK15] V. A. Mardiris, G. C. Sirakoulis, and I. G. Karafyllidis. Automated design architecture for 1-D cellular automata using quantum cellular automata. IEEE Transactions on Computers, 64(9): 2476–2489, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mosenia:2017:CCA

[MSKRJ17] A. Mosenia, S. Sur-Kolay, A. Raghunathan, and N. K. Jha. CABA: Continuous authentication based on BioAura. IEEE Transactions on Computers, 66(5): 759–772, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Majumder:2012:NBH

[MSPK12] Turbo Majumder, Souradip Sarkar, Partha Pratim Pande, and Ananth Kalyanaraman. NoC-based hardware accelerator for breakpoint phylogeny. IEEE Transactions on Computers, 61(6):857– 869, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Maitra:2017:DFA

[MSS17] Subhamoy Maitra, Akhilesh Siddhanti, and Santanu Sarkar. A differential fault attack on Plantlet. IEEE Transactions on Computers, 66(10):1804–1808, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7917296/.

Maitra:2018:TAA

[MSS^{+}18] Subhamoy Maitra, Nishant Sinha, Akhilesh Siddhanti, Ravi Anand, and Sugata Gangopadhyay. A TMDTO attack against Lizard. IEEE Transactions on Computers, 67(5):733–739, May 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8107499/.

Moser:2010:APM

[MTBB10] C. Moser, L. Thiele, D. Brunelli, and L. Benini. Adaptive power management for environmentally powered systems. IEEE Transactions on Computers, 59(4):478– 491, April 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5291691.

Mao:2019:AMC

[MTFK19] B. Mao, F. Tang, Z. M. Fadlullah, and N. Kato. An absorbing Markov chain based model to solve computation and communication tradeoff in GPU-accelerated MDRUs for safety confirmation in disaster scenarios. IEEE Transactions on Computers, 68(9):1256–1268, September 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Maniatakos:2012:GSV

[MTGM12] Michail Maniatakos, Chandrasekharan (Chandra) Tirumurti, Rajesh Galivanche, and Yiorgos Makris. Global signal vulnerability (GSV) analysis for selective state element hardening in modern microprocessors. IEEE Transactions on Computers, 61(10):1361–1370, October 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Milenkovic:2011:CPR

[MUMB11] Aleksandar Milenković, Vladimir Uzelac, Milena Milenković, and Martin Burtscher. Caches and predictors for realtime, unobtrusive, and costeffective program tracing in embedded systems. IEEE Transactions on Computers, 60(7):992–1005, July 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mangassarian:2010:RQE

[MVB10] H. Mangassarian, A. Veneris, and M. Benedetti. Robust QBF encodings for sequential circuits with applications to verification, debug, and test. IEEE Transactions on Computers, 59(7):981–994, July 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5441288.

Mao:2010:HSS

[MW10] Shufu Mao and T. Wolf.

Hardware support for secure processing in embedded systems. IEEE Transactions on Computers, 59 (6):847–854, June 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5406503.

Muller:2013:NXN

[MW13] Dirk Muller and Matthias Werner. A note on “New Strategies for Assigning Real-Time Tasks to Multiprocessor Systems”. IEEE Transactions on Computers, 62(9):1904–1905, September 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ma:2015:LOS

[MWLJ15] Sheng Ma, Zhiying Wang, Zonglin Liu, and N. E. Jerger. Leaving one slot empty: Flit bubble flow control for torus cache-coherent NoCs. IEEE Transactions on Computers, 64(3):763– 777, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ma:2014:DDP

[MWW14] Kai Ma, Xiaorui Wang, and Yefu Wang. DPPC: Dynamic power partitioning and control for improved chip multiprocessor performance. IEEE Transactions on Computers, 63 (7):1736–1750, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Maneesilp:2013:RSA

[MWWT13] Jullawadee Maneesilp, Chong Wang, Hongyi Wu, and Nian-Feng Tzeng. RFID support for accurate 3D localization. IEEE Transactions on Computers, 62 (7):1447–1459, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Meng:2016:SRA

[MWY^{+}16] T. Meng, F. Wu, Z. Yang, G. Chen, and A. V. Vasilakos. Spatial reusabilityaware routing in multi-hop wireless networks. IEEE Transactions on Computers, 65(1):244–255, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mao:2017:EEG

[MWZ^{+}17] Mengjie Mao, Wujie Wen, Yaojun Zhang, Yiran Chen, and Hai Li. An energyefficient GPGPU register file architecture using racetrack memory. IEEE Transactions on Computers, 66(9): 1478–1490, September 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7891951/.

Mokhov:2010:CPO

[MY10] A. Mokhov and A. Yakovlev.

Conditional partial order graphs: Model, synthesis, and application. IEEE Transactions on Computers, 59(11):1480–1493, November 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.

                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5432155.

Ma:2016:AFE

[MYHL16] J. Ma, G. Yan, Y. Han, and X. Li. An analytical framework for estimating scale-out and scale-up power efficiency of heterogeneous manycores. IEEE Transactions on Computers, 65 (2):367–381, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Min:2011:EDT

[MYW11] Jaehong Min, Daeyoung Yoon, and Youjip Won. Efficient deduplication techniques for modern backup operation. IEEE Transactions on Computers, 60 (6):824–840, June 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nannarelli:2016:PPS

[Nan16] Alberto Nannarelli. Performance/power space exploration for Binary64 division units. IEEE Transactions on Computers, 65 (5):1671–1677, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nannarelli:2019:TFP

[Nan19]  A. Nannarelli.  Tunable

floating-point adder. IEEE Transactions on Computers, 68(10):1553–1560, October 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nain:2017:SPE

[NBZP17] Ajay Kumar Nain, Jagadish Bandaru, Mohammed Abdullah Zubair, and Rajalakshmi Pachamuthu. A secure phase-encrypted IEEE 802.15.4 transceiver design. IEEE Transactions on Computers, 66(8):1421– 1427, August 2017. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7862216/.

Nisar:2011:GPC

[NC11] Muhammad M. Nisar and Abhijit Chatterjee. Guided probabilistic checksums for error control in low-power digital filters. IEEE Transactions on Computers, 60(9): 1313–1326, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5674023.

Napoli:2017:SRC

[NCD^{+}17] Ettore Napoli, Gerardo Castellano, Davide De Caro, Darjn Esposito, Nicola Petra, and Antonio G. M.

Strollo. A SISO register circuit tailored for input data with low transition probability. IEEE Transactions on Computers, 66(1):45–51, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Narasimhan:2013:HTD

[NDC^{+}13] Seetharam Narasimhan, Dongdong Du, Rajat Subhra Chakraborty, Somnath Paul, Francis G. Wolff, Christos A. Papachristou, Kaushik Roy, and Swarup Bhunia. Hardware Trojan detection by multiple-parameter sidechannel analysis. IEEE Transactions on Computers, 62(11):2183–2195, November 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ngo:2017:CSS

[NDG^{+}17] Xuan Thuy Ngo, Jean-Luc Danger, Sylvain Guilley, Tarik Graba, Yves Mathieu, Zakaria Najm, and Shivam Bhasin. Cryptographically secure shield for security IPs protection. IEEE Transactions on Computers, 66 (2):354–360, February 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Naithani:2018:OSE

[NEE18] Ajeya Naithani, Stijn Eyerman, and Lieven Eeckhout. Optimizing soft error reliability through scheduling on heterogeneous multicore processors. IEEE Transactions on Computers, 67(6):830–846, June 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8126846/.

Niyato:2010:MMH

[NH10] D. Niyato and E. Hossain.

A microeconomic model for hierarchical bandwidth sharing in dynamic spectrum access networks. IEEE Transactions on Computers, 59(7):865–877, July 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5453338.

Namba:2011:TSR

[NI11] Kazuteru Namba and Hideo Ito. Test sets for robust path delay fault testing on two-rail logic circuits. IEEE Transactions on Computers, 60(10):1459–1470, October 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5661767.

Nam:2011:OOO

[NKEM11] Eyee Hyun Nam, Bryan Suk Joon Kim, Hyeonsang Eom, and Sang Lyul Min. Ozone (O3): An out-of-order flash memory controller architecture. IEEE Transactions on Computers, 60(5):653–666, May 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nicolici:2014:NAa

[NL14] Nicola Nicolici and Zahra Lak. A novel algorithmic approach to aid postsilicon delay measurement and clock tuning. IEEE Transactions on Computers, 63(5):1074–1084, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Namba:2015:NBO

[NL15a] K. Namba and F. Lombardi. Non-binary orthogonal latin square codes for a multilevel phase charge memory (PCM). IEEE Transactions on Computers, 64(7):2092–2097, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Namba:2015:PDT

[NL15b] K. Namba and F. Lombardi. Parallel decodable two-level unequal burst error correcting codes. IEEE Transactions on Computers, 64(10): 2902–2911, October 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Namba:2016:PDM

[NL16a] Kazuteru Namba and Fabrizio Lombardi. Parallel decodable multi-level unequal burst error correcting codes for memories of approximate systems. IEEE Transactions on Computers, 65(12): 3794–3801, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Namba:2016:SMS

[NL16b] Kazuteru Namba and Fabrizio Lombardi. Single multiscale-symbol error correction codes for multiscale storage systems. IEEE Transactions on Computers, 65(6):2005–2009, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nikolov:2016:OLC

[NL16c] D. Nikolov and E. Larsson. Optimizing the level of confidence for multiple jobs. IEEE Transactions on Computers, 65(4):1239– 1252, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Namba:2018:SAE

[NL18] Kazuteru Namba and Fabrizio Lombardi. A single and adjacent error correction code for fast decoding of critical bits. IEEE Transactions on Computers, 67(10): 1525–1531, October 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8329222/.

Namba:2019:CWL

[NL19] Kazuteru Namba and Fabrizio Lombardi. Coding for write latency reduction in a Multi-Level Cell (MLC) Phase Change Memory (PCM). IEEE Transactions on Computers, 68 (2):301–306, February 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8454732/.

Nam:2014:GEE

[NLP^{+}14] Min-Young Nam, Jaemyoun Lee, Kyung-Joon Park, Lui Sha, and Kyungtae Kang. Guaranteeing the end-to-end latency of an IMA system with an increasing workload. IEEE Transactions on Computers, 63(6):1460– 1473, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Najafi:2017:PCE

[NLRB17] M. Hassan Najafi, David J. Lilja, Marc D. Riedel, and Kia Bazargan. Polysynchronous clocking: Exploiting the skew tolerance of stochastic circuits. IEEE Transactions on Computers, 66(10):1734–1746, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7911306/.

Neophytou:2010:TSG

[NM10] S. N. Neophytou and M. K.

Michael. Test set generation with a large number of unspecified bits using static and dynamic techniques. IEEE Transactions on Computers, 59(3):301–316, March 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5342413.

Negre:2015:NPA

[NR15] C. Negre and J.-M. Robert.

New parallel approaches for scalar multiplication in elliptic curve over fields of small characteristic. IEEE Transactions on Computers, 64(10):2875–2890, October 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nguyen:2015:FRT

[NRG15] Thi Huyen Chau Nguyen, P. Richard, and E. Grolleau. An FPTAS for response time analysis of fixed priority real-time tasks with resource augmentation. IEEE Transactions on Computers, 64(7):1805–1818, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nopper:2013:SMC

[NS13] Tobias Nopper and Christoph Scholl. Symbolic model checking for incomplete designs with flexible modeling of unknowns. IEEE Transactions on Computers, 62 (6):1234–1254, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nannarelli:2014:GEI

[NST14] Alberto Nannarelli, Peter-Michael Seidel, and Ping Tak Peter Tang. Guest Editors’ introduction: Special section on computer arithmetic. IEEE Transactions on Computers, 63 (8):1852–1853, August 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Neto:2014:PUP

[NTR14] Joao Carlos Neto, Alexandre Ferreira Tenca, and Wilson Vicente Ruggiero. A parallel and uniform k-partition method for Montgomery multiplication. IEEE Transactions on Computers, 63(9): 2122–2133, September 2014.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nieto:2016:PRS

[NVB16] Alejandro Nieto, David L.

Vilariño, and Victor M. Brea. PRECISION: A reconfigurable SIMD/MIMD coprocessor for computer vision systems-on-chip. IEEE Transactions on Computers, 65(8):2548–2561, August 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Namin:2011:WLF

[NWA11] Ashkan Hosseinzadeh Namin, Huapeng Wu, and Majid Ahmadi. A word-level finite field multiplier using normal basis. IEEE Transactions on Computers, 60(6):890– 895, June 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Namin:2012:HSA

[NWA12] Ashkan Hosseinzadeh Namin, Huapeng Wu, and Majid Ahmadi. High-speed architectures for multiplication using reordered normal basis. IEEE Transactions on Computers, 61 (2):164–172, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nunez-Yanez:2015:AVS

[NY15] J. L. Nunez-Yanez. Adaptive voltage scaling with insitu detectors in commercial FPGAs. IEEE Transactions on Computers, 64 (1):45–53, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nunez-Yanez:2019:EPN

[NY19] J. Nunez-Yanez. Energy proportional neural network inference with adaptive voltage and frequency scaling. IEEE Transactions on Computers, 68(5):676–687, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nunez-Yanez:2016:EOC

[NYHB16] J. Luis Nunez-Yanez, M. Hosseinabady, and A. Beldachi. Energy optimization in commercial FPGAs with voltage, frequency and logic scaling. IEEE Transactions on Computers, 65(5):1484– 1493, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Neishaburi:2014:NMT

[NZ14] M. H. Neishaburi and Zeljko Zilic. On a new mechanism of trigger generation for post-silicon debugging. IEEE Transactions on Computers, 63(9):

2330–2342, September 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nechma:2015:PSM

[NZ15] T. Nechma and M. Zwolinski. Parallel sparse matrix solution for circuit simulation on FPGAs. IEEE Transactions on Computers, 64(4):1090–1103, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nong:2011:TEA

[NZC11] Ge Nong, Sen Zhang, and Wai Hong Chan. Two efficient algorithms for linear time suffix array construction. IEEE Transactions on Computers, 60(10): 1471–1484, October 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5582081.

Nie:2014:LCB

[NZLK14] Weiran Nie, Sen Zhou, Kwei-Jay Lin, and Soo Dong Kim. An on-line capacity-based admission control for realtime service processes. IEEE Transactions on Computers, 63(9):2134–2145, September 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Oh:2017:DBE

[OCK17] Hyunggoy Oh, Inhyuk Choi, and Sungho Kang. DRAMbased error detection method to reduce the post-silicon debug time for multiple identical cores. IEEE Transactions on Computers, 66(9): 1504–1517, September 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7872459/.

Ozturk:2017:CAH

[ÖDSS17] Erdinç Öztürk, Yarkın Doröz, Erkay Savaş, and Berk Sunar. A custom accelerator for homomorphic encryption applications. IEEE Transactions on Computers, 66(1):3–16, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

On:2014:FBC

[OGH^{+}14] Sai Tung On, Shen Gao, Bingsheng He, Ming Wu, Qiong Luo, and Jianliang Xu. FD-Buffer: A cost-based adaptive buffer replacement algorithm for flash memory devices. IEEE Transactions on Computers, 63(9): 2288–2301, September 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ozsoy:2014:ESS

[OGPK14] Meltem Ozsoy, Nael Abu Ghazaleh, Dmitry Ponomarev, and Mehmet Kayaalp. Efficiently securing systems from code reuse attacks. IEEE Transactions on Computers, 63(5):1144– 1156, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Oh:2017:CED

[OHCK17] Hyunggoy Oh, Taewoo Han, Inhyuk Choi, and Sungho Kang. An on-chip error detection method to reduce the post-silicon debug time. IEEE Transactions on Computers, 66 (1):38–44, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ozturk:2013:CDE

[OKC13] Ozcan Ozturk, Mahmut Kandemir, and Guangyu Chen. Compiler-directed energy reduction using dynamic voltage scaling and voltage islands for embedded systems. IEEE Transactions on Computers, 62 (2):268–278, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ozsoy:2016:HBM

[OKD^{+}16] Meltem Ozsoy, Khaled N.

Khasawneh, Caleb Donovick, Iakov Gorelik, Nael Abu-Ghazaleh, and Dmitry Ponomarev. Hardware-based malware detection using lowlevel architectural features. IEEE Transactions on Computers, 65(11):3332–3344, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Oh:2019:ACP

[OKY^{+}19] Yunho Oh, Keunsoo Kim, Myung Kuk Yoon, Jong Hyun Park, Yongjun Park, Murali Annavaram, and Won Woo Ro. Adaptive cooperation of prefetching and warp scheduling on GPUs. IEEE Transactions on Computers, 68(4):609–616, April 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8515055/.

Onizawa:2014:HTC

[OMFH14] Naoya Onizawa, Atsushi Matsumoto, Tomoyoshi Funazaki, and Takahiro Hanyu. High-throughput compact delay-insensitive asynchronous NoC router. IEEE Transactions on Computers, 63 (3):637–649, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ozturk:2017:CHA

[OOD^{+}17] Ozcan Ozturk, Umut Orhan, Wei Ding, Praveen Yedlapalli, and Mahmut Taylan Kandemir. Cache hierarchyaware query mapping on emerging multicore architectures. IEEE Transactions on Computers, 66(3):403– 415, March 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Otoom:2015:MCM

[OP15] M. Otoom and J. M.

Paul. Multiprocessor capacity metric and analysis. IEEE Transactions on Computers, 64(11):3181–3196, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ozsoy:2014:SLC

[OPAGS14] Meltem Ozsoy, Dmitry Ponomarev, Nael Abu-Ghazaleh, and Tameesh Suri. SIFT: Low-complexity energy-efficient information flow tracking on SMT processors. IEEE Transactions on Computers, 63(2): 484–496, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Omana:2017:NAP

[OPV^{+}17] Martin Omana, Marco Padovani, Kreshnik Veliu, Cecilia Metra, Juergen Alt, and Rajesh Galivanche. New approaches for power binning of high performance microprocessors. IEEE Transactions on Computers, 66 (7):1159–1171, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/tc/2017/07/07822988-
                                        abs.html.

Oshin:2015:EER

[OPZ15] T. O. Oshin, S. Poslad, and Zelun Zhang. Energyefficient real-time human mobility state classification using Smartphones. IEEE Transactions on Computers, 64(6):1680–1693, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Omana:2013:LCN

[ORBM13] Martin Omana, Daniele Rossi, Nicolo Bosio, and Cecilia Metra. Low cost NBTI degradation detection and masking approaches. IEEE Transactions on Computers, 62(3):496–509, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Omana:2010:HPR

[ORM10] M. Omaña, D. Rossi, and C. Metra. High-performance robust latches. IEEE Transactions on Computers, 59 (11):1455–1465, November 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5396333.

Oseledets:2011:ITK

[Ose11] Ivan Oseledets. Improved nterm Karatsuba-like formulas in GF(2). IEEE Transactions on Computers, 60 (8):1212–1216, August 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Or:2016:MBH

[OWP16] N. L. Or, X. Wang, and D. Pao. MEMORY-based hardware architectures to detect ClamAV virus signatures with restricted regular expression features. IEEE Transactions on Computers, 65(4):1225–1238, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Oh:2018:WSD

[OYP^{+}18] Yunho Oh, Myung Kuk Yoon, Jong Hyun Park, Yongjun Park, and Won Woo Ro. WASP: Selective data prefetching with monitoring runtime warp progress on GPUs. IEEE Transactions on Computers, 67(9): 1366–1373, September 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8309426/.

Paterna:2012:VAT

[PAC^{+}12] Francesco Paterna, Andrea Acquaviva, Alberto Caprara, Francesco Papariello, Giuseppe Desoli, and Luca Benini. Variabilityaware task allocation for energy-efficient quality of service provisioning in embedded streaming multimedia applications. IEEE Transactions on Computers, 61(7):939–953, July 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Panda:2016:SSP

[Pan16] Biswabandan Panda. SPAC: A synergistic prefetcher aggressiveness controller for multi-core systems. IEEE Transactions on Computers, 65(12):3740–3753, December 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Patel:2013:NGN

[PAP13] Kimish Patel, Murali Annavaram, and Massoud Pedram. NFRA: Generalized network flow-based resource allocation for hosting centers. IEEE Transactions on Computers, 62(9): 1772–1785, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Piso:2011:VLG

[PB11] Daniel Piso and Javier D.

Bruguera. Variable latency Goldschmidt algorithm based on a new rounding method and a remainder estimate. IEEE Transactions on Computers, 60(11): 1535–1546, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5669291.

Parikh:2016:RCD

[PB16] Ritesh Parikh and Valeria Bertacco. Resource conscious diagnosis and reconfiguration for NoC permanent faults. IEEE Transactions on Computers, 65 (7):2241–2256, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Poursafaei:2017:NNA

[PBE17] Farimah R. Poursafaei, Mostafa Bazzaz, and Alireza Ejlali. NPAM: NVMaware page allocation for multi-core embedded systems. IEEE Transactions on Computers, 66(10): 1703–1716, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7926310/.

Park:2016:BIL

[PBL16] Jaesung Park, Heejung Byun, and Jung-Ryun Lee. Bio-inspired load-balancing framework for loosely coupled heterogeneous server systems. IEEE Transactions on Computers, 65(11): 3280–3292, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pontarelli:2013:TAD

[PBT13] Salvatore Pontarelli, Giuseppe Bianchi, and Simone Teofili. Traffic-aware design of a high-speed FPGA network intrusion detection system. IEEE Transactions on Computers, 62(11): 2322–2334, November 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Panda:2011:CCS

[PBV11] Preeti Ranjan Panda, M. Balakrishnan, and Anant Vishnoi. Compressing cache state for postsilicon processor debug. IEEE Transactions on Computers, 60(4):484– 497, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pellizzoni:2010:IPP

[PC10] R. Pellizzoni and M. Caccamo. Impact of peripheralprocessor interference on WCET analysis of real-time embedded systems. IEEE Transactions on Computers, 59(3):400–415, March 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5282496.

Paul:2016:TSO

[PC16] G. Paul and A. Chattopadhyay. Three snakes in one hole: The first systematic hardware accelerator design for SOSEMANUK with optional serpent and SNOW 2.0 modes. IEEE Transactions on Computers, 65 (2):640–653, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Park:2014:CXP

[PCHS14] Sun-Mi Park, Ku-Young Chang, Dowon Hong, and Changho Seo. Comments on “On the Polynomial Multiplication in Chebyshev Form”. IEEE Transactions on Computers, 63(12): 3162–3163, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [ACO12].

Park:2016:CMS

[PCHS16] S. Park, K. Chang, D. Hong, and C. Seo. Comments on “Multiway Splitting Method for Toeplitz Matrix Vector Product”. IEEE Transactions on Computers, 65 (1):332–333, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [HN13].

Park:2017:NBR

[PCHS17] Sun-Mi Park, Ku-Young Chang, Dowon Hong, and Changho Seo. New block recombination for subquadratic space complexity polynomial multiplication based on overlap-free approach. IEEE Transactions on Computers, 66(8):1396–1406, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7857684/.

Park:2018:SSC

[PCHS18] Sun-Mi Park, Ku-Young Chang, Dowon Hong, and Changho Seo. Subquadratic space complexity multiplier using even type GNB based on efficient Toeplitz matrixvector product. IEEE Transactions on Computers, 67(12):1794–1805, December 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8359428/.

Park:2015:IIB

[PCLN15] H. Park, J. Choi, D. Lee, and S. H. Noh. iBuddy: Inverse buddy for enhancing memory allocation/deallocation performance on multi-core systems. IEEE Transactions on Computers, 64 (3):720–732, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Paul:2011:RDE

[PCZB11] Somnath Paul, Fang Cai, Xinmiao Zhang, and Swarup Bhunia. Reliability-driven ECC allocation for multiple bit error resilience in processor cache. IEEE Transactions on Computers, 60 (1):20–34, January 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pourshaghaghi:2013:FCV

[PdG13] Hamid Reza Pourshaghaghi and Jose Pineda de Gyvez. Fuzzy-controlled voltage scaling based on supply current tracking. IEEE Transactions on Computers, 62(12): 2397–2410, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Patranabis:2019:SCS

[PDJ^{+}19] S. Patranabis, N. Datta, D. Jap, J. Breier, S. Bhasin, and D. Mukhopadhyay. SCADFA: Combined SCA + DFA attacks on block ciphers with practical validations. IEEE Transactions on Computers, 68(10): 1498–1510, October 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pan:2013:UQE

[PDXZ13] Yangyang Pan, Guiqiang Dong, Ningde Xie, and Tong Zhang. Using quasi-EZ-NAND flash memory to build large-capacity solidstate drives in computing systems. IEEE Transactions on Computers, 62 (5):1051–1057, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pravadelli:2014:TQS

[PFGB14] Graziano Pravadelli, Franco Fummi, Valerio Guarnieri, and Nicola Bombieri. Testbench qualification of SystemC TLM protocols through mutation analysis. IEEE Transactions on Computers, 63(5):1248–1261, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pedram:2014:AAF

[PGvdG14] Ardavan Pedram, Andreas Gerstlauer, and Robert A. van de Geijn. Algorithm, architecture, and floatingpoint unit codesign of a matrix factorization accelerator. IEEE Transactions on Computers, 63(8): 1854–1867, August 2014.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pippenger:2011:FAS

[Pip11] Nicholas Pippenger. On-the-fly algorithms and sequential machines. IEEE Transactions on Computers, 60(9): 1372–1375, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5710885.

Pagani:2017:TSP

[PKC^{+}17] Santiago Pagani, Heba Khdr, Jian-Jia Chen, Muhammad Shafique, Minming Li, and Jörg Henkel. Thermal safe power (TSP): Efficient power budgeting for heterogeneous manycore systems in dark silicon. IEEE Transactions on Computers, 66(1):147– 162, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Park:2019:GGC

[PKY19] SeongJae Park, Minchan Kim, and Heon Y. Yeom. GCMA: Guaranteed contiguous memory allocator. IEEE Transactions on Computers, 68(3):390– 401, March 2019. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8456561/.

Palangappa:2016:WCS

[PLM16] P. M. Palangappa, J. Li, and K. Mohanram. WOMcode solutions for low latency and high endurance in phase change memory. IEEE Transactions on Computers, 65(4):1025–1040, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Papagianni:2013:OAV

[PLP^{+}13] Chrysa Papagianni, Aris Leivadeas, Symeon Papavassiliou, Vasilis Maglaris, Cristina Cervello-Pastor, and Alvaro Monje. On the optimal allocation of virtual resources in cloud computing networks. IEEE Transactions on Computers, 62 (6):1060–1071, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Peng:2014:TSD

[PLZW14] Wei Peng, Feng Li, Xukai Zou, and Jie Wu. A twostage deanonymization attack against anonymized social networks. IEEE Transactions on Computers, 63 (2):290–303, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pricopi:2014:TSA

[PM14] M. Pricopi and T. Mitra.

Task scheduling on adaptive multi-core. IEEE Transactions on Computers, 63(10): 2590–2603, October 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Possa:2014:MRF

[PMH^{+}14] P. R. Possa, S. A. Mahmoudi, N. Harb, C. Valderrama, and P. Manneback. A multi-resolution FPGAbased architecture for realtime edge and corner detection. IEEE Transactions on Computers, 63(10): 2376–2388, October 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Papavramidou:2016:TAE

[PN16] Panagiota Papavramidou and Michael Nicolaidis. Test algorithms for ECCbased memory repair in ultimate CMOS and post-CMOS. IEEE Transactions on Computers, 65(7):2284– 2298, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pattabiraman:2013:SSP

[PNKI13] Karthik Pattabiraman, Nithin M.

Nakka, Zbigniew T. Kalbarczyk, and Ravishankar K. Iyer. SymPLFIED: Symbolic program-level fault injection and error detection framework. IEEE Transactions on Computers, 62(11): 2292–2307, November 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pontarelli:2013:EDC

[PO13] Salvatore Pontarelli and Marco Ottavi. Error detection and correction in content addressable memories by using Bloom filters. IEEE Transactions on Computers, 62(6):1111–1126, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pomeranz:2012:CFT

[Pom12a] Irith Pomeranz. Concatenation of functional test subsequences for improved fault coverage and reduced test length. IEEE Transactions on Computers, 61(6):899– 904, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pomeranz:2012:FIU

[Pom12b] Irith Pomeranz. Fast identification of undetectable transition faults under functional broadside tests. IEEE Transactions on Computers, 61(6):905–910, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pomeranz:2012:CCT

[Pom12c] Irith Pomeranz. On the computation of common test data for broadside and skewed-load tests. IEEE Transactions on Computers, 61(4):578–583, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pomeranz:2012:SAS

[Pom12d] Irith Pomeranz. On the switching activity and static test compaction of multicycle scan-based tests. IEEE Transactions on Computers, 61(8):1179–1188, August 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pomeranz:2013:ASA

[Pom13a] Irith Pomeranz. An adjacent switching activity metric under functional broadside tests. IEEE Transactions on Computers, 62 (2):404–410, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pomeranz:2013:STP

[Pom13b]  Irith Pomeranz.  Signal-

transition patterns of functional broadside tests. IEEE Transactions on Computers, 62(12):2544–2549, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pomeranz:2014:SLB

[Pom14] Irith Pomeranz. Sharing logic for built-in generation of functional broadside tests. IEEE Transactions on Computers, 63 (4):1048–1054, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pomeranz:2015:PFB

[Pom15a] I. Pomeranz. Piecewisefunctional broadside tests based on reachable states. IEEE Transactions on Computers, 64(8):2415–2420, August 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pomeranz:2015:TVO

[Pom15b] I. Pomeranz. Test vector omission for fault coverage improvement of functional test sequences. IEEE Transactions on Computers, 64(11):3317–3321, November 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pomeranz:2015:TDS

[Pom15c] I. Pomeranz. Two-dimensional static test compaction for functional test sequences. IEEE Transactions on Computers, 64(10):3009–3015, October 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pomeranz:2016:BGP

[Pom16a] Irith Pomeranz. LF SRbased generation of partiallyfunctional broadside tests. IEEE Transactions on Computers, 65(8):2659–2664, August 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pomeranz:2016:IAD

[Pom16b] Irith Pomeranz. Improving the accuracy of defect diagnosis with multiple sets of candidate faults. IEEE Transactions on Computers, 65(7):2332–2338, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pathak:2010:EET

[PP10] A. Pathak and V. K.

Prasanna. Energy-efficient task mapping for data-driven sensor network macroprogramming. IEEE Transactions on Computers, 59(7): 955–968, July 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5313800.

Park:2011:HPS

[PP11] Youngwoo Park and Kyu Ho Park. High-performance scalable flash file system using virtual metadata storage with phase-change RAM. IEEE Transactions on Computers, 60(3):321–334, March 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Park:2014:ETM

[PP14] Heemin Park and Moonju Park. An efficient test method for rate monotonic schedulability. IEEE Transactions on Computers, 63 (5):1309–1315, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Parfieniuk:2016:SIC

[PP16] Marek Parfieniuk and Sang Yoon Park. Sparse-iteration 4D CORDIC algorithms for multiplying quaternions. IEEE Transactions on Computers, 65(9):2859–2871, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Potharaju:2014:WSE

[PPB^{+}14] Rahul Potharaju, Dan Pei, Fabian E. Bustamante, Yan Chen, David R. Choffnes, Yao Zhao, and Kai Chen. Where the sidewalk ends: Extending the Internet AS graph using traceroutes from P2P users. IEEE Transactions on Computers, 63 (4):1021–1036, April 2014.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Park:2012:PAN

[PPKW12] Seung-Ho Park, Jung-Wook Park, Shin-Dug Kim, and Charles C. Weems. A pattern adaptive NAND flash memory storage structure. IEEE Transactions on Computers, 61(1):134–138, January 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Psarras:2017:DCM

[PPND17] Anastasios Psarras, Michalis Paschou, Chrysostomos Nicopoulos, and Giorgos Dimitrakopoulos. A dualclock multiple-queue shared buffer. IEEE Transactions on Computers, 66(10): 1809–1815, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7930511/.

Pham:2013:LOF

[PPP13] Hung-Manh Pham, Sebastien Pillement, and Stanisław J. Piestrak. Lowoverhead fault-tolerance technique for a dynamically reconfigurable softcore processor. IEEE Transactions on Computers, 62(6):1179– 1192, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pomeranz:2010:EDS

[PR10] I. Pomeranz and S. M. Reddy. Equivalence, dominance, and similarity relations between fault pairs and a fault pair collapsing process for fault diagnosis. IEEE Transactions on Computers, 59(2):150–158, February 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5184814.

Prasad:2014:MDA

[PR14] Abhinandan S. Prasad and Shrisha Rao. A mechanism design approach to resource procurement in cloud computing. IEEE Transactions on Computers, 63 (1):17–30, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pontarelli:2013:LCC

[PRBM13] Salvatore Pontarelli, Pedro Reviriego, Chris J. Bleakley, and Juan Antonio Maestro. Low complexity concurrent error detection for complex multiplication. IEEE Transactions on Computers, 62(9): 1899–1903, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Perez-Resa:2019:SSE

[PRGBSAC19] A. Pérez-Resa, M. Garcia-Bosque, C. Sánchez-Azqueta, and S. Celma. Selfsynchronized encryption for physical layer in 10Gbps optical links. IEEE Transactions on Computers, 68 (6):899–911, June 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pontarelli:2016:PDP

[PRM16] S. Pontarelli, P. Reviriego, and J. A. Maestro. Parallel d-pipeline: A cuckoo hashing implementation for increased throughput. IEEE Transactions on Computers, 65 (1):326–331, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Poudel:2019:MTU

[PRM19] Prawar Poudel, Biswajit Ray, and Aleksandar Milenkovic. Microcontroller TRNGs using perturbed states of NOR flash memory cells. IEEE Transactions on Computers, 68 (2):307–313, February 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8443106/.

Pontarelli:2015:LDS

[PROM15] S. Pontarelli, P. Reviriego, M. Ottavi, and J. A. Maestro. Low delay single symbol error correction codes based on Reed Solomon codes. IEEE Transactions on Computers, 64(5):1497– 1501, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pudi:2017:MLF

[PSL17] Vikramkumar Pudi, K. Sridharan, and Fabrizio Lombardi. Majority logic formulations for parallel adder designs at reduced delay and circuit complexity. IEEE Transactions on Computers, 66(10):1824–1830, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7909019/.

Patranabis:2017:PSK

[PSM17] S. Patranabis, Y. Shrivastava, and D. Mukhopadhyay. Provably secure key-aggregate cryptosystems with broadcast aggregate keys for online data sharing on the cloud. IEEE Transactions on Computers, 66(5):891–904, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Psarras:2016:SNC

[PSND16] Anastasios Psarras, Ioannis Seitanidis, Chrysostomos Nicopoulos, and Giorgos Dimitrakopoulos. Short-Path: A network-on-chip router with fine-grained pipeline bypassing. IEEE Transactions on Computers, 65(10):3136–3147, October 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pauwels:2012:CFG

[PTD^{+}12] Karl Pauwels, Matteo Tomasi, Javier Diaz, Eduardo Ros, and Marc M. Van Hulle. A comparison of FPGA and GPU for real-time phasebased optical flow, stereo, and local image features. IEEE Transactions on Computers, 61(7):999–1012, July 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pedram:2012:CTH

[PvdGG12] Ardavan Pedram, Robert A.

van de Geijn, and Andreas Gerstlauer. Codesign tradeoffs for high-performance, low-power linear algebra architectures. IEEE Transactions on Computers, 61(12): 1724–1736, December 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Park:2014:MXM

[PVKA14] Dongkook Park, Aniruddha Vaidya, Akhilesh Kumar, and Mani Azimi. MoDe-X: Microarchitecture of a layout-aware modular decoupled crossbar for on-chip interconnects. IEEE Transactions on Computers, 63 (3):622–636, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Papadopoulos:2016:SMO

[PWTS16] Lazaros Papadopoulos, Ivan Walulya, Philippas Tsigas, and Dimitrios Soudris. A systematic methodology for optimization of applications utilizing concurrent data structures. IEEE Transactions on Computers, 65 (7):2019–2031, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pao:2011:SSE

[PWW^{+}11] Derek Pao, Xing Wang, Xiaoran Wang, Cong Cao, and Yuesheng Zhu. String searching engine for virus scanning. IEEE Transactions on Computers, 60(11): 1596–1609, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5669261.

Pang:2019:PHS

[PZZQ19] Ning Pang, Jifu Zhang, Chaowei Zhang, and Xiao Qin. Parallel hierarchical subspace clustering of categorical data. IEEE Transactions on Computers, 68(4):542–555, April 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8520776/.

Quislant:2013:LSL

[QGPZ13] Ricardo Quislant, Eladio Gutierrez, Oscar Plata, and Emilio L. Zapata. LS-Sig: Locality-sensitive signatures for transactional memory. IEEE Transactions on Computers, 62(2):322–335, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qin:2010:CAL

[QJM^{+}10] Xiao Qin, Hong Jiang, A. Manzanares, Xiaojun Ruan, and Shu Yin. Communicationaware load balancing for parallel applications on clusters. IEEE Transactions on Computers, 59(1):42–52, January 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5184810.

Qiu:2016:WMA

[QLH^{+}16] Keni Qiu, Qingan Li, Jingtong Hu, Weigong Zhang, and Chun Jason Xue. Write mode aware loop tiling for high performance low power volatile PCM in embedded systems. IEEE Transactions on Computers, 65 (7):2313–2324, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qian:2011:AFT

[QLR^{+}11] Weikang Qian, Xin Li, Marc D. Riedel, Kia Bazargan, and David J. Lilja. An architecture for fault-tolerant computation with stochastic logic. IEEE Transactions on Computers, 60 (1):93–105, January 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qiu:2015:PCM

[QML^{+}15] M. Qiu, Z. Ming, J. Li, K. Gai, and Z. Zong. Phasechange memory optimization for green cloud with genetic algorithm. IEEE Transactions on Computers, 64(12): 3528–3540, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Quinones:2010:LRW

[QPG10] E. Quiñones, J. Parcerisa, and A. González. Leveraging register windows to reduce physical registers to the bare minimum. IEEE Transactions on Computers, 59(12): 1598–1610, December 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5453344.

Qian:2017:PPS

[QQW^{+}17] Jianwei Qian, Fudong Qiu, Fan Wu, Na Ruan, Guihai Chen, and Shaojie Tang. Privacy-preserving selective aggregation of online user behavior data. IEEE Transactions on Computers, 66 (2):326–338, February 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qiu:2016:LLM

[QSYS16] Chenxi Qiu, Haiying Shen, Lei Yu, and Sohraab Soltani. Low-latency multi-flow cooperative broadcast in fading wireless networks. IEEE Transactions on Computers, 65(6):1802–1815, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qiu:2013:PNT

[QWB^{+}13] Qinru Qiu, Qing Wu, Morgan Bishop, Robinson E. Pino, and Richard W. Linderman. A parallel neuromorphic text recognition system and its implementation on a heterogeneous high-performance computing cluster. IEEE Transactions on Computers, 62(5):886– 899, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qian:2015:MGL

[QZC15] J. Qian, Q. Zhu, and H. Chen. Multi-granularity locality-sensitive Bloom filter. IEEE Transactions on Computers, 64(12):3500– 3514, December 2015. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qi:2016:SPR

[QZL^{+}16] Saiyu Qi, Yuanqing Zheng, Mo Li, Li Lu, and Yunhao Liu. Secure and private RFID-enabled thirdparty supply chain systems. IEEE Transactions on Computers, 65(11):3413–3426, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rahimi:2014:AAG

[RBG14] Abbas Rahimi, Luca Benini, and Rajesh K. Gupta. Application-adaptive guardbanding to mitigate static and dynamic variability. IEEE Transactions on Computers, 63(9):2160–2173, September 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Roy:2019:CMF

[RBG^{+}19] Debapriya Basu Roy, Shivam Bhasin, Sylvain Guilley, Annelie Heuser, Sikhar Patranabis, and Debdeep Mukhopadhyay. CC meets FIPS: A hybrid test methodology for first order side channel analysis. IEEE Transactions on Computers, 68(3):347–361, March 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8490742/.

Randolph:2015:SCO

[RBIQ15] A. Randolph, H. Boucheneb, A. Imine, and A. Quintero. On synthesizing a consistent operational transformation approach. IEEE Transactions on Computers, 64 (4):1074–1089, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ratanaworabhan:2012:ERD

[RBK^{+}12] Paruj Ratanaworabhan, Martin Burtscher, Darko Kirovski, Benjamin Zorn, Rahul Nagpal, and Karthik Pattabiraman. Efficient runtime detection and toleration of asymmetric races. IEEE Transactions on Computers, 61(4):548–562, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Reviriego:2011:ODL

[RBMO11] Pedro Reviriego, Chris Bleakley, Juan Antonio Maestro, and Anne O’Donnell. Offset DMR: a low overhead soft error detection and correction technique for transform-based convolution. IEEE Transactions on Computers, 60(10): 1511–1516, October 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=6008538.

Ripoll:2013:PSM

[RBR13] Ismael Ripoll and Rafael Ballester-Ripoll. Period selection for minimal hyperperiod in periodic task systems. IEEE Transactions on Computers, 62(9): 1813–1822, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rezende:2015:RSU

[RBRL15] C. Rezende, A. Boukerche, H. S. Ramos, and A. A. F. Loureiro. A reactive and scalable unicast solution for video streaming over VANETs. IEEE Transactions on Computers, 64 (3):614–626, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Richter:2014:OTP

[RC14] Michael Richter and Krishnendu Chakrabarty. Optimization of test pin-count, test scheduling, and test access for NoC-based multicore SoCs. IEEE Transactions on Computers, 63 (3):691–702, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rahman:2014:CCO

[RCC14] Musfiq Rahman, Bruce R.

Childers, and Sangyeun Cho. COMeT+: Continuous online memory testing with multi-threading extension. IEEE Transactions on Computers, 63(7):1668– 1681, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ros:2012:EMC

[RCFP^{+}12] Alberto Ros, Blas Cuesta, Ricardo Fernandez-Pascual, Maria E. Gomez, Manuel E. Acacio, Antonio Robles, Jose M. Garcia, and Jose Duato. Extending Magny-Cours cache coherence. IEEE Transactions on Computers, 61(5):593–606, May 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rehman:2016:CLS

[RCK^{+}16] S. Rehman, K. Chen, F. Kriebel, A. Toma, M. Shafique, J. Chen, and J. Henkel. Cross-layer software dependability on unreliable hardware. IEEE Transactions on Computers, 65(1):80–94, January 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Radojkovic:2016:TAM

[RCM^{+}16] P. Radojkovic, P. M. Carpenter, M. Moreto, V. Cakarevic, J. Verdu, A. Pajuelo, F. J. Cazorla, M. Nemirovsky, and M. Valero. Thread assignment in multicore/multithreaded processors: A statistical approach. IEEE Transactions on Computers, 65(1):256–269, January 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Reda:2011:ITT

[RCN11] Sherief Reda, Ryan J. Cochran, and Abdullah Nazma Nowroz. Improved thermal tracking for processors using hard and soft sensor allocation techniques. IEEE Transactions on Computers, 60(6):841–851, June 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rottenstreich:2013:EWC

[RCRK13] Ori Rottenstreich, Rami Cohen, Danny Raz, and Isaac Keslassy. Exact worst case TCAM rule expansion. IEEE Transactions on Computers, 62(6):1127–1140, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Roohi:2018:NCN

[RD18] Arman Roohi and Ronald F.

DeMara. NV-clustering: Normally-off computing using non-volatile datapaths. IEEE Transactions on Computers, 67(7):949–959, July 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.
org/document/8263396/.

Refaei:2010:ARM

[RDEN10] M. T. Refaei, L. A. DaSilva, M. Eltoweissy, and T. Nadeem. Adaptation of reputation management systems to dynamic network conditions in ad hoc networks. IEEE Transactions on Computers, 59(5):707–719, May 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5406505.

Redinbo:2011:SWS

[Red11] G. Robert Redinbo. Systematic wavelet subcodes for data protection. IEEE Transactions on Computers, 60(6):904–909, June 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Redinbo:2014:CDC

[Red14] G. Robert Redinbo. Correcting DFT codes with a modified Berlekamp–Massey algorithm and Kalman recursive syndrome extension. IEEE Transactions on Computers, 63(1):196–203, January 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Redinbo:2018:DCD

[Red18] G. Robert Redinbo. Designing checksums for detecting errors in fast unitary transforms. IEEE Transactions on Computers, 67(4):566–572, April 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8039530/.

Redinbo:2019:TPD

[Red19] G. R. Redinbo. Tensor product DFT codes vs standard DFT codes. IEEE Transactions on Computers, 68(11): 1678–1688, November 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rattanatamrong:2014:DSR

[RF14] Prapaporn Rattanatamrong and Jose A. B. Fortes. Dynamic scheduling of real-time mixture-of-experts systems on limited resources. IEEE Transactions on Computers, 63(7):1751–1764, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rashid:2015:MAP

[RGK15] M. M. Rashid, I. Gondal, and J. Kamruzzaman. Mining associated patterns from wireless sensor networks. IEEE Transactions on Computers, 64(7):1998–2011, July 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ramamritham:2014:SAD

[RHC^{+}14] Krithi Ramamritham, Song Han, Deji Chen, Ming Xiong, Kam-Yiu Lam, and Aloysius K. Mok. Schedulability analysis of deferrable scheduling algorithms for maintaining real-time data freshness. IEEE Transactions on Computers, 63 (4):979–994, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Roy:2018:HFB

[RJV^{+}18] Sujoy Sinha Roy, Kimmo Järvinen, Jo Vliegen, Frederik Vercauteren, and Ingrid Verbauwhede. HEPCloud: An FPGA-based multicore processor for FV somewhat homomorphic function evaluation. IEEE Transactions on Computers, 67(11):

1637–1650, November 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8318681/.

Rodchenko:2018:TIE

[RKN^{+}18] Andrey Rodchenko, Christos Kotselidis, Andy Nisbet, Antoniu Pop, and Mikel Luján. Type information elimination from objects on architectures with tagged pointers support. IEEE Transactions on Computers, 67 (1):130–143, January 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7962268/.

Rajendran:2015:ITV

[RKR15] J. Rajendran, R. Karri, and G. S. Rose. Improving tolerance to variations in memristor-based applications using parallel memristors. IEEE Transactions on Computers, 64(3):733– 746, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rodriguez:2019:AMP

[RKT19] Gabriel Rodríguez, Mahmut T. Kandemir, and Juan Touriño. Affine modeling of program traces. IEEE Transactions on Computers, 68(2):294–300, February 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8408540/.

Ren:2016:FAL

[RKZ16] P. Ren, M. A. Kinsy, and N. Zheng. Fault-aware loadbalancing routing for 2Dmesh and torus on-chip network topologies. IEEE Transactions on Computers, 65(3):873–887, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ramanujam:2013:RTO

[RL13] Rohit Sunkam Ramanujam and Bill Lin. Randomized throughput-optimal oblivious routing for torus networks. IEEE Transactions on Computers, 62(3):561– 574, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Reviriego:2019:TBO

[RLRL19] P. Reviriego, S. Liu, O. Rottenstreich, and F. Lombardi. Two bit overlap: A class of double error correction one step majority logic decodable codes. IEEE Transactions on Computers, 68(5):798–803, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ryu:2018:CLP

[RLSK18] Junhee Ryu, Dongeun Lee, Kang G. Shin, and Kyungtae Kang. ClusterFetch: A lightweight prefetcher for intensive disk reads. IEEE Transactions on Computers, 67(2):284–290, February 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8025580/.

Ren:2015:DCM

[RLX15] Xiaojiang Ren, Weifa Liang, and Wenzheng Xu. Data collection maximization in renewable sensor networks via time-slot scheduling. IEEE Transactions on Computers, 64(7):1870–1883, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rebeiro:2015:MAA

[RM15a] C. Rebeiro and D. Mukhopadhyay. Micro-architectural analysis of time-driven cache attacks: Quest for the ideal implementation. IEEE Transactions on Computers, 64(3):778–790, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Reyhani-Masoleh:2015:CLL

[RM15b] A. Reyhani-Masoleh. Comments on “Low-Latency Digit-Serial Systolic Double Basis Multiplier over Using Subquadratic Toeplitz Matrix-Vector Product Approach”. IEEE Transactions on Computers, 64 (4):1215–1216, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [CLL^{+}14].

Rodriguez:2015:ECP

[RM15c] A. Rodriguez and F. Moreno.

Evolutionary computing and particle filtering: A hardwarebased motion estimation system. IEEE Transactions on Computers, 64(11): 3140–3152, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rafiev:2012:MRR

[RMB^{+}12] Ashur Rafiev, Andrey Mokhov, Frank P. Burns, Julian P. Murphy, Albert Koelmans, and Alex Yakovlev. Mixed radix Reed–Muller expansions. IEEE Transactions on Computers, 61(8):1189– 1202, August 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rahmati:2013:CAP

[RMB^{+}13] Dara Rahmati, Srinivasan Murali, Luca Benini, Federico Angiolini, Giovanni De Micheli, and Hamid Sarbazi-Azad. Computing accurate performance bounds for best effort networks-on-chip. IEEE Transactions on Computers, 62(3):452– 467, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rajapaksha:2015:VCA

[RMC^{+}15] N. Rajapaksha, A. Madanayake, R. J. Cintra, J. Adikari, and V. S. Dimitrov. VLSI computational architectures for the arithmetic cosine transform. IEEE Transactions on Computers, 64(9): 2708–2715, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Reyhani-Masoleh:2019:NMI

[RMERM19] A. Reyhani-Masoleh, H. El-Razouk, and A. Monfared. New multiplicative inverse architectures using Gaussian normal basis. IEEE Transactions on Computers, 68(7):991–1006, July 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rajendran:2012:EEM

[RMKR12] Jeyavijayan Rajendran, Harika Manem, Ramesh Karri, and Garrett S. Rose. An energyefficient memristive threshold logic circuit. IEEE Transactions on Computers, 61(4):474–487, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ruj:2013:PTK

[RNS13] Sushmita Ruj, Amiya Nayak, and Ivan Stojmenovic. Pairwise and triple key distribution in wireless sensor networks with applications. IEEE Transactions on Computers, 62(11): 2224–2237, November 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ratan:2011:SRQ

[RO11] Rahul Ratan and A. Yavuz Oruc. Self-routing quantum sparse crossbar packet concentrators. IEEE Transactions on Computers, 60(10): 1390–1405, October 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5487508.

Rodriguez-Olivares:2018:FBD

[ROGHNB^{+}18] N. A. Rodríguez-Olivares, A. Gómez-Hernández, L. Nava-Balanzar, H. Jiménez-Hernández, and J. A. Soto-Cajiga. FPGA-based data storage system on NAND flash memory in RAID 6 architecture for in-line pipeline inspection gauges. IEEE Transactions on Computers, 67 (7):1046–1053, July 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8315494/.

Rafferty:2017:ELI

[ROH17] Ciara Rafferty, Máire O’Neill, and Neil Hanley. Evaluation of large integer multiplication methods on hardware. IEEE Transactions on Computers, 66(8):1369– 1382, August 2017. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7869256/.

Raghavendra:2016:PPB

[RPM16] K. Raghavendra, Biswabandan Panda, and Madhu Mutyam. PBC: Prefetched blocks compaction. IEEE Transactions on Computers, 65(8):2534–2547, August 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Romano:2014:DEP

[RQ14] Paolo Romano and Francesco Quaglia. Design and evaluation of a parallel invocation protocol for transactional applications over the Web. IEEE Transactions on Computers, 63 (2):317–334, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Roy:2011:SRP

[RRK11] Soumyaroop Roy, Nagarajan Ranganathan, and Srinivas Katkoori. State-retentive power gating of register files in multicore processors featuring multithreaded inorder cores. IEEE Transactions on Computers, 60(11): 1547–1560, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5669257.

Ren:2016:DFC

[RRS^{+}16] P. Ren, X. Ren, S. Sane, M. A. Kinsy, and N. Zheng. A deadlock-free and connectivityguaranteed methodology for achieving fault-tolerance in on-chip networks. IEEE Transactions on Computers, 65(2):353–366, February 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rodrigues:2010:ACU

[RS10] T. K. Rodrigues and E. E.

Swartzlander. Adaptive CORDIC: Using parallel angle recoding to accelerate rotations. IEEE Transactions on Computers, 59 (4):522–531, April 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.

                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5374379.

Ruhrup:2013:OCO

[RS13] Stefan Ruhrup and Ivan Stojmenovic. Optimizing communication overhead while reducing path length in beaconless georouting with guaranteed delivery for wireless sensor networks. IEEE Transactions on Computers, 62(12):2440–2453, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rawat:2017:VIS

[RS17] Hemendra Rawat and Patrick Schaumont. Vector instruction set extensions for efficient computation of Keccak. IEEE Transactions on Computers, 66(10):1778– 1789, October 2017. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7918507/.

Roy:2019:EHP

[RSG^{+}19] I. Roy, A. Srivastava, M. Grimm, M. Nourian, M. Becchi, and S. Aluru. Evaluating high performance pattern matching on the automata processor. IEEE Transactions on Computers, 68(8):1201–1212, August 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Raha:2017:QCA

[RSJR17] Arnab Raha, Soubhagya Sutar, Hrishikesh Jayakumar, and Vijay Raghunathan. Quality configurable approximate DRAM. IEEE Transactions on Computers, 66 (7):1172–1187, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07784741-
abs.html.

Rodrigues:2018:CAS

[RSN^{+}18] Tiago Gama Rodrigues, Katsuya Suto, Hiroki Nishiyama, Nei Kato, and Katsuhiro Temma. Cloudlets activation scheme for scalable mobile edge computing with transmission power control and virtual machine migration. IEEE Transactions on Computers, 67(9): 1287–1300, September 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8322166/.

Rodrigues:2017:HMM

[RSNK17] T. G. Rodrigues, K. Suto, H. Nishiyama, and N. Kato. Hybrid method for minimizing service delay in edge cloud computing through VM migration and transmission power control. IEEE Transactions on Computers, 66(5):810–819, May 2017.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Reorda:2017:EDS

[RSU17] Matteo Sonza Reorda, Luca Sterpone, and Anees Ullah. An error-detection and selfrepairing method for dynamically and partially reconfigurable systems. IEEE Transactions on Computers, 66 (6):1022–1033, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7563888/.

Revol:2014:NRP

[RT14] Nathalie Revol and Philippe Theveny. Numerical reproducibility and parallel computations: Issues for interval algorithms. IEEE Transactions on Computers, 63 (8):1915–1924, August 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ren:2018:LHA

[RTL^{+}18] Shiru Ren, Le Tan, Chunqi Li, Zhen Xiao, and Weijia Song. Leveraging hardwareassisted virtualization for deterministic replay on commodity multi-core processors. IEEE Transactions on Computers, 67(1):45– 58, January 2018. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7982675/.

Ramos:2019:APM

[RTRM19] A. Ramos, R. G. Toral, P. Reviriego, and J. A. Maestro. An ALU protection methodology for soft processors on SRAM-based FPGAs. IEEE Transactions on Computers, 68(9): 1404–1410, September 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ramos:2018:EPR

[RURM18] Alexis Ramos, Anees Ullah, Pedro Reviriego, and Juan Antonio Maestro. Efficient protection of the register file in soft-processors implemented on Xilinx FP-GAs. IEEE Transactions on Computers, 67(2):299– 304, February 2018. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8008792/.

Russinoff:2013:CFV

[Rus13] David M. Russinoff. Computation and formal verification of SRT quotient and square root digit selection tables. IEEE Transactions on Computers, 62(5):900– 913, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Radonjic:2013:ICC

[RV13] Aleksandar Radonjic and Vladimir Vujicic. Integer codes correcting burst errors within a byte. IEEE Transactions on Computers, 62 (2):411–415, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ranieri:2015:NOT

[RVC^{+}15] J. Ranieri, A. Vincenzi, A. Chebira, D. Atienza, and M. Vetterli. Near-optimal thermal monitoring framework for many-core systemson-chip. IEEE Transactions on Computers, 64(11): 3197–3209, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rahulamathavan:2016:UCA

[RVH^{+}16] Yogachandran Rahulamathavan, Suresh Veluru, Jinguang Han, Fei Li, Muttukrishnan Rajarajan, and Rongxing Lu. User collusion avoidance scheme for privacy-preserving decentralized key-policy attributebased encryption. IEEE Transactions on Computers, 65(9):2939–2946, September 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rahmani:2014:HPF

[RVL^{+}14] Amir-Mohammad Rahmani, Kameswar Rao Vaddina, Khalid Latif, Pasi Liljeberg, Juha Plosila, and Hannu Tenhunen. Highperformance and fault-tolerant 3D NoC-bus hybrid architecture using ARB–NETbased adaptive monitoring platform. IEEE Transactions on Computers, 63 (3):734–747, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ramanathan:2018:SWC

[RWC18] Nadesh Ramanathan, John Wickerson, and George A. Constantinides. Scheduling weakly consistent C concurrency for reconfigurable hardware. IEEE Transactions on Computers, 67 (7):992–1006, July 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8241825/.

Ren:2014:QAD

[RWZZ14] Jinglei Ren, Yongwei Wu, Meiqi Zhu, and Weimin Zheng. Quatrain: Accelerating data aggregation between multiple layers. IEEE Transactions on Computers, 63(5):1207–1219, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ros:2015:ASC

[RXC^{+}15] A. Ros, P. Xekalakis, M. Cintra, M. E. Acacio, and J. M. Garcia. Adaptive selection of cache indexing bits for removing conflict misses. IEEE Transactions on Computers, 64(6):1534–1547, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ren:2019:PLL

[RZPX19] Shiru Ren, Yunqi Zhang, Lichen Pan, and Zhen Xiao. Phantasy: Low-latency virtualization-based fault tolerance via asynchronous prefetching. IEEE Transactions on Computers, 68 (2):225–238, February 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8438984/.

Rajendran:2015:FAB

[RZZ^{+}15] J. Rajendran, Huan Zhang, Chi Zhang, G. S. Rose, Youngok Pino, O. Sinanoglu, and R. Karri. Fault analysisbased logic encryption. IEEE Transactions on Computers, 64(2):410–424, February 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sanchez:2011:ISR

[SAR^{+}11] Pedro Sanchez, Diego Alonso, Francisca Rosique, Barbara Alvarez, and Juan A. Pastor. Introducing safety requirements traceability support in model-driven development of robotic applications. IEEE Transactions on Computers, 60 (8):1059–1071, August 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sarikaya:2010:UPM

[SB10] R. Sarikaya and A. Buyuktosunoglu. A unified prediction method for predicting program behavior. IEEE Transactions on Computers, 59 (2):272–282, February 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5184824.

S:2016:RSE

[SB16] Anjan Kumar Pudi N S and Maryam Shojaei Baghini. Robust soft error tolerant CMOS latch configurations. IEEE Transactions on Computers, 65(9): 2820–2834, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Samanta:2018:CCB

[SBB18] Jagannath Samanta, Jaydeb Bhaumik, and Soma Barman. Compact CA-based single byte error correcting codec. IEEE Transactions on Computers, 67 (2):291–298, February 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8010467/.

Samarah:2011:TAR

[SBH11] Samer Samarah, Azzedine Boukerche, and Alexander Shema Habyalimana. Target association rules: a new behavioral patterns for point of coverage wireless sensor networks. IEEE Transactions on Computers, 60(6):879–889, June 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Srinivasagopalan:2012:OST

[SBI12] Srivathsan Srinivasagopalan, Costas Busch, and S. S. Iyengar. An oblivious spanning tree for single-sink buyat-bulk in low doublingdimension graphs. IEEE Transactions on Computers, 61(5):700–712, May 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sarkar:2015:DFA

[SBM15] S. Sarkar, S. Banik, and S. Maitra. Differential fault attack against grain family with very few faults and minimal assumptions. IEEE Transactions on Computers, 64(6):1647–1657, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Swierczynski:2018:BFI

[SBMP18] Pawel Swierczynski, Georg T.

Becker, Amir Moradi, and Christof Paar. Bitstream fault injections (BiFI)-automated fault attacks against SRAMbased FPGAs. IEEE Transactions on Computers, 67 (3):348–360, March 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7809042/.

Saponara:2014:DNI

[SBP^{+}14] Sergio Saponara, Tony Bacchillone, Esa Petri, Luca Fanucci, Riccardo Locatelli, and Marcello Coppola. Design of an NoC interface macrocell with hardware support of advanced networking functionalities. IEEE Transactions on Computers, 63(3):609– 621, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sauer:2016:PSB

[SBP16] Matthias Sauer, Bernd Becker, and Ilia Polian. PHAETON: A SAT-based framework for timing-aware path sensitization. IEEE Transactions on Computers, 65(6):1869–1881, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2016:AOD

[SBW^{+}16] Zhenyu Sun, Xiuyuan Bi, Wenqing Wu, Sungjoo Yoo, and Hai (Helen) Li. Array organization and data management exploration in racetrack memory. IEEE Transactions on Computers, 65(4):1041–1054, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shamshiri:2011:MYC

[SC11] Saeed Shamshiri and Kwang-Ting (Tim) Cheng. Modeling yield, cost, and quality of a spare-enhanced multicore chip. IEEE Transactions on Computers, 60(9): 1246–1259, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5710882.

Saravanan:2018:PCE

[SC18] Karthikeyan P. Saravanan and Paul M. Carpenter. PerfBound: Conserving energy with bounded overheads in on/off-based HPC interconnects. IEEE Transactions on Computers, 67(7): 960–974, July 2018. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8248748/.

Shin:2016:QDQ

[SCJ^{+}16a] Wongyu Shin, Jungwhan Choi, Jaemin Jang, Jinwoong Suh, Yongkee Kwon, Youngsuk Moon, Hongsik Kim, and Lee-Sup Kim. Q-DRAM: Quick-access DRAM with decoupled restoring from row-activation. IEEE Transactions on Computers, 65(7):2213–2227, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shin:2016:DLO

[SCJ^{+}16b] Wongyu Shin, Jungwhan Choi, Jaemin Jang, Jinwoong Suh, Youngsuk Moon, Yongkee Kwon, and Lee-Sup Kim. DRAM-latency optimization inspired by relationship between row-access time and refresh timing. IEEE Transactions on Computers, 65(10):3027–3040, October 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sengupta:2010:GTF

[SCK10] S. Sengupta, M. Chatterjee, and K. A. Kwiat. A game theoretic framework for power control in wireless sensor networks. IEEE Transactions on Computers, 59(2):231–242, February 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5089316.

Sune:2010:CPA

[SCNS10] V. Sune, J. A. Carrasco, H. Nabli, and B. Sericola. Comment on “Performability Analysis: a New Algorithm”. IEEE Transactions on Computers, 59 (1):137–138, January 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5184816.

Shi:2012:VGA

[SCSL12] Lin Shi, Hao Chen, Jianhua Sun, and Kenli Li. vCUDA: GPU-accelerated high-performance computing in virtual machines. IEEE Transactions on Computers, 61(6):804–816, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sha:2016:NDM

[SCZ^{+}16] Edwin H.-M. Sha, Xianzhang Chen, Qingfeng Zhuge, Liang Shi, and Weiwen Jiang. A new design of in-memory file system based on file virtual address framework. IEEE Transactions on Computers, 65(10):

2959–2972, October 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Singh:2013:AGA

[SD13] Rajwinder Singh and Mayank Dave. Antecedence graph approach to checkpointing for fault tolerance in mobile agent systems. IEEE Transactions on Computers, 62 (2):247–258, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shirley:2014:CMC

[SD14] C. G. Shirley and W. R.

Daasch. Copula models of correlation: A DRAM case study. IEEE Transactions on Computers, 63(10): 2389–2401, October 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Streit:2018:PQK

[SD18] Silvan Streit and Fabrizio De Santis. Post-quantum key exchange on ARMv8-A: a new hope for NEON made simple. IEEE Transactions on Computers, 67(11): 1651–1662, November 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8107588/.

Schley:2017:MLD

[SDE^{+}17] G. Schley, A. Dalirsani, M. Eggenberger, N. Hatami, H. J. Wunderlich, and M. Radetzki. Multi-layer diagnosis for fault-tolerant networks-on-chip. IEEE Transactions on Computers, 66(5):848–861, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2019:APT

[SDL^{+}19] Chang-Ai Sun, Hepeng Dai, Huai Liu, Tsong Yueh Chen, and Kai-Yuan Cai. Adaptive partition testing. IEEE Transactions on Computers, 68(2):157–169, February 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8440117/.

Sethia:2012:CPE

[SDMM12] Ankit Sethia, Ganesh Dasika, Trevor Mudge, and Scott Mahlke. A customized processor for energy efficient scientific computing. IEEE Transactions on Computers, 61(12):1711–1723, December 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Strollo:2011:EFH

[SDP11] Antonio Giuseppe Maria Strollo, Davide De Caro, and Nicola Petra. Elementary functions hardware implementation using constrained piecewise-polynomial approximations. IEEE Transactions on Computers, 60 (3):418–432, March 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Savino:2012:SRE

[SDP^{+}12] Alessandro Savino, Stefano Di Carlo, Gianfranco Politano, Alfredo Benso, Alberto Bosio, and Giorgio Di Natale. Statistical reliability estimation of microprocessorbased systems. IEEE Transactions on Computers, 61 (11):1521–1534, November 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Saqib:2015:PDT

[SDP^{+}15] F. Saqib, A. Dutta, J. Plusquellic, P. Ortiz, and M. S. Pattichis. Pipelined decision tree classification accelerator implementation in FPGA (DT-CAIF). IEEE Transactions on Computers, 64 (1):280–285, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sundaresan:2015:ZKG

[SDZ15] S. Sundaresan, R. Doss, and Wanlei Zhou. Zero knowledge grouping proof protocol for RFID EPC C1G2 tags. IEEE Transactions on Computers, 64(10):2994– 3008, October 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Savas:2014:SMQ

[SEY14] Erkay Savas, Serdar Suer Erdem, and Kazim Yumbul. On selection of modulus of quadratic codes for the protection of cryptographic operations against fault attacks. IEEE Transactions on Computers, 63 (5):1182–1196, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Spinellis:2017:EUP

[SF17] Diomidis Spinellis and Marios Fragkoulis. Extending Unix pipelines to DAGs. IEEE Transactions on Computers, 66(9):1547–1561, September 2017. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7903579/.

Shen:2012:EAA

[SG12] Hong Shen and Longkun Guo. Efficient 2-approximation algorithms for computing 2connected Steiner minimal networks. IEEE Transactions on Computers, 61(7): 954–968, July 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2013:EAA

[SG13] Hong Shen and Longkun Guo. An eight-approximation algorithm for computing rooted three-vertex connected minimum Steiner networks. IEEE Transactions on Computers, 62(9): 1684–1693, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2012:LPC

[SH12] Haiying Shen and Kai Hwang. Locality-preserving clustering and discovery of resources in wide-area distributed computational grids. IEEE Transactions on Computers, 61(4):458– 473, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Schneider:2015:SDP

[SHGW15] S. Schneider, M. Hirzel, B. Gedik, and Kun-Lung Wu. Safe data parallelism for general streaming. IEEE Transactions on Computers, 64(2):504–517, February 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Suzuki:2016:DSD

[SHH^{+}16] Jun Suzuki, Yoichi Hidaka, Junichi Higuchi, Yuki Hayashi, Masaki Kan, and Takashi Yoshikawa. Disaggregation and sharing of I/O devices in cloud data centers. IEEE Transactions on Computers, 65(10): 3013–3026, October 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sarikaya:2013:RAB

[SIB13] Ruhi Sarikaya, Canturk Isci, and Alper Buyuktosunoglu. Runtime application behavior prediction using a statistical metric model. IEEE Transactions on Computers, 62(3):575–588, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shafique:2016:SPM

[SIVH16] Muhammad Shafique, Anton Ivanov, Benjamin Vogel, and Jörg Henkel. Scalable power management for onchip systems with malleable applications. IEEE Transactions on Computers, 65(11): 3398–3412, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shin:2017:BGL

[SJC^{+}17a] Wongyu Shin, Jaemin Jang, Jungwhan Choi, Jinwoong Suh, and Lee-Sup Kim. Bank-group level parallelism. IEEE Transactions on Computers, 66(8): 1428–1434, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7845573/.

Shin:2017:RLP

[SJC^{+}17b] Wongyu Shin, Jaemin Jang, Jungwhan Choi, Jinwoong Suh, Yongkee Kwon, Youngsuk Moon, and Lee-Sup Kim. Rank-level parallelism in DRAM. IEEE Transactions on Computers, 66 (7):1274–1280, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07820157-
abs.html.

Sha:2018:TDE

[SJD^{+}18] Edwin Hsing-Mean Sha, Weiwen Jiang, Hailiang Dong, Zhulin Ma, Runyu Zhang, Xianzhang Chen, and Qingfeng Zhuge. Towards the design of efficient and consistent index structure with minimal write activities for nonvolatile memory. IEEE Transactions on Computers, 67(3):432–448, March 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8046077/.

Stojkovic:2010:IAC

[SJS10] S. Stojkovic, D. Jankovic, and R. S. Stankovic. An improved algorithm for the construction of decision diagrams by rearranging and partitioning the input cube set. IEEE Transactions on Computers, 59(8):1105– 1119, August 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5396330.

Shen:2014:DIM

[SJS^{+}14] Yunfu Shen, Yi Jin, Rong Shen, Junjie Peng, and Sheng Luo. Design and implementation of modified signed-digit adder. IEEE Transactions on Computers, 63(5):1134–1143, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sem-Jacobsen:2011:DFT

[SJSLD11] Frank Olaf Sem-Jacobsen, Tor Skeie, Olav Lysne, and Jose Duato. Dynamic fault tolerance in fat trees. IEEE Transactions on Computers, 60(4):508–525, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sen:2019:SSA

[SJVR19] S. Sen, S. Jain, S. Venkataramani, and A. Raghunathan. SparCE: Sparsity aware general-purpose core extensions to accelerate deep neural networks. IEEE Transactions on Computers, 68(6):912–925, June 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sinanoglu:2010:ISA

[SKA10] O. Sinanoglu, M. H. Karaata, and B. AlBdaiwi. An inherently stabilizing algorithm for node-to-node routing over all shortest node-disjoint paths in hypercube networks. IEEE Transactions on Computers, 59(7):995–999, July 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5445078.

Seo:2014:IWC

[SKC^{+}14] Bumjoon Seo, Sooyong Kang, Jongmoo Choi, Jaehyuk Cha, Youjip Won, and Sungroh Yoon. IO workload characterization revisited: A data-mining approach. IEEE Transactions on Computers, 63(12): 3026–3038, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Salamat:2016:RRA

[SKEB16] Ronak Salamat, Misagh Khayambashi, Masoumeh Ebrahimi, and Nader Bagherzadeh. A resilient routing algorithm with formal reliability analysis for partially connected 3D-NoCs. IEEE Transactions on Computers, 65(11): 3265–3279, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Salamat:2018:LAN

[SKEB18] Ronak Salamat, Misagh Khayambashi, Masoumeh Ebrahimi, and Nader Bagherzadeh. LEAD: An adaptive 3D-NoC routing algorithm with queuing-theory based analytical verification. IEEE Transactions on Computers, 67(8):1153–1166, August 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8283712/.

Shafique:2016:CAL

[SKH16] Muhammad Shafique, Muhammad Usman Karim Khan, and Jörg Henkel. Contentaware low-power configurable aging mitigation for SRAM memories. IEEE Transactions on Computers, 65(12):3617–3630, December 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Soltiz:2013:MBN

[SKM^{+}13] Michael Soltiz, Dhireesha Kudithipudi, Cory Merkel, Garrett S. Rose, and Robinson E. Pino. Memristorbased neural logic blocks for nonlinearly separable functions. IEEE Transactions on Computers, 62(8):1597– 1606, August 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shim:2014:SWC

[SKM14] Hyotaek Shim, Jin-Soo Kim, and Seungryoul Maeng. System-wide cooperative optimization for NAND flashbased mobile systems. IEEE Transactions on Computers, 63(8):2052–2065, August 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sadiq:2015:SCO

[SKPC15] U. Sadiq, M. Kumar, A. Passarella, and M. Conti. Service composition in opportunistic networks: A load and mobility aware solution. IEEE Transactions on Computers, 64(8):2308– 2322, August 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sarkar:2010:NCH

[SKPK10] S. Sarkar, G. R. Kulkarni, P. P. Pande, and A. Kalyanaraman. Network-on-chip hardware accelerators for biological sequence alignment. IEEE Transactions on Computers, 59(1):29–41, January 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5255227.

Suzuki:2016:GGV

[SKYK16] Yusuke Suzuki, Shinpei Kato, Hiroshi Yamada, and Kenji Kono. GPUvm: GPU virtualization at the hypervisor. IEEE Transactions on Computers, 65(9): 2752–2766, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sakai:2013:DBE

[SKZS13] Kazuya Sakai, Wei-Shinn Ku, Roger Zimmermann, and Min-Te Sun. Dynamic bit encoding for privacy protection against correlation attacks in RFID backward channel. IEEE Transactions on Computers, 62 (1):112–123, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shieh:2010:WBM

[SL10] Ming-Der Shieh and Wen-Ching Lin. Word-based Montgomery modular multiplication algorithm for low-latency scalable architectures. IEEE Transactions on Computers, 59(8): 1145–1151, August 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5441286.

Shen:2013:LCM

[SL13] Haiying Shen and Guoxin Liu. A lightweight and cooperative multifactor considered file replication method in structured P2P systems. IEEE Transactions on Computers, 62(11):2115–2130, November 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2014:LSN

[SL14a] Haiying Shen and Ze Li.

Leveraging social networks for effective spam filtering. IEEE Transactions on Computers, 63(11):2743–2759, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shin:2014:PJE

[SL14b] Kang G. Shin and Jinkyu Lee. Preempt a job or not in EDF scheduling of uniprocessor systems. IEEE Transactions on Computers, 63(5):1197–1206, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2015:SIB

[SLC15a] Haiying Shen, Guoxin Liu, and H. Chandler. Swarm intelligence based file replication and consistency maintenance in structured P2P file sharing systems. IEEE Transactions on Computers, 64(10):2953–2967, October 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2015:SSA

[SLC^{+}15b] Haiying Shen, Jinwei Liu, Kang Chen, Jianwei Liu, and S. Moyer. SCPS: A socialaware distributed cyberphysical human-centric search engine. IEEE Transactions on Computers, 64(2):518– 532, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2015:CEF

[SLL15] Haiying Shen, Yuhua Lin, and Ting Li. Combining efficiency, fidelity, and flexibility in resource information services. IEEE Transactions on Computers, 64 (2):353–367, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2015:KSO

[SLLG15] Haiying Shen, Ze Li, Jinwei Liu, and J. E. Grant. Knowledge sharing in the online social network of Yahoo! answers and its implications. IEEE Transactions on Computers, 64 (6):1715–1728, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Schaumont:2018:SSS

[SLPB18] Patrick Schaumont, Ruby Lee, Ronald Perez, and Guido Bertoni. Special section on secure computer architectures. IEEE Transactions on Computers, 67 (3):305–306, March 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shi:2012:LBM

[SLS^{+}12] Lei Shi, Bin Liu, Changhua Sun, Zhengyu Yin, Laxmi N. Bhuyan, and H. Jonathan Chao. Load-balancing multipath switching system with flow slice. IEEE Transactions on Computers, 61 (3):350–365, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shang:2015:NPE

[SLZX15] Y. Shang, D. Li, J. Zhu, and M. Xu. On the network power effectiveness of data center architectures. IEEE Transactions on Computers, 64(11):3237–3248, November 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Salkhordeh:2019:AMP

[SMA19] R. Salkhordeh, O. Mutlu, and H. Asadi. An analytical model for performance and lifetime estimation of hybrid DRAM–NVM main memories. IEEE Transactions on Computers, 68 (8):1114–1130, August 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sanchez-Macian:2019:EIT

[SMAR^{+}19] Alfonso Sánchez-Macián, Luis Alberto Aranda, Pedro Reviriego, Vahdaneh Kiani, and Juan Antonio Maestro. Enhancing instruction TLB resilience to soft errors. IEEE Transactions on Computers, 68 (2):214–224, February 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8488602/.

Sarkar:2015:PAI

[SMB^{+}15] S. Sarkar, S. Misra, B. Bandyopadhyay, C. Chakraborty, and M. S. Obaidat. Performance analysis of IEEE 802.15.6 MAC protocol under non-ideal channel conditions and saturated traffic regime. IEEE Transactions on Computers, 64(10): 2912–2925, October 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sahoo:2018:MBA

[SMCN18] Durga Prasad Sahoo, Debdeep Mukhopadhyay, Rajat Subhra Chakraborty, and Phuong Ha Nguyen. A multiplexer-based arbiter PUF composition with enhanced reliability and security. IEEE Transactions on Computers, 67 (3):403–417, March 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8025790/.

Shreejith:2017:VHP

[SME^{+}17] Shanker Shreejith, Philipp Mundhenk, Andreas Ettner, Suhaib A. Fahmy, Sebastian Steinhorst, Martin Lukasiewycz, and Samarjit Chakraborty. VEGa: A high performance vehicular Ethernet gateway on hybrid FPGA. IEEE Transactions on Computers, 66(10):

1790–1803, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7917319/.

Stefan:2014:DTN

[SMG14] Radu Andrei Stefan, Anca Molnos, and Kees Goossens. dAElite: A TDM NoC supporting QoS, multicast, and fast connection set-up. IEEE Transactions on Computers, 63(3):583–594, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Seyedzadeh:2016:IBF

[SMK^{+}16] Seyed Mohammad Seyedzadeh, Rakan Maddah, Donald Kline, Alex K. Jones, and Rami Melhem. Improving bit flip reduction for biased and random data. IEEE Transactions on Computers, 65(11):3345–3356, November 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shamim:2017:WIF

[SMN^{+}17] Md Shahriar Shamim, Naseef Mansoor, Rounak Singh Narde, Vignesh Kothandapani, Amlan Ganguly, and Jayanti Venkataraman. A wireless interconnection framework for seamless inter and intra-chip communication in multichip systems. IEEE Transactions on Computers, 66(3):389–402, March 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Saha:2016:UBE

[SMP16] B. K. Saha, S. Misra, and S. Pal. Utility-based exploration for performance enhancement in opportunistic mobile networks. IEEE Transactions on Computers, 65(4):1310–1322, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sanchez-Macian:2017:CMK

[SMRM17] Alfonso Sánchez-Macián, Pedro Reviriego, and Juan Antonio Maestro. Combined modular key and data error protection for contentaddressable memories. IEEE Transactions on Computers, 66(6):1085–1090, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7763836/.

Sanchez-Macian:2017:SET

[SMRML17] Alfonso Sánchez-Macián, Pedro Reviriego, Juan Antonio Maestro, and Shanshan Liu. Single event transient tolerant Bloom filter implementations. IEEE Transactions on Computers, 66(10): 1831–1836, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7921607/.

Sarood:2012:XCL

[SMTK12] Osman Sarood, Phil Miller, Ehsan Totoni, and Laxmikant V. Kale. ‘cood’ load balancing for high performance computing data centers. IEEE Transactions on Computers, 61(12):1752–1764, December 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shi:2016:GCR

[SN16] Xiaobing Shi and Nicola Nicolici. Generating cyclicrandom sequences in a constrained space for in-system validation. IEEE Transactions on Computers, 65(12): 3676–3686, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Skitsas:2016:DEO

[SNM16] M. A. Skitsas, C. A. Nicopoulos, and M. K. Michael. DaemonGuard: Enabling O/S-orchestrated fine-grained software-based selective-testing in multi- /many-core microprocessors. IEEE Transactions on Computers, 65(5):1453– 1466, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Seong:2010:HBM

[SNY^{+}10] Yoon Jae Seong, Eyee Hyun Nam, Jin Hyuk Yoon, Hongseok Kim, Jin yong Choi, Sookwan Lee, Young Hyun Bae, Jaejin Lee, Yookun Cho, and Sang Lyul Min. Hydra: a block-mapped parallel flash memory solid-state disk architecture. IEEE Transactions on Computers, 59(7):905–921, July 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5432160.

Shukla:2010:MQS

[SO10] M. K. Shukla and A. Y.

Oruc. Multicasting in quantum switching networks. IEEE Transactions on Computers, 59(6):735–747, June 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5416694.

Salvador:2013:SRE

[SOM^{+}13] Ruben Salvador, Andres Otero, Javier Mora, Eduardo de la Torre, Teresa Riesgo, and Lukas Sekanina. Self-reconfigurable evolvable hardware system for adaptive image processing. IEEE Transactions on Computers, 62(8):1481–1493, August 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sousa:2015:RSE

[Sou15] L. Sousa. 2^{n} RNS scalers for extended 4-moduli sets. IEEE Transactions on Computers, 64(12):3322–3334, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Simao:2010:CCT

[SP10] Adenilso Simao and Alexandre Petrenko. Checking completeness of tests for finite state machines. IEEE Transactions on Computers, 59 (8):1023–1032, August 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5396326.

Sideris:2012:CEP

[SP12] Isidoros Sideris and Kiamal Pekmestzi. Cost effective protection techniques for TCAM memory arrays. IEEE Transactions on Computers, 61(12):1778–1788, December 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sakellariou:2016:ASL

[SP16] Panagiotis Sakellariou and Vassilis Paliouras. Applicationspecific low-power multipliers. IEEE Transactions on Computers, 65(10):2973– 2985, October 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Santos:2016:DDD

[SPC^{+}16] I. L. Santos, L. Pirmez, L. R.

Carmo, P. F. Pires, F. C. Delicato, S. U. Khan, and A. Y. Zomaya. A decentralized damage detection system for wireless sensor and actuator networks. IEEE Transactions on Computers, 65(5):1363–1376, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shin:2018:EEP

[SPC^{+}18] Ho Hyun Shin, Young Min Park, Duheon Choi, Byoung Jin Kim, Dae-Hyung Cho, and Eui-Young Chung. EXTREME: Exploiting page table for reducing refresh power of 3D-stacked DRAM memory. IEEE Transactions on Computers, 67 (1):32–44, January 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7968377/.

Sterpone:2013:NFT

[SPH13] Luca Sterpone, Mario Porrmann, and Jens Hagemeyer. A novel fault tolerant and runtime reconfigurable platform for satellite payload processing. IEEE Transactions on Computers, 62 (8):1508–1525, August 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shih:2015:RFN

[SPTC15] Y. Shih, A. Pang, M. Tsai, and C. Chai. A rewarding framework for network resource sharing in co-channel hybrid access femtocell networks. IEEE Transactions on Computers, 64(11): 3079–3090, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2015:REA

[SQJ^{+}15] Hui Sun, Xiao Qin, Hong Jiang, Jianzhong Huang, and Changsheng Xie. RB-Explorer: An accurate and practical approach to write amplification measurement for SSDs. IEEE Transactions on Computers, 64 (4):1133–1148, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shukla:2014:LLH

[SR14] R. Shukla and K. C. Ray. Low latency hybrid CORDIC algorithm. IEEE Transactions on Computers, 63(12):3066–3078, December 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Salehin:2015:SMP

[SRCbL^{+}15] K. M. Salehin, R. Rojas-Cessa, Chuan bi Lin, Ziqian Dong, and T. Kijkanjanarat. Scheme to measure packet processing time of a remote host through estimation of end-link capacity. IEEE Transactions on Computers, 64(1):205–218, January 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sanyal:2010:ETL

[SRCK10] A. Sanyal, A. Rastogi, Wei Chen, and S. Kundu. An efficient technique for leakage current estimation in nanoscaled CMOS circuits incorporating self-loading effects. IEEE Transactions on Computers, 59(7):922– 932, July 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5441289.

Sudan:2012:TMI

[SRHC12] Kshitij Sudan, Karthick Rajamani, Wei Huang, and John B. Carter. Tiered memory: An iso-power memory architecture to address the memory power wall. IEEE Transactions on Computers, 61(12):1697–1710, December 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Srinivasan:2010:ARC

[Sri10] Sudarshan K. Srinivasan. Automatic refinement checking of pipelines with outof-order execution. IEEE Transactions on Computers, 59(8):1138–1144, August 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5396327.

Shafique:2017:AGP

[SRK^{+}17] Muhammad Shafique, Semeen Rehman, Florian Kriebel, Muhammad Usman Karim Khan, Bruno Zatt, Arun Subramaniyan, Bruno Boessio Vizzotto, and Jörg Henkel. Applicationguided power-efficient fault tolerance for H.264 context adaptive variable length coding. IEEE Transactions on Computers, 66(4):560– 574, April 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shabani:2016:LAS

[SRR^{+}16] Hesam Shabani, Arman Roohi, Akram Reza, Midia Reshadi, Nader Bagherzadeh, and Ronald F. DeMara. Loss-aware switch design and non-blocking detection algorithm for intra-chip scale photonic interconnection networks. IEEE Transactions on Computers, 65 (6):1789–1801, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Swartzlander:2012:FIF

[SS12] Earl E. Swartzlander and Hani H. M. Saleh. FFT implementation with fused floating-point operations. IEEE Transactions on Computers, 61(2):284–288, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5669293.

Schuiki:2019:SNM

[SSGB19] Fabian Schuiki, Michael Schaffner, Frank K. Gürkaynak, and Luca Benini. A scalable near-memory architecture for training deep neural networks on large inmemory datasets. IEEE Transactions on Computers, 68(4):484–497, April 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8502059/.

Seol:2018:ERF

[SSJ^{+}18] Hoseok Seol, Wongyu Shin, Jaemin Jang, Jungwhan Choi, Hakseung Lee, and Lee-Sup Kim. Elaborate refresh: A fine granularity retention management for deep submicron DRAMs. IEEE Transactions on Computers, 67(10): 1403–1415, October 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8326524/.

Sakai:2016:NCS

[SSKL16] K. Sakai, M. Sun, W. Ku, and T. H. Lai. A novel coding scheme for secure communications in distributed RFID systems. IEEE Transactions on Computers, 65 (2):409–421, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sang:2012:ERD

[SST12] Yingpeng Sang, Hong Shen, and Hui Tian. Effective reconstruction of data perturbed by random projections. IEEE Transactions on Computers, 61(1):101– 117, January 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shang:2012:TET

[SSW12] Pengju Shang, Saba Sehrish, and Jun Wang. TRAID: Exploiting temporal redundancy and spatial redundancy to boost transaction processing systems performance. IEEE Transactions on Computers, 61(4):517– 529, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shukla:2011:GEI

[ST11a] Sandeep K. Shukla and Jean-Pierre Talpin. Guest Editors’ introduction: Special section on science of design for safety critical systems. IEEE Transactions on Computers, 60(8):1057– 1058, August 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Song:2011:TAF

[ST11b] Haoyu Song and Jonathan S.

Turner. Toward advocacyfree evaluation of packet classification algorithms. IEEE Transactions on Computers, 60(5):723–733, May 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Skoufis:2012:OFD

[ST12] Michael N. Skoufis and Spyros Tragoudas. An online failure detection method for data buses using multithreshold receiving logic. IEEE Transactions on Computers, 61(2):187–198, February 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sfikas:2016:TNC

[ST16] Yiorgos Sfikas and Yiorgos Tsiatouhas. Testing neighbouring cell leakage and transition induced faults in DRAMs. IEEE Transactions on Computers, 65 (7):2339–2345, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shu:2017:CST

[ST17] Wei Shu and Nian-Feng Tzeng. Compressed sharer tracking and relinquishment coherence for superior directory efficiency of chip multiprocessors. IEEE Transactions on Computers, 66(11): 1975–1981, November 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7911232/.

Sasaki:2018:TBD

[ST18a] Yu Sasaki and Yosuke Todo. Tight bounds of differentially and linearly active Sboxes and division property of Lilliput. IEEE Transactions on Computers, 67(5):717–732, May 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8126222/.

Shu:2018:NNU

[ST18b] Wei Shu and Nian-Feng Tzeng. NUDA: Non-uniform directory architecture for scalable chip multiprocessors. IEEE Transactions on Computers, 67(5):740– 747, May 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8107506/.

Stewart:2014:IND

[Ste14] I. A. Stewart. Interconnection networks of degree three obtained by pruning two-dimensional tori. IEEE Transactions on Computers, 63(10):2473–2486, October 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shahverdi:2017:LSC

[STE17] Aria Shahverdi, Mostafa Taha, and Thomas Eisenbarth. Lightweight side channel resistance: Threshold implementations of Simon. IEEE Transactions on Computers, 66(4):661– 671, April 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Soteriou:2016:HAT

[STK16] V. Soteriou, T. Theocharides, and E. Kakoulli. A holistic approach towards intelligent hotspot prevention in network-on-chip-based multicores. IEEE Transactions on Computers, 65(3):819– 833, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sheikh:2015:CFD

[STR15] R. Sheikh, J. Tuck, and E. Rotenberg. Control-flow decoupling: An approach for timely, non-speculative branching. IEEE Transactions on Computers, 64 (8):2182–2203, August 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sabarimuthu:2018:AMR

[SV18] Jasmine Madonna Sabarimuthu and T. G. Venkatesh. Analytical miss rate calculation of L2 cache from the RD profile of L1 cache. IEEE Transactions on Computers, 67(1):9–15, January 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7970187/.

Sridhar:2014:ICT

[SVAB14] A. Sridhar, A. Vincenzi, D. Atienza, and T. Brunschwiler. 3D-ICE: A compact thermal model for earlystage design of liquid-cooled ICs. IEEE Transactions on Computers, 63(10):2576– 2589, October 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Savino:2018:RCL

[SVD18] Alessandro Savino, Alessandro Vallero, and Stefano Di Carlo. ReDO: Crosslayer multi-objective designexploration framework for efficient soft error resilient systems. IEEE Transactions on Computers, 67(10): 1462–1477, October 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8323226/.

Shi:2019:LWW

[SWF^{+}19] Y. Shi, W. Wei, H. Fan, M. H. Au, and X. Luo. A light-weight white-box encryption scheme for securing distributed embedded devices. IEEE Transactions on Computers, 68(10): 1411–1427, October 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shoufan:2010:NCA

[SWM^{+}10] A. Shoufan, T. Wink, H. G. Molter, S. A. Huss, and E. Kohnert. A novel cryptoprocessor architecture for the McEliece public-key cryptosystem. IEEE Transactions on Computers, 59(11): 1533–1546, November 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5477413.

Sun:2011:NAP

[SWWC11] Hung-Min Sun, Hsun Wang, King-Hang Wang, and Chien-Ming Chen. A native APIs protection mechanism in the kernel mode against malicious code. IEEE Transactions on Computers, 60 (6):813–823, June 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shang:2011:NPI

[SWZG11] Pengju Shang, Jun Wang, Huijun Zhu, and Peng Gu. A new placement-ideal layout for multiway replication storage system. IEEE Transactions on Computers, 60 (8):1142–1156, August 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2015:ERE

[SWZG15] Y. Sun, T. Wu, G. Zhao, and M. Guizani. Efficient rule engine for smart building systems. IEEE Transactions on Computers, 64 (6):1658–1669, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2012:LCG

[SX12] Haiying Shen and Cheng-Zhong Xu. Leveraging a compound graph-based DHT for multi-attribute range queries with performance analysis. IEEE Transactions on Computers, 61(4):433– 447, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Song:2014:ARP

[SXCL14] Weijia Song, Zhen Xiao, Qi Chen, and Haipeng Luo. Adaptive resource provisioning for the cloud using online bin packing. IEEE Transactions on Computers, 63(11): 2647–2660, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Saifullah:2015:EEC

[SXLC15] A. Saifullah, You Xu, Chenyang Lu, and Yixin Chen. End-to-end communication delay analysis in industrial wireless networks. IEEE Transactions on Computers, 64(5):1361– 1374, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shafik:2018:RPC

[SYD18] Rishad Shafik, Alex Yakovlev, and Shidhartha Das. Realpower computing. IEEE Transactions on Computers, 67(10):1445–1461, October 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8330023/.

Son:2017:OOF

[SYH17] Yongseok Son, Heon Young Yeom, and Hyuck Han. Optimizing I/O operations in file systems for fast storage devices. IEEE Transactions on Computers, 66 (6):1071–1084, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7769226/.

Suei:2014:EAF

[SYK14] Pei-Lun Suei, Mi-Yen Yeh, and Tei-Wei Kuo. Enduranceaware flash-cache management for storage servers. IEEE Transactions on Computers, 63(10):2416–2430, October 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sui:2014:DAH

[SZDL14] Yan Sui, Xukai Zou, Eliza Y.

Du, and Feng Li. Design and analysis of a highly userfriendly, secure, privacypreserving, and revocable authentication method. IEEE Transactions on Computers, 63(4):902–916, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2018:ESM

[SZG^{+}18] Jiajie Shen, Kai Zhang, Jiazhen Gu, Yangfan Zhou, and Xin Wang. Efficient scheduling for multi-block updates in erasure coding based storage systems. IEEE Transactions on Computers, 67(4):573–581, April 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8094270/.

Sun:2016:SCB

[SZL^{+}16] Guangyu Sun, Chao Zhang, Peng Li, Tao Wang, and Yiran Chen. Statistical cache bypassing for non-volatile memory. IEEE Transactions on Computers, 65(11): 3427–3440, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Si:2014:GDR

[SZS14] Weisheng Si, Albert Y.

Zomaya, and Selvadurai Selvakennedy. A geometric deployment and routing scheme for directional wireless mesh networks. IEEE Transactions on Computers, 63(6):1323–1335, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Su:2016:SDP

[SZW^{+}16] Maomeng Su, Lei Zhang, Yongwei Wu, Kang Chen, and Keqin Li. Systematic data placement optimization in multi-cloud storage for complex requirements. IEEE Transactions on Computers, 65(6):1964–1977, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Song:2019:EEA

[SZZ^{+}19] Weining Song, Yang Zhou, Mengying Zhao, Lei Ju, Chun Jason Xue, and Zhiping Jia. EMC: Energyaware morphable cache design for non-volatile processors. IEEE Transactions on Computers, 68 (4):498–509, April 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8519610/.

Tarihi:2016:HNV

[TAH^{+}16] Mojtaba Tarihi, Hossein Asadi, Alireza Haghdoost, Mohammad Arjomand, and Hamid Sarbazi-Azad. A hybrid non-volatile cache design for solid-state drives using comprehensive I/O characterization. IEEE Transactions on Computers, 65 (6):1678–1691, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tsoumanis:2016:PEM

[TAM^{+}16] K. Tsoumanis, N. Axelos, N. Moschopoulos, G. Zervakis, and K. Pekmestzi. Pre-encoded multipliers based on non-redundant radix-4 signed-digit encoding. IEEE Transactions on Computers, 65(2):670–676, February 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2015:HSB

[TB15] Yi Tang and N. W. Bergmann.

A hardware scheduler based on task queues for FPGAbased embedded real-time systems. IEEE Transactions on Computers, 64 (5):1254–1267, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tsai:2017:ASR

[TBC^{+}17] Wei-Yu Tsai, Davis R.

Barch, Andrew S. Cassidy, Michael V. DeBole, Alexander Andreopoulos, Bryan L. Jackson, Myron D. Flickner, John V. Arthur, Dharmendra S. Modha, John Sampson, and Vijaykrishnan Narayanan. Alwayson speech recognition using TrueNorth, a reconfigurable, neurosynaptic processor. IEEE Transactions on Computers, 66 (6):996–1007, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7750640/.

Tian:2014:SDO

[TC14] Xiaohua Tian and Yu Cheng. A scalable destination-oriented multicast protocol with incremental deployability. IEEE Transactions on Computers, 63(4):793–806, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tay:2016:NIM

[TC16] T. Fatt Tay and C. Chang. A non-iterative multiple residue digit error detection and correction algorithm in RRNS. IEEE Transactions on Computers, 65 (2):396–408, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tsai:2018:STC

[TCHL18] Ting-Hao Tsai, Ya-Shu Chen, Xue-Xin He, and Cheng-Yu Li. STEM:

A thermal-constrained realtime scheduling for 3D heterogeneous-ISA multicore processors. IEEE Transactions on Computers, 67 (6):874–889, June 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8214221/.

Tran:2018:SIS

[TCK^{+}18] Kim-Anh Tran, Trevor E.

Carlson, Konstantinos Koukos, Magnus Själander, Vasileios Spiliopoulos, Stefanos Kaxiras, and Alexandra Jimborean. Static instruction scheduling for high performance on limited hardware. IEEE Transactions on Computers, 67(4):513–527, April 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8094900/.

Tsao:2015:EVB

[TCYH15] C. Tsao, Y. Chang, M. Yang, and P. Huang. Efficient victim block selection for flash storage devices. IEEE Transactions on Computers, 64(12):3444–3460, December 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tsai:2016:TSE

[TFCY16] T. H. Tsai, L. F. Fan, Y. S. Chen, and T. S. Yao. Triple speed: Energy-aware real-time task synchronization in homogeneous multicore systems. IEEE Transactions on Computers, 65 (4):1297–1309, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tsen:2011:HDB

[TGNSC11] Charles Tsen, Sonia Gonzalez-Navarro, Michael J. Schulte, and Katherine Compton. Hardware designs for binary integer decimal-based rounding. IEEE Transactions on Computers, 60(5): 614–627, May 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tavli:2011:EER

[TH11] Bulent Tavli and Wendi B. Heinzelman. Energy-efficient real-time multicast routing in mobile ad hoc networks. IEEE Transactions on Computers, 60(5):707– 722, May 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ttofis:2013:EDH

[THGT13] C. Ttofis, S. Hadjitheophanous, A. S. Georghiades, and T. Theocharides. Edgedirected hardware architecture for real-time disparity map computation. IEEE Transactions on Computers, 62(4):690–704, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tran:2014:QBS

[THM^{+}14] Hai Anh Tran, S. Hoceini, A. Mellouk, J. Perez, and S. Zeadally. QoE-based server selection for content distribution networks. IEEE Transactions on Computers, 63(11):2803–2815, November 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Thomasian:2012:CDH

[Tho12] Alexander Thomasian. Comment on “DACO: A High Performance Disk Architecture”. IEEE Transactions on Computers, 61(4):588– 590, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [LS10a].

Thornton:2015:SIU

[Tho15] M. A. Thornton. Simulation and implication using a transfer function model for switching logic. IEEE Transactions on Computers, 64(12):3580–3590, December 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Todo:2018:CAN

[TIHM18] Yosuke Todo, Takanori Isobe, Yonglin Hao, and Willi Meier. Cube attacks on non-blackbox polynomials based on division property. IEEE Transactions on Computers, 67(12): 1720–1736, December 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8357953/.

Tan:2015:DDS

[TJH^{+}15] Zhiyuan Tan, A. Jamdagni, Xiangjian He, P. Nanda, Ren Ping Liu, and Jiankun Hu. Detection of denialof-service attacks based on computer vision techniques. IEEE Transactions on Computers, 64(9): 2519–2533, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2017:NNI

[TJX^{+}17] Guoming Tang, Weixiang Jiang, Zhifeng Xu, Fangming Liu, and Kui Wu. NIPD: Non-intrusive power disaggregation in legacy datacenters. IEEE Transactions on Computers, 66(2): 312–325, February 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tziritas:2014:SGA

[TKL^{+}14] N. Tziritas, S. U. Khan, T. Loukopoulos, S. Lalis, Cheng-Zhong Xu, and P. Lampsas. Single and group agent migration: Algorithms, bounds, and optimality issues. IEEE Transactions on Computers, 63(12): 3143–3161, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ttofis:2016:LCR

[TKT16] Christos Ttofis, Christos Kyrkou, and Theocharis Theocharides. A low-cost real-time embedded stereo vision system for accurate disparity estimation based on guided image filtering. IEEE Transactions on Computers, 65(9): 2678–2693, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Trefzer:2017:HSE

[TLB^{+}17] Martin A. Trefzer, David M. R. Lawson, Simon J. Bale, James A. Walker, and Andy M. Tyrrell. Hierarchical strategies for efficient fault recovery on the reconfigurable PAnDA device. IEEE Transactions on Computers, 66(6):930–945, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7756359/.

Turakhia:2017:TPE

[TLGM17] Yatish Turakhia, Guangshuo Liu, Siddharth Garg, and Diana Marculescu. Thread progress equalization: Dynamically adaptive powerconstrained performance optimization of multi-threaded applications. IEEE Transactions on Computers, 66 (4):731–744, April 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2016:ASS

[TLH^{+}16] Shaohua Tang, Xiaoyu Li, Xinyi Huang, Yang Xiang, and Lingling Xu. Achieving simple, secure and efficient hierarchical access control in cloud computing. IEEE Transactions on Computers, 65(7):2325–2331, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tsai:2012:SRT

[TLL12] Wei-Lin Tsai, Wei-Chih Liu, and James Chien-Mo Li. Structural reduction techniques for logic-chain bridging fault diagnosis. IEEE Transactions on Computers, 61(7):928–938, July 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2013:AXP

[TLL^{+}13] Jie Tang, Shaoshan Liu, Chen Liu, Zhimin Gu, and Jean-Luc Gaudiot. Acceleration of XML parsing through prefetching. IEEE Transactions on Computers, 62 (8):1616–1628, August 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2017:RTM

[TLP17] Minghua Tang, Xiaola Lin, and Maurizio Palesi. The repetitive turn model for adaptive routing. IEEE Transactions on Computers, 66(1):138–146, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2018:SRA

[TLP18] Minghua Tang, Jing Lin, and Maurizio Palesi. The suboptimal routing algorithm for 2D mesh network. IEEE Transactions on Computers, 67(5):704–716, May 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8118167/.

Tang:2011:NSD

[TLZV11] Xiaoyong Tang, Kenli Li, Zeng Zeng, and Bharadwaj Veeravalli. A novel securitydriven scheduling algorithm for precedence-constrained tasks in heterogeneous distributed systems. IEEE Transactions on Computers, 60(7):1017–1029, July 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tsoutsos:2018:EDM

[TM18] Nektarios Georgios Tsoutsos and Michail Maniatakos. Efficient detection for malicious and random errors in additive encrypted computation. IEEE Transactions on Computers, 67 (1):16–31, January 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7967774/.

Take:2014:NIC

[TMS^{+}14] Yasuhiro Take, Hiroki Matsutani, Daisuke Sasaki, Michihiro Koibuchi, Tadahiro Kuroda, and Hideharu Amano. 3D NoC with inductive-coupling links for building-block SiPs. IEEE Transactions on Computers, 63(3):748–763, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Turcu:2016:OND

[TPR16] Alexandru Turcu, Roberto Palmieri, and Binoy Ravindran. On open nesting in distributed transactional memory. IEEE Transactions on Computers, 65 (6):1856–1868, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tomes:2018:TAP

[TRA18] Erica Tomes, Everett Neil Rush, and Nihat Altiparmak. Towards adaptive parallel storage systems. IEEE Transactions on Computers, 67(12):1840–1848, December 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8359084/.

Tumer:2011:IDS

[TS11] Irem Y. Tumer and Carol S.

Smidts. Integrated designstage failure analysis of software-driven hardware systems. IEEE Transactions on Computers, 60(8): 1072–1084, August 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tsai:2013:QPD

[Tsa13]  Chang-Hsiung Tsai.  A

quick pessimistic diagnosis algorithm for hypercube-like multiprocessor systems under the PMC model. IEEE Transactions on Computers, 62(2):259–267, February 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Testa:2019:MMB

[TSA^{+}19] E. Testa, M. Soeken, L. G. Amarù, W. Haaswijk, and G. De Micheli. Mapping monotone Boolean functions into majority. IEEE Transactions on Computers, 68(5): 791–797, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tse:2012:OBB

[Tse12]  Savio S. H. Tse.  On-

line bounds on balancing two independent criteria with replication and reallocation. IEEE Transactions on Computers, 61(11): 1601–1610, November 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Teymoori:2016:FFC

[TSK16] Peyman Teymoori, Khosrow Sohraby, and Kiseon Kim. Fair flow control and fairness evaluation in computer networks and systems. IEEE Transactions on Computers, 65(7):2090–2103, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2010:LHO

[TW10] Xiangyu Tang and Seongmoon Wang. A low hardware overhead selfdiagnosis technique using Reed–Solomon codes for self-repairing chips. IEEE Transactions on Computers, 59(10):1309–1319, October 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5374371.

Tian:2011:EMI

[TWTT11] Xin Tian, Lin Wu, Yi-Hua Tan, and Jin-Wen Tian. Efficient multi-input/multi-output VLSI architecture for two-dimensional liftingbased discrete wavelet transform. IEEE Transactions on Computers, 60(8):1207– 1211, August 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Trost:2016:OPC

[TX16] William R. Trost and Guangwu Xu. On the optimal pre-computation of window NAF for Koblitz curves. IEEE Transactions on Computers, 65(9): 2918–2924, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2011:STG

[TXL11] Yong Tang, Bin Xiao, and Xicheng Lu. Signature tree generation for polymorphic worms. IEEE Transactions on Computers, 60(4):565– 579, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2010:IST

[TYWC10] Shyue-Ming Tang, Jinn-Shyong Yang, Yue-Li Wang, and Jou-Ming Chang. Independent spanning trees on multidimensional torus networks. IEEE Transactions on Computers, 59 (1):93–102, January 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5184805.

Tang:2016:NOO

[TYY^{+}16] B. Tang, S. Yang, B. Ye, S. Guo, and S. Lu. Nearoptimal one-sided scheduling for coded segmented network coding. IEEE Transactions on Computers, 65 (3):929–939, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Teng:2014:SLS

[TZL^{+}14] Jin Teng, Boying Zhang, Xinfeng Li, Xiaole Bai, and Dong Xuan. E-Shadow: Lubricating social interaction using mobile phones. IEEE Transactions on Computers, 63(6):1422–1433, June 2014.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ugurdag:2017:HDS

[UdDG^{+}17] H. Fatih Ugurdag, Florent de Dinechin, Y. Serhan Gener, Sezer Gören, and Laurent-Stéphane Didier. Hardware division by small integer constants. IEEE Transactions on Computers, 66(12):2097–2110, December 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7933010/.

Ueno:2017:FAV

[UHSA17] Rei Ueno, Naofumi Homma, Yukihiro Sugawara, and Takafumi Aoki. Formal approach for verifying Galois field arithmetic circuits of higher degrees. IEEE Transactions on Computers, 66 (3):431–442, March 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Uchigaito:2018:EDA

[UMN18] Hiroshi Uchigaito, Seiji Miura, and Takumi Nito. Efficient data-allocation scheme for eliminating garbage collection during analysis of big graphs stored in NAND flash memory. IEEE Transactions on Computers, 67(5): 646–657, May 2018. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8118115/.

Ueno:2019:TBP

[USH19] R. Ueno, M. Suzuki, and N. Homma. Tackling biased PUFs through biased masking: a debiasing method for efficient fuzzy extractor. IEEE Transactions on Computers, 68(7):1091– 1104, July 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ubal:2013:HBG

[USP^{+}13] Rafael Ubal, Julio Sahuquillo, Salvador Petit, Pedro Lopez, and Jose Duato. Hardwarebased generation of independent subtraces of instructions in clustered processors. IEEE Transactions on Computers, 62(5):944– 955, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Upasani:2016:CAW

[UVG16] G. Upasani, X. Vera, and A. Gonzalez. A case for acoustic wave detectors for soft-errors. IEEE Transactions on Computers, 65(1): 5–18, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Unnikrishnan:2013:RDP

[UVL^{+}13] Deepak Unnikrishnan, Ramakrishna Vadlamani, Yong Liao, Jeremie Crenne, Lixin Gao, and Russell Tessier. Reconfigurable data planes for scalable network virtualization. IEEE Transactions on Computers, 62(12): 2476–2488, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vemu:2011:CCF

[VA11] Ramtilak Vemu and Jacob Abraham. CEDA: Controlflow error detection using assertions. IEEE Transactions on Computers, 60(9): 1233–1245, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5871589.

Vazquez:2014:FRM

[VAB14] Alvaro Vazquez, Elisardo Antelo, and Javier D. Bruguera. Fast radix-10 multiplication using redundant BCD codes. IEEE Transactions on Computers, 63 (8):1902–1914, August 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Venkatachalam:2019:DAA

[VALK19] S. Venkatachalam, E. Adams, H. J. Lee, and S. Ko. Design and analysis of area and power efficient approximate Booth multipliers. IEEE Transactions on Computers, 68(11):1697–1703, November 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vazquez:2010:IDH

[VAM10] A. Vazquez, E. Antelo, and P. Montuschi. Improved design of high-performance parallel decimal multipliers. IEEE Transactions on Computers, 59(5):679–693, May 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5313798.

Villegas:2018:LHT

[VAN^{+}18] Alejandro Villegas, Rafael Asenjo, Angeles Navarro, Oscar Plata, and David Kaeli. Lightweight hardware transactional memory for GPU scratchpad memory. IEEE Transactions on Computers, 67(6): 816–829, June 2018. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8119530/.

Vazquez:2013:IAA

[VB13] Alvaro Vazquez and Javier D. Bruguera. Iterative algorithm and architecture for exponential, logarithm, powering, and root extraction. IEEE Transactions on Computers, 62(9): 1721–1731, September 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Villas:2013:DLR

[VBR^{+}13] L. A. Villas, A. Boukerche, H. S. Ramos, H. A. B. F. de Oliveira, R. B. de Araujo, and A. A. F. Loureiro. DRINA: A lightweight and reliable routing approach for in-network aggregation in wireless sensor networks. IEEE Transactions on Computers, 62(4):676–689, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Valente:2010:HTD

[VC10] P. Valente and F. Checconi. High throughput disk scheduling with fair bandwidth distribution. IEEE Transactions on Computers, 59(9):1172–1186, September 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5467047.

Veronese:2013:EBF

[VCB^{+}13] Giuliana Santos Veronese, Miguel Correia, Alysson Neves Bessani, Lau Cheuk Lung, and Paulo Verissimo. Efficient Byzantine faulttolerance. IEEE Transactions on Computers, 62 (1):16–30, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vujic:2012:DFD

[VCG^{+}12] Nikola Vujic, Felipe Cabarcas, Marc Gonzalez, Alex Ramirez, Xavier Martorell, and Eduard Ayguade. DMA++: On the fly data realignment for on-chip memories. IEEE Transactions on Computers, 61(2):237–250, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Valero:2019:AAG

[VCSG^{+}19] Alejandro Valero, Francisco Candel, Darío Suárez-Gracia, Salvador Petit, and Julio Sahuquillo. An agingaware GPU register file design based on data redundancy. IEEE Transactions on Computers, 68 (1):4–20, January 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8395355/.

Vergos:2012:MAD

[VD12] Haridimos T. Vergos and Giorgos Dimitrakopoulos. On modulo 2^{n} + 1 adder design. IEEE Transactions on Computers, 61(2):173– 186, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

vandenBraak:2016:CXH

[vdBGLGL^{+}16] Gert-Jan van den Braak, Juan Gómez-Luna, José María González-Linares, Henk Corporaal, and Nicolás Guil. Configurable XOR hash functions for banked scratchpad memories in GPUs. IEEE Transactions on Computers, 65(7):2045–2058, July 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

VanWambeke:2013:AMA

[VECD13] Nicolas Van Wambeke, Ernesto Exposito, Christophe Chassot, and Michel Diaz. ATP: A microprotocol approach to autonomic communication. IEEE Transactions on Computers, 62(11):2131–2140, November 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

VandenSteen:2016:APP

[VED^{+}16] Sam Van den Steen, Stijn Eyerman, Sander De Pestel, Moncef Mechri, Trevor E. Carlson, David Black-Schaffer, Erik Hagersten, and Lieven Eeckhout. Analytical processor performance and power modeling using microarchitecture independent characteristics. IEEE Transactions on Computers, 65 (12):3537–3551, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vinco:2016:CMV

[VGF16] Sara Vinco, Valerio Guarnieri, and Franco Fummi. Code manipulation for virtual platform integration. IEEE Transactions on Computers, 65(9):2694–2708, September 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Volkova:2019:THI

[VIDH19] Anastasia Volkova, Matei Istoan, Florent De Dinechin, and Thibault Hilaire. Towards hardware IIR filters computing just right: Direct form I case study. IEEE Transactions on Computers, 68(4):597–608, April 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8521688/.

Voyiatzis:2015:GSP

[VK15] I. K. Voyiatzis and D. J. Kavvadias. On the generation of SIC pairs in optimal time. IEEE Transactions on Computers, 64(10): 2891–2901, October 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Venkatesan:2016:CDD

[VKS^{+}16] R. Venkatesan, V. J. Kozhikkottu, M. Sharad, C. Augustine, A. Raychowdhury, K. Roy, and A. Raghunathan. Cache design with domain wall memory. IEEE Transactions on Computers, 65 (4):1010–1024, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vogel:2019:ESV

[VMB19] Pirmin Vogel, Andrea Marongiu, and Luca Benini. Exploring shared virtual memory for FPGA accelerators with a configurable IOMMU. IEEE Transactions on Computers, 68(4):510–525, April 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8519631/.

Villalba-Moreno:2018:URH

[VMHGN18] Julio Villalba-Moreno, Javier Hormigo, and Sonia González-Navarro. Unbiased rounding for HUB floating-point addition. IEEE Transactions on Computers, 67(9): 1359–1365, September 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8300633/.

Veeravalli:2014:GEI

[VP14] Bharadwaj Veeravalli and Manish Parashar. Guest Editors’ introduction: Special issue on cloud of clouds. IEEE Transactions on Computers, 63(1):1–2, January 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Valero:2012:DPE

[VPS^{+}12] Alejandro Valero, Salvador Petit, Julio Sahuquillo, Pedro Lopez, and Jose Duato. Design, performance, and energy consumption of eDRAM/SRAM macrocells for L1 data caches. IEEE Transactions on Computers, 61(9):1231–1242, September 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vallero:2019:SES

[VSC^{+}19] A. Vallero, A. Savino, A. Chatzidimitriou, M. Kaliorakis, M. Kooli, M. Riera, M. Anglada, G. Di Natale, A. Bosio, R. Canal, A. Gonzalez, D. Gizopoulos, R. Mariani, and S. Di Carlo. SyRA: Early system reliability analysis for crosslayer soft errors resilience in memory arrays of microprocessor systems. IEEE Transactions on Computers, 68(5):765–783, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Viegas:2017:TEE

[VSF^{+}17] Eduardo Viegas, Altair O.

Santin, André França, Ricardo Jasinski, Volnei A. Pedroni, and Luiz S. Oliveira. Towards an energy-efficient anomaly-based intrusion detection engine for embedded systems. IEEE Transactions on Computers, 66 (1):163–177, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Valero:2015:DHS

[VSLD15] A. Valero, J. Sahuquillo, P. Lopez, and J. Duato. Design of hybrid secondlevel caches. IEEE Transactions on Computers, 64 (7):1884–1897, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Valadimas:2016:TET

[VTA16] S. Valadimas, Y. Tsiatouhas, and A. Arapoyanni. Timing error tolerance in small core designs for SoC applications. IEEE Transactions on Computers, 65 (2):654–663, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Venkatesan:2016:SCA

[VTW16] Vimalraj Venkatesan, Y. C. Tay, and Qingsong Wei. Sizing cleancache allocation for virtual machines’ transcendent memory. IEEE Transactions on Computers, 65 (6):1949–1963, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vazquez:2012:RFP

[VVMAZ12] Alvaro Vazquez, Julio Villalba-Moreno, Elisardo Antelo, and Emilio L. Zapata. Redundant floating-point decimal CORDIC algorithm. IEEE Transactions on Computers, 61(11):1551–1562, November 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Verbeek:2017:DVC

[VYEB17] Freek Verbeek, Pooria M. Yaghini, Ashkan Eghbal, and Nader Bagherzadeh. Deadlock verification of cache coherence protocols and communication fabrics. IEEE Transactions on Computers, 66(2):272–284, February 2017. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Verbeek:2018:CAV

[VYEB18] Freek Verbeek, Pooria M. Yaghini, Ashkan Eghbal, and Nader Bagherzadeh. A compositional approach for verifying protocols running on on-chip networks. IEEE Transactions on Computers, 67(7):905–919, July 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8239635/.

Whitham:2010:TPO

[WA10] J. Whitham and N. Audsley. Time-predictable out-of-order execution for hard realtime systems. IEEE Transactions on Computers, 59(9): 1210–1223, September 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5467051.

Wu:2017:AAD

[WAK^{+}17] Di Wu, Dmitri I. Arkhipov, Minyoung Kim, Carolyn L. Talcott, Amelia C. Regan, Julie A. McCann, and Nalini Venkatasubramanian. ADDSEN: Adaptive data processing and dissemination for drone swarms in urban sensing. IEEE Transactions on Computers, 66 (2):183–198, February 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Walther:2019:VNR

[Wal19] Christoph Walther. Verified Newton–Raphson iteration for multiplicative inverses modulo powers of any base. ACM Transactions on Mathematical Software, 45(1):9:1–9:7, March 2019. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3301317.  See

[Dum14].

Wang:2019:CAR

[WBGI19] Shibo Wang, Mahdi Nazm Bojnordi, Xiaochen Guo, and Engin Ipek. Content aware refresh: Exploiting the asymmetry of DRAM retention errors to reduce the refresh frequency of less vulnerable data. IEEE Transactions on Computers, 68 (3):362–374, March 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8456584/.

Wang:2015:FFT

[WBZ^{+}15] Ji Wang, Weidong Bao, Xiaomin Zhu, L. T. Yang, and Yang Xiang. FES-TAL: Fault-tolerant elastic scheduling algorithm for real-time tasks in virtualized clouds. IEEE Transactions on Computers, 64(9): 2545–2558, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2015:VAO

[WCH^{+}15] J. Wang, X. Chen, X. Huang, I. You, and Y. Xiang. Verifiable auditing for outsourced database in cloud computing. IEEE Transactions on Computers, 64(11): 3293–3303, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wei:2018:GCQ

[WCL^{+}18] Chun-Yan Wei, Xiao-Qiu Cai, Bin Liu, Tian-Yin Wang, and Fei Gao. A generic construction of quantumoblivious-key-transfer-based private query with ideal database security and zero failure. IEEE Transactions on Computers, 67(1): 2–8, January 2018. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7962191/.

Wang:2016:MCP

[WCLY16] B. Wang, J. Chen, W. Liu, and L. T. Yang. Minimum cost placement of bistatic radar sensors for belt barrier coverage. IEEE Transactions on Computers, 65 (2):577–588, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2016:TEE

[WCM^{+}16] J. Wang, Z. Cao, X. Mao, X. Li, and Y. Liu. Towards energy efficient dutycycled networks: Analysis, implications and improvement. IEEE Transactions on Computers, 65 (1):270–280, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2013:PPP

[WCW^{+}13] Cong Wang, Sherman S. M. Chow, Qian Wang, Kui Ren, and Wenjing Lou. Privacypreserving public auditing for secure cloud storage. IEEE Transactions on Computers, 62(2):362–375, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2012:HPM

[WDSP12] Guiming Wu, Yong Dou, Junqing Sun, and Gregory D. Peterson. A high performance and memory efficient LU decomposer on FPGAs. IEEE Transactions on Computers, 61(3):366– 378, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2012:RCC

[WE12] Dong Wang and Milos D. Ercegovac. A radix-16 combined complex division/square root unit with operand prescaling. IEEE Transactions on Computers, 61(9):1243–1255, September 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2019:EDT

[WEH^{+}19] Junshi Wang, Masoumeh Ebrahimi, Letian Huang, Xuan Xie, Qiang Li, Guangjun Li, and Axel Jantsch. Efficient design-for-test approach for networks-on-chip. IEEE Transactions on Computers, 68(2):198– 213, February 2019. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8440731/.

Wang:2014:CFA

[WEX14] Dong Wang, Miloš D. Ercegovac, and Yang Xiao. Complex function approximation using two-dimensional interpolation. IEEE Transactions on Computers, 63(12): 2948–2960, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2012:EMB

[WF12] Zhen Wang and Shuqin Fan. Efficient Montgomerybased semi-systolic multiplier for even-type GNB of GF(2^{m}). IEEE Transactions on Computers, 61 (3):415–419, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wei:2014:GTA

[WF14] Hongchuan Wei and S. Ferrari. A geometric transversals approach to analyzing the probability of track detection for maneuvering targets. IEEE Transactions on Computers, 63(11): 2633–2646, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wahba:2017:AEF

[WF17] Ahmed A. Wahba and Hossam A. H. Fahmy. Area efficient and fast combined binary/decimal floating point fused multiply add unit. IEEE Transactions on Computers, 66 (2):226–239, February 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wei:2019:NRS

[WFT^{+}19] X. Wei, D. Feng, W. Tong, J. LIU, and L. Ye. NICO: Reducing software-transparent crash consistency cost for persistent memory. IEEE Transactions on Computers, 68(9):1313–1324, September 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2017:PEH

[WFY^{+}17] Fenggang Wu, Ziqi Fan, Ming-Chang Yang, Baoquan Zhang, Xiongzi Ge, and David H. C. Du. Performance evaluation of host aware shingled magnetic recording (HA–SMR) drives. IEEE Transactions on Computers, 66(11): 1932–1945, November 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7942044/.

Weng:2013:HCM

[WGLL13] Chuliang Weng, Minyi Guo, Yuan Luo, and Minglu Li. Hybrid CPU management for adapting to the diversity of virtual machines. IEEE Transactions on Computers, 62(7):1332–1344, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2014:NSP

[WGR^{+}14] Yongwei Wu, Weichao Guo, Jinglei Ren, Xun Zhao, and Weimin Zheng. NO2: Speeding up parallel processing of massive compute-intensive tasks. IEEE Transactions on Computers, 63(10):2487– 2499, October 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2015:SSE

[WGW^{+}15] A. Wang, M. Gaudet, Peng Wu, M. Ohmacht, J. N. Amaral, C. Barton, R. Silvera, and M. M. Michael. Software support and evaluation of hardware transactional memory on Blue Gene/Q. IEEE Transactions on Computers, 64 (1):233–246, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wei:2015:ESW

[WGZ^{+}15] Kaimin Wei, Song Guo, Deze Zeng, Ke Xu, and Keqiu Li. Exploiting small world properties for message forwarding in delay tolerant networks. IEEE Transactions on Computers, 64(10): 2809–2818, October 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2016:CPS

[WHBR16] L. Wang, S. Hu, G. Betis, and R. Ranjan. A computing perspective on Smart City [guest editorial]. IEEE Transactions on Computers, 65(5):1337–1338, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2015:EFF

[WHC^{+}15a] Wei Wang, Yin Hu, Lianmu Chen, Xinming Huang, and B. Sunar. Exploring the feasibility of fully homomorphic encryption. IEEE Transactions on Computers, 64(3):698–706, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wen:2015:STE

[WHC^{+}15b] Sheng Wen, M. S. Haghighi, Chao Chen, Yang Xiang, Wanlei Zhou, and Weijia Jia. A sword with two edges: Propagation studies on both positive and negative information in online social networks. IEEE Transactions on Computers, 64(3):640–653, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wen:2012:NNU

[WhCCC12] Chi-Neng Wen, Shu hsuan Chou, Chien-Chih Chen, and Tien-Fu Chen. NUDA: a non-uniform debugging architecture and nonintrusive race detection for manycore systems. IEEE Transactions on Computers, 61 (2):199–212, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2012:NSI

[WHL^{+}12] Yue-Li Wang, Cheng-Ju Hsu, Jia-Jie Liu, Ming-Tat Ko, and Fu-Hsing Wang. A new subclass of integer linear programming problems and its applications. IEEE Transactions on Computers, 61(12):1813–1822, December 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wei:2017:PAA

[WHL17] Hengfeng Wei, Yu Huang, and Jian Lu. Probabilisticallyatomic 2-atomicity: Enabling almost strong consistency in distributed storage systems. IEEE Transactions on Computers, 66 (3):502–514, March 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2016:UMC

[WHYS16] Rui Wang, William N. N. Hung, Guowu Yang, and Xiaoyu Song. Uncertainty model for configurable hardware/software and resource partitioning. IEEE Transactions on Computers, 65(10): 3217–3223, October 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2015:EAB

[WHZ^{+}15] Z. Wang, D. Huang, Y. Zhu, B. Li, and C. Chung. Efficient attribute-based comparable data access control. IEEE Transactions on Computers, 64(12):3430–3443, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2011:IAR

[WJF^{+}11] Suzhen Wu, Hong Jiang, Dan Feng, Lei Tian, and Bo Mao. Improving availability of RAID-structured storage systems by workload outsourcing. IEEE Transactions on Computers, 60(1):64–79, January 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2012:SLM

[WJL^{+}12] Cheng Wang, Changjun Jiang, Xiang-Yang Li, Shaojie Tang, Yuan He, Xufei Mao, and Yunhao Liu. Scaling laws of multicast capacity for power-constrained wireless networks under Gaussian channel model. IEEE Transactions on Computers, 61(5):713–725, May 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2014:ACW

[WJL^{+}14] Cheng Wang, Changjun Jiang, Yunhao Liu, Xiang-Yang Li, and Shaojie Tang. Aggregation capacity of wireless sensor networks:

Extended network case. IEEE Transactions on Computers, 63(6):1351–1364, June 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2015:PDM

[WJM15] Suzhen Wu, Hong Jiang, and Bo Mao. Proactive data migration for improved storage availability in large-scale data centers. IEEE Transactions on Computers, 64(9): 2637–2651, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Woodruff:2019:CCP

[WJX^{+}19] J. Woodruff, A. Joannou, H. Xia, A. Fox, R. M. Norton, D. Chisnall, B. Davis, K. Gudka, N. W. Filardo, A. T. Markettos, M. Roe, P. G. Neumann, R. N. M. Watson, and S. W. Moore. CHERI Concentrate: Practical compressed capabilities. IEEE Transactions on Computers, 68(10):1455– 1469, October 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2017:HNA

[WJY^{+}17] Xiaohang Wang, Yingtao Jiang, Mei Yang, Hong Li, and Terrence Mak. HRC: A 3D NoC architecture with genuine support for runtime thermal-aware task management. IEEE Transactions on Computers, 66(10): 1676–1688, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7914637/.

Wang:2016:DRS

[WKB16] Zhen Wang, Mark Karpovsky, and Lake Bu. Design of reliable and secure devices realizing Shamir’s secret sharing. IEEE Transactions on Computers, 65 (8):2443–2455, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2013:MAS

[WL13] Yang Wang and Paul Lu.

Maximizing active storage resources with deadlock avoidance in workflowbased computations. IEEE Transactions on Computers, 62(11):2210–2223, November 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2015:BBM

[WLC^{+}15] Jiantao Wang, Kam-Yiu Lam, Yuan-Hao Chang, Jen-Wei Hsieh, and Po-Chun Huang. Block-based multiversion B-tree for flashbased embedded database systems. IEEE Transactions on Computers, 64(4):925– 940, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2019:EMC

[WLG^{+}19] Yi Wu, You Li, Xiangxuan Ge, Yuan Gao, and Weikang Qian. An efficient method for calculating the error statistics of block-based approximate adders. IEEE Transactions on Computers, 68 (1):21–38, January 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8419308/.

Wang:2016:SEP

[WLJ^{+}16] H. Wang, Y. Li, D. Jin, P. Hui, and J. Wu. Saving energy in partially deployed software defined networks. IEEE Transactions on Computers, 65(5):1578– 1592, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2015:WWA

[WLK15] Yang Wang, P. Lu, and K. B. Kent. WaFS: A workflowaware file system for effective storage utilization in the cloud. IEEE Transactions on Computers, 64(9): 2716–2729, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Weng:2016:CMV

[WLLZ16] Chuliang Weng, Qian Liu, Kenli Li, and Deqing Zou. CloudMon: Monitoring virtual machines in clouds. IEEE Transactions on Computers, 65(12):3787–3793, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Won:2015:MMC

[WLM15] Youjip Won, Kyeongyeol Lim, and Jaehong Min. MUCH: Multithreaded contentbased file chunking. IEEE Transactions on Computers, 64(5):1375–1388, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2013:ORI

[WLQS13] Yi Wang, Duo Liu, Zhiwei Qin, and Zili Shao. Optimally removing intercore communication overhead for streaming applications on MPSoCs. IEEE Transactions on Computers, 62 (2):336–350, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2018:HMR

[WLS18] Chun-Wei Wu, Kuen-Jong Lee, and Alan P. Su. A hybrid multicast routing approach with enhanced methods for mesh-based networks-on-chip. IEEE Transactions on Computers, 67(9):1231–1245, September 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8309347/.

Wang:2016:DEO

[WLT^{+}16] Y. Wang, Z. Li, G. Tyson, S. Uhlig, and G. Xie. Design and evaluation of the optimal cache allocation for contentcentric networking. IEEE Transactions on Computers, 65(1):95–107, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2014:DPA

[WLV^{+}14] Danyao Wang, Charles Lo, Jasmina Vasiljevic, Natalie Enright Jerger, and J. Gregory Steffan. DART: A programmable architecture for NoC simulation on FPGAs. IEEE Transactions on Computers, 63 (3):664–678, March 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

We:2014:HPN

[WLY^{+}14] Kyoung-Soo We, Chang-Gun Lee, Kyongsu Yi, Yun Sang Lee, and Kwei-Jay Lin. HRT-PLRU: A new paging scheme for executing hard real-time programs on NAND flash memory. IEEE Transactions on Computers, 63(4):927–940, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2016:MDG

[WLYY16] Cong Wang, Ji Li, Fan Ye, and Yuanyuan Yang. A mobile data gathering framework for wireless rechargeable sensor networks with vehicle movement costs and capacity constraints. IEEE Transactions on Computers, 65(8):2411–2427, August 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2010:GFP

[WLZ10] Jianjia Wu, Jyh-Charn Liu, and Wei Zhao. A general framework for parameterized schedulability bound analysis of real-time systems. IEEE Transactions on Computers, 59(6):776–783, June 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5406502.

Wang:2015:AST

[WLZ^{+}15] Chao Wang, Xi Li, Junneng Zhang, Peng Chen, Yunji Chen, Xuehai Zhou, and R. C. C. Cheung. Architecture support for task outof-order execution in MP-SoCs. IEEE Transactions on Computers, 64(5):1296– 1310, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wild:2018:GGF

[WMG18] Alexander Wild, Amir Moradi, and Tim Güneysu. GliFreD: Glitch-free duplication towards power-equalized circuits on FPGAs. IEEE Transactions on Computers, 67(3):375–387, March 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7827086/.

Wang:2012:CLC

[WMW12] Xiaorui Wang, Kai Ma, and Yefu Wang. Cache latency control for application fairness or differentiation in power-constrained chip multiprocessors. IEEE Transactions on Computers, 61(10): 1371–1385, October 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2017:CBS

[WNCH17] Yi Wu, Ge Nong, Wai Hong Chan, and Ling Bo Han. Checking big suffix and LCP arrays by probabilistic methods. IEEE Transactions on Computers, 66(10): 1667–1675, October 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7922513/.

Wei:2016:NST

[WNKL16] W. Wei, K. Namba, Y. Kim, and F. Lombardi. A novel scheme for tolerating single event/multiple bit upsets (SEU/MBU) in nonvolatile memories. IEEE Transactions on Computers, 65(3):781–790, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2016:MFR

[WP16] Yanzhi Wang and Massoud Pedram. Model-free reinforcement learning and Bayesian classification in system-level power management. IEEE Transactions on Computers, 65(12): 3713–3726, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2016:CBE

[WQZ^{+}16] Q. Wu, B. Qin, L. Zhang, J. Domingo-Ferrer, O. Farras, and J. A. Manjon. Contributory broadcast encryption with efficient encryption and short ciphertexts. IEEE Transactions on Computers, 65(2):466– 479, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2016:SOC

[WRW16] C. Wang, K. Ren, and J. Wang. Secure optimization computation outsourcing in cloud computing: A case study of linear programming. IEEE Transactions on Computers, 65 (1):216–229, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Waters:2010:RCW

[WS10] Ron S. Waters and Earl E. Swartzlander. A reduced complexity Wallace multiplier reduction. IEEE Transactions on Computers, 59 (8):1134–1137, August 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5467045.

Wang:2014:ANM

[WS14] Dawei Wang and Xian-He Sun. APC: A novel memory metric and measurement methodology for modern memory systems. IEEE Transactions on Computers, 63(7):1626–1639, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Won:2015:HMR

[WS15] M. Won and R. Stoleru. A hybrid multicast routing for large scale sensor networks with holes. IEEE Transactions on Computers, 64(12): 3362–3375, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2018:BBT

[WSL^{+}18] Xiaohang Wang, Amit Kumar Singh, Bing Li, Yang Yang, Hong Li, and Terrence Mak. Bubble budgeting: Throughput optimization for dynamic workloads by exploiting dark cores in many core systems. IEEE Transactions on Computers, 67 (2):178–192, February 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8006237/.

Wu:2013:UMP

[WSXZ13] Qi Wu, Fei Sun, Wei Xu, and Tong Zhang. Using multilevel phase change memory to build data storage: A time-aware system design perspective. IEEE Transactions on Computers, 62(10): 2083–2095, October 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2016:MNP

[WSZ^{+}16] Ruijun Wang, Pengju Shang, Junyao Zhang, Qingdong Wang, Ting Liu, and Jun Wang. MAR: A novel power management for CMP systems in data-intensive environment. IEEE Transactions on Computers, 65 (6):1816–1830, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Walker:2013:PRA

[WTBT13] James Alfred Walker, Martin A. Trefzer, Simon J. Bale, and Andy M. Tyrrell. PAnDA: A reconfigurable architecture that adapts to physical substrate variations. IEEE Transactions on Computers, 62(8):1584–1596, August 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2014:MLL

[WTY^{+}14] Jianxin Wang, Peiqiang Tan, Jinyi Yao, Qilong Feng, and Jianer Chen. On the minimum link-length rectilinear spanning path problem: Complexity and algorithms. IEEE Transactions on Computers, 63(12): 3092–3100, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2019:GGB

[WTZ^{+}19] Hai Wang, Diya Tang, Ming Zhang, Sheldon X. D. Tan, Chi Zhang, He Tang, and Yuan Yuan. GDP: A greedy based dynamic power budgeting method for multi/ many-core systems in dark silicon. IEEE Transactions on Computers, 68 (4):526–541, April 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8493277/.

Wu:2014:HBM

[WW14] Jie Wu and Yunsheng Wang.

Hypercube-based multipath social feature routing in human contact networks. IEEE Transactions on Computers, 63(2):383–396, February 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2016:TEW

[WW16] Chundong Wang and Weng-Fai Wong. TreeFTL: An efficient workload-adaptive algorithm for RAM buffer management of NAND flashbased devices. IEEE Transactions on Computers, 65 (8):2618–2630, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2016:ARA

[WWM16] Liang Wang, Xiaohang Wang, and Terrence Mak. Adaptive routing algorithms for lifetime reliability optimization in network-on-chip. IEEE Transactions on Computers, 65(9):2896– 2902, September 2016. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2018:FLA

[WWT^{+}18] Hai Wang, Jiachun Wan, Sheldon X.-D. Tan, Chi Zhang, He Tang, Yuan Yuan, Keheng Huang, and Zhenghong Zhang. A fast leakage-aware full-chip transient thermal estimation method. IEEE Transactions on Computers, 67(5): 617–630, May 2018. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8125149/.

Wang:2016:CSC

[WWY^{+}16] J. Wang, L. Wang, H. Yin, Z. Wei, Z. Yang, and N. Gong. cNV SRAM: CMOS technology compatible non-volatile SRAM based ultra-low leakage energy hybrid memory system. IEEE Transactions on Computers, 65(4):1055–1067, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2018:NDH

[WWY^{+}18] Chundong Wang, Qingsong Wei, Jun Yang, Cheng Chen, Yechao Yang, and Mingdi Xue. NV-Dedup: Highperformance inline deduplication for non-volatile memory. IEEE Transactions on Computers, 67(5): 658–671, May 2018. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8115169/.

Wang:2013:NNP

[WXLL13] Bang Wang, Han Xu, Wenyu Liu, and Hui Liang. A novel node placement for long belt coverage in wireless networks. IEEE Transactions on Computers, 62(12): 2341–2353, 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2015:ONP

[WXLY15] Bang Wang, Han Xu, Wenyu Liu, and L. T. Yang. The optimal node placement for long belt coverage in wireless networks. IEEE Transactions on Computers, 64 (2):587–592, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2012:SPC

[WXS12] Hongya Wang, Yingyuan Xiao, and LihChyun Shu.

Scheduling periodic continuous queries in realtime data broadcast environments. IEEE Transactions on Computers, 61(9): 1325–1340, September 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2014:FOF

[WXW^{+}14] Zhehui Wang, Jiang Xu, Xiaowen Wu, Yaoyao Ye, Wei Zhang, Mahdi Nikdast, Xuan Wang, and Zhe Wang. Floorplan optimization of fat-tree-based networks-on-chip for chip multiprocessors. IEEE Transactions on Computers, 63(6):1446– 1459, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2015:AMI

[WYL^{+}15] Yan Wu, Chun Gang Yan, Lu Liu, Zhi Jun Ding, and Chang Jun Jiang. An adaptive multilevel indexing method for disaster service discovery. IEEE Transactions on Computers, 64(9): 2447–2459, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2014:ORS

[WZ14] Qi Wu and Tong Zhang. OFWAR: Reducing SSD response time using ondemand fast-write-and-rewrite.

IEEE Transactions on Computers, 63(10):2500–2512, October 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2015:III

[WZBB15] Xinmu Wang, Yu Zheng, A. Basak, and S. Bhunia. IIPS: Infrastructure IP for secure SoC design. IEEE Transactions on Computers, 64(8):2226–2238, August 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2016:MGN

[WZCG16] Yangbing Wu, Jianfeng Zhao, Deming Chen, and Donghui Guo. Modeling of Gaussian network-based reconfigurable network-on-chip designs. IEEE Transactions on Computers, 65 (7):2134–2142, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2019:ICC

[WZK^{+}19] L. Wang, X. Zhao, D. Kaeli, Z. Wang, and L. Eeckhout. Intra-cluster coalescing and distributed-block scheduling to reduce GPU NoC pressure. IEEE Transactions on Computers, 68 (7):1064–1076, July 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Weng:2015:TEI

[WZL15] Chuliang Weng, Jianfeng Zhan, and Yuan Luo. TSAC: Enforcing isolation of virtual machines in clouds. IEEE Transactions on Computers, 64(5):1470–1482, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2017:IDF

[WZL^{+}17] H. Wang, K. Zhao, M. Lv, X. Zhang, H. Sun, and T. Zhang. Improving 3D DRAM fault tolerance through weak cell aware error correction. IEEE Transactions on Computers, 66(5): 820–833, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2016:WNS

[WZLS16] J. Wu, X. Zhu, T. Li, and X. Sui. WBSP: A novel synchronization mechanism for architecture parallel simulation. IEEE Transactions on Computers, 65 (3):992–1005, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2012:RPM

[WZLX12] Hao Wang, Haiquan (Chuck) Zhao, Bill Lin, and Jun (Jim)

Xu. Robust pipelined memory system with worst case performance guarantee for network processing. IEEE Transactions on Computers, 61(10):1386–1400, October 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2016:FGR

[WZM^{+}16] Xiaohang Wang, Baoxin Zhao, Terrence Mak, Mei Yang, Yingtao Jiang, and Masoud Daneshtalab. On fine-grained runtime power budgeting for networks-on-chip systems. IEEE Transactions on Computers, 65(9): 2780–2793, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2019:RAD

[WZW^{+}19] Chengwen Wu, Guangyan Zhang, Yang Wang, Xinyang Jiang, and Weimin Zheng. Redio: Accelerating diskbased graph processing by reducing disk I/Os. IEEE Transactions on Computers, 68(3):414–425, March 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8489961/.

Wei:2016:PDP

[WZY16] T. Wei, Q. Zhu, and N. Yu.

Proactive demand participation of smart buildings in smart grid. IEEE Transactions on Computers, 65 (5):1392–1406, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2010:EDC

[WZZ10] Chen Wang, Bing Bing Zhou, and A. Y. Zomaya. EvolvingSpace: a data centric framework for integrating bioinformatics applications. IEEE Transactions on Computers, 59 (6):721–734, June 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5416682.

Xia:2015:CBE

[XAYL15] Feng Xia, A. M. Ahmed, L. Tianruo Yang, and Zhongxuan Luo. Communitybased event dissemination with optimal load balancing. IEEE Transactions on Computers, 64(7):1857– 1869, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiang:2016:MBT

[XCF16] Dong Xiang, Krishnendu Chakrabarty, and Hideo Fujiwara. Multicast-based testing and thermal-aware test scheduling for 3D ICs with a stacked networkon-chip. IEEE Transactions on Computers, 65(9): 2767–2779, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiao:2010:CDR

[XCW^{+}10] Yang Xiao, Hui Chen, Kui Wu, Bo Sun, Ying Zhang, Xinyu Sun, and Chong Liu. Coverage and detection of a randomized scheduling algorithm in wireless sensor networks. IEEE Transactions on Computers, 59 (4):507–521, April 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5313803.

Xie:2011:ULD

[XDZ11] Ningde Xie, Guiqiang Dong, and Tong Zhang. Using lossless data compression in data storage systems: Not for saving space. IEEE Transactions on Computers, 60(3):335–345, March 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2016:SER

[XH16] Xin Xu and H. Howie Huang. On soft error reliability of virtualization infrastructure. IEEE Transactions on Computers, 65(12):

3727–3739, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xue:2019:DAM

[XHLW19] D. Xue, L. Huang, C. Li, and C. Wu. Dapper: An adaptive manager for large-capacity persistent memory. IEEE Transactions on Computers, 68(7):1019–1034, July 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xue:2017:CNC

[XHX^{+}17] Kaiping Xue, Jianan Hong, Yingjie Xue, David S. L. Wei, Nenghai Yu, and Peilin Hong. CABE: A new comparable attributebased encryption construction with 0-encoding and 1encoding. IEEE Transactions on Computers, 66(9): 1491–1503, September 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7896558/.

Xie:2014:ESP

[XHZ14] Jin Xie, Wei Hu, and Zhenghao Zhang. Efficient software partial packet recovery in 802.11 wireless LANs. IEEE Transactions on Computers, 63(10):2402–2415, October 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiao:2016:SSD

[XHZC16] K. Xiao, X. S. Hu, B. Zhou, and D. Z. Chen. Shell: A spatial decomposition data structure for ray traversal on GPU. IEEE Transactions on Computers, 65 (1):230–243, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xia:2015:SLB

[XJFH15] Wen Xia, Hong Jiang, Dan Feng, and Yu Hua. Similarity and locality based indexing for high performance data deduplication. IEEE Transactions on Computers, 64(4):1162–1176, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xia:2016:DDA

[XJFT16] Wen Xia, Hong Jiang, Dan Feng, and Lei Tian. DARE: A deduplication-aware resemblance detection and elimination scheme for data reduction with low overheads. IEEE Transactions on Computers, 65(6):1692– 1705, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2016:CIB

[XJW^{+}16] P. Xu, T. Jiao, Q. Wu, W. Wang, and H. Jin. Conditional identity-based broadcast proxy re-encryption and its application to cloud email. IEEE Transactions on Computers, 65(1):66–79, January 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2013:PKE

[XJWW13] Peng Xu, Hai Jin, Qianhong Wu, and Wei Wang. Publickey encryption with fuzzy keyword search: A provably secure scheme under keyword guessing attack. IEEE Transactions on Computers, 62(11):2266–2277, November 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2015:AMB

[XKT^{+}15] Dianxiang Xu, M. Kent, L. Thomas, T. Mouelhi, and Y. Le Traon. Automated model-based testing of role-based access control using predicate/transition nets. IEEE Transactions on Computers, 64(9):2490– 2505, September 2015. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xhafa:2016:GES

[XL16] Fatos Xhafa and Vincenzo Loia. Guest editorial for special section on advanced techniques for efficient and reliable cloud storage. IEEE Transactions on Computers, 65(8):2346–2347, August 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiao:2014:IMP

[XLC14] Zhen Xiao, Cheng Liu, and Qi Chen. Improving MapReduce performance using smart speculative execution strategy. IEEE Transactions on Computers, 63(4):954–967, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2015:ESP

[XLF15] Dan Xu, Xin Liu, and Bin Fan. Efficient server provisioning and offloading policies for Internet data centers with dynamic loaddemand. IEEE Transactions on Computers, 64 (3):682–697, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2016:HIA

[XLJ16] Fei Xu, Fangming Liu, and Hai Jin. Heterogeneity and interference-aware virtual machine provisioning for predictable performance in the cloud. IEEE Transactions on Computers, 65 (8):2470–2483, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2014:IML

[XLL^{+}14] Fei Xu, Fangming Liu, Linghui Liu, Hai Jin, Bo Li, and Baochun Li. iAware: Making live migration of virtual machines interferenceaware in the cloud. IEEE Transactions on Computers, 63(12):3012–3025, December 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2015:AAM

[XLL15] Wenzheng Xu, Weifa Liang, and Xiaola Lin. Approximation algorithms for min-max cycle cover problems. IEEE Transactions on Computers, 64(3):600–613, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xie:2018:CEC

[XLL^{+}18] Xiaolong Xie, Yun Liang, Xiuhong Li, Yudong Wu, Guangyu Sun, Tao Wang, and Dongrui Fan. CRAT: Enabling coordinated register allocation and threadlevel parallelism optimization for GPUs. IEEE Transactions on Computers, 67(6):890–897, June 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8119900/.

Xiao:2012:PMW

[XLS^{+}12] Yu Xiao, Wei Li, Matti Siekkinen, Petri Savolainen, Antti Yla-Jaaski, and Pan Hui. Power management for wireless data transmission using complex event processing. IEEE Transactions on Computers, 61(12): 1765–1777, December 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2011:ESS

[XLTZ11] Ping Xu, Xiang-Yang Li, Shaojie Tang, and JiZhong Zhao. Efficient and strategyproof spectrum allocations in multichannel wireless networks. IEEE Transactions on Computers, 60(4):580– 593, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2014:SLP

[XLW14] Han Xu, Wenyu Liu, and Bang Wang. Subarea localization performance of the divide-and-cover node deployment in a long-bounded belt scenario. IEEE Transactions on Computers, 63(11): 2842–2850, November 2014.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2014:SDF

[XLX^{+}14] Yinlong Xu, John C. S. Lui, Liping Xiang, Yunfeng Zhu, Runhui Li, Patrick P. C. Lee, and Silei Xu. Single disk failure recovery for X-code-based parallel storage systems. IEEE Transactions on Computers, 63(4):995–1007, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xie:2013:HER

[XMH13] Jiafeng Xie, Pramod Kumar Meher, and Jianjun He. Hardware-efficient realization of prime-length DCT based on distributed arithmetic. IEEE Transactions on Computers, 62 (6):1170–1178, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xia:2010:SNL

[XP10] Yinglong Xia and V. K. Prasanna. Scalable nodelevel computation kernels for parallel exact inference. IEEE Transactions on Computers, 59(1):103– 115, January 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=
&arnumber=5184808.

Xue:2015:OLI

[XSR15] Bin Xue, S. K. Shukla, and S. S. Ravi. Optimization of latency insensitive systems through back pressure minimization. IEEE Transactions on Computers, 64 (2):464–476, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiao:2012:CMP

[XTF^{+}12] Ying Xiao, Krishnaiyan Thulasiraman, Xi Fang, Dejun Yang, and Guoliang Xue. Computing a most probable delay constrained path: NP-hardness and approximation schemes. IEEE Transactions on Computers, 61(5):738–744, May 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2015:OEE

[XTW15] Z. Xu, Y. Tu, and X. Wang.

Online energy estimation of relational operations in database systems. IEEE Transactions on Computers, 64(11):3223–3236, November 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiao:2014:CAO

[XWH14] Mingjun Xiao, Jie Wu, and Liusheng Huang. Communityaware opportunistic routing in mobile social networks. IEEE Transactions on Computers, 63(7):1682– 1695, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2010:SSO

[XWL10] Ping Xu, ShiGuang Wang, and Xiang-Yang Li. SALSA: Strategyproof online spectrum admissions for wireless networks. IEEE Transactions on Computers, 59(12): 1691–1702, December 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5453346.

Xie:2016:IAC

[XWL^{+}16a] K. Xie, X. Wang, X. Liu, J. Wen, and J. Cao. Interference-aware cooperative communication in multiradio multi-channel wireless networks. IEEE Transactions on Computers, 65 (5):1528–1542, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2016:HSA

[XWL^{+}16b] Chang Xu, Gang Wang, Xiaoguang Liu, Dongdong Guo, and Tie-Yan Liu. Health status assessment and failure prediction for hard drives with recurrent neural networks. IEEE Transactions on Computers, 65(11):3502–3508, November 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiong:2017:ERT

[XWLX17] Qin Xiong, Fei Wu, Zhonghai Lu, and Changsheng Xie. Extending real-time analysis for wormhole NoCs. IEEE Transactions on Computers, 66(9):1532–1546, September 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7884964/.

Xiang:2010:SMM

[XWY10] Xiaojing Xiang, Xin Wang, and Yuanyuan Yang. Stateless multicasting in mobile ad hoc networks. IEEE Transactions on Computers, 59(8):1076–1090, August 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5467044.

Xie:2017:DUR

[XXBL17] Dingbao Xie, Wen Xiong, Lei Bu, and Xuandong Li. Deriving unbounded reachability proof of linear hybrid automata during bounded checking procedure. IEEE Transactions on Computers, 66(3):416–430, March 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xue:2015:USC

[XYF^{+}15] Wei Xue, Chao Yang, Haohuan Fu, Xinliang Wang, Yangtong Xu, Junfeng Liao, Lin Gan, Y. Lu, R. Ranjan, and L. Wang. Ultra-scalable CPU-MIC acceleration of mesoscale atmospheric modeling on Tianhe-2. IEEE Transactions on Computers, 64(8):2382–2393, August 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xue:2017:PCG

[XYHD17] J. Xue, Z. Yang, S. Hou, and Y. Dai. Processing concurrent graph analytics with decoupled computation model. IEEE Transactions on Computers, 66(5):876– 890, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiao:2019:NEA

[XZL^{+}19] C. Xiao, L. Zhang, W. Liu, L. Cheng, P. Li, Y. Pan, and N. Bergmann. NV-eCryptfs: Accelerating enterprise-level cryptographic file system with non-volatile memory. IEEE Transactions on Computers, 68(9):1338–1352, September 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yazdani:2019:LPH

[YAG19] R. Yazdani, J. Arnau, and A. González. A lowpower, high-performance speech recognition accelerator. IEEE Transactions on Computers, 68(12): 1817–1831, December 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yun:2017:BDM

[YAGB17] Heechul Yun, Waqar Ali, Santosh Gondi, and Siddhartha Biswas. BWLOCK: A dynamic memory access control framework for soft real-time applications on multicore platforms. IEEE Transactions on Computers, 66(7):1247–1252, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07784697-
abs.html.

Yamashita:2010:SOT

[Yam10] T. Yamashita. Stabilityoptimized time adjustment for a networked computer clock. IEEE Transactions on Computers, 59(10): 1320–1336, October 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5467048.

Yang:2014:PDB

[Yan14] Xu Yang. Path-dividing based scheduling algorithm for reducing energy consumption of clustered VLIW architectures. IEEE Transactions on Computers, 63 (10):2526–2539, October 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yildiz:2014:OCP

[YASS14] Enes Yildiz, Kemal Akkaya, Esra Sisikoglu, and Mustafa Y. Sir. Optimal camera placement for providing angular coverage in wireless video sensor networks. IEEE Transactions on Computers, 63(7):1812–1825, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2015:MUC

[YCCJ15] Lei Yang, Jiannong Cao, Hui Cheng, and Yusheng Ji. Multi-user computation partitioning for latency sensitive mobile cloud applications. IEEE Transactions on Computers, 64 (8):2253–2266, August 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu-Chun:2015:CTD

[YCCWC15] Cheng Yu-Chun, Cheng-Fu Chou, E. H. Wu, and Gen-Huey Chen. A cognitive TCP design for a cognitive radio network with an unstable-bandwidth link. IEEE Transactions on Computers, 64(10):2730–3740, October 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yun:2010:HSN

[YCK10] A. Yun, Jung Hee Cheon, and Yongdae Kim. On homomorphic signatures for network coding. IEEE Transactions on Computers, 59(9): 1295–1296, September 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5441287.

Yang:2016:VFC

[YCK16] Ming-Chang Yang, Yuan-Hao Chang, and Tei-Wei Kuo. Virtual flash chips: Reinforcing the hardware abstraction layer to improve data recoverability of flash devices. IEEE Transactions on Computers, 65(9): 2872–2883, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2016:CIA

[YCKH16] M. Yang, Y. Chang, T. Kuo, and P. Huang. Capacityindependent address mapping for flash storage devices with explosively growing capacity. IEEE Transactions on Computers, 65 (2):448–465, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2012:SCS

[YCL^{+}12] Jing Yang, Qiang Cao, Xu Li, Changsheng Xie, and Qing Yang. ST-CDP: Snapshots in TRAP for continuous data protection. IEEE Transactions on Computers, 61(6):753–766, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2016:CAS

[YCLH16] L. Yang, J. Cao, G. Liang, and X. Han. Cost aware service placement and load dispatching in mobile cloud systems. IEEE Transactions on Computers, 65 (5):1440–1452, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2011:EED

[YCW11] Ching-Nung Yang, Chih-Yang Chiu, and Gen-Chin Wu. Efficient encoding/ decoding for second-order spectral-null codes by reducing random walks. IEEE Transactions on Computers, 60(10):1503–1510, October 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5669269.

Yang:2019:IWR

[YCW^{+}19] Ming-Chang Yang, Yuan-Hao Chang, Fenggang Wu, Tei-Wei Kuo, and David H. C. Du. On improving the write responsiveness for host-aware SMR drives. IEEE Transactions on Computers, 68 (1):111–124, January 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8387485/.

Yoon:2010:MMF

[YCZ10] MyungKeun Yoon, Shigang Chen, and Zhan Zhang. Minimizing the maximum firewall rule set in a network with multiple firewalls. IEEE Transactions on Computers, 59(2):218–230, February 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5332225.

Yu:2015:GMA

[YEG^{+}15] Z. Yu, L. Eeckhout, N. Goswami, T. Li, L. K. John, H. Jin, C. Xu, and J. Wu. GPGPU-MiniBench: Accelerating GPGPU micro-architecture simulation. IEEE Transactions on Computers, 64(11): 3153–3166, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yaghini:2016:CIT

[YEY^{+}16] P. M. Yaghini, A. Eghbal, S. S. Yazdi, N. Bagherzadeh, and M. M. Green. Capacitive and inductive TSVto-TSV resilient approaches for 3D ICs. IEEE Transactions on Computers, 65 (3):693–705, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yao:2014:NRP

[YFCV14] Gavin Xiaoxu Yao, Junfeng Fan, Ray C. C. Cheung, and Ingrid Verbauwhede. Novel RNS parameter selection for fast modular multiplication. IEEE Transactions on Computers, 63 (8):2099–2105, August 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2014:PMP

[YFJ^{+}14] Baohua Yang, Jeffrey Fong, Weirong Jiang, Yibo Xue, and Jun Li. Practical multituple packet classification using dynamic discrete bit selection. IEEE Transactions on Computers, 63 (2):424–434, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yi:2010:NAS

[YG10] Kyueun Yi and J.-L. Gaudiot. Network applications on simultaneous multithreading processors. IEEE Transactions on Computers, 59(9):1200–1209, September 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5374374.

Yu:2015:FMI

[YGS15] Shui Yu, Song Guo, and I. Stojmenovic. Fool me if you can: Mimicking attacks and anti-attacks in cyberspace. IEEE Transactions on Computers, 64 (1):139–151, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2019:ROP

[YHC19] M. Yang, W. Huang, and J. Chen. Resource-oriented partitioning for multiprocessor systems with shared resources. IEEE Transactions on Computers, 68(6):882– 898, June 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yin:2012:MCS

[YHH^{+}12] Liangze Yin, Fei He, William N. N. Hung, Xiaoyu Song, and Ming Gu. Maxterm covering for satisfiability. IEEE Transactions on Computers, 61(3):420–426, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2016:ECA

[YHML16] Y. Yang, Y. Huang, X. Ma, and J. Lu. Enabling contextawareness by predicate detection in asynchronous environments. IEEE Transactions on Computers, 65 (2):522–534, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yue:2016:RLS

[YHT^{+}16] Y. Yue, B. He, L. Tian, H. Jiang, F. Wang, and D. Feng. Rotated logging storage architectures for data centers: Models and optimizations. IEEE Transactions on Computers, 65 (1):203–215, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2013:QDD

[YHV13] Heng Yu, Yajun Ha, and Bharadwaj Veeravalli. Qualitydriven dynamic scheduling for real-time adaptive applications on multiprocessor systems. IEEE Transactions on Computers, 62(10): 2026–2040, October 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2019:RYR

[YKDEV19] J. Yu, D. Kozhaya, J. Decouchant, and P. Esteves-Verissimo. RepuCoin: Your reputation is your power. IEEE Transactions on Computers, 68(8):1225–1237, August 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yaakob:2015:PIA

[YKK^{+}15] N. Yaakob, I. Khalil, H. Kumarage, M. Atiquzzaman, and Z. Tari. By-passing infected areas in wireless sensor networks using BPR. IEEE Transactions on Computers, 64 (6):1594–1606, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2014:REB

[YL14] Weihua Yang and Huiqiu Lin. Reliability evaluation of BC networks in terms of the extra vertex- and edgeconnectivity. IEEE Transactions on Computers, 63(10): 2540–2548, October 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Younis:2010:LAR

[YLA10] M. F. Younis, Sookyoung Lee, and A. A. Abbasi. A localized algorithm for restoring internode connectivity in networks of moveable sensors. IEEE Transactions on Computers, 59(12): 1669–1682, December 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5551121.

Yuen:2015:TAB

[YLA^{+}15] Tsz Hon Yuen, J. K. Liu, Man Ho Au, Xinyi Huang, W. Susilo, and Jianying Zhou. k-times attributebased anonymous access control for cloud computing. IEEE Transactions on Computers, 64(9):2595– 2608, September 2015. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2019:OAM

[YLGD19] L. Yang, W. Liu, N. Guan, and N. Dutt. Optimal application mapping and scheduling for network-on-chips with computation in STT–RAM based router. IEEE Transactions on Computers, 68(8):1174–1189, August 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yaguez:2014:SAS

[YLGE14] Alvaro Garcia Yaguez, Diego R.

Llanos, and Arturo Gonzalez-Escribano. Squashing alternatives for software-based speculative parallelization. IEEE Transactions on Computers, 63(7):1826–1839, July 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yoon:2010:MPJ

[YLH10] Man-Ki Yoon, Chang-Gun Lee, and Junghee Han. Migrating from per-job analysis to per-resource analysis for tighter bounds of end-to-end response times. IEEE Transactions on Computers, 59(7):933–942, July 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5342409.

Yoo:2013:PNB

[YLH13] Jonghun Yoo, Jaesoo Lee, and Seongsoo Hong. Petri net-based FTL architecture for parametric WCET estimation via FTL operation sequence derivation. IEEE Transactions on Computers, 62(11):2238–2251, November 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ye:2016:FPD

[YLL16] Liang-Cheng Ye, Jia-Rong Liang, and Hai-Xiang Lin. A fast pessimistic diagnosis algorithm for hypercube-like networks under the comparison model. IEEE Transactions on Computers, 65(9): 2884–2888, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2015:POU

[YLML15] Wangdong Yang, Kenli Li, Zeyao Mo, and Keqin Li. Performance optimization using partitioned SpMV on GPUs and multicore CPUs. IEEE Transactions on Computers, 64(9): 2623–2636, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2015:RCL

[YLP15] Ching-Nung Yang, Zih-Yang Lin, and Sheng-Lung Peng. Reducing code length of second-order spectral-null code. IEEE Transactions on Computers, 64(2):492– 503, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2015:NRF

[YLY^{+}15a] Xinyu Yang, Jie Lin, Wei Yu, P.-M. Moulema, Xinwen Fu, and Wei Zhao. A novel en-route filtering scheme against false data injection attacks in cyber-physical networked systems. IEEE Transactions on Computers, 64(1):4–18, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yuan:2015:DFM

[YLY15b] Dong Yuan, Xiao Liu, and Yun Yang. Dynamic onthe-fly minimum cost benchmarking for storing generated scientific datasets in the cloud. IEEE Transactions on Computers, 64(10): 2781–2795, October 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2011:PTE

[YM11] Heeyeol Yu and Rabi N.

Mahapatra. A power and throughput-efficient packet classifier with n Bloom filters. IEEE Transactions on Computers, 60(8):1182– 1193, August 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2017:ECB

[YMAG17] Bo Yang, Kalikinkar Mandal, Mark D. Aagaard, and Guang Gong. Efficient composited de Bruijn sequence generators. IEEE Transactions on Computers, 66 (8):1354–1368, August 2017.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7869363/.

Yavits:2015:CAA

[YMG15] L. Yavits, A. Morad, and R. Ginosar. Computer architecture with associative processor replacing last-level cache and SIMD accelerator. IEEE Transactions on Computers, 64(2):368–381, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yavits:2016:ETA

[YMG16] L. Yavits, A. Morad, and R. Ginosar. The effect of temperature on Amdahl Law in 3D multicore era. IEEE Transactions on Computers, 65(6):2010–2013, June 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yasudo:2017:SNC

[YMK^{+}17] Ryota Yasudo, Hiroki Matsutani, Michihiro Koibuchi, Hideharu Amano, and Tadao Nakamura. Scalable networkson-chip with elastic links demarcated by decentralized routers. IEEE Transactions on Computers, 66(4):702– 716, April 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yin:2013:ISM

[YMT13] Xiaoyan Yin, Xiaomin Ma, and Kishor S. Trivedi. An interacting stochastic models approach for the performance evaluation of DSRC vehicular safety communication. IEEE Transactions on Computers, 62(5):873– 885, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yin:2014:PRE

[YMTV14] Xiaoyan Yin, Xiaomin Ma, K. S. Trivedi, and A. Vinel. Performance and reliability evaluation of BSM broadcasting in DSRC with multichannel schemes. IEEE Transactions on Computers, 63(12):3101–3113, December 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2012:HPC

[YP12] Yi-Hua Edward Yang and Viktor K. Prasanna. Highperformance and compact architecture for regular expression matching on FPGA. IEEE Transactions on Computers, 61(7):1013–1025, July 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yao:2016:GRT

[YPB^{+}16] Gang Yao, Rodolfo Pellizzoni, Stanley Bak, Heechul Yun, and Marco Caccamo. Global real-time memorycentric scheduling for multicore systems. IEEE Transactions on Computers, 65(9): 2739–2751, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yi:2013:ICP

[YRG13] Kyueun Yi, Won W. Ro, and Jean-Luc Gaudiot. Importance of coherence protocols with network applications on multicore processors. IEEE Transactions on Computers, 62(1):6–15, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yi:2016:IPA

[YRT^{+}16] Xun Yi, Fang-Yu Rao, Zahir Tari, Feng Hao, Elisa Bertino, Ibrahim Khalil, and Albert Y. Zomaya. ID2S password-authenticated key exchange protocols. IEEE Transactions on Computers, 65(12):3687–3701, December 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yan:2016:CRS

[YSLL16] G. Yan, F. Sun, H. Li, and X. Li. CoreRank: Redeeming “Sick Silicon” by dynamically quantifying core-level healthy condition. IEEE Transactions on Computers, 65(3):716–729, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yi:2014:DIA

[YSZ^{+}14] Letian Yi, Jiwu Shu, Ying Zhao, Yingjin Qian, Youyou Lu, and Weimin Zheng. Design and implementation of an asymmetric block-based parallel file system. IEEE Transactions on Computers, 63(7):1723–1735, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yin:2017:AER

[YTD^{+}17] Jianwei Yin, Yan Tang, Shuiguang Deng, Ying Li, Wei Lo, Kexiong Dong, Albert Y. Zomaya, and Calton Pu. ASSER: An efficient, reliable, and costeffective storage scheme for object-based cloud storage systems. IEEE Transactions on Computers, 66(8): 1326–1340, August 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7859394/.

Yin:2018:DDD

[YTD^{+}18] Jianwei Yin, Yan Tang, Shuiguang Deng, Ying Li, and Albert Y. Zomaya. D^{3}: A dynamic dualphase deduplication framework for distributed primary storage. IEEE Transactions on Computers, 67 (2):193–207, February 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8015182/.

Yu:2016:NSA

[YTM16] Lei Yu, Fei Teng, and Frédéric Magoulès. Node scaling analysis for poweraware real-time tasks scheduling. IEEE Transactions on Computers, 65(8):2510– 2521, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2012:TPI

[YTND12] Joon-Sung Yang, Nur A. Touba, and Benoit Nadeau-Dostie. Test point insertion with control points driven by existing functional flip-flops. IEEE Transactions on Computers, 61(10): 1473–1483, October 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yuce:2014:FEC

[YUGD14] Bilgiday Yuce, H. Fatih Ugurdag, Sezer Gören, and Günhan Dündar. Fast and efficient circuit topologies for finding the maximum of n kbit numbers. IEEE Transactions on Computers, 63 (8):1868–1881, August 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yun:2012:ETB

[Yun12] SangKyun Yun. An efficient TCAM-based implementation of multipattern matching using covered state encoding. IEEE Transactions on Computers, 61 (2):213–221, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2012:UEP

[YW12] Kai-Chao Yang and Jia-Shung Wang. Unequal error protection for streaming media based on rateless codes. IEEE Transactions on Computers, 61(5): 666–675, May 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ye:2019:NCA

[YWM19] T. Ye, Y. Wei, and W. Meier.

A new cube attack on MORUS by using division property. IEEE Transactions on Computers, 68(12): 1731–1740, December 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2015:VSE

[YWQX15] Weikuan Yu, Yandong Wang, Xinyu Que, and Cong Xu. Virtual shuffling for efficient data movement in MapReduce. IEEE Transactions on Computers, 64 (2):556–568, February 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2016:NTC

[YWW^{+}16] Jun Yang, Qingsong Wei, Chundong Wang, Cheng Chen, Khai Leong Yong, and Bingsheng He. NV-tree: A consistent and workloadadaptive tree structure for non-volatile memory. IEEE Transactions on Computers, 65(7):2169–2183, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2012:RWE

[YWXZ12] Xuejun Yang, Zhiyuan Wang, Jingling Xue, and Yun Zhou. The reliability wall for exascale supercomputing. IEEE Transactions on Computers, 61(6):767– 779, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yin:2016:FSE

[YXWL16] Y. Yin, L. Xie, J. Wu, and S. Lu. Focus and shoot: Exploring auto-focus in RFID tag identification towards a specified area. IEEE Transactions on Computers, 65 (3):888–901, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2014:BBM

[YXZZ14] Hongliang Yu, Xiaojia Xiang, Ying Zhao, and Weimin Zheng. BIRDS: A baremetal recovery system for instant restoration of data services. IEEE Transactions on Computers, 63 (6):1392–1407, June 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2010:EHP

[YY10] Min Yang and Yuanyuan Yang. An efficient hybrid peer-to-peer system for distributed data sharing. IEEE Transactions on Computers, 59(9):1158–1171, September 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5342410.

Yang:2014:ANC

[YY14] Min Yang and Yuanyuan Yang. Applying network coding to peer-to-peer file sharing. IEEE Transactions on Computers, 63(8): 1938–1950, August 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yoo:2012:AMP

[YYC12] Junhee Yoo, Sungjoo Yoo, and Kiyoung Choi. Active memory processor for network-on-chip-based architecture. IEEE Transactions on Computers, 61(5):622– 635, May 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yan:2011:RSA

[YyHL11] Guihai Yan, yinhe Han, and Xiaowei Li. ReviveNet: a self-adaptive architecture for improving lifetime reliability via localized timing adaptation. IEEE Transactions on Computers, 60(9): 1219–1232, September 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5710881.

Yun:2016:MBM

[YYP^{+}16] H. Yun, G. Yao, R. Pellizzoni, M. Caccamo, and L. Sha. Memory bandwidth management for efficient performance isolation in multicore platforms. IEEE Transactions on Computers, 65 (2):562–576, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yao:2016:SAM

[YYW^{+}16] G. Yao, H. Yun, Z. P. Wu, R. Pellizzoni, M. Caccamo, and L. Sha. Schedulability analysis for memory bandwidth regulated multicore real-time systems. IEEE Transactions on Computers, 65(2):601–614, February 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ye:2015:SAS

[YZ15] Dayong Ye and Minjie Zhang. A self-adaptive strategy for evolution of cooperation in distributed networks. IEEE Transactions on Computers, 64(4):899– 911, April 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2010:LAI

[YZF^{+}10] Wei Yu, Nan Zhang, Xinwen Fu, R. Bettati, and Wei Zhao. Localization attacks to Internet threat monitors: Modeling and countermeasures. IEEE Transactions on Computers, 59(12): 1655–1668, December 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5453347.

Yu:2016:FIT

[YZGG16] S. Yu, W. Zhou, S. Guo, and M. Guo. A feasible IP traceback framework through dynamic deterministic packet marking. IEEE Transactions on Computers, 65(5):1418–1427, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yao:2015:OOD

[YZH^{+}15] H. Yao, D. Zeng, H. Huang, S. Guo, A. Barnawi, and I. Stojmenovic. Opportunistic offloading of deadlineconstrained bulk cellular traffic in vehicular DTNs. IEEE Transactions on Computers, 64(12):3515–3527, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2012:DCR

[YZHX12] Jian Yang, Ke Zeng, Han Hu, and Hongsheng Xi. Dynamic cluster reconfiguration for energy conservation in computation intensive service. IEEE Transactions on Computers, 61(10): 1401–1416, October 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zidaric:2019:HOA

[ZAG19] Nusa Zidaric, Mark Aagaard, and Guang Gong.

Hardware optimizations and analysis for the WG-16 cipher with tower field arithmetic. IEEE Transactions on Computers, 68 (1):67–82, January 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8409309/.

Zhang:2017:ARF

[ZBK^{+}17] Hongyan Zhang, Lars Bauer, Michael Andreas Kochte, Eric Schneider, Hans-Joachim Wunderlich, and Jörg Henkel. Aging resilience and fault tolerance in runtime reconfigurable architectures. IEEE Transactions on Computers, 66(6):957–970, June 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7588134/.

Zhao:2017:GFT

[ZBW17] Hongzhi Zhao, Nader Bagherzadeh, and Jie Wu. A general faulttolerant minimal routing for mesh architectures. IEEE Transactions on Computers, 66(7):1240–1246, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07814237-
abs.html.

Zheng:2013:MPC

[ZC13] Qiang Zheng and Guohong Cao. Minimizing probing cost and achieving identifiability in probe-based network link monitoring. IEEE Transactions on Computers, 62(3):510–523, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2014:MRH

[ZCC^{+}14] Weiping Zhu, Jiannong Cao, Henry C. B. Chan, Xuefeng Liu, and Vaskar Raychoudhury. Mobile RFID with a high identification rate. IEEE Transactions on Computers, 63(7):1778– 1792, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2019:EMP

[ZCK19] H. Zhang, D. Chen, and S. Ko. Efficient multipleprecision floating-point fused multiply-add with mixedprecision support. IEEE Transactions on Computers, 68(7):1035–1048, July 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2016:SMS

[ZCL^{+}16] Detian Zhang, Chi-Yin Chow, Qing Li, Xinming Zhang, and Yinlong Xu. A spatial mashup service for efficient evaluation of concurrent k-NN queries. IEEE Transactions on Computers, 65(8):2428–2442, August 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2016:PDA

[ZCR16] W. Zhu, J. Cao, and M. Raynal. Predicate detection in asynchronous distributed systems: A probabilistic approach. IEEE Transactions on Computers, 65 (1):173–186, January 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2016:DPP

[ZCS16] W. Zhang, A. M. K. Cheng, and J. Subhlok. Dwarf-Code: A performance prediction tool for parallel applications. IEEE Transactions on Computers, 65 (2):495–507, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2018:ACT

[ZCW18] Jian Zhang, Shuming Chen, and Yaohua Wang. Advancing CMOS-type Ising arithmetic unit into the domain of real-world applications. IEEE Transactions on Computers, 67(5): 604–616, May 2018. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8118124/.

Zhu:2016:PBM

[ZCY^{+}16] Di Zhu, Lizhong Chen, Siyu Yue, Timothy M. Pinkston, and Massoud Pedram. Providing balanced mapping for multiple applications in many-core chip multiprocessors. IEEE Transactions on Computers, 65(10): 3122–3135, October 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2015:AAB

[ZCYX15] X. Zhu, C. Chen, L. T. Yang, and Y. Xiang. ANGEL: Agent-based scheduling for real-time tasks in virtualized clouds. IEEE Transactions on Computers, 64(12): 3389–3403, December 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zabihi:2019:MPS

[ZCZ^{+}19] M. Zabihi, Z. I. Chowdhury, Z. Zhao, U. R. Karpuzcu, J. Wang, and S. S. Sapatnekar. In-memory processing on the spintronic CRAM: From hardware design to application mapping. IEEE Transactions on Computers, 68(8):1159– 1173, August 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhai:2016:PPL

[ZCZL16] Jidong Zhai, Wenguang Chen, Weimin Zheng, and Keqin Li. Performance prediction for large-scale parallel applications using representative replay. IEEE Transactions on Computers, 65(7):2184–2198, July 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zeng:2013:EFR

[ZD13] Haibo Zeng and Marco Di Natale. An efficient formulation of the real-time feasibility region for design optimization. IEEE Transactions on Computers, 62(4):644–661, April 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2015:PAL

[ZDP^{+}15] Xuyun Zhang, Wanchun Dou, Jian Pei, S. Nepal, Chi Yang, Chang Liu, and Jinjun Chen. Proximity-aware local-recoding anonymization with MapReduce for scalable big data privacy preservation in cloud. IEEE Transactions on Computers, 64(8):2293–2307, August 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2013:PVA

[ZDYZ13] Bo Zhao, Yu Du, Jun Yang, and Youtao Zhang. Process variation-aware nonuniform cache management in a 3D die-stacked multicore processor. IEEE Transactions on Computers, 62(11): 2252–2265, November 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See errata [ZDYZ14].

Zhao:2014:EPV

[ZDYZ14] Bo Zhao, Yu Du, Jun Yang, and Youtao Zhang. Errata to “Process Variation-Aware Nonuniform Cache Management in a 3D Die– Stacked Multicore Processor. IEEE Transactions on Computers, 63(2):525–526, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [ZDYZ13].

Zhang:2019:OFS

[ZFHC19] Xiaoyi Zhang, Dan Feng, Yu Hua, and Jianxi Chen. Optimizing file systems with a write-efficient journaling scheme on non-volatile memory. IEEE Transactions on Computers, 68 (3):402–413, March 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8466031/.

Zhang:2017:FAE

[ZFJ^{+}17] Yucheng Zhang, Dan Feng, Hong Jiang, Wen Xia, Min Fu, Fangting Huang, and Yukun Zhou. A fast asymmetric extremum content defined chunking algorithm for data deduplication in backup storage systems. IEEE Transactions on Computers, 66(2):199–211, February 2017. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zeng:2015:ISM

[ZGB^{+}15] Deze Zeng, Song Guo, A. Barnawi, Shui Yu, and I. Stojmenovic. An improved stochastic modeling of opportunistic routing in vehicular CPS. IEEE Transactions on Computers, 64 (7):1819–1829, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zeng:2016:JOT

[ZGG^{+}16] Deze Zeng, Lin Gu, Song Guo, Zixue Cheng, and Shui Yu. Joint optimization of task scheduling and image placement in fog computing supported software-defined embedded system. IEEE Transactions on Computers, 65(12):3702–3712, December 2016. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zang:2013:STL

[ZGR13] Wei Zang and Ann Gordon-Ross. T-SPaCS — a twolevel single-pass cache simulation methodology. IEEE Transactions on Computers, 62(2):390–403, February 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2014:RDS

[ZGW14] Qiang Zhu, Guodong Guo, and Dajin Wang. Relating diagnosability, strong diagnosability and conditional diagnosability of strong networks. IEEE Transactions on Computers, 63(7):1847– 1851, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2015:SSP

[ZGWC15] Zhenzhe Zheng, Yang Gui, Fan Wu, and Guihai Chen. STAR: Strategy-proof double auctions for multicloud, multi-tenant bandwidth reservation. IEEE Transactions on Computers, 64(7):2071–2083, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2013:EAA

[ZGY13] Zhemin Zhang, Zhiyang Guo, and Yuanyuan Yang. Efficient all-to-all broadcast in Gaussian on-chip networks. IEEE Transactions on Computers, 62(10): 1959–1971, October 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2014:BRO

[ZGY14] Zhemin Zhang, Zhiyang Guo, and Yuanyuan Yang. Bufferless routing in optical Gaussian macrochip interconnect. IEEE Transactions on Computers, 63(11): 2685–2700, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2010:OWS

[Zhe10] S. Q. Zheng. An optimal wide-sense nonblocking distributor. IEEE Transactions on Computers, 59(12): 1709–1714, December 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5483291.

Zhang:2014:BMC

[ZHM14] Baoxian Zhang, Jie Hao, and Hussein T. Mouftah. Bidirectional multi-constrained routing algorithms. IEEE Transactions on Computers, 63(9):2174–2186, September 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2019:IAM

[ZHM^{+}19] J. Zhou, X. S. Hu, Y. Ma, J. Sun, T. Wei, and S. Hu. Improving availability of multicore real-time systems suffering both permanent and transient faults. IEEE Transactions on Computers, 68(12):1785–1801, December 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2019:PPA

[ZHQX19] Panping Zhou, Jianzhong Huang, Xiao Qin, and Changsheng Xie. PaRS: A popularity-aware redundancy scheme for in-memory stores. IEEE Transactions on Computers, 68 (4):556–569, April 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8502118/.

Zhou:2015:EPP

[ZHW15] Zhibin Zhou, Dijiang Huang, and Zhijie Wang. Efficient privacy-preserving ciphertext-policy attribute based-encryption and broadcast encryption. IEEE Transactions on Computers, 64(1):126–138, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2016:PPV

[ZHW^{+}16] Lei Zhang, Chuanyan Hu, Qianhong Wu, Josep Domingo-Ferrer, and Bo Qin. Privacypreserving vehicular communication authentication with hierarchical aggregation and fast response. IEEE Transactions on Computers, 65 (8):2562–2574, August 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zadegan:2012:ATA

[ZICL12] Farrokh Ghani Zadegan, Urban Ingelsson, Gunnar Carlsson, and Erik Larsson. Access time analysis for IEEE P1687. IEEE Transactions on Computers, 61(10): 1459–1472, October 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2014:SAP

[ZJH^{+}14] Hehua Zhang, Yu Jiang, W. N. N. Hung, Xiaoyu Song, Ming Gu, and Jiaguang Sun. Symbolic analysis of programmable logic controllers. IEEE Transactions on Computers, 63(10): 2563–2575, October 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2016:FTO

[ZJL^{+}16] H. Zhang, H. Jiang, B. Li, F. Liu, A. V. Vasilakos, and J. Liu. A framework for truthful online auctions in cloud computing with heterogeneous user demands. IEEE Transactions on Computers, 65(3):805– 818, March 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhan:2014:CCO

[ZJS14] Dongyuan Zhan, Hong Jiang, and Sharad C. Seth. CLU: Co-optimizing locality and utility in threadaware capacity management for shared last level caches. IEEE Transactions on Computers, 63(7):1656– 1667, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2011:OPV

[ZJXL11] Jun Zhu, Zhefu Jiang, Zhen Xiao, and Xiaoming Li. Optimizing the performance of virtual machine synchronization for fault tolerance. IEEE Transactions on Computers, 60(12): 1718–1729, December 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5629326.

Zhong:2011:SSS

[ZL11] Zhenyu Zhong and Kang Li. Speed up statistical spam filter by approximation. IEEE Transactions on Computers, 60(1):120– 134, January 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2015:SOF

[ZL15] Zibin Zheng and M. R.

Lyu. Selecting an optimal fault tolerance strategy for reliable service-oriented systems with local and global constraints. IEEE Transactions on Computers, 64 (1):219–232, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2016:WAS

[ZL16] Qi Zhang and Ling Liu.

Workload adaptive shared memory management for high performance network I/O in virtualized cloud. IEEE Transactions on Computers, 65(11):3480–3494, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2018:DIV

[ZL18] Tianwei Zhang and Ruby B.

Lee. Design, implementation and verification of cloud architecture for monitoring a virtual machine’s security health. IEEE Transactions on Computers, 67(6):799–815, June 2018.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8169039/.

Zhang:2019:CCF

[ZL19] X. Zhang and Y. Lao. On the construction of composite finite fields for hardware obfuscation. IEEE Transactions on Computers, 68(9): 1353–1364, September 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zahaf:2019:PMM

[ZLBB19] H. Zahaf, G. Lipari, M. Bertogna, and P. Boulet. The parallel multi-mode digraph task model for energy-aware realtime heterogeneous multicore systems. IEEE Transactions on Computers, 68(10): 1511–1524, October 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zeng:2015:EMM

[ZLG^{+}15] D. Zeng, P. Li, S. Guo, T. Miyazaki, J. Hu, and Y. Xiang. Energy minimization in multi-task softwaredefined sensor networks. IEEE Transactions on Computers, 64(11):3128–3139, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2015:QCC

[ZLH^{+}15] Jiantao Zhou, Caihe Lan, W. N. N. Hung, Xinrui Guo, and Xiaoyu Song. A quantitative characterization of cross coverage. IEEE Transactions on Computers, 64 (8):2408–2414, August 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2017:HEA

[ZLJ^{+}17] Y. Zhang, X. Liao, H. Jin, L. Gu, G. Tan, and B. B. Zhou. HotGraph: Efficient asynchronous processing for real-world graphs. IEEE Transactions on Computers, 66(5):799–809, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2015:BDR

[ZLLX15] Yunfeng Zhu, Jian Lin, P. P. C. Lee, and Yinlong Xu. Boosting degraded reads in heterogeneous erasure-coded storage systems. IEEE Transactions on Computers, 64 (8):2145–2157, August 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2011:SPM

[ZLN11] Kai Zheng, Hongbin Lu, and Erich Nahum. Scalable pattern matching on multicore platform via dynamic differentiated distributed detection (D^{4}). IEEE Transactions on Computers, 60 (3):346–359, March 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2017:MSM

[ZLSI17] Qi Zhang, Ling Liu, Gong Su, and Arun Iyengar. Mem-Flex: A shared memory swapper for high performance VM execution. IEEE Transactions on Computers, 66(9):1645–1652, September 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7885535/.

Zhang:2017:RTO

[ZLW^{+}17] Xuebin Zhang, Jiangpeng Li, Hao Wang, Danni Xiong, Jerry Qu, Hyunsuk Shin, Jung Pill Kim, and Tong Zhang. Realizing transparent OS/apps compression in mobile devices at zero latency overhead. IEEE Transactions on Computers, 66 (7):1188–1199, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/www.computer.org/csdl/
trans/tc/2017/07/07843601-
abs.html.

Zhang:2015:ARR

[ZLWZ15] Guangyan Zhang, Keqin Li, Jingzhe Wang, and Weimin Zheng. Accelerate RDP RAID-6 scaling by reducing disk I/Os and XOR operations. IEEE Transactions on Computers, 64 (1):32–44, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2016:PPR

[ZLX^{+}16] W. Zhang, Y. Lin, S. Xiao, J. Wu, and S. Zhou. Privacy preserving ranked multikeyword search for multiple data owners in cloud computing. IEEE Transactions on Computers, 65 (5):1566–1577, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2015:DSG

[ZLXW15] Shuming Zhou, Limei Lin, Li Xu, and Dajin Wang. The t/k-diagnosability of star graph networks. IEEE Transactions on Computers, 64(2):547–555, February 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2015:TSP

[ZLY15] Gaofeng Zhang, Xiao Liu, and Yun Yang. Time-series pattern based effective noise generation for privacy protection on cloud. IEEE Transactions on Computers, 64(5):1456–1469, May 2015.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2015:CLB

[ZLYS15] Chunsheng Zhu, V. C. M. Leung, L. T. Yang, and Lei Shu. Collaborative locationbased sleep scheduling for wireless sensor networks integrated with mobile cloud computing. IEEE Transactions on Computers, 64 (7):1844–1856, July 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2010:CNA

[ZM10] Gang Zhou and H. Michalik. Comments on “A New Architecture for a Parallel Finite Field Multiplier with Low Complexity Based on Composite Field”. IEEE Transactions on Computers, 59(7):1007–1008, July 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5416689.

Zheng:2017:HFA

[ZM17] Nan Zheng and Pinaki Mazumder. Hardwarefriendly actor-critic reinforcement learning through modulation of spike-timing-dependent plasticity. IEEE Transactions on Computers, 66(2):299–311, February 2017. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zaccaria:2018:SFH

[ZMB18] V. Zaccaria, F. Melzani, and G. Bertoni. Spectral features of higher-order side-channel countermeasures. IEEE Transactions on Computers, 67(4):596–603, April 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8103813/.

Zhang:2013:BAF

[ZMR^{+}13] Yiwei Zhang, Joseph P.

Mcgeehan, Edward M. Regan, Stephen Kelly, and Jose Luis Nunez-Yanez. Biophysically accurate floating point neuroprocessors for reconfigurable logic. IEEE Transactions on Computers, 62(3):599–608, March 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zong:2011:EPT

[ZMRQ11] Ziliang Zong, Adam Manzanares, Xiaojun Ruan, and Xiao Qin. EAD and PEBD: Two energy-aware duplication scheduling algorithms for parallel tasks on homogeneous clusters. IEEE Transactions on Computers, 60(3):360–374, March 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zilic:2013:GEI

[ZMS13] Zeljko Zilic, Prabhat Mishra, and Sandeep K. Shukla. Guest Editors’ introduction: Special section on systemlevel design and validation of heterogeneous chip multiprocessors. IEEE Transactions on Computers, 62 (2):209–210, February 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2015:TSR

[ZMW15] W. Zheng, K. Ma, and X. Wang. TE-Shave: Reducing data center capital and operating expenses with thermal energy storage. IEEE Transactions on Computers, 64(11):3278–3292, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2019:CXC

[ZMW^{+}19] X. Zhao, S. Ma, Z. Wang, N. E. Jerger, and L. Eeckhout. CD-Xbar: A convergediverge crossbar network for high-performance GPUs. IEEE Transactions on Computers, 68(9):1283–1296, September 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2011:EDG

[ZMY11] Miao Zhao, Ming Ma, and Yuanyuan Yang. Efficient data gathering with mobile collectors and spacedivision multiple access technique in wireless sensor networks. IEEE Transactions on Computers, 60(3):400– 417, March 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zadegan:2018:CFM

[ZNL18] Farrokh Ghani Zadegan, Dimitar Nikolov, and Erik Larsson. On-chip fault monitoring using self-reconfiguring IEEE 1687 networks. IEEE Transactions on Computers, 67(2):237–251, February 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/7990262/.

Zhou:2013:EPR

[ZOD13] Bo Zhou, Hiroyuki Okamura, and Tadashi Dohi. Enhancing performance of random testing through Markov Chain Monte Carlo methods. IEEE Transactions on Computers, 62(1):186– 192, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2019:SMB

[ZOK^{+}19] Yinglin Zhao, Peng Ouyang, Wang Kang, Shouyi Yin, Youguang Zhang, Shaojun Wei, and Weisheng Zhao. An STT-MRAM based in memory architecture for low power integral computing. IEEE Transactions on Computers, 68(4):617–623, April 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8523808/.

Zomaya:2011:SJb

[Zom11a] Albert Zomaya. State of the journal. IEEE Transactions on Computers, 60(5): 609–613, May 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zomaya:2011:SJa

[Zom11b] Albert Y. Zomaya. State of the journal. IEEE Transactions on Computers, 60(1): 2, January 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zomaya:2012:MEB

[Zom12a] Albert Y. Zomaya. A message to the Editorial Board. IEEE Transactions on Computers, 61(12):1665, December 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zomaya:2012:SJa

[Zom12b] Albert Y. Zomaya. State of the journal. IEEE Transactions on Computers, 61(1): 1–2, January 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zomaya:2012:SJb

[Zom12c] Albert Y. Zomaya. State of the journal. IEEE Transactions on Computers, 61 (2):145–150, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zomaya:2013:SJ

[Zom13] Albert Y. Zomaya. State of the journal. IEEE Transactions on Computers, 62(1): 1–5, January 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zomaya:2015:FSJ

[Zom15a] A. Y. Zomaya. Farewell state of the journal editorial. IEEE Transactions on Computers, 64(1):1, January 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zomaya:2015:MIS

[Zom15b] A. Y. Zomaya. In memoriam [Ivan Stojmenovic]. IEEE Transactions on Computers, 64(1):3, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zotov:2010:DVB

[Zot10] I. V. Zotov. Distributed virtual bit-slice synchronizer: a scalable hardware barrier mechanism for ndimensional meshes. IEEE Transactions on Computers, 59(9):1187–1199, September 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5396324.

Zavattoni:2015:SIA

[ZPM^{+}15] E. Zavattoni, L. J. Dominguez Perez, S. Mitsunari, A. H. Sanchez-Ramirez, T. Teruya, and F. Rodriguez-Henriquez. Software implementation of an attribute-based encryption scheme. IEEE Transactions on Computers, 64 (5):1429–1441, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2011:QAF

[ZQQ11] Xiaomin Zhu, Xiao Qin, and Meikang Qiu. QoS-aware fault-tolerant scheduling for real-time tasks on heterogeneous clusters. IEEE Transactions on Computers, 60(6):800–812, June 2011.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2019:SCS

[ZQZ^{+}19] X. Zhang, Z. Qian, S. Zhang, X. Li, X. Wang, and S. Lu. Semi-clairvoyant scheduling in data analytics systems. IEEE Transactions on Computers, 68(9): 1376–1389, September 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhan:2015:PBT

[ZR15a] Xin Zhan and S. Reda.

Power budgeting techniques for data centers. IEEE Transactions on Computers, 64(8):2267–2278, August 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2015:SDL

[ZR15b] Z. Zhang and T. G. Robertazzi. Scheduling divisible loads in Gaussian, mesh and torus network of processors. IEEE Transactions on Computers, 64(11): 3249–3264, November 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhong:2015:DPM

[ZRL15] Ziming Zhong, V. Rychkov, and A. Lastovetsky. Data partitioning on multicore and multi-GPU platforms using functional performance models. IEEE Transactions on Computers, 64(9): 2506–2518, September 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2016:GDL

[ZRS^{+}16] Jiao Zhang, Fengyuan Ren, Ran Shu, Tao Huang, and Yunjie Liu. Guaranteeing delay of live virtual machine migration by determining and provisioning appropriate bandwidth. IEEE Transactions on Computers, 65(9): 2910–2917, September 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zha:2013:GGH

[ZS13] Xinyan Zha and Sartaj Sahni. GPU-to-GPU and host-to-host multipattern string matching on a GPU. IEEE Transactions on Computers, 62(6):1156–1169, June 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zanon:2019:FKC

[ZSP^{+}19] G. H. M. Zanon, M. A. Simplicio, G. C. C. F. Pereira, J. Doliskani, and P. S. L. M. Barreto. Faster key compression for isogenybased cryptosystems. IEEE Transactions on Computers, 68(5):688–701, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2015:ASP

[ZT15] Hanying Zheng and Xueyan Tang. Analysis of server provisioning for distributed interactive applications. IEEE Transactions on Computers, 64(10):2752–2766, October 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2014:DMA

[ZV14] Qin Zheng and Bharadwaj Veeravalli. On the design of mutually aware optimal pricing and load balancing strategies for grid computing systems. IEEE Transactions on Computers, 63 (7):1802–1811, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2013:REB

[ZWC13] Qiang Zhu, Xin-Ke Wang, and Guanglan Cheng. Reliability evaluation of BC networks. IEEE Transactions on Computers, 62(11): 2337–2340, November 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2018:DSS

[ZWC^{+}18] Yu Zhang, Qingsong Wei, Cheng Chen, Mingdi Xue, Xinkun Yuan, and Chundong Wang. Dynamic scheduling with service curve for QoS guarantee of largescale cloud storage. IEEE Transactions on Computers, 67(4):457–468, April 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8107532/.

Zheng:2016:BDD

[ZWD^{+}16] X. Zheng, J. Wang, W. Dong, Y. He, and Y. Liu. Bulk data dissemination in wireless sensor networks: Analysis, implications and improvement. IEEE Transactions on Computers, 65 (5):1428–1439, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2015:AAD

[ZWH^{+}15] Tao Zhang, Jianxin Wang, Jiawei Huang, Yi Huang, Jianer Chen, and Yi Pan. Adaptive-acceleration data center TCP. IEEE Transactions on Computers, 64 (6):1522–1533, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2015:CMC

[ZWL15] Sheng Zhang, Jie Wu, and Sanglu Lu. Collaborative mobile charging. IEEE Transactions on Computers, 64(3):654–667, March 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2019:FCS

[ZWL^{+}19] Q. Zhou, K. Wang, P. Li, D. Zeng, S. Guo, B. Ye, and M. Guo. Fast coflow scheduling via traffic compression and stage pipelining in datacenter networks. IEEE Transactions on Computers, 68(12):1755–1771, December 2019. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2015:NOP

[ZWLS15] Baoxian Zhang, Xili Wan, Junzhou Luo, and Xiaojun Shen. A nearly optimal packet scheduling algorithm for input queued switches with deadline guarantees. IEEE Transactions on Computers, 64(6):1548– 1563, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2016:CEC

[ZWW^{+}16] G. Zhang, G. Wu, S. Wang, J. Shu, W. Zheng, and K. Li. CaCo: An efficient Cauchy coding approach for cloud storage systems. IEEE Transactions on Computers, 65(2):435– 447, February 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2019:ADL

[ZWW19] Jian Zhou, Huafeng Wu, and Jun Wang. ApproxSSD: Data layout aware sampling on an array of SSDs. IEEE Transactions on Computers, 68(4):471–483, April 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8467988/.

Zhang:2012:NPM

[ZWX12] Ziyi Zhang, Xin Wang, and Qin Xin. A new performance metric for construction of robust and efficient wireless backbone network. IEEE Transactions on Computers, 61(10):1495–1506, October 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2015:SPR

[ZWYY15] Jin-Xin Zhou, Zhen-Lin Wu, Shi-Chen Yang, and Kui-Wu Yuan. Symmetric property and reliability of balanced hypercube. IEEE Transactions on Computers, 64(3):876–881, March 2015.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2014:SDF

[ZXX^{+}14] Yunfeng Zhu, Liping Xiang, Yinlong Xu, John C. S. Lui, Patrick P. C. Lee, Runhui Li, and Silei Xu. Single disk failure recovery for X-code-based parallel storage systems. IEEE Transactions on Computers, 63(4):995–1007, April 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2012:BRH

[ZY12] Miao Zhao and Yuanyuan Yang. Bounded relay hop mobile data gathering in wireless sensor networks. IEEE Transactions on Computers, 61(2):265– 277, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2016:PPD

[ZYC16] Q. Zhang, L. T. Yang, and Z. Chen. Privacy preserving deep computation model on cloud for big data feature learning. IEEE Transactions on Computers, 65 (5):1351–1362, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2016:CAE

[ZYHZ16] Yan Zhang, Guowu Yang, William N. N. Hung, and Juling Zhang. Computing affine equivalence classes of Boolean functions by group isomorphism. IEEE Transactions on Computers, 65 (12):3606–3616, December 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2015:RAS

[ZYL15] Ziming Zheng, Li Yu, and Zhiling Lan. Reliabilityaware speedup models for parallel applications with coordinated checkpointing/ restart. IEEE Transactions on Computers, 64(5):1402– 1415, May 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2016:HSC

[ZYW^{+}16] Weihua Zhang, Shiqiang Yu, Haojun Wang, Zhuofang Dai, and Haibo Chen. Hardware support for concurrent detection of multiple concurrency bugs on fused CPU–GPU architectures. IEEE Transactions on Computers, 65(10):3083– 3095, October 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2010:MMC

[ZYY10] Hao Zheng, Haiqiong Yao, and T. Yoneda. Modular model checking of large asynchronous designs with efficient abstraction refinement. IEEE Transactions on Computers, 59(4):561–573, April 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5374376.

Zhou:2018:RIB

[ZYY18] Quan Zhou, Liang Yang, and Xin Yan. Reconfigurable instruction-based multicore parallel convolution and its application in real-time template matching. IEEE Transactions on Computers, 67(12):1780– 1793, December 2018. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8375740/.

Zheng:2010:PPT

[ZZ10] Hongzhong Zheng and Zhichun Zhu. Power and performance trade-offs in contemporary DRAM system designs for multicore processors. IEEE Transactions on Computers, 59(8): 1033–1046, August 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=
                                        &arnumber=5467050.

Zarrinchian:2017:LBS

[ZZ17] Ghobad Zarrinchian and Morteza Saheb Zamani. Latch-based structure: A high resolution and selfreference technique for hardware Trojan detection. IEEE Transactions on Computers, 66(1):100–113, January 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2019:CUC

[ZZ19] Y. Zhao and H. Zeng. The concept of unschedulability core for optimizing realtime systems with fixedpriority scheduling. IEEE Transactions on Computers, 68(6):926–938, June 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2019:QDA

[ZZJ^{+}19] M. Zhang, L. Zhang, L. Jiang, F. T. Chong, and Z. Liu. Quick-and-Dirty: An architecture for high-performance temporary short writes in MLC PCM. IEEE Transactions on Computers, 68(9): 1365–1375, September 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2014:RRD

[ZZL14] Guangyan Zhang, Weimin Zheng, and Keqin Li. Rethinking RAID-5 data layout for better scalability. IEEE Transactions on Computers, 63(11):2816–2828, November 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2015:CMC

[ZZM^{+}15] Hao Zheng, Zhen Zhang, C. J. Myers, E. Rodriguez, and Yingying Zhang. Compositional model checking of concurrent systems. IEEE Transactions on Computers, 64(6):1607–1621, June 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2010:AND

[ZZS10] Guangyan Zhang, Weiman Zheng, and Jiwu Shu. ALV: a new data redistribution approach to RAID-5 scaling. IEEE Transactions on Computers, 59(3):345– 357, March 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5276795.

Zhang:2015:NPL

[ZZX^{+}15] Daqiang Zhang, Daqing Zhang, Haoyi Xiong, L. T. Yang, and V. Gauthier.

NextCell: Predicting location using social interplay from cell phone traces. IEEE Transactions on Computers, 64(2):452–463, February 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2014:TEP

[ZZYZ14] Ping Zhou, Bo Zhao, Jun Yang, and Youtao Zhang. Throughput enhancement for phase change memories. IEEE Transactions on Computers, 63(8):2080–2093, August 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).